A Complete Bibliography of Scientific American: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 November 2023 Version 4.10

Title word cross-reference

3 [666]. 4 [4011]. ^{3} [203]. _{2} [848, 859, 2343, 4125]. _{3} [2343]. _{x} [520]. CO_{2} [860].

-D [666, 4011]. -Rich [860, 859].

0 [4054].

1 [2959, 2963, 3975]. 1-2-3 [256]. 1793 [4535]. 183 [1956]. 1940s [4604]. 1990s [722]. 1992 [4769]. 1996 [3526]. 1997 [3764]. 1998 [4968].

2 [85]. 2000 [4641]. 20s [4778]. 20th [2921]. 21 [1193]. 21st [2646, 5296, 4268, 2629, 2623, 2636, 2650, 5303, 4323, 723]. 21st-Century [2629, 2623]. 273 [616]. 29 [3806].

3 [3560]. 3-D. [3560]. 30s [4788]. 3C [616].

40s [4788, 3339]. 486 [189].

50s [4800].

60s [4800].

70s [4808]. 747s [1689]. 763-mph [4053].

9 [2557]. ’90s [807].

A.D. [3464]. Aardvark [1572]. abductions [2977]. Abilities [4369]. Aborted [40, 5318]. Aborting [1088]. abortion [5319]. Abounding [2475]. above [2170]. Absorb [862]. Abu [2026]. Abuse [367, 98, 250, 2688]. Abyssal [889]. Acacia [392]. Academic [3430, 892]. Academy [1786]. accelerated [4053]. Access [3828, 4005, 4525, 5256, 4476]. Accidental [397]. Accidents [586, 3254]. According [2887]. Accounting [1033, 809]. Accretion [863]. Accurate [1524]. Achievement [892]. Achieving [1105]. Achille [3298]. Acid [109, 2709, 3507, 965, 1497, 4554]. Acoustic [487, 4637, 327, 336]. Acoustics [5293, 5300, 480]. Acquired [1300]. Acquisition [1331]. Acronyms [3944]. Act [2684, 3631, 2293]. Acting [2714]. activist [3873, 4358]. Activity [2170]. Ada [5069]. Adam [5062]. Adaptation [692]. Adapting [1325, 1320]. Adaptive [1946]. Add [923, 3881]. add-ons [3881]. Addiction [513, 3593, 520]. addictive [3705]. Addressing [4782]. adds [4137]. adhesion [3741]. Adjusted [319]. Administrator [1024]. Ado [3457, 3461]. Adoptive [157]. Adult [5066]. Advanced [130]. Advances [3325, 3338, 3336, 3335]. Advancing [3344, 3343]. Advantage [3104, 718, 1764]. Adventure [4848, 4606]. Adventures [5081]. Advertised [1919]. Advertisers [4058]. Advertising [1440, 4057, 1421]. advice [2755]. Advised [3880]. Aerial [5152]. Aerodynamics [3946]. Aerosol [1788]. Aerospace [226, 229]. affect [4230]. affects [4799]. afflict [4789]. Aflame [4565]. Africa [3034, 5157, 170, 3583, 3788, 3022, 4844, 3460, 3698, 5156]. African [3004, 4918, 4845, 959]. Africanized [1722]. After [4631, 849, 4484, 5201, 4772, 3091, 2918]. Aftershocks [3]. Again [1052, 2703, 4043, 341, 1533, 2261, 163, 3698, 5359, 3698, 3909]. Against [3171, 42, 402, 408, 3320, 586, 2745, 3478, 3909, 3832, 4053, 4476, 4348, 4914, 4954, 4994, 4895, 4859, 4784, 4797, 3469, 3177, 5215, 409, 3319, 4660, 5103, 4959]. Age [5126, 2616, 3886, 1221, 172, 4047, 1049, 1258, 1265, 1266, 5302, 1387, 4042, 2625, 1886, 1987, 885, 1289]. Agenda [319, 1193]. Agent [3573, 1603]. Agents [2269, 4155, 4868]. Ages [2992]. Aging [1408, 5341, 1142, 1912, 1913, 3882, 2900, 3560, 1906]. Ago [175, 2, 38, 69, 106, 139, 176, 211, 245, 276, 312, 351, 384, 427, 461, 497, 535, 570, 605, 638, 674, 708, 746, 781, 814, 846, 882, 914, 948, 982, 1017, 1047, 1087, 1121, 1159, 1160, 1197, 1198, 1240, 1241, 1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1304, 1341, 1378, 1396, 1436, 1512, 1548, 1620, 1658, 1701, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 1739, 1777, 1815, 1853, 1890, 1933, 1970, 2008, 2047, 2087, 2125, 2166, 2207]. Ago [2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2255, 2208, 2291, 2335, 2420, 2465, 2466, 2518, 2574, 2679, 2743, 2812, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 2873, 2929, 2979, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 3950, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623, 4675, 4719, 4499, 4674, 4720, 4837, 4838, 4889, 4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317, 385, 571, 1473, 4121, 4174, 4229, 4289, 4345, 4396]. ago [4448]. Agricultural [1037]. Agriculture [1980, 2651, 204, 3290]. Aground [1574]. Ahead [1283, 5028, 4551]. Aid [2128, 3097]. Aids [2707, 992, 2872, 3299, 4476, 4285, 472, 434, 4290, 3004, 1794, 3152, 4475, 1599, 945, 257, 3928, 1810, 712, 713]. AIDS-Related [257]. AIDS-Resistance [3928]. Aim [4346, 2529]. ain’t [2150]. Air [2052, 2772, 5189, 4330, 3685, 2283, 2279, 751, 1630, 4924, 3989, 3187, 1108, 903, 1107, 1906, 1913, 3186, 4410, 4908, 4125, 1103]. Air-Breathing [4924]. Air-Cooled [3989]. Airborne [4236]. Airliner [2627]. airplane [3089]. Airways [393, 1912, 1906, 1913]. Alan [3154, 3123, 3730, 4233, 5032, 4244, 3130]. Alarm [4046]. Alarming [3310]. Albert [2982]. Alcohol [3484, 3060, 4427, 4798]. Alert [782, 4018]. Alex [4824]. Algae [2745]. Algorithm [3661]. Algorithms [1064]. alias [2822]. Alien [442, 2977, 4588]. Aliens [3241]. Alive [1858, 405]. All-Optical [2608, 2615, 4722]. All-Star [1593]. Allen [4941]. Allergy [1601]. Alliance [1608]. Allies [1150]. Alloy [656]. Almost [2575]. Alone [2661]. Along [232, 660, 659, 655]. Alpha [1957]. Alphabet [2594, 3368]. alphabetical [3558]. Alphamagic [3557]. already [4756, 2833]. also [3741]. Altering [1961, 1951]. Alternative [1904, 3400, 3358]. Alternatives [3910, 4679]. although [4476]. Altruism [500]. Aluminum [2059, 2062]. Alzheimer [3140, 1667, 2257, 4811, 794, 790]. Amateur [271, 1232, 1334, 1391, 1466, 1579, 1650, 1731, 1807, 1883, 1962, 2039, 2118, 2197, 2790, 2860, 2562, 2369, 2455, 2917, 3121, 3512, 3462, 2906, 3015, 3284, 3173, 2964, 3364, 3606, 4024, 3556, 3662, 4072, 4536, 4584, 4378, 4487, 4431, 4211, 4273, 4331, 4936, 2038, 1231, 2789, 2859, 3227, 3234, 3129, 3283, 3511, 3014, 2963, 3181, 3461, 3068, 3363, 2916, 3414, 3661, 3555, 4023, 3800, 3890, 3846, 3704, 3605, 3986, 4071, 3935, 3747, 4210, 4545, 4332, 4271, 4706, 4707, 4661, 4157, 4330, 4272, 4377, 4592, 4747, 4430, 4486, 5218, 5036, 5145, 5107]. Amateur [5308, 4977, 5072, 4873, 5259, 5184, 4872, 4935, 2368, 1806, 1730, 2561, 2283, 1333, 2117, 1961, 1465, 1011, 1649, 1882, 810, 2454, 207, 270, 346, 943, 1578, 1080, 2196, 877, 1298, 33, 65, 135, 1154, 2284, 3705, 3847, 3801, 3891, 3936, 3748, 3610]. Amazing [1899]. Amazon [1290, 1388, 1531, 1527, 1530]. Amazonian [3178]. Amber [3064, 3062, 3529]. Ambivalent [2423]. America [4464, 1284, 1380, 898, 2217, 5217, 1796]. American [2866, 4076, 3517, 4750, 4053, 4961, 1787, 650, 3060, 1000, 315, 2863, 3515, 4075, 1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 2207, 2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623, 4675, 4719, 4838, 4889]. American [4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317, 4284]. American-Style [4961]. Americans [4918, 1214, 1220, 2329]. Ammonia [5187]. amniocentesis [4679]. amok [2933]. Among [2886, 3149, 4504, 3738, 1695, 4147, 4865]. Amphibian [2356]. Amphibians [4518, 3622]. Amplification [2005]. Amyloid [794]. Analysis [5194, 5274, 4509, 4900, 2428, 3477, 3298, 3192, 4039, 3996, 3908, 3571, 3719, 3862, 4397, 4291, 4993, 5228, 4231, 3819, 3379, 3429, 3620, 4677, 4449, 4346, 5050, 4953, 5134, 4843, 5318, 3141, 3084, 3766, 3244, 3030, 2479, 1521, 4176, 4556, 3527, 4124, 449, 3951, 3676, 4721, 4625, 5119, 4891, 5156, 5084, 4507, 3820, 4399, 4436]. Analytical [2203, 2150, 2814, 2746, 2520, 2681, 2250, 2330, 2374, 2380, 2461, 2507, 2576, 2931, 2981, 2874, 59, 129, 164, 269, 377, 417, 527, 597, 839, 1370, 1577, 1612, 2479, 2602, 2347, 2267, 26, 95, 200, 236, 307, 345, 452, 490, 564, 1539, 2707, 2839, 2219, 2306, 2890]. Anatomical [306]. ancestor [3823]. ancestors [3034]. Ancestral [4130]. Ancient [2003, 744, 4439, 1681, 1405, 656, 516, 518, 134, 3895, 3116, 3120, 3509, 2320]. Andaman [5070]. Andean [1947]. Andrei [4976]. Andro [4723]. Anergy [1560]. angel [4029]. Angeles [3895, 1630]. angina [3832]. Angst [3573, 4723]. Animal [3600, 3762, 3601, 3603, 1754, 2590, 812, 1219, 3604, 3568, 3599, 3368, 4878, 4531]. Animals [550, 2269, 4825, 4823, 4826]. Animating [4259]. Anniversaries [2423]. Annual [423, 422, 842, 1271, 1734, 2200, 2868, 3519, 4079, 4753, 5353, 2172]. Answer [2237, 2244]. Answers [4480]. Ant [3800]. Antarctic [2747, 1277]. Antarctica [1278]. Antelope [392]. anthropologist [4135]. anthropology [4453]. Anthropomorphic [2855, 2861]. Anti [3732, 2015, 4895]. anti-nuke [4895]. Anti-omniscience [2015]. Anti-Parkinson [3732]. Antibiotic [4257]. Antibodies [4009]. Antibody [57]. Antichaos [692]. Antidrug [151]. Antigens [2023]. Antigravity [4144, 4863]. Antigreen [2033]. Antihydrogen [3034]. antimalaria [3852]. Antimatter [4646, 8, 4313, 3038, 4839]. antinuke [5195]. antiporn [3881]. Antiprotons [1257]. Antirust [4735]. Antiscience [3553, 684, 3554]. Antisense [905, 21, 4689]. Antishoplifting [3757]. Ants [1175, 260, 4898, 1125, 315, 4900, 4895]. Anty [1999, 1990]. Anty-Particles [1999, 1990]. Anxiety [876]. Any [3391, 3720]. Anybody [4164, 4165]. Anyone [4751, 4752]. Anyons [581]. Anything [1311, 3374]. Anywhere [2631, 2620]. apart [3302]. Apartheid [3074]. apes [3823]. Apocalypse [5322]. Apocryphal [2950]. Apollo [2730, 5133, 5131, 1981]. Apple [2181]. applications [4320]. appreciated [4205, 4215]. Apprehensions [1644]. Approach [1950]. approved [4689]. Aquaculture [4612, 4424, 3193]. Arcadia [3851]. Archaeological [3709]. archaeologists [3193].

Archaeology [3886, 2408, 2413, 994, 1002, 62]. Archaeopteryx [158]. Architects [1789]. Architecture [475, 479, 4143]. Arctic [4606]. Aren’t [4632]. argues [3051, 4493]. Arguments [3193]. Arises [4479, 3322, 3323]. Arithmetic [3236, 3228]. Arithmetics [2446]. Arm [4784]. Armada [4776]. Armageddon [4641, 3300]. Armchair [3705, 3704]. armies [4265]. Arms [4453, 4459, 2577]. Aromatic [928]. Arrest [2605]. Arrhythmias [1525]. Arrivals [883]. Arrives [3099, 4562]. Arsenide [60]. Art [3008, 3007, 2368, 2361, 3287, 1965, 2841, 1671, 1676, 1677, 264, 263, 3808, 2295, 3169, 3172, 2057, 1357, 1358, 2363, 2362, 3460, 5146, 3288]. Artful [3148]. Arthritis [3779, 3783]. Arthur [623]. Artificial [17, 3311, 3443, 4827, 2841, 3314, 1668, 2635, 2617, 826, 833]. Artisans [971]. Artist [1348]. Artistic [2471]. Artistry [4647, 4648]. Artzybasheff [1671, 1676, 1677]. Asbestos [3843]. Ascent [5284]. Ascorbic [1497]. Ashes [609]. Asia [4844]. Asian [1256]. ask [3528]. Aspirin [441]. Assembling [2630, 2640]. Assembly [5366]. Assets [2936, 1033]. Assisted [4793]. Associates [3410]. Association [2173]. Asteroid [1406]. Asteroids [4768, 765, 3003, 4350]. Asthma [5272]. Astronomer [4649]. Astronomers [2340, 3610]. Astronomical [1908, 1907, 1022, 2169]. Astronomy [2019, 1221, 2799, 5371, 73, 467, 797, 802, 1192, 4381]. Astrophysics [1449, 1456, 930, 933]. Aswan [3832]. asymmetries [3021]. Asymmetry [4646, 4839]. AT&T [3632]. Athlete [3163, 3165]. athletes [4504]. Atlantic [4599, 4616]. Atlas [3851]. Atlases [2794]. Atmosphere [126, 4770]. Atmospheric [4332, 5218, 3507, 3748, 4378]. Atoll [4604]. Atom [4472, 1944, 3897]. Atomic [2809, 2410, 5259, 4770, 182, 3404, 783, 2474, 3138, 2929, 3397, 5238]. atomized [3179]. Atoms [4279, 4280, 3666]. Attack [4898, 503, 4989, 4063, 3032, 4606]. Attacking [3354, 3783]. Attacks [4789]. Attention [3245, 4576, 4581, 2260, 3244]. Attention-Deficit [4576, 4581, 3245]. Attitude [4649]. Attracted [2758]. Attraction [5049, 5050]. Attractive [2748, 1442]. Attractors [2212, 3733]. Audio [4639]. Augmenting [3048]. Austerity [1022]. Australia [1523]. Austronesian [657]. Autism [1490, 4097]. Autoimmune [4789, 4790, 1600]. Automata [32]. Automated [3927, 3972, 231]. automatic [3441, 4339]. Automobile [2626]. Automotive [2184]. autopilot [2676]. Avalanches [3667]. average [3420]. averages [4325, 4333]. Aviation [3739]. Avoiding [4484, 972]. Awaiting [3534]. Awards [2866, 4076, 2863, 3515, 3517, 4075, 4750]. Away [3161, 2541, 966, 977]. Ax [1068]. Axis [3915]. Aztec [3933].

B [3400, 4358, 4953, 4954, 3806, 552]. B-29 [3806]. B1 [3561]. Baa [3720]. Babbage [1361]. Babble [23]. Babies [3037, 1690, 834, 826, 833]. Baby [5363]. Back [5133, 1786, 4668, 4669, 4071, 4519, 4528, 2840, 5131, 1292, 2435, 5370, 1824, 4405, 2309, 1326, 2397, 4569, 2825, 4954, 5350]. Backfire [2053]. backflip [2690]. Background [2954, 31, 2553, 4227, 4235]. Backyard [30, 3063, 3068, 1333, 1324, 3625, 3364, 3069]. Bacteria [3909, 2957, 737, 3597, 3589, 4679, 4042, 4735]. Bacterial [4146, 1098]. Bad [5060, 5347, 3000, 4799, 2181, 5330, 3247, 3875]. badly [2150]. Bagging [3467]. Bags [468, 3187, 3693, 3186]. Balance [4290, 4291, 3173]. Balanced [1439]. Ball [3385, 2540, 4813, 3607, 2524]. Ballistic [255]. Ballistics [5007]. Balloon [5298, 5306, 3147]. Ballooning [5302]. balloons [2742, 4946]. Ballpark [974]. balls [3608]. Ban [3483, 2303]. Band [2707]. Band-Aids [2707]. Bandwidth [3543, 3854, 410, 4722]. Bang [1858, 2049, 2942, 3664, 4504, 4772, 787, 911, 3501, 1219, 4970, 3534]. Bangs [4692]. Bank [4168, 5299, 5307]. Banking [58]. Banks [1577]. Banner [1495]. Banning [3769]. Banzai [1516]. barometers [5151]. Baron [4077]. Barrel [3092]. barren [3997]. Barrier [5370, 1359]. Barriers [1983]. Bars [5165]. Baseball [3945]. Based [2319, 2966, 5384]. Basement [5143]. Bashing [2470]. Basic [1878, 1928]. Basics [1292]. Basis [2315, 1137, 1949]. Bass [4386, 4387]. Bat [5081, 5106, 5105]. Bats [191, 61]. Batteries [2155, 2150, 3690]. Battery [5144]. battle [4994]. Battlefields [1862]. Battling [3046]. Bay [3211]. Be [5022, 4779, 3974, 5341, 3168, 30, 2665, 3464, 3383, 3705, 4072, 3865, 4336, 4807, 4773, 4193, 4783, 4786, 3829, 4774, 4710, 2833, 4613, 5168, 3166, 2759]. Beam [390, 3925]. Beams [766]. Bean [2224]. Bears [1050]. Beast [2502, 4606]. Beastly [1845, 1834]. Beating [2150, 4736, 3589, 3730]. Beats [1491, 1503, 4053]. Beautiful [3676]. Beauty [2502, 1455, 3175, 2970]. Became [1099, 4439]. because [4603]. Become [2577, 1469, 4793, 4679]. Becomes [5017, 2394]. Bee [75, 3770]. Beer [3277, 4077]. Bees [1722]. beetle [2860]. Beetles [3664, 4350]. Beewolves [690]. Before [63, 2230, 5333, 2496, 4251, 2493]. begets [4797]. Beginning [482, 478, 481]. begins [3302]. Behavior [1413, 2359, 5015]. Behavioral [2316, 2321, 1488, 1493]. behaviors [4476]. Behind [110, 1551, 2957, 5165, 2285, 3521, 4539, 4643]. Being [641, 2429, 3210, 2439]. Believing [1790, 1791, 4500, 838]. Bell [2209, 4819, 2372, 5202, 5151, 4606]. Bellcore [2220]. Bellows [4488]. Bells [217, 1206]. Below [1827]. Belt [2305, 3103, 3111]. Belts [961]. Beluga [3114, 3119]. Ben [3852]. Bends [3913]. Benefits [3600, 3403, 2553, 3599]. Benjamin [1501]. Bent [1129]. Berners [4051]. Berners-Lee [4051]. Berry [3036]. Best [1163, 1713, 682, 2350, 2349, 4967, 4076]. Better [983, 2241, 2240, 904, 1464, 2430, 3746, 3745, 4637, 4744, 77, 1844, 3382, 3637, 2685, 3198, 4476, 4773, 2232]. Bettering [2155]. between [4336, 3885, 4646, 4839]. Beverage [2059, 2062]. Beware [1785, 3778, 3705]. Beyond [256, 837, 2544, 4511, 2696, 3483, 4328, 2686, 976]. Big [2148, 2147, 4039, 3719, 5288, 301, 247, 1513, 2760, 4002, 4517, 3749, 1799, 5009, 2266, 2336, 3664, 4772, 787, 2142, 911, 4460, 3501, 1219, 4970, 3534]. Big-League [5009]. Big-Time [2148, 2147, 301, 2142]. Bigfoot [2979]. Bigger [4748, 1844]. Biggest [4066, 4065, 648, 4504]. Bikini [4604]. Bill [2436, 3044, 2934, 2950, 4689]. Billion [4996, 4741]. Billions [1186].

Binaries [1675]. Binary [863, 2544, 2971, 2970, 2404, 2072, 999]. Binding [702]. Binoculars [5151]. Biocatalysts [1842]. Biochips [375]. Biodiversity [3083, 3865, 1333, 1324, 4539]. Biogenetics [3872]. biography [206]. Biological [2142, 2147, 4004, 1903, 3500, 1868, 1872, 1137, 1902, 4610, 1217, 1218, 5103, 993, 654, 1726]. Biologically [1901]. Biologist [3051]. biologists [4487]. Biology [691, 693, 2148, 2142, 2147, 1075, 4666, 4667, 1258, 1265]. biomass [4042]. Biomolecular [2231]. biopesticide [4360]. Biosensors [689]. biosolar [4236]. Biosonar [191]. Biosphere [2522]. Biotech [1037, 3634]. Biotechnology [1111, 3873]. Biowar [1662]. Biowarfare [1745]. bioweapons [4994]. Bird [4210, 4185, 4886]. Bird-Watching [4210]. Birds [3575, 3911, 3918, 1796, 4199, 643]. Birth [4795, 3055, 3037, 3307, 213, 4100, 4971, 2236, 2238, 401, 403, 3054, 3056, 3561, 4771, 396, 4724]. Birth-Control [3307, 213]. birthday [3560, 4433]. Bismarck [4617]. bison [2741]. Bit [3564, 614]. Bite [4132]. Bites [184]. Bits [5374, 2035, 1080, 1069]. Black [1285, 1381, 2579, 4464, 3170, 648, 2470, 1337, 5065, 363, 4200, 3697, 4771, 93]. Black-market [2579]. Blackballing [1611]. Blackbird [39]. Blade [1804]. blame [3282]. Bland [4310, 4311]. Blanket [2919, 2909]. Blast [4901, 2533]. Bleaching [1315]. Blessings [4717]. Blight [3770, 4405, 238]. Blind [3550, 997, 996, 2890, 3091]. Blindingly [4350]. blink [3531]. Bliss [888]. Blitzing [5374]. Blobs [4546]. Block [2220, 3312]. Blockvotia [2497, 2500]. Blood [1573, 4919, 177, 484, 3354, 3341, 3314, 4203, 4875, 4876, 4168, 1359, 3311, 4398, 4634]. Blood-Brain [1359]. Bloodless [4636]. Bloomer [3578]. Blooms [4032]. Blow [3628]. Blower [3919, 3920]. Blows [3493]. Blue [3893, 738, 1575, 3218, 12, 3980, 1995, 3864, 3923, 3922, 2697]. Blue-Laser [3218]. Blueprint [2449, 2448]. Blues [2013, 2439, 9]. Blumberg [172]. Blunder [4688]. Blurred [277]. Blustery [3252]. Board [2157, 2146, 3017, 3938]. Boat [107]. Bodies [4409]. Body [5100, 773, 225, 4789, 1570, 1597, 1789, 3628, 409, 402, 408, 4087, 2980, 4571, 4579, 3741]. Bohm [1904]. Bohr [2569, 2410, 2412, 2411]. Boldly [5241, 3572]. Boloney [3292]. bolts [4527]. Bomb [3125, 3752, 2412, 1375, 3302, 4501]. bombers [3806]. Bombs [3620, 2541, 4358, 3467]. Bonanza [57]. Bonds [3009, 3014, 3015]. Bone [3345, 1914, 702, 3382]. Bones [3756, 2239, 4905, 2026, 1979, 4041, 3089, 4788]. Bonestell [1908, 1907]. Bonobo [2317]. Bonus [5131, 1690, 853]. Booby [1313]. Book [1270, 1336, 1374, 1392, 1430, 1467, 1468, 1504, 1506, 1543, 1581, 1651, 1652, 1693, 1694, 1732, 1733, 2002, 1771, 1809, 1848, 1885, 1927, 1964, 2040, 2041, 2081, 2120, 2160, 2199, 2247, 2286, 2866, 4076, 1081, 1164, 2285, 1373, 1299, 634, 670, 743, 777, 811, 705, 878, 910, 978, 1044, 1191, 1233, 1269, 944, 1012, 1117, 1155, 1429, 1335, 1580, 1614, 1770, 1808, 1847, 1884, 1926, 1963, 2001, 2080, 2119, 2159, 2198, 2325, 2326, 2246, 2863, 3515, 3517, 4075, 4750, 1542, 2371, 2370, 5211, 4247, 4248, 4588]. Books [421, 456, 494, 601, 1234, 1192, 4878, 4942, 530, 493, 628, 172, 136, 35, 66, 103, 208, 241, 272, 309, 347, 379, 420, 670, 743, 777, 811, 841, 455, 566, 600, 878, 1191, 1233, 1269, 3291, 2730, 3072, 3125, 3752, 3560, 4712, 4710]. Boom [4013, 5234, 4186]. Boomerang [183]. Booming [3934]. Boosting [279]. Boot [2593]. Boots [1112]. Bopp [3606, 3556]. Borcherds [4687]. Bordeaux [4946]. border [3832]. Borders [5379]. Boris [1671, 1676, 1677]. Born [2242, 4194, 2829, 4618, 986]. Boron [337]. Borrowed [2056]. Bose [4255, 2935]. Bosnia [2984]. Boston [2578]. botanical [4206]. both [3664, 4617, 4773]. Bottle [5363, 4936]. Bottled [8, 3369]. bottles [4274]. Bottom [1202, 2867, 5377, 706]. bottomline [4539]. Bouncing [5191]. Boundary [1447]. Bourgeoisie [908]. Bouts [2454, 2447]. Bow [620]. Box [5161, 4186]. Boyk [2870]. Boys [4456, 1048]. Brace [586, 4858]. braces [3154]. Bracing [2336]. Brain [3504, 132, 949, 3511, 1674, 1138, 5342, 5356, 5343, 804, 1133, 4471, 1141, 5127, 1636, 2130, 5066, 1140, 1945, 4094, 4062, 1402, 13, 19, 1142, 1134, 2779, 1359, 4801, 3260, 1135, 2521, 2688, 3512, 3664]. Brainchildren [4493]. brains [4398, 4127, 3126]. brainstorm [5151]. Brakes [1424, 698, 4221, 4209, 4222]. Branching [2133, 4774]. Brando [3466]. Brass [3436]. Brawny [4240]. Brazil [2983]. Breaching [1359]. Break [1625, 2308, 5244, 4652]. Breakers [1364]. Breakfast [2922]. Breaking [3022, 1753]. breakthrough [4676]. Breakthroughs [3061, 3066, 3067]. Breakup [127]. Breast [4803, 2725, 2724, 2701, 4796, 2175, 2852, 4804, 4401]. Breath [5096, 4909, 1066]. breathe [2860, 4543]. Breathing [5209, 4924, 2383]. Bred [1914]. Breed [712]. Breeder [2891]. Breeze [145]. Brewing [2268]. Brick [5108]. Bridge [4084, 4067, 5372, 5113]. Bridges [1380, 991, 1284]. Bridging [5379]. Brief [3144, 3197, 3532, 4131, 4238, 4352, 4683, 4726, 4998, 5054, 5088, 5125, 5162, 5199, 5232, 5324, 3035, 3087, 3249, 3303, 3384, 3434, 3482, 3576, 3626, 3681, 3724, 3771, 3826, 3867, 3916, 3954, 4001, 4044, 4183, 4298, 4404, 4454, 4514, 4561, 4630, 2364, 634, 910, 978, 1044, 2325, 4847, 4899, 4957, 5279]. Briefing [4840, 3764, 3525, 4123]. Bright [1461, 159, 1938, 4517, 4370, 4371, 1706, 2116]. brighter [4741]. Bring [3431]. Bringing [3641, 1202, 3642, 4365, 5370, 4917, 3799, 3798]. brings [2982]. British [1537]. Broadband [5251, 5252, 5253, 5256]. Broadcast [74]. Broadcasting [3088]. Broadest [5253]. Broken [2259]. bronze [5304]. Brookhaven [4177, 4169, 4175, 4176]. Brooks [531]. Brother [1799]. Brouhaha [4175, 4176]. browsers [3881]. bruise [4254]. Brush [4734, 831]. Btfsplk [1740]. Bubble [2283, 2279, 4159]. Bubbles [4151, 3280, 4158, 4774]. Buck [3990, 3921]. Buckeyballs [450]. Buckyball [1593]. Buckyballs [3590]. Buckytubes [816, 1186]. Buddhist [3839]. Budding [3006]. Budgetary [2268]. budgets [2933, 3863]. Bufo [250]. Bug [4415, 4744, 1127, 3205, 5008, 374, 4413, 4641]. Bug-Eyed [374]. Bugs [3379, 941, 4867, 2564, 4510]. Builder [991]. Builders [4593]. Building [4066, 4082, 4585, 3930, 3947, 2443, 4545, 1224, 4065, 1055, 1061, 5367, 2117, 2108, 4069, 4744, 4035, 925, 438, 3382]. Buildings [291, 2246, 4068, 4086]. Builds [3700]. Built [5380, 30]. Bulldog [2017]. Bullet [3670].

Bullet-resistant [3670]. Bullets [3669]. Bumblebee [2472]. Bundling [1491, 1503]. Bunny [4307]. Burger [1495]. Burial [4145, 49]. Buried [1877]. Burke [2796]. Burn [2708, 3912]. Burning [4042, 639, 4554]. Burns [2132]. Burst [131, 4771]. Bursts [3842, 819]. Burying [4124]. Bus [1476]. bushel [4009]. BUSINES [3147]. Business [4614, 1146, 1182, 1225, 1327, 1365, 1390, 1420, 1459, 1496, 1534, 1571, 1606, 1642, 1685, 2150, 1762, 1800, 1838, 1875, 1915, 1955, 1992, 2031, 2071, 2111, 2205, 2755, 2692, 2529, 2825, 2252, 2376, 2382, 2463, 2509, 2587, 3311, 3443, 3047, 2934, 3198, 3394, 2984, 3261, 3089, 4009, 3634, 3540, 3586, 4053, 4464, 4137, 4360, 4562, 4413, 4184, 4510, 4299, 4236, 4908, 1008, 1072, 1641, 2149, 2070, 1914, 2268, 2481, 2603, 2708, 3266, 4855, 159, 23, 90, 124, 194, 265, 626, 521, 697, 835, 661, 1224, 970, 936, 1181]. Business [4904, 2840, 2801, 301, 373, 804, 484, 558, 871, 1036, 1145, 1109, 1292, 1570, 1605, 1533, 1874, 2220, 2348, 4191, 2110, 2353, 56, 4637, 5094, 1235, 231, 340, 410, 736, 591, 770, 446, 1004, 903, 1326, 1458, 1419, 1495, 1684, 1364, 1954, 1799, 1837, 1761, 1991, 2030, 2307, 4735, 4248, 1188, 2980, 2891, 3729, 4689, 411, 485, 522, 592, 2666, 3690, 3832, 3732, 3779, 3875, 3922, 4306]. Businessman [1351]. Bust [5234]. Busters [196]. butter [2862]. butter-side [2862]. Butterflies [3561, 3108, 4473, 3847]. Butterfly [3846, 282, 5167]. Buy [975, 2055, 2837]. buzz [3467]. bye [2692, 1956]. Byte [628]. Bytes [664, 470].

C [2932, 4493, 4954, 5246, 4953]. Cabin [3149]. Cable [3538, 5249, 5250, 5331]. Cable-Free [5331]. Cable-modem [3538]. cables [3901]. Cactus [4684]. Cadets [1954]. Cadmium [594]. Café [5238]. Cage [1127]. Caiman [4260]. Calculate [3366]. Calculating [5198, 445, 444, 5169, 2591, 439]. Calculation [1916]. Calculus [4073, 4376, 4375]. Call [561, 976]. Calls [315]. Caloric [2900]. Cambrian [1250]. Came [5090, 1124, 3574, 3612, 4336]. Camel [1043, 1032]. Camera [335]. Cameras [4441, 931, 3447, 4442]. Camp [2593]. Campaign [5283, 2067]. Can [550, 3309, 3288, 3521, 1826, 1507, 2725, 2724, 2059, 698, 4927, 4825, 1478, 1989, 1988, 2062, 3974, 1542, 4826, 3456, 3696, 5341, 3168, 3166, 2055, 30, 3705, 3556, 4476, 4793, 3612, 4781, 4783, 4613, 4734, 4531, 1985]. Can-do [3521]. Canal [1686, 4881]. Cancer [3355, 3875, 4803, 3358, 337, 1626, 1756, 1758, 3795, 1286, 1382, 2315, 2725, 2724, 2701, 2886, 3774, 3359, 3354, 3341, 1867, 4742, 3879, 4572, 3331, 2955, 3340, 3344, 3343, 3357, 3356, 3318, 4796, 3214, 3220, 2175, 894, 3329, 3171, 3177, 3352, 3351, 3353, 3110, 3112, 162, 3319, 3320, 157, 3324, 4804, 3336, 3335, 3327, 3328, 3322, 3323, 3330, 2749, 3394, 4800, 4348, 4562, 4401]. Cancers [3347, 3346]. Canines [1051]. Cannabis [322]. Cannot [2661]. Can’t [142]. Canyon [3551]. Cap [525]. Capelin [168]. capitalist [2931]. Capture [337, 4611]. Captured [3064, 3062]. Captures [335]. Capturing [3606, 4936]. Car [2516, 2196, 2190, 1004, 3730, 3679]. Carbohydrates [1317]. Carbon [4137, 537, 3094]. Carborundum [2036]. Card [851]. Cardiac [1525]. Cards [3269]. Care [3759, 983, 5002, 1204, 1212, 3360, 3745, 2360, 2366, 1975, 4147, 4865, 1539, 1881, 4476]. career [3873]. Caribbean [679, 3012]. Carry [2577, 3901]. Cars [3729, 2820, 2929]. Carson [4205, 4215]. Cart [3683, 4053]. Cartographer [1059]. Cartography [306]. Cascade [1279]. Case [979, 1670, 4653, 4061, 5215, 4660, 3450]. Cases [3152, 4804]. Cash [1822, 2543]. Cashcowosaurus [128]. Cassini [5194, 5195]. cast [3639, 4040, 5157]. Casualties [1662]. Cat [296, 4073, 3799, 4074, 5224, 3798]. Catalysis [1412]. Catalyst [1576]. Catalysts [965]. Catch [3555, 2540, 2524]. Catcher [162]. Catches [335]. Catching [3847, 4241, 935, 933, 2346, 3162, 3690, 930]. categorization [3095]. Caterpillar [315]. Caterpillars [1175]. Cathedral [438]. Cathodes [342]. Cats [3805]. Caught [4023, 4652, 4696, 4697]. Cauldron [932]. Caulerpa [2193]. Cause [2725, 2724, 4792, 5319, 4804]. Caused [331, 330, 332, 2701, 2763, 3484]. Causes [3326, 228, 3327, 3328, 4476]. Cautious [466]. Cave [5325]. Caveman [3911]. Cavity [1409]. CD [2794, 2668, 3512, 3125, 3466, 3610, 4590, 2287, 3218]. CD-ROM [2794, 3125, 3466, 3610]. CD-ROMs [2668, 2287]. Ceiling [4731]. Celebrating [4598]. Cell [4042, 3497, 3503, 827, 894, 511, 3505, 1560, 1317, 2063, 4299, 4236]. Cells [5141, 3930, 3947, 5126, 5140, 2023, 1714, 799, 2218, 5066, 2490, 862, 5021, 5023, 5016, 5018, 631, 1203, 3006, 5128, 400, 1448, 3054, 3056, 3875, 4398, 3055, 3741]. Cellular [32, 1251, 4320]. Cementing [4379]. Cemetery [4903]. Censor [3308]. Censorship [3044]. Census [4450, 952]. Cent [5098]. Centenary [3898]. Center [647, 120, 3820]. Centers [363]. Centrifugal [1285, 1381]. Centrifuge [4157]. Centrosome [1486]. Century [2646, 4529, 4534, 5296, 5349, 4769, 3606, 2913, 4268, 2629, 2623, 2636, 2650, 5303, 2921, 5350, 4323, 723]. CEO [3921]. Cepheid [1231]. Ceramic [2486]. Cerebrospinal [2136]. CERN [5121]. Certificates [4655, 4656]. Cesar [4086]. Cetacean [4846]. CFCs [4007, 4464]. CH [2343, 2343]. Chain [121, 3419]. Challenge [4706, 3598, 3652, 4257, 4073, 4707]. Challenged [1903]. Challenges [1243, 1704, 4483, 5246, 3357, 3356, 1242, 5028, 1702]. challenging [4616]. chambers [3415]. Chameleons [5178]. Champion [1520, 3083]. Chan [576]. Chance [2923, 2002]. Chances [4807]. Change [1507, 4969, 1788, 3442, 3962, 1352, 3455, 512, 771, 5171, 5190, 2934]. changes [4786, 4801]. Changing [3202, 2879, 1484, 3026, 2354, 486]. chango [3558]. Channel [771]. Channels [596, 433, 919]. Chaos [3641, 5308, 697, 632, 1210, 1161, 1563, 4904, 3642, 4737, 54, 53, 861, 1578, 1569, 2301, 877, 868, 1818, 2585, 2982, 3311, 696, 1260, 1267]. Chaotic [2780, 1818, 3851, 4735]. Characteristics [1300, 4758]. Charge [1864, 3890, 3891, 3730]. Charges [594]. Charging [3964]. Charles [4086, 2782, 3410, 3169, 3172, 1361]. Charlotte [970]. Charting [466]. chases [4453]. Chasing [1502, 5106, 5105, 3288]. Cheap [4264, 1111, 1627]. Cheaply [3636]. Check [2179, 1937, 1113, 3693]. Checkup [3539]. Cheers [4388]. Chemical [3009, 2313, 3014, 1674, 830, 3743, 4801, 3015, 4175]. chemist [4007, 3911]. Chemistry [3178, 1866, 3699, 1401, 3526]. Chemists [1355]. Chemoprevention [3331]. Chesley [1908, 1907]. Chess [2721, 325, 314, 324, 2157, 2146, 2727, 2728, 3083]. chessboard [3707]. Chestnut [238]. Chew [1746]. Chewing [3535]. Chicken [2682]. Child [4147, 4865, 4074]. Children [2144, 5181, 4481, 4820, 4076]. Chiller [1747]. Chilling [3541, 2270]. China} [4082, 3839, 3138, 3820, 3895, 3635, 290, 295, 4069, 3456, 3815, 1492]. Chinese [3634]. Chip [5164, 663, 3312]. Chip-Making [663]. Chipmakers [4040]. Chips [3541, 4010, 3732, 4184, 4191]. Chloride [131]. Chocolate [4662]. Choice [4715, 4716]. choke [3430]. choking [4634]. cholesterol [3690]. Choose [4317]. Choosing [4808]. Chornobyl [3052, 3053]. Christian [5122]. Christmas [406, 1260, 1267]. Chromoskedasic [796, 795]. Chronic [3960, 2069, 2066, 2068]. Chronicled [1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 2207, 2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623, 4675, 4719, 4838, 4889, 4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317]. Chronologically [1645]. Chub [41]. Chump [5171, 5190]. Chunnel [428, 1805]. Cicada [5076]. Cicadas [4368]. Cichlids [4920]. Cigarettes [4804]. Cigars [1551]. Circa [1975]. circle [4527]. Circuit [4107, 3938]. Circuits [1578, 1569]. circulating [3311]. Circumnavigators [5040]. Cities [703, 1947, 516, 518]. Citizen [1123, 1162, 1200, 1306, 1343, 1379, 1398, 1438, 1475, 1514, 1550, 1622, 1661, 1741, 1779, 1817, 1855, 1892, 1935, 1972, 2010, 2050, 2089, 2127, 2168, 2526, 2210, 2294, 2337, 2381, 2422, 2468, 2586, 2689, 2874, 3302, 2933, 3140, 3383, 2983, 3034, 3433, 3083, 3193, 3247, 3999, 3575, 3624, 3531, 4042, 4287, 4294, 4350, 4453, 4504, 4172, 4127, 4554, 4227, 4401, 4895, 3, 277, 313, 107, 140, 536, 782, 1088, 949, 983, 1199, 1586, 1854, 1816, 2467, 2293, 2209, 4627, 1161, 428, 177, 2167, 2379, 2575, 4178, 352, 815]. Citizen [915, 1018, 1513, 1621, 1397, 1934, 212, 386, 39, 70, 639, 498, 675, 572, 883, 1122, 1277, 2088, 1891, 2744, 2680, 2421, 4845, 4235, 4679, 1049, 847, 1342, 1740, 1778, 2009, 2048, 2126, 709, 747, 1305, 1437, 1549, 1971, 606, 1474, 462, 2813, 2930, 2878, 2519, 246, 1659, 2980, 2256, 2336, 4724, 387, 463, 499, 573, 640, 1587, 3769, 3823, 3679, 3722, 3865, 3911]. City [4638, 879, 3291, 1611, 2365, 5272, 232, 5196]. Civil [735, 727]. civilization [4710]. Civilizations [3075]. Civilizing [3617]. Clamp [926]. Clash [1556]. Class [951, 3741]. Classical [1944]. classroom [4553]. classrooms [4785]. Claude [4135, 16]. Claustrophobia [5282]. Clean [3448, 1921, 1708, 3991, 2626]. Cleaning [589, 590, 4907, 2903, 583]. cleanups [4908]. Clear [4, 7, 718, 3748]. Clearing [393, 2772, 1630]. Clementine [2201, 2011]. Cliffs [2185]. Climate [2780, 4969, 3742, 5340, 5357, 1484, 3657, 3672, 224, 2934]. Climatic [1788, 3455, 512]. Clinton [1478]. Clipper [1574]. Clock [2435, 4295, 251]. Clockmakers [4540]. Clone [5024, 3716]. cloned [3720]. Cloning [4137, 3797, 4740]. close [2795, 3245, 3433]. Closer [1856]. Closing [4247, 4248]. closure [4611]. Clot [630]. clothes [2931]. Cloud [1802, 2549, 4580, 4589]. Clouds [5036, 222, 618]. Clues [3110, 3112, 3754]. Clump [394]. Clumpiness [887]. Clusters [4048, 4059, 4060, 4739, 3318, 3329, 4775]. CO [848, 859, 4125, 124]. CO2 [3443]. Coal [589, 583, 590]. coastal [4613]. coastlines [3807]. Coat [3265]. COBE [1059, 1345]. Cocaine [3593, 650]. Cocconi [3730]. Cochran [4358]. cocoa [3516]. Code [3020, 5338, 560, 1364, 3694, 3695, 3896]. Codex [1028]. Coevolution [443]. Coffee [3810, 3728, 3752, 3811, 3073]. Cog [4855]. Cognitive [4369]. Coincidence [4432]. Cold [536, 2134, 4077, 4187, 1257, 3879, 990, 1916, 1823, 44, 342, 2930, 2534, 3875, 3566, 2384, 3444]. Coli} [4681, 1951, 1961]. collect [2906]. collecting [3847]. Collection [3024]. College [5237]. Collider [227]. Colliding [685, 5382, 1054]. Collins [4180, 4190, 4188]. Collisionless [551]. Collisions [3003]. Color [800, 4220, 810, 1561, 1562, 1707]. Colorblind [4790]. Colorectal [2886]. Colored [5035]. Colorful [4610]. coloring [3893]. Colors [3288, 3265]. Colossal [2956, 4771]. Columbian [898]. Columbus [1192, 1261, 1221]. Colwell [4732]. Combating [4742]. combinations [3299]. Combinatorial [3699]. combined [4804]. Combustion [526]. Come [5126, 5335, 5052]. Comes [3470, 5111, 1102, 1101, 3245]. Comet [3555, 3605, 3606, 3556, 2557]. Comets [4768, 3003, 2811]. Comic [4297]. Coming [2348, 3742, 5011, 2777, 2342, 4681, 2534, 3162]. Command [3874]. Commanding [2260]. Commentaries [4594, 2654, 2797, 2865, 3466, 2968, 3019, 2911, 2920, 2967, 3018, 3131, 3183, 3237, 3367, 3418, 3465, 3610, 3559, 3804, 4076, 3560, 3851, 3664, 4028, 3663, 3708, 3751, 3850, 3894, 3939, 4027, 4205, 4539, 4214, 4216, 4276, 4382, 4383, 4434, 4435, 4490, 4491, 4547, 4664, 4709, 4749, 4941, 4940, 5038, 5147, 5186, 5220, 3988, 2867, 4326, 4278, 3287, 4980, 4979, 4493, 3609, 5347, 5109, 4215, 4277, 4160, 5310, 2799, 2863, 3515, 4075, 4335, 4327, 4588, 4875, 4876, 5260, 2793, 5074, 4663, 4161, 4438, 3752, 3805, 3940, 3709, 3895, 3178, 3368]. Commentaries [3288, 3072, 3125, 3230]. Commentary [2638, 2631, 2621, 2646, 2619, 5263, 2652]. commerce [3267, 4015, 3020, 2802, 4875, 2692]. Commercial [2627]. Common [30, 734, 2599, 4546, 5119, 725]. Communicate [3597]. Communication [2856, 2555, 2559, 2556, 2188, 2194]. Communications [720, 867, 4322, 4934, 5178, 1564, 1567, 1831, 1835, 86, 96]. Communism [2931]. Community [3318, 3329]. commuter [3586]. Compact [3215, 3216, 4930, 4932, 5387]. Companies [1368]. Companions [2716]. Compendium [367]. Compensation [976]. compete [3621]. Competition [1842]. Competitiveness [457]. Complete [526]. Complex [1469, 3054, 3056, 4151, 3055, 4610, 4158]. Complexifying [2595]. Complexity [2453, 2445, 2452, 1325, 1320, 2673, 3125, 3752]. Complicated [4832]. Composite [5372, 4249]. composites [4236]. Compounds [928]. Comprehended [322]. Compton [1720]. Compulsive [1051]. Computation [5179, 2479, 191, 4559]. Computational [4827]. compute [4129, 4142, 3586]. Computer [442, 3458, 3832, 4650, 4498, 3257, 4654, 5032, 23, 34, 4651, 4656, 3620, 4657, 4141, 4522, 4842, 4020, 5069, 724, 2363, 2362, 13, 19, 3463, 975, 1361, 2962, 723, 5177, 3688, 4512, 4230, 4735]. Computerese [1582]. Computerized [2368, 2361]. Computers [732, 2150, 4102, 4464, 2319, 725, 720, 3845, 4828, 726, 475, 479, 3973, 727, 731, 2717, 4115, 3405, 729, 728, 1154, 1120, 3394, 3737, 4564, 4510, 4878, 5384]. Computing [4523, 4530, 3858, 439, 444, 5176, 5174, 4425, 2066, 2068, 2437, 3845, 4737, 1856, 2454, 2447, 2141, 2143, 722, 3787, 2580, 3492]. Con [3426]. concealed [3695]. Concentration [769, 776]. Concrete [1418, 1417, 1411]. Condensate [4255]. Condensed [2135]. Condensed-Matter [2135]. Conditioners [5189]. Conducting [4606]. Cone [5258]. confession [3366]. Confidential [2856]. Confidentiality [4413]. Confidentially [4416]. Configurable [3787]. Conflict [1352]. Confocal [2021, 2020]. Confronting [3411, 4781, 1825, 3052, 3168, 3124]. Confusing [885, 2839, 3125, 3752]. Confusion [1823, 4042]. Congestion [3974]. Congo [5196]. Congress [2755, 4685]. Conjecture [4488]. Connection [50, 340]. Connections [2864, 3467, 3369, 3020, 2969, 3420, 3179, 2912, 3073, 3516, 3231, 3126, 3289, 3561, 3852, 4077, 3665, 3896, 3806, 3753, 3941, 4029, 3612, 3710, 4165, 4326, 4540, 4436, 4668, 4206, 4219, 4587, 4492, 4713, 4384, 4336, 4751, 4281, 4265, 4596, 4946, 5151, 5352, 5222, 5305, 5079, 4880, 5043, 4881, 5113, 5267, 4985, 5330, 2796]. Conquer [3498]. Conscious [4855, 2853, 4089]. Consciousness [4545, 1144, 4088, 1989, 1985, 1988, 5285, 5294, 2932]. consensus [4450]. Consequences [5228, 4797]. consider [4617]. Considered [3345, 4815]. Constants [4678]. constellations [4588]. Constraints [2923]. Construction [2359, 852, 620, 4948]. Consumer [897, 901]. Consuming [4857]. contain [4774]. contamination [2683, 4875]. Contemplates [649]. Contented [1607]. Continental [89, 2905, 5222]. continue [4476]. Continued [3250]. continues [3311, 4285]. contraception [4778]. Contraceptives [2637, 3201]. Contrary [3401]. contributions [4122]. Control [335, 2313, 3883, 3307, 3874, 213, 1810, 654, 2274, 1031, 3561, 3884, 4613]. controlled [2562]. Controlling [4332, 3359, 4527, 4532, 3405]. Controls [511]. controversial [4892]. controversies [4122]. Controversy [2673, 3332, 2732, 2527, 3499]. Conundrums [1328, 3720]. Convention [3200]. Conventional [3342, 3343, 1588, 4531]. Convince [2248]. Conway [5005]. Cook [429]. Cooking [1039]. Cookstoves [2494]. Cool [4364, 1854, 3855, 145, 1226]. Coolants [4463]. Cooled [3989]. Cooling [2933]. Cooperation [3738]. cope [4781]. Copernican [3665]. Coping [4485, 876]. Copper [131, 656]. Copper-Alloy [656]. copter [3586]. Copyright [2270]. Coral [1315, 2706]. corals [4610]. Cord [3485, 5214]. Cordiale [5078]. Core [1710, 1447, 3383]. Cores [4204]. Corners [1662]. cornflakes [2912]. Corporation [729]. Corpus [1330]. Corral [2474]. Corroborated [1345]. Corruption [709]. Cosmic [2983, 2201, 3546, 4770, 1059, 5148, 4849, 5202, 1480, 1718, 3301, 1169, 927, 4313, 3767, 4227, 4845, 4235, 31]. Cosmological [3302, 4144, 4863, 4773]. Cosmologist [2687]. Cosmologists [4772]. Cosmology [1099, 4861, 825, 329, 338, 1049, 1664]. Cosmos [803, 797, 1457, 1449, 5034, 2566, 1400, 3723, 802, 1456, 5074]. Cost [546, 2567, 3690]. Costs [5002, 2674, 3805]. Cotton [3395, 5035, 5352, 1841]. Couch [3195]. Could [5089, 18, 20, 2802, 4230, 5208, 737, 4125]. Count [504, 2114, 4450]. counterfeiting [2736]. Counters [2224]. Counting [5218, 1090, 4717, 4593]. Countries [1719]. Country [247, 2768, 1960]. counts [2565, 3284]. County [2701, 2763]. Coupled [1726]. Course [1397]. Court [3956, 1082]. Courtroom [1045]. Cousins [46]. Cousteau [3914]. Couture [4988]. Cover [3437, 2589]. covered [4603]. Covert [3227, 3234]. Cow [355, 3478]. Cowpokes [2054]. Cows [3506, 149, 3513, 1607, 3514]. Crabby [4401]. Crabshoot [2076]. cracker [3695]. Cracking [3694]. CRADA [1641]. Cradle [4073, 4074]. Craig [4506]. Cranes [4947, 4948]. craps [4026]. Crash [3912, 4975]. Crater [4882, 4883]. Cratering [122]. Crawler [4947]. Crawling [2063]. Crayons [3667]. Create [32, 3743]. Creates [5342, 5356, 5343, 3895]. Creating [3932, 3502, 3180]. Creation [825, 4846, 5149]. Creationism [2805]. Creative [1624]. Creativity [2275, 4091]. Creature [3940]. creatures [4616]. credit [3664]. Creeps [1856]. Cretaceous [2280]. Cricket [3508]. Crime [2831, 4250, 4252, 4253, 5318, 1347, 3162, 2819, 2514, 3154, 2873, 5319]. Crimes [4959]. Criminal [2322, 2316, 2321, 4098]. criminologist [3154]. cripple [4779]. Crisis [2069, 2066, 2068, 5247, 2424]. critical [3198]. Criticality [436]. Criticism [358]. critics [3201, 4028]. Cropland [3387]. Crops [1027, 4531, 4537, 3291, 5103]. Cross [821]. Crossed [752, 5108]. crowded [2872]. Crowning [2815]. Cruising [3147]. Crumbling [1284, 1380]. Crun [4261]. Crunch [4262, 4263, 3489, 2989]. Crunching [1309]. Crusader [5059]. Crust [89, 2905]. Cryptography [1172, 4659]. cryptology [3709]. Crystal [3385, 1605, 3748]. crystal-clear [3748]. Crystallography [3607]. Crystals [1055, 1061, 2304, 4034]. cube [2530]. Cubes [4010, 3122]. Cuckoo [443]. Culprit [951]. Cults [1727]. Culture [2874, 3287, 1556, 2878]. Cultured [4398, 799]. Culturing [4397]. Cup [4464]. Cupholders [5329]. Cure [2241, 2232, 2240, 1474, 2352]. Cures [4476, 4988]. Curiosity [5224]. Curitiba [3002]. Current [873, 3332, 3344, 3343, 2017, 3859, 1729, 1728, 1723]. Curve [2209, 4819, 2372, 2285, 5202]. Customers [3023]. Customized [2626]. cutbacks [3863]. Cyanobacteria [1752]. Cyber [3088, 3044, 3195, 3309, 3145, 3258, 3392, 3441, 3927, 4058, 3695, 3785, 3881, 3639, 3592, 3737, 3545, 4015, 3838, 4420, 4366, 4564, 4129, 4421, 4254, 4312, 4512, 4914, 4859, 4310, 4738, 4418, 4641, 4252, 4142, 4197, 4693, 4967]. Cyberspace [2711, 4469, 4302, 4303, 735, 727, 2606, 5361, 1568, 1564, 1567, 1954, 4967, 2694]. Cycle [961, 511, 3844, 750]. Cycles [22]. Cygni [4769, 2236, 2238]. Cystic [2847].

D [3584, 666, 4011]. D. [3560]. D2 [954]. Dad [47]. Daily [3509]. Daisy [2237, 2244, 1444]. Dam [3868, 3831]. Damadian [3777]. Damage [1636, 3679]. Damaged [5214, 4398]. dammed [3865]. Damnable [4554]. Damned [1767]. Dan [3688]. Dance [1949, 1038]. Dances [4355, 5219]. Danger [2817, 751]. Dangerous [2951, 2590, 2256, 4410, 4527]. Dangers [98, 3368]. Daniel [2932, 4493]. Dare [2665]. Dark [3454, 3469, 5264, 4234, 1744, 2340, 4118, 4775, 3872, 246, 4235, 2396]. Darkly [150]. Darkness [3279, 3285]. Darling [2011]. Darwin [3410, 2932, 3125, 3909, 3752, 2467, 764, 2782, 2017]. Darwinists [2726, 2719]. Darwins [3999]. Data [3013, 5248, 3025, 3024, 3379, 1757, 641, 940, 3540]. databases [3145]. Date [2263]. Dating [3458, 3148, 3463]. David [1481, 719]. Day [4799, 4696, 4697, 3300, 4770]. Daydreaming [2094]. Days [4606, 3407, 4608, 246]. DDT [2740]. Dead [4242, 4243, 3510, 4527, 4452, 3637, 3997, 4248, 361]. Deadly [3477, 4507, 4783]. Dealer [76]. deals [2821]. Dear [2296, 3868]. Death [2428, 3996, 5123, 4955, 179, 212, 498, 1639, 1635, 1640, 2129, 1011, 998, 1727, 2042, 5201, 4609, 1594, 4876, 1446, 2236, 2238, 2690, 4028, 3911]. Deathbed [2211]. Deaths [2701, 2763, 3484, 3582, 4187]. Debate [17, 330, 957, 1669, 3762, 224]. debating [3940]. Debris [4527, 4532]. Debt [2442, 2707, 2681]. Decade [4890, 4310, 4311]. Decaffeinating [3810, 3811]. Dechronization [5211, 5212]. Deciphering [5338, 2449, 2448, 2175]. Declassified [1199]. Decline [5326, 1719]. declines [4724]. Declining [2356, 4611]. Decompression [2554]. Deep [3528, 3996, 5090, 3402, 2064, 4349, 3865, 4617]. deep-sea [4617]. Defeating [4476, 4475]. Defeats [2552]. Defend [5346]. Defender [5204]. Defending [3522]. defends [3584, 4859]. Defense [3394, 3690, 1717, 1787, 2851, 2857, 356, 5172]. Defensive [1151]. Deferred [5322]. Defibrillation [4428]. Defibrillator [4429]. Deficiency [759]. Deficit [4576, 4581, 3245]. defies [3095]. Defining [2512, 1057]. Definitions [1708]. deflate [3302]. Deforestation [119]. Degradation [2320]. Degree [5266]. Degrees [56]. Déjà [979, 5231]. Delayed [758]. deliberately [3909]. Delight [4281, 4282]. Deliverance [1709]. Delivering [2933]. delivers [4269]. Delivery [1983, 3875]. Deluge [4631]. Demand [4320, 5053]. Democracies [1695]. Demographic [3154]. Demolition [2718]. demon [4895, 4894]. Demonic [1007]. Demonizing [3785]. demonstrate [4611]. Demure [1442]. demystified [2862]. Dendrimer [2401, 2405]. Denied [2467, 4525]. Dennett [2932, 2951, 4493]. Dennis [4906]. Density [1864, 4864, 4301]. Department [2692]. dependable [3490]. Depletion [618]. Depressed [4786]. Depression [2426, 4092, 4422]. Depressive [2275, 4091]. depth [4154, 4167]. Deregulated [1426]. Deregulation [3043, 1577]. Derivatives [2219]. Dertouzos [3830]. Descartes [2121]. describe [4476]. Desert [1209, 2983]. Deserting [5235]. Deserts [1833]. Deserves [2596]. Design [1784, 1724, 1229, 620, 1789, 370]. Designer [5378, 4644, 3730]. Designers [449]. Designing [2657]. Desktop [971]. Desperados [281]. Desperate [1977]. despite [3499, 4285]. Despoiled [3486]. Details [1917]. detect [4081]. detected [4504, 4756]. Detecting [3121, 3278, 3129, 3283, 5036, 5145, 2561, 2558, 5180]. Detection [3334, 4680, 3336, 3335]. Detectives [5153]. Detector [2407, 3284]. Detectors [3714, 2548]. detonations [4348]. Detroit [3730]. Deuterium [3501]. deux [4265]. Develop [4922]. Developers [1147]. Developing [1980, 2621, 2614, 2494, 294, 1719, 1134]. Development [2952, 5181, 2138, 3272, 3268, 1193]. Developments [942]. Develops [1595]. Device [4497]. Devilish [1917]. Diabetes [228, 5086, 2490, 1167, 4637]. Diagnosis [1187, 1480]. Diagnostic [1114]. diameter [4773]. Diamond [1177, 4997]. Dice [4025]. Dicey [4026]. Dickens [406]. Dictatorship [2178]. Did [5319, 5157, 4627, 4819, 2412, 2411, 957]. Diet [1565]. Difference [3809, 2339, 323]. Differences [5122, 915, 1140]. Different [1183, 3438, 4774, 4613]. differently [4799]. Digesting [2816]. Digestive [3691]. Digital [695, 3288, 3261, 4230, 4231, 4655, 4656, 3791, 2658, 3645, 4372, 4441, 3223, 3229, 2227, 704, 1648, 281, 757, 2665, 3230, 4442]. Dignity [498]. Dilemma [4230, 4231, 2454, 2447]. Dilemmas [4485, 1867, 1832, 2872]. Dim [1092]. Dimensions [365, 372, 4103, 371]. Diminishing [371, 372, 4103, 365]. dimpled [3608]. Dining [5187]. Dinosaur [4063, 3072]. Dinosaurs [555, 2427, 1523, 3575, 3940]. Dioxide [3094]. Dioxin [543, 1748]. Dirac [1455]. Directed [1259]. Directional [1909, 3823]. Directly [3417]. Dirt [3877]. Dirty [3881, 2384]. Disabilities [4369]. disagreements [4230]. Disappear [3974]. Disappearance [908]. Disappearing [3722, 4242, 4243, 4401]. Disappointing [3135]. Disarming [942, 2058, 2060, 4149, 4866]. Disaster [809]. Discerning [3049, 5121]. disclaimer [3665]. Discord [3048]. discouragements [4285]. Discourse [1083, 1014]. Discovering [3653]. Discovery [1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 2207, 2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623, 4675, 4719, 4838, 4889, 4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317, 1950, 2787, 4508, 3929]. Discovery [3654, 4150, 3655]. Discreet [908]. Discretion [3880]. Discs [3215, 3216, 5387]. Discussion [574]. Disease [2763, 3960, 3497, 3503, 799, 4806, 962, 1905, 2058, 2060, 1026, 4701, 3505, 4022, 794, 790, 1600, 4087, 3882, 3547, 2752, 3478, 4800, 4811]. Disease-Resistant [4022]. Diseases [472, 4783, 1598, 2228, 4789]. disguise [2805]. Disinfectants [2072]. Disk [3215, 3610]. Disks [863, 2529]. Disliking [3784]. Dismay [747]. dismembered [4527]. Disorder [4576, 4581, 3245]. Disorders [1141, 4790, 4779]. Dispersal [657, 260]. Display [2304, 4034]. Displays [1287, 1384, 3102]. Disposable [467]. Dispute [434]. Disruption [922]. Dissecting [3504, 3511]. Disservice [3591]. Dissidents [1249]. Distant [1485, 4757]. Distinguished [3095]. Distributed [2141, 2143]. Disturbance [677]. Dive [4003]. diver [4604]. divers [4487]. diverse [4759]. Diversions [32]. Diversity [2013, 1755]. Divide [1714]. Divided [3738, 2687]. divine [3896]. Division [4874]. Divorce [4961]. DNA [4523, 4530, 2580, 4584, 4506, 4127, 4390, 1052, 676, 1187, 2591, 4155, 4868, 1681, 1389, 1291, 1383, 3882, 21]. Do [2237, 3072, 4083, 1469, 233, 676, 3627, 607, 2661, 5239, 4070, 1258, 1266, 3537, 3082, 3521, 1626, 4793, 1464, 2439, 1265, 2244]. Do-It-Yourself [5239]. Doctor [254, 4305, 2939]. Doctors [4632, 2933]. Documents [666, 2227]. Dodecahedra [5219]. Does [1669, 4668, 4669, 1014, 3340, 2860, 3380, 4431]. Dogma [4677]. Doing [2801, 3563]. Dollar [2433]. Dolly [3720]. Dolphins [3280]. Domain [2351, 3966]. Domesticating [1568, 1567, 1564]. Donald [1323]. Done [284]. Don’t [3282, 4128, 3868, 771, 2668, 4299, 4429, 1344, 4014, 4354]. Doomsday [1114]. Door [5156, 2807]. Doors [3930, 3947]. Dope [4573]. Doppler [4436]. Dot [1660]. Dots [1322]. Double [3230, 4966, 4012, 391, 2426, 4092, 4386, 4387, 612, 4159]. Double-Gate [4966]. Double-Sticky [4012]. Doubling [1079]. Doubts [3823, 5157]. Doughnut [488]. doughnuts [2792]. Down [2933, 2539, 78, 2547, 161, 267, 4680, 4365, 1478, 1090, 3766, 1819, 5278, 504, 2345, 4691, 3480, 2862, 4603, 4921]. Downloading [4252, 4253]. Downsizing [4360]. Downtown [3140]. Dr [3466]. Dr. [2822, 395, 1799]. Dr.P [4777]. Drag [4457, 939]. Dragon [4974]. Dragons [4743]. drags [4453]. Drain [2645, 949, 2196, 2190, 2662, 3480]. Draw [2390, 3550]. Dream [4536, 744, 1803, 4604]. Dreams [3518, 5057, 366, 4093]. Dreamtime [2259]. Dreck [2880]. dredge [2578]. dredgers [3231]. Dredging [3691]. Dressed [5020, 3670]. Dresses [4219, 4206]. Dressmaker [1540, 1532]. drift [5222]. drifting [4527]. Drill [360]. Drilling [1909]. Drink [5240, 2055]. Drinking [4005]. Drinks [47]. Driven [5327]. Drives [22, 4722]. Driving [231, 774, 1151, 3975]. drop [5319, 2802]. Dropping [3581]. Drug [1950, 4988, 1983, 4479, 3549, 3443, 4476, 4679, 4689]. Drugnet [3690]. Drugs [1724, 1794, 24, 4821, 2312, 3699, 2114, 3299, 3261, 3634, 4954, 4624, 4724]. Dry [815, 3146, 3374]. DSL [5251, 5252]. Duality [2191]. Dubos [582]. Dugong [4558]. dumping [4175]. dumps [3481]. Dunhuang [3839]. during [4795, 4604]. Dust [3507, 196]. Duty [976]. Dwindle [213]. Dying [4779, 2392]. Dynamics [1209, 1097, 1832, 1320]. Dynamo [2215, 3270]. Dyslexia [2526, 784, 1057, 3459, 2519, 757]. Dyson [1557].

E-Mail [984, 1886]. E. [4438]. Ear [4984, 2985, 2983]. Earcons [1536]. Earful [373]. Earlier [3334, 3823, 3335]. Earliest [4315, 4316, 5139, 5138, 1318, 3897]. Early [1025, 101, 883, 1555, 3788, 620, 1947, 4886, 651]. Ears [1410, 2032]. Earth [2101, 1194, 2687, 3383, 4762, 4287, 4294, 4453, 5195, 2539, 1827, 3289, 5145, 4873, 4872, 4770, 2230, 286, 288, 2187, 2186, 3609, 3402, 2103, 122, 4603, 2109, 2100, 1171, 3885, 2107, 2104, 3021, 2102, 2262, 1318]. Earthcam.com [4413]. earthquake [4287, 4294]. Earthquakes [2064, 2850, 89, 3069]. Easier [3540, 284, 2263]. Easing [1399, 187]. East [1093, 1910, 588]. Eastern [290, 295]. easy [3705, 4151, 4158]. Eat [3999]. Eating [4779]. Eavesdropping [1893]. Ebola [2582]. Ebonics [3624]. Echidna [477]. Eclipse [5262]. Eclipsed [5263]. ECM [1689]. Ecolocation [1477]. Ecological [2642, 2652]. Ecologist [3584]. Ecology [2968, 2650, 3406, 3412, 4172, 4178]. Economic [1153, 1370, 2156, 775, 2945, 2981, 4855, 4617]. Economics [1997, 4109, 2899, 2901, 1446, 1263, 4375, 3526]. Economist [2203, 2150, 2814, 2746, 2520, 2681, 2250, 2330, 2374, 2380, 2461, 2507, 2576, 2931, 2981, 2874, 59, 129, 164, 269, 377, 417, 527, 597, 839, 1370, 1577, 1612, 2479, 2602, 2347, 2267, 26, 95, 200, 236, 307, 345, 452, 490, 564, 1539, 2707, 2839, 2219, 2306, 2890]. Economy [64, 925, 2659, 2642, 1502, 2037, 2652, 4539]. Ectoplasm [2592]. Eddington [623]. Eden [1093]. Edge [5364, 1161]. Edison [2723, 2722, 3596]. Editor [1239, 1377, 1618, 1656, 2513, 4989, 4833, 3188, 3426, 3135, 2871, 2924, 2975, 3859, 3522, 3815, 3716, 3905, 3761, 3673, 4717, 4284, 4391, 4340, 4168, 4223, 4671, 4989, 5153, 5191, 5115, 4886]. Editorial [5059]. Editors [1699, 2289, 2333, 2872, 2977, 2927, 4712, 4814, 4878, 4942, 4834, 2570, 4341, 4392, 4444, 4777, 5224, 603, 637, 673, 707, 745, 780, 813, 844, 881, 913, 947, 981, 1016, 1046, 1085, 1119, 1157, 1195, 1238, 1275, 1302, 1339, 1394, 1434, 1471, 1510, 1546, 1584, 1617, 1655, 1698, 1737, 1775, 1813, 1851, 1888, 1931, 1968, 2006, 2045, 2085, 2123, 2164, 2206, 2254, 2290, 2332, 2377, 2418, 2464, 2511, 2568, 2672, 2735, 2804, 2876, 2926, 2976, 3028, 3137, 3189, 3242, 3296, 3376, 3427, 3475, 3762, 3523, 3569, 3618, 3674]. Editors [3717, 3817, 3860, 3906, 3949, 3994, 4037, 4665, 4711, 4119, 4171, 4226, 4286, 4343, 4394, 4446, 4502, 4552, 4622, 4672, 4718, 4951, 4877, 4981, 5039, 5075, 5110, 5261, 5311, 5348, 4835, 4887, 4950, 4990, 5047, 5082, 5116, 5154, 5192, 5225, 5270, 5316, 3080, 4832, 2734, 2803, 3079, 3241, 3026, 3473, 3375, 3295, 3568, 4035, 3617, 3993, 3948, 4443, 4500, 4551, 4621, 4118, 5081, 4949, 5315, 5046, 5269]. Education [1176, 1179, 5247, 4785, 731, 879, 1188, 712]. Educational [4858]. Edward [172]. Edwin [1529]. Effect [4843, 328, 5167, 3005, 4125, 4153, 4853, 2602]. Effects [574, 3856, 414, 4236]. Efficiency [2184]. Efficient [289, 2314]. Egil [2239]. Egypt [3509]. Egyptian [2067]. Einstein [4054, 3999, 4457, 3170, 4255, 3552]. either [4771]. Elastic [2231]. Election [2497, 2500]. Electric [2810, 3492, 3730, 3679, 2398, 3450, 1004, 3970, 3097, 3729]. Electric-car [3730]. Electrical [2170, 3891]. Electricity [4747, 5145, 3883, 3904, 289, 3732]. Electrobalances [3174, 3181].

Electrochemical [5142]. Electrodynamics [1409]. Electromagnetic [2683, 3129, 3121, 5352]. Electron [305]. Electronic [3758, 4360, 4562, 734, 725, 1105, 2802, 373, 1730, 1725, 4283, 2117, 2108, 1349, 2180, 3074]. Electronics [897, 901, 5144, 1678, 1030, 1829, 3937, 3938, 5051, 1673]. Electrons [3011, 3005, 4569]. Electrorheological [1633]. Elegant [5146]. Element [2322, 2316, 2321, 4098, 449]. Elemental [2296]. Elementary [4577, 4583, 687, 688]. Elements [4577, 4583, 2100]. elephant [5352]. Elephants [2183, 2908, 2910, 4895]. Elevators [3985, 3128]. ELF [4335, 4334]. Eliminating [1558]. Elite [4949]. Elkin [3499]. Ellis [2687, 2714]. Eloquent [2026, 1003, 1002, 994]. Elsewhere [3755, 5344]. Elusive [710, 4476]. Embargo [2303]. embraces [3830]. Embryo [3030, 3272, 3268]. Embryonic [5016, 5018]. Embryos [4973, 4972, 3031]. Emergence [2106, 4822, 407]. Emergency [4778]. Emerging [2715]. Emissions [3094]. Emits [131]. Emotion [1945, 4094, 3805]. Empathize [4825, 4826]. Emperor [4245, 3827, 4244, 2660]. Emperors [3938]. Empire [3933, 5346]. Empires [3892, 3937]. Employment [2439]. Empowerment [2299]. Encapsulated [5021, 5023]. Enclosure [1772]. Encoding [719]. encounters [2795]. Encryption [3311, 4660, 2984, 4413]. encyclopedias [3466]. End [5355, 5257, 3768, 4264, 2338, 4642, 2732, 887, 1242, 761, 1805, 1702, 4196, 5339, 2733]. Endangered [2684, 3125, 2703, 2293, 2869, 4293, 2990, 1307, 1928, 2293]. Endgame [789]. Ending [2721, 2764, 2727, 2728]. Endless [789]. Endometriosis [4792]. Endotoxins [1098]. Enemy [3046, 1445]. Energizer [4307]. Energy [2647, 291, 287, 293, 290, 295, 2061, 583, 590, 3546, 4770, 286, 288, 297, 298, 2648, 300, 546, 294, 292, 2196, 2190, 299, 4064, 2872]. Engendering [3042]. Engine [30, 5142]. Engineer [545]. Engineered [1147, 5025, 5026]. Engineering [5364, 5012, 5013, 2641, 2355, 588, 5028, 5381]. Engineers [278]. Engines [4924]. English [778, 1272, 1570]. Enhancement [1816]. Enigma [3477, 2306]. enjoying [4813]. Enormous [3040]. Enough [1544, 2707, 4794]. enrich [1014]. Enriching [4609]. Ensuring [2227]. Entente [5078]. Enter [5206]. entertaining [4941]. Entomophagy [1089]. entrepreneur [3267]. entrepreneurs [4617]. Envelope [3213]. Envelopes [1349]. Environment [1669, 2647, 4424, 2272, 4685, 2442, 4786, 4679]. Environmental [2481, 2725, 2724, 2192, 763, 1527, 1352, 574, 1033, 2320, 1530, 4510]. Environmentalism [1060]. Envisioning [3317, 3316]. Envy [4874]. enwraps [3807]. Enzyme [1971]. Eon [761]. epic [4099]. Epidemic [5274, 3004, 4535, 1859, 2402, 4544]. Epidemics [2136]. Epidemiologist [4410]. Epidemiology [1756, 1758]. epiphanies [4099]. Episode [2280]. Epistolary [1886]. Eponymous [1735]. Epsilon [1309]. EQ [3679, 3682, 1743]. Equal [2769, 3490]. Equation [923, 793]. equatorial [5267]. Equilibrium [83]. Era [472, 4100, 3053, 3052]. erase [2529]. Eraser [2888]. erasers [2874]. Erosion [3387, 3700]. Err [5242]. Errata [243, 349, 381, 175, 459, 534, 779, 845, 880, 912, 946, 980, 1015, 1086, 1433, 1509, 1774, 1812, 1930, 2163, 2253, 3027, 3136, 3816, 4036, 4170, 4342, 4393, 4445, 4503, 4673, 4836, 4991, 4839]. Erratum [274, 425, 604, 636, 672, 1237, 1274, 1301, 1697, 1967, 2044, 2084, 2510, 2875, 2925, 3377, 3474, 4225, 4951, 4888, 5117, 5226, 5271]. Error [2121, 3739, 2968]. Errors [3100]. Eruption [334, 332]. Escaping [2549]. Escher [2983, 2140, 2139, 2985]. Escherichia [1961, 1951]. espionage [2569]. Essay [2004, 2003, 1194, 1811, 2870, 2670, 2288, 2505, 1470, 1469, 36, 979, 2373, 2372, 1616, 1615, 2869, 1508, 1507, 911, 1376, 1375, 2202, 2201, 2459, 2122, 2121, 2083, 2082, 778, 1273, 1272, 1013, 1083, 1735, 1582, 1431, 1432, 1583, 1736, 1965, 1886, 2161, 1887, 1966, 2506, 2801, 2802, 2504, 2287, 3025, 3024, 2162, 602, 209, 1545, 1544, 2460, 2458, 137, 812, 2732, 2733, 1045, 458, 457, 348, 1338, 1337, 635, 879, 1393, 1300, 744, 173, 67, 310, 532, 2043, 2042, 945, 424, 1156, 2417, 1773, 1772, 495, 2671, 2669, 1193, 1810, 567, 2973, 2974, 843, 273, 1696, 1695, 1236]. Essay [1235, 2327, 1118, 671, 2249, 2248, 104, 1654, 1653, 1929, 1928, 380, 706, 242, 1850, 1849, 2567, 2923]. Essential [1245]. estimates [4773]. Estrogen [1020]. Estrogens [2725, 2724, 4644, 2927]. Eternally [1400]. Eternity [386]. Ethical [4485]. Ethics [3600, 4793, 3599, 3720]. Ethnic [4568]. Ethnobotanical [1950]. Ethnologist [4302, 4303]. Eugenics [1494, 1493, 3634, 1488]. euro [4510]. Europa [3679, 5243, 3686]. Europe [3392, 4914, 290, 295, 5165, 895]. European [101]. Europeans [2399, 1207]. Evaluation [2312]. Eve [113, 4845]. Even [232, 1625]. Event [873, 2945]. Events [3859, 1729, 1723, 1728]. Everglades [5158]. every [4770]. everybody [4588]. Everyday [3050, 4208]. Everything [2915, 2904, 2914]. Everywhere [4353]. Evidence [1903, 1375, 1902, 3380]. evil [3999]. Evolution [2101, 3908, 3451, 1178, 2627, 1414, 1624, 2103, 4739, 478, 481, 4619, 1259, 1219, 1565, 4155, 4868, 4701, 2099, 4772, 4578, 4586, 1102, 1101, 2905, 958, 960, 2927, 3611, 3073, 4385, 4277, 2918, 3560]. Evolutionarily [1632]. Evolutionary [3050]. evolve [5157]. Evolving [3613, 3908, 2934]. exam [4778]. Excellent [4848]. exceptions [4489]. Excerpt [5211]. Excessive [4187]. Excitement [14, 4035]. Executive [937]. exercise [4794]. exile [3920]. exit [4512]. exonerated [3449]. Exorcise [646]. Exoskeletal [3225, 3233]. Exotic [1973]. expanding [4453, 4773, 1529]. expanse [3997]. Expansion [1208, 1213, 4773]. Expectations [2828, 507]. Expeditions [5182, 4871, 5034, 5105, 4869]. Expendable [1245]. Expensive [3090]. Experience [2853, 4089, 1143]. Experienced [3473]. Experiment [2229, 3512]. Experimental [3551, 964, 1263]. Experimentation [3762]. experimenters [3415]. Experimenting [1115]. Experiments [4330]. Expert [5107, 3688]. experts [2150, 4777]. expire [4299]. Explain [1989, 1985, 1988, 4381, 4710]. Explained [406]. Explaining [2915, 2914, 2904]. explains [2932]. Exploiting [4064]. exploration [3710]. Explorations [4890]. Explore [4896, 4897]. Explorer [31]. Exploring [3009, 3348, 3266, 3014, 3363, 4271, 4620, 5337]. Explosions [3141, 2956, 4771]. Explosive [5352, 872, 5084, 5265]. Export [2393, 2691].

Exporting [675]. Expose [98]. Exposure [4484, 4208]. Express [4520]. extend [4476]. Extinction [331, 332, 330]. Extinctions [3221, 3226]. Extinguished [1498]. extinguisher [2929]. Extra [1801, 4809, 1801]. Extracting [4584]. extraction [3811]. Extraterrestrial [333, 331, 2105, 4942]. Extreme [2019, 4126, 5381]. Extremely [2795, 1257]. Extremophiles [3701]. Eye [221, 132, 4417, 3049, 5056, 3246, 5034, 3531]. Eyed [374]. Eyes [3313, 2431, 2483, 4500]. eyewitnesses [3709].

F [2687, 2714]. FAA [2818]. Fabré [1986]. Fabrics [3841]. Face [91, 2835, 771, 2425, 4455, 3031]. Faces [3042, 1441, 4007]. Facing [4993]. Fact [3346, 4439]. Factor [4818, 1020]. Factories [3549, 2979]. Factorland [3802, 3803]. Factors [1153]. Factory [1665, 369, 5108, 3300]. Facts [4117, 3153, 151, 3566]. Fads [1592]. Fail [1891]. Fair [3065, 3070, 1048, 3071]. fair-haired [1048]. fairness [3417]. Fairy [2969]. Faith [4958]. Fake [3089, 4251]. Fakes [4031, 4030]. Faking [3722]. Fall [703, 3738, 1427, 1415]. Fallacies [5222]. Fallacy [3349, 3365]. Falling [5308]. Fallout [5194, 2523, 81, 4901, 4175]. Falls [3248, 2862]. False [5247, 3932]. fame [4438]. Familiar [3188]. Families [824, 892, 473]. Family [1897, 1096]. Famous [3961]. fan [3540]. Fancy [165, 3527]. Fantastic [770]. Far [5135, 388]. Far-Flung [5135]. farm [3140]. Farmer [3688]. Farming [3878]. Farway [4756]. fashion [4129, 4142]. Fast [4117, 3613, 3841, 3926, 3978, 1554, 890, 2543, 2860, 4350, 4413, 4773]. fastball [4770]. fastener [3078]. Faster [4299, 1559, 3983, 2312, 4195, 4184, 4191]. Fat [5233, 431, 4807, 3282, 3276, 3281, 3535, 4309, 4724]. Fatal [818]. Fate [5281, 5291]. Father [3255, 3170, 3777]. Fault [1254]. Faults [1877]. Fauna [165]. Faux [1346]. Favorite [1053]. Fear [3488, 2217, 1454, 4095, 5230]. Fears [4857, 2749, 3832]. Feast [5029, 2764, 2307]. Feathered [3713]. Feathers [4958]. Feature [2396, 3940]. Fecundity [3488]. Federal [2821, 3047, 3863]. Feds [1592]. Feed [149, 1826, 3456]. Feedbacks [64]. Feeding [99]. feel [4464]. Feeling [5360]. Feet [2705]. Female [3725, 4787, 2746]. Females [4317]. feminists [5305]. Fenway [1993]. Fermat [406, 1590, 1824, 2261, 4017, 1680, 1691]. Fertile [1783]. Fertility [170, 3630, 1719, 3247]. Fertilizing [4299, 4308]. Fetal [3539, 283]. Feud [177, 4224]. Fever [2497, 4535, 4544, 643, 2500]. Few [5150, 4320, 3401]. Feynman [3288]. Fiber [3575, 2759, 3901]. Fiber-optic [3575, 3901]. Fiberglass [2749]. Fibers [3900]. Fibonacci [2409, 2415]. Fibrosis [2847]. Fickle [1077]. Fiction [4439, 2798]. Field [658, 669, 5057, 3953, 4666, 4667, 414, 3951, 4777]. Fields [5009, 5352, 4248]. Fiend [5276]. Fifth [867]. Fifty [4099]. Fight [1286, 1382, 2310, 1905, 3381, 2581, 3758, 4476, 4450, 4842, 4125]. Fighting [2149, 2344, 3354, 3341, 4020, 2858, 2857, 4361, 2851]. Fights [1824, 2074, 4358]. fill [2792, 3415]. Film [1177, 2730, 3419, 4442]. Films [1575]. Filtering [3646]. Filters [5354]. fin [4487]. Final [4071, 1825, 3572].

Finance [5283]. Financial [1399, 5239]. Finches [764]. Find [2263, 1288, 1385, 2582, 2585, 4058]. Finding [3362, 4489, 3361, 3127, 1254, 3644]. finds [4941]. Fine [1185, 3803]. Finessing [2261]. Finest [1069]. Fingerprinting [1052, 3420, 3261]. Fingerprints [676]. Fingers [920, 1354, 2403, 2406]. Finite [5030, 449]. Fire [2929, 5233, 2437, 4898, 956, 2265, 4696, 4697, 3715, 4900]. Firearms [792]. fireball [4772]. Fireflies [1730, 1725]. Firefly [4978]. fires [4244]. Firewalls [4653, 4654]. Firewater [699]. Fireworks [5152]. First [5025, 1757, 3574, 5069, 5111, 5026, 951, 4689, 4562, 4778, 4756]. First-Class [951]. first-detected [4756]. Fischer [93]. Fish [1074, 4049, 4610, 3381, 3992, 4391, 2778, 4611, 2771, 699]. fisheries [4611]. Fishermen [3311]. Fishing [3639, 4613, 4611]. Fishy [2353, 2093]. Fission [2897, 4150]. Fit [2704]. Fittest [2234, 1080]. fitting [3230]. Fix [506, 937, 413, 1648, 4413]. Fixing [3801, 3616]. Flag [3614]. Flames [572]. Flaming [2185]. flamingos [4172, 4178]. Flammable [5289, 5295]. Flaps [1426]. Flash [112]. Flashing [4978]. Flat [626, 1287, 1384, 1401]. Flat-Panel [1287, 1384]. flatland [3005, 3011]. Flatter [4195, 4184, 4191]. Flaw [818]. Flea [4172, 4178]. Flea-size [4172, 4178]. Fleas [3108]. flecks [4527]. fleet [3186]. Flesh [5160]. Flew [5298, 5306]. Flexible [3690]. Flickering [2116]. Flies [1386, 5109, 3773, 1298, 3769]. Flight [3528, 5044, 1386, 4958, 4199, 3527, 593, 1298, 2818, 3267, 2979, 5045]. Flight-Testing [1386, 1298]. Flights [165, 2842]. Flip [4560]. Floating [3832, 4706, 4707, 5301, 5302, 3780]. flood [4944, 4943]. Flooded [1290, 1388]. Flooding [3551]. floor [4617]. Flora [165, 3125]. Floral [5276]. Florence [438]. Flotsam [1205]. Flower [2138]. flowing [3122]. Flu [4149, 4866]. Fluids [1633, 1649, 1638, 1077]. flummoxed [3769]. Flung [5135]. Fluoridation [2938, 2933]. Flutter [3564]. Fly [2872, 5328, 2540, 5087, 5124, 2524]. flyby [5195]. Flying [2581, 2604, 2425, 2890, 3913, 3779, 183]. Flynn [4852, 4853]. Flywheels [3971]. Focus [3430, 3138, 3031, 3380, 3082, 4040, 3997, 3528, 3572, 3621, 4169, 4450, 4398, 4501, 4348, 4224, 4285, 4553, 5272, 5195, 5319, 5157, 4954, 5085, 5051, 5229, 4892, 4994, 4290, 4795, 4230, 4676, 5049, 4842, 4785, 4175, 4125, 4722, 4624, 3720, 3909, 3767, 3863, 3677, 3820, 4232, 4292, 4347, 4451, 3191]. Focused [766]. Folate [4800]. Fold [3955]. Folded [115]. Folding [437]. Folies [3239]. Folkman [4634]. Follow [2485, 5191]. Food [3634, 415, 3033, 804, 394, 2225, 4516, 2074, 61, 2482, 4531, 4000, 4855]. Food-poisoning [3634]. Foods [1147]. Fool [1491, 1503]. footing [2685]. Footprints [4567, 4575]. Footwear [1205]. forage [4531]. Forbidden [985]. Force [1285, 1381, 4724]. Forecast [3146]. Forecasting [3140, 3531]. Forest [3437, 4301, 2989, 2695]. Forestalling [4556]. Forests [1290, 1388, 3696]. foretold [2516]. Forever [1663, 5269, 665]. Forgeries [2409, 2415, 5222]. Forgotten [3709, 5032, 4016]. form [4756]. format [4230]. Formed [1316]. former [2583]. Formerly [4202]. Formula [1167]. Fortran [665]. Forty [4606]. Forum [3600, 1204, 3599]. Fossil [297]. Fossils [3788, 2185]. Foucault [4474]. Found [483, 2016, 3288]. Foundation [4732]. Fountain [182]. Four [3125, 3752, 1662]. Fowl [317]. Fox [3154]. Fractal [165, 3122, 1538]. Fractals [2896, 54, 53, 258, 5130]. Fractured [1711]. France [2933]. Francis [4180, 4190, 4188]. Frankenstein [3179, 5305]. Franklin [3852]. Frankly [3868]. Fred [2311]. Free [1669, 1670, 2892, 1672, 2842, 153, 5331, 4967, 315, 2818]. Free-for-All [2842]. Freedom [5059]. Freeman [1557]. Freewheeling [2893]. freeze [3374]. freeze-dry [3374]. Freezer [1280]. Freezing [577, 259]. frenzies [4855]. frequency [3121]. Frequent [1426]. frequently [4789]. Fresh [4909]. Freshwater [4049]. fret [2934]. Freud [2977, 2595, 3510]. Friction [3397, 3404]. Friend [1507, 1713, 682, 2349]. Friends [3713]. Fright [2472]. Fringe [2710]. Frisbee [5045, 5044]. Frog [2556]. Frogfishes [202, 201]. Froggy [4563]. Frogs [1067, 1833, 2571, 3852]. Front [5278]. Frontier [734, 725, 4071, 3546, 4770, 1840, 3572]. Frothing [1005]. frown [4385]. Frozen [4204, 3031, 2936, 405]. Fruit [3769, 1386, 1298]. Frustrations [1132]. Fry [3775]. Fuel [5141, 5140, 4299]. Fuelish [1040]. Fuels [830, 4270, 297]. Full [4045, 3462, 3632]. Fullerenes [760, 1346, 754]. fumed [4462]. Fun [5004, 3462, 1686, 3660]. Function [2783, 2785]. Functioning [100]. Functions [1711, 3741]. Fundamental [3321, 3322, 1900, 3741]. fundamentally [4773]. funds [2582]. Furry [4045]. Further [907]. Fusion [3419, 4464, 279, 2649, 4468, 990, 4515, 4526]. Future [2637, 3350, 5141, 3215, 2638, 2628, 159, 2121, 5176, 5174, 4356, 5140, 4805, 4116, 2355, 3983, 1487, 4106, 5029, 5033, 1787, 5372, 2657, 3351, 3815, 435, 3968, 2309, 447, 2858, 2851, 2857, 5360, 1706, 3019, 3083, 4350]. Futures [58]. Fuzzy [1526, 1528].

G [172, 2959, 2963, 1063, 1062]. G. [3201]. Gag [1399, 3430]. Gags [574]. Gaining [3282, 3281, 3276]. Galactic [3580, 363, 2956, 4771, 4775]. Galaxies [685, 5382, 3595, 3602, 4643, 3744, 1485, 4775, 4771]. Galaxy [4739, 120]. Galileo [3140, 2846, 5076, 5077]. Gallery [1911, 1924]. Galling [1050]. Gallium [60]. Gallons [3542]. Galore [5000, 3854]. galvanic [3852]. galvanized [3179]. Game [2721, 505, 3371, 2169, 2727, 3749, 853, 2728, 3017, 5005]. Games [5004, 4828, 755, 668]. Gamester [2822, 2845]. Gamma [3842, 1720]. Gamma-Ray [3842]. gap [4855]. Garbage [1463, 1612, 416]. Garden [772]. Gardens [5139, 5138, 4206]. Gardner [2822, 2845]. Gargantuan [5367]. Gas [3185, 562, 4270, 4725, 2525, 4759, 4436]. Gasket [5185]. gaslight [3806]. gasoline [2873, 3179]. Gate [4966]. Gates [2934, 2950, 4689]. Gateway [4082, 4069]. Gatherer [5009]. Gathering [1936]. Gay [2750, 2766]. Gears [737]. Geese [248]. Gels [1450]. Gem [1576]. Gender [4663, 4811]. Gene [2633, 3261, 3793, 254, 3795, 1828, 448, 3794, 3792, 3796, 2175, 548, 4146, 3797, 5170, 1953, 1822, 1948, 1952, 1766, 4691, 364, 1310, 2588]. General [4818]. Generate [877, 868]. Generation [3215, 3216, 5071, 867, 5115, 446, 3401, 5387, 3256]. generations [4758]. Genes [218, 694, 691, 693, 2351, 801, 225, 1036, 1174, 3653, 1347, 2766, 2600, 4513, 2471, 3928, 2398, 185, 575, 1305, 4340, 1307, 326, 2274, 3991, 2750, 3089, 4811, 4804]. Genesis [585, 959]. Genetic [2315, 2189, 2359, 1064, 3337, 5136, 4845]. Genetically [1147, 1961, 1951, 3394]. Geneticist [3449]. Genetics [3013, 587, 963, 968, 5015, 1488, 1493, 4796, 2138, 4369, 1291, 1383, 1948, 1952, 2819, 4624]. Genie [4357]. Genius [1749, 3761, 5115, 3709]. Genome [4180, 4190, 4000, 4520, 4188, 969, 963, 968]. Genomes [5378]. Genomic [4506]. Gentler [539]. gently [3857]. Geometry [4373, 4374, 4212, 523, 1331, 4489]. Geophysics [619, 624]. George [2687, 3123, 2714, 3130]. Gerard [1194, 1811]. geriatrics [4845]. Germain [832]. German [3516]. Germanium [1829]. Germany [1375]. Germplasm [756]. Gestalt [1458, 404]. Get [4794, 2744, 1578, 1569, 142, 2498, 2499, 4114, 2495, 3433, 3875, 4224, 4498, 1626, 4627, 2343]. Gets [3834, 4886, 3730, 3664, 4350, 4938, 151, 2483]. Getting [2983, 4466, 5092, 3800, 3890, 1037, 4832, 5361, 941, 2263, 2991, 1742, 3877]. Ghastleigh [1158, 1189]. Ghost [5106, 5105]. Ghostbusters [3410]. Ghostliest [3595, 3602]. Ghosts [4604]. Giant [2850, 4756, 172, 3469, 4947, 3433, 3737, 3750, 3468]. Giants [3780, 4759]. GIF [2308]. Gifted [4820]. Gifts [3887]. Gigabit [340, 1458]. GigaPOP [3489]. Gilded [1127]. girds [4994]. Girl [1713]. Girls [4785]. Girth [3998]. Give [2237, 1625, 1297, 2244]. gives [4809]. Giving [2702, 4154, 4167]. Glacial [22]. Glance [955]. Glandular [3887]. Glantz [4462]. Glass [4731, 150, 4274, 1041, 554, 553]. Glazes [134]. Glen [1460]. Glenn [4848]. Glimpse [1092]. Glimpses [3188]. Global [923, 4627, 4969, 3437, 3630, 698, 726, 2028, 2025, 2027, 2176, 261, 3455, 3652, 1782, 4477, 2128, 3968, 2258, 1425, 3844, 97, 86, 96, 2816, 2402, 733, 2953]. Global-warming [4627]. Globally [2714]. glory [2581]. Glow [4234, 4235]. Glucose [862]. Glue [5114, 3741]. Glueballs [4700]. Gluons [2271]. Glut [2939]. gluttony [3282]. Go [3417, 4923, 2631, 2620, 3627, 3984, 3486, 3948, 4551, 144, 2157, 2146, 167, 156, 166, 3419, 4028, 4053]. go-cart [4053]. Goals [4827]. Goats [2991, 2983]. Gobbler [2009]. God [3466, 2783, 2785]. Goddess [1094]. Gödel [5104]. Gods [3661]. Goes [3726, 3470, 4057, 266, 1311, 3125, 3752, 4476, 2309]. Going [3198, 2547, 389, 3645, 919, 5156, 2942, 3572]. Gold [3199]. Goldberg [4339]. Golden [2616, 172, 1049, 736, 2625]. Goldilocks [1664]. Golf [3607, 3683, 3608]. Golfers [2820]. Golygon [232]. Gombe [3961]. Gone [4456, 112, 2049, 3383]. Good [2692, 2563, 2596, 4964, 5094, 3158, 4402, 1956, 5168, 3247, 4476, 2759]. Good-bye [2692, 1956]. Goods [741]. Goose [3753]. GOP [2840]. Gorilla [4458]. Gossip [3433]. got [4501]. Gothic [220]. Gotta [3955]. Gourmet [3712]. governments [3247]. Grace [1899]. Gradients [3272, 3268]. Grading [1953, 1952, 1948]. Gradiometry [4426]. Graft [709]. Graham [3632]. Grain [1185, 144]. Grameen [5299, 5307]. grammar [3586]. Grand [4885, 3231, 3551]. Grande [1608]. Grandeur [3239]. Grandmaster [325, 314, 324]. Grange [1158, 1189]. Grapevine [524]. graph [3938]. Graphics [442]. Grass [45, 1804]. Gratification [758]. Grave [11]. Gravel [1463]. Gravity [4426, 1122, 711, 1295, 4627]. Gray [1922]. Great [2828, 352, 2212, 3291, 4021, 468, 2645, 1566, 2662, 1492, 224]. Greatest [5380]. Greatness [4408]. Greece [2320]. Greek [3193, 3116, 3120]. Green [131, 114, 1997, 2070, 4206, 4219, 874, 1252, 1923, 1819, 2883, 2439, 4964, 266, 185, 3659, 1684, 2946, 3660]. Greenback [2489]. Greenberger [4777]. Greenhouse [4401, 748, 4402, 4125, 395]. Greenies [2033]. Greening [1612, 3686]. Greenland [4204, 4698, 4699]. Greens [2844]. greeted [1192]. grew [4772]. Grim [1659, 507]. Grin [296]. Grind [3915]. grooming [3433]. Ground [1073, 1783, 5255, 5254, 4300, 5119, 4287, 4294]. grounds [4611]. Group [3204]. Groups [4568]. Grow [5241, 2964]. Growing [2959, 1826, 2963, 871, 5014, 3951]. grows [3478]. Growth [1153]. Gruff [2135]. guarantees [4914]. Guards [1911, 1924, 3832]. Guessed [5310]. Guessing [505, 613]. Guide [3458, 2372, 4666, 4667, 3463, 4759, 4776]. Guided [5102]. guides [2968]. Guilty [3366, 15]. Guinea [4194, 4617]. Guitar [3856, 3857]. Guitar-Effects [3856]. Gulf [3997, 2345]. Gulliver [2602]. Gun [5184]. Gundestrup [932]. Guppy [3481]. Guru [2135, 2606]. Gusher [748]. Gushers [3580]. Gutenberg [2713]. Gutless [3238]. Gutting [2874]. Guy [1407, 975]. gym [3865]. gymnastics [3516].

H [5005, 2226]. H. [4777]. Habeas [1330]. Habit [1171]. Habitat [3403]. Habits [5347]. Hack [3440]. Hackers [4652, 4621]. Hacking [969, 968, 963]. hacks [3688]. hacktivists [4842]. Hag [4645]. Hair [4018, 4430, 4431]. Hair-Trigger [4018]. haired [1048]. Haiti [2178]. HAL [3560]. Hale [3556, 3606]. Half [5112, 5310, 4748]. Hall [3005]. Halls [2427]. Halo [2450]. halogen [3235]. hamper [5085]. Hand [1115, 4637]. handed [3722]. Handedness [27]. Handful [591]. hands [3632]. handwriting [4436]. Hanford [3124, 3118]. Happening [120, 2993]. Happiness [3113, 3117, 4090]. Happy [3560, 4433]. Harbor [4481]. Hard [1621, 3487, 557, 556, 3566, 4553, 549]. Harder [5375]. Harebrained [2944]. Harm [1669]. Harpsichord [480]. Harrison [3267]. Harvard [2737]. Harvest [608]. Hatchet [280]. haunting [3178]. Have-Nots [2379]. havens [4757]. Having [4813, 2996]. Haystack [3444]. haystacks [4390]. haze [3748]. Hazy [3747]. Head [3800, 2921, 5358, 4136, 4732]. headache [3941]. Headaches [4564, 4780]. Heading [1881]. Heads [1802]. Headsets [1296]. Headshrinker [3200]. Headway [3319, 3320]. Healing [4572, 1519]. Health [3759, 4777, 4814, 5002, 3497, 3503, 1204, 1212, 2025, 2027, 4782, 3745, 3741, 3740, 3760, 2360, 2366, 4856, 5094, 3505, 5334, 1539, 4236, 4855]. Healthy [4032, 2092, 4813]. Hear [235, 2226]. heard [4384]. Hearing [1519]. Heart [2763, 4806, 1703, 1026, 1580, 2438, 2752, 4800, 4771, 3443]. Heartache [3835]. heartbeats [3284]. Heat [884, 2073, 4626, 773, 4187, 71]. Heaven [695, 3609, 704, 1653]. Heavenly [1335]. Heaviside [206, 193, 205]. Heavy [4496, 3963, 3173]. Hebb [1323]. He’d [1551]. Heel [3298]. Heike [3654, 3655]. Heisenberg [2412, 1001]. Heli [79]. Heli-Hopper [79]. Helicopter [3585]. Heliosphere [2357]. Helium [2983]. Hello [4738]. Help [1109, 4781, 2446, 3677, 4793, 4125]. Helping [3835, 975]. helps [3394]. Henri [1986]. Hepatitis [5246, 552, 4954]. Her [1683, 1679, 1682]. Herb [4557]. Herbal [4184, 4191]. Herbicide [353]. Here [4854, 3990, 4372, 5052, 1294, 142]. heroes [4099]. Hidden [3224, 3219, 5148, 2680, 5243, 1629]. Hide [2598, 4775]. Hide-and-Seek [2598]. High [1353, 5383, 4919, 170, 4977, 2634, 2644, 5248, 2061, 2616, 2625, 1673, 1678, 815, 2095, 3275, 1829, 5255, 5254, 169, 3977, 2698, 2813, 1419, 2537, 1295, 2837, 2567, 2873, 3820, 4030, 3271, 3180]. High-Energy [2061]. high-octane [2873]. High-Performance [169]. High-Power [1678, 1673]. High-Precision [4977]. High-Speed [5248, 2616, 1829, 3977, 2625]. high-tech [3820]. High-Temperature [1353, 5383, 2634, 2644, 3271]. Highbrow [3133]. Higher [4803]. Highways [972, 4926, 4928, 3972, 3927]. Hill [4684, 2770]. Hint [5310]. Hints [2358, 5385]. Histones [1174]. Historian [2363, 2362]. History [2934, 3610, 101, 2278, 2096, 3108, 544, 2950, 2364, 650, 3060, 1542, 3905, 984, 1318, 3556, 3709, 4875]. hitched [4224]. HIV [4476, 4483, 4478, 3298, 4485, 4290, 4484, 4482, 4477, 2552, 4481]. Hoax [4694, 4693]. HOH [3755]. Hol [576]. Hold [3924, 3364, 3380, 4562, 4775]. Hole [1149, 648, 488, 4771]. Holes [1285, 1381, 3170, 2470, 5065, 2216, 363, 3697]. Holey [936, 5132]. Holograms [4154, 4167]. Holographic [4562, 2781]. Holography [4573]. Home [232, 2546, 4661, 5248, 2830, 1743, 877, 868, 3537]. Homecoming [858]. Homeless [824]. Homemade [131, 3174, 3415, 3181, 4977]. Homeobox [225, 608]. Homes [291]. Hominid [3788]. Homosexuality [1901, 1902]. Honest [1440, 3426]. Honeybee [1949]. Hong [3820]. Honor [5238, 3082]. honors [3526]. Hookworm [2444]. Hoop [4513]. Hop [3952]. Hope [3295, 390]. Hopper [79]. Hopping [3142]. Hopscotch [1110]. Horizon [3776, 2945]. Horizons [626]. Hormone [3922, 3332, 3333, 4802, 4724]. Hormones [3924]. Horror [3098, 3109]. Horse [921, 5304]. Horseback [828]. Horses [587]. hospitable [4757]. Hospitals [1881]. host [4741]. Hosts [443, 1288, 1385]. Hot [3679, 2052, 1879, 1854, 2134, 4463, 3516, 2789, 2784, 5036, 5072, 916, 1145, 2838, 248, 1125, 2258, 3432, 3678, 1253, 2997, 3073]. Hot-Spotting [3678]. Hour [2894]. house [2933, 1260, 1267]. How-To [3846]. Howls [747]. Hoyle [2311]. Hubble [3198, 178, 1025, 1529, 468]. Hubbub [41]. Hubris [5381]. hues [4610]. Huffman [719]. Hullabaloo [3388]. Human [2685, 3034, 3560, 3013, 1178, 1826, 2121, 4728, 54, 53, 3739, 4259, 962, 5340, 5357, 1721, 4495, 686, 1408, 5341, 1262, 4973, 4972, 808, 4571, 4579, 3368, 4476, 4506, 4296, 4676, 4842, 4277, 4494, 4099, 4180, 4190, 4188]. human-machine [4296]. humanities [4244, 4438]. humanity [3230, 3421]. Humankind [1910]. Humans [957, 1416, 4962, 4896, 4897, 407, 958, 960, 959, 4916, 3823, 5157, 4277]. humidity [4431]. Hundred [4996]. Hunger [917, 4722]. hungry [4504]. hunt [3383]. Hunter [1132]. Hunters [4698, 4699, 1406, 4696, 4697]. Hunting [3750, 1987, 3749]. Hunts [3390]. Hureyra [2026]. Hurricane [3388]. Hurricanes [3383]. hurts [4127, 2329]. Hybrid [3971, 3970]. Hyde [2762]. hydrates [2525]. Hydraulic [4221, 4209, 4222]. Hydraulics [2912]. hydrogen [2820, 4326]. Hygrometer [4430]. Hyperactivity [4576, 4581]. Hypersearching [5101]. Hypertension [4918, 1310]. Hypertext [2911].

I.T. [2546]. Icarus [3981]. Ice [4204, 1987, 885, 5289, 5295, 1387, 2686, 2530, 3481, 4606, 1289]. icy [5149]. Idea [4364, 2951, 1370, 2116]. Ideal [3360]. Ideas [5188, 5032, 683]. Identifying [2927]. Idols [4223]. If [4538, 4429, 92, 1344]. Ig [3481, 2172, 817, 2174, 2834]. Igneous [1631]. IgNobility [2815]. Ignored [1859, 3318, 3329]. II [1048, 4698, 3124]. Ill [4993, 1572]. Illegitimi [2036]. Illiteracy [3725]. Illness [2275, 4091, 2554, 2874]. ills [4724]. Illuminating [3235, 1404, 4741]. Illuminator [1170]. Illusion [775, 2473]. Illusions [3943, 3942]. illustrated [3805]. Image [753, 1816, 5102, 2879, 2354, 1135]. Image-Guided [5102]. Images [607, 872]. Imaging [687, 688, 3325, 3338]. Imanishi [3449]. Imanishi-Kari [3449]. Immeasurable [215]. Immigration [5203]. Immortal [1971]. Immune [3451, 1286, 1382, 1599, 1596, 1601, 1597, 1594, 2552, 1598, 2779, 1595, 1603, 762]. Immunity [5198, 4912, 3451, 3452, 3453, 1439]. Immuno [1958]. Immuno-Logistics [1958]. immunodeficiency [4476]. Immunology [1615, 740, 402, 408]. Immunotherapy [3593, 3352, 3351, 157]. Impact [333, 331, 122, 5340, 5357, 4349, 4858]. impede [1014]. imperial [3895, 1272]. Imperiled [2778, 4611, 2771]. implant [3732]. implants [3047]. Implosion [2718]. Importance [4777, 2067, 4782, 2429]. Important [4769, 1194]. imports [4453]. impossible [4616]. Impressions [3372]. Imprinting [326]. Improve [394, 4348]. Improved [2628, 2638]. Improving [3342, 4478, 2184]. Imroving [3343]. Incentives [1539]. Income [5093, 3401]. Inconstant [4678, 1457, 1456, 4679, 1449]. Incontinence [4812]. Incredible [1660]. Indecent [49]. Indeed [3988]. Index [422, 842, 1271, 1734, 2200, 2868, 3519, 4079, 4753, 5353, 2839, 423]. Indexing [2814]. India [4501]. Indian [4601]. Indictment [1748]. Indigo [2482]. Individuality [1137]. Indivisible [5022]. Indo [101, 1207]. Indo-European [101]. Indo-Europeans [1207]. Indochinese [892]. indomitable [3777]. Industrial [1029, 1034, 740, 2650, 3430, 4539, 3282]. Industries [1464, 4224]. Industry [4109, 2899, 2901, 292, 1849, 4462, 5051]. inefficiencies [3047]. Inequality [5093]. Infamy [5238]. infant [4028]. Infants [98]. Infection [4484, 4482, 2444]. Infections [4476, 257]. Infectious [1598]. Inferno [849, 4771]. Infertility [4791, 4792]. infinite [2512]. Infinitesimal [3834, 4774]. Infinity [829, 2364, 840]. Inflation [4453, 4864, 3766, 4452, 2814, 3186, 3767]. Inflationary [2137, 4774]. inflexible [4489]. Influence [1902, 3753]. Influenced [1901]. Influential [1118]. Information [2610, 3362, 3361, 2379, 3646, 3697, 2659, 4358]. Infrared [943, 934, 931]. Infrastructure [1332, 733, 2656, 343, 726]. Inherit [2107]. Inheritance [1300]. Injected [1794]. Ink [2549]. Inn [5325]. Inner [4638, 2400, 879, 1611, 5272]. Inner-City [4638, 879]. Innocent [676, 3366]. Innovation [1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 2207, 2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623, 4675, 4719, 4838, 4889, 4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317, 2607, 2609]. Innovative [4912]. Inquiring [2487]. Ins [303]. inscrutable [4564]. Insect [370, 3284]. Insectoids [658, 669]. Insects [487, 823, 1663, 3256, 4147, 4865, 3801]. Insignificance [2347]. inspiration [2585]. inspire [3138]. Instant [5114]. instants [4772]. Institute [4506]. Instruments [4629, 4758]. Insult [1965]. Insurance [4587]. Insurers [2934, 4637]. Integrated [4107]. Integrins [3741, 3740, 3760]. Intellectual [2952, 1211]. Intelligence [17, 2106, 4822, 4827, 4818, 1668, 2617, 4829, 4816, 826, 833, 4815, 4878, 4942]. Intelligences [4817]. Intelligent [4830, 2618, 1450, 2643, 4816]. intense [3180]. Intensive [1881]. interact [4811]. Interacting [863]. Interaction [4087]. Intercepted [1228]. Interests [938]. Interfaces [3647]. Interferometry [3594, 653]. Interferons [1905]. Interleukin [85]. Interleukin-2 [85]. International [4904, 4893, 964]. Internet [2666, 3309, 2961, 2934, 3785, 3737, 4015, 3641, 4650, 4366, 4312, 4269, 2856, 3308, 3099, 3828, 3642, 4651, 2161, 3784, 2153, 4365, 4842, 4283, 3651, 3643, 5249, 5250, 3649, 3617, 3646, 2965, 4659, 2150]. Interplanetary [1106, 1116]. Interrogator [3349, 3365]. Interstellar [4927]. Interview [3039, 4777]. Intolerance [712]. Intractability [1753]. Intrigues [2410]. Introducing [4099]. Introduction [4598, 4987, 2607, 3319, 4777]. Intuitive [1229]. Invade [658, 669]. Invaders [116, 1596, 3870]. invades [4287, 4294]. Invading [4900]. Invasion [894, 2224, 3025]. Inventing [3779]. Invention [4156]. Inventions [4949]. Inventor [183, 5005]. Invertebrates [3451, 3452]. Investigating [3883, 3904, 2973]. Investigation [2410]. investigators [4476]. Investment [1118]. Investments [3819]. Invisible [5274, 442, 1115, 4878, 3127, 4661]. Inward [625, 619, 624]. Ion [766]. Ions [5218]. IQ [2754, 4852]. IQs [2977]. Iran [3786]. Iraq [2003, 462]. Iron [759, 4266]. Irregular [3043]. Irrigating [4531, 4537]. Irving [1520]. Isaac [4492]. Island [4618, 3466, 895]. islanders [1192, 5070]. islands [3779, 3807]. Isn’t [3510, 3683, 4606]. Isotopes [1184]. Issue [2097, 4248, 5355]. Italian [2973]. Italian-Style [2973]. Itself [4911, 1826, 3308, 4789, 3456, 409, 402, 4859, 4230, 1542, 408]. Ivy [1690]. Izaak [4492].

J [4082, 4083, 1557]. J. [4506, 2496]. jackal [3584]. Jackpot [3294]. Jacques [3914]. Jacques-Yves [3914]. Jail [3417]. James [2870, 3154, 4852, 2487]. Jamison [2753]. Jamming [4735]. Jan [3920]. Jane [550]. Jankowski [3920]. January [3464]. Japan [2525, 4745, 663, 698, 990, 5009]. Japanese [4373, 4374]. jar [3462]. Java [2873, 4657, 4658]. Jean [1986]. Jeff [716]. Jekyll [2762]. Jeremy [3873]. Jerry [1203]. jet [4287, 4294]. jetmobile [4053]. Jigsaw [3987]. Job [1854, 2477, 857, 3426, 280, 4799, 4606]. Jobs [805, 1863, 2093]. jocks [4724]. Joe [3420, 1740]. Joel [4410]. John [4848, 4082, 5005, 4941]. Jolt [1250, 1589]. Jonas [3400]. Jones [4777]. Joseph [4086, 2897]. Josephson [2400]. Journey [232]. Jove [2088]. Jovian [1589]. Joy [1039]. Joycean [1716]. joyful [3940]. Joys [3704]. Judah [4634]. judgment [5263]. Juggling [2788, 3631, 3632]. Jules [582, 3703]. Junction [2400]. Jungle [550, 3865, 3479, 4637]. Juniper [3659, 3660]. Junk [1045, 2874, 4527]. Jupiter [3140, 2557]. Jupiters [3432]. Jurassic [1628]. Just [5004, 91, 3389, 4391, 1940, 2476, 3193, 4606, 4774, 2264, 984]. Justifying [3805]. Juvenile [2368, 2361].

Kaczynski [3149]. Kalahari [123, 133]. Kamerlingh [3654, 3655]. kangaroo [1082]. Kangaroos [5068]. Kansas [5263]. Kari [3449]. Kay [2753]. Keck [2536]. Keeping [3145, 1537, 4616, 1582, 1096, 2930, 3211]. Kefir [2042]. Key [4890, 3145, 2984]. Keyhole [2201, 1749]. Kick [2343]. Kicking [1210]. Kidding [3673]. Kidnaps [356]. kids [2823, 1203]. Kilauea [1097]. Kill [392, 4366]. killed [3940]. Killer [3034, 679, 5068, 3032]. kills [4287, 4294, 4634]. Kin [2451]. Kinder [1518]. Kinetic [1357, 1358]. King [3439, 3251, 1841]. Kingdom [5204]. Kings [3150, 3140]. Kissing [46, 900, 909]. Kitchen [4157, 1882, 1866, 1871, 4584]. Kitty [1402]. Klein [4274]. knee [3616]. Knees [4287, 4294]. Knight [3706, 3707, 3169, 3172]. Knot [4074, 368]. Knots [1768, 2786, 1760, 2791, 2792]. Knotty [3228]. Know [4642, 4783, 3955, 2966, 4772]. Knowledge [3521, 3235, 3294, 3128, 3472, 3670, 3566, 3616, 4034, 4081, 4498, 4154, 4543, 4590, 4390, 4442, 4209, 4269, 5045, 4948, 3373, 3471, 4166, 4167, 5189, 5044, 5114, 5268, 5354, 4670, 4550, 3565, 3714, 5314, 4389, 5080, 4497, 4338, 4339, 3293, 4283, 3424, 3077, 3810, 3900, 3856, 4441, 3240, 4884, 4885, 4495, 4033, 3945, 3615, 4597, 4754, 4755, 3992, 4080, 3187, 3520, 3757, 3669, 4947, 3134, 5223, 4546, 4221, 4222, 5152, 4986, 3946, 3758, 3857, 3901, 3811, 3715, 4494, 3186, 3374]. Known [3596, 4202, 2458, 4336]. Knows [2935, 4538]. Komodo [4974]. Kong [3820]. Kosowatz [4082]. Kuiper [3103, 3111]. Kung [123, 133]. Kuru [249]. Kuwait [751, 641].

L [3830, 3632]. Lab [1683, 1679, 1682, 1882, 1871, 4666, 4667, 4679, 2265, 4464]. Label [1923, 4524]. Labels [3757, 3758]. Labor [5058, 4795]. Laboratory [2723, 2722, 2693, 4177, 4169, 4175]. Labs [3632]. Labs-Research [3632]. Labyrinth [483]. Lace [3236]. Ladders [3992]. Lady [5347, 4326]. Laetoli [4567, 4575]. lag [4287, 4294]. Lagoon [2396]. Lags [3978]. Laid [1786]. Laid-Back [1786]. Laima [4682]. Lakes [4920]. Lamarck [3909]. Lambic [3277]. Lamps [1289, 1387]. Land [92, 3098, 3109, 2513, 3769]. Landau [3888]. Landauer [4559]. Landing [2348]. Landscape [335]. Lane [890]. Langevin [4336]. Language [1138, 1272, 1013, 3629, 258, 1949, 2824, 4421, 1084]. Languages [657, 801, 4242, 4243, 101, 4227]. Lanka [4613]. Large [1631, 3598]. LaRouche [990]. Laser [131, 4911, 621, 4318, 5268, 2313, 896, 4138, 4472, 3218, 4735]. Lasers [3832, 1210, 854, 3883, 3884, 3922]. Last [1942, 4372, 2782, 4017, 1680, 1691, 1840, 1068, 3449, 3492]. Late [2430, 1781, 3578, 1987]. Latin [315]. Lattice [540]. laughing [4436]. Law [734, 725, 4108, 283, 4619, 2270, 4325, 2329, 4333, 2862, 3702]. Lawrence [3119, 3114]. laws [4510]. Lay [2392]. Layer [4006, 4336]. layers [4336]. layoff [2150]. Lead [966, 977]. Leader [1786]. Leading [503, 4476, 4180, 4190, 4188]. Leads [232]. Leaf [3212]. Leafy [4743]. League [1112, 5009, 1690]. Leak [4725]. leaks [5085, 4175]. Leaky [433]. lean [2150]. Leaning [2849, 2848]. Leapfrog [663]. Leaping [730, 742, 4727]. Learn [2358, 5385, 1143]. learned [4757]. Learning [3308, 1368, 1137, 737, 4621, 1256]. Learns [4418, 4419, 762]. Leave [1551]. leaves [3198]. Leaving [5060]. Lee [4051]. Left [3722, 5100, 4783, 4548]. Left-handed [3722]. Leg [4730]. Legacy [5245, 178, 1976, 404, 3052, 1981, 3168, 3124, 2523, 3074]. legalizing [5319]. Legend [1712]. Lego [1329]. Legs [2563, 2843]. Leisure [741]. Lemurs [1319, 1321]. Lend [4984, 4487]. Lending [3676]. lends [4230]. Léon [4474]. Leonard [4086]. Leonardo [4156, 5304]. LEP [227]. Leptin [4727]. Less [442, 2959, 983, 2963, 4359, 344, 3490, 3125, 3752, 2974]. Lessening [4795]. Lesser [3596]. lesson [2934]. Lessons [544, 3290]. Let [4193]. Lethal [1976, 3254, 1279, 4531]. Letter [1016, 2734, 2803, 2871, 2924, 2975]. Letters [383, 1, 37, 68, 105, 138, 174, 210, 244, 275, 311, 350, 382, 603, 637, 673, 707, 745, 780, 813, 426, 460, 496, 533, 568, 569, 1196, 844, 881, 913, 947, 981, 1046, 1085, 1119, 1157, 1195, 1238, 1239, 1275, 1302, 1339, 1377, 1394, 1434, 1471, 1510, 1546, 1584, 1617, 1618, 1655, 1656, 1698, 1699, 1737, 1775, 1813, 1851, 1888, 1931, 1968, 2006, 2045, 2085, 2123, 2164, 2513, 2206, 2254, 2289, 2290, 2332, 2333, 2377, 2418, 2464, 2511, 2568, 2570, 2672, 2735, 2804, 2872, 2977, 2927, 2876, 2926, 2976]. Letters [3028, 3137, 3189, 3242, 3296, 3376, 3427, 3475, 3762, 3523, 3569, 3618, 3674, 3717, 3817, 3860, 3906, 3949, 3994, 4037, 4119, 4171, 4226, 4286, 4343, 4394, 4446, 4502, 4552, 4622, 4672, 4718, 4834, 4835, 4887, 4950, 4990, 5047, 5082, 5116, 5154, 5192, 5225, 5270, 5316, 3080, 2594]. Lev [3888]. Levant [1416]. Level [2469, 3741]. Levi [4135, 1314]. Levi-Montalcini [1314]. Levi-Strauss [4135]. Libchaber [2982]. Liberties [735, 727]. Liberty [4326]. Lie [108, 4409]. Lies [1767, 4081]. Life [3057, 4758, 3380, 2977, 3140, 509, 848, 859, 860, 2390, 4397, 5368, 5096, 5135, 3986, 4592, 5338, 2121, 3202, 4428, 484, 5209, 4704, 4061, 3888, 2103, 3050, 987, 4143, 4757, 1337, 2109, 5281, 5291, 4830, 3509, 4917, 3755, 1594, 4615, 2102, 4876, 890, 3058, 2105, 4628, 1446, 3933, 651, 5344, 1102, 1101, 2098, 2097, 3799, 3798, 5386, 3230, 2874, 3193, 4627, 4617, 4813, 3232, 3754, 2918, 5005]. Lifeblood [4635]. Lift [1646, 5008]. Light [131, 5257, 2073, 3876, 4945, 1559, 521, 1246, 2191, 1423, 1605, 2895, 130, 4929, 4193, 2283, 2279, 1805, 4926, 4928, 1129, 2994, 2273, 2491, 376, 318, 357, 593, 770, 3207, 4722, 4721, 5009, 985, 2171, 4362, 4040, 3901, 4184, 4590, 4206, 4191]. Light-Water [130]. Lighten [627]. Lightness [1896, 3210]. Lightning [3883, 3884, 1367, 3885]. Lights [661, 4517, 3235]. Lightwave [867]. Lightweights [1006]. Like [2134, 2601, 2711, 2838]. Likely [90]. Lilac [4029]. Limbo [4400, 4172, 4401, 4178]. Limbs [967, 4096, 4922]. Limit [5376, 1825, 1944, 4773]. Limits [3411, 5104, 2357, 4559]. Lindy [3516]. Line [1202, 2399, 265, 5134, 4057, 1500, 706, 3178, 4236, 4914, 3622, 5334, 4248]. Linéaire [5389]. Linear [3193, 4053]. linear-accelerated [4053]. Lines [752, 5108, 3540]. Linguistic [1214, 1220, 1755]. Linguistics [549, 556, 3624]. Lining [1148]. link [2873]. linked [4724]. Linking [2751]. Links [1768, 1760]. Lint [3878]. Linus [3230]. Lions [3738]. Lipoprotein [1026]. Liquid [4034, 4908, 1793, 1806, 1795, 4270, 4909, 359, 2304, 4326]. Lisa [2572]. Lise [4150]. list [3140]. Listening [1410, 2878]. Literacy [2658]. literally [4779]. Lithium [2150]. Lithography [2394]. Litigating [3047]. Litigation [90]. Little [3262, 5089, 3279, 5288, 4945, 1163, 219, 1959, 4970, 1053, 1781, 3285, 975, 4692, 2389, 4867, 4536, 885]. Livable [1342]. Live [4452, 4809, 647, 5269, 4290]. liver [2983]. Lives [4290, 4291, 2687, 4476]. Livestock [3549]. Living [2654, 2653, 2655, 3355, 1784, 3356, 1712, 5027, 2669, 870, 869, 1474, 3006, 2567, 3082, 4610, 4611, 865]. Lizard [3140, 3150]. LMDS [5256]. Load [4480]. Loathing [2217]. Lobster [2597]. lobsters [2578]. Local [4220, 2272]. lock [3145]. Locomotion [2902]. Logic [2720, 5104, 1526, 1528, 4824]. Logical [3411, 3514]. Logique [5389]. Logistics [1958]. logy [2993]. Lollipop [1920]. Lonely [924]. Lonesome [2054]. Long [63, 4452, 1998, 3941, 4813]. Longer [4809]. Longest [4084, 4067]. Longevity [2227]. Longship [4201]. Look [4423, 4524, 4122, 4757]. Looking [5133, 4854, 5131, 4811, 5350, 541, 3910]. looks [4296]. Loops [2048]. Loose [3153]. loosen [4493]. Lore [4025]. Lorenz [282]. Losing [2513]. Loss [2741, 4007]. Lost [1201, 2560, 2559, 3116, 3120, 4469, 2152, 222, 4671, 3209, 2983, 2555]. Lot [3292, 4793]. Lots [5129]. Lotteries [98]. Loud [4]. Louis [172]. Love [2564, 558, 1993, 4363, 5163, 3481, 3851]. Loves [823, 2958, 2960]. Low [4864, 2061, 4519, 4528, 3037, 2837, 3121, 3401]. Low-Back [4519, 4528]. Low-Birth-Weight [3037]. Low-Density [4864]. Low-Energy [2061]. low-frequency [3121]. low-income [3401]. Luck [3194]. Lucky [1894, 3899, 4682, 3516]. Luddites [2770]. Lunar [98, 1427, 1415]. Lunatics [5043]. Lung [4803, 3774]. lungs [4908]. Lupus [4400]. Lure [3981, 1367, 4025, 3178]. Lurking [5173]. lurks [1084]. Lyapunov [730, 742].

Lying [1164]. Lyme [2058, 2060]. Lyndon [990].

M [4777, 1455, 2947, 617, 1427, 1415]. M-Pire [1415, 1427]. M.D. [4777]. M.I.T. [3830]. MacGuffin [1590]. Mach [5079, 3975]. Machinations [1076, 1223, 1216, 1222]. Machine [18, 20, 874, 3287, 1078, 325, 314, 324, 5011, 2669, 3836, 4975, 2314, 2773, 3491, 3294, 4296, 3852, 3753]. Machinery [3889, 3903, 2063, 5388]. Machines [2358, 5385, 2639, 836, 2641, 2604, 2274, 2231, 2150]. Machos [1312]. Mad [149, 3478]. Madagascar [1319, 1321]. maddening [3478]. Made [3633, 3286, 3566, 4205, 2705, 4215]. Madness [2777, 2753]. Madonnas [2974]. Mag [1646]. Magic [4048, 4059, 4060, 5132, 5309, 3558]. Maglev [3979]. Magnet [3260]. Magnetic [4872, 1644, 5049, 5050, 4705, 5051]. magnetically [3261]. magnetism [5079, 4873]. Magnets [2443, 1649, 1638, 645]. Magnificent [2604, 4030]. Magnitude [3422]. Magruder [5211, 5212]. Mail [4283, 984, 4269, 1886]. Mainstream [1557]. Maintain [400]. Maintaining [4788]. Major [1147, 3347, 4831, 3346, 1141, 353, 4792]. Majority [2940, 2933]. Make [442, 4024, 3940, 4368, 1507, 3105, 55, 3340, 5120, 3977, 4251, 1443, 2243, 3125, 3556, 3752, 4151, 4158]. Makers [1068]. makes [2528, 3051, 3820]. Making [4931, 4741, 2873, 3793, 3586, 4129, 4390, 4577, 4583, 905, 1350, 3876, 2892, 3423, 4330, 663, 1128, 1806, 1795, 2192, 1499, 3588, 4142, 5001, 1649, 1638, 5207, 2114, 3319, 3320, 3905, 4022, 5159, 1837, 5292, 5297, 4349, 3419]. Maladies [3304]. Malaria [3498, 642, 3499]. Male [2763, 1902]. Males [3774, 4453]. Malignant [1591]. Malnutrition [2952]. Malta [1727]. Mammals [3536, 895]. Mammograms [4788, 3339]. mammoth [4906]. Man [1854, 4979, 3498, 823, 184, 682, 2226, 3566]. man-made [3566]. Managed [1975]. Management [1521, 3696]. Managing [3739, 1194]. Manatees [1984]. Mandelbrot [118]. Mangrove [3012]. Mania [1641]. Manic [2275, 4091]. Manic-Depressive [2275, 4091]. Mantle [1827, 1447]. Manuel [3499]. Manufacturing [2639, 4107, 2277, 2282]. Many [4594, 2243, 4786, 1171, 3265, 1305, 1911, 1924, 2389, 4867, 2933, 4759, 4787, 4771]. MAP [700]. mapmakers [4540]. Mapper [4217, 4218]. Mapping [2752, 215, 1465, 1451, 5099]. Maps [1453, 1452, 3893]. marbles [4587]. Margie [3051]. Marianas [893]. Marijuana [650, 4554]. Marine [3403, 4487, 4611]. Mario [4007]. Mark [3423, 3393]. Market [3964, 925, 1263, 2579, 4464, 4722]. Marks [3086, 1403]. Marriage [5326, 2057]. Married [213]. Marrow [3345]. Mars [2964, 3380, 4763, 4287, 4294, 107, 4061, 4470, 3455, 4917, 3375, 4910, 4916]. Marsupials [5068]. Martin [2822, 2845]. Marx [843, 843]. Masers [2276]. Mass [331, 332, 3221, 2397, 4504, 330, 3226]. masses [4775]. Massive [4508, 5180]. Master [3966]. Mastering [1563]. Mastermind [1668]. Masters [542]. Match [3494, 3963, 3492]. Mate [5362]. Material [1109, 1764]. Materials [30, 2639, 5135, 365, 372, 2643, 3502, 1411, 1417, 2630, 2640, 2499, 2495, 4710].

Mates [4317, 4401]. Math [367, 98, 876, 646, 3558, 4151, 4941, 4785, 4588, 4158]. Mathematical [419, 517, 454, 492, 529, 599, 1190, 1268, 1372, 1428, 1505, 1541, 1692, 1768, 1769, 1846, 1925, 2000, 2079, 2158, 2728, 2792, 2862, 2245, 2324, 2501, 2663, 2822, 3228, 3514, 2909, 3017, 3286, 3464, 2966, 3893, 4074, 3608, 4026, 3938, 3750, 3558, 3660, 3803, 3707, 3849, 4151, 4585, 4489, 4433, 4538, 4325, 4275, 4380, 4938, 102, 171, 240, 308, 378, 669, 742, 453, 528, 598, 418, 776, 633, 840, 704, 491, 565, 977, 1267, 1116, 909, 1189, 1043, 1503, 1691, 1371, 1427, 1613, 1540, 1924, 1845, 2157, 2078, 1999, 2861]. Mathematical [2244, 2500, 2415, 2662, 2727, 2323, 2791, 3236, 3513, 3463, 3070, 3365, 2919, 3416, 2965, 3016, 3130, 3285, 3182, 3557, 3749, 3607, 4073, 3892, 3937, 3659, 3706, 4025, 3802, 3848, 3987, 4158, 4488, 4379, 4593, 4159, 4212, 4274, 4546, 4662, 4333, 4708, 4213, 4432, 4748, 5146, 5108, 5258, 4874, 5219, 5346, 5309, 4937, 5073, 4978, 5185, 5037, 2845, 3122, 3175, 3366]. Mathematician [5005, 4687, 1942, 4074]. Mathematics [4594, 4529, 4534, 1635, 1640, 4941]. Mating [4702, 4703, 4473]. Matrix [2822]. Matter [3959, 5266, 1751, 2191, 473, 1922, 2895, 3629, 1703, 2135, 3422, 2016, 1168, 1744, 2994, 4646, 4839, 667, 4775, 5287]. Matters [4577, 4583, 1897, 4850, 1580, 1552, 4192]. mature [3560]. maven [3538]. Maverick [2311, 223]. Maxwell [4895, 4894]. May [63, 712, 3383, 3927, 5085, 394, 4786, 4774, 2833, 4804, 4801, 2759]. Maya [264, 263]. Maybe [5221, 4826, 4793, 3375]. Mayhem [3250]. Maze [3506, 3513, 3514]. MBAs [1042]. Me [2237, 3431, 4536, 5328, 4984, 2244, 2434]. Meal [315]. Mean [4614, 2150, 5208]. Meaning [1264, 366, 4093]. Means [3797, 3767]. Measure [1977]. Measurement [1524]. Measuring [2790, 2860, 3015, 3193, 3891, 4378, 2789, 2859, 2854, 2784, 386, 2196, 2190]. Meat [4127]. Mechanical [588, 2717, 4115, 1361, 5242, 370]. Mechanics [1904, 368]. Medal [3765, 4968, 3764]. Media [3206]. Medical [3083, 5059, 3782]. Medications [4299]. Medicinal [4554]. Medicine [2632, 2232, 2240, 4793, 3603, 3479, 3276, 3281, 5016, 5018, 1066, 2075, 3399, 4740, 2528, 3526, 4184, 4191, 3400]. Medicines [2189, 3653]. Medieval [588]. Medin [3538]. medium [3088]. Meet [4189]. Meeting [2993]. Meets [3963, 2557, 2773]. megabits [3853]. Meitner [4150]. melding [4230]. Melt [3487]. Meltdown [1277, 2747, 4177, 4169]. Melting [2480, 259]. Memorable [662]. Memories [3633, 4705, 3932, 2781, 321, 4562, 4781]. Memory [3634, 1136, 1139, 1945, 4094, 2370, 3722, 5051]. Men [1705, 2886, 2992, 5237, 125, 4134, 4351, 5109, 4782, 4809, 5081, 2604, 4799]. Menagerie [4112]. Mendoza [1028]. Meningitis [2136]. Menopause [4801]. Mental [100, 2874]. Menu [517]. Mephistopheles [891]. Mercury [4760, 4016]. mercy [4029]. Merging [5011]. Merit [2104]. Merology [1845, 1834]. Merrier [3096]. Mesmerizing [2873]. Mesopotamian [2365]. Mesozoic [4403]. Mess [2092]. Message [151]. Messages [1228, 302]. Messengers [1255]. Messy [1278]. Meta [2774]. Meta-Virus [2774].

Metabolism [2859, 2854]. Metacomputing [3737]. Metal [4048, 4059, 4060, 2789, 2784, 3963, 2548, 1869]. Metallurgy [656]. metals [4617]. Metamorphosis [4976]. Metaphor [1013, 1014]. Metaphors [2140, 2139]. Metaphysics [1798, 1792, 1797]. Metastasis [894]. Metastatic [1076]. Meteor [4698, 4699]. Meteoric [302]. Meteorite [4698, 4699, 4696, 4697, 2929, 3722]. Meteorites [98]. Meter [1643]. Metering [2150]. methane [4401]. Methuselah [5001]. Metropolis [5031]. Mexico [3997, 1163, 2345]. MHC [1721]. Mice [1705, 125, 714, 216, 1019]. Michael [3830]. Mickey [5280]. Micro [3198, 3540, 4299]. Micro-fan [3540]. micro-fuel [4299]. Micro-optics [3198]. Microbes [3402, 3375, 2874]. microchip [3198]. Microcreditors [3677]. Microdiamonds [4533, 4542]. Microexplosions [4689]. microfossils [4287, 4294]. Microgears [3875]. Microgram [3174, 3181]. Microlasers [4256, 798]. Micromachine [1500]. Micromechanical [3922]. Micromechanics [1216, 1222]. micrometabolism [2860]. Micromicrophones [5063]. micromotion [3284]. Micron [3278, 3283, 1223, 1222, 1216]. Micron-Size [3278, 3283]. Microorganisms [4647, 4648]. Microplasmas [29]. micropond [4271]. Microprocessors [2611, 4113]. Microquasars [2131]. Microrotors [4895]. Microsatellites [4155, 4868]. Microscopes [476]. Microscopic [5072, 2641, 3801]. Microscopy [2021, 2020]. Microsoft [4197, 4198]. Microstrip [2407]. Microsubs [3984]. Microwave [1169]. Mid [2280, 192]. Mid-Cretaceous [2280]. Mid-Ocean [192]. Middle [1869]. Midnight [2894]. Midst [4458]. Midwaters [2491]. might [4398, 4773]. Mighty [5365]. Migraine [4780]. Migrating [5213]. Mile [3542]. miles [4613]. Military [2984, 4859]. Milk [2852]. Milky [4643, 1316]. Millennium [5355, 4415]. Milliken [124]. millimeters [3670]. Million [376]. Millions [5171, 5190, 4779]. Milo [3538]. Mimic [278]. Mimicry [1591]. Miming [2525]. Mind [3147, 563, 3959, 5342, 5356, 5343, 1133, 1141, 3157, 1136, 1139, 649, 5011, 1156, 1323, 1865, 13, 19, 1142, 1264, 4087, 2773, 2487, 1135, 5287, 4227, 4493, 3940]. Mind-Body [4087]. Minds [3473, 610]. Mineral [4617]. Mines [3098, 3109, 3769]. Minimizing [3679, 3330]. Mining [4267, 4617, 3025]. Minion [3251]. minority [2933]. minting [5079]. Mir [4367]. Miracle [362]. Miracles [2973, 2393]. miraculous [2974]. Mirage [1169, 1502]. Miriam [3095]. Mirror [1806, 1795, 1874, 4541, 3636, 3196]. Mirroring [803, 802, 797]. Mirrors [1793, 3286]. Miscarried [4453]. Mischief [1821]. misdiagnosed [2526]. Misery [675]. misguided [2329]. Misleading [3601]. Mismatch [3782]. Misreading [2519]. Missed [3899, 1019]. Missile [5172]. Missiles [255]. Missing [537, 2016, 2215, 2873, 4787]. Mission [2846, 1870, 4148, 486, 4470]. missions [4776]. Mistake [3577]. Misunderstood [3703, 3716, 3584]. Mitchell [4083]. Mitochondria [4955]. Mitochondrial [3882]. MMT [4414]. Mobile [1632]. Mobility [3968]. Möbius [3849, 4274]. Model [1024, 714, 1443]. Modeling [5259]. Models [1767, 4464, 87]. modem [3538]. modems [4299, 4320]. Modern [2003, 957, 1416, 5371, 407, 5157, 4265]. Modest [500, 223]. modify [4290]. Modules [1632]. Moehlman [3584]. molasses [2517]. Molding [3443, 4050]. Molds [3446]. Mole [2213, 1100]. Molecular [2720, 278, 1821, 644, 1055, 1061, 1259, 1789, 548, 3667, 3353, 994, 1002, 2274, 3446]. Molecules [4747, 4425, 4704, 1217, 1218, 1982, 2401, 2405, 401, 403, 3443, 4627, 3741, 3666, 4541, 396]. Moles [739]. Molina [4007]. Mona [2572]. Monarchs [3935, 3936]. Money [4931, 1573, 2485, 3638, 1609, 4239, 1837, 2520, 3639, 4875]. Monitor [262, 3648, 4735]. Monitoring [3069, 3936, 4873, 4527, 4532]. monitors [4637]. monkey [1082]. Monkeying [3072]. Monkeys [1264]. Monks [4546]. Monogamy [1489]. Monolith [5365]. Monopoly [3071, 4198, 4197, 3417, 3065, 3409, 3070, 3416]. monsters [2731]. Monstrous [4686, 4687]. Montalcini [1314]. Months [4367]. Monticello [1614]. Moo [3514]. Moo-ving [3514]. moods [2753]. moon [3893, 3922, 84, 1943, 680, 3892]. Moonball [1515]. moons [4757]. Moonshine [4686, 4687]. Moor [3920]. Moor-Jankowski [3920]. Moose [3575]. Moose-suit [3575]. Moral [5181]. morality [3368]. Morass [3824]. morose [3466]. Moroto [3824]. Morrison [2796, 4076]. Morse [3896, 560]. Morsels [517]. Mortality [2886, 2992]. Mortem [465]. Most [2590, 1485, 4769, 5309, 1118, 2870, 3741]. Most-Perfect [5309]. Mostly [2604]. Mother [278, 3221, 3226, 1053, 2919, 2909, 3753, 5347]. Mothers [3401]. moths [2983]. Motif [1423, 357]. Motion [4670, 4259, 737]. Motion-Picture [4670]. Motor [287, 293]. Mound [4882, 4883, 3946]. Mountain [1173, 961, 3168, 3166]. Mountains [3700]. mounting [3801]. Mouse [3255, 5280, 2032]. Mouth [2180]. Move [3085, 3985, 234]. Moved [2262]. movement [3147]. Movements [3278, 3283, 2943]. Moves [1554]. Movies [1431, 3125, 3752, 4661]. Moving [4794, 4328]. mph [4053]. Mr. [1478, 234]. MRI [3777, 2309]. MS [216]. Much [3262, 3457, 3461, 2872, 2983, 4794, 4793, 4789, 1544, 4772]. Mud [893, 4551, 2478]. Muffling [304]. Mug [3836]. Multicultural [3435]. Multifractal [4921]. Multilingualism [3649]. multiple [3893]. Multiplicity [4817]. Multiregional [958, 960]. Multiverse [4774]. mummies [4878]. Munk [2226]. Murder [786, 1158, 1189, 4028]. Murderous [5068]. Murky [143]. Murphy [2862, 3702]. Muscle [3105]. Muscles [4227, 3089]. Muscular [5359]. museum [3230]. museums [2329]. Music [3865, 5064, 2768, 2757, 3940, 3911]. Musical [1104]. Mussel [3250]. Musseling [1165]. Mussels [394, 3247]. Mustard [3667]. Mutable [988]. Mutant [4504]. mutating [3909]. Mutation [1716, 988]. Mutations [5000]. Mutt [716]. Mutual [2446]. My [4430, 3868, 3625, 4748, 774, 3729, 3857, 1923, 2264, 2296]. myriad [4759]. Mysteries [252]. Mysterious [4698, 4699, 3304, 4724]. Mystery [786, 3277, 2538, 607, 2215, 4517, 5137, 220, 2306, 4403]. myth [4205, 4215].

n [2947]. Nada [1513]. Naked [4134, 2213, 1100]. Nambu [2271]. Name [3391, 3076, 3371, 2177, 2091]. Named [2078]. Naming [2169]. Nanofuture [5]. Nanolasers [4258]. Nanophase [3502]. Nanotechnology [3061, 3066]. napkin [4339]. Napoleon [2067]. Narrow [410, 3088]. NASA [2983, 3851]. Nasty [219]. Nation [5022]. National [4177, 4169, 4968, 1328, 3764, 3765, 4732, 5172]. Nations [1980, 4613]. Native [1214, 1220, 2329]. Natural [3129, 4270, 1610, 4194, 764, 3108, 4198, 4695, 3587, 4437]. Natural-Born [4194]. naturalist [3364, 3095]. Nature [2907, 278, 3104, 2916, 3217, 3222, 1873, 1868, 1872, 4146, 4959, 5339, 2906, 3201, 4277, 5265, 4679, 4775, 5347]. Naughtiest [2391]. nausea [2983]. nautical [4613]. Navigate [864]. Navy [4689]. Neanderthal [4127, 3917]. Neanderthals [4277]. Near [588, 4617]. Nearly [4603]. Neatness [719]. Nebulae [995]. necessary [3911]. Neck [63]. Need [4083, 4783, 3976, 4070, 3560, 4493, 4014]. Needles [3444]. Needless [51]. Needs [1235]. Nefertari [5245]. Negative [70]. Neglected [3176, 3182]. Negotiating [1018]. Negroponte [2606]. neighboring [4757]. nemesis [3873]. Neptune [2686, 4766, 3371, 2696]. Nerve [5066]. nerves [2751]. Nervous [3796, 14]. Nesting [3227, 3234]. Net [3490, 2801, 2802, 2504, 3024, 5017, 4843, 1143, 115, 3288, 3639, 3025]. Nets [3310]. Netted [3101, 3259, 3315, 3398, 3496]. Network [5290]. Networked [722]. Networking [3258, 3392]. Networks [732, 3115, 725, 721, 2608, 2615, 1224, 720, 726, 2902, 4324, 1757, 727, 731, 729, 724, 728, 717, 668, 2612]. Neural [2902, 1143, 430, 1893, 3405, 191]. Neurobiology [891, 1454, 4095, 4422, 3805, 3624]. Neuron [2117, 2108]. Neurons [1154, 1120, 2983, 3531, 4676]. Neuroscience [3902, 3889, 1985, 1988, 5388]. Neutral [896]. Neutrino [4504, 155, 2799]. Neutrinos [5180, 3032, 3034]. Neutron [337, 2404, 3999]. Never [5380, 2721, 1625, 2430, 2764, 2727, 2870, 2728, 3895]. Never-Ending [2721, 2764, 2727, 2728]. New-World [4134]. Newborns [2852]. Newcomb [4649]. News [5194, 5274, 3820, 4399, 4509, 4900, 3477, 3298, 3192, 4039, 3996, 3908, 3571, 3719, 3862, 4397, 4291, 4993, 5228, 4231, 3819, 3379, 3429, 3620, 4677, 4449, 4346, 5050, 4953, 5134, 4843, 5318, 3141, 3084, 3766, 3244, 3030, 4176, 4556, 3527, 4402, 4124, 3951, 3676, 4721, 4625, 5119, 4891, 5156, 5084, 4507, 4476]. Newton [4492, 4054]. Next [4262, 4261, 4263, 4890, 3571, 3215, 3216, 5387, 3952, 3085, 2336, 4910, 2807]. Next-Generation [3215, 5387]. Nice [3689, 2843]. Nicest [1407]. Nicholas [2606]. Niels [2410]. Night [884, 4626]. Nightmare [4460]. Nikola [3709]. Nikolai [899]. nine [3670]. Nit [6]. Nitric [993]. Nitrogen [3769, 506, 5264, 5265, 3844]. No [2582, 2929, 5332, 2711, 3544, 2167, 5290, 2177, 1782, 2994, 1294, 875, 2266, 4041, 809, 4193, 4012, 4014, 2343, 4734, 3389]. No-Name [2177]. No-Polluting [2167]. Noah [4943, 4944, 1280]. Nobel [3526, 2172, 817, 2174, 4007, 4840, 4122, 4841, 1778, 3525, 4123]. Nobels [3481]. Noise [4368, 2954, 2561, 2558, 2553, 3940]. Noisy [1126]. Non [2213, 2036]. Non-Naked [2213]. Noncontrol [4004]. Nonimaging [514]. Nonlinear [1320]. Nonviral [3794]. Normal [1549]. North [5267]. Northwest [2850]. noses [2761]. Nosing [5362]. Note [3864]. Notebooks [2723, 2722]. Notes [5150, 5323, 1778, 3917]. noteworthy [4712]. Nothing [3457, 3461, 1091, 2264, 2439, 541, 4722, 4721, 3462]. Notice [3809]. Nots [2379]. Notworking [3257]. Nova [2132, 4769, 2236, 2238]. Novelty [3045]. Novgorod [62]. Nowhere [2223, 3969, 2733]. Noxious [3772]. Nuclear [615, 3779, 4348, 471, 4357, 2411, 397, 4018, 793, 298, 3141, 2550, 2299, 5215, 4930, 4932, 462, 3786, 3052, 4150, 1558, 3168, 2885, 2898, 3124, 3118, 4604, 2569, 2523, 3481, 2933, 3621, 4358, 5085, 4437, 3166]. Nuclei [2450]. Nucleon [5137]. Nucleotides [1126]. nucleus [87]. nuke [4895]. Number [3176, 4941, 4945, 473, 900, 909, 3182, 412]. Numbers [471, 4048, 4059, 4060, 4210, 3598, 4243, 214, 2971, 4715, 4716, 1845, 1834, 48, 3072]. numerical [3175]. Numskull [3072]. Nurture [5339]. Nurturing [1873, 1872, 1868]. Nutrients [3772]. Nyet [606].

O [1323, 4438, 2947, 5351]. Obesity [3282]. Objective [940]. Objects [1335, 1869]. Oblivion [3979]. Observations [3227, 3234, 4773, 396]. Observatories [84]. Observatory [1720]. Observe [2061, 198]. Obstacle [1397]. Obstacles [3792, 3360]. Obstructive [3960]. Obvious [2603]. Ocean [4619, 4620, 4615, 5243, 2491, 2403, 2406, 327, 336, 2226, 2699, 4299, 4617, 4599, 4601, 4600, 192]. Oceans [4598, 4602, 1827, 4486, 4607, 4603]. octane [2873]. Odd [232, 3230]. odds [3366, 4026]. odors [2748]. Odysseus [924]. Odyssey [3135]. Oeuvre [5236]. Off [4018, 341, 5017, 1555, 647, 700, 4455, 4136, 3017, 3311, 4236, 1037, 4634, 3312, 560, 2386]. Off-Center [647]. Off-the-Net [5017]. Off-the-Rack [5017]. offer [3299]. Office [1918, 4799, 3400]. Often [3318, 3329]. Oh [3612, 2343]. Oil [4268, 2583, 4262, 4261, 4263, 4264, 4267, 4963, 4695, 4011, 1294, 3710, 4009]. oilseed [4531]. Oily [2821]. okay [4807]. Old [4356, 5347, 542, 3312, 1862, 1979, 2233, 185, 4965, 3236, 4350, 3290, 3422, 3421, 4052]. Older [4813]. Oldest [2233]. Oliver [206, 193, 205]. Olympic [3163, 3165]. Olympics [1116, 1106]. omniscience [2015]. On-Line [2399, 5134, 4057, 3178, 4914, 3622, 5334]. Once [4941, 2267, 3909]. Oncogene [50]. One [5022, 406, 2293, 911, 1074, 4979, 3264, 92, 4057, 3581, 2596, 94, 4443, 2383, 2300, 2277, 2282, 1960, 2336, 4936, 4774, 2281, 4110]. One-on-One [4057]. One-Step [94]. Ones [1894, 719]. Only [3313, 911, 4320, 5157, 4902]. Onnes [3654, 3655]. ons [3881]. onset [4801]. onstage [2870, 3178]. Oops [4713, 4714]. Oort [4580, 4589]. Ooze [3396]. Opaque [4574]. Open [596, 141]. OpenDoc [2113]. Opens [5034]. opera [3231]. Opinion [140]. Opium [650]. Oppenheimer [2496, 2493]. Opportunity [1461]. optic [3575, 3901]. Optical [2529, 2608, 2615, 972, 1765, 3900, 653, 4722]. Optics [1946, 999, 514, 3198]. Options [213]. Oral [580]. Oravske [249]. orbit [3247]. Orbiting [221, 4756]. Order [3641, 313, 3642, 1036, 4134, 429, 2982]. Ordered [4305].

Ordering [696, 697]. Ores [585]. Organisms [2859, 2854]. Organize [3272, 3268]. Organized [436]. organizing [3433]. Organs [2635, 5014, 4137, 4360]. Orgasm [4787]. Origami [4938, 4937]. Origin [828, 657, 1910, 121, 2102, 4199, 2874, 4385]. Originate [957]. Origins [765, 1214, 1220, 1060, 4956, 4603, 1721, 3453, 4701, 2685, 3368]. Ornithology [3704, 3705]. Orphan [1184, 301]. Oscillators [1726]. Osteoporosis [4810, 4808]. Other [3057, 4758, 440, 2269, 3391, 4927, 4522, 3058, 2174, 2834, 5068, 5386, 32, 4040, 4756, 4724]. Ounce [2082, 2154]. Our [4757, 4774, 4458, 5337, 4717, 4340, 120, 4759, 4277, 4603]. Outer [3399]. Outhouse [3142]. outlet [2826]. Outline [2642, 2652]. outsmart [3927]. Outsmarting [4340]. outstanding [3526]. Outstrips [2479]. Ovarian [4796]. overblown [3767]. Overcoming [1010, 3792]. overload [3261, 3769]. Overlooked [2806, 3175]. Overpopulation [3030]. overrated [3370]. Overreaction [5085]. overseas [2577]. Overtaking [5315]. overtaxed [4611]. overthrowing [4756]. overtones [3710]. Overturned [4677]. Overview [188, 285, 361, 3526]. Own [1683, 1679, 1682, 3512, 4024, 4053, 4584]. Owns [3223, 3229]. Oxide [993]. Oxygen [55, 5175, 5174]. Ozone [923, 2216, 4006, 618, 2579, 4007].

P [3201]. P. [1455]. P.R. [2935]. Pace [4983]. Pacific [4600, 2850]. Pack [1647, 4606]. packing [4770]. Pain [1895, 4519, 4528, 3359, 51, 591, 4464, 4795]. painful [4781]. Painkiller [701]. paint [4527]. Painted [4185]. Painter [4279, 4280]. Painting [800, 810, 796, 795]. Pakistan [4501]. Paleoanthropologist [4906]. paleontology [3072]. Panacea [3209]. Panama [5320]. Pandemic [992]. Panel [1287, 1384]. Pangea [2230]. Panoramas [3790, 3789]. Panos [172]. Paper [4570, 3827, 2077, 4562]. Papers [3149]. Papua [4617]. Parable [2417, 5077]. Parachutes [169, 4755]. Paradise [1201]. Paradox [1285, 1381, 2064, 4047, 3697, 4663]. Paradoxes [1104, 2145]. Parasite [3036]. Parasites [1535]. Parasitic [4019, 1288, 1385]. Parasitology [865, 869]. Parched [578]. Parental [326, 3880]. parents [4793]. Parkinson [3732, 3547]. Parlor [2417]. Parrot [3039]. Parrots [4824, 3825, 3823]. Parsing [5128]. part [4774, 4698, 3052, 4697, 3124]. Particle [924, 1798, 1797, 1094, 1792]. Particles [687, 688, 896, 1999, 1990, 4410, 4770, 3433]. Partly [4563, 1371, 1360]. Parts [4913]. Pass [2298, 2995]. Passage [82, 3821]. Passenger [2605]. Past [10, 2414, 2408, 2413, 3815, 3968, 2533, 2933, 4757]. Pasteur [2677]. Patarroyo [3499]. Patch [926, 5163]. Patent [970, 197, 2484, 4688, 2588, 3261]. Patenting [509]. Patently [2603]. Patents [1072, 4689]. Path [466]. Pathfinder [4470]. pathological [4028]. Paths [966, 977]. Patient [3253]. Patricia [3584]. Pauling [3230]. Paulos [4941]. Pay [2892, 1464]. Paying [3244, 1805]. Payoffs [1079]. PC [4116]. Pcb [4182, 4172, 4178]. PCR [4390]. PCs [4224]. Peace [2897, 1695]. Peaceful [3141, 3138]. Peaked [713]. Peary [180]. pebbles [4759]. peculiarities [4759]. Pedals [3856]. Peering [625, 624, 619]. Pelli [4086]. pelvic [4778]. Penalty [212]. Pendulums [4881]. Penny [2670]. People [820, 841, 1269, 3734, 4793, 4611]. Peoples [801]. Perception [474]. Perennial [3967]. Perfect [3605, 902, 897, 901, 5309]. perform [2870]. Performance [169]. peril [3911]. Perils [4612, 3043, 5173, 1672, 5229]. Periodic [4578, 4586]. Perishing [5208]. Permanent [3102]. Permanently [2529]. permeates [4235]. Perpendicular [1557]. Perplexity [2453, 2445, 2452]. Persistence [5051]. Persistently [2421]. Person [2372]. Personal [4322, 4934, 2264, 4625, 2802]. Personalizing [4624]. Persuasive [2229]. Peru [656]. Perverse [1539]. Peso [2433]. Pest [2596]. Pfiesteria} [5173]. Ph.D. [5280]. Phantom [967, 4096, 4640]. Pharm [267]. Pharmacology [519, 513]. Pharmacy [1009, 4808]. Phase [3155, 1940]. phases [4936]. Phenomena [2061, 4741]. Philadelphia [4535, 4544]. Philip [4076]. philosophers [4493]. Philosophy [1071, 1065, 1070]. Phone [2310, 2546, 5251, 5252, 1743, 4042]. phones [2516, 4320]. phonetic [3126]. Photochromic [554, 553]. photocopiers [3425]. photoelectricity [2969]. photographic [396]. photography [3230]. Photoreconnaissance [3008, 3007]. Photorefractive [328]. Photos [4442]. Photosensitive [554, 553]. Photovores [1131]. Phrenologists [5043]. Phylis [4076]. Phyllis [4777]. physical [4559, 4774]. physician [3419]. Physicians [3082]. Physicist [4233, 1351, 1348, 4244]. Physicists [1048, 1054, 4745]. Physics [2927, 3814, 1830, 4511, 924, 1065, 1070, 2135, 1792, 1797, 1455, 1866, 2971, 5349, 2904, 2914, 974, 5336, 1900, 3798, 4064, 3466, 3019, 3526, 4942, 2913, 5350, 172]. Physiology [474, 54, 53, 2554]. Piano [2869, 4885, 2870]. Pick [4905]. Picker [6, 2181]. Picking [3395, 2214]. picnic [3369]. Picosecond [4152]. PICS [3309]. pictorial [4759]. Picture [4670, 3605, 902, 901, 4477, 4002, 897]. Picture-Perfect [3605]. Pictures [3224, 3812, 2722, 4648, 687, 3007, 2186, 4316, 3791, 5138, 53, 479, 3062, 4743, 2448, 2848, 1676, 263, 795, 2020, 3273, 1561, 1907, 1790, 1217, 3172, 201, 3789, 4703, 996, 1357, 398, 2139, 2908, 2362, 4973, 3447, 3644, 895, 1101, 553, 1452, 4154, 4167]. Pie [5169]. pieced [4772]. Piecemeal [1072]. Pieces [2214, 2580]. Pied [791]. Piel [1194, 1811]. Pigments [800, 810]. Pigs [4194, 3083]. piles [4381]. Pill [4407]. Pills [4625]. Pinball [788]. Pinch [4515, 4526]. Pink [3199]. Pinker [5127]. Pinn [4777]. Pinning [3766]. pioneer [3201, 2979, 1870]. Pioneering [4725]. Pipe [5257, 1803, 4309]. Piper [791]. Piracy [4254]. Pirate [5230]. Pirates [5073, 1831, 1835, 1836]. Pire [1415, 1427]. Pisa [2849, 2848]. pit [3909, 4914]. Pitch [1104, 4436]. pitcher [4770]. Pitches [3945]. Pitohui [1308]. Pits [2035]. Pizza [2417]. Place [2711, 2765, 3741]. Placebo [4153]. Places [918, 2881, 2790, 4207]. Plague [559, 2588]. plagues [3282]. Plan [663, 225, 5320, 4290]. Plane [5205, 2986, 523, 3585]. planes [2818]. Planet [1373, 4620, 3132, 4016, 1973, 5191, 4603, 1194, 286, 288]. Planet-Tude [3132]. Planetary [4759, 995]. Planets [3057, 5113, 5213, 4756, 1342, 3058, 5386, 2874, 3433, 3893, 3722, 4757, 3127, 3562]. Plankton [950]. Planning [5239, 3002]. plans [3138]. Plant [848, 859, 860, 5204, 5143, 4192]. Plantibodies [4008]. Plants [5107, 1634, 3871, 2776, 55, 2958, 2960, 3656, 2964, 4531]. Plaques [559]. plasma [4464]. Plastic [3692, 807, 3495, 266, 2030, 4172, 4178, 3610]. Plastics [3743, 2498, 2499, 4114, 1501, 2495]. Plateau [512]. Play [317, 1687, 2161, 163, 3241, 2827, 3869, 668, 2728, 4074]. players [4590]. Playful [3807]. Playing [3010, 3153, 3689, 2896, 4179, 2157, 2146, 3016, 4662, 48, 3660]. Please [2298, 2436, 4012, 875, 2995]. Pleasures [3348, 3363, 4272]. Plenty [1091, 3364]. Plotting [3085]. plowshares [2150]. Plug [2827]. Plugging [2580]. Plumage [749]. plump [4807]. Plums [517]. Plunder [3825]. plus [4777]. Pluto [4767, 190]. Plutonium [2298, 2984]. pneumatic [3231]. Pneumonia [3582, 3575]. Poetry [4336, 4337, 671]. poets [3830]. Poincaré [3563]. Point [4935, 2281, 2277, 2282, 577, 4064, 1194, 4498, 4110]. pointillist [3369]. Pointing [920, 4497]. points [259]. Poison [352]. poisoning [3634]. Poland [925]. Polar [4602, 618, 1523, 71]. Polarized [4628]. Pole [2758, 2933, 5267]. Police [75, 3309]. Policies [2946]. Policy [732, 578, 725, 194, 1110, 726, 1717, 727, 792]. Polio [4314]. Polishing [4184, 4191]. Political [545, 2394, 4450]. Politics [4133, 4182, 2176, 2819, 3624]. Pollinated [1634]. Pollutants [4208]. Polluting [2167]. Pollution [3302, 3685, 751, 2882, 4056, 1666, 2150, 3368, 4053, 4908, 1666]. Pollution-Purging [4056]. Polly [3034]. Polyester [3586]. Polygraph [5229, 4080]. polygraphs [4081]. Polymer [63]. Polymerase [121]. polymers [3492]. Polymorphism [1721]. Pond [4272, 4903]. ponders [4852]. Ponds [3348, 3363, 3364]. Pool [1004]. Poor [2372, 1822, 3592]. Pop [3520, 3521]. Poplars [4056]. popular [3401]. Population [856, 1826, 2272, 3630, 1130, 1934, 4729, 1810, 3844]. Populations [2356]. Pork [5286, 3092, 2394, 2997]. porpoises [3311]. Portable [5144]. Portrait [2782]. Positioning [2953]. Positive [64, 681]. Possibility [199]. Possible [5364]. Post [4314]. Post-Polio [4314]. Postal [465]. Poster [5131]. postmodernist [4028]. Postponed [5341]. Pot [389, 3194]. Potato [1145]. potatoes [3195]. potential [4617]. Potholes [972]. Potpourri [73]. Pound [5160]. Poverty [2952, 2272, 3677]. powder [3875, 4381]. Power [615, 1687, 3692, 197, 298, 1678, 3495, 5143, 4849, 3734, 1647, 2075, 2365, 1453, 1452, 2809, 4236, 1673, 4248]. power-line [4236, 4248]. powered [4771]. Powerful [131, 1483]. Powerhouse [1248]. Pox [2700, 2682]. Practical [1538, 5130, 4938]. Prairie [1489]. Praise [4031]. prankster [4244]. Pre [898]. Pre-Columbian [898]. Precious [1869, 4611]. Precision [4977, 4327]. prediction [2809, 3370]. predictions [4287, 4294]. Preeclampsia [4795]. Pregnancy [4795]. Prehistoric [1727]. Prelude [4995]. Premature [3337]. Prenatal [4788, 3540]. Preservation [2414, 2408, 2413, 1373]. Preserve [1465, 1451]. Preserves [820]. Preserving [4567, 4575, 5245, 5107, 3651, 4161]. Presidents [3439]. Press [543]. Pressure [856, 4919, 4332, 3442, 1130, 3184, 3911]. pressures [3180]. Presto [3558]. Pretesting [4892, 4891]. Pretty [91]. prevent [4553]. Preventable [1293]. Prevented [5055]. Preventing [2150, 4262, 4261, 4263, 4482, 1293]. Prevention [3326, 2082, 2154, 4407, 2367, 2360, 2366, 3623, 5055, 3327]. Preventive [2232, 2240]. Price [1921, 3429, 2387, 2367, 2360, 2366, 2839, 3173]. Pricing [2153]. Primate [3961, 1565]. Prime [3660, 2484, 4715, 4716, 4611]. Prime-time [3660]. Primeland [3749]. primers [2666]. primes [3750]. Primordial [3396, 3501]. Principle [2855, 2156, 2861]. Printed [2077]. Printers [5268]. Prion [2228]. priorities [3863]. Prisoner [2454, 2447]. Privacy [2775, 1105, 4851, 3024, 4845, 610, 4015, 4914, 2802, 3025]. Private [4913, 1994]. Privatizing [2065]. Privileged [1645]. Prize [2897, 3577, 2174]. Prizes [3526, 4122, 1313, 4840, 4841, 2172, 817, 3525, 4123]. Probability [3071]. Probably [4724]. Probe [595, 5195]. probes [4724]. Probing [3271, 3275, 3399, 4995]. Problem [155, 1144, 4088, 4522, 784, 437, 344, 4124, 3662]. Problematic [2488]. Problems [3967, 247]. Process [2023, 3811]. processing [2984]. Prodigies [4820]. Producing [2283, 2279]. Production [4268]. Productivity [2152, 3083, 4510]. Products [5025, 724, 5026]. professional [3173, 3610]. professional-quality [3173]. Professor [1849]. Profet [3051]. Profile [545, 2204, 2687, 2753, 2822, 2251, 2328, 2375, 2462, 2508, 2584, 2982, 3051, 3201, 2932, 3154, 3095, 3449, 3267, 3400, 3873, 3538, 3688, 3730, 3584, 3920, 3777, 3830, 4007, 3632, 4051, 4410, 4180, 4190, 4506, 4233, 4559, 4296, 4462, 4358, 4246, 4304, 4411, 4461, 4852, 5005, 4906, 395, 254, 434, 1520, 4188, 4732, 4634, 2714, 3316, 4135, 4687, 2531, 323, 1314, 4303, 579, 1557, 2095, 2311, 508, 2897, 2606, 2271, 4244, 223, 719, 1593, 2487, 2845, 2226]. profit [2150]. Profitable [1297]. profits [4254, 4299]. Program [13, 19]. programmers [3830]. Programming [3394, 3396]. Programs [32]. Progress [4893, 60, 580, 5286, 56, 1861]. Project [4180, 4190, 5175, 4000, 4188, 5174]. Projectors [4670]. Projects [5380, 4831]. Promise [4612, 5012, 5013]. prompts [5195]. Proof [2961, 1639, 1635, 1640, 2965]. Proofreading [575]. Proper [919, 4476]. Property [1879]. Prophet [3872]. Proposal [889, 500]. Prospectors [3727]. Prospects [4483, 2981, 4476]. Prostate [1867, 4742, 3340]. Prosthetic [1843, 4637]. Protect [676]. Protecting [2489]. Protects [2852]. Protein [2319, 5384, 617, 115, 437, 794, 790, 595]. Protein-Based [2319, 5384]. Proteins [1632, 1063, 642, 1062, 2580]. Proterozoic [761]. protesters [4895]. protests [5195]. prototypes [2984]. Prove [3993]. proves [4553]. Provide [4480]. Provinces [1631, 3933]. pseudoscience [3178]. Psychiatry [3652]. Psychic [715]. Psychological [3357, 3356]. Psychologists [2585]. Psychology [3510, 404, 2927]. Pterosaurs [3729]. Public [732, 2984, 2351, 725, 2065, 726, 727, 3246, 792, 4177, 4254, 4169, 4175]. Public-key [2984]. public-relations [4177, 4169, 4175]. Publishing [5208]. Puffery [3301]. pull [3195]. pulling [4724]. Pulmonary [3960]. Pulse [4872]. Pump [3445, 2562]. Pumping [4125]. Punctuated [83]. Punishment [2831]. Purchase [2961, 2965]. Purging [4056]. purpose [2977]. Pursuit [3113, 3117, 4090, 2669, 4327]. push [4954]. Pushing [3213]. Put [698, 1478, 3306]. puts [2818, 3078, 4212]. Putting [1424, 1535, 2257, 161, 5064, 4207, 2541, 2844, 2397]. Puzzle [2356, 2853, 4089, 4918, 3786, 5073, 2909]. Puzzles [3987]. Puzzling [517, 1248]. Pyramid [4593]. pyramids [3895, 4585]. Pyrotechnics [237]. Pythagoras [217].

QCD [1517]. QED [1517]. Quads [3010, 3016, 3017]. Quakes [1876]. Quality [3202, 3173, 3193]. Quandary [4130, 3634]. quantity [4775]. Quantized [1122]. Quantum [1904, 2874, 1249, 1172, 1001, 5282, 5231, 5061, 1830, 778, 4425, 861, 825, 1409, 1071, 1070, 1856, 3454, 2717, 4115, 1322, 3426, 645, 788, 1780, 2888, 4942, 4627, 3005, 1065]. Quantum-Mechanical [2717, 4115]. Quark [2387, 3929, 1132, 1282]. Quarks [2962]. Quarry [710]. Quarter [4529, 4534]. Quarter-Century [4529, 4534]. quarters [4603]. quartz [4326, 4336]. Quasar [616]. Quasars [3575, 4423]. Quasicrystal [7]. Quasicrystals [547]. Quasimodal [2535]. Quasispecies [1522]. Quazars [3010, 3016]. Qubit [5164]. Queens [5292, 5297]. Queer [3988]. Quest [2357, 1282]. Question [3568, 2682, 3083]. Questioning [614]. Questions [353, 1710, 639, 4772]. Quick [5041]. Quickens [3314]. quite [4613].

R [2687, 4852, 4732, 2714, 3169, 3172, 4953, 520]. R&D [1717]. R.I.P [2677]. R.I.P. [39]. Rabies [1031]. RAC [1037]. Race [2173, 230, 226, 229, 1943, 4160, 939]. racers [4464]. Races [5248, 4506]. Rachel [4205, 4215]. racial [4855]. Racing [1246, 3979]. Rack [5017]. Radar [2933, 3594, 2937]. Radiation [4741, 262, 152, 2571, 4227, 4235]. Radical [153, 3586]. radicalism [4312]. Radio [4329, 1485]. Radioactive [4145, 3624, 4175]. Radioactivities [80, 88, 87]. radioactivity [4881]. radios [3922, 1344]. radiotherapies [4348]. Radium [1566]. Radon [2012]. rages [4771]. Rail [2616, 2625]. railroad [5113]. Rain [3890, 3507, 2989]. Rainbow [4979, 711, 2940, 4266]. rainbows [4610]. Raindrop [1005]. raindrops [3891]. rained [4603]. Raising [1637, 3847]. RAM [5051]. Raman [3316]. Rampaging [4009]. Ramsey [239]. Ran [555]. Randi [2487]. Range [2937]. ranging [4616]. Ranking [3466]. Rape [4781]. Rapid [1058]. Rapids [1941]. Rare [1092]. Rat [3255]. Rate [1208, 1213, 4773, 4775, 4724]. Ratings [3471]. Rational [24]. Rats [2201, 2213, 1100]. Rattles [1786]. ravaged [4604]. Raw [5179, 5135, 4099]. Ray [3842, 1720, 4055, 476, 502, 4915, 1675]. Raymond [3777]. Rays [4741, 4241, 3546, 4770, 3579, 3393, 2983, 4845, 2979, 2787]. re [3895]. re-creates [3895]. Reach [147, 4734, 489]. Reaches [3399]. Reaching [4933, 2945]. Reaction [821, 121, 3419]. Reactions [2313, 5084]. Reactor [964]. Reactors [130, 4437]. Read [4418, 4419]. Reader [563, 1643]. Readers [2866, 4076, 2863, 3515, 3517, 4075, 4750, 3473, 1766]. readies [3528]. Reading [628, 3147]. Readings [3157, 3748]. Ready [5003, 2112]. Real [1180, 1176, 1179, 4466, 5097, 4828, 546, 1943, 5361, 4223, 2219, 2885, 2898, 4003, 2674, 2981, 3089]. realities [4774]. Reality [445, 439, 444, 2179, 1856, 2619, 2613, 2730, 3895, 4613]. Really [2209, 3829, 2266, 2933, 5319]. Reap [3093]. Rearrange [771]. Reason [3522, 5262]. Reasonably [3522]. Reasoning [4823, 4237]. Reasons [3295]. Reassessing [4127]. rebellion [3289]. Rebellious [2823]. Rebottling [4357]. Recall [1058]. Receive [1297]. Recently [3101, 3259, 3315, 3398, 3496]. recipients [3401]. Reciprocity [1615]. Recognition [2451, 1317]. Recognizes [1596, 1597]. Recognizing [3761]. Recollections [2550]. Recombinant [2307]. Recommend [4712, 4665, 4711, 4878, 4942, 4877, 4981, 5039, 5075, 5110, 5261, 5311, 5348]. Recomputing [3545]. Reconnaissance [435]. Reconsidered [1052]. Record [2443, 543, 264, 263, 5373, 2906, 4053]. Recording [2907, 2916, 3986]. Records [1484]. recover [4175]. recovers [4177, 4169]. Recovery [360, 3530]. Recreation [1769]. Recreational [4487, 4529, 4534]. Recreations [1190, 1268, 1372, 1428, 1505, 1541, 1692, 1768, 1846, 1925, 2000, 2079, 2158, 2728, 2792, 2862, 2245, 2324, 2501, 2663, 3228, 3514, 2909, 3017, 3286, 3464, 2966, 4074, 3608, 4026, 3558, 3660, 4151, 4585, 4489, 4433, 4538, 4325, 4938, 34, 102, 171, 240, 308, 378, 669, 742, 453, 528, 598, 418, 776, 633, 840, 704, 491, 565, 977, 1267, 1116, 909, 1189, 1043, 1503, 1691, 1371, 1427, 1613, 1540, 1924, 1845, 2157, 2078, 1999, 2861, 2244, 2500, 2415, 2662, 2727, 2323, 2791, 3236, 3513, 3463, 3070, 3365, 2919, 3416, 2965, 3016]. Recreations [3130, 3285, 3182, 3557, 3749, 3607, 4073, 3892, 3937, 3659, 3706, 4025, 3802, 3848, 3987, 4158, 4488, 4379, 4593, 4159, 4212, 4274, 4546, 4662, 4333, 4708, 4213, 4432, 4748, 5146, 5108, 5258, 4874, 5219, 5346, 5309, 4937, 5073, 4978, 5185, 5037, 419, 454, 492, 529, 599, 3893, 3938, 3750, 3803, 3707, 3849, 4275, 4380, 3122, 3175, 3366]. rectangle [3849]. Red [2024, 2738, 782, 904, 2889, 1495, 2488, 3637, 4077]. Red-Banner [1495]. Redeeming [1361]. Redesigning [1035, 1034, 1029]. Redfield [2753]. redirects [3400]. Rediscovering [5091]. Redoubled [4386, 4387]. Redshift [9]. Redux [2432, 1614, 180]. Reef [4610, 3288]. References [907]. Reflecting [915]. Reflections [277, 4997, 1780]. Reform [1212]. Refrigerators [3552]. Refugee [892]. Regain [1073]. regenerating [4676]. Regrowing [3531]. regulate [3741]. Regulating [4015, 4660, 2980]. Regulation [4014, 548]. Regulations [2327]. Regulators [4550, 1174]. regulatory [4509]. rehabilitated [2756]. Rehydration [580]. Reigns [2592]. Reinvention [4570]. Reivews [1885]. Related [257]. relations [4177, 4169, 4296, 4175]. Relationships [4379]. Relative [917, 1006]. Relatively [3090]. relativism [4233]. relativity [4233]. Relaxation [2504]. Release [543]. Reliable [3115]. Relic [4061]. Relics [2182, 2327]. Relief [2947, 4297, 117, 3299, 3302, 3941]. Religion [4189, 5217]. Relinquishing [2182]. Reluctant [3170, 579, 2248]. remain [4476, 4772]. Remains [1003, 994, 1002, 4892]. Remedy [4557]. Remote [335, 199, 3106].

Remote-Control [335]. Rendezvous [850]. René [582]. Repair [2093, 3106, 4398]. Repairing [5214]. Repealing [4325, 4333]. Repeating [1542]. Replacement [1828, 3332, 3333, 4802]. Replacing [5144]. Replicating [1982]. Replumbing [5158]. Report [3671, 3813, 4085, 4261, 4263, 4319, 5175, 5012, 851, 4653, 5248, 4065, 3642, 4651, 4475, 4861, 3792, 5252, 5250, 5255, 4676]. Reported [3152]. reporting [3851]. represent [4320]. repression [3820]. Reproducing [1400, 2137, 4774]. Reproduction [4793]. Reproductive [1067, 1413]. Requiem [1957]. Required [5366]. Rescuing [2990, 4006]. Research [3600, 4285, 3601, 40, 1088, 2142, 2147, 3603, 2065, 835, 1035, 1029, 1034, 2897, 3604, 3599, 3657, 3672, 3191, 2874, 4172, 4606, 4178, 3632, 4506]. Researchers [4290, 3430, 3394]. reshuffling [3562]. residing [4610]. Resistance [2884, 1353, 5383, 4257, 4479, 4361, 1244, 3928]. Resistant [2874, 4022, 3670, 4679]. Resisting [2884]. Resolving [4554, 2145]. resource [4611]. Respond [1448]. Response [681, 1251]. Ressovsky [899]. Rest [2504]. Restoration [2368, 2361, 1868, 1872]. Restriction [2900]. restrictions [4254]. Result [115]. Results [1025, 5157]. Résumé [5312]. Resurrect [296]. Resurrection [4708]. retained [4603]. Rethinking [835, 4827]. Retina [584]. Retire [1998]. Retractable [5223]. Retrying [1082]. Return [1332, 2311, 2889, 3866, 2891]. Revamps [183]. Revealed [4598, 847, 4944]. Revealing [1629]. Reveals [3640]. Revelations [2211]. Revenge [4960]. Reversed [5293, 5300]. Review [1374, 1392, 1543, 1081, 2285, 1373, 1299, 172, 1542, 2371, 2286]. Reviews [4594, 1270, 1336, 1430, 1467, 1468, 1504, 1506, 1581, 1651, 1652, 1693, 1694, 1732, 1733, 2002, 1771, 1809, 1848, 1927, 1964, 2040, 2041, 2081, 2120, 2160, 2199, 2247, 2730, 2502, 2797, 2865, 2564, 2416, 2456, 2457, 2503, 2664, 2667, 2729, 3466, 2968, 3019, 2911, 2920, 2967, 3018, 3131, 3183, 3237, 3367, 3418, 3465, 3752, 3805, 3610, 3559, 3804, 3940, 4076, 3560, 3851, 3709, 3664, 3895, 4028, 3663, 3708, 3751, 3850, 3894, 3939, 4027, 4205, 4539, 4214, 4216, 4276, 4382, 4383, 4434, 4435, 4490, 4491, 4547, 4664, 4709, 4749]. Reviews [4941, 4940, 5038, 5147, 5186, 5220, 3988, 4278, 3287, 4980, 4979, 4493, 3609, 5347, 1081, 5109, 4215, 4277, 4160, 2563, 5310, 634, 670, 743, 777, 811, 705, 878, 910, 978, 1044, 1191, 1233, 1269, 944, 1012, 1117, 1155, 1429, 1335, 1580, 1614, 1770, 1808, 1847, 1884, 1926, 1963, 2001, 2080, 2119, 2159, 2198, 2325, 2326, 2246, 2863, 3515, 4075, 4588, 4875, 4876, 5260, 2370, 2793, 5074, 4663, 4161, 4438, 3178, 3368, 3288, 3072, 3125, 3230]. Revisited [3843, 3409, 152, 4471, 1494, 1488, 1493, 2766, 2941, 3416]. Revisiting [1862, 3417]. Revival [662, 5035, 4539]. Revolution [1001, 4861]. Revolutionary [3668, 579]. Reward [683]. rex [4042, 5209]. Rhine [2903]. Rib [5062]. Rice [4022]. Rich [848, 859, 1822, 2150, 860]. Richard [4687, 1593]. riches [3178]. Richter [4464]. Richter-scale [4464]. Riddle [2014, 502, 1043, 1032, 2306]. Ride [3982, 660, 655, 659, 3679]. Ridge [192]. Riding [828, 903, 4569]. Rifkin [3873]. Rift [4920]. Right [5100, 108, 3084, 1023, 3263, 4548]. Rights [2775, 4728, 2221, 2327, 1900, 3821, 4842]. Rigidity [586]. Ring [3280, 183]. Rings [1842]. Ringside [2341]. Rio [1008, 2432]. Rise [703, 5345, 1427, 1415, 1742, 3191, 5272, 4852]. rise-and-shine [3191]. Rising [2754, 2977, 3832, 3747, 3658, 3657, 3672, 4605, 4436]. Risk [3332, 3333, 4049, 3871, 4518, 1230, 4811, 1521, 3110, 3112, 4376, 4375, 3330, 2984]. Risks [4803, 152, 2012, 1215]. Rita [4732, 1314]. River [4566, 3114, 2903, 2478, 4554, 3119]. riveting [3896]. Rms [2699]. RNA [2874, 4908, 2887, 21]. Roach [1996]. Roaches [4137, 4140]. Road [2223, 2338, 2949, 97, 86, 96, 2820, 3941, 2800]. Roads [232]. Roam [550]. Robbery [2645, 2662]. Robert [2496, 2493]. Robo [4989]. Robo-Editor [4989]. Robot [3508]. Robotics [2646, 2636]. Robots [658, 5206, 669, 2107, 2104, 5345, 4896, 4897, 2581, 4009, 4942]. RoboTuna [1761]. Rock [3460]. Rocket [5375]. rocketeer [4072]. rocketing [3611]. Rockets [4930, 4932, 4527]. Rocking [3041]. Rocks [3041]. Rocky [850]. Rogers [3921]. Rogue [2395, 3823]. Role [1024]. Roles [993]. Rolf [4559]. ROM [2794, 3125, 3466, 3610]. Roman [2449, 2448, 369, 5346]. Romance [1918]. Romantic [3710]. Rome [3895, 966, 977]. ROMs [2668, 2287]. Ronald [3632]. Ronde [1939]. Roofs [5223]. Rookeries [3846]. Room [5112]. Rooms [3059, 3286]. Root [1686]. Roots [161, 1326, 3195]. Rosetta [3440, 3753]. Rotaxanes [2034]. Rotblat [2897]. Rothschild [3095]. Rough [4690]. ’round [4384, 3762, 5306, 4213, 5298]. roundup [3779]. Route [4520, 3737]. Routing [2395]. Roving [43]. Rowing [3116, 3120]. Ru [189]. Rub [611]. Rubbed [2888]. rubber [5267]. rubble [5149]. Rube [4339]. Ruckus [5148]. Rugged [622, 633]. Rugworld [32]. Rule [4105, 4489, 4775]. Rules [4521, 2949, 3430]. Run [1128, 5290, 4982, 1627, 1998]. Running [3931, 4414, 1253]. Runs [2478, 1574]. Rush [448]. Rushing [4966]. Russia [3481]. Russian [1356, 1362]. Rx [513].

S [4777, 4180, 4190, 1520, 4188]. Saboteurs [1623]. Safe [4005, 5197, 82, 470, 3128]. Safeguarding [3478]. Safely [3168, 198, 3512, 3166]. safer [4476]. Safety [3134, 3831]. Saffron [3680, 3679]. Sahara [5235]. Saharan [170]. Said [284]. sail [4464]. Sailing [4690]. Sails [3841, 4929]. Sakharov [4976]. Sale [1163, 1111, 1960, 3491, 4042]. salesmen [3662]. Salivating [3680]. Salmon [4871, 2389, 4869, 4870]. Salt [2403, 2406, 2995, 4531]. salt-tolerant [4531]. salvo [4244]. Sam [5211, 5212]. Same [3026, 259]. Samurai [2566]. Sanctifying [5074]. Sand [4107, 3934]. Sandbox [4657, 4658, 4467, 1715]. Sands [3273, 3274, 3802, 4696, 4697, 3803]. sandwich [4326, 4336]. sanguine [4875]. Sanity [1937]. Sardines [4213, 4212]. Satan [3687]. Satellite [3594, 4616]. Satellites [2621, 2614, 4322, 4934, 5255, 5254, 4527]. Satisfaction [758]. Satisfied [3023]. Saturn [4764]. Saucer [3781, 3779]. Sausage [1665]. savage [3072]. Savagery [2056]. Savants [4820]. Save [1577, 4871, 3696, 4869, 4870, 4476, 4350]. Saving [4360, 4168]. Saw [1261, 2271]. Say [4043, 3389, 3612]. Scale [3404, 4464, 3397]. Scandal [1566]. Scandinavia [5379]. Scanning [2038, 2029, 3776]. Scarlet [4985]. Scars [3958, 2680]. Scary [2986]. Scavenging [1178]. Scene [4250]. Scent [3722, 5284, 3727]. Scents [2599, 3711, 5207]. schedule [4585]. Scheme [2944]. Schizophrenia [2933, 188, 2941]. Schlock [5033]. Schooling [1079]. Schools [1256]. Schrödinger [5224, 3799, 3798]. Schwartz [4410]. Science [3224, 411, 387, 485, 522, 592, 463, 499, 573, 640, 1146, 1182, 1225, 1123, 1162, 1200, 1327, 1365, 1390, 1420, 1459, 1496, 1534, 1571, 1606, 1642, 1685, 1306, 1343, 1379, 1398, 1438, 1475, 1514, 1550, 1587, 1622, 1661, 2150, 1762, 1800, 1838, 1875, 1915, 1955, 1992, 2031, 2071, 2111, 1741, 1779, 1817, 1855, 1892, 1935, 1972, 2010, 2050, 2089, 2127, 2168, 2522, 2526, 2210, 2294, 2337, 2381, 2422, 2468, 2514, 2586, 2689, 2798, 2874, 3302, 2933, 3140, 3383, 2983, 3034, 3433, 3083, 3193, 3247, 3769, 3823, 3999]. Science [3575, 3679, 3624, 3722, 3531, 3865, 4042, 3911, 3812, 3553, 4287, 4294, 4350, 4453, 4504, 4172, 4127, 4554, 4227, 4401, 4840, 4895, 5275, 110, 2722, 3, 277, 313, 107, 140, 536, 782, 1202, 1072, 1088, 949, 983, 1199, 1180, 1176, 1179, 1520, 1641, 1586, 1356, 1362, 2149, 2070, 1914, 1854, 1816, 2467, 2293, 2209, 3159, 3167, 3862, 4732, 4627, 4731, 4189, 2788, 4648, 3988, 5288, 687, 3008, 3007, 1099, 3411, 5032, 159, 23, 90, 124, 194, 265, 626, 521, 697, 835, 661, 1224, 970, 936, 1181, 1161, 428, 2724]. Science [3287, 1272, 1013, 1083, 1735, 301, 373, 177, 804, 484, 558, 871, 1036, 1145, 1109, 2186, 4316, 5138, 1292, 1570, 1605, 1533, 1874, 2167, 2560, 2379, 2575, 2316, 2321, 2555, 2559, 4126, 4178, 5247, 53, 1544, 479, 3062, 4743, 2448, 2848, 352, 815, 649, 915, 1018, 1513, 1621, 1397, 1553, 1527, 1679, 1682, 1934, 3413, 3406, 3412, 2732, 212, 386, 39, 70, 639, 498, 675, 572, 883, 1122, 1277, 1445, 1989, 2088, 1891, 1985, 1988, 2744, 2719, 5375, 1045, 1676, 635, 263, 795, 2020, 3273, 5335, 1561, 3702, 1907, 3403, 3092, 3581, 1790, 841, 1269, 3808, 2110]. Science [2680, 2421, 3436, 4845, 4235, 4679, 5120, 56, 2216, 2295, 154, 1217, 3172, 201, 181, 1049, 847, 1342, 2057, 1740, 1778, 2009, 2048, 2126, 3956, 3789, 4703, 996, 709, 747, 1305, 1437, 1549, 1971, 3525, 1357, 362, 606, 1474, 462, 398, 2139, 2908, 2813, 2362, 231, 4973, 340, 410, 736, 591, 770, 446, 1004, 903, 1326, 1458, 1419, 1495, 1684, 1364, 1954, 1799, 1837, 1761, 1991, 2030, 2930, 2878, 3554, 4123, 5166, 895, 2519, 1101, 671, 553, 246, 1659, 2980, 1928, 1452, 1366, 2256, 2499, 2495, 2336, 4724, 1530, 2806, 3072, 3047, 3920, 3805, 4076, 4512, 4712, 4205, 5085]. science [1084, 4785, 4215, 2733, 4710, 3180, 3526, 3159, 4122, 4841]. Science-in-fiction [2798]. Sciences [2342, 4438]. Scientific [3492, 1965, 2555, 2559, 2067, 2188, 2194, 1981, 4741, 2002, 1014, 3370, 2866, 4076, 4053, 4776, 3517, 4750, 1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 2207, 2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623]. Scientific [4675, 4719, 4838, 4889, 4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317, 2863, 3515, 4075, 4284]. Scientist [1232, 1334, 1391, 1466, 1579, 1650, 1731, 1807, 1883, 1962, 2039, 2118, 2197, 2790, 2860, 2562, 2369, 3121, 3512, 3462, 2906, 3015, 3284, 3173, 2964, 3606, 4024, 3556, 3662, 4072, 4536, 4584, 4378, 4487, 4431, 4462, 4936, 1024, 2038, 1231, 2789, 2859, 3227, 3234, 3129, 3283, 3511, 3014, 2963, 3181, 3461, 3068, 3363, 2916, 3414, 3661, 3555, 4023, 3800, 3890, 3846, 3704, 3605, 3986, 4071, 3935, 3747, 4210, 4545, 4332, 4271, 4706, 4707, 4661, 4157, 4330, 4272, 4377, 4592, 4747, 4430, 4486, 5218, 5036, 5145, 5107, 5308, 4977, 5072, 4873, 5259]. Scientist [5184, 4872, 4935, 2368, 1806, 1730, 2561, 2283, 1333, 2117, 1961, 1465, 1337, 1011, 1649, 1882, 810, 2454, 207, 270, 346, 943, 1578, 1080, 2196, 877, 1298, 2248, 2302, 33, 65, 135, 1154, 271, 2284, 2455, 2917, 3705, 3847, 3801, 3891, 3936, 3748, 4211, 4273, 4331, 3364]. Scientists [1507, 1431, 2287, 1278, 5217, 4200, 1235, 3863, 3820, 4028, 5229, 4617]. Scissors [4318]. scoop [2984]. scores [4852]. Scoring [3590]. Scots [4985]. Scottish [3289]. Scotty [1476]. Scratch [629, 2476]. Scrawny [4240]. Screen [5170, 4223]. Screening [4800, 3340, 1994]. Screws [736]. Screwy [607]. Scribble [4440]. Scrolls [361]. scrutiny [2750, 3245, 4184, 4191]. Scuba [4550]. Sculpting [5061, 3666]. sculpture [3122]. Sculptures [3123, 3130]. Scum [4272]. scurvy [4985]. Sea [2731, 4598, 253, 3407, 4608, 1095, 4743, 3984, 864, 4308, 4609, 5377, 2813, 2817, 4617, 3853, 4617, 361]. Seabed [4145]. Seaborg [2296]. Seafloor [3624, 3790, 3789, 3528, 3232]. Search [1751, 4698, 4699, 4203, 3928, 2105, 4476]. Searching [3057, 5386, 4758, 3286, 3791, 3647, 4757, 3643, 3058]. Seas [3658, 3657, 3672, 4605]. Season [1860, 4554]. Seawater [4531, 4537]. Second [140, 613, 5071, 955, 446, 706, 4504, 4227, 3401, 3853, 4235]. second-biggest [4504]. Second-Generation [5071, 446, 3401]. Secret [2691, 4019, 772, 3888, 5085]. Secrets [4539, 2384, 2767, 2481, 5332, 2411, 3935, 5107, 2222, 3640, 4645, 4507]. Section [1242, 1702]. Secure [2141, 2143]. Security [4650, 4654, 4651, 4656, 4657, 3688]. See [233, 12, 2218, 1152, 4240, 146, 4354, 3612]. See-Through [1152]. Seed [260, 3667]. Seedlings [2959, 2963]. Seeing [5096, 2954, 2218, 3454, 1790, 1791, 943, 934, 3942, 3001, 318, 2262, 838, 2695, 4362]. Seek [2598]. Seeking [538, 2322, 2321, 4098, 4821, 3746, 3745, 5119, 2316]. seeks [4559]. Seem [442]. Seen [5333]. Seer [2271]. sees [3267]. Seismology [3063, 3068]. Selection [1610, 764]. Self [3433, 436, 2217, 1400, 2137, 4774, 1982, 285, 2630, 2640, 762]. Self-Assembling [2630, 2640]. Self-Loathing [2217]. Self-Organized [436]. Self-organizing [3433]. Self-Replicating [1982]. Self-Reproducing [1400, 2137, 4774]. Sell [2837, 3922]. Seller [3778]. Selling [3732, 1363, 1362, 631, 1356]. sellout [3779]. Sells [268]. Semi [1460]. Semi-Success [1460]. Semiconducting [4112]. Semiconductor [1073, 4109, 2899, 2901, 3156, 3164, 2277, 2282]. Semiconductors [60, 1177]. Semiready [1227]. Semper [5087]. Sending [4916]. Senescence [538].

Senile [3151]. Sensation [5003]. Sense [3588, 3965, 3340, 2302, 3051]. sensibility [3711]. Sensible [1075]. Sensing [4377]. sensitive [4786]. Sensor [487, 3634]. Sensors [3225, 3233]. Sensory [1949]. Sentries [1623]. Separate [3682]. Sequence [1629, 3999, 4506]. series [3175]. serious [4327]. Seriously [3673]. Services [724, 3639]. Set [118, 737, 449, 3863]. Set-To [118]. SETI [4831]. Sets [543]. Setting [4295, 2440]. Seven [2566, 5371, 1112]. Seven-League [1112]. Sewage [2776, 2745]. Sex [2565, 2690, 2927, 3485, 2317, 501, 2129, 3773, 1140, 2042, 2256, 2911, 4785]. Sexism [1084]. Sexual [2688]. Sexuality [1754]. Sexually [472, 4783]. Shading [4733]. shadows [3286]. shakes [5051]. Shaking [1588, 1876, 1247]. Shaky [4287, 4294, 4300]. Shannon [16]. Shape [4370, 167, 156, 166, 3289]. Shapiro [1520]. Share [2411, 4813]. Sharing [756, 61, 3690]. Shark [184]. Sharks [4614, 1626, 3453]. Sharper [753]. Shear [888]. Shedding [3279, 3285]. Sheep [1613, 1602, 3720]. Sheer [4336, 4337]. Sheet [63, 3346, 808]. Shell [1369]. Shelley [3267]. Shepherd [1613, 1602]. Sherry [4296]. shine [3191]. Shining [3198, 1537]. Ships [3983, 4690, 3896]. Shirt [1497]. Shock [2115, 551, 447, 4052]. Shock-Wave [4052]. Shocked [1369]. Shocking [3911, 218]. shoelaces [3228]. Shoot [3922]. Shooting [1941, 4026]. Shoots [3958]. shoplifting [3758]. Shopping [2998]. Shores [767, 768, 763]. Short [829, 1010, 840, 429]. Short-Order [429]. Short-Term [1010]. shortage [2983]. Shot [1613, 1602]. Should [1355, 3339, 4896, 4897, 4758]. Shout [3822]. Show [469, 2972, 4138]. Showdown [4052]. shows [3071, 4042]. shreds [4906]. Shrimp [4424, 2386, 356]. Shrink [4039]. Shrinking [1715]. Shrubbery [3431]. shrunk [1402]. Shuffle [4708, 3563]. Shutting [4691]. Sicán [1869]. Sick [195, 1095, 2051]. Sickness [1173, 3051]. Side [5264, 4560, 2862, 1910]. sidewalk [2517]. SIDS [2527, 2538]. Sierpinski [5185]. Sieve [2601, 3803]. Sifting [3803, 3802]. Sight [5205, 4196, 2733]. Sighting [1231]. Sign [2824, 2829]. signal [4801]. Signaling [1674]. Signals [916, 2561, 2558, 3405]. Significance [2347]. Signing [560]. Silence [3429, 2600]. Silent [1627]. Silicon [4107, 661, 936, 4139, 2407, 584, 1829, 3712, 446, 834, 833, 2751, 4137, 1460, 826]. Silicone [5197]. Silk [4219, 4206, 2800]. Silks [929]. Silly [1860]. Silver [1148]. Silversides [501]. Simon [2670, 4649]. Simpler [3982]. Simply [2350]. simulate [3621]. Simulating [2964, 4704, 1636]. Singers [5216]. Singing [1175]. Single [3401, 4472, 4785, 1030]. Single-Atom [4472]. Single-sex [4785]. Sinister [3723, 2943]. Site [4860]. Six [4367, 2563]. Size [3278, 3283, 1208, 1213, 4773, 4172, 4178]. Sizing [3999, 4746]. Skeptic [715]. Skeptically [2487]. skeptics [4007]. Sketch [4249]. Skewing [5202]. Skies [3747, 1426, 141, 2872]. Skin [5025, 303, 3214, 3220, 5026]. Skinside [806]. Skip [4982]. Skipper [1688]. Skullduggery [186]. Sky [3556, 3159, 3167, 4417, 3883, 3904, 2276, 2187, 2186, 359, 5376, 2678, 4235]. skyrocketing [4320]. Skyscrapers [4083, 4070]. Slartibartfast [2896]. Slash [2708]. Slate [3448]. Slave [4919, 5292, 5297]. Slave-Making [5292, 5297]. Slaying [4047]. Sleep [15, 3191]. Sleepers [1405]. Slide [4729]. slightly [4042]. Slime [4645]. slippery [2909]. slot [3294]. sloth [3282]. Slow [107]. Slowly [5206]. Slushballs [3865]. Small [3364, 2859, 2854, 3264, 3775, 181, 677, 451, 3676, 5369, 3801, 3639]. Smaller [3834, 822, 4195]. Smalley [1593]. Smart [694, 693, 5019, 2998, 3269, 4690, 4821, 2744, 2246, 3059, 1166, 691]. smarter [3875]. Smashing [4731]. Smattering [3038]. Smell [2720, 3253]. Smelling [77]. Smile [3097]. Smile-Aid [3097]. smiles [4385]. Smog [2483, 3305]. Smoke [4046, 3714, 3715]. Smoking [2674, 4803, 4778, 4804]. smuggler [2577]. Smuggling [2885, 2898]. Snail [148]. Snake [1294]. Snakes [45]. Snap [3489]. snapping [3467]. Snaps [63, 1479]. SNAREs [1479]. Sneaky [2429]. Snow [4351, 2302, 3566]. Snowballs [3866]. Snowflake [3001]. snub [4637]. SNUs [1718]. Soaking [3579]. Social [2724, 1832, 2726, 2719, 4244, 2874]. Socializing [2213]. Society [3553, 2317, 3567]. Sociobiology [4438]. Sociologist [4296]. Sociology [1679, 1682]. Sock [3306]. Soft [1329, 2712]. Software [3115, 2892, 5367, 2069, 2068, 4746, 1215, 2618, 1688, 2577, 3047, 3875, 4254, 2066]. Softwars [1880]. SOHO [3640]. Soil [3387, 3865, 3290]. Soiled [767, 768, 763]. Sokal [4233, 4244]. Sol [749]. Solar [4758, 3851, 155, 2767, 268, 2648, 4757, 4759, 4756]. Solar-Neutrino [155]. Solid [1983, 965]. Solids [1649, 1638]. Solitude [4996]. Solo [508]. Solutions [1418, 1411, 1417]. Solve [607]. Solving [2064]. Some [2870, 2134, 2838, 5366, 2769, 3639, 3528, 1464, 4757, 4786]. Something [3719, 1746]. sometimes [4793]. Somewhere [4775]. song [3612]. Songbirds [1000]. Sonic [1857]. Sonoluminescence [2273]. Sophie [832]. sorely [4611]. Sort [2773]. Sorting [4747, 632]. Soul [1370]. Sound [3504, 5188, 3511, 664, 4055, 1553, 4045, 2383, 1056, 1707, 3311, 4590, 4384, 2273]. Sounding [3413, 3412, 3406]. Sounds [2907, 2916, 3986, 1226, 2906]. Soup [2594, 4509]. South [2933, 1796, 3022]. Southern [3460]. Soviet [2583, 290, 295, 635]. Sowing [3093]. Soyuz [5124]. Space [615, 4931, 2786, 3267, 2874, 3560, 4890, 3266, 3159, 3167, 4893, 5368, 4923, 1025, 230, 226, 229, 466, 1181, 742, 3886, 4925, 116, 3217, 3222, 4862, 4527, 4532, 4939, 4160, 5030, 3982, 3870, 5124, 3866, 435, 1421, 5129, 3305, 4896, 5301, 5302, 4897, 2791, 3869, 730, 2792, 3662, 3865, 4453, 4776, 4627, 4603, 3754, 4775, 4724, 3247, 3885, 4845, 3135, 4965]. Space-Age [5302]. Space-time [4862]. Spacecraft [5089, 2629, 2623]. Spaceflight [4927]. spacetime [4774]. Spanish [2929]. Spark [4428]. Sparrows [3227, 3234]. Spawning [168]. Speaking [5275]. Spec [808]. Special [2410, 3671, 3813, 4085, 4261, 4263, 4319, 5355, 4840, 5175, 5012, 5248, 3764, 4065, 3642, 4651, 4475, 4861, 3792, 5252, 1242, 5250, 5255, 1702, 3525, 4123, 2097]. Species [2684, 2293, 1171, 1408, 785, 3096, 2738]. Specious [3368, 785]. Specter [3500]. Spectrum [4329]. Speech [3317, 3316, 4824]. Speed [4053, 469, 5248, 2616, 3976, 1829, 5373, 3977, 2195, 2188, 2194, 2790, 4206, 2625]. Spell [2947, 1979]. Sperm [4955]. Spider [929]. Spiders [4702, 4703]. Spies [487, 4735]. Spills [4695, 2583]. Spin [1535, 1864, 5137, 1023, 750]. Spinal [3485, 5214]. Spinning [1806, 1795, 3385, 1820]. spins [3175]. spirals [3122, 4464]. Splicing [3679]. Split [4471]. splitters [3897]. Spock [234]. spoke [4462]. sponge [3575]. Sponging [2386]. Spooling [4592]. spoonful [2528]. Spot [12, 1938, 1403]. Spotlight [4639, 4637]. Spots [3086, 997, 996, 3679]. Spotter [630]. Spotting [2339, 3678]. Spread [4329]. Spread-Spectrum [4329]. Spreads [3324]. Springs [2864]. Spying [3781]. Square [3848, 3849]. Squares [3557, 5309, 3558]. Squaring [3017, 3848]. Squeeze [5064, 3863, 4212]. Squeezing [3540]. squid [3468]. SQUIDs [2022]. Squirt [219]. Sri [4613]. St [3123, 3130]. St. [3114, 3119]. Stab [4118]. Stability [400, 775]. Stable [928, 89]. Stadium [5223]. Stage [2295]. Stained [1041, 3126]. Stalkers [950]. Stalking [4558]. stall [4908]. stamps [4587]. Stand [303, 4730, 154, 4017, 3492]. Stand-Ins [303]. Standard [2426, 4092, 2440]. standards [4230]. Standing [2090]. Stanford [4906]. Stanley [623, 172]. Stanton [4462]. Star [2968, 3019, 3571, 4412, 3142, 1011, 998, 4217, 4218, 3998, 819, 2009, 2880, 1593, 3999, 2825, 4205, 5097, 4215]. Star-Hopping [3142]. Stardust [321]. Stars [3531, 5328, 440, 2716, 3206, 863, 4927, 4179, 1170, 4915, 4933, 4756, 2404, 4223, 651, 2341, 3467, 4207, 3433, 3911]. Start [536, 3719, 1555, 2921, 4627]. Starting [3448, 629, 1194]. starts [3311]. Starving [4635]. State [2115, 793, 635, 4476]. static [2562]. station [3560, 4893, 3135, 4965]. Statistical [4449, 368, 2347]. Statistics [1659, 4029]. Stay [3026, 4948]. STDs [4783, 4784]. Steal [5171, 5190]. Steel [94, 2864, 4336]. Steering [3261]. Stellar [3270, 1206, 4771]. Stem [5126, 827, 5016, 5018, 3677, 4398]. Step [550, 3264, 94, 451]. Steps [3530]. Stereogram [2222]. stereotypes [3401]. Sterilization [213]. Stew [2597]. Stickleback [1413]. Sticky [4012, 3941]. Still [3441, 4083, 3940, 1018, 4070, 2478, 1844, 3082, 4007, 1084, 2258]. Stinging [358]. Stirling [30]. stolen [4587]. Stone [1068, 3753]. Stones [43, 3756]. Stop [4209, 1542, 4222]. Stoppard [3851]. Stopping [3684, 3669]. Stops [3990]. Storage [1181, 44, 3624]. Store [4805]. Stored [3168, 3166]. stories [4941]. Storm [2268, 5007, 2130, 3288]. Story [483, 2436, 1910, 1149, 1839, 4391, 1371, 1360]. Straight [3980]. Strange [1751, 3733, 2881, 1552, 5272, 4852]. stranger [4274]. Stranglers [952]. Strategic [3819, 5255, 5254]. Strategies [1067, 3794, 1878, 4702, 4703, 4473, 3330, 4637]. Strategy [769, 3862, 776]. Stratospheric [618]. Strauss [4135]. Stream [3953, 4536]. Streams [4545]. Street [632, 4921, 2110]. Strength [214, 3015]. Streptococcal [617]. Stress [5058, 5321, 4128, 28, 1448, 4127, 4799]. Stretched [63]. Stretching [5267]. Strife [5201]. Strike [5200, 4770]. Strikes [2840]. String [1936, 4629]. Strings [4202, 2271]. strip [3849]. strips [4274]. Stroke [2992, 4806, 652, 3679]. Strokes [3684, 1183, 3438]. Strong [3841, 871, 4788]. Structure [927, 547]. Struggle [2836, 2385]. Stubbornly [1283]. student [3034]. Studies [2445, 2452, 3435, 2982, 4236]. studs [2874]. Study [1040]. studying [3556]. Stuff [5019, 3133, 3855, 3668, 4496, 4592, 2999, 2221, 4710]. Style [4961, 2973]. Sub [170]. Sub-Saharan [170]. Subatomic [3898]. Submarine [3528, 3851, 4606]. Submarines [2018]. Subsidies [3156, 3164]. Subspecies [4956]. Substitutes [4203]. Subterranean [5373]. Subtle [4377, 649]. Suburban [3529]. Subway [2078]. Success [1460, 2545]. successful [4205, 4215]. Suck [2434]. Sucker [1625]. Sudden [4028]. Suez [4881]. Sugar [413, 2129, 2528]. Suicide [2832, 3497, 3503, 3623, 3505, 3624]. suit [3575]. suitable [4531]. Sulfate [1788]. Sulfates [2691]. Sum [4495, 4494]. Sumer [2003]. Summer [1042]. Summit [1934]. Sun [5184, 4741, 262, 1537, 830, 52, 3640, 2339, 299]. Sundial [695, 704]. Sunless [3232]. Sunlight [3214, 3220]. Suns [5052, 1281, 4756]. Sunspots [198]. Super [3147, 4104, 1586, 2048, 1857, 1861]. Super-transistor [4104]. Super-ultrasound [3147]. Superantigens [962]. Supercar [5006]. Supercomputer [508]. Supercomputers [3548, 3621]. Superconducting [469]. Superconductivity [3275, 3654, 1723, 1728, 3655, 3271]. Superconductor [607]. Superconductors [1353, 5383, 256, 2634, 2644, 1742]. Supercontinent [961]. Supercritical [3443, 3811]. Superfluids [203]. Superhack [1974]. Supernovae [4862, 4915, 4172, 4178]. Superplume [2280]. Supersoft [4915]. superstrength [3443]. Superstrings [791]. Supersymmetry [3193]. Superwire [2486]. Supply [3354, 3341, 394, 5053, 4634]. Support [1928, 4845]. supporting [4758]. Suppressing [4895]. Surefire [5312]. Surface [398, 399]. Surfaces [2038, 2029, 1412, 653, 4274]. surfers [4058]. Surgeons [2151, 2593]. Surgery [3224, 3219, 621, 5102]. Surgical [1525]. Surprisingly [3741]. Surreal [1366]. survey [4076]. Surveying [516, 518, 4862]. Survival [2234, 2149, 3862, 1763, 1080, 1069, 5166]. Survive [1363, 1356, 1362, 1604]. Survives [503, 2927]. Surviving [3019]. Survivors [5046]. Suspension [4084, 4067, 99]. Suspension-Feeding [99]. Sustainable [805, 1193, 2651, 204, 3696]. Sustaining [1531, 2109, 1530, 1527]. Swamps [3012]. Swap [3999]. Swapping [4146]. sweep [4476]. Sweet [2545, 3518, 2830]. Swift [622, 633]. Swifter [3875]. Swim [737]. Swimming [2314]. Swing [2300]. Swings [1462]. Switch [5086, 1023]. Switches [305]. Swords [2032, 2150]. Symbiosis [1175]. Symmetries [3075, 3608]. Symmetry [3022]. Symmetry-Breaking [3022]. Sympathy [837]. Sync [1578, 1569]. Synchronicity [4978]. Synchronization [1726]. Syndrome [3635, 1010, 4314, 4062]. Synesthesia [3965]. Synthetic [2278, 1982]. Synthetics [3587]. System [3851, 3451, 1286, 1382, 3691, 1599, 3796, 4757, 1596, 1601, 1597, 1594, 2552, 1598, 4578, 4586, 2779, 1595, 1603, 1056, 762, 4759, 2953]. Systematic [3100]. Systems [4758, 3650, 4328, 4756]. Szilard [3552].

T [1560, 3382]. T-Bone [3382]. T. [4042, 3316]. Table [4831, 3752]. Tables [2318, 2323]. Tacking [3341]. tackles [2755]. Tackling [4935, 4111, 3548]. Tacky [854]. Tactics [1763]. Tadpoles [1653]. tag [3173]. Tail [3555, 3386]. tailless [3089]. Tainted [149]. Take [3748, 4476, 1895, 3944, 4668, 4669, 4475, 2744]. Takes [154, 1613, 1602, 3201, 3368, 3230]. Taki [1956]. Taking [3302, 3707, 4018, 4071, 4872, 5007, 3845, 4346, 4307, 1230, 4494, 3267]. Tale [406, 148, 253, 11, 2979]. talent [2870]. Talents [4820]. Tales [2002, 3176, 3182, 2969]. Talk [2144, 3728, 2126]. Talking [194, 4824, 2297, 5177]. talks [2753]. Tall [2090, 2979]. Tallest [4086, 4068]. Talwar [3201]. Taming [3735, 4894]. Tamoxifen [4401]. Tangling [5037]. Tankers [1424]. tanks [3420]. Tap [1038]. Tape [4012, 2529]. Tapestry [2365]. Tapping [4939]. target [4058]. Targeted [1828]. Targets [3353]. Tarzan [550]. Task [3845]. Tasmanian [3383]. Taste [5060, 4850, 667, 5277]. Taxes [2746]. Taxoids [3171, 3177]. Tea [4752, 4751]. Teachers [320]. teaches [3290]. Teaching [1180, 1179, 1286, 1382, 320, 1176]. Teacup [4465]. Team [464]. tech [3820]. technical [4741]. Technician [2869]. Technique [926]. Technological [2607, 2609, 3761]. Technologies [2610, 2653, 4319, 4321, 4320, 2654, 2655]. Technology [2205, 2755, 2692, 2529, 2825, 2252, 2376, 2382, 2463, 2509, 2587, 2331, 3311, 3443, 3047, 2934, 3198, 3394, 2984, 3261, 3089, 4009, 3634, 3690, 3540, 3832, 3732, 3779, 3586, 3875, 3922, 4053, 4306, 4464, 4137, 4360, 4562, 4413, 4184, 4510, 4299, 4236, 4908, 3613, 1537, 2268, 2481, 2603, 2708, 4855, 3215, 2840, 2220, 2348, 3218, 3116, 3120, 4191, 763, 4109, 2899, 2901, 457, 5027, 1987, 2661, 4637, 2651, 2656, 1357, 1358, 4660, 362, 4370, 4371, 1419, 2307, 2851, 2857, 4735, 4248, 2891, 4689, 3611, 4616, 3807, 4722, 4099, 3147, 3765, 4968, 3764]. Tectonics [4467]. teenagers [3401]. Teens [2391, 4076]. Teeth [5210, 5352]. Telecommunications [4323]. Telecoms [3044]. Teleconomics [2948]. Telegraph [2742]. telescope [3198, 1025]. Telescopes [3662, 4758]. Television [3471, 3099, 3391, 4372, 4370, 4371, 3472, 4230]. Tell [4081, 2412, 2436, 561]. Teller [172]. Tells [132, 5100]. Telomerase [2955]. Telomere [686]. Telomeres [2955]. Temperature [1353, 5383, 2634, 2644, 3275, 3271]. Temperatures [2226]. Tempest [4736, 4465, 2813]. Temple [4373, 4374]. Ten [3407, 4608, 3052, 3053]. Tender [4132]. Tens [4778]. Tensaiji [25]. Tension [2698]. Term [1010]. Terrain [622, 633]. Terrestrial [4324]. Terror [558, 1796]. Terrorist [356, 4689]. Tesla [3709]. Tessellations [4937]. Test [109, 2709, 855, 70, 987, 2744, 3483, 5321, 4902, 2820, 3767]. Testing [3767, 4808, 2542, 3160, 4627, 1386, 4636, 5215, 4816, 3337, 1298, 4788, 4604]. Testosterone [2278]. Tests [2257, 4480, 1066, 1953, 1948, 1952, 3138, 3499, 2933, 4476, 3540]. Tethers [4925]. Tevatron [510]. Text} [4244]. Textures [927]. Thalidomide [2756]. Thaw [2760]. Their [4368, 4317, 4635, 3339, 4199, 2604, 1385, 3195, 4224, 4759, 469, 1507, 2287, 4634, 394, 4207, 2879, 1288]. Them [3212, 198, 4757, 2271]. Themselves [2144, 3656, 3927]. theology [4504]. Theorem [3040, 406, 1824]. Theoretical [2904, 2914, 1244]. Theories [3050, 4773, 4756]. Theory [586, 4202, 239, 503, 368, 2267, 432, 3302, 3767, 4613]. Therapeutic [1603]. Therapies [3350, 3351, 4892]. Therapy [2633, 3342, 3793, 337, 4478, 539, 3795, 3332, 3333, 3794, 3792, 3343, 580, 3796, 4802, 5021, 5023, 3797, 3353, 364, 652, 2588]. There [4164, 4165, 4830, 611, 3715, 4430, 4193, 2267, 142, 5344]. Theres [2711]. Thereza [3449]. Thermometer [4977]. Thermonuclear [964]. Thermophotovoltaics [4582, 4591]. These [4476, 2705, 2974, 3715, 4616]. Theseus [483]. Thin [4330, 265, 4779]. Thing [2219]. Things [1469, 5244, 4967, 167, 156, 166, 2668, 3175, 3419]. Think [18, 20, 1355, 3204, 3083]. thinkers [4244]. Thinking [836, 1252, 2714, 812, 5161, 525, 3957, 3368, 3521]. Third [2560, 2555, 2559, 255, 2018, 1648, 2707]. Thirsty [2983]. Thomas [4358, 2723, 2722]. Thornton [4086]. Thoroughbred [587]. Those [2834, 2564, 4476, 4610]. Thought [3889, 415, 3903, 2225, 4516, 5388]. Thoughts [2922]. Thousand [3447]. Threat [1877, 462, 2885, 2898, 2736, 4617]. Threatened [3918, 3536]. Three [2235, 1462, 4710, 1441, 4777, 4936, 4603]. Thriller [1747]. Thrive [2235, 3288]. Throat [5216]. throughout [4728, 4133]. Thrust [1877, 4408]. Thumbs [114]. Thunder [5111]. Thunderstorms [2170]. Thyself [3253]. Tick [4549]. Tidbits [517]. Tides [2024]. Tidings [2537]. Tiger [3386, 3211, 3383]. Tight [2704, 4213, 2790]. Tightens [2305]. Tile [2786, 2791]. Tiling [3849, 5146]. Till [5236]. Tim [4051]. Time [4204, 2148, 2147, 2344, 3726, 4684, 2972, 741, 1830, 301, 5293, 5300, 3217, 3222, 1283, 1524, 2267, 1785, 1680, 1691, 3869, 1482, 4867, 3191, 3660, 4862, 2142]. Time-Reversed [5293, 5300]. Time-Trip [1680, 1691]. Time-Trippers [1785]. Timekeeping [4335, 4327, 4334]. Times [1621, 4741]. Timing [2532, 1168, 3533, 784, 3000, 4971, 2521, 4795]. Timoféeff [899]. Timoféeff-Ressovsky [899]. Tiniest [854]. Tins [4213]. Tiny [3251, 2713]. Tipping [875]. tips [4813]. Tissue [5025, 5012, 5013, 5028, 5026]. Tissue-Engineered [5025, 5026]. TNTenderizer [4127]. Toads [1833]. toast [2862]. Tobacco [2402, 4462, 4460]. Tock [4549]. Today [4604, 4476, 4757]. Together [394, 870, 865, 869, 4772, 4775]. Toil [4159]. Toilet [3827]. tolerant [4531]. Tolerating [285]. Toll [2542, 4819]. Tolls [2209]. Tom [3851]. Tomography [1765, 327, 336]. Tomorrow [4563, 5315, 2433, 3031]. Tone [117]. Tongue [160, 4363, 4360]. Too [3262, 1278, 1781, 5168, 4793]. took [3526]. Tool [4684, 3610]. Toolmakers [1073]. Tools [442, 3743]. Toothpaste [5314]. Top [2867, 3888, 2387, 3929, 1978, 3521]. Top-Secret [3888]. Topic [2258]. Topological [517, 1540, 1532]. Topology [5037]. Tops [3520]. Tornadoes [2551]. Total [5262]. Touch [147, 3263, 4640]. Touchpad [4497]. touchpads [4498]. Tough [2999]. Tour [4759, 4403, 3707]. Tours [3706]. Tower [2849, 2848]. Town [3203]. Toxic [2173, 2421, 4208]. Toxin [1007]. Toxins [1752, 2475]. trace [2983]. Tracers [2988]. Track [4078, 4616]. Tracking [4009, 537, 687, 688, 4063]. Tracks [3977]. Trade [1669, 1670, 4260, 4919, 1672, 4459, 4521, 1863, 1572, 4510]. Trading [741]. Tradition [2425]. traditional [4756]. Traffic [521, 3973]. Tragedy [51, 1772, 5095]. Trail [72, 2183, 3468]. Trailing [5182, 5183]. Training [3163, 3165]. Trains [3978, 3977, 1108, 1103, 1107]. Tranquillity [4904]. Trans [431]. transactions [2966]. Transform [165]. Transformation [123, 133]. Transforming [2762]. transfusions [4875]. Transgenic [1027, 3549]. Transistor [4117, 1487, 4106, 4101, 4104]. Transistors [4102, 469]. Transit [2397]. Transition [300]. translators [3441]. transmission [4230]. Transmitted [472, 4783]. Transmitters [2678]. Transmutation [989]. Transparency [4574]. transplant [3247]. Transplanted [2490]. Transplants [3345, 5358]. Transportation [2622, 2624, 2623, 3967, 655, 659, 1103, 1107, 1906, 1913, 156, 166]. Transvaal [392]. Trapped [2171]. Trapping [4184, 896, 4191]. Trash [2297]. Trauma [187, 4781]. Travel [1830, 4927, 3974]. traveler [2002]. traveling [3662]. Treasures [3839]. Treaties [2192, 1891, 4510]. Treating [2490]. Treatment [2776, 2305, 799, 1525, 3779]. Treatments [3358, 3344, 3343, 3922, 4476]. Treaty [2575]. Treble [3208]. Trebuchet [2492, 2807]. Tree [5327, 2990, 1247, 2383]. Trees [4053, 251, 2695, 392, 4627]. Trek [829, 3019, 3722, 3571, 840]. Tremor [3735]. trench [3665]. Trends [3858, 3759, 3902, 3889, 3814, 3553, 691, 693, 1176, 1179, 1356, 1362, 2142, 2147, 1756, 1758, 2232, 2240, 3013, 3159, 3167, 5388, 226, 229, 365, 372, 439, 444, 583, 590, 897, 901, 1029, 1034, 2724, 963, 968, 2066, 2068, 2316, 2321, 2555, 2559, 3276, 3281, 3845, 5247, 763, 519, 513, 1527, 1679, 1682, 1868, 1872, 2025, 2027, 2408, 2413, 3406, 3412, 329, 338, 478, 481, 1065, 1070, 1488, 1493, 1635, 1640, 1985, 1988, 1792, 1797, 2445, 2452, 2719, 3510, 3745, 261, 684, 2360, 2366, 2904, 2914, 3604, 797, 802, 619, 624]. Trends [1449, 1456, 402, 408, 865, 869, 1291, 1383, 1948, 1952, 549, 556, 994, 1002, 1258, 1265, 930, 933, 1320, 3657, 3672, 655, 659, 1216, 1222, 1103, 1107, 1564, 1567, 1411, 1417, 1906, 1913, 2188, 2194, 2851, 2857, 2277, 2282, 3061, 3066, 3567, 4375, 826, 833, 1831, 1835, 86, 96, 156, 166, 1723, 1728, 2499, 3798, 4064, 2495, 1530]. Très [1608]. trial [1082]. trials [3083]. Triassic [3205]. Tribal [866, 2385]. Trick [160]. trickery [3942]. Tricks [4139, 3881, 4040]. Trickster [644]. tries [4175, 4438]. Trigger [4018]. Trip [622, 633, 1680, 1691, 3732]. Triple [4935, 1437]. Trippers [1785]. trips [5041]. triumph [4099]. Trojan [4265, 355, 921]. Tropical [3696]. Tropics [119]. Trouble [1586, 391, 1156, 4159, 2706, 2578]. Troubled [858]. True [483, 738, 4686, 1371, 1360, 5168, 3923, 3366]. Trusted [3650]. Truth [5228, 4081]. Truths [329, 338, 339]. Try [1533, 3537, 4290]. Trying [989]. Tsunami [5067]. Tsunamis [4377, 4378]. Tube [987]. Tubes [4105]. Tude [3132]. Tuesday [3464]. Tumbleweed [515]. Tumor [3325, 3338, 3393, 4892]. Tumors [4635, 4346, 1983, 4891, 4634]. Tune [4911]. Tuning [916]. Tunnel [4023, 3833, 4024]. tunnels [3832]. Turbulence [3548]. Turing [5032, 855, 2078]. Turkey [831]. Turkish [4281, 4282]. Turkle [4296]. Turmoil [4904]. Turn [3772, 213, 1842, 449, 783, 4590].

Turn-on [783]. Turning [2318, 3865, 4236, 2070, 2435, 2441, 2323, 1684]. turns [3122]. Turquoise [898]. Turtle [253, 5095]. Turtles [4616, 3627, 864]. Tuva [5216]. TVs [4224]. Tweaking [5015]. Tweezers [4318]. Twelve [3347, 3346]. Twinkle [4733]. Twist [4468, 3822, 1389, 5258, 3558]. Twists [1291, 1383]. Two [3762, 3560, 3944, 1074, 1410, 3987, 610, 3999, 4336, 4230]. Two-Way [3987]. Type [2264]. Tyrannosaurs [5210]. Tyrannosaurus [5209]. Tyranny [4111].

U [5134]. U.S [3978]. U.S. [2691, 3572, 3779, 4180, 4190, 4539, 4914, 4994, 4188, 1073, 3387, 3685, 4049, 3774, 3871, 3582, 4301, 4406, 5165, 5093, 5203, 4859, 213, 574, 1722, 3694, 4689]. U.S.A. [2242]. U.S.S.R. [684]. Ubiquitous [5185]. ugly [3247]. Uh [2826]. Ulcers [2957]. Ultimate [5363, 1999, 1990, 2530]. Ultrabright [4741]. ultrafast [396]. ultramassive [4771]. Ultrasonics [4152]. Ultrasound [3155, 3147]. Ultraviolet [262, 2019, 4040]. Ulysses [4148]. Umklapp [304]. Unabomber [3140, 4312, 4310, 4311]. unanswered [4772]. Unbearable [1896]. uncertain [3031]. Uncertainties [2607, 2609]. Uncertainty [1001, 126, 4449, 2156]. Uncommon [4820]. Unconscious [100]. Uncountable [2811]. Uncovering [3110, 3112]. Underground [1150, 5323, 1484, 4009]. Undersea [2693]. Understanding [992, 2359, 3547]. Understandings [3321, 3322]. Underwater [2038, 2029, 2954, 3403, 3210, 4543]. undiagnosed [4783]. Undressing [4244, 4245]. Unearthing [2096]. unease [3138]. Unequal [4856]. Unexpected [5335, 4845, 4724]. Unicorn [3390]. Unified [5336]. unify [4438]. Union [4524, 2583, 290, 295]. Unions [4510, 5058]. unique [4277]. Unite [4962]. Universal [3591, 338, 339, 1942, 329]. Universally [2714]. Universe [1208, 4773, 5281, 4774, 1562, 4772, 4775, 2098, 3823, 4042, 4453, 4554, 4864, 5259, 1213, 825, 27, 648, 503, 5291, 5099, 2137, 3744, 1561, 4560, 1529, 2099, 5337, 5344, 2097]. University [3621, 1849]. Unjamming [3973]. Unknowable [2458]. Unknown [2458, 3188]. unleashes [2932]. Unlikely [586, 1255, 918, 2918]. Unlimited [3543]. Unmasking [5065]. Unmeltable [3481]. Unmet [5246]. Unnatural [2173]. unpredictable [3562]. Unraveling [1667, 3935]. unravels [3228]. Unrequited [5313]. unrobed [2002]. Unscientific [1520]. Unsolved [252]. Unsound [4237]. Unusual [121]. Unveiled [316]. Unzipping [3077]. upcoming [4776, 4230]. updated [3709]. Uphill [1128]. Uplift [512]. upon [4941, 2267, 4408]. upright [4948]. Uranus [4765]. Urbaculture [3143]. Urban [3290, 3002, 4637]. URE3 [2014]. Urge [744]. Urinary [4812]. Urine [4239]. Use [1794, 3307, 289, 4230]. used [4042]. Using [2562]. utero [2823]. utilities [3732]. Utility [2783, 2785, 4610]. UV [2571].

V [3777, 2947, 3316, 899]. V-e-n-o-m [2947]. V1974 [2236, 2238, 4769]. Vaccine [2243, 5123, 562, 2682, 3499, 4285, 4562, 4290]. Vaccines [4808, 4483, 3213, 2930, 5136, 2933, 4476, 4954, 4908]. Vacuum [3408, 3414, 3415]. Valuable [4480, 4617]. value [4539, 3370, 4030]. Vampire [61]. Vanishing [4227, 3721, 1043, 1032, 1000, 2571]. Vapor [72]. Variable [52]. Variables [1231]. Various [5313]. Vasa [1637]. Vector [430]. vegetables [3999]. veggies [3369]. Vehicles [5142, 287, 293, 3969, 3971, 3450, 3970]. Velcro [3077]. Velikovsky [3371]. velocity [2520]. Venter [4506]. Venus [2002, 4761, 4969, 1870, 847, 1441, 316, 398, 399]. Verne [3703]. versa [3466]. Version [30]. Versus [3976, 2566, 3553, 3554, 5339]. Vertebrate [225, 2902, 3453]. Vertebrates [99]. Vertical [3134]. Vesicles [3006]. Vested [938]. vests [3670]. Vexing [953]. via [5249, 5250]. vice [3466]. Victims [4781, 1662]. Victors [3486]. Victory [1481]. Video [1995, 931, 2529]. video-tape [2529]. Videotape [1768, 3773, 1760]. View [3088, 3044, 3195, 3309, 3145, 3258, 3392, 3441, 4058, 3639, 3592, 3545, 4015, 4366, 4564, 4129, 4421, 4254, 4312, 4512, 4914, 2151, 5112, 4859, 4310, 4738, 4418, 4641, 4252, 4142, 1749, 2028, 2025, 2027, 4197, 1425, 1152, 2341, 4693, 4967, 2699, 3801, 3927, 3695, 3785, 3881, 3737, 3838, 4420]. Viewing [2694]. Views [5072, 4315, 4316]. Viking [4201]. Village [726, 1009, 97, 86, 96, 733]. Vinci [2572, 3610]. vine [3586]. ving [3514]. Violated [1898]. Violence [4553, 4638, 4797, 4556, 792]. Violent [1352, 4771]. Viral [1522, 2988, 4480, 5170]. Viral-Load [4480]. Virtual [3895, 4453, 2179, 2619, 2613, 3974, 2882, 5215]. Virtually [3974]. virtuoso [2870]. Virtuous [1422]. Virulence [1414]. Virus [2933, 5182, 5183, 2177, 1628, 552, 2774, 4476]. Viruses [2983, 1422, 4020, 4149, 4866, 2715]. Vishnu [1898]. Vision [3664, 3208, 428, 1078, 1843, 5285, 5294, 612]. Visionary [3703]. Visions [1908, 1907, 3833]. Visit [1973]. Visual [1135]. Visualizing [1217, 1218, 1865, 4973, 4972]. Vital [3603, 2312, 3013]. vitamin [3561]. Vitriol [4566]. Vitro [1609]. Vive [323]. Vivian [4777]. Voice [1483, 1262, 4328]. Void [3252, 790, 3415]. Volcanic [334, 332, 922]. Volcano [2757, 3531, 1097]. Volcanoes [893]. Vole [1489]. volume [2562]. Volunteering [3911]. Vortices [953]. Voted [4685]. Voting [2352, 4679]. vs [715, 4896, 4897]. Vu [979, 5231].

W [4777]. W. [4777]. W.C. [2864]. WABOT [163]. Wages [1863, 2178]. Waiter [4430]. Waiting [3066, 3067, 5006, 3441, 3061]. Waking [3192]. Walk [550, 4921, 4982, 4551]. Walking [2705, 3225, 3233]. Wall [110, 42, 1991, 3248, 632, 4921, 2110]. Wallaroo [1572]. wallop [4770]. Walter [2226]. Walton [4492]. Wanamaker [4540]. Wand [1643]. Wanda [4777]. Wandering [3089, 4616, 4759]. Wanna [5124]. Want [3868, 4072]. Wanted [1717]. Wants [5269, 2933, 3034]. War [1048, 1756, 1758, 1759, 3768, 397, 574, 1518, 2550, 5120, 2493, 2384, 4265, 3444, 2496]. Warfare [866, 5103, 3665]. Warheads [1558]. Warm [3656, 2150]. Warming [923, 698, 261, 1782, 2258, 2816, 4627]. Warned [4412]. Warning [74, 2740]. warns [4410]. Warp [1482]. Warrior [1211, 5020]. Wars [1745, 2128, 3919, 2370, 1996, 2858, 2851, 2857, 2825, 4205, 5097, 4215]. Wartime [4745]. Was [3911, 1943, 3895, 2267, 2266, 2918]. Washington [2983, 1407]. Wasn’t [843]. Wasp [4019]. Wasps [1288, 1385]. Waste [2173, 973, 4145, 2353, 3168, 3481, 3624, 3166]. Wasteful [3601]. Wasteland [3124, 3118]. Watch [5010, 179, 2180, 3472]. Watchdogging [3247]. Watchers [5010, 1502, 3556]. Watching [3472, 4210, 1011, 998, 4633]. Water [2933, 5354, 1128, 1634, 4005, 4704, 3931, 130, 956, 4963, 4603, 4353, 143, 4936]. Water-Pollinated [1634]. Waterfall [1348, 2473]. Watering [1478]. Waters [4939]. Watershed [1465, 1451]. Watery [3754]. Watts [376]. Wave [5318, 935, 930, 933, 4052, 2346, 3162, 3154]. Wavelet [432]. Wavelets [1499]. Waves [1350, 1864, 3129, 678, 551, 5159, 678, 3121]. Waving [3614]. Way [483, 4923, 2947, 5290, 3746, 3745, 3948, 3987, 2668, 3258, 3383, 4299, 4643, 1316]. Wayne [3400]. Ways [2786, 2061, 2791, 2243, 4058, 3664]. weaken [4510]. Wealth [4617]. Weapon [4019, 4063, 3695]. Weapons [4018, 3500, 3171, 3177, 3621, 4689]. Wear [2712, 5020]. Wearable [4829]. Wearing [4141]. Weather [4880, 4879, 2126, 4607, 3247, 4554]. Web [4058, 3881, 3838, 3688, 4421, 5071, 3287, 4859, 4738, 4418, 4419, 3647, 4050, 5101, 3644]. Webs [929]. Webspinner [4051]. Websurfing [3648]. Weeds [654]. weeping [2974]. weeps [3857]. weevil [3999]. weigh [4617]. Weight [3037, 4633]. Weightlessness [4571, 4579, 5277]. Weird [1021]. weirdly [3997]. Welding [3492, 3494]. Welfare [3401, 2030]. Well [5020, 1492, 1780, 4710]. Well-Dressed [5020]. Went [3969]. We’re [3673, 1662, 2705, 4277, 1207]. West [4052]. Western [1060, 362, 4427]. Wetlands [4406, 1750]. Whale [4633]. Whales [3114, 3119, 4845]. Whammy [1437]. Whatchamacallit [4181]. Wheat [506]. Wheel [4140, 4137]. Wheellock [4156]. Wheels [1166, 3838]. Whence [4845]. Where [4456, 957, 4028, 4189, 4632, 3638, 3627, 3493, 4439, 4105, 4734, 4409, 2826, 5329, 3612, 4710, 3093, 3514]. Which [3574, 4774]. While [231, 4224]. Whistle [3919, 3920]. Whistle-Blower [3919, 3920]. White [468, 5153]. Whiteout [4999]. Who [3664, 3498, 3223, 3229, 76, 1549, 4671, 5269, 1207, 4896, 4897, 1192, 2235, 823, 2226]. Whom [2209, 4819]. Whoof [886]. Whorf [886]. Wide [4051, 4859, 3736, 2300, 4616]. wide-ranging [4616]. Widget [2349]. Widgets [3736]. Wiener [2833]. Wild [2183, 2883, 4558, 28]. Wildfire [4163, 4162]. Wildlife [1031, 3364]. Wilds [3583]. Will [5022, 3464, 4476, 4475, 1577, 2107, 2104, 1604, 3974, 5020, 4637, 3829]. William [4083]. Williams [4062]. wills [3082]. Wilson [4438]. Wimps [4960, 1312]. Wind [335, 4993, 2789, 2784, 4023, 3493, 3962, 2270, 2790, 3383, 4024]. Window [5285, 5294]. Windy [3146]. wing [3368]. Winging [3107]. Wings [4827, 370]. Wink [5056]. Winner [5046]. Winners [1959, 2174, 2834]. Winning [769, 776, 4476, 4122]. Wire [78, 3731, 1835, 1836, 3592, 1831]. Wired [5092, 2498, 2499, 4114, 2495]. Wireless [4319, 4321, 4320, 4324, 5256, 4328, 2612]. wires [3732]. Wireworld [32]. Wisdom [1588]. Withheld [641]. Within [3046, 1124, 2836, 2301]. Without [3768, 3648, 800, 2755, 3173, 4450, 4413, 4442, 810, 4710, 709, 4874, 1539]. Witness [154]. Woes [343]. Wolf [2889, 2488]. Wolves [4401, 2738]. Women [2981, 4512, 4777, 4803, 4133, 5237, 1083, 213, 1483, 2025, 2027, 4783, 4786, 3339, 945, 4797, 4809, 4792, 2769, 4798, 2737, 4789, 4779, 4813, 4799, 4787]. Won [1756, 1758, 1759]. Wonderful [3203]. Wonders [3611, 1021, 3370, 4030, 3074, 2970, 2913, 3021, 3127, 3180, 3421, 3232, 3468, 3290, 3897, 3853, 3807, 3666, 3711, 3942, 3562, 3754, 4327, 4588, 4387, 4667, 4266, 4541, 4437, 4716, 4280, 4207, 4218, 4334, 4494, 4163, 4595, 5149, 5304, 5350, 4883, 5265, 4944, 5077, 4983, 5041, 5371, 2796]. Won’t [5172]. Wood [3158]. Word [755]. Words [2161, 557, 549, 556, 3151]. Wore [5153]. Work [3793, 3876, 4893, 195, 2192, 5172, 728, 739, 3207, 2668, 3425, 3394, 3545, 3526, 4442, 4209, 4122, 4222]. Workaholic [2037]. Worked [2575]. workforce [2746]. Working [3521, 3235, 3294, 3128, 3472, 3408, 3946, 3670, 3566, 3758, 3616, 3857, 3901, 4034, 3811, 4081, 3715, 4498, 4154, 4543, 4590, 4390, 4442, 4209, 4269, 5045, 4948, 3373, 3471, 4166, 4167, 5189, 5044, 5114, 5268, 5354, 4670, 4550, 3565, 3714, 5314, 3414, 4389, 5080, 4497, 1136, 4338, 4339, 1139, 3293, 4283, 3424, 3077, 3810, 3900, 3856, 4441, 3240, 4884, 4885, 4033, 3945, 3615, 4597, 4754, 4755, 3992, 4080, 3187, 3520, 3757, 2908, 2910, 3669, 4947, 3134, 5223, 4221, 4222, 5152, 4986, 3186, 3374]. Workplace [4851, 4845, 2660]. Works [4104, 4654, 4656, 4657, 1143, 3223, 3229, 4284, 3664]. World [3224, 3219, 367, 442, 4086, 848, 859, 860, 2443, 5072, 2621, 2614, 3307, 4568, 4728, 4133, 2560, 2391, 2555, 3736, 4134, 4828, 4642, 2887, 2494, 3273, 3274, 3001, 4565, 4068, 1092, 3721, 294, 1755, 2778, 4611, 2771, 5298, 5306, 822, 5369, 4161, 2681, 4476, 4350, 4384, 3282, 5041, 2987, 1048, 4051, 4661, 4859, 2559, 3422, 3421, 255, 2018, 1648, 5091, 4427, 2707]. World-Record [2443]. Worlds [4758, 440, 4759]. Worldwide [3725]. Worm [2909, 4886, 2919]. Worms [4355]. Worries [2588, 4233]. Worry [2388, 3820]. Worst [4460, 1445]. Worth [1230, 3447, 2077, 5098, 3720]. Would [3498, 2964, 4710, 2226]. Wrap [906, 1491, 1503]. Wrapping [3808]. wreck [4604]. wreck-diver [4604]. wreckers [2579]. Writ [181]. Write [2195, 2188, 2194]. Writing [2710, 3126]. Wrong [151, 543, 683, 843, 3258, 3383, 3419, 4028]. wrong-way [3383]. www.batmobile.car [3837].

X [4953, 4741, 2979, 2787, 4055, 476, 4915, 3086, 502, 1403, 3256, 3393, 1675]. X-ray [1675, 4055, 476, 502, 4915]. X-rays [2979, 2787, 4741]. Xenotransplantation [3840]. XML [5071].

Y2K [4413, 4642, 4867]. Yale [449]. yang [3178]. Yard [1163]. Year [4486, 1242, 1702, 4641]. Year-End [1242, 1702]. Years [385, 175, 2, 38, 69, 106, 139, 176, 211, 245, 276, 312, 351, 384, 427, 461, 497, 535, 570, 605, 638, 674, 708, 746, 781, 814, 846, 882, 914, 948, 982, 1017, 1047, 1087, 1121, 1159, 1160, 1197, 1198, 1240, 1241, 1276, 1303, 1340, 1395, 1435, 1472, 1511, 1547, 1585, 1619, 1657, 1700, 1304, 1341, 1378, 1396, 1436, 1473, 1512, 1548, 1620, 1658, 1701, 1738, 1776, 1814, 1852, 1889, 1932, 1969, 2007, 2046, 2086, 2124, 2165, 1739, 1777, 1815, 1853, 1890, 1933, 1970, 2008, 2047, 2087, 2125]. Years [2166, 2207, 2292, 2334, 2378, 2419, 2515, 2573, 2675, 2739, 2808, 2255, 2208, 2291, 2335, 2420, 2465, 2466, 2518, 2574, 2679, 2743, 2812, 2877, 2928, 2978, 3029, 3081, 3139, 3190, 3243, 3297, 3378, 3428, 3476, 2873, 2929, 2979, 3524, 3570, 3619, 3675, 3718, 3763, 3818, 3861, 3907, 3995, 4038, 3950, 4120, 4173, 4228, 4288, 4344, 4395, 4447, 4505, 4555, 4623, 4675, 4719, 4499, 4674, 4720, 4837, 4838, 4889, 4952, 4992, 5048, 5083, 5118, 5155, 5193, 5227, 5273, 5317, 4705, 4996, 3052, 3053, 571, 4121, 4174]. years [4229, 4289, 4345, 4396, 4448, 4809, 4099]. Yellow [131, 4535, 4544]. Yesterday [986, 2433]. yield [4773, 4611]. yin [3178]. Yoichiro [2271]. Yonder [2883]. Young [2866, 4076, 2716, 1351, 4782, 841, 1269, 2863, 3515, 3517, 4075, 4750, 1281, 3473, 4908, 4779, 3744]. you’re [1858, 2264, 4689, 1978]. Yours [4416]. Yourself [586, 4784, 5239, 4983]. Yucca [3168, 3166]. Yves [3914].

Z [4515, 4526]. Z-Pinch [4515]. Zap [111]. Zapper [4415]. Zebra [3247]. Zen [1192, 3687]. Zeno [2145]. Zeppelin [5296, 5303]. Zero [2966, 1404, 4064]. Zero-based [2966]. Zero-Point [4064]. Zeroes [719]. Zimbabwe [4021]. Zinc [1354]. Zip [1364, 3078, 4137]. Zippers [548]. Zone [2167, 4613, 5200, 3997, 5351]. Zoo [2993]. Zoos [5139, 5138, 3896]. Zzzzzzz [5042].

References

Anonymous:1990:La

[1] Anonymous. Letters. Scientific American, 262(1):10, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

      URL http://www.nature.com/scientificamerican/journal/v262/
      n1/pdf/scientificamerican0190-10.pdf.

Anonymous:1990:YAa

[2] Anonymous. 50 and 100 years ago. Scientific American, 262(1):12, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

      (electronic). URL http://www.nature.com/scientificamerican/
      journal/v262/n1/pdf/scientificamerican0190-12.pdf.

Beardsley:1990:SCA

[3] Tim Beardsley. Science and the citizen: Aftershocks. Scientific American, 262(1):14, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
14a.pdf.

B:1990:LC

[4] T. M. B. Loud and clear. Scientific American, 262(1):14–15, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-14b.pdf.

B:1990:N

[5] T. M. B. Nanofuture. Scientific American, 262(1):15–16, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-15b.pdf.

Ross:1990:NP

[6] Philip E. Ross. Nit picker. Scientific American, 262(1):15–16, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-15a.pdf.

Horgan:1990:QC

[7] John Horgan. Quasicrystal clear. Scientific American, 262(1):16–17, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-16.pdf.

Ruthen:1990:BA

[8] Russell Ruthen. Bottled antimatter. Scientific American, 262(1):17, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-17a.pdf.

Powell:1990:RB

[9] Corey S. Powell. The redshift blues. Scientific American, 262(1):17–18, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-17b.pdf.

B:1990:P

[10] T. M. B. On to the past. Scientific American, 262(1):18, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-18a.pdf.

H:1990:GT

[11] J. H. A grave tale. Scientific American, 262(1):18–20, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-18b.pdf.

H:1990:SSS

[12] J. H. See spot see blue. Scientific American, 262(1):20–21, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-20.pdf.

Searle:1990:BMCa

[13] John R. Searle. Is the brain’s mind a computer program? Scientific American, 262(1):20–25, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1990:NE

[14] John Rennie. Nervous excitement. Scientific American, 262(1):21, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-21a.pdf.

H:1990:SG

[15] J. H. Sleep of the guilty. Scientific American, 262(1):21–22, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-21b.pdf.

Horgan:1990:CS

[16] John Horgan. Claude E. Shannon. Scientific American, 262(1):22–22B, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-22.pdf.

Anonymous:1990:AID

[17] Anonymous. Artificial intelligence: a debate. Scientific American, 262(1):25, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
25.pdf.

Churchland:1990:CMTa

[18] Paul M. Churchland and Patricia Smith Churchland. Could a machine think? Scientific American, 262(1):26–31, January 1990. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Searle:1990:BMCb

[19] John R. Searle. Is the brain’s mind a computer program? Scientific American, 262(1):26–31, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
26.pdf.

Churchland:1990:CMTb

[20] Paul M. Churchland and Patricia Smith Churchland. Could a machine think? Scientific American, 262(1):32–37, January 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-32.pdf.

Weintraub:1990:ARD

[21] Harold M. Weintraub. Antisense RNA and DNA. Scientific American, 262(1):40–46, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
40.pdf.

Broecker:1990:WDG

[22] Wallace S. Broecker and George H. Denton. What drives glacial cycles? Scientific American, 262(1):48–56, January 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-48.pdf.

Corcoran:1990:SBC

[23] Elizabeth Corcoran. Science and business: Computer babble. Scientific American, 262(1):100–102, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
100.pdf.

Erickson:1990:RD

[24] Deborah Erickson. Rational drugs. Scientific American, 262(1):102–104, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-102.pdf.

Dambrot:1990:T

[25] Stuart M. Dambrot. Tensaiji. Scientific American, 262(1):104, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-104a.pdf.

Wallich:1990:AEa

[26] Paul Wallich and Elizabeth Corcoran. The analytical economist. Scientific American, 262(1):104–105, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
104b.pdf.

Hegstrom:1990:HU

[27] Roger A. Hegstrom and Dilip K. Kondepudi. The handedness of the Universe. Scientific American, 262(1):108–115, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n1/pdf/scientificamerican0190-108.pdf.

Sapolsky:1990:SW

[28] Robert M. Sapolsky. Stress in the wild. Scientific American, 262(1):116–123, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
116.pdf.

Bollinger:1990:M

[29] John J. Bollinger and David J. Wineland. Microplasmas. Scientific American, 262(1):124–130, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
124.pdf.

Anonymous:1990:BVS

[30] Anonymous. A backyard version of a Stirling engine can be built with common materials. Scientific American, 262(1):130–??, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gulkis:1990:CBE

[31] Samuel Gulkis, Philip M. Lubin, Stephan S. Meyer, and Robert F. Silverberg. The Cosmic Background Explorer. Scientific American, 262(1):132–139, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
132.pdf.

Anonymous:1990:CAP

[32] Anonymous. The cellular automata programs that create wireworld, rugworld and other diversions. Scientific American, 262(1):136–??, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Walker:1990:ASa

[33] Jearl Walker. The amateur scientist. Scientific American, 262 (1):140–144, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
140.pdf.

Dewdney:1990:CR

[34] A. K. Dewdney. Computer recreations. Scientific American, 262 (1):146–149, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n1/pdf/scientificamerican0190-
146.pdf.

Morrison:1990:Ba

[35] Philip Morrison. Books. Scientific American, 262(1):150–153, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-150.pdf.

Benedick:1990:E

[36] Richard Elliot Benedick. Essay. Scientific American, 262(1):154, January 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n1/pdf/scientificamerican0190-154.pdf.

Anonymous:1990:Lb

[37] Anonymous. Letters. Scientific American, 262(2):8, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n2/pdf/scientificamerican0290-8.pdf.

Anonymous:1990:YAb

[38] Anonymous. 50 and 100 years ago. Scientific American, 262(2):10, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-10.pdf.

Horgan:1990:SCR

[39] John Horgan. Science and the citizen: R.I.P. Blackbird. Scientific American, 262(2):12–16, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
12.pdf.

Beardsley:1990:AR

[40] Tim Beardsley. Aborted research. Scientific American, 262(2):16, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-16a.pdf.

H:1990:CH

[41] J. H. Chub hubbub. Scientific American, 262(2):16–18, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n2/pdf/scientificamerican0290-16b.pdf.

Powell:1990:AW

[42] Corey S. Powell. Up against the wall. Scientific American, 262(2):18–19, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-18.pdf.

Appenzeller:1990:RS

[43] Tim Appenzeller. Roving stones. Scientific American, 262(2):19–20, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-19.pdf.

Ross:1990:CS

[44] Philip E. Ross. Cold storage. Scientific American, 262(2):20–22, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-20.pdf.

M:1990:SG

[45] T. B. M. Snakes in the grass. Scientific American, 262(2):22, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-22a.pdf.

Rennie:1990:KC

[46] John Rennie. Kissing cousins. Scientific American, 262(2):22–23, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-22b.pdf.

H:1990:WDD

[47] J. H. When Dad drinks. Scientific American, 262(2):23, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n2/pdf/scientificamerican0290-23a.pdf.

Wallich:1990:PN

[48] Paul Wallich. Playing the numbers. Scientific American, 262(2):23–24, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-23b.pdf.

Horgan:1990:IB

[49] John Horgan. Indecent burial. Scientific American, 262(2):24, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-24D.pdf.

Kinoshita:1990:OC

[50] June Kinoshita. The oncogene connection. Scientific American, 262(2):24, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
24.pdf.

Melzack:1990:TNP

[51] Ronald Melzack. The tragedy of needless pain. Scientific American, 262(2):27–33, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
27.pdf.

Foukal:1990:VS

[52] Peter V. Foukal. The variable Sun. Scientific American, 262(2):34– 41 (Intl. ed. 26–??), February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
34.pdf.

Goldberger:1990:SPC

[53] Ary L. Goldberger, David R. Rigney, and Bruce J. West. Science in pictures: Chaos and fractals in human physiology. Scientific American, 262(2):42–49, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
42.pdf.

Goldberger:1990:CFH

[54] Ary L. Goldberger, David R. Rigney, and Bruce J. West. Chaos and fractals in human physiology. Scientific American, 262(2):42–?? (Intl. ed. 34–??), February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Govindjee:1990:HPM

[55] Govindjee and William J. Coleman. How plants make oxygen. Scientific American, 262(2):50–58, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
50.pdf.

Myers:1990:SBP

[56] Frederick S. Myers and E. Corcoran. Science and business: Progress by degrees. Scientific American, 262(2):60–66, February 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n2/pdf/scientificamerican0290-60.pdf.

Rennie:1990:AB

[57] John Rennie. Antibody bonanza. Scientific American, 262(2):62–65, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-62.pdf.

C:1990:BF

[58] E. C. Banking futures. Scientific American, 262(2):65–66, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n2/pdf/scientificamerican0290-65.pdf.

Corcoran:1990:AEa

[59] Elizabeth Corcoran. The analytical economist. Scientific American, 262(2):66, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
66.pdf.

Brodsky:1990:PGA

[60] Marc H. Brodsky. Progress in gallium arsenide semiconductors. Scientific American, 262(2):68–75 (Intl. ed. 56–??), February 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n2/pdf/scientificamerican0290-68.pdf.

Wilkinson:1990:FSV

[61] Gerald S. Wilkinson. Food sharing in vampire bats. Scientific American, 262(2):76–82, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
76.pdf.

Yanin:1990:AN

[62] Valentin L. Yanin. The archaeology of Novgorod. Scientific American, 262(2):84–91, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
84.pdf.

Anonymous:1990:WPS

[63] Anonymous. When a polymer sheet is stretched, it may “neck” long before it snaps. Scientific American, 262(2):86–??, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Arthur:1990:PFE

[64] W. Brian Arthur. Positive feedbacks in the economy. Scientific American, 262(2):92–99, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
92.pdf.

Walker:1990:ASb

[65] Jearl Walker. The amateur scientist. Scientific American, 262 (2):100–105, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n2/pdf/scientificamerican0290-
100.pdf.

Morrison:1990:Bb

[66] Philip Morrison. Books. Scientific American, 262(2):106–111, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-106.pdf.

Malcom:1990:E

[67] Shirley M. Malcom. Essay. Scientific American, 262(2):112, February 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n2/pdf/scientificamerican0290-112.pdf.

Anonymous:1990:Lc

[68] Anonymous. Letters. Scientific American, 262(3):12, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-12.pdf.

Anonymous:1990:YAc

[69] Anonymous. 50 and 100 years ago. Scientific American, 262(3):16, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-16.pdf.

Horgan:1990:SCT

[70] John Horgan. Science and the citizen: Test negative. Scientific American, 262(3):18–22, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
18.pdf.

Wallich:1990:PH

[71] Paul Wallich. Polar heat. Scientific American, 262(3):22–24, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-22.pdf.

H:1990:VT

[72] J. H. Vapor trail. Scientific American, 262(3):24–26, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-24.pdf.

Powell:1990:AP

[73] Corey S. Powell. Astronomy potpourri. Scientific American, 262(3):26– 27, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-26.pdf.

H:1990:BW

[74] J. H. Broadcast warning. Scientific American, 262(3):27–28, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-27.pdf.

H:1990:BP

[75] J. H. Bee police. Scientific American, 262(3):28–30, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-28.pdf.

Rennie:1990:WD

[76] John Rennie. Who’s the dealer? Scientific American, 262(3):30–32, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-30.pdf.

Ross:1990:SB

[77] Philip E. Ross. Smelling better. Scientific American, 262(3):32, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-32a.pdf.

Corcoran:1990:W

[78] Elizabeth Corcoran. Down to the wire. Scientific American, 262(3):32– 34, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-32b.pdf.

H:1990:HH

[79] J. H. Heli-hopper. Scientific American, 262(3):34–35, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-34.pdf.

Greiner:1990:NRa

[80] Walter Greiner and Aurel Sandulescu. New radioactivities. Scientific American, 262(3):34–??, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1990:F

[81] Tim Beardsley. Fallout. Scientific American, 262(3):35–36, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-35.pdf.

Rusting:1990:SP

[82] Ricki Rusting. Safe passage? Scientific American, 262(3):36, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-36a.pdf.

Beardsley:1990:PE

[83] Tim Beardsley. Punctuated equilibrium. Scientific American, 262(3):36– 38, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-36b.pdf.

Burns:1990:OM

[84] Jack O. Burns, Nebojsa Duric, G. Jeffrey Taylor, and Stewart W. Johnson. Observatories on the Moon. Scientific American, 262(3): 42–49 (Intl. ed. 18–??), March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
42.pdf.

Smith:1990:I

[85] Kendall A. Smith. Interleukin-2. Scientific American, 262(3):50–57, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-50.pdf.

Wright:1990:TCRa

[86] Karen Wright. Trends in communications: The road to the global village. Scientific American, 262(3):57–??, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Greiner:1990:NRN

[87] Walter Greiner and Aurel Sandulescu. New radioactivities (nucleus models). Scientific American, 262(3):58–63, 66, 67, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Greiner:1990:NRb

[88] Walter Greiner and Aurel Sandulescu. New radioactivities. Scientific American, 262(3):58–67, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
58.pdf.

Johnston:1990:ESC

[89] Arch C. Johnston and Lisa R. Kanter. Earthquakes in stable continental crust. Scientific American, 262(3):68–75 (Intl. ed. 42–??), March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-68.pdf.

Corcoran:1990:SBL

[90] Elizabeth Corcoran. Science and business: Likely litigation. Scientific American, 262(3):76–77, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
76.pdf.

C:1990:JPF

[91] E. C. Not just a pretty face. Scientific American, 262(3):77–78, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-77.pdf.

Guterl:1990:OIL

[92] Fred Guterl. One if by land. Scientific American, 262(3):78, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-78a.pdf.

C:1990:FB

[93] E. C. Fischer Black. Scientific American, 262(3):78–79, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-78b.pdf.

Pope:1990:OSS

[94] Gregory T. Pope. One-step steel. Scientific American, 262(3):79–80, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-79.pdf.

Wallich:1990:AEb

[95] Paul Wallich and Elizabeth Corcoran. The analytical economist. Scientific American, 262(3):80–81, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
80.pdf.

Wright:1990:TCRb

[96] Karen Wright. Trends in communications: The road to the global village. Scientific American, 262(3):83–94, March 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-83.pdf.

Wright:1990:RGV

[97] Karen Wright. The road to the global village. Scientific American, 262 (3):83–?? (Intl. ed. 57–??), March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1990:LIL

[98] Anonymous. Lunar infants, lotteries and meteorites expose the dangers of math abuse. Scientific American, 262(3):90–??, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sanderson:1990:SFV

[99] S. Laurie Sanderson and Richard Wassersug. Suspension-feeding vertebrates. Scientific American, 262(3):96–101, March 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-96.pdf.

Weiss:1990:UMF

[100] Joseph Weiss. Unconscious mental functioning. Scientific American, 262(3):103–109, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
103.pdf.

Gamkrelidze:1990:EHI

[101] Thomas V. Gamkrelidze and V. V. Ivanov. The early history of Indo-European languages. Scientific American, 262(3):110–116, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-110.pdf.

Dewdney:1990:MRa

[102] A. K. Dewdney. Mathematical recreations. Scientific American, 262(3):118–121, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n3/pdf/scientificamerican0390-
118.pdf.

Morrison:1990:Bc

[103] Philip Morrison. Books. Scientific American, 262(3):122–127, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n3/pdf/scientificamerican0390-122.pdf.

Tucker:1990:E

[104] Richard F. Tucker. Essay. Scientific American, 262(3):128, March 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n3/pdf/scientificamerican0390-128.pdf.

Anonymous:1990:Ld

[105] Anonymous. Letters. Scientific American, 262(4):10, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-10.pdf.

Anonymous:1990:YAd

[106] Anonymous. 50 and 100 years ago. Scientific American, 262(4):12, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-12.pdf.

Beardsley:1990:SCSa

[107] Tim Beardsley. Science and the citizen: Slow boat to Mars. Scientific American, 262(4):14, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
14a.pdf.

Horgan:1990:RL

[108] John Horgan. Right to lie? Scientific American, 262(4):14–18, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-14b.pdf.

B:1990:AT

[109] T. M. B. Acid test. Scientific American, 262(4):18–20, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-18.pdf.

B:1990:SBW

[110] T. M. B. Science behind the wall. Scientific American, 262(4):20, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-20a.pdf.

Wallich:1990:Z

[111] Paul Wallich. Zap! Scientific American, 262(4):20–22, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-20b.pdf.

Powell:1990:GF

[112] Corey S. Powell. Gone in a flash. Scientific American, 262(4):22, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-22.pdf.

Ross:1990:AAE

[113] Philip E. Ross. All about eve. Scientific American, 262(4):22–24, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-22D.pdf.

B:1990:GT

[114] T. M. B. Green thumbs. Scientific American, 262(4):24, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-24a.pdf.

Kinoshita:1990:NRF

[115] June Kinoshita. Net result: Folded protein. Scientific American, 262(4):24–26, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
24b.pdf.

H:1990:SI

[116] J. H. Space invaders. Scientific American, 262(4):26–30, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-26.pdf.

Rennie:1990:TR

[117] John Rennie. Tone of relief. Scientific American, 262(4):30, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-30a.pdf.

Horgan:1990:MS

[118] John Horgan. Mandelbrot set-to. Scientific American, 262(4):30–34, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-30b.pdf.

Repetto:1990:DT

[119] Robert Repetto. Deforestation in the tropics. Scientific American, 262(4):36–42, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
36.pdf.

Townes:1990:WHC

[120] Charles H. Townes and Reinhard Genzel. What is happening at the center of our galaxy? Scientific American, 262(4):46–55 (Intl. ed. 26–??), April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-46.pdf.

Mullis:1990:UOP

[121] Kary B. Mullis. The unusual origin of the polymerase chain reaction. Scientific American, 262(4):56–65, April 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-56.pdf.

Grieve:1990:ICE

[122] Richard A. F. Grieve. Impact cratering on the Earth. Scientific American, 262(4):66–73 (Intl. ed. 44–??), April 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-66.pdf.

Yelled:1990:TKK

[123] John E. Yelled. The transformation of the Kalahari !Kung. Scientific American, 262(4):72–??, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1990:SBM

[124] Elizabeth Corcoran. Science and business: Milliken & Co. Scientific American, 262(4):74–76, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
74.pdf.

Erickson:1990:MM

[125] Deborah Erickson. Of mice and men. Scientific American, 262(4):76–77, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-76.pdf.

E:1990:AU

[126] D. E. Atmosphere of uncertainty. Scientific American, 262(4):77–78, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-77.pdf.

Stokes:1990:B

[127] Henry Scott Stokes. Breakup? Scientific American, 262(4):78, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-78a.pdf.

E:1990:C

[128] D. E. Cashcowosaurus. Scientific American, 262(4):78–79, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-78b.pdf.

Corcoran:1990:AEb

[129] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 262(4):79–80, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
79.pdf.

Golay:1990:ALW

[130] Michael W. Golay and Neil E. Todreas. Advanced light-water reactors. Scientific American, 262(4):82–89 (Intl. ed. 58–??), April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-82.pdf.

Anonymous:1990:HCC

[131] Anonymous. A homemade copper chloride laser emits powerful burst of green and yellow light. Scientific American, 262(4):88–??, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Barlow:1990:WBT

[132] Robert B. Barlow, Jr. What the brain tells the eye. Scientific American, 262(4):90–95 (Intl. ed. 66–??), April 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-90.pdf.

Yellen:1990:TKK

[133] John E. Yellen. The transformation of the Kalahari !Kung. Scientific American, 262(4):96–105, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
96.pdf.

Vandiver:1990:AG

[134] Pamela B. Vandiver. Ancient glazes. Scientific American, 262(4):106– 113, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-106.pdf.

Walker:1990:ASc

[135] Jearl Walker. The amateur scientist. Scientific American, 262(4):114– 117, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-114.pdf.

Morgan:1990:B

[136] M. Granger Morgan. Books. Scientific American, 262(4):118–123, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n4/pdf/scientificamerican0490-118.pdf.

Gore:1990:E

[137] Al Gore. Essay. Scientific American, 262(4):124, April 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n4/pdf/scientificamerican0490-124.pdf.

Anonymous:1990:Le

[138] Anonymous. Letters. Scientific American, 262(5):13, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-13.pdf.

Anonymous:1990:YAe

[139] Anonymous. 50 and 100 years ago. Scientific American, 262(5):16, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-16.pdf.

Beardsley:1990:SCSb

[140] Tim Beardsley. Science and the citizen: Second opinion. Scientific American, 262(5):19–20, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n5/pdf/scientificamerican0590-
19.pdf.

Horgan:1990:OSa

[141] John Horgan. Open skies. Scientific American, 262(5):20–24, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-20.pdf.

Powell:1990:CGT

[142] Corey S. Powell. Can’t get there from here. Scientific American, 262(5): 24–25, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-24.pdf.

Wallich:1990:MW

[143] Paul Wallich. Murky water. Scientific American, 262(5):25–26, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-25.pdf.

Ruthen:1990:GG

[144] Russell Ruthen. Go with the grain. Scientific American, 262(5):26– 30, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-26.pdf.

Epstein:1990:CB

[145] Ari W. Epstein. Cool breeze. Scientific American, 262(5):30, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-30.pdf.

H:1990:NYS

[146] J. H. Now you see it …. Scientific American, 262(5):30–32, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-30b.pdf.

B:1990:RTL

[147] T. M. B. Reach out and touch …. Scientific American, 262(5):32, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-32a.pdf.

B:1990:ST

[148] T. M. B. Snail tale. Scientific American, 262(5):32–34, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-32b.pdf.

B:1990:TFM

[149] T. M. B. Tainted feed, mad cows. Scientific American, 262(5):34, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-34a.pdf.

H:1990:TGD

[150] J. H. Through a glass, darkly. Scientific American, 262(5):34–36, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-34b.pdf.

H:1990:AMG

[151] J. H. An antidrug message gets its facts wrong. Scientific American, 262 (5):36, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-36a.pdf.

B:1990:RRR

[152] T. M. B. Radiation risks revisited. Scientific American, 262(5):36– 40, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-36.pdf.

Horgan:1990:FR

[153] John Horgan. Free radical. Scientific American, 262(5):40–44, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-40.pdf.

Neufeld:1990:WST

[154] Peter J. Neufeld and Neville Colman. When science takes the witness stand. Scientific American, 262(5):46–53, May 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-46.pdf.

Bahcall:1990:SNP

[155] John N. Bahcall. The solar-neutrino problem. Scientific American, 262 (5):54–61 (Intl. ed. 26–??), May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n5/pdf/scientificamerican0590-
54.pdf.

Wright:1990:TTSa

[156] Kared Wright. Trends in transportation: The shape of things to go. Scientific American, 262(5):58–??, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rosenberg:1990:AIC

[157] Steven A. Rosenberg. Adoptive immunotherapy for cancer. Scientific American, 262(5):62–69 (Intl. ed. 34–??), May 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-62.pdf.

Wellnhofer:1990:A

[158] Peter Wellnhofer. Archaeopteryx. Scientific American, 262(5):70–77, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-70.pdf.

Corcoran:1990:SBB

[159] Elizabeth Corcoran. Science and business: Bright future. Scientific American, 262(5):78–80, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n5/pdf/scientificamerican0590-
78.pdf.

Erickson:1990:TT

[160] Deborah Erickson. Trick of the tongue. Scientific American, 262(5):80– 81, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-80.pdf.

E:1990:PR

[161] D. E. Putting down roots. Scientific American, 262(5):81–84, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-81.pdf.

Rennie:1990:CC

[162] John Rennie. Cancer catcher. Scientific American, 262(5):84, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-84a.pdf.

Myers:1990:PIA

[163] Frederick S. Myers. Play it again, WABOT. Scientific American, 262(5): 84–85, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-84b.pdf.

Corcoran:1990:AEc

[164] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 262(5):86–88, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n5/pdf/scientificamerican0590-
86.pdf.

Anonymous:1990:HTF

[165] Anonymous. How to transform flights of fancy into fractal flora or fauna. Scientific American, 262(5):90–??, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wright:1990:TTSb

[166] Karen Wright. Trends in transportation: The shape of things to go. Scientific American, 262(5):92–101, May 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-92.pdf.

Wright:1990:STG

[167] Karen Wright. The shape of things to go. Scientific American, 262(5): 92–??, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leggett:1990:SC

[168] William C. Leggett and Kenneth T. Frank. The spawning of the capelin. Scientific American, 262(5):102–107, May 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-102.pdf.

Peterson:1990:HPP

[169] Carl W. Peterson. High-performance parachutes. Scientific American, 262(5):108–116, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n5/pdf/scientificamerican0590-
108.pdf.

Caldwell:1990:HFS

[170] John C. Caldwell and Pat Caldwell. High fertility in Sub-Saharan Africa. Scientific American, 262(5):118–125, May 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-118.pdf.

Dewdney:1990:MRb

[171] A. K. Dewdney. Mathematical recreations. Scientific American, 262(5):126–129, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n5/pdf/scientificamerican0590-
126.pdf.

McMillan:1990:BRE

[172] Priscilla Johnson McMillan. Books: Review of Edward Teller: Giant of the Golden Age of Physics, by Stanley A. Blumberg and Louis G. Panos. Scientific American, 262(5):130–135, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-130.pdf.

Mahoney:1990:E

[173] Richard J. Mahoney. Essay. Scientific American, 262(5):136, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n5/pdf/scientificamerican0590-136.pdf.

Anonymous:1990:Lf

[174] Anonymous. Letters. Scientific American, 262(6):10, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-10.pdf.

Anonymous:1990:YAEa

[175] Anonymous. 50 and 100 years ago: Errata. Scientific American, 262(6): 10, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-10a.pdf.

Anonymous:1990:YAf

[176] Anonymous. 50 and 100 years ago. Scientific American, 262(6):12, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-12.pdf.

Erickson:1990:SCB

[177] Deborah Erickson. Science and the citizen: Blood feud. Scientific American, 262(6):17–18, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n6/pdf/scientificamerican0690-
17.pdf.

Beardsley:1990:HL

[178] Tim Beardsley. Hubble’s legacy. Scientific American, 262(6):18–22, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-18.pdf.

Horgan:1990:DW

[179] John Horgan. Death watch. Scientific American, 262(6):22–25, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-22.pdf.

Wallich:1990:PR

[180] Paul Wallich. Peary redux. Scientific American, 262(6):25–26, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-25.pdf.

Powell:1990:SWS

[181] Corey S. Powell. Science writ small. Scientific American, 262(6):26, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-26a.pdf.

Kinoshita:1990:AF

[182] June Kinoshita. Atomic fountain. Scientific American, 262(6):26–31, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-26b.pdf.

H:1990:FRI

[183] J. H. “Flying Ring” inventor revamps the boomerang. Scientific American, 262(6):30, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n6/pdf/scientificamerican0690-
30.pdf.

Ross:1990:MBS

[184] Philip E. Ross. Man bites shark. Scientific American, 262(6):31, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-31a.pdf.

Rennie:1990:OGG

[185] John Rennie. Old green genes. Scientific American, 262(6):31–34, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-31b.pdf.

H:1990:S

[186] J. H. Skullduggery. Scientific American, 262(6):34, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-34a.pdf.

Rusting:1990:ET

[187] Ricki Rusting. Easing the trauma. Scientific American, 262(6):34–36, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-34b.pdf.

Horgan:1990:OSb

[188] John Horgan. Overview: Schizophrenia. Scientific American, 262(6):37– 40, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-37.pdf.

Ulmann:1990:R

[189] André Ulmann, Georges Teutsch, and Daniel Philibert. Ru 486. Scientific American, 262(6):42–48, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n6/pdf/scientificamerican0690-
42.pdf.

Binzel:1990:P

[190] Richard P. Binzel. Pluto. Scientific American, 262(6):50–58 (Intl. ed. 26– ??), June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-50.pdf.

Suga:1990:BNC

[191] Nobuo Suga. Biosonar and neural computation in bats. Scientific American, 262(6):60–68, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n6/pdf/scientificamerican0690-
60.pdf.

Macdonald:1990:MOR

[192] Kenneth C. Macdonald and Paul J. Fox. The Mid-Ocean Ridge. Scientific American, 262(6):72–79 (Intl. ed. 42–??), June 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-72.pdf.

Nahin:1990:OHa

[193] Paul J. Nahin. Oliver Heaviside. Scientific American, 262(6):80–??, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1990:SBTa

[194] Elizabeth Corcoran. Science and business: Talking policy. Scientific American, 262(6):82–84, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n6/pdf/scientificamerican0690-
82.pdf.

Erickson:1990:SW

[195] Deborah Erickson. Sick of work. Scientific American, 262(6):84–86, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-84.pdf.

Ross:1990:DB

[196] Philip E. Ross. Dust busters. Scientific American, 262(6):86–88, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-86.pdf.

E:1990:PP

[197] D. E. Patent power. Scientific American, 262(6):88–93, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-88.pdf.

Anonymous:1990:SHO

[198] Anonymous. Sunspots and how to observe them safely. Scientific American, 262(6):88–??, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

E:1990:RP

[199] D. E. Remote possibility. Scientific American, 262(6):93, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-93.pdf.

Wallich:1990:AEc

[200] Paul Wallich. The analytical economist. Scientific American, 262(6):94– 95, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-94.pdf.

Pietsch:1990:SPF

[201] Theodore W. Pietsch and David B. Grobecker. Science in pictures: Frogfishes. Scientific American, 262(6):96–103, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-96.pdf.

Pietsch:1990:F

[202] Theodore W. Pietsch and David B. Grobecker. Frogfishes. Scientific American, 262(6):96–?? (Intl. ed. 56–??), June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lounasmaa:1990:S

[203] Olli V. Lounasmaa and George Pickett. The ^{3}He superfluids. Scientific American, 262(6):104–111 (Intl. ed. 64–??), June 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-104.pdf.

Reganold:1990:SA

[204] John P. Reganold, Robert I. Papendick, and James F. Parr. Sustainable agriculture. Scientific American, 262(6):112–120, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-112.pdf.

Nahin:1990:OHb

[205] Paul J. Nahin. Oliver Heaviside. Scientific American, 262(6):122–129, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n6/pdf/scientificamerican0690-122.pdf.

Nahin:1990:OHBa

[206] Paul J. Nahin. Oliver Heaviside (biography). Scientific American, 262 (6):122–129, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mims:1990:ASa

[207] Forrest M. Mims III. The amateur scientist. Scientific American, 262(6):130–133, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n6/pdf/scientificamerican0690-
130.pdf.

Morrison:1990:Bd

[208] Philip Morrison. Books. Scientific American, 262(6):134–139, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-134.pdf.

Gomory:1990:E

[209] Ralph Gomory. Essay. Scientific American, 262(6):140, June 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v262/
n6/pdf/scientificamerican0690-140.pdf.

Anonymous:1990:Lg

[210] Anonymous. Letters. Scientific American, 263(1):12, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-12.pdf.

Anonymous:1990:YAg

[211] Anonymous. 50 and 100 years ago. Scientific American, 263(1):14, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-14.pdf.

Horgan:1990:SCD

[212] John Horgan. Science and the citizen: The death penalty. Scientific American, 263(1):17–20, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n1/pdf/scientificamerican0790-
17.pdf.

H:1990:BCO

[213] J. H. As birth-control options in the U.S. dwindle, more married women turn to sterilization. Scientific American, 263(1):18, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-18.pdf.

H:1990:SN

[214] J. H. Strength in numbers. Scientific American, 263(1):20, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-20.pdf.

Janowski:1990:MI

[215] Pat Janowski. Mapping the immeasurable. Scientific American, 263(1): 24–25, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-24.pdf.

Rennie:1990:MM

[216] John Rennie. Of mice and MS. Scientific American, 263(1):25, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-25a.pdf.

H:1990:PB

[217] J. H. Pythagoras’s bells. Scientific American, 263(1):25–26, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-25b.pdf.

Beardsley:1990:SG

[218] Tim Beardsley. Shocking genes. Scientific American, 263(1):26, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-26.pdf.

H:1990:NLS

[219] J. H. A nasty little squirt. Scientific American, 263(1):28, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-28a.pdf.

Ross:1990:GM

[220] Philip E. Ross. Gothic mystery. Scientific American, 263(1):28, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-28b.pdf.

B:1990:OE

[221] T. M. B. Orbiting eye. Scientific American, 263(1):30, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-30a.pdf.

Powell:1990:LC

[222] Corey S. Powell. Lost in the clouds. Scientific American, 263(1):30– 32, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-30b.pdf.

Ruthen:1990:PMM

[223] Russell Ruthen. Profile: Modest maverick. Scientific American, 263(1): 32–35, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-32.pdf.

White:1990:GCD

[224] Robert M. White. The Great Climate Debate. Scientific American, 263(1):36–43, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n1/pdf/scientificamerican0790-
36.pdf.

DeRobertis:1990:HGV

[225] Eddy M. De Robertis, Guillermo Oliver, and Christopher V. E. Wright. Homeobox genes and the vertebrate body plan. Scientific American, 263(1):46–52 (Intl. ed. 26–??), July 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-46.pdf.

Corcoran:1990:TANa

[226] Elizabeth Corcoran and Tim Beardsley. Trends in aerospace: The new space race. Scientific American, 263(1):50–??, July 1990. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Myers:1990:LC

[227] Stephen Myers and Emilio Picasso. The LEP collider. Scientific American, 263(1):54–61 (Intl. ed. 34–??), July 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-54.pdf.

Atkinson:1990:WCD

[228] Mark A. Atkinson and Noel K. Maclaren. What causes diabetes? Scientific American, 263(1):62–71, July 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-62.pdf.

Corcoran:1990:TANb

[229] Elizabeth Corcoran and Tim Beardsley. Trends in aerospace: The new space race. Scientific American, 263(1):72–85, July 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-72.pdf.

Corcoran:1990:NSR

[230] Elizabeth Corcoran and Tim Beardsley. The new space race. Scientific American, 263(1):72–??, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sheldrick:1990:SBD

[231] Michael G. Sheldrick. Science and business: Driving while automated. Scientific American, 263(1):86–88, July 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-86.pdf.

Anonymous:1990:OJA

[232] Anonymous. An odd journey along even roads leads to home in Golygon City. Scientific American, 263(1):86–??, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1990:DYS

[233] Deborah Erickson. Do you see what I see? Scientific American, 263(1): 88–89, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-88.pdf.

Rennie:1990:MMS

[234] John Rennie. Move over, Mr. Spock. Scientific American, 263(1):89– 92, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-89.pdf.

E:1990:NH

[235] D. E. Now hear this. Scientific American, 263(1):92–93, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-92.pdf.

Wallich:1990:AEd

[236] Paul Wallich. The analytical economist. Scientific American, 263(1): 93, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-93.pdf.

Conkling:1990:P

[237] John A. Conkling. Pyrotechnics. Scientific American, 263(1):96–102, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-96.pdf.

Newhouse:1990:CB

[238] Joseph R. Newhouse. Chestnut blight. Scientific American, 263(1):106– 111, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n1/pdf/scientificamerican0790-106.pdf.

Graham:1990:RT

[239] Ronald L. Graham and Joel H. Spencer. Ramsey theory. Scientific American, 263(1):112–117, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n1/pdf/scientificamerican0790-
112.pdf.

Dewdney:1990:MRc

[240] A. K. Dewdney. Mathematical recreations. Scientific American, 263(1):118–121, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n1/pdf/scientificamerican0790-
118.pdf.

Morrison:1990:Be

[241] Philip Morrison. Books. Scientific American, 263(1):122–127, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-122.pdf.

Zinder:1990:E

[242] Norton D. Zinder. Essay. Scientific American, 263(1):128, July 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n1/pdf/scientificamerican0790-128.pdf.

Anonymous:1990:Ea

[243] Anonymous. Errata. Scientific American, 263(2):10, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-10a.pdf.

Anonymous:1990:Lh

[244] Anonymous. Letters. Scientific American, 263(2):10, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-10.pdf.

Anonymous:1990:YAh

[245] Anonymous. 50 and 100 years ago. Scientific American, 263(2):12, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-12.pdf.

Wallich:1990:SCD

[246] Paul Wallich. Science and the citizen: Dark days. Scientific American, 263(2):16–20, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
16.pdf.

Guterl:1990:BCB

[247] Fred Guterl. Big country, big problems. Scientific American, 263(2):20, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-20.pdf.

Holloway:1990:HG

[248] Marguerite Holloway. Hot geese. Scientific American, 263(2):22, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-22.pdf.

Beardsley:1990:OK

[249] Tim Beardsley. Oravske Kuru. Scientific American, 263(2):24–26, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-24.pdf.

Horgan:1990:BA

[250] John Horgan. Bufo abuse. Scientific American, 263(2):26–27, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-26.pdf.

Rusting:1990:CT

[251] Ricki Rusting. A clock in the trees. Scientific American, 263(2):30, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-30a.pdf.

Horgan:1990:UM

[252] John Horgan. Unsolved mysteries. Scientific American, 263(2):30–32, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-30b.pdf.

B:1990:STT

[253] T. M. B. The sea turtle’s tale. Scientific American, 263(2):32, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-32.pdf.

Beardsley:1990:PGD

[254] Tim Beardsley. Profile: Gene doctor. Scientific American, 263 (2):33–33B, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
33.pdf.

Nolan:1990:TWB

[255] Janne E. Nolan and Albert D. Wheelon. Third World ballistic missiles. Scientific American, 263(2):34–40, August 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-34.pdf.

Cava:1990:SB

[256] Robert J. Cava. Superconductors beyond 1-2-3. Scientific American, 263 (2):42–49 (Intl. ed. 24–??), August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
42.pdf.

Mills:1990:ARI

[257] John Mills and Henry Masur. AIDS-related infections. Scientific American, 263(2):50–57, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
50.pdf.

Jurgens:1990:LF

[258] Hartmut Jürgens, Heinz-Otto Peitgen, and Dietmar Saupe. The language of fractals. Scientific American, 263(2):60–67, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-60.pdf.

Berry:1990:WMF

[259] R. Stephen Berry. When the melting and freezing points are not the same. Scientific American, 263(2):68–74, August 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-68.pdf.

Handel:1990:SDA

[260] Steven N. Handel and Andrew J. Beattie. Seed dispersal by ants. Scientific American, 263(2):76–83A, August 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-76.pdf.

Jones:1990:GWT

[261] Philip D. Jones and Tom M. L. Wigley. Global warming trends. Scientific American, 263(2):84–91, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
84.pdf.

Anonymous:1990:HMU

[262] Anonymous. How to monitor ultraviolet radiation from the Sun. Scientific American, 263(2):86–??, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kinoshita:1990:SPM

[263] June Kinoshita. Science in pictures: Maya art for the record. Scientific American, 263(2):92–97, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
92.pdf.

Kinoshita:1990:MAR

[264] June Kinoshita. Maya art for the record. Scientific American, 263(2):92– ?? (Intl. ed. 74–??), August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1990:SBTb

[265] Elizabeth Corcoran. Science and business: a thin line. Scientific American, 263(2):98–100, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
98.pdf.

Powell:1990:PGG

[266] Corey S. Powell. Plastic goes green. Scientific American, 263(2):101, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-101.pdf.

Erickson:1990:P

[267] Deborah Erickson. Down on the pharm. Scientific American, 263(2):102–103, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
102.pdf.

E:1990:SS

[268] D. E. Solar sells. Scientific American, 263(2):103, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-103.pdf.

Corcoran:1990:AEd

[269] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 263(2):104, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
104.pdf.

Mims:1990:ASb

[270] Forrest M. Mims III. The amateur scientist. Scientific American, 263(2):106–109, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n2/pdf/scientificamerican0890-
106.pdf.

Anon:1990:AS

[271] Anon. Amateur scientist. Scientific American, 263(2):106–??, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1990:Bf

[272] Philip Morrison. Books. Scientific American, 263(2):110–115, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n2/pdf/scientificamerican0890-110.pdf.

Rothschild:1990:E

[273] Miriam Rothschild. Essay. Scientific American, 263(2):116, August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n2/pdf/scientificamerican0890-116.pdf.

Anonymous:1990:Eb

[274] Anonymous. Erratum. Scientific American, 263(3):12, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-12a.pdf.

Anonymous:1990:Li

[275] Anonymous. Letters. Scientific American, 263(3):12, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-12.pdf.

Anonymous:1990:YAi

[276] Anonymous. 50 and 100 years ago. Scientific American, 263(3):16, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-16.pdf.

Beardsley:1990:SCB

[277] Tim Beardsley. Science and the citizen: Blurred reflections. Scientific American, 263(3):21–29, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
21.pdf.

B:1990:MEM

[278] T. M. B. Molecular engineers mimic mother nature. Scientific American, 263(3):30, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
30b.pdf.

B:1990:BF

[279] T. M. B. Boosting fusion. Scientific American, 263(3):30–31, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-30a.pdf.

Rusting:1990:HJ

[280] Ricki Rusting. Hatchet job. Scientific American, 263(3):34, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-34a.pdf.

Wallich:1990:DD

[281] Paul Wallich. Digital desperados. Scientific American, 263(3):34–38, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-34b.pdf.

Ross:1990:LB

[282] Philip E. Ross. Lorenz’s butterfly. Scientific American, 263(3):42, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-42.pdf.

Holloway:1990:FL

[283] Marguerite Holloway. Fetal law. Scientific American, 263(3):46–48, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-46.pdf.

B:1990:EST

[284] T. M. B. Easier said than done. Scientific American, 263(3):48–50, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-48.pdf.

Rennie:1990:OTS

[285] John Rennie. Overview: Tolerating self. Scientific American, 263(3):50–53, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
50.pdf.

Davis:1990:EPEa

[286] Ged R. Davis. Energy for Planet Earth. Scientific American, 263(3): 54–60, 62 (Intl. ed. 20–27), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bleviss:1990:EMVa

[287] Deborah L. Bleviss and Peter Walzer. Energy for motor vehicles. Scientific American, 263(3):54–61 (or 102–109??), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Davis:1990:EPEb

[288] Ged R. Davis. Energy for Planet Earth. Scientific American, 263(3):54–62, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
54.pdf.

Fickett:1990:EUE

[289] Arnold P. Fickett, Clark W. Gellings, and Amory B. Lovins. Efficient use of electricity. Scientific American, 263(3):64–74, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-64.pdf.

Chandler:1990:ESUa

[290] William U. Chandler, Alexei A. Makarov, and Zhou Dadi. Energy for the Soviet Union, Eastern Europe and China. Scientific American, 263 (3):74–80 (or 120–127??), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bevington:1990:EBH

[291] Rick Bevington and Arthur H. Rosenfeld. Energy for buildings and homes. Scientific American, 263(3):76–86, September 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-76.pdf.

Ross:1990:EI

[292] Marc H. Ross and Daniel Steinmeyer. Energy for industry. Scientific American, 263(3):88–98, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
88.pdf.

Bleviss:1990:EMVb

[293] Deborah L. Bleviss and Peter Walzer. Energy for motor vehicles. Scientific American, 263(3):102–109, September 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-102.pdf.

Reddy:1990:EDW

[294] Amulya K. N. Reddy and José Goldemberg. Energy for the developing world. Scientific American, 263(3):110–118 (or 62–72??) (Intl. ed. 62–72), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
110.pdf.

Chandler:1990:ESUb

[295] William U. Chandler, Alexei A. Makarov, and Zhou Dadi. Energy for the Soviet Union, Eastern Europe and China. Scientific American, 263(3):120–127, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
120.pdf.

Anonymous:1990:HRC

[296] Anonymous. How to resurrect a cat from its grin. Scientific American, 263(3):124–??, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fulkerson:1990:EFF

[297] William Fulkerson, Roddie R. Judkins, and Manoj K. Sanghvi. Energy from fossil fuels. Scientific American, 263(3):128–135 (Intl. ed. 82–89), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-128.pdf.

Hafele:1990:ENP

[298] Wolf Häfele. Energy from nuclear power. Scientific American, 263(3):136–144, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
136.pdf.

Weinberg:1990:ES

[299] Carl J. Weinberg and Robert H. Williams. Energy from the Sun. Scientific American, 263(3):146–155 (Intl. ed. 98–106), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-146.pdf.

Holdren:1990:ET

[300] John P. Holdren. Energy in transition. Scientific American, 263(3):156– 163 (Intl. ed. 108–115), September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
156.pdf.

Erickson:1990:SBB

[301] Deborah Erickson. Science and business: Big-time orphan. Scientific American, 263(3):164–166, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
164.pdf.

Stix:1990:MM

[302] Gary Stix. Meteoric messages. Scientific American, 263(3):167, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-167.pdf.

E:1990:SSI

[303] D. E. Skin stand-ins. Scientific American, 263(3):168, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-168.pdf.

S:1990:MU

[304] G. S. Muffling Umklapp. Scientific American, 263(3):169, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-169a.pdf.

Corcoran:1990:ES

[305] Elizabeth Corcoran. Electron switches. Scientific American, 263 (3):169–170, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
169b.pdf.

Rennie:1990:AC

[306] John Rennie. Anatomical cartography. Scientific American, 263(3):170, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-170.pdf.

Wallich:1990:AEe

[307] Paul Wallich and Elizabeth Corcoran. The analytical economist. Scientific American, 263(3):171, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
171.pdf.

Dewdney:1990:MRd

[308] A. K. Dewdney. Mathematical recreations. Scientific American, 263(3):174–177, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n3/pdf/scientificamerican0990-
174.pdf.

Morrison:1990:Bg

[309] Philip Morrison. Books. Scientific American, 263(3):178–183, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n3/pdf/scientificamerican0990-178.pdf.

Malpas:1990:E

[310] Robert Malpas. Essay. Scientific American, 263(3):184, September 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n3/pdf/scientificamerican0990-184.pdf.

Anonymous:1990:Lj

[311] Anonymous. Letters. Scientific American, 263(4):10, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-10.pdf.

Anonymous:1990:YAj

[312] Anonymous. 50 and 100 years ago. Scientific American, 263(4):14–16, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-14.pdf.

Beardsley:1990:SCN

[313] Tim Beardsley. Science and the citizen: New order. Scientific American, 263(4):18–24, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
18.pdf.

Hsu:1990:GCMa

[314] Feng hsiung Hsu, Thomas Anantharaman, Murray Campbell, and Andreas Nowatzyk. A grandmaster chess machine. Scientific American, 263 (4):18–24, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wittekindt:1990:LAC

[315] Holger Wittekindt. A Latin American caterpillar calls ants to a free meal. Scientific American, 263(4):24, October 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-24.pdf.

Ross:1990:VU

[316] Philip E. Ross. Venus unveiled. Scientific American, 263(4):26, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-26.pdf.

B:1990:FP

[317] T. M. B. Fowl play. Scientific American, 263(4):28, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-28a.pdf.

Rusting:1990:SL

[318] Ricki Rusting. Seeing the light. Scientific American, 263(4):28–30, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-28b.pdf.

B:1990:AA

[319] T. M. B. Adjusted agenda. Scientific American, 263(4):30–32, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-30.pdf.

Cooke:1990:TT

[320] David Cooke. Teaching teachers. Scientific American, 263(4):34, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-34a.pdf.

Yam:1990:SM

[321] Philip Yam. Stardust memories. Scientific American, 263(4):34–37, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-34b.pdf.

B:1990:CC

[322] T. M. B. Cannabis comprehended. Scientific American, 263(4):38, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-38.pdf.

Holloway:1990:PVD

[323] Marguerite Holloway. Profile: Vive la différence. Scientific American, 263(4):40–42, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
40.pdf.

Hsu:1990:GCMb

[324] Feng hsiung Hsu, Thomas Anantharaman, Murray Campbell, and Andreas Nowatzyk. A grandmaster chess machine. Scientific American, 263(4):44–50, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
44.pdf.

Feng-hsiung:1990:GCM

[325] Hsu Feng-hsiung, Thomas Anantharaman, Murray Campbell, and Andreas Nowatzyk. A grandmaster chess machine. Scientific American, 263 (4):44–??, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sapienza:1990:PIG

[326] Carmen Sapienza. Parental imprinting of genes. Scientific American, 263(4):52–60, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
52.pdf.

Spindel:1990:OATa

[327] Robert C. Spindel and Peter F. Worcester. Ocean acoustic tomography. Scientific American, 263(4):62–67 (or 94–??), October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pepper:1990:PE

[328] David M. Pepper, Jack Feinberg, and Nicolai V. Kukhtarev. The photorefractive effect. Scientific American, 263(4):62–74, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-62.pdf.

Horgan:1990:TCUa

[329] John Horgan. Trends in cosmology: Universal truths. Scientific American, 263(4):74–??, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1990:DWC

[330] Anonymous. Debate: What caused the Mass Extinction? Scientific American, 263(4):76–77, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
76.pdf.

Alvarez:1990:WCM

[331] Walter Alvarez and Frank Asaro. What caused the mass extinction? an extraterrestrial impact. Scientific American, 263(4):76–?? (Intl. ed.

44–??), October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Courtillot:1990:WCM

[332] Vincent E. Courtillot. What caused the mass extinction? a volcanic eruption. Scientific American, 263(4):76–?? (Intl. ed. 53–??), October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Alvarez:1990:EI

[333] Walter Alvarez and Frank Asaro. An extraterrestrial impact. Scientific American, 263(4):78–84, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
78.pdf.

Courtillot:1990:VE

[334] Vincent E. Courtillot. A volcanic eruption. Scientific American, 263(4):85–92, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
85.pdf.

Anonymous:1990:RCC

[335] Anonymous. A remote-control camera that catches the wind and captures the landscape. Scientific American, 263(4):92–??, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Spindel:1990:OATb

[336] Robert C. Spindel and Peter F. Worcester. Ocean acoustic tomography. Scientific American, 263(4):94–99, October 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-94.pdf.

Barth:1990:BNC

[337] Rolf F. Barth, Albert H. Soloway, and Ralph G. Fairchild. Boron neutron capture therapy for cancer. Scientific American, 263(4):100–107, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-100.pdf.

Horgan:1990:TCUb

[338] John Horgan. Trends in cosmology: Universal truths. Scientific American, 263(4):108–117, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
108.pdf.

Horgan:1990:UT

[339] John Horgan. Universal truths. Scientific American, 263(4):108–??, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1990:SBG

[340] Gary Stix. Science and business: Gigabit connection. Scientific American, 263(4):118–120, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
118.pdf.

Erickson:1990:AA

[341] Deborah Erickson. On again, off again. Scientific American, 263(4):121–122, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
121.pdf.

S:1990:CC

[342] G. S. Cold cathodes. Scientific American, 263(4):122–123, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-122.pdf.

S:1990:IW

[343] G. S. Infrastructure woes. Scientific American, 263(4):123, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-123.pdf.

S:1990:LP

[344] G. S. Less of a problem? Scientific American, 263(4):124, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-124.pdf.

Wallich:1990:AEf

[345] Paul Wallich. The analytical economist. Scientific American, 263(4):125, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-125.pdf.

Mims:1990:ASc

[346] Forrest M. Mims III. The amateur scientist. Scientific American, 263(4):126–129, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n4/pdf/scientificamerican1090-
126.pdf.

Morrison:1990:Bh

[347] Philip Morrison. Books. Scientific American, 263(4):130–134, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n4/pdf/scientificamerican1090-130.pdf.

Johnson:1990:E

[348] James L. Johnson. Essay. Scientific American, 263(4):135, October 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n4/pdf/scientificamerican1090-135.pdf.

Anonymous:1990:Ec

[349] Anonymous. Errata. Scientific American, 263(5):10, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-10b.pdf.

Anonymous:1990:Lk

[350] Anonymous. Letters. Scientific American, 263(5):10, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-10a.pdf.

Anonymous:1990:YAk

[351] Anonymous. 50 and 100 years ago. Scientific American, 263(5):14, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-14.pdf.

Holloway:1990:SCG

[352] Marguerite Holloway. Science and the citizen: “a great poison”. Scientific American, 263(5):16–20, November 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-16.pdf.

H:1990:QAM

[353] M. H. Questions about a major herbicide. Scientific American, 263(5):20, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-20.pdf.

Stix:1990:I

[354] Gary Stix. The ø’s have it. Scientific American, 263(5):24, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-24.pdf.

Erickson:1990:TC

[355] Deborah Erickson. Trojan cow. Scientific American, 263(5):26, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-26a.pdf.

Yam:1990:TSK

[356] Philip Yam. Terrorist shrimp kidnaps for defense. Scientific American, 263(5):26, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
26b.pdf.

Ruthen:1990:LM

[357] Russell Ruthen. Light motif. Scientific American, 263(5):29, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-29a.pdf.

Horgan:1990:SC

[358] John Horgan. Stinging criticism. Scientific American, 263(5):29–32, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-29b.pdf.

H:1990:LS

[359] J. H. Liquid sky. Scientific American, 263(5):32, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-32.pdf.

Beardsley:1990:RD

[360] Tim Beardsley. Recovery drill. Scientific American, 263(5):34, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-34.pdf.

Ross:1990:ODS

[361] Philip E. Ross. Overview: Dead Sea scrolls. Scientific American, 263(5):36–38, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
36.pdf.

Rosenberg:1990:STW

[362] Nathan Rosenberg and L. E. Birdzell, Jr. Science, technology and the Western Miracle. Scientific American, 263(5):42–54, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-42.pdf.

Rees:1990:BHG

[363] Martin J. Rees. Black holes in galactic centers. Scientific American, 263(5):56–66, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
56.pdf.

Verma:1990:GT

[364] Inder M. Verma. Gene therapy. Scientific American, 263(5):68–84, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-68.pdf.

Corcoran:1990:TMDa

[365] Elizabeth Corcoran. Trends in materials: Diminishing dimensions. Scientific American, 263(5):74–??, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Winson:1990:MD

[366] Jonathan Winson. The meaning of dreams. Scientific American, 263(5):86–96, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
86.pdf.

Anonymous:1990:CMA

[367] Anonymous. A compendium of math abuse from around the world. Scientific American, 263(5):92–??, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jones:1990:KTS

[368] Vaughan F. R. Jones. Knot theory and statistical mechanics. Scientific American, 263(5):98–103, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
98.pdf.

Hodge:1990:RF

[369] A. Trevor Hodge. A Roman factory. Scientific American, 263(5):106–111, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-106.pdf.

Wootton:1990:MDI

[370] Robin J. Wootton. The mechanical design of insect wings. Scientific American, 263(5):114–120 (Intl. ed. 66–??), November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-114.pdf.

Corcoran:1990:DD

[371] Elizabeth Corcoran. Diminishing dimensions. Scientific American, 263 (5):122–126, 128–131, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1990:TMDb

[372] Elizabeth Corcoran. Trends in materials: Diminishing dimensions. Scientific American, 263(5):122–131, November 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n5/pdf/scientificamerican1190-122.pdf.

Erickson:1990:SBE

[373] Deborah Erickson. Science and business: Electronic earful. Scientific American, 263(5):132–134, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
132.pdf.

Stix:1990:BE

[374] Gary Stix. Bug-eyed. Scientific American, 263(5):134–135, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-134.pdf.

E:1990:B

[375] D. E. Not biochips? Scientific American, 263(5):136–138, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-136.pdf.

Ross:1990:MWL

[376] Philip E. Ross. A million watts of light. Scientific American, 263(5):138–139, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
138.pdf.

Corcoran:1990:AEe

[377] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 263(5):140, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
140.pdf.

Dewdney:1990:MRe

[378] A. K. Dewdney. Mathematical recreations. Scientific American, 263(5):142–145, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n5/pdf/scientificamerican1190-
142.pdf.

Morrison:1990:Bi

[379] Philip Morrison. Books. Scientific American, 263(5):146–151, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-146.pdf.

Westland:1990:E

[380] J. Christopher Westland. Essay. Scientific American, 263(5):152, November 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n5/pdf/scientificamerican1190-152.pdf.

Anonymous:1990:Ed

[381] Anonymous. Errata. Scientific American, 263(6):12, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n6/pdf/scientificamerican1290-12a.pdf.

Anonymous:1990:Ll

[382] Anonymous. Letters. Scientific American, 263(6):12, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n6/pdf/scientificamerican1290-12.pdf.

Anon:1990:L

[383] Anon. Letters. Scientific American, 263(6):12–??, December 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1990:YAl

[384] Anonymous. 50 and 100 years ago. Scientific American, 263(6):14, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-14.pdf.

Anon:1990:YA

[385] Anon. 50 and 100 years ago. Scientific American, 263(6):14–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1990:SCM

[386] John Horgan. Science and the citizen: Measuring eternity. Scientific American, 263(6):16–20, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
16.pdf.

Anon:1990:SC

[387] Anon. Science and the citizen. Scientific American, 263(6):16–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1990:F

[388] Corey S. Powell. Far out. Scientific American, 263(6):20–23, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-20.pdf.

Horgan:1990:GP

[389] John Horgan. Going to pot. Scientific American, 263(6):23–24, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-23.pdf.

Stix:1990:BH

[390] Gary Stix. Beam of hope. Scientific American, 263(6):24–25, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-24.pdf.

Horgan:1990:DT

[391] John Horgan. Double trouble. Scientific American, 263(6):25–28, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-25.pdf.

Yam:1990:ATK

[392] Philip Yam. Acacia trees kill antelope in the transvaal. Scientific American, 263(6):28, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
28a.pdf.

Beardsley:1990:CA

[393] Tim Beardsley. Clearing the airways. Scientific American, 263(6):28–33, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-28b.pdf.

Holloway:1990:MMC

[394] Marguerite Holloway. Mussels may clump together to improve their food supply. Scientific American, 263(6):30, December 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n6/pdf/scientificamerican1290-30.pdf.

Beardsley:1990:PDG

[395] Tim Beardsley. Profile: Dr. Greenhouse. Scientific American, 263(6):33–36, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
33.pdf.

Zewail:1990:BMU

[396] A. H. Zewail. The birth of molecules (ultrafast photographic observations). Scientific American, 263(6):40–46, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Blair:1990:ANW

[397] Bruce G. Blair and Henry W. Kendall. Accidental nuclear war. Scientific American, 263(6):53–58, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
53.pdf.

Saunders:1990:SPS

[398] R. Stephen Saunders. Science in pictures: The surface of Venus. Scientific American, 263(6):60–62, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
60.pdf.

Saunders:1990:SV

[399] R. Stephen Saunders. The surface of Venus. Scientific American, 263 (6):60–?? (Intl. ed. 26–31), December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Todorov:1990:HCM

[400] Igor N. Todorov. How cells maintain stability. Scientific American, 263(6):66–75, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
66.pdf.

Zewail:1990:BMa

[401] Ahmed H. Zewail. The birth of molecules. Scientific American, 263(6):76–82, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
76.pdf.

Rennie:1990:TIBa

[402] John Rennie. Trends in immunology: The body against itself. Scientific American, 263(6):76–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zewail:1990:BMb

[403] Ahmed H. Zewail. The birth of molecules. Scientific American, 263 (6):76–?? (Intl. ed. 40–??), December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rock:1990:LGP

[404] Irvin Rock and Stephen Palmer. The legacy of Gestalt psychology. Scientific American, 263(6):84–90, December 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n6/pdf/scientificamerican1290-84.pdf.

Storey:1990:FA

[405] Kenneth B. Storey and Janet M. Storey. Frozen and alive. Scientific American, 263(6):92–97, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
92.pdf.

Anonymous:1990:FCT

[406] Anonymous. Fermat’s Christmas Theorem is explained in one Dickens of a tale. Scientific American, 263(6):94–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stringer:1990:EMH

[407] Christopher B. Stringer. The emergence of modern humans. Scientific American, 263(6):98–104, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
98.pdf.

Rennie:1990:TIBb

[408] John Rennie. Trends in immunology: The body against itself. Scientific American, 263(6):106–115, December 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v263/
n6/pdf/scientificamerican1290-106.pdf.

Rennie:1990:BAI

[409] John Rennie. The body against itself. Scientific American, 263(6):106– ??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1990:SBN

[410] Gary Stix. Science and business: Narrow bandwidth. Scientific American, 263(6):116–118, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
116.pdf.

Anon:1990:SB

[411] Anon. Science and business. Scientific American, 263(6):116–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1990:NY

[412] Gary Stix. This number is you. Scientific American, 263(6):118–119, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-118.pdf.

Erickson:1990:SF

[413] Deborah Erickson. Sugar fix? Scientific American, 263(6):119–122, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-119.pdf.

Stix:1990:FE

[414] Gary Stix. Field effects. Scientific American, 263(6):122–123, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-122.pdf.

Beardsley:1990:FT

[415] Tim Beardsley. Food for thought. Scientific American, 263(6):125, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-125.pdf.

Wallich:1990:GG

[416] Paul Wallich. Garbage in, garbage out. Scientific American, 263(6):126, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-126.pdf.

Corcoran:1990:AEf

[417] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 263(6):127, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
127.pdf.

Stewart:1990:MR

[418] Ian Stewart. Mathematical recreations. Scientific American, 263 (6):128–131, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
128.pdf.

Anon:1990:MR

[419] Anon. Mathematical recreations. Scientific American, 263(6):128–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1990:Bj

[420] Philip Morrison and Phylis Morrison. Books. Scientific American, 263(6):132–139, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
132.pdf.

Anon:1990:B

[421] Anon. Books. Scientific American, 263(6):132–??, December 1990. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1990:AI

[422] Anonymous. Annual index 1990. Scientific American, 263(6):140–143, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v263/n6/pdf/scientificamerican1290-140.pdf.

Anon:1990:AI

[423] Anon. Annual index. Scientific American, 263(6):140–??, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mckay:1990:E

[424] Christopher P. Mckay and Robert H. Haynes. Essay. Scientific American, 263(6):144, December 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v263/n6/pdf/scientificamerican1290-
144.pdf.

Anonymous:1991:Ea

[425] Anonymous. Erratum. Scientific American, 264(1):12, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-12b.pdf.

Anonymous:1991:La

[426] Anonymous. Letters. Scientific American, 264(1):12, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-12a.pdf.

Anonymous:1991:YAa

[427] Anonymous. 50 and 100 years ago. Scientific American, 264(1):14, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-14.pdf.

Cuterl:1991:SCC

[428] Fred Cuterl and Russell Ruthen. Science and the citizen: Chunnel vision. Scientific American, 264(1):22–28, January 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-22.pdf.

Yam:1991:SOC

[429] Philip Yam. Short-order cook. Scientific American, 264(1):28, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-28.pdf.

Holloway:1991:NV

[430] Marguerite Holloway. Neural vector. Scientific American, 264(1):32, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-32.pdf.

Beardsley:1991:TF

[431] Tim Beardsley. Trans fat. Scientific American, 264(1):34, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-34a.pdf.

Wallich:1991:WT

[432] Paul Wallich. Wavelet theory. Scientific American, 264(1):34–35, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-34b.pdf.

Rennie:1991:LC

[433] John Rennie. Leaky channels. Scientific American, 264(1):35–36, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-35.pdf.

Beardsley:1991:PAD

[434] Tim Beardsley. Profile: AIDS dispute. Scientific American, 264(1):36– 37, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-36.pdf.

Richelson:1991:FSR

[435] Jeffrey T. Richelson. The future of space reconnaissance. Scientific American, 264(1):38–44, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
38.pdf.

Bak:1991:SOC

[436] Per Bak and Kan Chen. Self-organized criticality. Scientific American, 264(1):46–53 (Intl. ed. 26–??), January 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-46.pdf.

Richards:1991:PFP

[437] Frederic M. Richards. The protein folding problem. Scientific American, 264(1):54–63, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
54.pdf.

Scaglia:1991:BCF

[438] Gustina Scaglia. Building the Cathedral in Florence. Scientific American, 264(1):66–72, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
66.pdf.

Corcoran:1991:TCCa

[439] Elizabeth Corcoran. Trends in computing: Calculating reality. Scientific American, 264(1):74–??, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Black:1991:WAO

[440] David C. Black. Worlds around other stars. Scientific American, 264(1):76–83, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
76.pdf.

Weissmann:1991:A

[441] Gerald Weissmann. Aspirin. Scientific American, 264(1):84–90, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-84.pdf.

Anonymous:1991:TCG

[442] Anonymous. Tools for computer graphics make an invisible world seem less alien. Scientific American, 264(1):90–??, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Davies:1991:CCH

[443] Nicholas B. Davies and Michael Brooke. Coevolution of the cuckoo and its hosts. Scientific American, 264(1):92–98, January 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-92.pdf.

Corcoran:1991:TCCb

[444] Elizabeth Corcoran. Trends in computing: Calculating reality. Scientific American, 264(1):100–109, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
100.pdf.

Corcoran:1991:CR

[445] Elizabeth Corcoran. Calculating reality. Scientific American, 264(1): 100–?? (Intl. ed. 74–??), January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1991:SBS

[446] Gary Stix. Science and business: Second-generation silicon. Scientific American, 264(1):110–111, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
110.pdf.

Stix:1991:FS

[447] Gary Stix. Future shock. Scientific American, 264(1):111–112, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-111.pdf.

Erickson:1991:GR

[448] Deborah Erickson. Gene rush. Scientific American, 264(1):112–113, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-112.pdf.

Stix:1991:YSD

[449] Gary Stix. Yale set designers turn to finite element analysis. Scientific American, 264(1):113, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
113.pdf.

Ross:1991:Ba

[450] Philip E. Ross. Buckeyballs. Scientific American, 264(1):114–116, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-114.pdf.

Stix:1991:ASS

[451] Gary Stix. Another small step? Scientific American, 264(1):116, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-116.pdf.

Wallich:1991:AEa

[452] Paul Wallich and Elizabeth Corcoran. The analytical economist. Scientific American, 264(1):117, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
117.pdf.

Dewdney:1991:MRa

[453] A. K. Dewdney. Mathematical recreations. Scientific American, 264(1):118–121, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n1/pdf/scientificamerican0191-
118.pdf.

Anonymous:1991:MRa

[454] Anonymous. Mathematical recreations. Scientific American, 264(1):118– ??, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1991:Ba

[455] Philip Morrison. Books. Scientific American, 264(1):122–125, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n1/pdf/scientificamerican0191-122.pdf.

Anonymous:1991:Ba

[456] Anonymous. Books. Scientific American, 264(1):122–??, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Inman:1991:ETC

[457] B. R. Inman and Daniel F. Burton. Essay: Technology and competitiveness. Scientific American, 264(1):126, January 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n1/pdf/scientificamerican0191-126.pdf.

Inman:1991:E

[458] B. R. Inman and Daniel F. Burton. Essay. Scientific American, 264 (1):126–??, January 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1991:Eb

[459] Anonymous. Errata. Scientific American, 264(2):10, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-10b.pdf.

Anonymous:1991:Lb

[460] Anonymous. Letters. Scientific American, 264(2):10, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-10a.pdf.

Anonymous:1991:YAb

[461] Anonymous. 50 and 100 years ago. Scientific American, 264(2):12, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-12.pdf.

Ruthen:1991:SCI

[462] Russell Ruthen. Science and the citizen: Iraq’s nuclear threat. Scientific American, 264(2):14–16, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
14.pdf.

Anonymous:1991:SCa

[463] Anonymous. Science and the citizen. Scientific American, 264(2):14–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1991:T

[464] Tim Beardsley. The A team. Scientific American, 264(2):16–19, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-16.pdf.

Horgan:1991:PM

[465] John Horgan. Postal mortem. Scientific American, 264(2):22–23, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-22.pdf.

Corcoran:1991:CMC

[466] Elizabeth Corcoran. Charting a more cautious path to space. Scientific American, 264(2):24, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
24b.pdf.

Powell:1991:DA

[467] Corey S. Powell. Disposable astronomy. Scientific American, 264(2):24–26, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
24a.pdf.

Powell:1991:HBG

[468] Corey S. Powell. Hubble bags a great white. Scientific American, 264(2):26, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
26.pdf.

Beardsley:1991:STS

[469] Tim Beardsley. Superconducting transistors show their speed. Scientific American, 264(2):28, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
28b.pdf.

Wallich:1991:SBa

[470] Paul Wallich. Safe bytes. Scientific American, 264(2):28–29, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-28a.pdf.

Beardsley:1991:NN

[471] Tim Beardsley. Nuclear numbers. Scientific American, 264(2):30, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-30.pdf.

Aral:1991:STD

[472] Sevgi O. Aral and King K. Holmes. Sexually transmitted diseases in the AIDS era. Scientific American, 264(2):62–69, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-62.pdf.

Feldman:1991:NFM

[473] Gary J. Feldman and Jack Steinberger. The number of families of matter. Scientific American, 264(2):70–75, February 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-70.pdf.

Freeman:1991:PP

[474] Walter J. Freeman. The physiology of perception. Scientific American, 264(2):78–85 (Intl. ed. 34–??), February 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-78.pdf.

Greenberg:1991:CA

[475] Donald P. Greenberg. Computers and architecture. Scientific American, 264(2):88–93 (or 104–??), February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Howells:1991:XRM

[476] Malcolm R. Howells, Janos Kirz, and David Sayre. X-ray microscopes. Scientific American, 264(2):88–94 (Intl. ed. 42–48), February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-88.pdf.

Rismiller:1991:E

[477] Peggy D. Rismiller and Roger S. Seymour. The echidna. Scientific American, 264(2):96–103, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
96.pdf.

Horgan:1991:TEBa

[478] John Horgan. Trends in evolution: In the beginning. … Scientific American, 264(2):100–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Greenberg:1991:SPC

[479] Donald P. Greenberg. Science in pictures: Computers and architecture. Scientific American, 264(2):104–109, February 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-104.pdf.

Kottick:1991:AH

[480] Edward L. Kottick, Kenneth D. Marshall, and Thomas J. Hendrickson. The acoustics of the harpsichord. Scientific American, 264(2):110–115, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-110.pdf.

Horgan:1991:TEBb

[481] John Horgan. Trends in evolution: in the beginning. … Scientific American, 264(2):116–125, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
116.pdf.

Horgan:1991:B

[482] John Horgan. In the beginning. Scientific American, 264(2):116–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1991:TSH

[483] Anonymous. The true story of how Theseus found his way out of the labyrinth. Scientific American, 264(2):118–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1991:SBLa

[484] Deborah Erickson. Science and business: Life in the blood. Scientific American, 264(2):126–127, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
126.pdf.

Anonymous:1991:SBa

[485] Anonymous. Science and business. Scientific American, 264(2):126–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1991:CM

[486] Gary Stix. Changing mission. Scientific American, 264(2):130–131, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-130.pdf.

Erickson:1991:ASS

[487] Deborah Erickson. An acoustic sensor spies on insects. Scientific American, 264(2):131, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
131.pdf.

Stix:1991:HD

[488] Gary Stix. Hole or doughnut. Scientific American, 264(2):132–133, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-132.pdf.

Stix:1991:R

[489] Gary Stix. Reach out. Scientific American, 264(2):134, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-134.pdf.

Wallich:1991:AEb

[490] Paul Wallich and Elizabeth Corcoran. The analytical economist. Scientific American, 264(2):135, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
135.pdf.

Stewart:1991:MRa

[491] Ian Stewart. Mathematical recreations. Scientific American, 264 (2):136–139, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n2/pdf/scientificamerican0291-
136.pdf.

Anonymous:1991:MRb

[492] Anonymous. Mathematical recreations. Scientific American, 264(2): 136–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Klass:1991:B

[493] Philip J. Klass. Books. Scientific American, 264(2):140–143, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n2/pdf/scientificamerican0291-140.pdf.

Anonymous:1991:Bb

[494] Anonymous. Books. Scientific American, 264(2):140–??, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nelson:1991:E

[495] Bill Nelson. Essay. Scientific American, 264(2):144, February 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n2/pdf/scientificamerican0291-144.pdf.

Anonymous:1991:Lc

[496] Anonymous. Letters. Scientific American, 264(3):10, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n3/pdf/scientificamerican0391-10.pdf.

Anonymous:1991:YAc

[497] Anonymous. 50 and 100 years ago. Scientific American, 264(3):12, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-12.pdf.

Horgan:1991:SCD

[498] John Horgan. Science and the citizen: Death with dignity. Scientific American, 264(3):17–20, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
17.pdf.

Anonymous:1991:SCb

[499] Anonymous. Science and the citizen. Scientific American, 264(3):17–??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1991:MPA

[500] John Horgan. A modest proposal on altruism. Scientific American, 264(3):20, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
20.pdf.

Holloway:1991:SSa

[501] Marguerite Holloway. Sex and silversides. Scientific American, 264(3): 24, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-24.pdf.

Powell:1991:XRR

[502] Corey S. Powell. X-ray riddle. Scientific American, 264(3):26, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-26.pdf.

Horgan:1991:LTU

[503] John Horgan. The leading theory of the Universe survives another attack. Scientific American, 264(3):27, March 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n3/pdf/scientificamerican0391-27b.pdf.

Rennie:1991:C

[504] John Rennie. Down for the count. Scientific American, 264(3):27–30, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-27a.pdf.

Beardsley:1991:GG

[505] Tim Beardsley. Guessing game. Scientific American, 264(3):30–33, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-30.pdf.

Beardsley:1991:NFW

[506] Tim Beardsley. A nitrogen fix for wheat. Scientific American, 264(3):32, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-32.pdf.

Yam:1991:GE

[507] Philip Yam. Grim expectations. Scientific American, 264(3):33–34, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-33.pdf.

Koppel:1991:PSS

[508] Tom Koppel. Profile: Supercomputer solo. Scientific American, 264(3):34–36, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
34.pdf.

Barton:1991:PL

[509] John H. Barton. Patenting life. Scientific American, 264(3):40–46, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-40.pdf.

Lederman:1991:T

[510] Leon M. Lederman. The Tevatron. Scientific American, 264(3):48– 55 (Intl. ed. 26–??), March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
48.pdf.

Murray:1991:WCC

[511] Andrew W. Murray and Marc W. Kirschner. What controls the cell cycle. Scientific American, 264(3):56–63, March 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n3/pdf/scientificamerican0391-56.pdf.

Ruddiman:1991:PUC

[512] William F. Ruddiman and John E. Kutzbach. Plateau uplift and climatic change. Scientific American, 264(3):66–75, March 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n3/pdf/scientificamerican0391-66.pdf.

Holloway:1991:TPR

[513] Marguerite Holloway. Trends in pharmacology: Rx for addiction. Scientific American, 264(3):70–??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Winston:1991:NO

[514] Roland Winston. Nonimaging optics. Scientific American, 264(3):76– 81 (Intl. ed. 52–??), March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
76.pdf.

Young:1991:T

[515] James A. Young. Tumbleweed. Scientific American, 264(3):82–87, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-82.pdf.

Snodgrass:1991:SACa

[516] Anthony M. Snodgrass and John L. Bintliff. Surveying ancient cities. Scientific American, 264(3):88–93, March 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n3/pdf/scientificamerican0391-88.pdf.

Anonymous:1991:MMM

[517] Anonymous. A menu of mathematical morsels, topological tidbits and puzzling plums. Scientific American, 264(3):88–??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Snodgrass:1991:SACb

[518] Anthony M. Snodgrass and John L. Bintliff. Surveying ancient cities. Scientific American, 264(3):88–?? (Intl. ed. 64–??), March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1991:TP

[519] Marguerite Holloway. Trends in pharmacology. Scientific American, 264(3):94–103, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
94.pdf.

Holloway:1991:RA

[520] Marguerite Holloway. R_{x} for addiction. Scientific American, 264(3):94– ??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:SBL

[521] Elizabeth Corcoran. Science and business: Light traffic. Scientific American, 264(3):106–108, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
106.pdf.

Anonymous:1991:SBb

[522] Anonymous. Science and business. Scientific American, 264(3):106–??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1991:PG

[523] Gary Stix. Plane geometry. Scientific American, 264(3):110–111, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-110.pdf.

Erickson:1991:TG

[524] Deborah Erickson. Through the grapevine. Scientific American, 264(3):111–112, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
111.pdf.

Stix:1991:TC

[525] Gary Stix. Thinking cap. Scientific American, 264(3):112–113, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-112.pdf.

Erickson:1991:CCa

[526] Deborah Erickson. Complete combustion. Scientific American, 264(3): 113, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-113.pdf.

Corcoran:1991:AEa

[527] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 264(3):114, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
114.pdf.

Dewdney:1991:MRb

[528] A. K. Dewdney. Mathematical recreations. Scientific American, 264(3):116–118, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n3/pdf/scientificamerican0391-
116.pdf.

Anonymous:1991:MRc

[529] Anonymous. Mathematical recreations. Scientific American, 264(3):116– ??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ferris:1991:B

[530] Timothy Ferris. Books. Scientific American, 264(3):120–125, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-120.pdf.

Anonymous:1991:Bc

[531] Anonymous. Brooks. Scientific American, 264(3):120–??, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mann:1991:E

[532] Jonathan Mann. Essay. Scientific American, 264(3):126, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n3/pdf/scientificamerican0391-126.pdf.

Anonymous:1991:Ld

[533] Anonymous. Letters. Scientific American, 264(4):10–11, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-10.pdf.

Anonymous:1991:Ec

[534] Anonymous. Errata. Scientific American, 264(4):11, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-11.pdf.

Anonymous:1991:YAd

[535] Anonymous. 50 and 100 years ago. Scientific American, 264(4):12, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-12.pdf.

Beardsley:1991:SCC

[536] Tim Beardsley. Science and the citizen: Cold start. Scientific American, 264(4):14–16, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
14.pdf.

Beardsley:1991:TMC

[537] Tim Beardsley. Tracking the missing carbon. Scientific American, 264(4): 16, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-16.pdf.

Erickson:1991:SS

[538] Deborah Erickson. Seeking senescence. Scientific American, 264(4):18, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-18.pdf.

Beardsley:1991:GT

[539] Tim Beardsley. A gentler therapy? Scientific American, 264(4):20, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-20.pdf.

Yam:1991:L

[540] Philip Yam. Into the lattice. Scientific American, 264(4):21, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-21.pdf.

Powell:1991:LN

[541] Corey S. Powell. Looking for nothing. Scientific American, 264(4):22– 23, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-22.pdf.

Geller:1991:OM

[542] Nisa Geller. Old masters. Scientific American, 264(4):23, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-23.pdf.

Holloway:1991:PRD

[543] Marguerite Holloway. A press release on dioxin sets the record wrong. Scientific American, 264(4):24, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
24b.pdf.

Horgan:1991:HL

[544] John Horgan. History lessons. Scientific American, 264(4):24–26, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-24a.pdf.

Anonymous:1991:PPE

[545] Anonymous. Profile: Political engineer. Scientific American, 264(4):26– 35, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-26.pdf.

Hubbard:1991:RCE

[546] Harold M. Hubbard. The real cost of energy. Scientific American, 264(4):36–42, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
36.pdf.

Stephens:1991:SQ

[547] Peter W. Stephens and Alan I. Goldman. The structure of quasicrystals. Scientific American, 264(4):44–53, April 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-44.pdf.

McKnight:1991:MZG

[548] Steven Lanier McKnight. Molecular zippers in gene regulation. Scientific American, 264(4):54–64, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
54.pdf.

Ross:1991:TLHa

[549] Philip E. Ross. Trends in linguistics: Hard words. Scientific American, 264(4):70–??, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1991:WTJ

[550] Anonymous. Why Tarzan and Jane can walk in step with the animals that roam the jungle. Scientific American, 264(4):88–??, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sagdeev:1991:CSW

[551] Roald Z. Sagdeev and Charles F. Kennel. Collisionless shock waves. Scientific American, 264(4):106–113, April 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-106.pdf.

Tiollais:1991:HBV

[552] Pierre Tiollais and Marie-Annick Buendia. Hepatitis B virus. Scientific American, 264(4):116–123, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
116.pdf.

Trotter:1991:SPP

[553] Donald M. Trotter, Jr. Science in pictures: Photochromic and photosensitive glass. Scientific American, 264(4):124–129, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-124.pdf.

Trotter:1991:PPG

[554] Donald M. Trotter, Jr. Photochromic and photosensitive glass. Scientific American, 264(4):124–?? (Intl. ed. 56–61), April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Alexander:1991:HDR

[555] R. McNeill Alexander. How dinosaurs ran. Scientific American, 264(4):130–136, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
130.pdf.

Ross:1991:TLHb

[556] Philip E. Ross. Trends in linguistics: Hard words. Scientific American, 264(4):138–147, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
138.pdf.

Ross:1991:HW

[557] Philip E. Ross. Hard words. Scientific American, 264(4):138–?? (Intl. ed. 70–??), April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1991:SBLb

[558] Deborah Erickson. Science and business: Love and terror. Scientific American, 264(4):148–150, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
148.pdf.

Rennie:1991:PP

[559] John Rennie. A plague of plaques. Scientific American, 264(4):150, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-150.pdf.

Stix:1991:MCS

[560] Gary Stix. Is Morse code signing off? Scientific American, 264(4):152, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-152b.pdf.

Stix:1991:CT

[561] Gary Stix. Call and tell. Scientific American, 264(4):152–153, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-152a.pdf.

Fischetti:1991:GV

[562] Mark Fischetti. Gas vaccine. Scientific American, 264(4):153–154, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-153.pdf.

Beardsley:1991:MR

[563] Tim Beardsley. Mind reader. Scientific American, 264(4):154–156, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-154.pdf.

Wallich:1991:AEc

[564] Paul Wallich and Elizabeth Corcoran. The analytical economist. Scientific American, 264(4):157, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n4/pdf/scientificamerican0491-
157.pdf.

Stewart:1991:MRb

[565] Ian Stewart. Mathematical recreations. Scientific American, 264(4):158– 161, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n4/pdf/scientificamerican0491-158.pdf.

Morrison:1991:Bb

[566] Philip Morrison. Books. Scientific American, 264(4):162–167, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-162.pdf.

Porter:1991:E

[567] Michael E. Porter. Essay. Scientific American, 264(4):168, April 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n4/pdf/scientificamerican0491-168.pdf.

Anonymous:1991:Le

[568] Anonymous. Letters. Scientific American, 264(5):10, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-10.pdf.

Anonymous:1991:Lf

[569] Anonymous. Letters. Scientific American, 264(5):10–??, May 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1991:YAe

[570] Anonymous. 50 and 100 years ago. Scientific American, 264(5):14, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-14.pdf.

Anonymous:1991:YAf

[571] Anonymous. 50 and 100 years ago. Scientific American, 264(5):14– ??, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1991:SCF

[572] John Horgan. Science and the citizen: Up in flames. Scientific American, 264(5):17–24, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
17.pdf.

Anonymous:1991:SCc

[573] Anonymous. Science and the citizen. Scientific American, 264(5):17– ??, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1991:UGD

[574] John Horgan. U.S. gags discussion of war’s environmental effects. Scientific American, 264(5):24, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
24.pdf.

Rennie:1991:PG

[575] John Rennie. Proofreading genes. Scientific American, 264(5):28–32, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-28.pdf.

Holloway:1991:HC

[576] Marguerite Holloway. Hol chan. Scientific American, 264(5):32, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-32.pdf.

Yam:1991:FP

[577] Philip Yam. Freezing point. Scientific American, 264(5):35, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-35.pdf.

Beardsley:1991:PP

[578] Tim Beardsley. Parched policy. Scientific American, 264(5):36–37, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-36.pdf.

Horgan:1991:PRR

[579] John Horgan. Profile: Reluctant revolutionary. Scientific American, 264(5):40–49, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
40.pdf.

Hirschhorn:1991:POR

[580] Norbert Hirschhorn and William B. Greenough III. Progress in oral rehydration therapy. Scientific American, 264(5):50–56, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-50.pdf.

Wilczek:1991:A

[581] Frank Wilczek. Anyons. Scientific American, 264(5):58–65 (Intl. ed. 24– ??), May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-58.pdf.

Moberg:1991:RJD

[582] Carol L. Moberg and Zanvil A. Cohn. René Jules Dubos. Scientific American, 264(5):66–74, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
66.pdf.

Corcoran:1991:TECa

[583] Elizabeth Corcoran. Trends in energy: Cleaning up coal. Scientific American, 264(5):70–??, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mahowald:1991:SR

[584] Misha A. Mahowald and Carver Mead. The silicon retina. Scientific American, 264(5):76–82 (Intl. ed. 40–??), May 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-76.pdf.

Brimhall:1991:GO

[585] George Brimhall. The genesis of ores. Scientific American, 264(5): 84–91 (Intl. ed. 48–??), May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
84.pdf.

Anonymous:1991:TRH

[586] Anonymous. The theory of rigidity, or how to brace yourself against unlikely accidents. Scientific American, 264(5):89–??, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cunningham:1991:GTH

[587] Patrick Cunningham. The genetics of thoroughbred horses. Scientific American, 264(5):92–98, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
92.pdf.

Hill:1991:MEM

[588] Donald R. Hill. Mechanical engineering in the Medieval Near East. Scientific American, 264(5):100–105 (Intl. ed. 64–??), May 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-100.pdf.

Corcoran:1991:CC

[589] Elizabeth Corcoran. Cleaning up coal. Scientific American, 264(5):106– 116, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:TECb

[590] Elizabeth Corcoran. Trends in energy: Cleaning up coal. Scientific American, 264(5):106–116, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
106.pdf.

Stix:1991:SBH

[591] Gary Stix. Science and business: Handful of pain. Scientific American, 264(5):118–120, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
118.pdf.

Anonymous:1991:SBc

[592] Anonymous. Science and business. Scientific American, 264(5):118– ??, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1991:LF

[593] Gary Stix. Light flight. Scientific American, 264(5):120–121, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-120.pdf.

Erickson:1991:CCb

[594] Deborah Erickson. Cadmium charges. Scientific American, 264(5):122, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-122.pdf.

Stix:1991:PP

[595] Gary Stix. Protein probe. Scientific American, 264(5):123, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-123.pdf.

Erickson:1991:OC

[596] Deborah Erickson. Open channels. Scientific American, 264(5):124, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-124.pdf.

Corcoran:1991:AEb

[597] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 264(5):125, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
125.pdf.

Dewdney:1991:MRc

[598] A. K. Dewdney. Mathematical recreations. Scientific American, 264(5):126–128, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n5/pdf/scientificamerican0591-
126.pdf.

Anonymous:1991:MRd

[599] Anonymous. Mathematical recreations. Scientific American, 264(5):126– ??, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1991:Bc

[600] Philip Morrison. Books. Scientific American, 264(5):130–135, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n5/pdf/scientificamerican0591-130.pdf.

Anonymous:1991:Bd

[601] Anonymous. Books. Scientific American, 264(5):130–??, May 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Galbraith:1991:E

[602] John Kenneth Galbraith. Essay. Scientific American, 264(5):136, May 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n5/pdf/scientificamerican0591-136.pdf.

Anonymous:1991:LEa

[603] Anonymous. Letters to the Editors. Scientific American, 264(6):10–11, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-10.pdf.

Anonymous:1991:Ed

[604] Anonymous. Erratum. Scientific American, 264(6):11, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n6/pdf/scientificamerican0691-11.pdf.

Anonymous:1991:YAg

[605] Anonymous. 50 and 100 years ago. Scientific American, 264(6):14, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-14.pdf.

Ross:1991:SCS

[606] Philip E. Ross. Science and the citizen: Science? Nyet. Scientific American, 264(6):17–20, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
17.pdf.

Horgan:1991:DSI

[607] John Horgan. Do screwy images solve a superconductor mystery? Scientific American, 264(6):20, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
20.pdf.

Rennie:1991:HH

[608] John Rennie. Homeobox harvest. Scientific American, 264(6):24, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-24.pdf.

Horgan:1991:A

[609] John Horgan. From the ashes. Scientific American, 264(6):26–27, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-26.pdf.

Wallich:1991:TMA

[610] Paul Wallich. Of two minds about privacy. Scientific American, 264(6): 27, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-27.pdf.

Yam:1991:TR

[611] Philip Yam. There’s the rub. Scientific American, 264(6):30, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n6/pdf/scientificamerican0691-30.pdf.

Ruthen:1991:DV

[612] Russell Ruthen. Double vision. Scientific American, 264(6):32, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-32.pdf.

Beardsley:1991:SGa

[613] Tim Beardsley. Second guessing. Scientific American, 264(6):34–35, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-34.pdf.

Horgan:1991:QIB

[614] John Horgan. Questioning the “it from bit”. Scientific American, 264(6):36–38, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
36.pdf.

Aftergood:1991:NPS

[615] Steven Aftergood, David W. Hafemeister, Oleg F. Prilutsky, Joel R. Primack, and Stanislav N. Rodionov. Nuclear power in space. Scientific American, 264(6):42–47, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
42.pdf.

Courvoisier:1991:Q

[616] Thierry J.-L. Courvoisier and E. Ian Robson. The Quasar 3C 273. Scientific American, 264(6):50–57, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
50.pdf.

Fischetti:1991:SMP

[617] Vincent A. Fischetti. Streptococcal M protein. Scientific American, 264(6):58–65, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
58.pdf.

Toon:1991:PSC

[618] Owen B. Toon and Richard P. Turco. Polar stratospheric clouds and ozone depletion. Scientific American, 264(6):68–74, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n6/pdf/scientificamerican0691-68.pdf.

Powell:1991:TGPa

[619] Corey S. Powell. Trends in geophysics: Peering inward. Scientific American, 264(6):72–81, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McEwen:1991:EBD

[620] Edward McEwen, Robert L. Miller, and Christopher A. Bergman. Early bow design and construction. Scientific American, 264(6):76–82, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-76.pdf.

Berns:1991:LS

[621] Michael W. Berns. Laser surgery. Scientific American, 264(6):84–90, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-84.pdf.

Anonymous:1991:STR

[622] Anonymous. A swift trip over rugged terrain. Scientific American, 264 (6):89–??, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McCrea:1991:ASE

[623] Sir William McCrea. Arthur Stanley Eddington. Scientific American, 264(6):92–97, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
92.pdf.

Powell:1991:TGPb

[624] Corey S. Powell. Trends in geophysics: Peering inward. Scientific American, 264(6):100–111, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
100.pdf.

Powell:1991:PI

[625] Corey S. Powell. Peering inward. Scientific American, 264(6):100–??, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:SBF

[626] Elizabeth Corcoran. Science and business: Flat horizons. Scientific American, 264(6):112–114, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
112.pdf.

Erickson:1991:L

[627] Deborah Erickson. Lighten up. Scientific American, 264(6):116–117, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-116a.pdf.

Koppel:1991:RBB

[628] Tom Koppel. Reading books byte by byte. Scientific American, 264(6):116–117, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
116b.pdf.

Corcoran:1991:SS

[629] Elizabeth Corcoran. Starting from scratch. Scientific American, 264(6): 118, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-118b.pdf.

Erickson:1991:CS

[630] Deborah Erickson. Clot spotter. Scientific American, 264(6):118–119, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-118a.pdf.

Rennie:1991:SC

[631] John Rennie. Selling cells. Scientific American, 264(6):120, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n6/pdf/scientificamerican0691-120.pdf.

Corcoran:1991:SCW

[632] Elizabeth Corcoran. Sorting out chaos on Wall Street. Scientific American, 264(6):121, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
121.pdf.

Stewart:1991:MRSa

[633] Ian Stewart. Mathematical recreations: a swift trip over rugged terrain. Scientific American, 264(6):123–125, June 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v264/
n6/pdf/scientificamerican0691-123.pdf.

Morrison:1991:BBRa

[634] Philip Morrison. Brief: Book reviews. Scientific American, 264(6):127– 131, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v264/n6/pdf/scientificamerican0691-127.pdf.

Kapitza:1991:ESS

[635] Sergei Kapitza. Essay: The state of Soviet science. Scientific American, 264(6):132, June 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v264/n6/pdf/scientificamerican0691-
132.pdf.

Anonymous:1991:Ee

[636] Anonymous. Erratum. Scientific American, 265(1):12, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-12b.pdf.

Anonymous:1991:LEb

[637] Anonymous. Letters to the Editors. Scientific American, 265(1):12, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-12a.pdf.

Anonymous:1991:YAh

[638] Anonymous. 50 and 100 years ago. Scientific American, 265(1):15, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-15.pdf.

Horgan:1991:SCB

[639] John Horgan. Science and the citizen: Burning questions. Scientific American, 265(1):17–24, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
17.pdf.

Anonymous:1991:SCd

[640] Anonymous. Science and the citizen. Scientific American, 265(1):17– ??, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1991:WDK

[641] John Horgan. Why are data from Kuwait being withheld? Scientific American, 265(1):20, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
20.pdf.

Rennie:1991:PM

[642] John Rennie. Proteins 2, malaria 0. Scientific American, 265(1):24– 25, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-24.pdf.

Rennie:1991:BF

[643] John Rennie. Birds of a fever. Scientific American, 265(1):25, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-25.pdf.

Erickson:1991:MT

[644] Deborah Erickson. Molecular trickster. Scientific American, 265(1): 26, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-26.pdf.

Ruthen:1991:QM

[645] Russell Ruthen. Quantum magnets. Scientific American, 265(1):28, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-28.pdf.

Yam:1991:ME

[646] Philip Yam. Math exorcise. Scientific American, 265(1):29, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-29a.pdf.

Powell:1991:LC

[647] Corey S. Powell. Live from off-center. Scientific American, 265(1):29– 32, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-29b.pdf.

Horgan:1991:BBH

[648] John Horgan. Biggest black hole in the Universe? Scientific American, 265(1):32, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
32.pdf.

Holloway:1991:SMC

[649] Marguerite Holloway. A subtle mind contemplates science. Scientific American, 265(1):33–34, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
33.pdf.

Musto:1991:OCM

[650] David F. Musto. Opium, cocaine and marijuana in American history. Scientific American, 265(1):40–47, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
40.pdf.

Stahler:1991:ELS

[651] Steven W. Stahler. The early life of stars. Scientific American, 265(1): 48–55, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-48.pdf.

Zivin:1991:ST

[652] Justin A. Zivin and Dennis W. Choi. Stroke therapy. Scientific American, 265(1):56–63, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
56.pdf.

Robinson:1991:OIS

[653] Glen M. Robinson, David M. Perry, and Richard W. Peterson. Optical interferometry of surfaces. Scientific American, 265(1):66–71, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-66.pdf.

Strobel:1991:BCW

[654] Gary A. Strobel. Biological control of weeds. Scientific American, 265(1): 72–78, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-72.pdf.

Stix:1991:TTAa

[655] Gary Stix. Trends in transportation: Along for the ride? Scientific American, 265(1):76–85, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shimada:1991:CAM

[656] Izumi Shimada and John F. Merkel. Copper-alloy metallurgy in ancient Peru. Scientific American, 265(1):80–86, July 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-80.pdf.

Bellwood:1991:ADO

[657] Peter Bellwood. The Austronesian dispersal and the origin of languages. Scientific American, 265(1):88–93, July 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-88.pdf.

Anonymous:1991:IIF

[658] Anonymous. Insectoids invade a field of robots. Scientific American, 265(1):93–??, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1991:TTAb

[659] Gary Stix. Trends in transportation: Along for the ride? Scientific American, 265(1):94–106, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
94.pdf.

Stix:1991:AR

[660] Gary Stix. Along for the ride? Scientific American, 265(1):94–??, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:SBS

[661] Elizabeth Corcoran. Science and business: Silicon lights up. Scientific American, 265(1):108–109, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
108.pdf.

Corcoran:1991:MR

[662] Elizabeth Corcoran. Memorable revival. Scientific American, 265 (1):109–110, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
109.pdf.

Corcoran:1991:CMP

[663] Elizabeth Corcoran. A chip-making plan to leapfrog Japan. Scientific American, 265(1):110, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
110.pdf.

Corcoran:1991:SB

[664] Elizabeth Corcoran. Sound bytes. Scientific American, 265(1):111– 112, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-111.pdf.

Wallich:1991:FF

[665] Paul Wallich. Fortran forever. Scientific American, 265(1):112, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-112.pdf.

Erickson:1991:DD

[666] Deborah Erickson. 3-D documents. Scientific American, 265(1):113– 116, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-113.pdf.

Ross:1991:MT

[667] Philip E. Ross. A matter of taste. Scientific American, 265(1):116, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-116.pdf.

Wallich:1991:GNP

[668] Paul Wallich and Elizabeth Corcoran. Games that networks play. Scientific American, 265(1):117, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
117.pdf.

Dewdney:1991:MRI

[669] A. K. Dewdney. Mathematical recreations: Insectoids invade a field of robots. Scientific American, 265(1):118–121, July 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n1/pdf/scientificamerican0791-118.pdf.

Morrison:1991:BBRb

[670] Philip Morrison. Books: Book reviews. Scientific American, 265(1):122– 127, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n1/pdf/scientificamerican0791-122.pdf.

Timpane:1991:EPS

[671] John Timpane. Essay: The poetry of science. Scientific American, 265(1):128, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n1/pdf/scientificamerican0791-
128.pdf.

Anonymous:1991:Ef

[672] Anonymous. Erratum. Scientific American, 265(2):10, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n2/pdf/scientificamerican0891-10b.pdf.

Anonymous:1991:LEc

[673] Anonymous. Letters to the Editors. Scientific American, 265(2):10, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-10a.pdf.

Anonymous:1991:YAi

[674] Anonymous. 50 and 100 years ago. Scientific American, 265(2):14, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-14.pdf.

Horgan:1991:SCE

[675] John Horgan. Science and the citizen: Exporting misery. Scientific American, 265(2):16–18, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
16.pdf.

Erickson:1991:DDF

[676] Deborah Erickson. Do DNA fingerprints protect the innocent? Scientific American, 265(2):18, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
18.pdf.

Rennie:1991:SD

[677] John Rennie. A small disturbance. Scientific American, 265(2):20, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-20.pdf.

Ruthen:1991:WW

[678] Russell Ruthen. Waves are waves. … Scientific American, 265(2):21–22, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-21.pdf.

Horgan:1991:CK

[679] John Horgan. Caribbean killer. Scientific American, 265(2):22–23, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-22.pdf.

Powell:1991:NM

[680] Corey S. Powell. The new Moon. Scientific American, 265(2):24, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-24.pdf.

Beardsley:1991:PR

[681] Tim Beardsley. Positive response. Scientific American, 265(2):26, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-26.pdf.

Ross:1991:MBF

[682] Philip E. Ross. Man’s best friend. Scientific American, 265(2):28, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-28.pdf.

Holloway:1991:RIW

[683] Marguerite Holloway. The reward of ideas that are wrong. Scientific American, 265(2):29–30, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
29.pdf.

Kapitza:1991:ATU

[684] Sergei Kapitza. Antiscience trends in the U.S.S.R. Scientific American, 265(2):32–38, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
32.pdf.

Barnes:1991:CGa

[685] Joshua Barnes, Lars Hernquist, and François Schweizer. Colliding galaxies. Scientific American, 265(2):40–47, August 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n2/pdf/scientificamerican0891-40.pdf.

Moyzis:1991:HT

[686] Robert K. Moyzis. The human telomere. Scientific American, 265(2):48– 55, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-48.pdf.

Breuker:1991:SPT

[687] Horst Breuker, Hans Drevermann, Christoph Grab, Alphonse A. Rademakers, and Howard Stone. Science in pictures: Tracking and imaging elementary particles. Scientific American, 265(2):58–63, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n2/pdf/scientificamerican0891-58.pdf.

Breuker:1991:TIE

[688] Horst Breuker, Hans Drevermann, Christoph Grab, Alphonse A. Rademakers, and Howard Stone. Tracking and imaging elementary particles. Scientific American, 265(2):58–?? (Intl. ed. 42–47), August 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schultz:1991:B

[689] Jerome S. Schultz. Biosensors. Scientific American, 265(2):64–69, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-64.pdf.

Evans:1991:B

[690] Howard E. Evans and Kevin M. O’Neill. Beewolves. Scientific American, 265(2):70–76, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
70.pdf.

Beardsley:1991:TBSa

[691] Tim Beardsley. Trends in biology: Smart genes. Scientific American, 265(2):72–??, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kauffman:1991:AA

[692] Stuart A. Kauffman. Antichaos and adaptation. Scientific American, 265(2):78–84, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
78.pdf.

Beardsley:1991:TBSb

[693] Tim Beardsley. Trends in biology: Smart genes. Scientific American, 265(2):86–95, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
86.pdf.

Beardsley:1991:SGb

[694] Tim Beardsley. Smart genes. Scientific American, 265(2):86–??, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1991:WHD

[695] Anonymous. What in heaven is a digital sundial? Scientific American, 265(2):89–??, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:OC

[696] Elizabeth Corcoran. Ordering chaos. Scientific American, 265(2):96, 98, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:SBO

[697] Elizabeth Corcoran. Science and business: Ordering chaos. Scientific American, 265(2):96–98, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
96.pdf.

Erickson:1991:CJP

[698] Deborah Erickson. Can Japan put the brakes on global warming? Scientific American, 265(2):98, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
98.pdf.

Stix:1991:FF

[699] Gary Stix. Firewater fish. Scientific American, 265(2):99–100, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-99.pdf.

Stix:1991:M

[700] Gary Stix. Off the MAP. Scientific American, 265(2):100–101, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-100.pdf.

Erickson:1991:Pa

[701] Deborah Erickson. Painkiller. Scientific American, 265(2):101, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-101a.pdf.

Erickson:1991:BB

[702] Deborah Erickson. Binding bone. Scientific American, 265(2):101–102, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-101b.pdf.

Corcoran:1991:RFC

[703] Elizabeth Corcoran and Paul Wallich. The rise and fall of cities. Scientific American, 265(2):103, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
103.pdf.

Stewart:1991:MRW

[704] Ian Stewart. Mathematical recreations: What in heaven is a digital sundial? Scientific American, 265(2):104–106, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n2/pdf/scientificamerican0891-104.pdf.

Morrison:1991:BR

[705] Philip Morrison. Book reviews. Scientific American, 265(2):107–111, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n2/pdf/scientificamerican0891-107.pdf.

Ziegler:1991:ESB

[706] Charles E. Ziegler. Essay: The second bottom line. Scientific American, 265(2):112, August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n2/pdf/scientificamerican0891-
112.pdf.

Anonymous:1991:LEd

[707] Anonymous. Letters to the Editors. Scientific American, 265(3):12, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-12.pdf.

Anonymous:1991:YAj

[708] Anonymous. 50 and 100 years ago. Scientific American, 265(3):16, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-16.pdf.

Rennie:1991:SCG

[709] John Rennie. Science and the citizen: Graft without corruption. Scientific American, 265(3):18–22, September 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-18.pdf.

Beardsley:1991:EQ

[710] Tim Beardsley. Elusive quarry. Scientific American, 265(3):22–26, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-22.pdf.

Ruthen:1991:GR

[711] Russell Ruthen. Gravity’s rainbow. Scientific American, 265(3):26, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-26.pdf.

Yam:1991:AEM

[712] Philip Yam. AIDS education may breed intolerance. Scientific American, 265(3):30, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
30b.pdf.

Yam:1991:AP

[713] Philip Yam. Has AIDS peaked? Scientific American, 265(3):30–34, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-30a.pdf.

Erickson:1991:MM

[714] Deborah Erickson. Model mice. Scientific American, 265(3):34, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-34.pdf.

Rennie:1991:PVS

[715] John Rennie. Psychic vs. skeptic. Scientific American, 265(3):39–40, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-39.pdf.

Ross:1991:MJ

[716] Philip E. Ross. Mutt & Jeff. Scientific American, 265(3):40–48, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-40.pdf.

Vinton:1991:N

[717] G. C. Vinton. Networks. Scientific American, 265(3):42–51, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ruthen:1991:CA

[718] Russell Ruthen. Clear advantage. Scientific American, 265(3):48–53, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-48.pdf.

Stix:1991:PDH

[719] Gary Stix. Profile: David A. Huffman: Encoding the “neatness” of ones and zeroes. Scientific American, 265(3):54–58, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-54.pdf.

Dertouzos:1991:CCN

[720] Michael L. Dertouzos. Communications, computers and networks. Scientific American, 265(3):62–69, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
62.pdf.

Cerf:1991:N

[721] Vinton G. Cerf. Networks. Scientific American, 265(3):72–81, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-72.pdf.

Tesler:1991:NC

[722] Lawrence G. Tesler. Networked computing in the 1990s. Scientific American, 265(3):86–93, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
86.pdf.

Weiser:1991:CC

[723] Mark Weiser. The computer for the 21st Century. Scientific American, 265(3):94–104, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
94.pdf.

Negroponte:1991:PSC

[724] Nicholas P. Negroponte. Products and services for computer networks. Scientific American, 265(3):106–113, September 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-106.pdf.

Branscomb:1991:CNP

[725] Anne W. Branscomb. Computers, networks and public policy: Common law for the electronic frontier. Scientific American, 265(3):108–??, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gore:1991:CNP

[726] Al Gore. Computers, networks and public policy: Infrastructure for the global village. Scientific American, 265(3):108–??, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kapor:1991:CNP

[727] Mitchell Kapor. Computers, networks and public policy: Civil liberties in cyberspace. Scientific American, 265(3):108–??, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sproull:1991:CNW

[728] Lee Sproull and Sara Kiesler. Computers, networks and work. Scientific American, 265(3):116–123, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
116.pdf.

Malone:1991:CNC

[729] Thomas W. Malone and John F. Rockart. Computers, networks and the corporation. Scientific American, 265(3):128–136, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-128.pdf.

Anonymous:1991:LLS

[730] Anonymous. Leaping into Lyapunov space. Scientific American, 265(3): 130–??, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kay:1991:CNE

[731] Alan C. Kay. Computers, networks and education. Scientific American, 265(3):138–148, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
138.pdf.

Anonymous:1991:CNP

[732] Anonymous. Computers, networks and public policy. Scientific American, 265(3):150, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
150a.pdf.

Gore:1991:IGV

[733] Al Gore. Infrastructure for the Global Village. Scientific American, 265(3):150–153, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
150b.pdf.

Branscomb:1991:CLE

[734] Anne W. Branscomb. Common law for the electronic frontier. Scientific American, 265(3):154–158, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
154.pdf.

Kapor:1991:CLC

[735] Mitchell Kapor. Civil liberties in cyberspace. Scientific American, 265(3):158–164, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
158.pdf.

Stix:1991:SBG

[736] Gary Stix. Science and business: Golden screws. Scientific American, 265(3):166–169, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
166.pdf.

Koppel:1991:LHB

[737] Tom Koppel. Learning how bacteria swim could set new gears in motion. Scientific American, 265(3):168–169, September 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-168.pdf.

Corcoran:1991:TB

[738] Elizabeth Corcoran. True blue. Scientific American, 265(3):171–172, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-171.pdf.

Stix:1991:MW

[739] Gary Stix. Moles at work. Scientific American, 265(3):173, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n3/pdf/scientificamerican0991-173.pdf.

Erickson:1991:II

[740] Deborah Erickson. Industrial immunology. Scientific American, 265(3):174–175, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
174.pdf.

Corcoran:1991:TLT

[741] Elizabeth Corcoran and Paul Wallich. Trading leisure time for more goods? Scientific American, 265(3):176, September 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-176.pdf.

Dewdney:1991:MRL

[742] A. K. Dewdney. Mathematical recreations: Leaping into Lyapunov space. Scientific American, 265(3):178–180, September 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n3/pdf/scientificamerican0991-178.pdf.

Morrison:1991:BBRc

[743] Philip Morrison. Books: Book reviews. Scientific American, 265 (3):182–189, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
182.pdf.

Lucky:1991:EUA

[744] R. W. Lucky. Essay: The urge of an ancient dream. Scientific American, 265(3):190, September 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n3/pdf/scientificamerican0991-
190.pdf.

Anonymous:1991:LEe

[745] Anonymous. Letters to the Editors. Scientific American, 265(4):12, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-12.pdf.

Anonymous:1991:YAk

[746] Anonymous. 50 and 100 years ago. Scientific American, 265(4):16, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-16.pdf.

Rennie:1991:SCH

[747] John Rennie. Science and the citizen: Howls of dismay. Scientific American, 265(4):18–20, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
18.pdf.

Powell:1991:GG

[748] Corey S. Powell. Greenhouse gusher. Scientific American, 265(4):20, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-20.pdf.

Yam:1991:SP

[749] Philip Yam. Sol’s plumage. Scientific American, 265(4):24, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n4/pdf/scientificamerican1091-24.pdf.

Yam:1991:SC

[750] Philip Yam. Spin cycle. Scientific American, 265(4):26, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n4/pdf/scientificamerican1091-26.pdf.

Horgan:1991:DKA

[751] John Horgan. The danger from Kuwait’s air pollution. Scientific American, 265(4):30, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
30b.pdf.

Ross:1991:CL

[752] Philip E. Ross. Crossed lines. Scientific American, 265(4):30–32, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-30a.pdf.

Beardsley:1991:SI

[753] Tim Beardsley. Sharper image. Scientific American, 265(4):32, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-32.pdf.

Curl:1991:Fa

[754] Robert F. Curl and Richard E. Smalley. Fullerenes. Scientific American, 265(4):32–??, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1991:WG

[755] John Horgan. Word games. Scientific American, 265(4):34, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-34.pdf.

Beardsley:1991:SGc

[756] Tim Beardsley. Sharing germplasm. Scientific American, 265(4):36, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-36b.pdf.

Wallich:1991:DD

[757] Paul Wallich. Digital dyslexia. Scientific American, 265(4):36, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-36a.pdf.

Holloway:1991:SDG

[758] Marguerite Holloway. The satisfaction of delayed gratification. Scientific American, 265(4):40–44, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
40.pdf.

Scrimshaw:1991:ID

[759] Nevin S. Scrimshaw. Iron deficiency. Scientific American, 265(4):46–52, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-46.pdf.

Curl:1991:Fb

[760] Robert F. Curl and Richard E. Smalley. Fullerenes. Scientific American, 265(4):54–63, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
54.pdf.

Knoll:1991:EPE

[761] Andrew H. Knoll. End of the proterozoic eon. Scientific American, 265(4):64–73, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
64.pdf.

vonBoehmer:1991:HIS

[762] Harald von Boehmer and Pawel Kisielow. How the immune system learns about self. Scientific American, 265(4):74–81, October 1991.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n4/pdf/scientificamerican1091-74.pdf.

Holloway:1991:TET

[763] Marguerite Holloway and John Horgan. Trends in environmental technology: Soiled shores. Scientific American, 265(4):80–??, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Grant:1991:NSD

[764] Peter R. Grant. Natural selection and Darwin’s finches. Scientific American, 265(4):82–87, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
82.pdf.

Binzel:1991:OA

[765] Richard P. Binzel, M. Antonietta Barucci, and Marcello Fulchignoni. The origins of the asteroids. Scientific American, 265(4):88–94, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n4/pdf/scientificamerican1091-88.pdf.

Orloff:1991:FIB

[766] Jon Orloff. Focused ion beams. Scientific American, 265(4):96–101, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-96.pdf.

Holloway:1991:SSb

[767] Marguerite Holloway. Soiled shores. Scientific American, 265(4):102–116, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-102.pdf.

Holloway:1991:SSc

[768] Marguerite Holloway and John Horgan. Soiled shores. Scientific American, 265(4):102–??, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1991:CWS

[769] Anonymous. Concentration: A winning strategy. Scientific American, 265(4):103–??, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1991:SBL

[770] Gary Stix. Science and business: The light fantastic. Scientific American, 265(4):118–121, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
118.pdf.

Stix:1991:DCC

[771] Gary Stix. Don’t change the channel, rearrange that face. Scientific American, 265(4):120–121, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
120.pdf.

Erickson:1991:SG

[772] Deborah Erickson. Secret garden. Scientific American, 265(4):121–122, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-121.pdf.

Corcoran:1991:BH

[773] Elizabeth Corcoran. Body heat. Scientific American, 265(4):123, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-123.pdf.

Stix:1991:HMD

[774] Gary Stix. How’s my driving? Scientific American, 265(4):124, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n4/pdf/scientificamerican1091-124.pdf.

Wallich:1991:IES

[775] Paul Wallich and Elizabeth Corcoran. An illusion of economic stability? Scientific American, 265(4):125, October 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n4/pdf/scientificamerican1091-125.pdf.

Stewart:1991:MRC

[776] Ian Stewart. Mathematical recreations: Concentration: a winning strategy. Scientific American, 265(4):126–128, October 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n4/pdf/scientificamerican1091-126.pdf.

Morrison:1991:BBRd

[777] Philip Morrison. Books: Book reviews. Scientific American, 265(4):129–133, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
129.pdf.

Eisenberg:1991:EQE

[778] Anne Eisenberg. Essay: Quantum English. Scientific American, 265(4):134, October 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n4/pdf/scientificamerican1091-
134.pdf.

Anonymous:1991:Eg

[779] Anonymous. Errata. Scientific American, 265(5):10, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n5/pdf/scientificamerican1191-10b.pdf.

Anonymous:1991:LEf

[780] Anonymous. Letters to the Editors. Scientific American, 265(5):10, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-10a.pdf.

Anonymous:1991:YAl

[781] Anonymous. 50 and 100 years ago. Scientific American, 265(5):14, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-14.pdf.

Beardsley:1991:SCR

[782] Tim Beardsley. Science and the citizen: Red alert. Scientific American, 265(5):17–18, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
17.pdf.

Yam:1991:AT

[783] Philip Yam. Atomic turn-on. Scientific American, 265(5):20, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-20.pdf.

Rennie:1991:DPT

[784] John Rennie. Dyslexia: a problem of timing. Scientific American, 265(5):26, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
26b.pdf.

Rennie:1991:SS

[785] John Rennie. Are species specious? Scientific American, 265(5):26, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-26a.pdf.

Beardsley:1991:MM

[786] Tim Beardsley. Murder mystery. Scientific American, 265(5):29, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-29.pdf.

Beardsley:1991:BB

[787] Tim Beardsley. Big Bang. Scientific American, 265(5):30, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-30.pdf.

Ruthen:1991:QP

[788] Russell Ruthen. Quantum pinball. Scientific American, 265(5):36–38, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-36.pdf.

Ross:1991:EE

[789] Philip E. Ross. Endless endgame? Scientific American, 265(5):38, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-38.pdf.

Selkoe:1991:VPA

[790] Dennis J. Selkoe. Void protein and Alzheimer’s disease. Scientific American, 265(5):40–??, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1991:PPS

[791] John Horgan. The pied piper of superstrings. Scientific American, 265(5):42–47, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
42.pdf.

Zimring:1991:FVP

[792] Franklin E. Zimring. Firearms, violence and public policy. Scientific American, 265(5):48–54, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
48.pdf.

Gutbrod:1991:NES

[793] Hans Gutbrod and Horst Stöcker. The nuclear equation of state. Scientific American, 265(5):58–66, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
58.pdf.

Selkoe:1991:APA

[794] Dennis J. Selkoe. Amyloid protein and Alzheimer’s disease. Scientific American, 265(5):68–78, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
68.pdf.

Lam:1991:SPC

[795] Dominic Man-Kit Lam and Bryant W. Rossiter. Science in pictures: Chromoskedasic painting. Scientific American, 265(5):80–85, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-80.pdf.

Lam:1991:CP

[796] Dominic Man-Kit Lam and Bryant W. Rossiter. Chromoskedasic painting. Scientific American, 265(5):80–?? (Intl. ed. 48–??), November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1991:TAMa

[797] Corey S. Powell. Trends in astronomy: Mirroring the cosmos. Scientific American, 265(5):80–??, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jewell:1991:M

[798] Jack L. Jewell, James P. Harbison, and Axel Scherer. Microlasers. Scientific American, 265(5):86–94, November 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n5/pdf/scientificamerican1191-86.pdf.

Green:1991:CCT

[799] Howard Green. Cultured cells for the treatment of disease. Scientific American, 265(5):96–102, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
96.pdf.

Anonymous:1991:PCP

[800] Anonymous. Painting in color without pigments. Scientific American, 265(5):98–??, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cavalli-Sforza:1991:GPL

[801] Luigi Luca Cavalli-Sforza. Genes, peoples and languages. Scientific American, 265(5):104–110, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
104.pdf.

Powell:1991:TAMb

[802] Corey S. Powell. Trends in astronomy: Mirroring the Cosmos. Scientific American, 265(5):112–123, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
112.pdf.

Powell:1991:MC

[803] Corey S. Powell. Mirroring the cosmos. Scientific American, 265(5): 112–??, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1991:SBB

[804] Deborah Erickson. Science and business: Brain, food. Scientific American, 265(5):124–126, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
124.pdf.

Erickson:1991:SJ

[805] Deborah Erickson. Sustainable jobs. Scientific American, 265 (5):127–128, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
127.pdf.

Erickson:1991:S

[806] Deborah Erickson. Skinside out. Scientific American, 265(5):128–130, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-128.pdf.

Erickson:1991:Pb

[807] Deborah Erickson. Plastic for the ’90s. Scientific American, 265 (5):130–131, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
130.pdf.

Stix:1991:HSS

[808] Gary Stix. Human spec sheet. Scientific American, 265(5):132–133, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-132.pdf.

Wallich:1991:NAD

[809] Paul Wallich. No accounting for disaster. Scientific American, 265(5):134, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
134.pdf.

Lam:1991:ASP

[810] Dominic Man-Kit Lam and Alexandra J. Baran. The amateur scientist: Painting in color without pigments. Scientific American, 265(5):136–137, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n5/pdf/scientificamerican1191-136.pdf.

Morrison:1991:BBRe

[811] Philip Morrison. Books: Book reviews. Scientific American, 265 (5):139–143, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
139.pdf.

Griffin:1991:EAT

[812] Donald R. Griffin. Essay: Animal thinking. Scientific American, 265(5):144, November 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n5/pdf/scientificamerican1191-
144.pdf.

Anonymous:1991:LEg

[813] Anonymous. Letters to the Editors. Scientific American, 265(6):10, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-10.pdf.

Anonymous:1991:YAm

[814] Anonymous. 50 and 100 years ago. Scientific American, 265(6):12, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-12.pdf.

Holloway:1991:SCH

[815] Marguerite Holloway. Science and the citizen: High and dry. Scientific American, 265(6):16–20, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
16.pdf.

Ross:1991:Bb

[816] Philip E. Ross. Buckytubes. Scientific American, 265(6):24, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-24.pdf.

Rennie:1991:INP

[817] John Rennie. The Ig Nobel Prizes. Scientific American, 265(6):26, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-26.pdf.

Beardsley:1991:FF

[818] Tim Beardsley. Fatal flaw. Scientific American, 265(6):28–29, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-28.pdf.

Powell:1991:SB

[819] Corey S. Powell. Star bursts. Scientific American, 265(6):32, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-32.pdf.

Koch:1991:PP

[820] Eddie Koch. People preserves. Scientific American, 265(6):54, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-54.pdf.

Beardsley:1991:CR

[821] Tim Beardsley. Cross reaction. Scientific American, 265(6):56–57, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-56.pdf.

Stix:1991:SW

[822] Gary Stix. Smaller world. Scientific American, 265(6):58–59, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-58.pdf.

Horgan:1991:MWL

[823] John Horgan. The man who loves insects. Scientific American, 265(6):60–64, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
60.pdf.

Bassuk:1991:HF

[824] Ellen L. Bassuk. Homeless families. Scientific American, 265(6):66–74, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-66.pdf.

Halliwell:1991:QCC

[825] Jonathan J. Halliwell. Quantum cosmology and the creation of the Universe. Scientific American, 265(6):76–85, December 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n6/pdf/scientificamerican1291-76.pdf.

Wallich:1991:TAIa

[826] P. Wallich. Trends in artificial intelligence: Silicon babies. Scientific American, 265(6):82–91, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Golde:1991:SC

[827] David W. Golde. The stem cell. Scientific American, 265(6):86–93, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-86.pdf.

Anthony:1991:OHR

[828] David Anthony, Dimitri Y. Telegin, and Dorcas Brown. The origin of horseback riding. Scientific American, 265(6):94–100, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n6/pdf/scientificamerican1291-94.pdf.

Anonymous:1991:STI

[829] Anonymous. A short trek to infinity. Scientific American, 265(6):100–??, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dostrovsky:1991:CFS

[830] Israel Dostrovsky. Chemical fuels from the Sun. Scientific American, 265(6):102–107, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
102.pdf.

Seymour:1991:BT

[831] Roger S. Seymour. The brush turkey. Scientific American, 265 (6):108–114, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
108.pdf.

Dalmedico:1991:SG

[832] Amy Dahan Dalmédico. Sophie Germain. Scientific American, 265(6):116–122, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
116.pdf.

Wallich:1991:TAIb

[833] Paul Wallich. Trends in artificial intelligence: Silicon babies. Scientific American, 265(6):124–134, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
124.pdf.

Wallich:1991:SBb

[834] Paul Wallich. Silicon babies. Scientific American, 265(6):124–??, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1991:SBR

[835] Elizabeth Corcoran. Science and business: Rethinking research. Scientific American, 265(6):136–139, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
136.pdf.

Corcoran:1991:TM

[836] Elizabeth Corcoran. Thinking of machines. Scientific American, 265(6):140–141, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
140.pdf.

Erickson:1991:BS

[837] Deborah Erickson. Beyond sympathy. Scientific American, 265(6):141, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-141.pdf.

Stix:1991:SB

[838] Gary Stix. Seeing is believing. Scientific American, 265(6):142, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v265/n6/pdf/scientificamerican1291-142.pdf.

Corcoran:1991:AEc

[839] Elizabeth Corcoran and Paul Wallich. The analytical economist. Scientific American, 265(6):143, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
143.pdf.

Stewart:1991:MRSb

[840] Ian Stewart. Mathematical recreations: a short trek to infinity. Scientific American, 265(6):144–147, December 1991. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n6/pdf/scientificamerican1291-144.pdf.

Morrison:1991:BSB

[841] Philip Morrison and Phylis Morrison. Books: Science books for young people. Scientific American, 265(6):148–153, December 1991.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n6/pdf/scientificamerican1291-148.pdf.

Anonymous:1991:AIA

[842] Anonymous. Annual index: Annual index 1991. Scientific American, 265(6):154–157, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v265/n6/pdf/scientificamerican1291-
154.pdf.

Rosenberg:1991:EMW

[843] Nathan Rosenberg. Essay: Marx wasn’t all wrong: Essay: Marx wasn’t all wrong. Scientific American, 265(6):158, December 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v265/
n6/pdf/scientificamerican1291-158.pdf.

Anonymous:1992:LEa

[844] Anonymous. Letters to the Editors. Scientific American, 266(1):9–10, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-9.pdf.

Anonymous:1992:Ea

[845] Anonymous. Errata. Scientific American, 266(1):10, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n1/pdf/scientificamerican0192-10.pdf.

Anonymous:1992:YAa

[846] Anonymous. 50 and 100 years ago. Scientific American, 266(1):14, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-14.pdf.

Powell:1992:SCV

[847] Corey S. Powell. Science and the citizen: Venus revealed. Scientific American, 266(1):16–17, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
16.pdf.

Bazzaz:1992:PLCa

[848] Fakhri A. Bazzaz and Eric D. Fajer. Plant life in a CO_{2} rich world. Scientific American, 266(1):18–??, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1992:AI

[849] John Horgan. After the inferno. Scientific American, 266(1):20, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-20b.pdf.

Powell:1992:RR

[850] Corey S. Powell. Rocky rendezvous. Scientific American, 266(1):20, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-20a.pdf.

Beardsley:1992:RC

[851] Tim Beardsley. Report card. Scientific American, 266(1):22–23, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-22.pdf.

Holloway:1992:UC

[852] Marguerite Holloway. Under construction. Scientific American, 266(1):25–26, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
25.pdf.

Yam:1992:BG

[853] Philip Yam. Bonus game. Scientific American, 266(1):27–28, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-27.pdf.

Corcoran:1992:TLT

[854] Elizabeth Corcoran. Tacky lasers are the tiniest yet. Scientific American, 266(1):28, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
28.pdf.

Dewdney:1992:TT

[855] A. K. Dewdney. Turing test. Scientific American, 266(1):30–32, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-30.pdf.

Beardsley:1992:PP

[856] Tim Beardsley. Population pressure. Scientific American, 266(1):32, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-32b.pdf.

Rennie:1992:IJ

[857] John Rennie. Inside job. Scientific American, 266(1):32, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n1/pdf/scientificamerican0192-32a.pdf.

Beardsley:1992:TH

[858] Tim Beardsley. A troubled homecoming. Scientific American, 266(1):33– 36, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-33.pdf.

Bazzaz:1992:PLCb

[859] Fakhri A. Bazzaz and Eric D. Fajer. Plant life in a CO_{2}-rich world. Scientific American, 266(1):68–74, January 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n1/pdf/scientificamerican0192-68.pdf.

Bazzaz:1992:PLR

[860] Fakhri A. Bazzaz and Eric D. Fajer. Plant life in a CO_{2}-Rich world. Scientific American, 266(1):68–??, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gutzwiller:1992:QC

[861] Martin C. Gutzwiller. Quantum chaos. Scientific American, 266(1):78– 84, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-78.pdf.

Lienhard:1992:HCA

[862] Gustav E. Lienhard, Jan W. Slot, David E. James, and Mike M. Mueckler. How cells absorb glucose. Scientific American, 266(1):86–91, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-86.pdf.

Cannizzo:1992:ADI

[863] John K. Cannizzo and Ronald H. Kaitchuck. Accretion disks in interacting binary stars. Scientific American, 266(1):92–99, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n1/pdf/scientificamerican0192-92.pdf.

Lohmann:1992:HST

[864] Kenneth J. Lohmann. How sea turtles navigate. Scientific American, 266(1):100–106, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
100.pdf.

Rennie:1992:TPLa

[865] John Rennie. Trends in parasitology: Living together. Scientific American, 266(1):104–??, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ferguson:1992:TW

[866] R. Brian Ferguson. Tribal warfare. Scientific American, 266(1):108–113, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-108.pdf.

Desurvire:1992:LCF

[867] Emmanuel Desurvire. Lightwave communications: The fifth generation. Scientific American, 266(1):114–121, January 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n1/pdf/scientificamerican0192-114.pdf.

Smith:1992:HGC

[868] Douglas Smith. How to generate chaos at home. Scientific American, 266 (1):121–??, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1992:TPLb

[869] John Rennie. Trends in parasitology: Living together. Scientific American, 266(1):122–133, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
122.pdf.

Rennie:1992:LT

[870] John Rennie. Living together. Scientific American, 266(1):122–??, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1992:SBGa

[871] Deborah Erickson. Science and business: Growing strong. Scientific American, 266(1):134–142, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
134.pdf.

Stix:1992:EI

[872] Gary Stix. Explosive images. Scientific American, 266(1):137, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-137.pdf.

Beardsley:1992:CE

[873] Tim Beardsley. Current event. Scientific American, 266(1):138–139, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-138.pdf.

Corcoran:1992:GM

[874] Elizabeth Corcoran. Green machine. Scientific American, 266 (1):140–141, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
140.pdf.

Stix:1992:NTP

[875] Gary Stix. No tipping, please. Scientific American, 266(1):141, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n1/pdf/scientificamerican0192-141.pdf.

Corcoran:1992:CMA

[876] Elizabeth Corcoran and Paul Wallich. Coping with math anxiety. Scientific American, 266(1):142, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
142.pdf.

Smith:1992:ASH

[877] Douglas Smith. The amateur scientist: How to generate chaos at home. Scientific American, 266(1):144–146, January 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n1/pdf/scientificamerican0192-144.pdf.

Morrison:1992:BBRa

[878] Philip Morrison. Books: Book reviews. Scientific American, 266(1):147–150, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
147.pdf.

Lach:1992:EIC

[879] Michael C. Lach. Essay: An inner-city education. Scientific American, 266(1):151, January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n1/pdf/scientificamerican0192-
151.pdf.

Anonymous:1992:Eb

[880] Anonymous. Errata. Scientific American, 266(2):10, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n2/pdf/scientificamerican0292-10b.pdf.

Anonymous:1992:LEb

[881] Anonymous. Letters to the Editors. Scientific American, 266(2):10, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-10a.pdf.

Anonymous:1992:YAb

[882] Anonymous. 50 and 100 years ago. Scientific American, 266(2):12, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-12.pdf.

Horgan:1992:SCE

[883] John Horgan. Science and the citizen: Early arrivals. Scientific American, 266(2):17–20, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
17.pdf.

Beardsley:1992:NH

[884] Tim Beardsley. Night heat. Scientific American, 266(2):21, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-21a.pdf.

Powell:1992:CLI

[885] Corey S. Powell. A confusing Little Ice Age. Scientific American, 266(2):21, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
21b.pdf.

Ross:1992:NWW

[886] Philip E. Ross. New whoof in whorf. Scientific American, 266(2):24–26, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-24.pdf.

Horgan:1992:EC

[887] John Horgan. The end of clumpiness? Scientific American, 266(2):26, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-26.pdf.

Beardsley:1992:SB

[888] Tim Beardsley. Shear bliss. Scientific American, 266(2):27–29, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-27.pdf.

Holloway:1992:AP

[889] Marguerite Holloway. Abyssal proposal. Scientific American, 266 (2):30–31, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
30.pdf.

Rennie:1992:LFL

[890] John Rennie. Life in the fast lane. Scientific American, 266(2):31, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-31.pdf.

Horgan:1992:MN

[891] John Horgan. The Mephistopheles of neurobiology. Scientific American, 266(2):32–35, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
32.pdf.

Caplan:1992:IRF

[892] Nathan Caplan, Marcella H. Choy, and John K. Whitmore. Indochinese refugee families and academic achievement. Scientific American, 266(2):36–42, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
36.pdf.

Fryer:1992:MVM

[893] Patricia Fryer. Mud volcanoes of the Marianas. Scientific American, 266(2):46–52, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
46.pdf.

Liotta:1992:CCI

[894] Lance A. Liotta. Cancer cell invasion and metastasis. Scientific American, 266(2):54–63, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
54.pdf.

Storch:1992:SPM

[895] Gerhard Storch. Science in pictures: The mammals of Island Europe. Scientific American, 266(2):64–69 (Intl. ed. 42–??), February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n2/pdf/scientificamerican0292-64.pdf.

Chu:1992:LTN

[896] Steven Chu. Laser trapping of neutral particles. Scientific American, 266(2):70–76, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
70.pdf.

Corcoran:1992:TCEa

[897] Elizabeth Corcoran. Trends in consumer electronics: Picture perfect. Scientific American, 266(2):72–??, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Harbottle:1992:TPC

[898] Garman Harbottle and Phil C. Weigand. Turquoise in pre-Columbian America. Scientific American, 266(2):78–85, February 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n2/pdf/scientificamerican0292-78.pdf.

Paul:1992:NVT

[899] Diane B. Paul and Costas B. Krimbas. Nikolai V. Timoféeff-Ressovsky. Scientific American, 266(2):86–92, February 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n2/pdf/scientificamerican0292-86.pdf.

Stewart:1992:KN

[900] Ian Stewart. The kissing number. Scientific American, 266(2):90–??, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:TCEb

[901] Elizabeth Corcoran. Trends in consumer electronics: Picture perfect. Scientific American, 266(2):94–103, February 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n2/pdf/scientificamerican0292-94.pdf.

Corcoran:1992:PP

[902] Elizabeth Corcoran. Picture perfect. Scientific American, 266(2):94–??, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1992:SBR

[903] Gary Stix. Science and business: Riding on air. Scientific American, 266(2):104–105, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
104.pdf.

Erickson:1992:BR

[904] Deborah Erickson. A better red. Scientific American, 266(2):106, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-106.pdf.

Beardsley:1992:MA

[905] Tim Beardsley. Making antisense. Scientific American, 266(2):107–108, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-107.pdf.

Erickson:1992:W

[906] Deborah Erickson. It’s a wrap. Scientific American, 266(2):108–109, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-108.pdf.

Stix:1992:FRa

[907] Gary Stix. Further references. Scientific American, 266(2):110, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n2/pdf/scientificamerican0292-110.pdf.

Wallich:1992:DDB

[908] Paul Wallich and Elizabeth Corcoran. The discreet disappearance of the bourgeoisie. Scientific American, 266(2):111, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n2/pdf/scientificamerican0292-111.pdf.

Stewart:1992:MRK

[909] Ian Stewart. Mathematical recreations: The kissing number. Scientific American, 266(2):112–115, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
112.pdf.

Morrison:1992:BBRb

[910] Philip Morrison. Brief: Book reviews. Scientific American, 266 (2):116–119, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
116.pdf.

Burbidge:1992:EWO

[911] Geoffrey Burbidge. Essay: Why only one Big Bang? Scientific American, 266(2):120, February 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n2/pdf/scientificamerican0292-
120.pdf.

Anonymous:1992:Ec

[912] Anonymous. Errata. Scientific American, 266(3):8, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n3/pdf/scientificamerican0392-8b.pdf.

Anonymous:1992:LEc

[913] Anonymous. Letters to the Editors. Scientific American, 266(3):8, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-8a.pdf.

Anonymous:1992:YAc

[914] Anonymous. 50 and 100 years ago. Scientific American, 266(3):10, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-10.pdf.

Holloway:1992:SCR

[915] Marguerite Holloway and Philip Yam. Science and the citizen: Reflecting differences. Scientific American, 266(3):13–18, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n3/pdf/scientificamerican0392-13.pdf.

Corcoran:1992:THS

[916] Elizabeth Corcoran. Tuning in hot signals. Scientific American, 266(3): 16, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-16.pdf.

Beardsley:1992:RH

[917] Tim Beardsley. Relative hunger. Scientific American, 266(3):18, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-18.pdf.

Powell:1992:UP

[918] Corey S. Powell. Unlikely places. Scientific American, 266(3):22, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-22.pdf.

Wallich:1992:GTP

[919] Paul Wallich. Going through proper channels. Scientific American, 266(3):26, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
26b.pdf.

Beardsley:1992:PF

[920] Tim Beardsley. Pointing fingers. Scientific American, 266(3):26–27, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-26a.pdf.

Rennie:1992:TH

[921] John Rennie. Trojan horse. Scientific American, 266(3):27–28, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-27.pdf.

Horgan:1992:VD

[922] John Horgan. Volcanic disruption. Scientific American, 266(3):28–29, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-28.pdf.

Beardsley:1992:AOG

[923] Tim Beardsley. Add ozone to the global warming equation. Scientific American, 266(3):29, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
29.pdf.

Horgan:1992:LOP

[924] John Horgan. The lonely Odysseus of particle physics. Scientific American, 266(3):30–32, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
30.pdf.

Sachs:1992:BME

[925] Jeffrey Sachs. Building a market economy in Poland. Scientific American, 266(3):34–40, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
34.pdf.

Neher:1992:PCT

[926] Erwin Neher and Bert Sakmann. The patch clamp technique. Scientific American, 266(3):44–51, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
44.pdf.

Spergel:1992:TCS

[927] David N. Spergel and Neil G. Turok. Textures and cosmic structure. Scientific American, 266(3):52–59, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
52.pdf.

Aihara:1992:WAC

[928] Jun ichi Aihara. Why aromatic compounds are stable. Scientific American, 266(3):62–68, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
62.pdf.

Vollrath:1992:SWS

[929] Fritz Vollrath. Spider webs and silks. Scientific American, 266(3):70–76, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-70.pdf.

Ruthen:1992:TACa

[930] Russell Ruthen. Trends in astrophysics: Catching the wave. Scientific American, 266(3):72–??, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Silverman:1992:IVC

[931] Jerry Silverman, Jonathan M. Mooney, and Freeman D. Shepherd. Infrared video cameras. Scientific American, 266(3):78–83, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n3/pdf/scientificamerican0392-78.pdf.

Taylor:1992:GC

[932] Timothy Taylor. The Gundestrup cauldron. Scientific American, 266(3):84–89, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
84.pdf.

Ruthen:1992:TACb

[933] Russell Ruthen. Trends in astrophysics: Catching the wave. Scientific American, 266(3):90–99, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
90.pdf.

Mooney:1992:SI

[934] Donald G. Mooney. Seeing infrared. Scientific American, 266(3):90–??, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ruthen:1992:CW

[935] Russell Ruthen. Catching the wave. Scientific American, 266(3):90–??, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:SBH

[936] Elizabeth Corcoran. Science and business: Holey silicon. Scientific American, 266(3):102–104, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
102.pdf.

Beardsley:1992:EF

[937] Tim Beardsley. Executive fix. Scientific American, 266(3):104–106, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-104.pdf.

Beardsley:1992:VI

[938] Tim Beardsley. Vested interests. Scientific American, 266(3):106, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-106.pdf.

Stix:1992:DR

[939] Gary Stix. Drag race. Scientific American, 266(3):107–108, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n3/pdf/scientificamerican0392-107.pdf.

Stix:1992:OD

[940] Gary Stix. Objective data. Scientific American, 266(3):108, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n3/pdf/scientificamerican0392-108.pdf.

Rennie:1992:GB

[941] John Rennie. Getting out the bugs. Scientific American, 266(3):109, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-109.pdf.

Corcoran:1992:DD

[942] Elizabeth Corcoran. Disarming developments. Scientific American, 266(3):110, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
110.pdf.

Mooney:1992:ASS

[943] Donald G. Mooney. The amateur scientist: Seeing infrared. Scientific American, 266(3):112–113, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n3/pdf/scientificamerican0392-
112.pdf.

Morrison:1992:BRa

[944] Philip Morrison. Book reviews. Scientific American, 266(3):114–117, March 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n3/pdf/scientificamerican0392-114.pdf.

Mayer:1992:EWA

[945] Kenneth H. Mayer and Charles C. J. Carpenter. Essay: Women and AIDS. Scientific American, 266(3):118–119, March 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n3/pdf/scientificamerican0392-118.pdf.

Anonymous:1992:Ed

[946] Anonymous. Errata. Scientific American, 266(4):10, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-10b.pdf.

Anonymous:1992:LEd

[947] Anonymous. Letters to the Editors. Scientific American, 266(4):10, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-10a.pdf.

Anonymous:1992:YAd

[948] Anonymous. 50 and 100 years ago. Scientific American, 266(4):14, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-14.pdf.

Beardsley:1992:SCBa

[949] Tim Beardsley. Science and the citizen: Brain drain. Scientific American, 266(4):17–20, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
17.pdf.

Holloway:1992:PS

[950] Marguerite Holloway. The plankton stalkers. Scientific American, 266(4): 20, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-20.pdf.

Rennie:1992:FCC

[951] John Rennie. First-class culprit. Scientific American, 266(4):24–25, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-24.pdf.

Rennie:1992:CS

[952] John Rennie. A census of stranglers. Scientific American, 266(4):25, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-25.pdf.

Ruthen:1992:VV

[953] Russell Ruthen. Vexing vortices. Scientific American, 266(4):28–29, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-28.pdf.

Horgan:1992:DD

[954] John Horgan. D2 or not D2. Scientific American, 266(4):29–32, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-29.pdf.

Powel:1992:SG

[955] Corey S. Powel. A second glance. Scientific American, 266(4):36, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-36.pdf.

Holloway:1992:FW

[956] Marguerite Holloway. Fire in water. Scientific American, 266(4):37–40, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-37.pdf.

Anonymous:1992:DWD

[957] Anonymous. Debate: Where did modern humans originate? Scientific American, 266(4):66–67, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
66.pdf.

Thorne:1992:MEHa

[958] Alan G. Thorne and Milford H. Wolpoff. The multiregional evolution of humans. Scientific American, 266(4):66–?? (Intl. ed. 28–??), April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wilson:1992:RAG

[959] Allan C. Wilson and Rebecca L. Cann. The recent African genesis of humans. Scientific American, 266(4):68–73, April 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-68.pdf.

Thorne:1992:MEHb

[960] Alan G. Thorne and Milford H. Wolpoff. The multiregional evolution of humans. Scientific American, 266(4):76–83, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-76.pdf.

Murphy:1992:MBS

[961] J. Brendan Murphy and R. Damian Nance. Mountain belts and the supercontinent cycle. Scientific American, 266(4):84–91, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-84.pdf.

Johnson:1992:SHD

[962] Howard M. Johnson, Jeffry K. Russell, and Carol H. Pontzer. Superantigens in human disease. Scientific American, 266(4):92–101, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-92.pdf.

Erickson:1992:TGHa

[963] Deborah Erickson. Trends in genetics: Hacking the genome. Scientific American, 266(4):98–??, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Conn:1992:ITE

[964] Robert W. Conn, Valery A. Chuyanov, Nobuyuki Inoue, and Donald R. Sweetman. The International Thermonuclear Experimental Reactor. Scientific American, 266(4):102–110 (Intl. ed. 74–80), April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-102.pdf.

Thomas:1992:SAC

[965] Sir John Meurig Thomas. Solid acid catalysts. Scientific American, 266 (4):112–118 (Intl. ed. 82–??), April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
112.pdf.

Stewart:1992:APL

[966] Ian Stewart. All paths lead away from Rome. Scientific American, 266 (4):113(??)–??, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Melzack:1992:PL

[967] Ronald Melzack. Phantom limbs. Scientific American, 266(4):120–125, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-120.pdf.

Erickson:1992:TGHb

[968] Deborah Erickson. Trends in genetics: Hacking the genome. Scientific American, 266(4):128–137, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
128.pdf.

Erickson:1992:HG

[969] Deborah Erickson. Hacking the genome. Scientific American, 266(4): 128–??, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:SBC

[970] Elizabeth Corcoran. Science and business: Charlotte’s patent. Scientific American, 266(4):138–140, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
138.pdf.

Stix:1992:DA

[971] Gary Stix. Desktop artisans. Scientific American, 266(4):141–142, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-141.pdf.

Corcoran:1992:APO

[972] Elizabeth Corcoran. Avoiding the potholes on optical highways. Scientific American, 266(4):143, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
143b.pdf.

Beardsley:1992:WN

[973] Tim Beardsley. Waste not. … Scientific American, 266(4):143–144, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-143a.pdf.

Stix:1992:BP

[974] Gary Stix. Ballpark physics. Scientific American, 266(4):145–146, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-145.pdf.

Stix:1992:HLG

[975] Gary Stix. Helping the little guy buy a computer. Scientific American, 266(4):146, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
146.pdf.

Wallich:1992:CBC

[976] Paul Wallich and Elizabeth Corcoran. Compensation beyond the call of duty. Scientific American, 266(4):147, April 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-147.pdf.

Stewart:1992:MRA

[977] Ian Stewart. Mathematical recreations: All paths lead away from Rome. Scientific American, 266(4):150–152, April 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n4/pdf/scientificamerican0492-150.pdf.

Morrison:1992:BBRc

[978] Philip Morrison. Brief: Book reviews. Scientific American, 266(4):154– 159, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n4/pdf/scientificamerican0492-154.pdf.

Benedick:1992:ECD

[979] Richard Elliot Benedick. Essay: a case of déjà vu. Scientific American, 266(4):160, April 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n4/pdf/scientificamerican0492-
160.pdf.

Anonymous:1992:Ee

[980] Anonymous. Errata. Scientific American, 266(5):8, May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-8b.pdf.

Anonymous:1992:LEe

[981] Anonymous. Letters to the Editors. Scientific American, 266(5):8, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-8a.pdf.

Anonymous:1992:YAe

[982] Anonymous. 50 and 100 years ago. Scientific American, 266(5):12, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-12.pdf.

Beardsley:1992:SCBb

[983] Tim Beardsley. Science and the citizen: Better care, less care. Scientific American, 266(5):17–20, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
17.pdf.

Wallich:1992:IHJ

[984] Paul Wallich. Is it history or just E-mail? Scientific American, 266(5): 20, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-20.pdf.

Yam:1992:FL

[985] Philip Yam. Forbidden light. Scientific American, 266(5):24–28, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-24.pdf.

Powell:1992:BY

[986] Corey S. Powell. Born yesterday. Scientific American, 266(5):28–30, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-28.pdf.

Horgan:1992:LTT

[987] John Horgan. Life in a test tube? Scientific American, 266(5):30, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-30.pdf.

Rennie:1992:MMa

[988] John Rennie. Mutable mutation. Scientific American, 266(5):34, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-34.pdf.

Horgan:1992:TT

[989] John Horgan. Trying transmutation. Scientific American, 266(5):36– 53, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-36.pdf.

Horgan:1992:JCF

[990] John Horgan. Japan, cold fusion and Lyndon LaRouche. Scientific American, 266(5):53, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
53.pdf.

Holloway:1992:BB

[991] Marguerite Holloway. The builder of bridges. Scientific American, 266(5): 56–57, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-56.pdf.

Anderson:1992:UAP

[992] Roy M. Anderson and Robert M. May. Understanding the AIDS pandemic. Scientific American, 266(5):58–66, May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-58.pdf.

Snyder:1992:BRN

[993] Solomon H. Snyder and David S. Bredt. Biological roles of nitric oxide. Scientific American, 266(5):68–77, May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-68.pdf.

Ross:1992:TMAa

[994] Philip E. Ross. Trends in molecular archaeology: Eloquent remains. Scientific American, 266(5):72–??, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Soker:1992:PN

[995] Noam Soker. Planetary nebulae. Scientific American, 266(5):78–88, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-78.pdf.

Ramachandran:1992:SPB

[996] Vilayanur S. Ramachandran. Science in pictures: Blind spots. Scientific American, 266(5):86–94, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
86.pdf.

Ramachandran:1992:BS

[997] Vilayanur S. Ramachandran. Blind spots. Scientific American, 266(5): 86–?? (Intl. ed. 44–49), May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kaler:1992:WDS

[998] James B. Kaler. Watching the death of a star. Scientific American, 266(5):89–??, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Veldkamp:1992:BO

[999] Wilfrid B. Veldkamp and Thomas J. McHugh. Binary optics. Scientific American, 266(5):92–97 (Intl. ed. 50–55), May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-92.pdf.

Terborgh:1992:WAS

[1000] John Terborgh. Why American songbirds are vanishing. Scientific American, 266(5):98–104, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
98.pdf.

Cassidy:1992:HUQ

[1001] David C. Cassidy. Heisenberg, uncertainty and the quantum revolution. Scientific American, 266(5):106–112 (Intl. ed. 64–??), May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-106.pdf.

Ross:1992:TMAb

[1002] Philip E. Ross. Trends in molecular archaeology: Eloquent remains. Scientific American, 266(5):114–125, May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-114.pdf.

Ross:1992:ER

[1003] Philip E. Ross. Eloquent remains. Scientific American, 266(5):114– ??, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1992:SBE

[1004] Gary Stix. Science and business: Electric car pool. Scientific American, 266(5):126–127, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
126.pdf.

Stix:1992:FRb

[1005] Gary Stix. Frothing a raindrop. Scientific American, 266(5):128, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-128b.pdf.

Erickson:1992:RL

[1006] Deborah Erickson. Relative lightweights. Scientific American, 266(5):128–129, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
128a.pdf.

Erickson:1992:DT

[1007] Deborah Erickson. Demonic toxin. Scientific American, 266(5):129– 130, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-129.pdf.

Beardsley:1992:BR

[1008] Tim Beardsley. Business at rio. Scientific American, 266(5):130–131, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-130.pdf.

Stix:1992:VP

[1009] Gary Stix. Village pharmacy. Scientific American, 266(5):132, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-132.pdf.

Corcoran:1992:OST

[1010] Elizabeth Corcoran and Paul Wallich. Overcoming the shortterm syndrome. Scientific American, 266(5):133, May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-133.pdf.

Kaler:1992:ASW

[1011] James B. Kaler. The amateur scientist: Watching the death of a star. Scientific American, 266(5):134–136, May 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n5/pdf/scientificamerican0592-134.pdf.

Morrison:1992:BRb

[1012] Philip Morrison. Book reviews. Scientific American, 266(5):138–142, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n5/pdf/scientificamerican0592-138.pdf.

Eisenberg:1992:EML

[1013] Anne Eisenberg. Essay: Metaphor in the language of science. Scientific American, 266(5):144, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
144.pdf.

Eisenberg:1992:DMI

[1014] Anne Eisenberg. Does metaphor impede or enrich scientific discourse? Scientific American, 266(5):144–??, May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:Ef

[1015] Anonymous. Errata. Scientific American, 266(6):10, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n6/pdf/scientificamerican0692-10b.pdf.

Anonymous:1992:LEf

[1016] Anonymous. Letter to the Editors. Scientific American, 266(6):10, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-10a.pdf.

Anonymous:1992:YAf

[1017] Anonymous. 50 and 100 years ago. Scientific American, 266(6):14, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-14.pdf.

Holloway:1992:SCS

[1018] Marguerite Holloway. Science and the citizen: Still negotiating. Scientific American, 266(6):17–19, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
17.pdf.

Rennie:1992:MMb

[1019] John Rennie. The mice that missed. Scientific American, 266(6):20–26, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-20.pdf.

Holloway:1992:EF

[1020] Marguerite Holloway. The estrogen factor. Scientific American, 266(6): 26, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-26.pdf.

Beardsley:1992:WW

[1021] Tim Beardsley. Weird wonders. Scientific American, 266(6):30–34, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-30.pdf.

Powell:1992:AA

[1022] Corey S. Powell. Astronomical austerity. Scientific American, 266(6):34– 36, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-34.pdf.

Ruthen:1992:SRS

[1023] Russell Ruthen. A switch with the right spin. Scientific American, 266(6):36, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
36.pdf.

Beardsley:1992:SAR

[1024] Tim Beardsley. Scientist, administrator, role model. Scientific American, 266(6):40–41, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
40.pdf.

Chaisson:1992:ERH

[1025] Eric J. Chaisson. Early results from the Hubble Space Telescope. Scientific American, 266(6):44–51, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
44.pdf.

Lawn:1992:LHD

[1026] Richard M. Lawn. Lipoprotein (a) in heart disease. Scientific American, 266(6):54–60, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
54.pdf.

Gasser:1992:TC

[1027] Charles S. Gasser and Robert T. Fraley. Transgenic crops. Scientific American, 266(6):62–69, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
62.pdf.

Anawalt:1992:CM

[1028] Patricia Rieff Anawalt and Frances F. Berdan. The Codex Mendoza. Scientific American, 266(6):70–79, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
70.pdf.

Corcoran:1992:TIRa

[1029] Elizabeth Corcoran. Trends in industrial research: Redesigning research. Scientific American, 266(6):72–??, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Likharev:1992:SE

[1030] Konstantin K. Likharev and Tord Claeson. Single electronics. Scientific American, 266(6):80–85 (Intl. ed. 50–55), June 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n6/pdf/scientificamerican0692-80.pdf.

Winkler:1992:CRW

[1031] William G. Winkler and Konrad Bögel. Control of rabies in wildlife. Scientific American, 266(6):86–91, June 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n6/pdf/scientificamerican0692-86.pdf.

Stewart:1992:RVC

[1032] Ian Stewart. The riddle of the vanishing camel. Scientific American, 266(6):88–??, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Repetto:1992:AEA

[1033] Robert Repetto. Accounting for environmental assets. Scientific American, 266(6):94–100, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
94.pdf.

Corcoran:1992:TIRb

[1034] Elizabeth Corcoran. Trends in industrial research: Redesigning research. Scientific American, 266(6):102–110, June 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n6/pdf/scientificamerican0692-102.pdf.

Corcoran:1992:RR

[1035] Elizabeth Corcoran. Redesigning research. Scientific American, 266(6): 102–??, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1992:SBGb

[1036] Deborah Erickson. Science and business: Genes to order. Scientific American, 266(6):112–114, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
112.pdf.

Erickson:1992:GAB

[1037] Deborah Erickson. Getting agricultural biotech off the RAC. Scientific American, 266(6):115, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
115b.pdf.

Wallich:1992:TD

[1038] Paul Wallich. Tap dance. Scientific American, 266(6):115–116, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-115a.pdf.

stix:1992:JC

[1039] Gary stix. Joy of cooking. Scientific American, 266(6):116–117, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-116.pdf.

Beardsley:1992:FS

[1040] Tim Beardsley. Fuelish study? Scientific American, 266(6):117, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-117.pdf.

Stix:1992:SG

[1041] Gary Stix. Stained glass. Scientific American, 266(6):118–119, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-118.pdf.

Wallich:1992:MS

[1042] Paul Wallich and Elizabeth Corcoran. The MBAs of summer. Scientific American, 266(6):120, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
120.pdf.

Stewart:1992:MRR

[1043] Ian Stewart. Mathematical recreations: The riddle of the vanishing camel. Scientific American, 266(6):122–124, June 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v266/
n6/pdf/scientificamerican0692-122.pdf.

Morrison:1992:BBRd

[1044] Philip Morrison. Brief: Book reviews. Scientific American, 266(6):126– 131, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v266/n6/pdf/scientificamerican0692-126.pdf.

Huber:1992:EJS

[1045] Peter W. Huber. Essay: Junk science in the courtroom. Scientific American, 266(6):132, June 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n6/pdf/scientificamerican0692-
132.pdf.

Anonymous:1992:LEg

[1046] Anonymous. Letters to the Editors. Scientific American, 267(1):10, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-10.pdf.

Anonymous:1992:YAg

[1047] Anonymous. 50 and 100 years ago. Scientific American, 267(1):14, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-14.pdf.

Anonymous:1992:PFH

[1048] Anonymous. 1942: Physicists are the fair-haired boys of World War II. Scientific American, 267(1):14–??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1992:SCG

[1049] Corey S. Powell. Science and the citizen: The golden age of cosmology. Scientific American, 267(1):17–22, July 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n1/pdf/scientificamerican0792-17.pdf.

Beardsley:1992:GB

[1050] Tim Beardsley. Galling the bears. Scientific American, 267(1):24, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-24b.pdf.

Ross:1992:CC

[1051] Philip E. Ross. Compulsive canines. Scientific American, 267(1):24– 26, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-24a.pdf.

Beardsley:1992:DFR

[1052] Tim Beardsley. DNA fingerprinting reconsidered (again). Scientific American, 267(1):26, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
26a.pdf.

Rennie:1992:MLF

[1053] John Rennie. Mother’s little favorite. Scientific American, 267(1):26– 28, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-26b.pdf.

Beardsley:1992:CP

[1054] Tim Beardsley. Colliding physicists. Scientific American, 267(1):28– 30, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-28.pdf.

Fagan:1992:BMCa

[1055] Paul J. Fagan and Michael D. Ward. Building molecular crystals. Scientific American, 267(1):28–??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1992:SS

[1056] Philip Yam. Sound system. Scientific American, 267(1):30, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n1/pdf/scientificamerican0792-30.pdf.

Rennie:1992:DD

[1057] John Rennie. Defining dyslexia. Scientific American, 267(1):31–32, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-31.pdf.

Wallich:1992:RR

[1058] Paul Wallich. Rapid recall. Scientific American, 267(1):32, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n1/pdf/scientificamerican0792-32.pdf.

Hargon:1992:CCC

[1059] John Hargon. COBE’s cosmic cartographer. Scientific American, 267(1): 34–41, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-34.pdf.

Grove:1992:OWE

[1060] Richard H. Grove. Origins of Western environmentalism. Scientific American, 267(1):42–47, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
42.pdf.

Fagan:1992:BMCb

[1061] Paul J. Fagan and Michael D. Ward. Building molecular crystals. Scientific American, 267(1):48–54, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
48.pdf.

Under:1992:GP

[1062] Maurine E. Under and Alfred G. Gilman. G proteins. Scientific American, 267(1):56–65, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
56.pdf.

Linder:1992:GP

[1063] Maurine E. Linder and Alfred G. Gilman. G proteins. Scientific American, 267(1):56–?? (Intl. ed. 36–??), July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holland:1992:GA

[1064] John H. Holland. Genetic algorithms. Scientific American, 267(1): 66–72 (Intl. ed. 44–50), July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
66.pdf.

Horgan:1992:TPQa

[1065] John Horgan. Trends in physics: Quantum philosophy. Scientific American, 267(1):72–??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Phillips:1992:BTM

[1066] Michael Phillips. Breath tests in medicine. Scientific American, 267(1): 74–79, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-74.pdf.

Duellman:1992:RSF

[1067] William E. Duellman. Reproductive strategies of frogs. Scientific American, 267(1):80–87, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
80.pdf.

Toth:1992:LSA

[1068] Nicholas Toth, Desmond Clark, and Giancarlo Ligabue. The last stone ax makers. Scientific American, 267(1):88–93, July 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n1/pdf/scientificamerican0792-88.pdf.

Riolo:1992:SFB

[1069] Rick L. Riolo. Survival of the finest bits. Scientific American, 267(1):89– ??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1992:TPQb

[1070] John Horgan. Trends in physics: Quantum philosophy. Scientific American, 267(1):94–104, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
94.pdf.

Horgan:1992:QP

[1071] John Horgan. Quantum philosophy. Scientific American, 267(1):94– ??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1992:SBP

[1072] Tim Beardsley. Science and business: Piecemeal patents. Scientific American, 267(1):106–108, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
106.pdf.

Corcoran:1992:UST

[1073] Elizabeth Corcoran. U.S. semiconductor toolmakers regain ground. Scientific American, 267(1):108, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
108.pdf.

Corcoran:1992:OFT

[1074] Elizabeth Corcoran. One fish, two fish. Scientific American, 267(1): 109, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-109.pdf.

Erickson:1992:SB

[1075] Deborah Erickson. Sensible biology. Scientific American, 267(1):110, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-110a.pdf.

Erickson:1992:MM

[1076] Deborah Erickson. Metastatic machinations. Scientific American, 267(1):110–111, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
110b.pdf.

Ruthen:1992:FF

[1077] Russell Ruthen. Fickle fluids. Scientific American, 267(1):111–112, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-111.pdf.

Erickson:1992:MV

[1078] Deborah Erickson. Machine vision. Scientific American, 267(1):112, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n1/pdf/scientificamerican0792-112.pdf.

Corcoran:1992:DPS

[1079] Elizabeth Corcoran and Paul Wallich. Doubling up on payoffs for schooling. Scientific American, 267(1):113, July 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n1/pdf/scientificamerican0792-113.pdf.

Riolo:1992:ASS

[1080] Rick L. Riolo. The amateur scientist: Survival of the fittest bits. Scientific American, 267(1):114–116, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
114.pdf.

Gould:1992:BRB

[1081] Stephen Jay Gould. Book reviews: Book review. Scientific American, 267(1):118–121, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
118.pdf.

Gould:1992:RMT

[1082] Stephen Jay Gould. Retrying the monkey trial in a kangaroo court. Scientific American, 267(1):118–??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1992:EWD

[1083] Anne Eisenberg. Essay: Women and the discourse of science. Scientific American, 267(1):122, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n1/pdf/scientificamerican0792-
122.pdf.

Eisenberg:1992:SSL

[1084] Anne Eisenberg. Sexism still lurks in the language of science. Scientific American, 267(1):122–??, July 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:LEh

[1085] Anonymous. Letters to the Editors. Scientific American, 267(2):10–12, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-10.pdf.

Anonymous:1992:Eg

[1086] Anonymous. Errata. Scientific American, 267(2):12, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n2/pdf/scientificamerican0892-12.pdf.

Anonymous:1992:YAh

[1087] Anonymous. 50 and 100 years ago. Scientific American, 267(2):14, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-14.pdf.

Beardsley:1992:SCA

[1088] Tim Beardsley. Science and the citizen: Aborting research. Scientific American, 267(2):17–18, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
17.pdf.

Rennie:1992:E

[1089] John Rennie. Entomophagy. Scientific American, 267(2):20, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-20a.pdf.

Horgan:1992:C

[1090] John Horgan. Counting down. Scientific American, 267(2):20–22, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-20b.pdf.

Horgan:1992:PN

[1091] John Horgan. Plenty of nothing. Scientific American, 267(2):23–26, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-23.pdf.

Powell:1992:RGD

[1092] Corey S. Powell. A rare glimpse of a dim world. Scientific American, 267(2):24, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
24.pdf.

Ross:1992:EE

[1093] Philip E. Ross. East of Eden. Scientific American, 267(2):26, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-26b.pdf.

Ruthen:1992:PG

[1094] Russell Ruthen. Particle goddess. Scientific American, 267(2):26–30, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-26a.pdf.

Erickson:1992:SS

[1095] Deborah Erickson. Sea sick. Scientific American, 267(2):30, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n2/pdf/scientificamerican0892-30.pdf.

Rennie:1992:KIF

[1096] John Rennie. Keeping it in the family. Scientific American, 267(2):43– 44, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-43.pdf.

Dvorak:1992:DKV

[1097] John J. Dvorak, Carl Johnson, and Robert I. Tilling. Dynamics of Kilauea volcano. Scientific American, 267(2):46–53, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n2/pdf/scientificamerican0892-46.pdf.

Rietschel:1992:BE

[1098] Ernst Theodor Rietschel and Helmut Brade. Bacterial endotoxins. Scientific American, 267(2):54–61, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
54.pdf.

Brush:1992:HCB

[1099] Stephen G. Brush. How cosmology became a science. Scientific American, 267(2):62–70, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
62.pdf.

Sherman:1992:NMR

[1100] Paul W. Sherman, Jennifer U. M. Jarvis, and Stanton H. Braude. Naked mole rats. Scientific American, 267(2):72–78, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n2/pdf/scientificamerican0892-72.pdf.

Tattersall:1992:SPE

[1101] Ian Tattersall. Science in pictures: Evolution comes to life. Scientific American, 267(2):80–87, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
80.pdf.

Tattersall:1992:ECL

[1102] Ian Tattersall. Evolution comes to life. Scientific American, 267(2):80– ?? (Intl. ed. 62–??), August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1992:TTAa

[1103] Gary Stix. Trends in transportation: Air trains. Scientific American, 267(2):84–??, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Deutsch:1992:PMP

[1104] Diana Deutsch. Paradoxes of musical pitch. Scientific American, 267(2):88–95, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
88.pdf.

Chaum:1992:AEP

[1105] David Chaum. Achieving electronic privacy. Scientific American, 267(2):96–101, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
96.pdf.

Stewart:1992:IO

[1106] Ian Stewart. The interplanetary Olympics. Scientific American, 267 (2):101–??, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1992:TTAb

[1107] Gary Stix. Trends in transportation: Air trains. Scientific American, 267(2):102–113, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
102.pdf.

Stix:1992:AT

[1108] Gary Stix. Air trains. Scientific American, 267(2):102–??, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Erickson:1992:SBM

[1109] Deborah Erickson. Science and business: Material help. Scientific American, 267(2):114–116, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
114.pdf.

Corcoran:1992:HP

[1110] Elizabeth Corcoran. Hopscotch policy. Scientific American, 267 (2):116–118, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
116.pdf.

Erickson:1992:BSC

[1111] Deborah Erickson. Biotechnology for sale, cheap. Scientific American, 267(2):117, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
117.pdf.

Ross:1992:SLB

[1112] Philip E. Ross. Seven-league boots. Scientific American, 267(2):118, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-118.pdf.

Stix:1992:CI

[1113] Gary Stix. Check it out. Scientific American, 267(2):119–120, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-119.pdf.

Erickson:1992:DDa

[1114] Deborah Erickson. Doomsday diagnostic? Scientific American, 267(2):120, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
120.pdf.

Wallich:1992:EIH

[1115] Paul Wallich. Experimenting with the invisible hand. Scientific American, 267(2):121, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
121.pdf.

Stewart:1992:MRI

[1116] Ian Stewart. Mathematical recreations: The interplanetary olympics. Scientific American, 267(2):122–124, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
122.pdf.

Morrison:1992:BRc

[1117] Philip Morrison. Book reviews. Scientific American, 267(2):126–131, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n2/pdf/scientificamerican0892-126.pdf.

Summers:1992:EMI

[1118] Lawrence Summers. Essay: The most influential investment. Scientific American, 267(2):132, August 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n2/pdf/scientificamerican0892-
132.pdf.

Anonymous:1992:LEi

[1119] Anonymous. Letters to the Editors. Scientific American, 267(3):12, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-12.pdf.

vanCamp:1992:NC

[1120] Drew van Camp. Neurons for computers. Scientific American, 267(3): 12S–??, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:YAi

[1121] Anonymous. 50 and 100 years ago. Scientific American, 267(3):16, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-16.pdf.

Horgan:1992:SCG

[1122] John Horgan. Science and the citizen: Gravity quantized? Scientific American, 267(3):18–20, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
18.pdf.

Anonymous:1992:SCa

[1123] Anonymous. Science and the citizen. Scientific American, 267(3):18–??, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1992:ICW

[1124] John Horgan. It came from within. Scientific American, 267(3):20, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-20.pdf.

Holloway:1992:HA

[1125] Marguerite Holloway. Hot ants. Scientific American, 267(3):23, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-23b.pdf.

Yam:1992:NN

[1126] Philip Yam. Noisy nucleotides. Scientific American, 267(3):23–27, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-23a.pdf.

Rennie:1992:BGC

[1127] John Rennie. Bug in a gilded cage. Scientific American, 267(3):27–28, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-27.pdf.

Corcoran:1992:MWR

[1128] Elizabeth Corcoran. Making water run uphill. Scientific American, 267(3):28, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
28.pdf.

Powell:1992:BL

[1129] Corey S. Powell. Bent light. Scientific American, 267(3):30–32, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-30.pdf.

Holloway:1992:PP

[1130] Marguerite Holloway. Population pressure. Scientific American, 267(3):32–38, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
32.pdf.

Dewdney:1992:P

[1131] A. K. Dewdney. Photovores. Scientific American, 267(3):42, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-42.pdf.

Ruthen:1992:FQH

[1132] Russell Ruthen. The frustrations of a quark hunter. Scientific American, 267(3):44–46, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
44.pdf.

Fischbach:1992:MB

[1133] Gerald D. Fischbach. Mind and brain. Scientific American, 267(3):48–57, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-48.pdf.

Shatz:1992:DB

[1134] Carla J. Shatz. The developing brain. Scientific American, 267(3):60–67, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-60.pdf.

Zeki:1992:VIM

[1135] Semir Zeki. The visual image in mind and brain. Scientific American, 267(3):68–76, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
68.pdf.

Gold-Rakic:1992:WMM

[1136] Patricia S. Gold-Rakic. Working memory and the mind. Scientific American, 267(3):72–??, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kandel:1992:BBL

[1137] Eric R. Kandel and Robert D. Hawkins. The biological basis of learning and individuality. Scientific American, 267(3):78–86, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n3/pdf/scientificamerican0992-78.pdf.

Damasio:1992:BL

[1138] Antonio R. Damasio and Hanna Damasio. Brain and language. Scientific American, 267(3):88–95, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
88.pdf.

Goldman-Rakic:1992:WMM

[1139] Patricia S. Goldman-Rakic. Working memory and the mind. Scientific American, 267(3):110–117, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
110.pdf.

Kimura:1992:SDB

[1140] Doreen Kimura. Sex differences in the brain. Scientific American, 267(3):118–125, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
118.pdf.

Gershon:1992:MDM

[1141] Elliot S. Gershon and Ronald O. Rieder. Major disorders of mind and brain. Scientific American, 267(3):126–133, September 1992.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n3/pdf/scientificamerican0992-126.pdf.

Selkoe:1992:ABA

[1142] Dennis J. Selkoe. Aging brain, aging mind. Scientific American, 267(3):134–142, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
134.pdf.

Hinton:1992:HNN

[1143] Geoffrey E. Hinton. How neural net works learn from experience. Scientific American, 267(3):144–151, September 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n3/pdf/scientificamerican0992-144.pdf.

Crick:1992:PC

[1144] Francis Crick and Christof Koch. The problem of consciousness. Scientific American, 267(3):152–159, September 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n3/pdf/scientificamerican0992-152.pdf.

Erickson:1992:SBH

[1145] Deborah Erickson. Science and business: Hot potato. Scientific American, 267(3):160–162, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
160.pdf.

Anonymous:1992:SBa

[1146] Anonymous. Science and business. Scientific American, 267(3):160–??, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:MDG

[1147] Anonymous. Major developers of genetically engineered foods. Scientific American, 267(3):162, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
162.pdf.

Erickson:1992:SL

[1148] Deborah Erickson. Silver lining. Scientific American, 267(3):163, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-163b.pdf.

Erickson:1992:HS

[1149] Deborah Erickson. The hole story. Scientific American, 267(3):163–164, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-163a.pdf.

Erickson:1992:UA

[1150] Deborah Erickson. Underground allies. Scientific American, 267(3):164, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-164.pdf.

Stix:1992:DD

[1151] Gary Stix. Defensive driving. Scientific American, 267(3):164–166, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-164b.pdf.

Stix:1992:STV

[1152] Gary Stix. See-through view. Scientific American, 267(3):166, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-166.pdf.

Corcoran:1992:EGF

[1153] Elizabeth Corcoran. Economic growth factors. Scientific American, 267(3):167, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
167.pdf.

vanCamp:1992:ASN

[1154] Drew van Camp. The amateur scientist: Neurons for computers. Scientific American, 267(3):170–172, September 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n3/pdf/scientificamerican0992-170.pdf.

Morrison:1992:BRd

[1155] Philip Morrison. Book reviews. Scientific American, 267(3):174–179, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n3/pdf/scientificamerican0992-174.pdf.

Miller:1992:ETM

[1156] Jonathan Miller. Essay: Trouble in mind. Scientific American, 267(3):180, September 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n3/pdf/scientificamerican0992-
180.pdf.

Anonymous:1992:LEj

[1157] Anonymous. Letters to the Editors. Scientific American, 267(4):8–10, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-8.pdf.

Stewart:1992:MGG

[1158] Ian Stewart. Murder at Ghastleigh Grange. Scientific American, 267 (4):9S–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:YAj

[1159] Anonymous. 50 and 100 years ago. Scientific American, 267(4):12–14, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-12.pdf.

Anonymous:1992:YAk

[1160] Anonymous. 50 and 100 years ago. Scientific American, 267(4):12–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:SCE

[1161] Elizabeth Corcoran. Science and the citizen: The edge of chaos. Scientific American, 267(4):17–22, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
17.pdf.

Anonymous:1992:SCb

[1162] Anonymous. Science and the citizen. Scientific American, 267(4):17–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:BLY

[1163] Elizabeth Corcoran. The best little yard sale in New Mexico. Scientific American, 267(4):18, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
18.pdf.

Horgan:1992:LB

[1164] John Horgan. Lying by the book. Scientific American, 267(4):20, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-20.pdf.

Holloway:1992:M

[1165] Marguerite Holloway. Musseling in. Scientific American, 267(4):22–23, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-22.pdf.

Wallich:1992:SW

[1166] Paul Wallich. Smart wheels. Scientific American, 267(4):23, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-23.pdf.

Rennie:1992:FD

[1167] John Rennie. Formula for diabetes? Scientific American, 267(4):24–26, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-24.pdf.

Powell:1992:MT

[1168] Corey S. Powell. A matter of timing. Scientific American, 267(4):26–31, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-26.pdf.

Ruthen:1992:CMM

[1169] Russell Ruthen. The cosmic microwave mirage? Scientific American, 267(4):30, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
30.pdf.

Horgan:1992:IS

[1170] John Horgan. Illuminator of the stars. Scientific American, 267(4):32–40, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-32.pdf.

May:1992:HMS

[1171] Robert M. May. How many species in habit the Earth? Scientific American, 267(4):42–48, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
42.pdf.

Bennett:1992:QC

[1172] Charles H. Bennett, Gilles Brassard, and Artur K. Ekert. Quantum cryptography. Scientific American, 267(4):50–57, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n4/pdf/scientificamerican1092-50.pdf.

Houston:1992:MS

[1173] Charles S. Houston. Mountain sickness. Scientific American, 267(4):58– 66, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-58.pdf.

Grunstein:1992:HRG

[1174] Michael Grunstein. Histones as regulators of genes. Scientific American, 267(4):68–74, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
68.pdf.

DeVries:1992:SCA

[1175] Philip J. DeVries. Singing caterpillars, ants and symbiosis. Scientific American, 267(4):76–82, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
76.pdf.

Beardsley:1992:TSEa

[1176] Tim Beardsley. Trends in science education: Teaching real science. Scientific American, 267(4):78–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Geis:1992:DFS

[1177] Michael W. Geis and John C. Angus. Diamond film semiconductors. Scientific American, 267(4):84–89, October 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n4/pdf/scientificamerican1092-84.pdf.

Blumenschine:1992:SHE

[1178] Robert J. Blumenschine and John A. Cavallo. Scavenging and human evolution. Scientific American, 267(4):90–96, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n4/pdf/scientificamerican1092-90.pdf.

Beardsley:1992:TSEb

[1179] Tim Beardsley. Trends in science education: Teaching real science. Scientific American, 267(4):98–108, October 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n4/pdf/scientificamerican1092-98.pdf.

Beardsley:1992:TRS

[1180] Tim Beardsley. Teaching real science. Scientific American, 267(4):98–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:SBS

[1181] Elizabeth Corcoran. Science and business: Storage space. Scientific American, 267(4):110–112, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
110.pdf.

Anonymous:1992:SBb

[1182] Anonymous. Science and business. Scientific American, 267(4):110–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ross:1992:DS

[1183] Philip E. Ross. Different strokes. Scientific American, 267(4):112–113, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-112.pdf.

Corcoran:1992:OI

[1184] Elizabeth Corcoran. Orphan isotopes. Scientific American, 267(4):113, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-113.pdf.

Beardsley:1992:FG

[1185] Tim Beardsley. Fine grain. Scientific American, 267(4):114–115, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-114.pdf.

Ross:1992:BB

[1186] Philip E. Ross. Billions of buckytubes. Scientific American, 267(4):115–116, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
115.pdf.

Erickson:1992:DDb

[1187] Deborah Erickson. Diagnosis by DNA. Scientific American, 267(4):116, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-116.pdf.

Wallich:1992:BE

[1188] Paul Wallich. The business of education. Scientific American, 267(4):117, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n4/pdf/scientificamerican1092-117.pdf.

Stewart:1992:MRM

[1189] Ian Stewart. Mathematical recreations: Murder at Ghastleigh Grange. Scientific American, 267(4):118–120, October 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n4/pdf/scientificamerican1092-118.pdf.

Anonymous:1992:MRa

[1190] Anonymous. Mathematical recreations. Scientific American, 267(4):118– ??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1992:BBRe

[1191] Philip Morrison. Books: Book reviews. Scientific American, 267(4):122–127, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
122.pdf.

Anonymous:1992:BIW

[1192] Anonymous. Books: The islanders who greeted Columbus; zen astronomy. Scientific American, 267(4):122–??, October 1992. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Piel:1992:EAS

[1193] Gerard Piel. Essay: Agenda 21: Sustainable development. Scientific American, 267(4):128, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n4/pdf/scientificamerican1092-
128.pdf.

Anonymous:1992:EGP

[1194] Anonymous. Essay: Gerard Piel: An important starting point for managing Planet Earth. Scientific American, 267(4):128–??, October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:LEk

[1195] Anonymous. Letters to the Editors. Scientific American, 267(5):10, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-10.pdf.

Anonymous:1992:L

[1196] Anonymous. Letters. Scientific American, 267(5):10–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:YAl

[1197] Anonymous. 50 and 100 years ago. Scientific American, 267(5):14, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-14.pdf.

Anonymous:1992:YAm

[1198] Anonymous. 50 and 100 years ago. Scientific American, 267(5):14–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1992:SCD

[1199] Tim Beardsley. Science and the citizen: Declassified. Scientific American, 267(5):16–18, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
16.pdf.

Anonymous:1992:SCc

[1200] Anonymous. Science and the citizen. Scientific American, 267(5):16–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1992:PL

[1201] Tim Beardsley. Paradise lost? Scientific American, 267(5):18–20, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-18.pdf.

Beardsley:1992:BSB

[1202] Tim Beardsley. Bringing science to the bottom line. Scientific American, 267(5):20, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
20.pdf.

Rennie:1992:CJK

[1203] John Rennie. Cells for Jerry’s Kids. Scientific American, 267(5):22–24, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-22.pdf.

Fein:1992:HCF

[1204] Rashi Fein. Health care forum. Scientific American, 267(5):22–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1992:FF

[1205] Corey S. Powell. Flotsam footwear. Scientific American, 267(5):26, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-26b.pdf.

Powell:1992:SB

[1206] Corey S. Powell. Stellar bells. Scientific American, 267(5):26–28, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-26a.pdf.

Ross:1992:WWI

[1207] Philip E. Ross. Who were the Indo-Europeans? Scientific American, 267(5):28, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
28.pdf.

Freed:1992:ERS

[1208] Wendy L. Freed. The expansion, rate and size of the universe. Scientific American, 267(5):30–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1992:DD

[1209] Tim Beardsley. Desert dynamics. Scientific American, 267(5):32–36, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-32.pdf.

Corcoran:1992:KCL

[1210] Elizabeth Corcoran. Kicking chaos out of lasers. Scientific American, 267(5):36, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
36.pdf.

Horgan:1992:IW

[1211] John Horgan. The intellectual warrior. Scientific American, 267 (5):38–44, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
38.pdf.

Fein:1992:HCR

[1212] Rashi Fein. Health care reform. Scientific American, 267(5):46–53, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-46.pdf.

Freedman:1992:ERS

[1213] Wendy L. Freedman. The expansion rate and size of the Universe. Scientific American, 267(5):54–60, November 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n5/pdf/scientificamerican1192-54.pdf.

Greenberg:1992:LONa

[1214] Joseph H. Greenberg and Merritt Ruhlen. Linguistic origins of Native Americans. Scientific American, 267(5):60–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Littlewood:1992:RS

[1215] Bev Littlewood and Lorenzo Strigini. The risks of software. Scientific American, 267(5):62–75, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
62.pdf.

Stix:1992:TMMa

[1216] Gary Stix. Trends in micromechanics: Micron machinations. Scientific American, 267(5):72–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Olson:1992:SPV

[1217] Arthur J. Olson and David S. Goodsell. Science in pictures: Visualizing biological molecules. Scientific American, 267(5):76–81, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n5/pdf/scientificamerican1192-76.pdf.

Olson:1992:VBM

[1218] Arthur J. Olson and David S. Goodsell. Visualizing biological molecules. Scientific American, 267(5):76–?? (Intl. ed. 44–47, 50–51), November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Levinton:1992:BBA

[1219] Jeffrey S. Levinton. The Big Bang of animal evolution. Scientific American, 267(5):84–91, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
84.pdf.

Greenberg:1992:LONb

[1220] Joseph H. Greenberg and Merritt Ruhlen. Linguistic origins of Native Americans. Scientific American, 267(5):94–99, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n5/pdf/scientificamerican1192-94.pdf.

Gingerich:1992:AAC

[1221] Owen Gingerich. Astronomy in the age of Columbus. Scientific American, 267(5):100–105, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
100.pdf.

Stix:1992:TMMb

[1222] Gary Stix. Trends in micromechanics: Micron machinations. Scientific American, 267(5):106–117, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
106.pdf.

Stix:1992:MM

[1223] Gary Stix. Micron machinations. Scientific American, 267(5):106–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1992:SBB

[1224] Elizabeth Corcoran. Science and business: Building networks. Scientific American, 267(5):118–120, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
118.pdf.

Anonymous:1992:SBc

[1225] Anonymous. Science and business. Scientific American, 267(5):118–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1992:CS

[1226] Philip Yam. Cool sounds. Scientific American, 267(5):120–121, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-120.pdf.

Corcoran:1992:S

[1227] Elizabeth Corcoran. Semiready. Scientific American, 267(5):121–122, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-121.pdf.

Erickson:1992:IM

[1228] Deborah Erickson. Intercepted messages. Scientific American, 267 (5):122–123, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
122.pdf.

Erickson:1992:ID

[1229] Deborah Erickson. Intuitive design. Scientific American, 267(5):124–125, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n5/pdf/scientificamerican1192-124.pdf.

Holloway:1992:RWT

[1230] Marguerite Holloway and Paul Wallich. A risk worth taking. Scientific American, 267(5):126, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
126.pdf.

Carlson:1992:ASS

[1231] George A. Carlson. The amateur scientist: Sighting Cepheid variables. Scientific American, 267(5):128–130, November 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n5/pdf/scientificamerican1192-128.pdf.

Anonymous:1992:AS

[1232] Anonymous. The amateur scientist. Scientific American, 267(5):128–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1992:BBRf

[1233] Philip Morrison. Books: Book reviews. Scientific American, 267 (5):132–137, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
132.pdf.

Anonymous:1992:B

[1234] Anonymous. Books. Scientific American, 267(5):132–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schulhof:1992:EWB

[1235] Michael Schulhof. Essay: Why business needs scientists. Scientific American, 267(5):138, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n5/pdf/scientificamerican1192-
138.pdf.

Schulhof:1992:E

[1236] Michael Schulhof. Essay. Scientific American, 267(5):138–??, November 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:Eh

[1237] Anonymous. Erratum. Scientific American, 267(6):10, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n6/pdf/scientificamerican1292-10b.pdf.

Anonymous:1992:LEl

[1238] Anonymous. Letters to the Editors. Scientific American, 267(6):10, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-10a.pdf.

Anonymous:1992:LEm

[1239] Anonymous. Letters to the editor. Scientific American, 267(6):10–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:YAn

[1240] Anonymous. 50 and 100 years ago. Scientific American, 267(6):14, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-14.pdf.

Anonymous:1992:YAo

[1241] Anonymous. 50 and 100 years ago. Scientific American, 267(6):14–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1992:SYE

[1242] John Horgan. Special year-end section: The new challenges. Scientific American, 267(6):16–22, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
16.pdf.

Anonymous:1992:NC

[1243] Anonymous. New challenges for 1993. Scientific American, 267(6):16–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1992:TR

[1244] Philip Yam. Theoretical resistance. Scientific American, 267(6):24, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-24.pdf.

Rennie:1992:EE

[1245] John Rennie. Essential but expendable. Scientific American, 267(6):25–28, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
25.pdf.

Corcoran:1992:RL

[1246] Elizabeth Corcoran. Racing light. Scientific American, 267(6):32–36, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-32.pdf.

Ross:1992:ST

[1247] Philip E. Ross. Shaking the tree. Scientific American, 267(6):36–38, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-36.pdf.

Ruthen:1992:PP

[1248] Russell Ruthen. Puzzling powerhouse. Scientific American, 267(6):38–39, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-38.pdf.

Beardsley:1992:QD

[1249] Tim Beardsley. Quantum dissidents. Scientific American, 267(6):39–40, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-39.pdf.

Beardsley:1992:CJ

[1250] Tim Beardsley. Cambrian jolt. Scientific American, 267(6):40–42, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-40.pdf.

Beardsley:1992:CR

[1251] Tim Beardsley. Cellular response. Scientific American, 267(6):42–44, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-42.pdf.

Corcoran:1992:TG

[1252] Elizabeth Corcoran. Thinking green. Scientific American, 267(6):44–46, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-44.pdf.

Stix:1992:RH

[1253] Gary Stix. Running hot. Scientific American, 267(6):46–48, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-46.pdf.

Stix:1992:FF

[1254] Gary Stix. Finding fault. Scientific American, 267(6):48–52, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v267/n6/pdf/scientificamerican1292-48.pdf.

Holloway:1992:UM

[1255] Marguerite Holloway. Unlikely messengers. Scientific American, 267(6):52–56, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
52.pdf.

Stevenson:1992:LAS

[1256] Harold W. Stevenson. Learning from Asian schools. Scientific American, 267(6):70–76, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
70.pdf.

Gabrielse:1992:ECA

[1257] Gerald Gabrielse. Extremely cold antiprotons. Scientific American, 267(6):78–89, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
78.pdf.

Rusting:1992:TBWa

[1258] Ricki L. Rusting. Trends in biology: Why do we age? Scientific American, 267(6):86–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Joyce:1992:DME

[1259] Gerald F. Joyce. Directed molecular evolution. Scientific American, 267(6):90–97, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
90.pdf.

Stewart:1992:CHC

[1260] Ian Stewart. Christmas in the House of Chaos. Scientific American, 267 (6):96–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cohen:1992:WCS

[1261] I. Bernard Cohen. What Columbus “saw” in 1492. Scientific American, 267(6):100–106, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
100.pdf.

Sataloff:1992:HV

[1262] Robert T. Sataloff. The human voice. Scientific American, 267 (6):108–115, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
108.pdf.

Smith:1992:EME

[1263] Vernon L. Smith and Arlington W. Williams. Experimental market economics. Scientific American, 267(6):116–121, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n6/pdf/scientificamerican1292-116.pdf.

Seyfarth:1992:MMM

[1264] Robert M. Seyfarth and Dorothy L. Cheney. Meaning and mind in monkeys. Scientific American, 267(6):122–128, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n6/pdf/scientificamerican1292-122.pdf.

Rusting:1992:TBWb

[1265] Ricki L. Rusting. Trends in biology: Why do we age? Scientific American, 267(6):130–141, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
130.pdf.

Rusting:1992:WDW

[1266] Ricki L. Rusting. Why do we age? Scientific American, 267(6):130–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1992:MRC

[1267] Ian Stewart. Mathematical recreations: Christmas in the House of Chaos. Scientific American, 267(6):144–146, December 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n6/pdf/scientificamerican1292-144.pdf.

Anonymous:1992:MRb

[1268] Anonymous. Mathematical recreations. Scientific American, 267(6): 144–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1992:BRS

[1269] Philip Morrison and Phylis Morrison. Book reviews: Science books for young people. Scientific American, 267(6):148, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n6/pdf/scientificamerican1292-148.pdf.

Anonymous:1992:BR

[1270] Anonymous. Book reviews. Scientific American, 267(6):148–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1992:AIA

[1271] Anonymous. Annual index: Annual index 1992. Scientific American, 267(6):158–161, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v267/n6/pdf/scientificamerican1292-
158.pdf.

Eisenberg:1992:EIE

[1272] Anne Eisenberg. Essay: Imperial English: The language of science? Scientific American, 267(6):162, December 1992. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v267/
n6/pdf/scientificamerican1292-162.pdf.

Eisenberg:1992:E

[1273] Anne Eisenberg. Essay. Scientific American, 267(6):162–??, December 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:Ea

[1274] Anonymous. Erratum. Scientific American, 266(3):12, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-12b.pdf.

Anonymous:1993:LEa

[1275] Anonymous. Letters to the Editors. Scientific American, 266(3):12, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-12a.pdf.

Anonymous:1993:YAIa

[1276] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 266(3):16, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
16.pdf.

Horgan:1993:SCA

[1277] John Horgan. Science and the citizen: Antarctic meltdown. Scientific American, 266(3):19–28, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
19.pdf.

Horgan:1993:STM

[1278] John Horgan. Are scientists too messy for Antarctica? Scientific American, 266(3):22–23, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
22.pdf.

Holloway:1993:LC

[1279] Marguerite Holloway. Lethal cascade. Scientific American, 266(3):28–30, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-28.pdf.

Rennie:1993:NF

[1280] John Rennie. Noah’s freezer. Scientific American, 266(3):30, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-30a.pdf.

Powell:1993:YS

[1281] Corey S. Powell. Young suns. Scientific American, 266(3):30–32, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-30b.pdf.

Ruthen:1993:QQ

[1282] Russell Ruthen. Quark quest. Scientific American, 266(3):32–34, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-32.pdf.

Horgan:1993:SAH

[1283] John Horgan. Stubbornly ahead of his time. Scientific American, 266(3):36–40, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
36.pdf.

Dunker:1993:WABa

[1284] Kenneth F. Dunker and Basile G. Rabbat. Why America’s bridges are crumbling. Scientific American, 266(3):66–72, March 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-66.pdf.

Abramowicz:1993:BHCa

[1285] Marek Artur Abramowicz. Black holes and the centrifugal force paradox. Scientific American, 266(3):74–81, March 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-74.pdf.

Boon:1993:TISa

[1286] Thierry Boon. Teaching the immune system to fight cancer. Scientific American, 266(3):82–89, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
82.pdf.

Depp:1993:FPDa

[1287] Steven W. Depp and Webster E. Howard. Flat-panel displays. Scientific American, 266(3):90–97, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
90.pdf.

Tumlinson:1993:HPWa

[1288] James H. Tumlinson, W. Joe Lewis, and Louise E. M. Vet. How parasitic wasps find their hosts. Scientific American, 266(3):100–106, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-100.pdf.

deBeaune:1993:IALa

[1289] Sophie A. de Beaune and Randall White. Ice Age lamps. Scientific American, 266(3):108–113, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
108.pdf.

Goulding:1993:FFAa

[1290] Michael Goulding. Flooded forests of the Amazon. Scientific American, 266(3):114–120, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
114.pdf.

Rennie:1993:TGDa

[1291] John Rennie. Trends in genetics: DNA’s new twists. Scientific American, 266(3):122–132, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
122.pdf.

Gibbs:1993:SBBa

[1292] W. Wayt Gibbs. Science and business: Back to basics. Scientific American, 266(3):134–136, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
134.pdf.

Gibbs:1993:PP

[1293] W. Wayt Gibbs. Preventing the preventable. Scientific American, 266(3): 135, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-135.pdf.

Rennie:1993:NSO

[1294] John Rennie. No snake oil here. Scientific American, 266(3):136–138, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-136.pdf.

Yam:1993:HG

[1295] Philip Yam. High on gravity. Scientific American, 266(3):138–140, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-138.pdf.

Stix:1993:H

[1296] Gary Stix and Tom Koppel. Headsets. Scientific American, 266(3):141, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-141.pdf.

Wallich:1993:MPG

[1297] Paul Wallich and Marguerite Holloway. More profitable to give than to receive? Scientific American, 266(3):142, March 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-142.pdf.

Strong-Gunderson:1993:ASF

[1298] Janet M. Strong-Gunderson. The amateur scientist: Flight-testing fruit flies. Scientific American, 266(3):144–145, March 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n3/pdf/scientificamerican0393-144.pdf.

May:1993:BR

[1299] Robert M. May. Book review. Scientific American, 266(3):146–149, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n3/pdf/scientificamerican0393-146.pdf.

Landman:1993:EIA

[1300] Otto E. Landman. Essay: Inheritance of acquired characteristics. Scientific American, 266(3):150, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n3/pdf/scientificamerican0393-
150.pdf.

Anonymous:1993:Eb

[1301] Anonymous. Erratum. Scientific American, 268(1):12, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n1/pdf/scientificamerican0193-12b.pdf.

Anonymous:1993:LEb

[1302] Anonymous. Letters to the Editors. Scientific American, 268(1):12, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-12a.pdf.

Anonymous:1993:YAIb

[1303] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 268(1):14, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
14.pdf.

Anonymous:1993:YAa

[1304] Anonymous. 50 and 100 years ago. Scientific American, 268(1):14–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1993:SCH

[1305] John Rennie. Science and the citizen: How many genes and Y. Scientific American, 268(1):16–17, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
16.pdf.

Anonymous:1993:SCa

[1306] Anonymous. Science and the citizen. Scientific American, 268(1):16–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ross:1993:EG

[1307] Philip E. Ross. Endangered genes. Scientific American, 268(1):17, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-17.pdf.

Holloway:1993:P

[1308] Marguerite Holloway. Pitohui! Scientific American, 268(1):20–22, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-20.pdf.

Wallich:1993:CE

[1309] Paul Wallich. Crunching epsilon. Scientific American, 268(1):22–24, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-22.pdf.

Yam:1993:GH

[1310] Philip Yam. A gene for hypertension. Scientific American, 268(1):24, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-24.pdf.

Stix:1993:AG

[1311] Gary Stix. Anything goes. Scientific American, 268(1):26–27, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-26.pdf.

Powell:1993:MW

[1312] Corey S. Powell. Machos or wimps? Scientific American, 268(1):27–28, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-27.pdf.

Vogel:1993:BP

[1313] Shawna Vogel and J. Rennie. Booby prizes. Scientific American, 268(1): 28, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-28.pdf.

Holloway:1993:PRL

[1314] Marguerite Holloway. Profile: Rita Levi-Montalcini. Scientific American, 268(1):32–36, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
32.pdf.

Brown:1993:CB

[1315] Barbara E. Brown and John C. Ogden. Coral bleaching. Scientific American, 268(1):64–70, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
64.pdf.

vandenBergh:1993:HMW

[1316] Sidney van den Bergh and James E. Hesser. How the Milky Way formed. Scientific American, 268(1):72–78, January 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n1/pdf/scientificamerican0193-72.pdf.

Sharon:1993:CCR

[1317] Nathan Sharon and Halina Lis. Carbohydrates in cell recognition. Scientific American, 268(1):82–89, January 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n1/pdf/scientificamerican0193-82.pdf.

York:1993:EHE

[1318] Derek York. The earliest history of the Earth. Scientific American, 268(1):90–96, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
90.pdf.

Tattersall:1993:MLa

[1319] Ian Tattersall. Madagascar’s lemurs. Scientific American, 268 (1):110–117, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
110.pdf.

Ruthen:1993:TND

[1320] Russell Ruthen. Trends in nonlinear dynamics: Adapting to complexity. Scientific American, 268(1):110–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Tattersall:1993:MLb

[1321] Ian Tattersall. Madagascar’s lemurs. Scientific American, 268(1):110–?? (Intl. ed. 90–??), January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Reed:1993:QD

[1322] Mark A. Reed. Quantum dots. Scientific American, 268(1):118–123, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-118.pdf.

Milner:1993:MDH

[1323] Peter M. Milner. The mind and Donald O. Hebb. Scientific American, 268(1):124–129, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
124.pdf.

Horn:1993:BB

[1324] Henry S. Horn. Biodiversity in the backyard. Scientific American, 268 (1):125–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ruthen:1993:AC

[1325] Russell Ruthen. Adapting to complexity. Scientific American, 268(1):130–140, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
130.pdf.

Stix:1993:SBB

[1326] Gary Stix. Science and business: Back to roots. Scientific American, 268(1):142–143, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
142.pdf.

Anonymous:1993:SBa

[1327] Anonymous. Science and business. Scientific American, 268(1):142–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1993:NC

[1328] Elizabeth Corcoran. National conundrums. Scientific American, 268(1):144–145, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
144.pdf.

Corcoran:1993:SL

[1329] Elizabeth Corcoran. Soft Lego. Scientific American, 268(1):145–146, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-145.pdf.

Stix:1993:HC

[1330] Gary Stix. Habeas corpus. Scientific American, 268(1):146–147, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-146.pdf.

Stix:1993:GA

[1331] Gary Stix. Geometry acquisition. Scientific American, 268(1):147, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n1/pdf/scientificamerican0193-147.pdf.

Corcoran:1993:RI

[1332] Elizabeth Corcoran. The return on infrastructure. Scientific American, 268(1):148, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
148.pdf.

Horn:1993:ASB

[1333] Henry S. Horn. The amateur scientist: Biodiversity in the backyard. Scientific American, 268(1):150–152, January 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n1/pdf/scientificamerican0193-150.pdf.

Anonymous:1993:ASa

[1334] Anonymous. The amateur scientist. Scientific American, 268(1):150–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1993:BRH

[1335] Philip Morrison. Book reviews: Heavenly objects. Scientific American, 268(1):154–159, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
154.pdf.

Anonymous:1993:BRa

[1336] Anonymous. Book reviews. Scientific American, 268(1):154–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Johnson:1993:ELB

[1337] Howard M. Johnson. Essay: The life of a black scientist. Scientific American, 268(1):160, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n1/pdf/scientificamerican0193-
160.pdf.

Johnson:1993:E

[1338] Howard M. Johnson. Essay. Scientific American, 268(1):160–??, January 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEc

[1339] Anonymous. Letters to the Editors. Scientific American, 268(2):12, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-12.pdf.

Anonymous:1993:YAIc

[1340] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 268(2):16, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
16.pdf.

Anonymous:1993:YAb

[1341] Anonymous. 50 and 100 years ago. Scientific American, 268(2):16–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1993:SCL

[1342] Corey S. Powell. Science and the citizen: Livable planets. Scientific American, 268(2):18–20, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
18.pdf.

Anonymous:1993:SCb

[1343] Anonymous. Science and the citizen. Scientific American, 268(2):18–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1993:WIT

[1344] John Horgan. What if they don’t have radios? Scientific American, 268(2):20, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
20.pdf.

Horgan:1993:CCa

[1345] John Horgan. COBE corroborated. Scientific American, 268(2):22, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-22.pdf.

Ross:1993:FF

[1346] Philip E. Ross. Faux fullerenes. Scientific American, 268(2):24, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-24b.pdf.

Horgan:1993:GC

[1347] John Horgan. Genes and crime. Scientific American, 268(2):24–29, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-24a.pdf.

Horgan:1993:APW

[1348] John Horgan. The artist, the physicist and the waterfall. Scientific American, 268(2):30, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
30a.pdf.

Wallich:1993:EE

[1349] Paul Wallich. Electronic envelopes? Scientific American, 268(2):30–32, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-30b.pdf.

Beardsley:1993:MW

[1350] Tim Beardsley. Making waves. Scientific American, 268(2):32, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-32.pdf.

Corcoran:1993:PYB

[1351] Elizabeth Corcoran. The physicist as a young businessman. Scientific American, 268(2):34–35, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
34.pdf.

Homer-Dixon:1993:ECV

[1352] Thomas F. Homer-Dixon, Jeffrey H. Boutwell, and George W. Rathjens. Environmental change and violent conflict. Scientific American, 268(2):38–45, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
38.pdf.

Bishop:1993:RHTa

[1353] David J. Bishop, Peter L. Gammel, and David A. Huse. Resistance in high-temperature superconductors. Scientific American, 268(2):48–55, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-48.pdf.

Rhodes:1993:ZF

[1354] Daniela Rhodes and Aaron Klug. Zinc fingers. Scientific American, 268(2):56–65, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
56.pdf.

Hoffmann:1993:HSC

[1355] Roald Hoffmann. How should chemists think? Scientific American, 268(2):66–73, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
66.pdf.

Beardsley:1993:TRSa

[1356] Tim Beardsley. Trends in Russian science: Selling to survive. Scientific American, 268(2):68–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rickey:1993:SPT

[1357] George Rickey. Science in pictures: a technology of kinetic art. Scientific American, 268(2):74–79, February 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n2/pdf/scientificamerican0293-74.pdf.

Rickey:1993:TKA

[1358] George Rickey. A technology of kinetic art. Scientific American, 268 (2):74–?? (Intl. ed. 50–??), February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Tuomanen:1993:BBB

[1359] Elaine Tuomanen. Breaching the blood-brain barrier. Scientific American, 268(2):80–84, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
80.pdf.

Stewart:1993:PTS

[1360] Ian Stewart. A partly true story. Scientific American, 268(2):85–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Swade:1993:RCB

[1361] Doron D. Swade. Redeeming Charles Babbage’s mechanical computer. Scientific American, 268(2):86–91, February 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n2/pdf/scientificamerican0293-86.pdf.

Beardsley:1993:TRSb

[1362] Tim Beardsley. Trends in Russian science: Selling to survive. Scientific American, 268(2):92–100, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
92.pdf.

Beardsley:1993:SS

[1363] Tim Beardsley. Selling to survive. Scientific American, 268(2):92–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:SBZ

[1364] Gary Stix. Science and business: Zip code breakers. Scientific American, 268(2):102–103, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
102.pdf.

Anonymous:1993:SBb

[1365] Anonymous. Science and business. Scientific American, 268(2):102–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1993:SS

[1366] Philip Yam. Surreal science. Scientific American, 268(2):103–104, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-103.pdf.

Koppel:1993:LL

[1367] Tom Koppel. Lightning lure. Scientific American, 268(2):105, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-105b.pdf.

Corcoran:1993:LC

[1368] Elizabeth Corcoran. Learning companies. Scientific American, 268(2):105–108, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
105a.pdf.

Stix:1993:SSa

[1369] Gary Stix. Shell shocked. Scientific American, 268(2):108, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n2/pdf/scientificamerican0293-108.pdf.

Corcoran:1993:AES

[1370] Elizabeth Corcoran. The analytical economist: Soul of a new economic idea? Scientific American, 268(2):109, February 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n2/pdf/scientificamerican0293-109.pdf.

Stewart:1993:MRP

[1371] Ian Stewart. Mathematical recreations: a partly true story. Scientific American, 268(2):110–112, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
110.pdf.

Anonymous:1993:MRa

[1372] Anonymous. Mathematical recreations. Scientific American, 268(2): 110–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Keyfitz:1993:BRP

[1373] Nathan Keyfitz. Book review: The preservation of the planet. Scientific American, 268(2):114–119, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v268/n2/pdf/scientificamerican0293-
114.pdf.

Anonymous:1993:BRb

[1374] Anonymous. Book review. Scientific American, 268(2):114–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cassidy:1993:EGB

[1375] David C. Cassidy. Essay: Germany and the bomb: New evidence. Scientific American, 268(2):120, February 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n2/pdf/scientificamerican0293-120.pdf.

Cassidy:1993:E

[1376] David C. Cassidy. Essay. Scientific American, 268(2):120–??, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEd

[1377] Anonymous. Letters to the editor. Scientific American, 268(3):12–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:YAc

[1378] Anonymous. 50 and 100 years ago. Scientific American, 268(3):16–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:SCc

[1379] Anonymous. Science and the citizen. Scientific American, 268(3):19–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dunker:1993:WABb

[1380] Kenneth F. Dunker and Basile G. Rabbat. Why America’s bridges are crumbling. Scientific American, 268(3):66–?? (Intl. ed. 18–??), March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Abramowicz:1993:BHCb

[1381] Marek Artur Abramowicz. Black holes and the centrifugal force paradox. Scientific American, 268(3):74–?? (Intl. ed. 26–??), March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Boon:1993:TISb

[1382] Thierry Boon. Teaching the immune system to fight cancer. Scientific American, 268(3):82–?? (Intl. ed. 32–??), March 1993. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1993:TGDb

[1383] John Rennie. Trends in genetics: DNA’s new twists. Scientific American, 268(3):88–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Depp:1993:FPDb

[1384] Steven W. Depp and Webster E. Howard. Flat-panel displays. Scientific American, 268(3):90–?? (Intl. ed. 40–??), March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Tumlinson:1993:HPWb

[1385] James H. Tumlinson, W. Joe Lewis, and Louise E. M. Vet. How parasitic wasps find their hosts. Scientific American, 268(3):100–?? (Intl. ed. 46– ??), March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gunderson:1993:FTF

[1386] Janet M. Strong Gunderson. Flight-testing fruit flies. Scientific American, 268(3):106–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

deBeaune:1993:IALb

[1387] Sophie A. de Beaune and Randall White. Ice age lamps. Scientific American, 268(3):108–?? (Intl. ed. 74–??), March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Goulding:1993:FFAb

[1388] Michael Goulding. Flooded forests of the Amazon. Scientific American, 268(3):114–?? (Intl. ed. 80–??), March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1993:DNT

[1389] John Rennie. DNA’s new twist. Scientific American, 268(3):122–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:SBc

[1390] Anonymous. Science and business. Scientific American, 268(3):134–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:ASb

[1391] Anonymous. The amateur scientist. Scientific American, 268(3):144–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:BRc

[1392] Anonymous. Book review. Scientific American, 268(3):146–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Landman:1993:E

[1393] Otto E. Landman. Essay. Scientific American, 268(3):150–??, March 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEe

[1394] Anonymous. Letters to the Editors. Scientific American, 268(4):12, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-12.pdf.

Anonymous:1993:YAId

[1395] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 268(4):16, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
16.pdf.

Anonymous:1993:YAd

[1396] Anonymous. 50 and 100 years ago. Scientific American, 268(4):16– ??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1993:SCO

[1397] Marguerite Holloway. Science and the citizen: Obstacle course. Scientific American, 268(4):18–24, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
18.pdf.

Anonymous:1993:SCd

[1398] Anonymous. Science and the citizen. Scientific American, 268(4):18– ??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

H:1993:EFG

[1399] M. H. Easing a financial gag. Scientific American, 268(4):22, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-22.pdf.

Horgan:1993:ESR

[1400] John Horgan. An eternally self-reproducing Cosmos? Scientific American, 268(4):24, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
24.pdf.

Ruthen:1993:FC

[1401] Russell Ruthen. Flat chemistry. Scientific American, 268(4):26, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-26.pdf.

Rennie:1993:KWS

[1402] John Rennie. Kitty, we shrunk your brain. Scientific American, 268(4): 29, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-29b.pdf.

Rennie:1993:SMX

[1403] John Rennie. Spot marks the X. Scientific American, 268(4):29–30, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-29a.pdf.

Ruthen:1993:IZ

[1404] Russell Ruthen. Illuminating zero. Scientific American, 268(4):32–33, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-32.pdf.

Ross:1993:AS

[1405] Philip E. Ross. Ancient sleepers. Scientific American, 268(4):33, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-33.pdf.

Powell:1993:AH

[1406] Corey S. Powell. Asteroid hunters. Scientific American, 268(4):34–40, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-34.pdf.

Horgan:1993:NGW

[1407] John Horgan. The nicest guy in Washington. Scientific American, 268(4):42–43, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
42.pdf.

Olshansky:1993:AHS

[1408] S. Jay Olshansky, Bruce A. Carnes, and Christine K. Cassel. The aging of the human species. Scientific American, 268(4):46–52, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-46.pdf.

Haroche:1993:CQE

[1409] Serge Haroche and Jean Michel Raimond. Cavity quantum electrodynamics. Scientific American, 268(4):54–62, April 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-54.pdf.

Konishi:1993:LTE

[1410] Masakazu Konishi. Listening with two ears. Scientific American, 268(4):66–73, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
66.pdf.

Stix:1993:TMCa

[1411] Gary Stix. Trends in materials: Concrete solutions. Scientific American, 268(4):72–??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Friend:1993:CS

[1412] Cynthia M. Friend. Catalysis on surfaces. Scientific American, 268(4): 74–79 (Intl. ed. 42–??), April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
74.pdf.

FitzGerald:1993:RBS

[1413] Gerard J. FitzGerald. The reproductive behavior of the stickleback. Scientific American, 268(4):80–85, April 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-80.pdf.

Ewald:1993:EV

[1414] Paul W. Ewald. The evolution of virulence. Scientific American, 268(4):86–93, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
86.pdf.

Stewart:1993:RFL

[1415] Ian Stewart. The rise and fall of the Lunar M-Pire. Scientific American, 268(4):90–??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bar-Yosef:1993:MHL

[1416] Ofer Bar-Yosef and Bernard Vandermeersch. Modern humans in the Levant. Scientific American, 268(4):94–100, April 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-94.pdf.

Stix:1993:TMCb

[1417] Gary Stix. Trends in materials: Concrete solutions. Scientific American, 268(4):102–112, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
102.pdf.

Stix:1993:CS

[1418] Gary Stix. Concrete solutions. Scientific American, 268(4):102–??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:SBH

[1419] Gary Stix. Science and business: High technology. Scientific American, 268(4):113–114, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
113.pdf.

Anonymous:1993:SBd

[1420] Anonymous. Science and business. Scientific American, 268(4):113– ??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

S:1993:AS

[1421] G. S. Advertising space. Scientific American, 268(4):114, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-114.pdf.

Beardsley:1993:VV

[1422] Tim Beardsley. Virtuous viruses. Scientific American, 268(4):115–116, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-115.pdf.

Gibbs:1993:LM

[1423] W. Wayt Gibbs. Light motif. Scientific American, 268(4):116–117, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-116.pdf.

Beardsley:1993:PBT

[1424] Tim Beardsley. Putting the brakes on tankers. Scientific American, 268(4):117, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
117.pdf.

Service:1993:GV

[1425] Robert Service. Global view. Scientific American, 268(4):118, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-118.pdf.

Corcoran:1993:FFD

[1426] Elizabeth Corcoran and Paul Wallich. Frequent flaps in the deregulated skies. Scientific American, 268(4):119, April 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-119.pdf.

Stewart:1993:MRR

[1427] Ian Stewart. Mathematical recreations: The rise and fall of the Lunar M-pire. Scientific American, 268(4):120–121, April 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n4/pdf/scientificamerican0493-120.pdf.

Anonymous:1993:MRb

[1428] Anonymous. Mathematical recreations. Scientific American, 268(4):120– ??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1993:BR

[1429] Philip Morrison. Book reviews. Scientific American, 268(4):123–127, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n4/pdf/scientificamerican0493-123.pdf.

Anonymous:1993:BRd

[1430] Anonymous. Book reviews. Scientific American, 268(4):123–??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1993:ESM

[1431] Anne Eisenberg. Essay: Scientists in the movies. Scientific American, 268(4):128, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n4/pdf/scientificamerican0493-
128.pdf.

Eisenberg:1993:Ea

[1432] Anne Eisenberg. Essay. Scientific American, 268(4):128–??, April 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:Ec

[1433] Anonymous. Errata. Scientific American, 268(5):12, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n5/pdf/scientificamerican0593-12b.pdf.

Anonymous:1993:LEf

[1434] Anonymous. Letters to the Editors. Scientific American, 268(5):12, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-12a.pdf.

Anonymous:1993:YAIe

[1435] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 268(5):16, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
16.pdf.

Anonymous:1993:YAe

[1436] Anonymous. 50 and 100 years ago. Scientific American, 268(5):16– ??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1993:SCT

[1437] John Rennie. Science and the citizen: Triple whammy. Scientific American, 268(5):18–22, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
18.pdf.

Anonymous:1993:SCe

[1438] Anonymous. Science and the citizen. Scientific American, 268(5):18– ??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1993:BI

[1439] John Rennie. Balanced immunity. Scientific American, 268(5):22–24, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-22.pdf.

Beardsley:1993:HA

[1440] Tim Beardsley. Honest advertising. Scientific American, 268(5):24– 27, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-24.pdf.

Powell:1993:TFV

[1441] Corey S. Powell. Three faces of Venus. Scientific American, 268(5):26– 27, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-26.pdf.

Ruthen:1993:AD

[1442] Russell Ruthen. Attractive and demure. Scientific American, 268(5): 30, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-30b.pdf.

Wallich:1993:MMA

[1443] Paul Wallich. Make, model and. … Scientific American, 268(5):30– 32, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-30a.pdf.

Yam:1993:DD

[1444] Philip Yam. “Daisy, Daisy”. Scientific American, 268(5):32–33, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-32.pdf.

Horgan:1993:WES

[1445] John Horgan. The worst enemy of science. Scientific American, 268(5): 36–37, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-36.pdf.

Sen:1993:ELD

[1446] Amartya Sen. The economics of life and death. Scientific American, 268(5):40–47, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
40.pdf.

Jeanloz:1993:CMB

[1447] Raymond Jeanloz and Thorne Lay. The core mantle boundary. Scientific American, 268(5):48–55, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
48.pdf.

Welch:1993:HCR

[1448] William J. Welch. How cells respond to stress. Scientific American, 268(5):56–64, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
56.pdf.

Powell:1993:TAIa

[1449] Corey S. Powell. Trends in astrophysics: Inconstant cosmos. Scientific American, 268(5):68–??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Osada:1993:IG

[1450] Yoshihito Osada and Simon B. Ross-Murphy. Intelligent gels. Scientific American, 268(5):82–87, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
82.pdf.

Jescavage-Bernard:1993:MPW

[1451] Karen Jescavage-Bernard and Anders Crofoot. Mapping to preserve a watershed. Scientific American, 268(5):85–??, May 1993. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wood:1993:SPP

[1452] Denis Wood. Science in pictures: The power of maps. Scientific American, 268(5):88–93, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
88.pdf.

Wood:1993:PM

[1453] Denis Wood. The power of maps. Scientific American, 268(5):88–?? (Intl. ed. 48–??), May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kalin:1993:NF

[1454] Ned H. Kalin. The neurobiology of fear. Scientific American, 268(5):94– 101, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-94.pdf.

Hovis:1993:PMD

[1455] R. Corby Hovis and Helge Kragh. P. A. M. Dirac and the beauty of physics. Scientific American, 268(5):104–109, May 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n5/pdf/scientificamerican0593-104.pdf.

Powell:1993:TAIb

[1456] Corey S. Powell. Trends in astrophysics: Inconstant Cosmos. Scientific American, 268(5):110–118, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
110.pdf.

Powell:1993:IC

[1457] Corey S. Powell. Inconstant cosmos. Scientific American, 268(5):110– ??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:SBG

[1458] Gary Stix. Science and business: Gigabit gestalt. Scientific American, 268(5):122–126, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
122.pdf.

Anonymous:1993:SBe

[1459] Anonymous. Science and business. Scientific American, 268(5):122– ??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1993:SGS

[1460] Tim Beardsley. Silicon Glen’s semi-success. Scientific American, 268(5): 124, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-124.pdf.

Beardsley:1993:BO

[1461] Tim Beardsley. Bright opportunity. Scientific American, 268(5):126– 128, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-126.pdf.

Gibbs:1993:TS

[1462] W. Wayt Gibbs. Three swings. Scientific American, 268(5):128–129, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-128.pdf.

Gibbs:1993:GG

[1463] W. Wayt Gibbs. Garbage in, gravel out. Scientific American, 268(5): 130, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-130.pdf.

Fields:1993:WDS

[1464] Judith Fields and Paul Wallich. Why do some industries pay better? Scientific American, 268(5):131, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
131.pdf.

Jescavage-Bernard:1993:ASM

[1465] Karen Jescavage-Bernard and Anders Crofoot. The amateur scientist: Mapping to preserve a watershed. Scientific American, 268(5):134–136, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-134.pdf.

Anonymous:1993:ASc

[1466] Anonymous. The amateur scientist. Scientific American, 268(5):134– ??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:BRe

[1467] Anonymous. Book reviews. Scientific American, 268(5):138–142, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n5/pdf/scientificamerican0593-138.pdf.

Anonymous:1993:BRf

[1468] Anonymous. Book reviews. Scientific American, 268(5):138–??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Arthur:1993:EWD

[1469] W. Brian Arthur. Essay: Why do things become more complex? Scientific American, 268(5):144, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n5/pdf/scientificamerican0593-
144.pdf.

Arthur:1993:E

[1470] W. Brian Arthur. Essay. Scientific American, 268(5):144–??, May 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEg

[1471] Anonymous. Letters to the Editors. Scientific American, 268(6):9–13, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-9.pdf.

Anonymous:1993:YAIf

[1472] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 268(6):16, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
16.pdf.

Anonymous:1993:YAf

[1473] Anonymous. 50 and 100 years ago. Scientific American, 268(6):16– ??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ross:1993:SCL

[1474] Philip E. Ross. Science and the citizen: Living cure. Scientific American, 268(6):18–23, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
18.pdf.

Anonymous:1993:SCf

[1475] Anonymous. Science and the citizen. Scientific American, 268(6):18– ??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1993:BS

[1476] Philip Yam. A bus for Scotty. Scientific American, 268(6):26, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-26a.pdf.

Beardsley:1993:E

[1477] Tim Beardsley. Ecolocation. Scientific American, 268(6):26–32, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-26b.pdf.

Guterl:1993:MCP

[1478] Fred Guterl. Mr. Clinton, put down that watering can. Scientific American, 268(6):28, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
28.pdf.

Rennie:1993:SS

[1479] John Rennie. SNAPs and SNAREs. Scientific American, 268(6):32–33, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-32.pdf.

Powell:1993:CD

[1480] Corey S. Powell. Cosmic diagnosis. Scientific American, 268(6):33, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-33.pdf.

Rennie:1993:DV

[1481] John Rennie. David’s victory. Scientific American, 268(6):34–35, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-34.pdf.

Wallich:1993:TW

[1482] Paul Wallich. Time warp. Scientific American, 268(6):35, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-35.pdf.

Holloway:1993:PVW

[1483] Marguerite Holloway. A powerful voice for women. Scientific American, 268(6):36–40, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
36.pdf.

Pollack:1993:URC

[1484] Henry N. Pollack and David S. Chapman. Underground records of changing climate. Scientific American, 268(6):44–50, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-44.pdf.

Miley:1993:MDR

[1485] George K. Miley and Kenneth C. Chambers. The most distant radio galaxies. Scientific American, 268(6):54–61, June 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-54.pdf.

Glover:1993:C

[1486] David M. Glover, Cayetano Gonzalez, and Jordan W. Raff. The centrosome. Scientific American, 268(6):62–68, June 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-62.pdf.

Keyes:1993:FT

[1487] Robert W. Keyes. The future of the transistor. Scientific American, 268(6):70–78, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
70.pdf.

Horgan:1993:TBGa

[1488] John Horgan. Trends in behavioral genetics: Eugenics revisited. Scientific American, 268(6):92–??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carter:1993:MPV

[1489] C. Sue Carter and Lowell L. Getz. Monogamy and the prairie vole. Scientific American, 268(6):100–106, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
100.pdf.

Frith:1993:A

[1490] Uta Frith. Autism. Scientific American, 268(6):108–114, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-108.pdf.

Stewart:1993:BFB

[1491] Ian Stewart. A bundling fool beats the wrap. Scientific American, 268 (6):109–??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Vogel:1993:GWC

[1492] Hans Ulrich Vogel. The Great Well of China. Scientific American, 268(6):116–121, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
116.pdf.

Horgan:1993:TBGb

[1493] John Horgan. Trends in behavioral genetics: Eugenics revisited. Scientific American, 268(6):122–131, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
122.pdf.

Horgan:1993:ER

[1494] John Horgan. Eugenics revisited. Scientific American, 268(6):122–??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:SBR

[1495] Gary Stix. Science and business: Red-banner burger. Scientific American, 268(6):132–135, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
132.pdf.

Anonymous:1993:SBf

[1496] Anonymous. Science and business. Scientific American, 268(6):132– ??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1993:SAA

[1497] Philip Yam. Is that shirt ascorbic acid? Scientific American, 268(6):134– 135, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-134.pdf.

Gibbs:1993:E

[1498] W. Wayt Gibbs. Extinguished. Scientific American, 268(6):136, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-136.pdf.

Gibbs:1993:MW

[1499] W. Wayt Gibbs. Making wavelets. Scientific American, 268(6):137–138, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-137.pdf.

Stix:1993:ML

[1500] Gary Stix. Micromachine line. Scientific American, 268(6):138, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-138.pdf.

Yam:1993:PB

[1501] Philip Yam. “Plastics, Benjamin… ”. Scientific American, 268(6):140, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-140.pdf.

Wallich:1993:EWC

[1502] Paul Wallich. Are economy watchers chasing a mirage? Scientific American, 268(6):141, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v268/n6/pdf/scientificamerican0693-
141.pdf.

Stewart:1993:MRB

[1503] Ian Stewart. Mathematical recreations: a bundling fool beats the wrap. Scientific American, 268(6):142–144, June 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-142.pdf.

Anonymous:1993:BRg

[1504] Anonymous. Book reviews. Scientific American, 268(6):142–??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:MRc

[1505] Anonymous. Mathematical recreations. Scientific American, 268(6):142– ??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:BRh

[1506] Anonymous. Book reviews. Scientific American, 268(6):146–151, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v268/n6/pdf/scientificamerican0693-146.pdf.

Brown:1993:ECS

[1507] George E. Brown, Jr. Essay: Can scientists “make change their friend”? Scientific American, 268(6):152, June 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v268/
n6/pdf/scientificamerican0693-152.pdf.

Brown:1993:E

[1508] George E. Brown, Jr. Essay. Scientific American, 268(6):152–??, June 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:Ed

[1509] Anonymous. Errata. Scientific American, 269(1):10, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-10a.pdf.

Anonymous:1993:LEh

[1510] Anonymous. Letters to the Editors. Scientific American, 269(1):10, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-10.pdf.

Anonymous:1993:YAIg

[1511] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 269(1):12, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
12.pdf.

Anonymous:1993:YAg

[1512] Anonymous. 50 and 100 years ago. Scientific American, 269(1):12– ??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1993:SCB

[1513] Marguerite Holloway. Science and the citizen: The big nada? Scientific American, 269(1):20–21, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
20.pdf.

Anonymous:1993:SCg

[1514] Anonymous. Science and the citizen. Scientific American, 269(1):20– ??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1993:M

[1515] Corey S. Powell. Moonball. Scientific American, 269(1):21–23, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-21.pdf.

Gibbs:1993:B

[1516] W. Wayt Gibbs. Banzai! Scientific American, 269(1):22, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-22.pdf.

Yam:1993:QQ

[1517] Philip Yam. QED for QCD. Scientific American, 269(1):23–24, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-23.pdf.

Horgan:1993:KW

[1518] John Horgan. A kinder war. Scientific American, 269(1):24–26, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-24.pdf.

Rennie:1993:HH

[1519] John Rennie. Healing hearing. Scientific American, 269(1):26–27, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-26.pdf.

Beardsley:1993:PIS

[1520] Tim Beardsley. Profile: Irving S. Shapiro: Science’s unscientific champion. Scientific American, 269(1):28–29, July 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-28.pdf.

Morgan:1993:RAM

[1521] M. Granger Morgan. Risk analysis and management. Scientific American, 269(1):32–41, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
32.pdf.

Eigen:1993:VQ

[1522] Manfred Eigen. Viral quasispecies. Scientific American, 269(1):42–49, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-42.pdf.

Vickers-Rich:1993:APD

[1523] Patricia Vickers-Rich and Thomas Hewitt Rich. Australia’s polar dinosaurs. Scientific American, 269(1):50–55, July 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-50.pdf.

Itano:1993:AMT

[1524] Wayne M. Itano and Norman F. Ramsey. Accurate measurement of time. Scientific American, 269(1):56–65, July 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-56.pdf.

Harken:1993:STC

[1525] Alden H. Harken. Surgical treatment of cardiac arrhythmias. Scientific American, 269(1):68–74, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
68.pdf.

Kosko:1993:FLa

[1526] Bart Kosko and Satoru Isaka. Fuzzy logic. Scientific American, 269(1): 76–81, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-76.pdf.

Holloway:1993:TES

[1527] Marguerite Holloway. Trends in environmental science: Sustaining the Amazon. Scientific American, 269(1):76–??, July 1993. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kosko:1993:FLb

[1528] Bart Kosko and Satoru Isaka. Fuzzy logic. Scientific American, 269(1): 76–?? (Intl. ed. 62–??), July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Osterbrock:1993:EHE

[1529] Donald E. Osterbrock, Joel A. Gwinn, and Ronald S. Brashear. Edwin Hubble and the Expanding Universe. Scientific American, 269(1):84– 89, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-84.pdf.

writer:1993:TES

[1530] Marguerite Hollowaystaff writer. Trends in environmental science: Sustaining the Amazon. Scientific American, 269(1):90–99, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-90.pdf.

Holloway:1993:SA

[1531] Marguerite Holloway. Sustaining the Amazon. Scientific American, 269 (1):90–??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1993:TD

[1532] Ian Stewart. The topological dressmaker. Scientific American, 269(1): 93–??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1993:SBT

[1533] W. Wayt Gibbs. Science and business: Try, try again. Scientific American, 269(1):101–103, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
101.pdf.

Anonymous:1993:SBg

[1534] Anonymous. Science and business. Scientific American, 269(1):101– ??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1993:PSP

[1535] Tim Beardsley. Putting a spin on parasites. Scientific American, 269(1): 103, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-103b.pdf.

Stix:1993:Ea

[1536] Gary Stix. Earcons. Scientific American, 269(1):103–107, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-103a.pdf.

Beardsley:1993:KSS

[1537] Tim Beardsley. Keeping the Sun shining on British technology. Scientific American, 269(1):106–107, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
106.pdf.

Gibbs:1993:PF

[1538] W. Wayt Gibbs. Practical fractal. Scientific American, 269(1):107– 108, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-107.pdf.

Wallich:1993:AEH

[1539] Paul Wallich and Marguerite Holloway. The analytical economist: Health care without perverse incentives. Scientific American, 269(1): 109, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n1/pdf/scientificamerican0793-109.pdf.

Stewart:1993:MRT

[1540] Ian Stewart. Mathematical recreations: The topological dressmaker. Scientific American, 269(1):110–112, July 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-110.pdf.

Anonymous:1993:MRd

[1541] Anonymous. Mathematical recreations. Scientific American, 269(1):110– ??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Piel:1993:BRC

[1542] Gerard Piel. Book review: Can history stop repeating itself? Scientific American, 269(1):114–118, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
114.pdf.

Anonymous:1993:BRi

[1543] Anonymous. Book review. Scientific American, 269(1):114–??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gomory:1993:ESH

[1544] Ralph Gomory and Hirsh Cohen. Essay: Science: How much is enough? Scientific American, 269(1):120, July 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n1/pdf/scientificamerican0793-120.pdf.

Gomory:1993:E

[1545] Ralph Gomory and Hirsh Cohen. Essay. Scientific American, 269(1): 120–??, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEi

[1546] Anonymous. Letters to the Editors. Scientific American, 269(2):10–11, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-10.pdf.

Anonymous:1993:YAIh

[1547] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 269(2):12, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
12.pdf.

Anonymous:1993:YAh

[1548] Anonymous. 50 and 100 years ago. Scientific American, 269(2):12–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1993:SCW

[1549] John Rennie. Science and the citizen: Who is normal? Scientific American, 269(2):14–17, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
14.pdf.

Anonymous:1993:SCh

[1550] Anonymous. Science and the citizen. Scientific American, 269(2):14–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1993:HLC

[1551] Tim Beardsley. But he’d have to leave the cigars behind. Scientific American, 269(2):16, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
16.pdf.

Ruthen:1993:SM

[1552] Russell Ruthen. Strange matters. Scientific American, 269(2):17, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-17.pdf.

Holloway:1993:SS

[1553] Marguerite Holloway. Sound science? Scientific American, 269(2):20, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-20.pdf.

Powell:1993:FM

[1554] Corey S. Powell. Fast moves. Scientific American, 269(2):22–24, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-22.pdf.

Horgan:1993:ES

[1555] John Horgan. Off to an early start. Scientific American, 269(2):24, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-24.pdf.

Horgan:1993:CCb

[1556] John Horgan. Culture clash. Scientific American, 269(2):26, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-26.pdf.

Horgan:1993:PFJ

[1557] John Horgan. Profile: Freeman J. Dyson: Perpendicular to the mainstream. Scientific American, 269(2):27–28, August 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n2/pdf/scientificamerican0893-27.pdf.

VonHippel:1993:ENW

[1558] Frank Von Hippel, Marvin Miller, Harold Feiveson, Anatoli Diakov, and Frans Berkhout. Eliminating nuclear warheads. Scientific American, 269(2):44–49, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
44.pdf.

Chiao:1993:FTL

[1559] Raymond Y. Chiao, Paul G. Kwiat, and Aephraim M. Steinberg. Faster than light? Scientific American, 269(2):52–60, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n2/pdf/scientificamerican0893-52.pdf.

Schwartz:1993:CA

[1560] Ronald H. Schwartz. T cell anergy. Scientific American, 269(2):62–71, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-62.pdf.

Malin:1993:SPU

[1561] David F. Malin. Science in pictures: a Universe of color. Scientific American, 269(2):72–77, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
72.pdf.

Malin:1993:UC

[1562] David F. Malin. A universe of color. Scientific American, 269(2):72– ?? (Intl. ed. 56–??), August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ditto:1993:MC

[1563] William L. Ditto and Louis M. Pecora. Mastering chaos. Scientific American, 269(2):78–84, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
78.pdf.

Stix:1993:TCDa

[1564] Gary Stix. Trends in communications: Domesticating cyberspace. Scientific American, 269(2):84–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Milton:1993:DPE

[1565] Katharine Milton. Diet and primate evolution. Scientific American, 269(2):86–93, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
86.pdf.

Macklis:1993:GRS

[1566] Roger M. Macklis. The Great Radium Scandal. Scientific American, 269(2):94–99, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
94.pdf.

Stix:1993:TCDb

[1567] Gary Stix. Trends in communications: Domesticating cyberspace. Scientific American, 269(2):100–110, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
100.pdf.

Stix:1993:DC

[1568] Gary Stix. Domesticating cyberspace. Scientific American, 269(2):100– ??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Neff:1993:CGC

[1569] Joseph Neff and Thomas L. Carroll. Circuits that get chaos in sync. Scientific American, 269(2):101–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1993:SBBb

[1570] W. Wayt Gibbs. Science and business: Body English. Scientific American, 269(2):112–115, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
112.pdf.

Anonymous:1993:SBh

[1571] Anonymous. Science and business. Scientific American, 269(2):112–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:ITY

[1572] Gary Stix. “I’ll trade you a wallaroo for an aardvark… ”. Scientific American, 269(2):114, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
114.pdf.

Beardsley:1993:BM

[1573] Tim Beardsley. Blood money? Scientific American, 269(2):115–117, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-115.pdf.

Wallich:1993:CRA

[1574] Paul Wallich. Clipper runs aground. Scientific American, 269(2):116, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-116.pdf.

Gibbs:1993:BF

[1575] W. Wayt Gibbs. Blue films. Scientific American, 269(2):117–118, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-117.pdf.

Yam:1993:GC

[1576] Philip Yam. A gem of a catalyst. Scientific American, 269(2):118, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-118.pdf.

Fields:1993:AEW

[1577] Judith Fields and Paul Wallich. The analytical economist: Will deregulation save the banks? Scientific American, 269(2):119, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n2/pdf/scientificamerican0893-119.pdf.

Neff:1993:ASC

[1578] Joseph Neff and Thomas L. Carroll. The amateur scientist: Circuits that get chaos in sync. Scientific American, 269(2):120–122, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n2/pdf/scientificamerican0893-120.pdf.

Anonymous:1993:ASd

[1579] Anonymous. The amateur scientist. Scientific American, 269(2):120–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1993:BRMa

[1580] Philip Morrison. Book reviews: Matters of the heart. Scientific American, 269(2):124–127, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
124.pdf.

Anonymous:1993:BRj

[1581] Anonymous. Book reviews. Scientific American, 269(2):124–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1993:EKC

[1582] Anne Eisenberg. Essay: Keeping up with computerese. Scientific American, 269(2):128, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n2/pdf/scientificamerican0893-
128.pdf.

Eisenberg:1993:Eb

[1583] Anne Eisenberg. Essay. Scientific American, 269(2):128–??, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEj

[1584] Anonymous. Letters to the Editors. Scientific American, 269(3):12, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-12.pdf.

Anonymous:1993:YAIi

[1585] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 269(3):16, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
16.pdf.

Beardsley:1993:SCS

[1586] Tim Beardsley and Russell Ruthen. Science and the citizen: Super trouble. Scientific American, 269(3):20–24, September 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-20.pdf.

Anonymous:1993:SCi

[1587] Anonymous. Science and the citizen. Scientific American, 269(3):20–??, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leutwyler:1993:SCW

[1588] Kristin Leutwyler. Shaking conventional wisdom. Scientific American, 269(3):24, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
24.pdf.

Powell:1993:JJ

[1589] Corey S. Powell. Jovian jolt. Scientific American, 269(3):26–30, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-26.pdf.

Horgan:1993:FM

[1590] John Horgan. Fermat’s MacGuffin. Scientific American, 269(3):30–34, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-30.pdf.

Rennie:1993:MM

[1591] John Rennie. Malignant mimicry. Scientific American, 269(3):34–38, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-34.pdf.

Beardsley:1993:FF

[1592] Tim Beardsley. Fads and feds. Scientific American, 269(3):39–44, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-39.pdf.

Yam:1993:PRS

[1593] Philip Yam. Profile: Richard E. Smalley: The all-star of buckyball. Scientific American, 269(3):46–48, September 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-46.pdf.

Nossal:1993:LDI

[1594] Sir Gustav J. V. Nossal. Life, death and the immune system. Scientific American, 269(3):52–62, September 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-52.pdf.

Weissman:1993:HIS

[1595] Irving L. Weissman and Max D. Cooper. How the immune system develops. Scientific American, 269(3):64–71, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-64.pdf.

Janeway:1993:HIS

[1596] Charles A. Janeway, Jr. How the immune system recognizes invaders. Scientific American, 269(3):72–79, September 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-72.pdf.

Marrack:1993:HIS

[1597] Philippa Marrack and John W. Kappler. How the immune system recognizes the body. Scientific American, 269(3):80–89, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-80.pdf.

Paul:1993:IDI

[1598] William E. Paul. Infectious diseases and the immune system. Scientific American, 269(3):90–97, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
90.pdf.

Greene:1993:AIS

[1599] Warner C. Greene. AIDS and the immune system. Scientific American, 269(3):98–105, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
98.pdf.

Steinman:1993:AD

[1600] Lawrence Steinman. Autoimmune disease. Scientific American, 269 (3):106–114, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
106.pdf.

Lichtenstein:1993:AIS

[1601] Lawrence M. Lichtenstein. Allergy and the immune system. Scientific American, 269(3):116–124, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
116.pdf.

Stewart:1993:STS

[1602] Ian Stewart. A shepherd takes a sheep shot. Scientific American, 269(3): 118–??, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wigzell:1993:IST

[1603] Hans Wigzell. The immune system as a therapeutic agent. Scientific American, 269(3):126–134, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
126.pdf.

Mitchison:1993:WWS

[1604] Avrion Mitchison. Will we survive? Scientific American, 269(3):136–144, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-136.pdf.

Gibbs:1993:SBC

[1605] W. Wayt Gibbs. Science and business: Crystal light. Scientific American, 269(3):146–147, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
146.pdf.

Anonymous:1993:SBi

[1606] Anonymous. Science and business. Scientific American, 269(3):146–??, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:CC

[1607] Gary Stix. Contented cows? Scientific American, 269(3):147, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-147.pdf.

Guterl:1993:TGA

[1608] Fred Guterl, Robert Pattonin Tokyo, and Gary Stix. Très grande alliance. Scientific American, 269(3):148–149, September 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-148.pdf.

Gibbs:1993:VM

[1609] W. Wayt Gibbs. In vitro, in the money. Scientific American, 269(3):150, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-150.pdf.

Gibbs:1993:NS

[1610] W. Wayt Gibbs. Natural selection. Scientific American, 269(3):151–152, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-151.pdf.

Stix:1993:BIC

[1611] Gary Stix. Blackballing the inner city. Scientific American, 269(3):152, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n3/pdf/scientificamerican0993-152.pdf.

Holloway:1993:AEG

[1612] Marguerite Holloway and Paul Wallich. The analytical economist: The greening of garbage. Scientific American, 269(3):153, September 1993.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-153.pdf.

Stewart:1993:MRS

[1613] Ian Stewart. Mathematical recreations: a shepherd takes a sheep shot. Scientific American, 269(3):154–156, September 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n3/pdf/scientificamerican0993-154.pdf.

Morrison:1993:BRMb

[1614] Philip Morrison. Book reviews: Monticello redux. Scientific American, 269(3):158–163, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
158.pdf.

Bloom:1993:EIR

[1615] Barry R. Bloom. Essay: Immunology and reciprocity. Scientific American, 269(3):164, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n3/pdf/scientificamerican0993-
164.pdf.

Bloom:1993:E

[1616] Barry R. Bloom. Essay. Scientific American, 269(3):164–??, September 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEk

[1617] Anonymous. Letters to the Editors. Scientific American, 269(4):8, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-8.pdf.

Anonymous:1993:LEl

[1618] Anonymous. Letters to the editor. Scientific American, 269(4):8–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:YAIj

[1619] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 269(4):12, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
12.pdf.

Anonymous:1993:YAj

[1620] Anonymous. 50 and 100 years ago. Scientific American, 269(4):12–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1993:SCH

[1621] Marguerite Holloway. Science and the citizen: Hard times. Scientific American, 269(4):14–16, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
14.pdf.

Anonymous:1993:SCj

[1622] Anonymous. Science and the citizen. Scientific American, 269(4):14–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1993:SS

[1623] W. Wayt Gibbs. Sentries and saboteurs. Scientific American, 269(4):16– 24, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-16.pdf.

Gibbs:1993:CE

[1624] W. Wayt Gibbs. Creative evolution. Scientific American, 269(4):20–21, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-20.pdf.

Beardsley:1993:NGS

[1625] Tim Beardsley. Never give a sucker an even break. Scientific American, 269(4):22, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
22.pdf.

Beardsley:1993:SDG

[1626] Tim Beardsley. Sharks do get cancer. Scientific American, 269(4):24–26, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-24.pdf.

Stix:1993:RSR

[1627] Gary Stix. Run silent, run (not so) cheap. Scientific American, 269(4):26, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-26.pdf.

Ross:1993:JV

[1628] Philip E. Ross. Jurassic virus? Scientific American, 269(4):28, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-28.pdf.

Ruthen:1993:RHS

[1629] Russell Ruthen. Revealing the hidden sequence. Scientific American, 269(4):30–31, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
30.pdf.

Lents:1993:CAA

[1630] James M. Lents and William J. Kelly. Clearing the air in Los Angeles. Scientific American, 269(4):32–39, October 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n4/pdf/scientificamerican1093-32.pdf.

Coffin:1993:LIP

[1631] Millard F. Coffin and Olav Eldholm. Large igneous provinces. Scientific American, 269(4):42–49, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
42.pdf.

Doolittle:1993:EMM

[1632] Russell F. Doolittle and Peer Bork. Evolutionarily mobile modules in proteins. Scientific American, 269(4):50–56, October 1993.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n4/pdf/scientificamerican1093-50.pdf.

Halsey:1993:EF

[1633] Thomas C. Halsey and James E. Martin. Electrorheological fluids. Scientific American, 269(4):58–64, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
58.pdf.

Cox:1993:WPP

[1634] Paul Alan Cox. Water-pollinated plants. Scientific American, 269(4):68– 74, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-68.pdf.

Horgan:1993:TMDa

[1635] John Horgan. Trends in mathematics: The death of proof. Scientific American, 269(4):74–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hinton:1993:SBD

[1636] Geoffrey E. Hinton, David C. Plaut, and Tim Shallice. Simulating brain damage. Scientific American, 269(4):76–82, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n4/pdf/scientificamerican1093-76.pdf.

Kvarning:1993:RV

[1637] Lars-Åke Kvarning. Raising the Vasa. Scientific American, 269(4):84–91, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-84.pdf.

Klingenberg:1993:MFS

[1638] Daniel J. Klingenberg. Making fluids into solids with magnets. Scientific American, 269(4):90–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1993:DP

[1639] John Horgan. The death of proof. Scientific American, 269(4):92–103, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1993:TMDb

[1640] John Horgan. Trends in mathematics: The death of proof. Scientific American, 269(4):92–103, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
92.pdf.

Beardsley:1993:SBC

[1641] Tim Beardsley. Science and business: CRADA mania. Scientific American, 269(4):104–106, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
104.pdf.

Anonymous:1993:SBj

[1642] Anonymous. Science and business. Scientific American, 269(4):104–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shapiro:1993:WMR

[1643] Joshua Shapiro. A wand for the meter reader. Scientific American, 269(4):105, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
105.pdf.

Fitzgerald:1993:MA

[1644] Karen Fitzgerald. Magnetic apprehensions. Scientific American, 269(4):106–107, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
106.pdf.

Gibbs:1993:CP

[1645] W. Wayt Gibbs. Chronologically privileged. Scientific American, 269(4):107, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
107b.pdf.

Patton:1993:ML

[1646] Robert Patton. Mag lift. Scientific American, 269(4):107–108, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-107a.pdf.

Stix:1993:PP

[1647] Gary Stix. Power pack. Scientific American, 269(4):108–110, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-108.pdf.

Stix:1993:DFT

[1648] Gary Stix and Paul Wallich. A digital fix for the Third World? Scientific American, 269(4):111, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
111.pdf.

Klingenberg:1993:ASM

[1649] Daniel J. Klingenberg. The amateur scientist: Making fluids into solids with magnets. Scientific American, 269(4):112–113, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n4/pdf/scientificamerican1093-112.pdf.

Anonymous:1993:ASe

[1650] Anonymous. The amateur scientist. Scientific American, 269(4):112–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:BRk

[1651] Anonymous. Book reviews. Scientific American, 269(4):114–119, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n4/pdf/scientificamerican1093-114.pdf.

Anonymous:1993:BRl

[1652] Anonymous. Book reviews. Scientific American, 269(4):114–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wassersug:1993:ETH

[1653] Richard Wassersug. Essay: Tadpoles from heaven. Scientific American, 269(4):120, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n4/pdf/scientificamerican1093-
120.pdf.

Wassersug:1993:E

[1654] Richard Wassersug. Essay. Scientific American, 269(4):120–??, October 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:LEm

[1655] Anonymous. Letters to the Editors. Scientific American, 269(5):8, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-8.pdf.

Anonymous:1993:LEn

[1656] Anonymous. Letters to the editor. Scientific American, 269(5):8–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:YAIk

[1657] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 269(5):12, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
12.pdf.

Anonymous:1993:YAk

[1658] Anonymous. 50 and 100 years ago. Scientific American, 269(5):12–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wallich:1993:SCG

[1659] Paul Wallich. Science and the citizen: Grim statistics. Scientific American, 269(5):14, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
14a.pdf.

Yam:1993:DI

[1660] Philip Yam. Dot’s incredible. Scientific American, 269(5):14–16, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-14b.pdf.

Anonymous:1993:SCk

[1661] Anonymous. Science and the citizen. Scientific American, 269(5):14–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1993:WFC

[1662] John Horgan. Were four corners victims biowar casualties? Scientific American, 269(5):16, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
16.pdf.

Rennie:1993:IF

[1663] John Rennie. Insects are forever. Scientific American, 269(5):18–20, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-18.pdf.

Powell:1993:GC

[1664] Corey S. Powell. Goldilocks cosmology. Scientific American, 269(5):20–23, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
20.pdf.

Beardsley:1993:SF

[1665] Tim Beardsley. Sausage factory. Scientific American, 269(5):23, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-23a.pdf.

Leutwyler:1993:PP

[1666] Kristin Leutwyler. “Pollution, pollution… ”. Scientific American, 269(5):23–28, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
23b.pdf.

Beardsley:1993:UA

[1667] Tim Beardsley. Unraveling Alzheimer’s. Scientific American, 269 (5):28–30, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
28.pdf.

Horgan:1993:MAI

[1668] John Horgan. The mastermind of artificial intelligence. Scientific American, 269(5):35–38, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
35.pdf.

Anonymous:1993:DDF

[1669] Anonymous. Debate: Does free trade harm the environment? Scientific American, 269(5):41, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
41.pdf.

Bhagwati:1993:CFT

[1670] Jagdish Bhagwati. The case for free trade. Scientific American, 269(5):42–49, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
42.pdf.

Iacono:1993:ABA

[1671] Domenic J. Iacono. The art of Boris Artzybasheff. Scientific American, 269(5):46–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Daly:1993:PFT

[1672] Herman E. Daly. The perils of free trade. Scientific American, 269(5):50–57, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
50.pdf.

Hingorani:1993:HPEa

[1673] N. G. Hingorani and K. E. Stahlkopf. High-power electronics. Scientific American, 269(5):52–59, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Changeux:1993:CSB

[1674] Jean-Pierre Changeux. Chemical signaling in the brain. Scientific American, 269(5):58–62, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
58.pdf.

vandenHeuvel:1993:XRB

[1675] Edward P. J. van den Heuvel and Jan Van Paradijs. X-ray binaries. Scientific American, 269(5):64–71, November 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n5/pdf/scientificamerican1193-64.pdf.

Iacono:1993:SPA

[1676] Domenic J. Iacono. Science in pictures: The art of Boris Artzybasheff. Scientific American, 269(5):72–77, November 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n5/pdf/scientificamerican1193-72.pdf.

Jacono:1993:ABA

[1677] Domenic J. Jacono. The art of Boris Artzybasheff. Scientific American, 269(5):72–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hingorani:1993:HPEb

[1678] Narain G. Hingorani and Karl E. Stahlkopf. High-power electronics. Scientific American, 269(5):78–85 (Intl. ed. 52–??), November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n5/pdf/scientificamerican1193-78.pdf.

Holloway:1993:TSSa

[1679] Marguerite Holloway. Trends in the sociology of science: A lab of her own. Scientific American, 269(5):84–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1993:FLT

[1680] Ian Stewart. Fermat’s last time-trip. Scientific American, 269(5):85–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Paabo:1993:AD

[1681] Svante Pääbo. Ancient DNA. Scientific American, 269(5):86–92, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-86.pdf.

Holloway:1993:TSSb

[1682] Marguerite Holloway. Trends in the sociology of science: a lab of her own. Scientific American, 269(5):94–103, November 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n5/pdf/scientificamerican1193-94.pdf.

Holloway:1993:LHO

[1683] Marguerite Holloway. A lab of her own. Scientific American, 269(5): 94–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1993:SBT

[1684] Gary Stix. Science and business: Turning green. Scientific American, 269(5):104–107, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
104.pdf.

Anonymous:1993:SBk

[1685] Anonymous. Science and business. Scientific American, 269(5):104–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1993:MFT

[1686] W. Wayt Gibbs. More fun than a root canal. Scientific American, 269(5):106, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
106.pdf.

Beardsley:1993:PP

[1687] Tim Beardsley. Power play. Scientific American, 269(5):107–108, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-107.pdf.

Patton:1993:SS

[1688] Robert Patton. Software skipper. Scientific American, 269(5):108–109, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-108.pdf.

Stix:1993:Eb

[1689] Gary Stix. ECM for 747s. Scientific American, 269(5):109–110, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-109.pdf.

Fields:1993:ILB

[1690] Judith Fields. Ivy League bonus babies. Scientific American, 269(5):111, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-111.pdf.

Stewart:1993:MRF

[1691] Ian Stewart. Mathematical recreations: Fermat’s last time-trip. Scientific American, 269(5):112–115, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
112.pdf.

Anonymous:1993:MRe

[1692] Anonymous. Mathematical recreations. Scientific American, 269(5): 112–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:BRm

[1693] Anonymous. Book reviews. Scientific American, 269(5):116–118, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n5/pdf/scientificamerican1193-116.pdf.

Anonymous:1993:BRn

[1694] Anonymous. Book reviews. Scientific American, 269(5):116–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Russett:1993:EPA

[1695] Bruce Russett. Essay: Peace among democracies. Scientific American, 269(5):120, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n5/pdf/scientificamerican1193-
120.pdf.

Russett:1993:E

[1696] Bruce Russett. Essay. Scientific American, 269(5):120–??, November 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:Ee

[1697] Anonymous. Erratum. Scientific American, 269(6):10, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n6/pdf/scientificamerican1293-10b.pdf.

Anonymous:1993:LEo

[1698] Anonymous. Letters to the Editors. Scientific American, 269(6):10, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-10a.pdf.

Anonymous:1993:LEp

[1699] Anonymous. Letters to the editors. Scientific American, 269(6):10–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:YAIl

[1700] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 269(6):14, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n6/pdf/scientificamerican1293-
14.pdf.

Anonymous:1993:YAl

[1701] Anonymous. 50 and 100 years ago. Scientific American, 269(6):14–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Piel:1993:SYE

[1702] Jonathan Piel. Special year-end section: Challenges for 1994. Scientific American, 269(6):16, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n6/pdf/scientificamerican1293-
16a.pdf.

Horgan:1993:HM

[1703] John Horgan. Heart of the matter. Scientific American, 269(6):16–17, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-16b.pdf.

Anonymous:1993:C

[1704] Anonymous. Challenges for 1994. Scientific American, 269(6):16–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1993:MM

[1705] Tim Beardsley. From mice to men. Scientific American, 269(6):18–22, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-18.pdf.

Wallich:1993:BF

[1706] Paul Wallich. Bright future. Scientific American, 269(6):22–24, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-22.pdf.

Yam:1993:CS

[1707] Philip Yam. The color of sound. Scientific American, 269(6):24–26, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-24.pdf.

Stix:1993:CD

[1708] Gary Stix. Clean definitions. Scientific American, 269(6):26–30, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-26.pdf.

Gibbs:1993:D

[1709] W. Wayt Gibbs. Deliverance. Scientific American, 269(6):30–33, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-30.pdf.

Holloway:1993:CQ

[1710] Marguerite Holloway. Core questions. Scientific American, 269(6):34–36, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-34.pdf.

Horgan:1993:FF

[1711] John Horgan. Fractured functions. Scientific American, 269(6):36–37, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-36.pdf.

Holloway:1993:LL

[1712] Marguerite Holloway. Living legend. Scientific American, 269(6):40, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-40a.pdf.

Leutwyler:1993:GBF

[1713] Kristin Leutwyler. A girl’s best friend. Scientific American, 269(6):40–44, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-40b.pdf.

Gibbs:1993:WCD

[1714] W. Wayt Gibbs. When cells divide. Scientific American, 269(6):44–45, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-44.pdf.

Stix:1993:SSb

[1715] Gary Stix. Shrinking sandbox. Scientific American, 269(6):45–46, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-45.pdf.

Beardsley:1993:JM

[1716] Tim Beardsley. A Joycean mutation. Scientific American, 269(6):47, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-47a.pdf.

Horgan:1993:WDR

[1717] John Horgan. Wanted: a defense R&D policy. Scientific American, 269(6):47–50, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n6/pdf/scientificamerican1293-
47b.pdf.

Powell:1993:CS

[1718] Corey S. Powell. Cosmic SNUs. Scientific American, 269(6):50, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-50.pdf.

Robey:1993:FDD

[1719] Bryant Robey, Shea O. Rutstein, and Leo Morris. The fertility decline in developing countries. Scientific American, 269(6):60–67, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n6/pdf/scientificamerican1293-60.pdf.

Gehrels:1993:CGR

[1720] Neil Gehrels, Carl E. Fichtel, Gerald J. Fishman, James D. Kurfess, and Volker Schönfelder. The Compton Gamma Ray Observatory. Scientific American, 269(6):68–77, December 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n6/pdf/scientificamerican1293-68.pdf.

Klein:1993:MPH

[1721] Jan Klein, Naoyuki Takahata, and Francisco J. Ayala. MHC polymorphism and human origins. Scientific American, 269(6):78–83, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-78.pdf.

Rinderer:1993:ABU

[1722] Thomas E. Rinderer, Benjamin P. Oldroyd, and Walter S. Sheppard. Africanized bees in the U.S. Scientific American, 269(6):84–90, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-84.pdf.

Yam:1993:TSCa

[1723] Philip Yam. Trends in superconductivity: Current events. Scientific American, 269(6):84–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bugg:1993:DD

[1724] Charles E. Bugg, William M. Carson, and John A. Montgomery. Drugs by design. Scientific American, 269(6):92–98, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n6/pdf/scientificamerican1293-92.pdf.

Garver:1993:EF

[1725] Wayne Garver and Frank Moss. Electronic fireflies. Scientific American, 269(6):94–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Strogatz:1993:COB

[1726] Steven H. Strogatz and Ian Stewart. Coupled oscillators and biological synchronization. Scientific American, 269(6):102–109, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n6/pdf/scientificamerican1293-102.pdf.

Malone:1993:DCP

[1727] Caroline Malone, Anthony Bonanno, Tancred Gouder, Simon Stoddart, and David Trump. The death cults of prehistoric Malta. Scientific American, 269(6):110–117, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n6/pdf/scientificamerican1293-
110.pdf.

Yam:1993:TSCb

[1728] Philip Yam. Trends in superconductivity: Current events. Scientific American, 269(6):118–126, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v269/n6/pdf/scientificamerican1293-
118.pdf.

Yam:1993:CE

[1729] Philip Yam. Current events. Scientific American, 269(6):118–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Garver:1993:ASE

[1730] Wayne Garver and Frank Moss. The amateur scientist: Electronic fireflies. Scientific American, 269(6):128–130, December 1993. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v269/
n6/pdf/scientificamerican1293-128.pdf.

Anonymous:1993:ASf

[1731] Anonymous. Amateur scientist. Scientific American, 269(6):128–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:BRo

[1732] Anonymous. Book reviews. Scientific American, 269(6):132–138, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-132.pdf.

Anonymous:1993:BRp

[1733] Anonymous. Book reviews. Scientific American, 269(6):132–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1993:AI

[1734] Anonymous. Annual index 1993. Scientific American, 269(6):140–142, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v269/n6/pdf/scientificamerican1293-140.pdf.

Eisenberg:1993:EES

[1735] Anne Eisenberg. Essay: Eponymous science. Scientific American, 269(6):144, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n6/pdf/scientificamerican1293-
144.pdf.

Eisenberg:1993:Ec

[1736] Anne Eisenberg. Essay. Scientific American, 269(6):144–??, December 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:LEa

[1737] Anonymous. Letters to the Editors. Scientific American, 270(1):9, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-9.pdf.

Anonymous:1994:YAIa

[1738] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 270(1):12, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
12.pdf.

Anonymous:1994:YAa

[1739] Anonymous. 50 and 100 years ago. Scientific American, 270(1):12–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:SCJ

[1740] Corey S. Powell. Science and the citizen: Joe Btfsplk. Scientific American, 270(1):17–18, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
17.pdf.

Anonymous:1994:SCa

[1741] Anonymous. Science and the citizen. Scientific American, 270(1):17–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1994:GNR

[1742] Philip Yam. Getting a new rise out of superconductors. Scientific American, 270(1):18, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
18.pdf.

Powell:1994:EPH

[1743] Corey S. Powell. “EQ, phone home”. Scientific American, 270(1):19, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-19.pdf.

Powell:1994:DM

[1744] Corey S. Powell. A dark matter. Scientific American, 270(1):20, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-20.pdf.

Horgan:1994:BW

[1745] John Horgan. Biowarfare wars. Scientific American, 270(1):22, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-22.pdf.

Leutwyler:1994:SC

[1746] Kristin Leutwyler. Something to chew on. Scientific American, 270(1):24, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-24b.pdf.

Leutwyler:1994:CT

[1747] Kristin Leutwyler. Chiller thriller. Scientific American, 270(1):24–25, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-24a.pdf.

Holloway:1994:DI

[1748] Marguerite Holloway. Dioxin indictment. Scientific American, 270(1):25, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-25.pdf.

Guterl:1994:KVG

[1749] Fred Guterl. Keyhole view of a genius. Scientific American, 270(1):26–28, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-26.pdf.

Kusler:1994:W

[1750] Jon A. Kusler, William J. Mitsch, and Joseph S. Larson. Wetlands. Scientific American, 270(1):64–70, January 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n1/pdf/scientificamerican0194-64B.pdf.

Crawford:1994:SSM

[1751] Henry J. Crawford and Carsten H. Greiner. The search for strange matter. Scientific American, 270(1):72–77, January 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n1/pdf/scientificamerican0194-72.pdf.

Carmichael:1994:TC

[1752] Wayne W. Carmichael. The toxins of cyanobacteria. Scientific American, 270(1):78–86, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
78.pdf.

Traub:1994:BI

[1753] Joseph F. Traub and Henryk Woźniakowski. Breaking intractability. Scientific American, 270(1):102–107, January 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n1/pdf/scientificamerican0194-102.pdf.

Crews:1994:AS

[1754] David Crews. Animal sexuality. Scientific American, 270(1):108–114, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-108.pdf.

Renfrew:1994:WLD

[1755] Colin Renfrew. World linguistic diversity. Scientific American, 270(1):116–123, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
116.pdf.

Beardsley:1994:TCEa

[1756] Tim Beardsley. Trends in cancer epidemiology: A war not won. Scientific American, 270(1):118–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holzmann:1994:FDN

[1757] Gerard J. Holzmann and Björn Pehrson. The first data networks. Scientific American, 270(1):124–129, January 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n1/pdf/scientificamerican0194-124.pdf.

Beardsley:1994:TCEb

[1758] Tim Beardsley. Trends in cancer epidemiology: a war not won. Scientific American, 270(1):130–138, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
130.pdf.

Beardsley:1994:WW

[1759] Tim Beardsley. A war not won. Scientific American, 270(1):130–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1994:KLV

[1760] Ian Stewart. Knots, links and videotape. Scientific American, 270(1): 134–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SBR

[1761] Gary Stix. Science and business: RoboTuna. Scientific American, 270(1):142, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
142.pdf.

Anonymous:1994:SBa

[1762] Anonymous. Science and business. Scientific American, 270(1):142–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Patton:1994:ST

[1763] Robert Patton. Survival tactics. Scientific American, 270(1):144–146, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-144.pdf.

Stix:1994:MA

[1764] Gary Stix. Material advantage. Scientific American, 270(1):146–147, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-146.pdf.

Leutwyler:1994:OT

[1765] Kristin Leutwyler. Optical tomography. Scientific American, 270 (1):147–149, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
147.pdf.

Stix:1994:GR

[1766] Gary Stix. Gene readers. Scientific American, 270(1):149–150, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-149.pdf.

Wallich:1994:LDL

[1767] Paul Wallich. Lies, damned lies and models. Scientific American, 270(1):151, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
151.pdf.

Anonymous:1994:MRK

[1768] Anonymous. Mathematical recreations: Knots, links and videotape. Scientific American, 270(1):152–154, January 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n1/pdf/scientificamerican0194-152.pdf.

Anonymous:1994:MRa

[1769] Anonymous. Mathematical recreation. Scientific American, 270(1):152– ??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRa

[1770] Philip Morrison. Book reviews. Scientific American, 270(1):155–158, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n1/pdf/scientificamerican0194-155.pdf.

Anonymous:1994:BRa

[1771] Anonymous. Book reviews. Scientific American, 270(1):155–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Monbiot:1994:ETE

[1772] George Monbiot. Essay: The tragedy of enclosure. Scientific American, 270(1):159, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n1/pdf/scientificamerican0194-
159.pdf.

Monbiot:1994:E

[1773] George Monbiot. Essay. Scientific American, 270(1):159–??, January 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Ea

[1774] Anonymous. Errata. Scientific American, 270(2):10, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n2/pdf/scientificamerican0294-10b.pdf.

Anonymous:1994:LEb

[1775] Anonymous. Letters to the Editors. Scientific American, 270(2):10, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-10a.pdf.

Anonymous:1994:YAIb

[1776] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 270(2):12, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
12.pdf.

Anonymous:1994:YAb

[1777] Anonymous. 50 and 100 years ago. Scientific American, 270(2):12–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:SCN

[1778] Corey S. Powell. Science and the citizen: Nobel notes. Scientific American, 270(2):16–20, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
16.pdf.

Anonymous:1994:SCb

[1779] Anonymous. Science and the citizen. Scientific American, 270(2):16–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1994:RQW

[1780] Philip Yam. Reflections in a quantum well. Scientific American, 270(2):20–21, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
20.pdf.

Rennie:1994:TLT

[1781] John Rennie. Too little, too late? Scientific American, 270(2):21–24, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-21.pdf.

Leutwyler:1994:NGW

[1782] Kristin Leutwyler. No global warming? Scientific American, 270(2):24, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-24.pdf.

Gibbs:1994:FG

[1783] W. Wayt Gibbs and Tim Beardsley. Fertile ground. Scientific American, 270(2):26–28, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
26.pdf.

Beardsley:1994:DL

[1784] Tim Beardsley. Design for living. Scientific American, 270(2):28–32, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-28.pdf.

Mukerjee:1994:TTB

[1785] Madhusree Mukerjee. Time-trippers beware. Scientific American, 270(2):32, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
32.pdf.

Beardsley:1994:LBL

[1786] Tim Beardsley. Laid-back leader rattles the Academy. Scientific American, 270(2):34–35, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
34.pdf.

Morrison:1994:FAD

[1787] Philip Morrison, Kosta Tsipis, and Jerome Wiesner. The future of American defense. Scientific American, 270(2):38–45, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n2/pdf/scientificamerican0294-38.pdf.

Charlson:1994:SAC

[1788] Robert J. Charlson and Tom M. L. Wigley. Sulfate aerosol and climatic change. Scientific American, 270(2):48–57, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n2/pdf/scientificamerican0294-48.pdf.

McGinnis:1994:MAB

[1789] William McGinnis and Michael Kuziora. The molecular architects of body design. Scientific American, 270(2):58–66, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n2/pdf/scientificamerican0294-58.pdf.

Mitchell:1994:SPW

[1790] William J. Mitchell. Science in pictures: When is seeing believing? Scientific American, 270(2):68–73, February 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n2/pdf/scientificamerican0294-68.pdf.

Mitchell:1994:WSB

[1791] William J. Mitchell. When is seeing believing? Scientific American, 270 (2):68–?? (Intl. ed. 44–??), February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:TPPa

[1792] John Horgan. Trends in physics: Particle metaphysics. Scientific American, 270(2):70–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Borra:1994:LM

[1793] Ermanno F. Borra. Liquid mirrors. Scientific American, 270(2):76–81, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-76.pdf.

DesJarlais:1994:AUI

[1794] Don C. Des Jarlais and Samuel R. Friedman. AIDS and the use of injected drugs. Scientific American, 270(2):82–88 (Intl. ed. 56–??), February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-82.pdf.

Dragovan:1994:MMS

[1795] Mark Dragovan and Don Alvarez. Making a mirror by spinning a liquid. Scientific American, 270(2):86–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Marshall:1994:TBS

[1796] Larry G. Marshall. The terror birds of South America. Scientific American, 270(2):90–95, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
90.pdf.

Horgan:1994:TPPb

[1797] John Horgan. Trends in physics: Particle metaphysics. Scientific American, 270(2):96–106, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
96.pdf.

Horgan:1994:PM

[1798] John Horgan. Particle metaphysics. Scientific American, 270(2):96–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SBD

[1799] Gary Stix. Science and business: Dr. Big Brother. Scientific American, 270(2):108–110, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
108.pdf.

Anonymous:1994:SBb

[1800] Anonymous. Science and business. Scientific American, 270(2):108–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:EE

[1801] Gary Stix. Extra! Extra! Scientific American, 270(2):110–111, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-110.pdf.

Gibbs:1994:HC

[1802] W. Wayt Gibbs. Heads in the cloud. Scientific American, 270(2):112–113, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-112.pdf.

Stix:1994:PD

[1803] Gary Stix. Pipe dream. Scientific American, 270(2):113–114, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-113.pdf.

Stix:1994:BG

[1804] Gary Stix. A blade of grass. Scientific American, 270(2):114, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-114.pdf.

London:1994:PLE

[1805] Fred Guterl London. Paying for light at the end of the chunnel. Scientific American, 270(2):115, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
115.pdf.

Dragovan:1994:ASM

[1806] Mark Dragovan and Don Alvarez. The amateur scientist: Making a mirror by spinning a liquid. Scientific American, 270(2):116–117, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-116.pdf.

Anonymous:1994:ASa

[1807] Anonymous. The amateur scientist. Scientific American, 270(2):116–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRb

[1808] Philip Morrison. Book reviews. Scientific American, 270(2):118–123, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n2/pdf/scientificamerican0294-118.pdf.

Anonymous:1994:BRb

[1809] Anonymous. Book reviews. Scientific American, 270(2):118–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Piel:1994:EAP

[1810] Gerard Piel. Essay: AIDS and population “control”. Scientific American, 270(2):124, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n2/pdf/scientificamerican0294-
124.pdf.

Anonymous:1994:EGP

[1811] Anonymous. Essay: Gerard Piel. Scientific American, 270(2):124–??, February 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Eb

[1812] Anonymous. Errata. Scientific American, 270(3):10, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-10b.pdf.

Anonymous:1994:LEc

[1813] Anonymous. Letters to the Editors. Scientific American, 270(3):10, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-10a.pdf.

Anonymous:1994:YAIc

[1814] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 270(3):12, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
12.pdf.

Anonymous:1994:YAc

[1815] Anonymous. 50 and 100 years ago. Scientific American, 270(3):12–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1994:SCI

[1816] Tim Beardsley. Science and the citizen: Image enhancement. Scientific American, 270(3):14–18, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
14.pdf.

Anonymous:1994:SCc

[1817] Anonymous. Science and the citizen. Scientific American, 270(3):14–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1994:CC

[1818] Philip Yam. Chaotic chaos. Scientific American, 270(3):16, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-16.pdf.

Leutwyler:1994:G

[1819] Kristin Leutwyler. Down the green. Scientific American, 270(3):18–19, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-18.pdf.

Mukerjee:1994:S

[1820] Madhusree Mukerjee. Spinning out. Scientific American, 270(3):19–20, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-19.pdf.

Beardsley:1994:MM

[1821] Tim Beardsley. Molecular mischief. Scientific American, 270(3):20–21, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-20.pdf.

Rennie:1994:GRC

[1822] John Rennie. Gene rich, cash poor. Scientific American, 270(3):21–22, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-21.pdf.

Powell:1994:CC

[1823] Corey S. Powell. Cold confusion. Scientific American, 270(3):22–28, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-22.pdf.

Horgan:1994:FTF

[1824] John Horgan. Fermat’s Theorem fights back. Scientific American, 270(3): 28, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-28.pdf.

Horgan:1994:CFL

[1825] John Horgan. Confronting the final limit. Scientific American, 270(3):32– 33, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-32.pdf.

Bongaarts:1994:CGH

[1826] John Bongaarts. Can the growing human population feed itself? Scientific American, 270(3):36–42, March 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-36.pdf.

Bonatti:1994:EMB

[1827] Enrico Bonatti. The Earth’s mantle below the oceans. Scientific American, 270(3):44–51, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
44.pdf.

Capecchi:1994:TGR

[1828] Mario R. Capecchi. Targeted gene replacement. Scientific American, 270(3):52–59, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
52.pdf.

Meyerson:1994:HSS

[1829] Bernard S. Meyerson. High-speed silicon germanium electronics. Scientific American, 270(3):62–67, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
62.pdf.

Deutsch:1994:QPT

[1830] David Deutsch and Michael Lockwood. The quantum physics of time travel. Scientific American, 270(3):68–74, March 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-68.pdf.

Wallich:1994:TCWa

[1831] Paul Wallich. Trends in communications: Wire pirates. Scientific American, 270(3):72–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Glance:1994:DSD

[1832] Natalie S. Glance and Bernardo A. Huberman. The dynamics of social dilemmas. Scientific American, 270(3):76–81, March 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-76.pdf.

McClanahan:1994:FTD

[1833] Lon L. McClanahan, Rodolfo Ruibal, and Vaughan H. Shoemaker. Frogs and toads in deserts. Scientific American, 270(3):82–88, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-82.pdf.

Stewart:1994:NMB

[1834] Ian Stewart. The new merology of beastly numbers. Scientific American, 270(3):88–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wallich:1994:TCWb

[1835] Paul Wallich. Trends in communications: Wire pirates. Scientific American, 270(3):90–101, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
90.pdf.

Wallich:1994:WP

[1836] Paul Wallich. Wire pirates. Scientific American, 270(3):90–?? (Intl. ed. 72–??), March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SBM

[1837] Gary Stix. Science and business: Making money. Scientific American, 270(3):102–105, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
102.pdf.

Anonymous:1994:SBc

[1838] Anonymous. Science and business. Scientific American, 270(3):102–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1994:IS

[1839] W. Wayt Gibbs. Inside story. Scientific American, 270(3):104–105, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-104.pdf.

Stix:1994:LF

[1840] Gary Stix. The last frontier. Scientific American, 270(3):105–106, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-105.pdf.

Gibbs:1994:KC

[1841] W. Wayt Gibbs. King Cotton. Scientific American, 270(3):106–107, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-106.pdf.

Rennie:1994:BTR

[1842] John Rennie. Biocatalysts turn rings around the competition. Scientific American, 270(3):107, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
107.pdf.

Leutwyler:1994:PV

[1843] Kristin Leutwyler. Prosthetic vision. Scientific American, 270(3):108, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-108.pdf.

Stix:1994:BSB

[1844] Gary Stix and Paul Wallich. Is bigger still better? Scientific American, 270(3):109, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n3/pdf/scientificamerican0394-
109.pdf.

Stewart:1994:MRN

[1845] Ian Stewart. Mathematical recreations: The new merology of beastly numbers. Scientific American, 270(3):110–113, March 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-110.pdf.

Anonymous:1994:MRb

[1846] Anonymous. Mathematical recreations. Scientific American, 270(3):110– ??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRc

[1847] Philip Morrison. Book reviews. Scientific American, 270(3):114–119, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n3/pdf/scientificamerican0394-114.pdf.

Anonymous:1994:BRc

[1848] Anonymous. Book reviews. Scientific American, 270(3):114–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zolla-Pazner:1994:EPU

[1849] Susan Zolla-Pazner. Essay: The professor, the university and industry. Scientific American, 270(3):120, March 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n3/pdf/scientificamerican0394-120.pdf.

Zolla-Pazner:1994:E

[1850] Susan Zolla-Pazner. Essay. Scientific American, 270(3):120–??, March 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:LEd

[1851] Anonymous. Letters to the Editors. Scientific American, 270(4):10, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-10.pdf.

Anonymous:1994:YAId

[1852] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 270(4):12, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
12.pdf.

Anonymous:1994:YAd

[1853] Anonymous. 50 and 100 years ago. Scientific American, 270(4):12– ??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1994:SCC

[1854] Tim Beardsley. Science and the citizen: Cool man, hot job. Scientific American, 270(4):14–18, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
14.pdf.

Anonymous:1994:SCd

[1855] Anonymous. Science and the citizen. Scientific American, 270(4):14– ??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:QCC

[1856] John Horgan. Quantum computing creeps closer to reality. Scientific American, 270(4):18, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
18.pdf.

Rennie:1994:SS

[1857] John Rennie. Super sonic. Scientific American, 270(4):20, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-20.pdf.

Horgan:1994:BYA

[1858] John Horgan. Bang! you’re alive. Scientific American, 270(4):22–24, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-22.pdf.

Holloway:1994:EI

[1859] Marguerite Holloway. An epidemic ignored. Scientific American, 270(4):24–27, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
24.pdf.

Leutwyler:1994:SS

[1860] Kristin Leutwyler. Silly season. Scientific American, 270(4):28–29, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-28.pdf.

Yam:1994:SP

[1861] Philip Yam. Super progress. Scientific American, 270(4):29–30, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-29.pdf.

Horgan:1994:ROB

[1862] John Horgan. Revisiting old battlefields. Scientific American, 270(4):36– 41, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-36.pdf.

Krugman:1994:TJW

[1863] Paul R. Krugman and Robert Z. Lawrence. Trade, jobs and wages. Scientific American, 270(4):44–49, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
44.pdf.

Brown:1994:CSD

[1864] Stuart Brown and George Grüner. Charge and spin density waves. Scientific American, 270(4):50–56, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
50.pdf.

Raichle:1994:VM

[1865] Marcus E. Raichle. Visualizing the mind. Scientific American, 270(4):58– 64, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-58.pdf.

Kurti:1994:CPK

[1866] Nicholas Kurti and Hervé This-Benckhard. Chemistry and physics in the kitchen. Scientific American, 270(4):66–71, April 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-66.pdf.

Garnick:1994:DPC

[1867] Marc B. Garnick. The dilemmas of prostate cancer. Scientific American, 270(4):72–81, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
72.pdf.

Holloway:1994:TBRa

[1868] Marguerite Holloway. Trends in biological restoration: Nurturing nature. Scientific American, 270(4):76–??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shimada:1994:PMO

[1869] Izumi Shimada and Jo Ann Griffin. Precious metal objects of the middle sicán. Scientific American, 270(4):82–89, April 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-82.pdf.

Luhmann:1994:PMV

[1870] Janet G. Luhmann, James B. Pollack, and Lawrence Colin. The Pioneer mission to Venus. Scientific American, 270(4):90–97, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-90.pdf.

Kurti:1994:KL

[1871] Nicholas Kurti and Herve This-Benckhard. The kitchen as a lab. Scientific American, 270(4):94–??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1994:TBRb

[1872] Marguerite Holloway. Trends in biological restoration: Nurturing nature. Scientific American, 270(4):98–108, April 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-98.pdf.

Holloway:1994:NN

[1873] Marguerite Holloway. Nurturing nature. Scientific American, 270(4):98– ??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1994:SBM

[1874] W. Wayt Gibbs. Science and business: Mirror, mirror. Scientific American, 270(4):110–112, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
110.pdf.

Anonymous:1994:SBd

[1875] Anonymous. Science and business. Scientific American, 270(4):110– ??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:SQ

[1876] Corey S. Powell. Shaking quakes. Scientific American, 270(4):112–113, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-112.pdf.

Horgan:1994:TBT

[1877] John Horgan. The threat of buried thrust faults. Scientific American, 270(4):113, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
113.pdf.

Gibbs:1994:BS

[1878] W. Wayt Gibbs. Basic strategies. Scientific American, 270(4):114–115, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-114.pdf.

Beardsley:1994:HP

[1879] Tim Beardsley. Hot property. Scientific American, 270(4):115, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-115.pdf.

Patton:1994:S

[1880] Robert Patton. Softwars. Scientific American, 270(4):116, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-116.pdf.

Wallich:1994:HHI

[1881] Paul Wallich. Are hospitals heading for intensive care? Scientific American, 270(4):118, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
118.pdf.

Kurti:1994:ASK

[1882] Nicholas Kurti and Hervé This-Benckhard. The amateur scientist: The kitchen as a lab. Scientific American, 270(4):120–123, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n4/pdf/scientificamerican0494-120.pdf.

Anonymous:1994:ASb

[1883] Anonymous. The amateur scientist. Scientific American, 270(4):120– ??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRd

[1884] Philip Morrison. Book reviews. Scientific American, 270(4):124–127, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n4/pdf/scientificamerican0494-124.pdf.

Anonymous:1994:BRd

[1885] Anonymous. Book reivews. Scientific American, 270(4):124–??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1994:EMN

[1886] Anne Eisenberg. Essay: E-mail and the new Epistolary Age. Scientific American, 270(4):128, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n4/pdf/scientificamerican0494-
128.pdf.

Eisenberg:1994:Ea

[1887] Anne Eisenberg. Essay. Scientific American, 270(4):128–??, April 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:LEe

[1888] Anonymous. Letters to the Editors. Scientific American, 270(5):10, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-10.pdf.

Anonymous:1994:YAIe

[1889] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 270(5):12, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
12.pdf.

Anonymous:1994:YAe

[1890] Anonymous. 50 and 100 years ago. Scientific American, 270(5):12– ??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:SCW

[1891] John Horgan. Science and the citizen: When treaties fail. Scientific American, 270(5):14–20, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
14.pdf.

Anonymous:1994:SCe

[1892] Anonymous. Science and the citizen. Scientific American, 270(5):14– ??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:NE

[1893] John Horgan. Neural eavesdropping. Scientific American, 270(5):16, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-16.pdf.

Beardsley:1994:LO

[1894] Tim Beardsley. The lucky ones. Scientific American, 270(5):20–25, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-20.pdf.

Beardsley:1994:TP

[1895] Tim Beardsley. Take the pain? Scientific American, 270(5):25–26, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-25.pdf.

Powell:1994:UL

[1896] Corey S. Powell. Unbearable lightness. Scientific American, 270(5):27– 28, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-27.pdf.

Holloway:1994:FM

[1897] Marguerite Holloway. Family matters. Scientific American, 270(5):28– 32, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-28.pdf.

Mukerjee:1994:VV

[1898] Madhusree Mukerjee. Vishnu violated. Scientific American, 270(5):32– 34, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-32.pdf.

Leutwyler:1994:AG

[1899] Kristin Leutwyler. Amazing grace. Scientific American, 270(5):34, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-34.pdf.

Yam:1994:FRF

[1900] Philip Yam. Fundamental rights, fundamental physics. Scientific American, 270(5):39–40, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
39.pdf.

Anonymous:1994:HBI

[1901] Anonymous. Is homosexuality biologically influenced? Scientific American, 270(5):43, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
43.pdf.

LeVay:1994:EBI

[1902] Simon LeVay and Dean H. Hamer. Evidence for a biological influence in male homosexuality. Scientific American, 270(5):44–49, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n5/pdf/scientificamerican0594-44.pdf.

Byne:1994:BEC

[1903] William Byne. The biological evidence challenged. Scientific American, 270(5):50–55, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
50.pdf.

Albert:1994:BAQ

[1904] David Z. Albert. Bohm’s alternative to quantum mechanics. Scientific American, 270(5):58–67, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
58.pdf.

Johnson:1994:HIF

[1905] Howard M. Johnson, Fuller W. Bazer, Brian E. Szente, and Michael A. Jarpe. How interferons fight disease. Scientific American, 270(5):68– 75, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-68.pdf.

Stix:1994:TATa

[1906] Gary Stix. Trends in air transportation: Aging airways. Scientific American, 270(5):70–??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Miller:1994:SPC

[1907] Ron Miller. Science in pictures: Chesley Bonestell’s astronomical visions. Scientific American, 270(5):76–81, May 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n5/pdf/scientificamerican0594-76.pdf.

Miller:1994:CBA

[1908] Ron Miller. Chesley Bonestell’s astronomical visions. Scientific American, 270(5):76–?? (Intl. ed. 50–??), May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cooper:1994:DD

[1909] George A. Cooper. Directional drilling. Scientific American, 270(5): 82–87 (Intl. ed. 56–??), May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
82.pdf.

Coppens:1994:ESS

[1910] Yves Coppens. East Side story: The origin of humankind. Scientific American, 270(5):88–95, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
88.pdf.

Stewart:1994:HMG

[1911] Ian Stewart. How many guards in the gallery? Scientific American, 270(5):88–??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:AA

[1912] Gary Stix. Aging airways. Scientific American, 270(5):96–104 (Intl. ed. 70–78), May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:TATb

[1913] Gary Stix. Trends in air transportation: Aging airways. Scientific American, 270(5):96–104, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
96.pdf.

Beardsley:1994:SBW

[1914] Tim Beardsley. Science and business: What’s bred in the bone. Scientific American, 270(5):106–108, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
106.pdf.

Anonymous:1994:SBe

[1915] Anonymous. Science and business. Scientific American, 270(5):106– ??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:CC

[1916] John Horgan. Cold calculation. Scientific American, 270(5):108, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-108b.pdf.

Beardsley:1994:DD

[1917] Tim Beardsley. Devilish details. Scientific American, 270(5):108–110, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-108a.pdf.

Patton:1994:OR

[1918] Robert Patton. Office romance. Scientific American, 270(5):110–112, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-110.pdf.

Leutwyler:1994:A

[1919] Kristin Leutwyler. As advertised. Scientific American, 270(5):112, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-112.pdf.

Stix:1994:LL

[1920] Gary Stix. Lollipop, lollipop. Scientific American, 270(5):113, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n5/pdf/scientificamerican0594-113a.pdf.

Gibbs:1994:PC

[1921] W. Wayt Gibbs and Gary Stix. The price of clean. Scientific American, 270(5):113–114, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
113b.pdf.

Gibbs:1994:GM

[1922] W. Wayt Gibbs. Gray matter. Scientific American, 270(5):114, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-114.pdf.

Holloway:1994:HGM

[1923] Marguerite Holloway and Paul Wallich. How green is my label? Scientific American, 270(5):115, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
115.pdf.

Stewart:1994:MRH

[1924] Ian Stewart. Mathematical recreations: How many guards in the gallery? Scientific American, 270(5):118–120, May 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n5/pdf/scientificamerican0594-118.pdf.

Anonymous:1994:MRc

[1925] Anonymous. Mathematical recreations. Scientific American, 270(5):118– ??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRe

[1926] Philip Morrison. Book reviews. Scientific American, 270(5):122–127, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n5/pdf/scientificamerican0594-122.pdf.

Anonymous:1994:BRe

[1927] Anonymous. Book reviews. Scientific American, 270(5):122–??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Weisskopf:1994:EES

[1928] Victor F. Weisskopf. Essay: Endangered support of basic science. Scientific American, 270(5):128, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n5/pdf/scientificamerican0594-
128.pdf.

Weisskopf:1994:E

[1929] Victor F. Weisskopf. Essay. Scientific American, 270(5):128–??, May 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Ec

[1930] Anonymous. Errata. Scientific American, 270(6):10, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-10b.pdf.

Anonymous:1994:LEf

[1931] Anonymous. Letters to the Editors. Scientific American, 270(6):10, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-10a.pdf.

Anonymous:1994:YAIf

[1932] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 270(6):12, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
12.pdf.

Anonymous:1994:YAf

[1933] Anonymous. 50 and 100 years ago. Scientific American, 270(6):12– ??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1994:SCP

[1934] Marguerite Holloway. Science and the citizen: Population summit. Scientific American, 270(6):14–16, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
14.pdf.

Anonymous:1994:SCf

[1935] Anonymous. Science and the citizen. Scientific American, 270(6):14– ??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mukerjee:1994:GS

[1936] Madhusree Mukerjee. Gathering string. Scientific American, 270(6):16, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-16.pdf.

Powell:1994:SC

[1937] Corey S. Powell. Sanity check. Scientific American, 270(6):19–22, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-19.pdf.

Leutwyler:1994:BS

[1938] Kristin Leutwyler. Bright spot. Scientific American, 270(6):22, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-22.pdf.

Beardsley:1994:R

[1939] Tim Beardsley. La ronde. Scientific American, 270(6):26–29, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-26.pdf.

Yam:1994:JP

[1940] Philip Yam. It’s just a phase. Scientific American, 270(6):29, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-29.pdf.

Beardsley:1994:SR

[1941] Tim Beardsley. Shooting the rapids. Scientific American, 270(6):32, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-32.pdf.

Horgan:1994:LUM

[1942] John Horgan. The last universal mathematician. Scientific American, 270(6):33–34, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
33.pdf.

Logsdon:1994:WRM

[1943] John M. Logsdon and Alain Dupas. Was the race to the Moon real? Scientific American, 270(6):36–43, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
36.pdf.

Nauenberg:1994:CLA

[1944] Michael Nauenberg, Carlos Stroud, and John Yeazell. The classical limit of an atom. Scientific American, 270(6):44–49, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-44.pdf.

LeDoux:1994:EMB

[1945] Joseph E. LeDoux. Emotion, memory and the brain. Scientific American, 270(6):50–57, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
50.pdf.

Hardy:1994:AO

[1946] John W. Hardy. Adaptive optics. Scientific American, 270(6):60–65, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-60.pdf.

Pozorski:1994:EAC

[1947] Shelia Pozorski and Thomas Pozorski. Early Andean cities. Scientific American, 270(6):66–72, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
66.pdf.

Rennie:1994:TGGa

[1948] John Rennie. Trends in genetics: Grading the gene tests. Scientific American, 270(6):66–??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kirchner:1994:SBH

[1949] Wolfgang H. Kirchner and William F. Towne. The sensory basis of the honeybee’s dance language. Scientific American, 270(6):74–80, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-74.pdf.

Cox:1994:EAD

[1950] Paul Alan Cox and Michael J. Balick. The ethnobotanical approach to drug discovery. Scientific American, 270(6):82–87, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-82.pdf.

Iovine:1994:GAE

[1951] John Iovine. Genetically altering Escherichia Coli. Scientific American, 270(6):84–??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1994:TGGb

[1952] John Rennie. Trends in genetics: Grading the gene tests. Scientific American, 270(6):88–97, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
88.pdf.

Rennie:1994:GGT

[1953] John Rennie. Grading the gene tests. Scientific American, 270(6):88– ??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SBC

[1954] Gary Stix. Science and business: Cyberspace cadets. Scientific American, 270(6):98–101, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
98.pdf.

Anonymous:1994:SBf

[1955] Anonymous. Science and business. Scientific American, 270(6):98–??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:GBT

[1956] Gary Stix. Good-bye, Taki 183. Scientific American, 270(6):100–101, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-100.pdf.

Beardsley:1994:RA

[1957] Tim Beardsley. Requiem for alpha? Scientific American, 270(6):101– 102, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-101.pdf.

Stix:1994:IL

[1958] Gary Stix. Immuno-logistics. Scientific American, 270(6):102–105, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-102.pdf.

Mukerjee:1994:LW

[1959] Madhusree Mukerjee. Little winners. Scientific American, 270(6):105– 106, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-105.pdf.

Wallich:1994:SOC

[1960] Paul Wallich and Marguerite Holloway. For sale: One country, as is. Scientific American, 270(6):107, June 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-107.pdf.

Iovine:1994:ASG

[1961] John Iovine. The amateur scientist: Genetically altering Escherichia coli. Scientific American, 270(6):108–111, June 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v270/
n6/pdf/scientificamerican0694-108.pdf.

Anonymous:1994:ASc

[1962] Anonymous. The amateur scientist. Scientific American, 270(6):108– ??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRf

[1963] Philip Morrison. Book reviews. Scientific American, 270(6):112–115, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v270/n6/pdf/scientificamerican0694-112.pdf.

Anonymous:1994:BRf

[1964] Anonymous. Book reviews. Scientific American, 270(6):112–??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1994:EAS

[1965] Anne Eisenberg. Essay: The art of the scientific insult. Scientific American, 270(6):116, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v270/n6/pdf/scientificamerican0694-
116.pdf.

Eisenberg:1994:Eb

[1966] Anne Eisenberg. Essay. Scientific American, 270(6):116–??, June 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Ed

[1967] Anonymous. Erratum. Scientific American, 271(1):10, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n1/pdf/scientificamerican0794-10b.pdf.

Anonymous:1994:LEg

[1968] Anonymous. Letters to the Editors. Scientific American, 271(1):10, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-10a.pdf.

Anonymous:1994:YAIg

[1969] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 271(1):12, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
12.pdf.

Anonymous:1994:YAg

[1970] Anonymous. 50 and 100 years ago. Scientific American, 271(1):12– ??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1994:SCI

[1971] John Rennie. Science and the citizen: Immortal’s enzyme. Scientific American, 271(1):14–16, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
14.pdf.

Anonymous:1994:SCg

[1972] Anonymous. Science and the citizen. Scientific American, 271(1):14– ??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:VEP

[1973] Corey S. Powell. A visit to an exotic planet. Scientific American, 271(1): 16, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-16.pdf.

Leutwyler:1994:S

[1974] Kristin Leutwyler. Superhack. Scientific American, 271(1):17–20, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-17.pdf.

Stix:1994:MCC

[1975] Gary Stix. Managed care, circa 1300. Scientific American, 271(1):20, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-20a.pdf.

Beardsley:1994:LL

[1976] Tim Beardsley. Lethal legacy. Scientific American, 271(1):20–22, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-20b.pdf.

Beardsley:1994:DM

[1977] Tim Beardsley. Desperate measure. Scientific American, 271(1):24, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-24a.pdf.

Yam:1994:YT

[1978] Philip Yam. “You’re the top… ”. Scientific American, 271(1):24–26, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-24b.pdf.

Mukerjee:1994:SOB

[1979] Madhusree Mukerjee. Spell for old bones. Scientific American, 271(1): 28–29, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-28.pdf.

Bray:1994:ADN

[1980] Francesca Bray. Agriculture for developing nations. Scientific American, 271(1):30–37, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
30.pdf.

Taylor:1994:SLA

[1981] G. Jeffrey Taylor. The scientific legacy of Apollo. Scientific American, 271(1):40–47, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
40.pdf.

Rebek:1994:SSR

[1982] Julius Rebek, Jr. Synthetic self-replicating molecules. Scientific American, 271(1):48–55, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
48.pdf.

Jain:1994:BDD

[1983] Rakesh K. Jain. Barriers to drug delivery in solid tumors. Scientific American, 271(1):58–65, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
58.pdf.

OShea:1994:M

[1984] Thomas J. O’Shea. Manatees. Scientific American, 271(1):66–72, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-66.pdf.

Horgan:1994:TNCa

[1985] John Horgan. Trends in neuroscience: Can science explain consciousness? Scientific American, 271(1):72–??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pasteur:1994:JHF

[1986] Georges Pasteur. Jean Henri Fabré. Scientific American, 271(1):74– 80, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-74.pdf.

Knecht:1994:LIA

[1987] Heidi Knecht. Late Ice Age hunting technology. Scientific American, 271(1):82–87, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
82.pdf.

Horgan:1994:TNCb

[1988] John Horgan. Trends in neuroscience: Can science explain consciousness? Scientific American, 271(1):88–94, July 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n1/pdf/scientificamerican0794-88.pdf.

Horgan:1994:CSE

[1989] John Horgan. Can science explain consciousness? Scientific American, 271(1):88–??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1994:UAP

[1990] Ian Stewart. The ultimate in anty-particles. Scientific American, 271(1): 88–??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SBWa

[1991] Gary Stix. Science and business: The wall. Scientific American, 271(1): 96–98, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-96.pdf.

Anonymous:1994:SBg

[1992] Anonymous. Science and business. Scientific American, 271(1):96–??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1994:FL

[1993] W. Wayt Gibbs. To Fenway, with love. Scientific American, 271(1): 98, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-98.pdf.

Stix:1994:PS

[1994] Gary Stix. Private screening. Scientific American, 271(1):99–100, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-99.pdf.

PattonTokyo:1994:BV

[1995] Robert PattonTokyo. Blue video. Scientific American, 271(1):100–101, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-100.pdf.

Stix:1994:RW

[1996] Gary Stix. Roach wars. Scientific American, 271(1):101, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n1/pdf/scientificamerican0794-101.pdf.

Beardsley:1994:GE

[1997] Tim Beardsley. Green economics. Scientific American, 271(1):102, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-102.pdf.

Wallich:1994:LRW

[1998] Paul Wallich. In the long run, we all retire. Scientific American, 271(1): 103, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-103.pdf.

Stewart:1994:MRU

[1999] Ian Stewart. Mathematical recreations: The ultimate in antyparticles. Scientific American, 271(1):104–107, July 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n1/pdf/scientificamerican0794-104.pdf.

Anonymous:1994:MRd

[2000] Anonymous. Mathematical recreations. Scientific American, 271(1):104– ??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRg

[2001] Philip Morrison. Book reviews. Scientific American, 271(1):108–111, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n1/pdf/scientificamerican0794-108.pdf.

Anonymous:1994:BRS

[2002] Anonymous. Book reviews: The scientific traveler … tales of chance … Venus unrobed. Scientific American, 271(1):108–??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Adams:1994:EAS

[2003] Robert McC. Adams. Essay: Ancient Sumer, modern Iraq. Scientific American, 271(1):112, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n1/pdf/scientificamerican0794-
112.pdf.

Adams:1994:E

[2004] Robert McC Adams. Essay. Scientific American, 271(1):112–??, July 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:A

[2005] Anonymous. Amplification. Scientific American, 271(2):9, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n2/pdf/scientificamerican0894-9b.pdf.

Anonymous:1994:LEh

[2006] Anonymous. Letters to the Editors. Scientific American, 271(2):9, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-9a.pdf.

Anonymous:1994:YAIh

[2007] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 271(2):10, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
10.pdf.

Anonymous:1994:YAh

[2008] Anonymous. 50 and 100 years ago. Scientific American, 271(2):10–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:SCSa

[2009] Corey S. Powell. Science and the citizen: Star gobbler. Scientific American, 271(2):12–13, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
12.pdf.

Anonymous:1994:SCh

[2010] Anonymous. Science and the citizen. Scientific American, 271(2):12–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:DC

[2011] Corey S. Powell. Darling Clementine? Scientific American, 271(2):13, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-13.pdf.

Horgan:1994:RR

[2012] John Horgan. Radon’s risks. Scientific American, 271(2):14–16, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-14.pdf.

Holloway:1994:DB

[2013] Marguerite Holloway. Diversity blues. Scientific American, 271(2):16–18, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-16.pdf.

Beardsley:1994:RU

[2014] Tim Beardsley. The riddle of [ure3]. Scientific American, 271(2):19–20, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-19.pdf.

Horgan:1994:AO

[2015] John Horgan. Anti-omniscience. Scientific American, 271(2):20–22, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-20.pdf.

Mukerjee:1994:MMF

[2016] Madhusree Mukerjee. Missing matter found? Scientific American, 271(2):22, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
22.pdf.

Rennie:1994:DCB

[2017] John Rennie. Darwin’s current bulldog. Scientific American, 271(2):24– 25, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-24.pdf.

Revelle:1994:TWS

[2018] Daniel J. Revelle and Lora Lumpe. Third World submarines. Scientific American, 271(2):26–31, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
26.pdf.

Bowyer:1994:EUA

[2019] Stuart Bowyer. Extreme ultraviolet astronomy. Scientific American, 271(2):32–39, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
32.pdf.

Lichtman:1994:SPC

[2020] Jeff W. Lichtman. Science in pictures: Confocal microscopy. Scientific American, 271(2):40–45, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
40.pdf.

Lichtman:1994:CM

[2021] Jeff W. Lichtman. Confocal microscopy. Scientific American, 271(2):40– ?? (Intl. ed. 30–35), August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Clarke:1994:S

[2022] John Clarke. SQUIDs. Scientific American, 271(2):46–53, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n2/pdf/scientificamerican0894-46.pdf.

Engelhard:1994:HCP

[2023] Victor H. Engelhard. How cells process antigens. Scientific American, 271(2):54–61, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
54.pdf.

Anderson:1994:RT

[2024] Donald M. Anderson. Red tides. Scientific American, 271(2):62–68, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-62.pdf.

Holloway:1994:TWHa

[2025] Marguerite Holloway. Trends in women’s health: A global view. Scientific American, 271(2):66–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Molleson:1994:EBA

[2026] Theya Molleson. The eloquent bones of Abu Hureyra. Scientific American, 271(2):70–75, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
70.pdf.

Holloway:1994:TWHb

[2027] Marguerite Holloway. Trends in women’s health: a global view. Scientific American, 271(2):76–83, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
76.pdf.

Holloway:1994:GV

[2028] Marguerite Holloway. A global view. Scientific American, 271(2):76–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Becker:1994:SUS

[2029] Kyle M. Becker. Scanning underwater surfaces. Scientific American, 271 (2):80–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SBWb

[2030] Gary Stix. Science and business: Welfare plastic. Scientific American, 271(2):84–86, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
84.pdf.

Anonymous:1994:SBh

[2031] Anonymous. Science and business. Scientific American, 271(2):84–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:SME

[2032] Gary Stix. From swords to mouse ears. Scientific American, 271(2):85, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-85.pdf.

Stix:1994:AG

[2033] Gary Stix. Antigreen greenies. Scientific American, 271(2):86–87, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-86.pdf.

Leutwyler:1994:R

[2034] Kristin Leutwyler. Rotaxanes. Scientific American, 271(2):87, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-87a.pdf.

Horgan:1994:MBP

[2035] John Horgan and Robert Patton. More bits from pits. Scientific American, 271(2):87–88, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n2/pdf/scientificamerican0894-
87b.pdf.

Beardsley:1994:INC

[2036] Tim Beardsley. Illegitimi non carborundum. Scientific American, 271(2): 88, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-88.pdf.

Wallich:1994:WE

[2037] Paul Wallich. A workaholic economy. Scientific American, 271(2):89, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-89.pdf.

Becker:1994:ASS

[2038] Kyle M. Becker. The amateur scientist: Scanning underwater surfaces. Scientific American, 271(2):90–92, August 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n2/pdf/scientificamerican0894-90.pdf.

Anonymous:1994:ASd

[2039] Anonymous. The amateur scientist. Scientific American, 271(2):90–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:BRg

[2040] Anonymous. Book reviews. Scientific American, 271(2):93–95, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-93.pdf.

Anonymous:1994:BRh

[2041] Anonymous. Book reviews. Scientific American, 271(2):93–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Margulis:1994:ESD

[2042] Lynn Margulis. Essay: Sex, death and kefir. Scientific American, 271(2): 96, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n2/pdf/scientificamerican0894-96.pdf.

Margulis:1994:E

[2043] Lynn Margulis. Essay. Scientific American, 271(2):96–??, August 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Ee

[2044] Anonymous. Erratum. Scientific American, 271(3):10, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-10b.pdf.

Anonymous:1994:LEi

[2045] Anonymous. Letters to the Editors. Scientific American, 271(3):10, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-10a.pdf.

Anonymous:1994:YAIi

[2046] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 271(3):12, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
12.pdf.

Anonymous:1994:YAj

[2047] Anonymous. 50 and 100 years ago. Scientific American, 271(3):12–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:SCSb

[2048] Corey S. Powell. Science and the citizen: Super loops. Scientific American, 271(3):14, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
14a.pdf.

Powell:1994:GB

[2049] Corey S. Powell. Gone with a bang. Scientific American, 271(3):14–17, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-14b.pdf.

Anonymous:1994:SCi

[2050] Anonymous. Science and the citizen. Scientific American, 271(3):14–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leutwyler:1994:SSS

[2051] Kristin Leutwyler. Sick, sick, sick. Scientific American, 271(3):17–18, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-17.pdf.

Beardsley:1994:HA

[2052] Tim Beardsley. Hot air. Scientific American, 271(3):18–22, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-18.pdf.

Nemecek:1994:B

[2053] Sasha Nemecek. Backfire. Scientific American, 271(3):22–23, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-22.pdf.

Mukerjee:1994:LC

[2054] Madhusree Mukerjee. Lonesome cowpokes. Scientific American, 271(3):23–24, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
23.pdf.

Yam:1994:CBY

[2055] Philip Yam. Can I buy you a drink? Scientific American, 271(3):26, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-26b.pdf.

Rennie:1994:BS

[2056] John Rennie. Borrowed savagery. Scientific American, 271(3):26–27, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-26a.pdf.

Powell:1994:MSA

[2057] Corey S. Powell. A marriage of science and art. Scientific American, 271(3):28–31, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
28.pdf.

Kantor:1994:DLDa

[2058] Fred S. Kantor. Disarming Lyme disease. Scientific American, 271(3):34–39, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
34.pdf.

Duncan:1994:ABC

[2059] John L. Duncan and William F. Hosford. The aluminum beverage can. Scientific American, 271(3):34–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kantor:1994:DLDb

[2060] Fred S. Kantor. Disarming Lyme disease. Scientific American, 271(3): 34–?? (Intl. ed. 20–??), September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cline:1994:LEW

[2061] David B. Cline. Low-energy ways to observe high-energy phenomena. Scientific American, 271(3):40–47, September 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-40.pdf.

Hosford:1994:ABC

[2062] William F. Hosford and John L. Duncan. The aluminum beverage can. Scientific American, 271(3):48–53, September 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-48.pdf.

Stossel:1994:MCC

[2063] Thomas P. Stossel. The machinery of cell crawling. Scientific American, 271(3):54–63, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
54.pdf.

Green:1994:SPD

[2064] Harry W. Green, II. Solving the paradox of deep earthquakes. Scientific American, 271(3):64–71, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
64.pdf.

Cohen:1994:PPR

[2065] Linda R. Cohen and Roger G. Noll. Privatizing public research. Scientific American, 271(3):72–77, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
72.pdf.

Gibbs:1994:TCSa

[2066] W. Wayt Gibbs. Trends in computing: Software’s chronic crisis. Scientific American, 271(3):72–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gillispie:1994:SIN

[2067] Charles C. Gillispie. The scientific importance of Napoleon’s Egyptian campaign. Scientific American, 271(3):78–85, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-78.pdf.

Gibbs:1994:TCSb

[2068] W. Wayt Gibbs. Trends in computing: Software’s chronic crisis. Scientific American, 271(3):86–95, September 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-86.pdf.

Gibbs:1994:SCC

[2069] W. Wayt Gibbs. Software’s chronic crisis. Scientific American, 271(3): 86–?? (Intl. ed. 72–81), September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1994:SBT

[2070] Tim Beardsley. Science and business: Turning green. Scientific American, 271(3):96–97, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
96.pdf.

Anonymous:1994:SBi

[2071] Anonymous. Science and business. Scientific American, 271(3):96–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:BD

[2072] Gary Stix. Binary disinfectants. Scientific American, 271(3):97–98, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-97.pdf.

Beardsley:1994:HL

[2073] Tim Beardsley. Heat and light. Scientific American, 271(3):98–99, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-98.pdf.

Stix:1994:FF

[2074] Gary Stix. Food fights. Scientific American, 271(3):99–101, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-99.pdf.

Stix:1994:PM

[2075] Gary Stix. Power medicine. Scientific American, 271(3):100, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-100.pdf.

Beardsley:1994:C

[2076] Tim Beardsley. Crabshoot. Scientific American, 271(3):102, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-102.pdf.

Wallich:1994:WPP

[2077] Paul Wallich. Not worth the paper it’s printed on. Scientific American, 271(3):103, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n3/pdf/scientificamerican0994-
103.pdf.

Stewart:1994:MRS

[2078] Ian Stewart. Mathematical recreations: a subway named Turing. Scientific American, 271(3):104–107, September 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-104.pdf.

Anonymous:1994:MRe

[2079] Anonymous. Mathematical recreations. Scientific American, 271(3): 104–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRh

[2080] Philip Morrison. Book reviews. Scientific American, 271(3):108–111, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n3/pdf/scientificamerican0994-108.pdf.

Anonymous:1994:BRi

[2081] Anonymous. Book reviews. Scientific American, 271(3):108–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Davis:1994:EOP

[2082] Devra Lee Davis and Harold P. Freeman. Essay: An ounce of prevention. Scientific American, 271(3):112, September 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n3/pdf/scientificamerican0994-112.pdf.

Davis:1994:E

[2083] Devra Lee Davis and Harold Freeman. Essay. Scientific American, 271 (3):112–??, September 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Ef

[2084] Anonymous. Erratum. Scientific American, 271(4):10, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n4/pdf/scientificamerican1094-10b.pdf.

Anonymous:1994:LEj

[2085] Anonymous. Letters to the Editors. Scientific American, 271(4):10, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-10a.pdf.

Anonymous:1994:YAIj

[2086] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 271(4):12, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
12.pdf.

Anonymous:1994:YAk

[2087] Anonymous. 50 and 100 years ago. Scientific American, 271(4):12–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:SCJ

[2088] John Horgan. Science and the citizen: By Jove! Scientific American, 271(4):16–20, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
16.pdf.

Anonymous:1994:SCj

[2089] Anonymous. Science and the citizen. Scientific American, 271(4):16–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1994:ST

[2090] John Horgan. Standing tall. Scientific American, 271(4):22, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-22.pdf.

Mukerjee:1994:WN

[2091] Madhusree Mukerjee. What’s in a name? Scientific American, 271(4):26, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-26a.pdf.

Leutwyler:1994:HM

[2092] Kristin Leutwyler. A healthy mess. Scientific American, 271(4):26–31, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-26b.pdf.

Rennie:1994:FRJ

[2093] John Rennie. Fishy repair jobs. Scientific American, 271(4):31–32, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-31.pdf.

Horgan:1994:D

[2094] John Horgan. Daydreaming. Scientific American, 271(4):32–33, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-32.pdf.

Horgan:1994:HP

[2095] John Horgan. High profile. Scientific American, 271(4):33–34, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-33.pdf.

Holloway:1994:UH

[2096] Marguerite Holloway. Unearthing history. Scientific American, 271(4):37–40, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
37.pdf.

Weinberg:1994:SIL

[2097] Steven Weinberg. Special issue: Life in the Universe. Scientific American, 271(4):44–49, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
44.pdf.

Weinberg:1994:LU

[2098] Steven Weinberg. Life in the universe. Scientific American, 271(4):44–?? (Intl. ed. 22–??), October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Peebles:1994:EU

[2099] P. James E. Peebles, David N. Schramm, Edwin L. Turner, and Richard G. Kron. The evolution of the Universe. Scientific American, 271(4):52–57, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
52.pdf.

Kirshner:1994:EE

[2100] Robert P. Kirshner. The Earth’s elements. Scientific American, 271(4):58–65, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
58.pdf.

Allegre:1994:EE

[2101] Claude J. Allègre and Stephen H. Schneider. The evolution of the Earth. Scientific American, 271(4):66–75, October 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n4/pdf/scientificamerican1094-66.pdf.

Orgel:1994:OLE

[2102] Leslie E. Orgel. The origin of life on the Earth. Scientific American, 271(4):76–83, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
76.pdf.

Gould:1994:ELE

[2103] Stephen Jay Gould. The evolution of life on the Earth. Scientific American, 271(4):84–91, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
84.pdf.

Minsky:1994:WRM

[2104] Marvin Minsky. Will robots merit the Earth? Scientific American, 271 (4):86–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sagan:1994:SEL

[2105] Carl Sagan. The search for extraterrestrial life. Scientific American, 271(4):92–99, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
92.pdf.

Calvin:1994:EI

[2106] William H. Calvin. The emergence of intelligence. Scientific American, 271(4):100–107, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
100.pdf.

Minsky:1994:WRI

[2107] Marvin Minsky. Will robots inherit the Earth? Scientific American, 271(4):108–113, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
108.pdf.

Iovine:1994:BEN

[2108] John Iovine. Building an electronic neuron. Scientific American, 271 (4):110–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kates:1994:SLE

[2109] Robert W. Kates. Sustaining life on the Earth. Scientific American, 271(4):114–122, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
114.pdf.

Mukerjee:1994:SBW

[2110] Madhusree Mukerjee. Science and business: Wall Street. Scientific American, 271(4):126–128, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n4/pdf/scientificamerican1094-
126.pdf.

Anonymous:1994:SBj

[2111] Anonymous. Science and business. Scientific American, 271(4):126–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1994:R

[2112] W. Wayt Gibbs. Ready or not. Scientific American, 271(4):128–129, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-128.pdf.

Gibbs:1994:O

[2113] W. Wayt Gibbs. Opendoc. Scientific American, 271(4):130–132, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-130.pdf.

Rennie:1994:MDC

[2114] John Rennie. Making drugs count. Scientific American, 271(4):132, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-132.pdf.

Gibbs:1994:SS

[2115] W. Wayt Gibbs. State of shock. Scientific American, 271(4):133–134, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-133.pdf.

Wallich:1994:BIF

[2116] Paul Wallich. Is a bright idea flickering? Scientific American, 271(4):135, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-135.pdf.

Iovine:1994:ASB

[2117] John Iovine. The amateur scientist: Building an electronic neuron. Scientific American, 271(4):136–138, October 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n4/pdf/scientificamerican1094-136.pdf.

Anonymous:1994:ASe

[2118] Anonymous. The amateur scientist. Scientific American, 271(4):136–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRi

[2119] Philip Morrison. Book reviews. Scientific American, 271(4):140–143, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n4/pdf/scientificamerican1094-140.pdf.

Anonymous:1994:BRj

[2120] Anonymous. Book reviews. Scientific American, 271(4):140–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Damasio:1994:EDE

[2121] Antonio R. Damasio. Essay: Descartes’ error and the future of human life. Scientific American, 271(4):144, October 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n4/pdf/scientificamerican1094-144.pdf.

Damasio:1994:E

[2122] Antonio R. Damasio. Essay. Scientific American, 271(4):144–??, October 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:LEk

[2123] Anonymous. Letters to the Editors. Scientific American, 271(5):10, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-10.pdf.

Anonymous:1994:YAIk

[2124] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 271(5):12, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
12.pdf.

Anonymous:1994:YAl

[2125] Anonymous. 50 and 100 years ago. Scientific American, 271(5):12–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:SCT

[2126] Corey S. Powell. Science and the citizen: Talk about the weather. Scientific American, 271(5):14–20, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
14.pdf.

Anonymous:1994:SCk

[2127] Anonymous. Science and the citizen. Scientific American, 271(5):14–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mukerjee:1994:GAW

[2128] Madhusree Mukerjee. Global aid wars. Scientific American, 271(5):16, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-16.pdf.

Horgan:1994:SDS

[2129] John Horgan. Sex, death and sugar. Scientific American, 271(5):20–24, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-20.pdf.

Horgan:1994:BS

[2130] John Horgan. Brain storm. Scientific American, 271(5):24, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-24.pdf.

Horgan:1994:M

[2131] John Horgan. Microquasars. Scientific American, 271(5):24, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-24Ba.pdf.

Mukerjee:1994:NB

[2132] Madhusree Mukerjee. A nova burns out. Scientific American, 271(5):24, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-24Bb.pdf.

Yam:1994:B

[2133] Philip Yam. Branching out. Scientific American, 271(5):26–30, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-26.pdf.

Beardsley:1994:SLI

[2134] Tim Beardsley. Some like it hot — and cold. Scientific American, 271(5):30, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
30.pdf.

Horgan:1994:GGC

[2135] John Horgan. Gruff guru of condensed-matter physics. Scientific American, 271(5):34–35, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
34.pdf.

Moore:1994:CME

[2136] Patrick S. Moore and Claire V. Broome. Cerebrospinal meningitis epidemics. Scientific American, 271(5):38–45, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n5/pdf/scientificamerican1194-38.pdf.

Linde:1994:SRI

[2137] Andrei Linde. The self-reproducing inflationary Universe. Scientific American, 271(5):48–55, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
48.pdf.

Meyerowitz:1994:GFD

[2138] Elliot M. Meyerowitz. The genetics of flower development. Scientific American, 271(5):56–65, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
56.pdf.

Schattschneider:1994:SPE

[2139] Doris Schattschneider. Science in pictures: Escher’s metaphors. Scientific American, 271(5):66–71, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
66.pdf.

Schattschneider:1994:EM

[2140] Doris Schattschneider. Escher’s metaphors. Scientific American, 271 (5):66–?? (Intl. ed. 48–??), November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schiller:1994:SDCa

[2141] Jeffrey I. Schiller. Secure distributed computing. Scientific American, 271(5):72–76, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
72.pdf.

Beardsley:1994:TBRa

[2142] Tim Beardsley. Trends in biological research: Big-Time biology. Scientific American, 271(5):72–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schiller:1994:SDCb

[2143] Jeffrey I. Schiller. Secure distributed computing. Scientific American, 271(5):72–?? (Intl. ed. 54–58), November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Berk:1994:WCT

[2144] Laura E. Berk. Why children talk to themselves. Scientific American, 271(5):78–83, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
78.pdf.

McLaughlin:1994:RZP

[2145] William I. McLaughlin. Resolving Zeno’s paradoxes. Scientific American, 271(5):84–89, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
84.pdf.

Stewart:1994:PCG

[2146] Ian Stewart. Playing chess on a go board. Scientific American, 271(5): 88–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1994:TBRb

[2147] Tim Beardsley. Trends in biological research: Big-time biology. Scientific American, 271(5):90–97, November 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n5/pdf/scientificamerican1194-90.pdf.

Beardsley:1994:BTB

[2148] Tim Beardsley. Big-time biology. Scientific American, 271(5):90–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1994:SBF

[2149] Tim Beardsley. Science and business: Fighting for survival. Scientific American, 271(5):98–100, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
98.pdf.

Anonymous:1994:SBBa

[2150] Anonymous. Science and business: Beating swords into plowshares — badly… computers: the lean, mean layoff machines?… metering the Internet… preventing pollution for profit… lithium batteries warm up… the analytical economist: Why ain’t the experts rich? Scientific American, 271(5):98–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1994:NVS

[2151] Tim Beardsley. A new view for surgeons. Scientific American, 271(5):100, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-100.pdf.

Leutwyler:1994:PL

[2152] Kristin Leutwyler. Productivity lost. Scientific American, 271 (5):101–102, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
101.pdf.

Gibbs:1994:PI

[2153] W. Wayt Gibbs. Pricing Internet. Scientific American, 271(5):102–103, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-102.pdf.

Gibbs:1994:OP

[2154] W. Wayt Gibbs. Ounce of prevention. Scientific American, 271 (5):103–105, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
103.pdf.

Nemecek:1994:BB

[2155] Sasha Nemecek. Bettering batteries. Scientific American, 271(5):106, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-106.pdf.

McCloskey:1994:EUP

[2156] Donald N. McCloskey. An economic uncertainty principle. Scientific American, 271(5):107, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
107.pdf.

Stewart:1994:MRP

[2157] Ian Stewart. Mathematical recreations: Playing chess on a go board. Scientific American, 271(5):108–111, November 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n5/pdf/scientificamerican1194-108.pdf.

Anonymous:1994:MRf

[2158] Anonymous. Mathematical recreations. Scientific American, 271(5): 108–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRj

[2159] Philip Morrison. Book reviews. Scientific American, 271(5):112–115, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n5/pdf/scientificamerican1194-112.pdf.

Anonymous:1994:BRk

[2160] Anonymous. Book reviews. Scientific American, 271(5):112–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1994:EWP

[2161] Anne Eisenberg. Essay: Words at play on the Internet. Scientific American, 271(5):116, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n5/pdf/scientificamerican1194-
116.pdf.

Eosemberg:1994:E

[2162] Anne Eosemberg. Essay. Scientific American, 271(5):116–??, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:Eg

[2163] Anonymous. Errata. Scientific American, 271(6):8, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n6/pdf/scientificamerican1294-8b.pdf.

Anonymous:1994:LEl

[2164] Anonymous. Letters to the Editors. Scientific American, 271(6):8, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-8a.pdf.

Anonymous:1994:YAIl

[2165] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 271(6):10, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
10.pdf.

Anonymous:1994:YAm

[2166] Anonymous. 50 and 100 years ago. Scientific American, 271(6):10–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1994:SCN

[2167] W. Wayt Gibbs. Science and the citizen: No-polluting zone. Scientific American, 271(6):14–16, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
14.pdf.

Anonymous:1994:SCl

[2168] Anonymous. Science and the citizen. Scientific American, 271(6):14–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1994:ANG

[2169] Corey S. Powell. The astronomical naming game. Scientific American, 271(6):16–20, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
16.pdf.

Anonymous:1994:EAA

[2170] Anonymous. Electrical activity above thunderstorms. Scientific American, 271(6):20, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
20b.pdf.

Yam:1994:TL

[2171] Philip Yam. Trapped in the light. Scientific American, 271(6):20–22, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-20a.pdf.

Mirsky:1994:AIN

[2172] Steve Mirsky. The annual Ig Nobel Prizes. Scientific American, 271(6):22–26, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
22.pdf.

Anonymous:1994:TWR

[2173] Anonymous. Toxic waste and race: An unnatural association. Scientific American, 271(6):26, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
26c.pdf.

Stykes:1994:OIN

[2174] Mervin Stykes. And the other 1994 Ig Nobel Prize winners are. Scientific American, 271(6):26, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
26a.pdf.

Leutwyler:1994:DBC

[2175] Kristin Leutwyler. Deciphering the breast cancer gene. Scientific American, 271(6):26–30, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
26b.pdf.

Horgan:1994:GP

[2176] John Horgan. Global politics. Scientific American, 271(6):32–34, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-32.pdf.

Horgan:1994:NNV

[2177] John Horgan. The no-name virus. Scientific American, 271(6):34–36, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-34.pdf.

Wallich:1994:WHD

[2178] Paul Wallich. The wages of Haiti’s dictatorship. Scientific American, 271(6):36, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
36.pdf.

Gibbs:1994:VRC

[2179] W. Wayt Gibbs. Virtual reality check. Scientific American, 271(6):40–42, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-40.pdf.

Wallich:1994:WYE

[2180] Paul Wallich. Watch your electronic mouth. Scientific American, 271(6):43–44, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
43.pdf.

Stix:1994:BAP

[2181] Gary Stix. Bad apple picker. Scientific American, 271(6):44–46, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-44.pdf.

Powell:1994:RR

[2182] Corey S. Powell. Relinquishing relics. Scientific American, 271(6):46–47, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-46.pdf.

Holloway:1994:TWE

[2183] Marguerite Holloway. On the trail of wild elephants. Scientific American, 271(6):48–50, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
48.pdf.

DeCicco:1994:IAE

[2184] John DeCicco and Marc Ross. Improving automotive efficiency. Scientific American, 271(6):52–57, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
52.pdf.

Novacek:1994:FFC

[2185] Michael J. Novacek, Mark Norell, Malcolm C. McKenna, and James Clark. Fossils of the Flaming Cliffs. Scientific American, 271(6):60–69, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-60.pdf.

Evans:1994:SPE

[2186] Diane L. Evans, Ellen R. Stofan, Thomas D. Jones, and Linda M. Godwin. Science in pictures: Earth from sky. Scientific American, 271(6):70–75, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
70.pdf.

Evans:1994:ES

[2187] Diane L. Evans, Ellen R. Stofan, Thomas D. Jones, and Linda M. Godwin. Earth from sky. Scientific American, 271(6):70–?? (Intl. ed. 44–??), December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1994:TSCa

[2188] Gary Stix. Trends in scientific communication: The speed of write. Scientific American, 271(6):72–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cohen:1994:NGM

[2189] Jack S. Cohen and Michael E. Hogan. The new genetic medicines. Scientific American, 271(6):76–82, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
76.pdf.

Ross:1994:MED

[2190] Marc Ross and John Decicco. Measuring the energy drain on your car. Scientific American, 271(6):80–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Englert:1994:DML

[2191] Berthold-Georg Englert, Marlan O. Scully, and Herbert Walther. The duality in matter and light. Scientific American, 271(6):86–92, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-86.pdf.

French:1994:MET

[2192] Hilary F. French. Making environmental treaties work. Scientific American, 271(6):94–97, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
94.pdf.

Jacobs:1994:C

[2193] William P. Jacobs. Caulerpa. Scientific American, 271(6):100–105, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-100.pdf.

Stix:1994:TSCb

[2194] Gary Stix. Trends in scientific communication: The speed of write. Scientific American, 271(6):106–111, December 1994. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v271/
n6/pdf/scientificamerican1294-106.pdf.

Stix:1994:SW

[2195] Gary Stix. The speed of write. Scientific American, 271(6):106–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ross:1994:ASM

[2196] Marc Ross and John DeCicco. The amateur scientist: Measuring the energy drain on your car. Scientific American, 271(6):112–115, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-112.pdf.

Anonymous:1994:ASf

[2197] Anonymous. The amateur scientist. Scientific American, 271(6):112–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1994:BRk

[2198] Philip Morrison and Phylis Morrison. Book reviews. Scientific American, 271(6):116–123, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
116.pdf.

Anonymous:1994:BRl

[2199] Anonymous. Book reviews. Scientific American, 271(6):116–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:AI

[2200] Anonymous. Annual index 1994. Scientific American, 271(6):124–126, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v271/n6/pdf/scientificamerican1294-124.pdf.

Chaisson:1994:ECK

[2201] Eric J. Chaisson. Essay: Clementine, keyhole and cosmic rats. Scientific American, 271(6):127, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v271/n6/pdf/scientificamerican1294-
127.pdf.

Chaisson:1994:E

[2202] Eric J. Chaisson. Essay. Scientific American, 271(6):127–??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:AE

[2203] Anonymous. The analytical economist. Scientific American, 271(6):??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:P

[2204] Anonymous. Profile. Scientific American, 271(6):??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1994:TB

[2205] Anonymous. Technology and business. Scientific American, 271(6):??, December 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEa

[2206] Anonymous. Letters to the Editors. Scientific American, 272(1):8, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-8.pdf.

Anonymous:1995:YAIa

[2207] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 272(1):10, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
10.pdf.

Anonymous:1995:YAa

[2208] Anonymous. 50 and 100 years ago. Scientific American, 272(1):10–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1995:SCW

[2209] Tim Beardsley. Science and the citizen: For whom the bell curve really tolls. Scientific American, 272(1):14–17, January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-14.pdf.

Anonymous:1995:SCa

[2210] Anonymous. Science and the citizen. Scientific American, 272(1):14–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1995:DR

[2211] Corey S. Powell. Deathbed revelations. Scientific American, 272(1):16, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-16.pdf.

Leutwyler:1995:GA

[2212] Kristin Leutwyler. The great attractors. Scientific American, 272(1):17– 20, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-17.pdf.

Holloway:1995:SNN

[2213] Marguerite Holloway. Socializing with non-naked mole rats. Scientific American, 272(1):20, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
20.pdf.

Powell:1995:PP

[2214] Corey S. Powell. Picking up the pieces. Scientific American, 272(1):21, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-21.pdf.

Mukerjee:1995:MMD

[2215] Madhusree Mukerjee. Mystery of the missing dynamo. Scientific American, 272(1):24–26, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
24.pdf.

Nemecek:1995:HOS

[2216] Sasha Nemecek. Holes in ozone science. Scientific American, 272(1):26– 27, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-26.pdf.

Holloway:1995:FSL

[2217] Marguerite Holloway. Fear and self-loathing in America. Scientific American, 272(1):27, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
27a.pdf.

Holloway:1995:SCS

[2218] Marguerite Holloway. Seeing the cells that see. Scientific American, 272(1):27, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
27b.pdf.

Wallich:1995:AED

[2219] Paul Wallich. The analytical economist: Derivatives: Not the real thing. Scientific American, 272(1):28, January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-28.pdf.

Gibbs:1995:TBB

[2220] W. Wayt Gibbs. Technology and business: Bellcore on the block. Scientific American, 272(1):29–30, January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-29.pdf.

Powell:1995:RS

[2221] Corey S. Powell. The rights stuff. Scientific American, 272(1):30–31, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-30.pdf.

Holloway:1995:SS

[2222] Marguerite Holloway. Secrets in stereogram. Scientific American, 272(1): 31, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-31.pdf.

Beardsley:1995:RN

[2223] Tim Beardsley. On the road to nowhere? Scientific American, 272(1):32– 33, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-32.pdf.

Wallich:1995:IBC

[2224] Paul Wallich. Invasion of the bean counters. Scientific American, 272(1):33–36, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
33.pdf.

Holloway:1995:FT

[2225] Marguerite Holloway. Food for thought. Scientific American, 272(1):36, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-36.pdf.

Yam:1995:PWH

[2226] Philip Yam. Profile: Walter H. Munk: The man who would hear ocean temperatures. Scientific American, 272(1):38–40, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-38.pdf.

Rothenberg:1995:ELD

[2227] Jeff Rothenberg. Ensuring the longevity of digital documents. Scientific American, 272(1):42–47 (Intl. ed. 24–29), January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-42.pdf.

Prusiner:1995:PD

[2228] Stanley B. Prusiner. The Prion diseases. Scientific American, 272(1):48– 57, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-48.pdf.

Anonymous:1995:PE

[2229] Anonymous. A persuasive experiment. Scientific American, 272(1):57, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-57.pdf.

Dalziel:1995:EBP

[2230] Ian W. D. Dalziel. Earth before Pangea. Scientific American, 272(1):58– 63, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-58.pdf.

Urry:1995:EBM

[2231] Dan W. Urry. Elastic biomolecular machines. Scientific American, 272 (1):64–69 (Intl. ed. 44–??), January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
64.pdf.

Beardsley:1995:TPMa

[2232] Tim Beardsley. Trends in preventive medicine: Better than a cure. Scientific American, 272(1):68–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Perls:1995:OO

[2233] Thomas T. Perls. The oldest old. Scientific American, 272(1):70–75, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-70.pdf.

Anonymous:1995:SFa

[2234] Anonymous. Survival of the fittest. Scientific American, 272(1):75, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-75a.pdf.

Anonymous:1995:TWT

[2235] Anonymous. Three who thrive. Scientific American, 272(1):75, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-75b.pdf.

Starrfield:1995:BDNa

[2236] Sumner Starrfield and Steven N. Shore. The birth and death of Nova V1974 Cygni. Scientific American, 272(1):76–81, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-76.pdf.

Anonymous:1995:DDG

[2237] Anonymous. Daisy, Daisy, give me your answer, do. Scientific American, 272(1):76–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Starrfield:1995:BDNb

[2238] Sumner Starrfield and Steven N. Shore. The birth and death of Nova V1974 Cygni. Scientific American, 272(1):76–?? (Intl. ed. 56–61), January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Byock:1995:EB

[2239] Jesse L. Byock. Egil’s bones. Scientific American, 272(1):82–87, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-82.pdf.

Beardsley:1995:TPMb

[2240] Tim Beardsley. Trends in preventive medicine: Better than a cure. Scientific American, 272(1):88–95, January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-88.pdf.

Beardsley:1995:BTC

[2241] Tim Beardsley. Better than a cure. Scientific American, 272(1):88–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:BU

[2242] Anonymous. Born in the U.S.A. Scientific American, 272(1):95, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-95b.pdf.

Anonymous:1995:MWM

[2243] Anonymous. Many ways to make a vaccine. Scientific American, 272(1): 95, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n1/pdf/scientificamerican0195-95a.pdf.

Stewart:1995:MRD

[2244] Ian Stewart. Mathematical recreations: Daisy, Daisy, give me your answer, do. Scientific American, 272(1):96–99, January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n1/pdf/scientificamerican0195-96.pdf.

Anonymous:1995:MRa

[2245] Anonymous. Mathematical recreations. Scientific American, 272(1):96– ??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1995:BRS

[2246] Philip Morrison. Book reviews: Smart buildings. Scientific American, 272(1):100–103, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
100.pdf.

Anonymous:1995:BRa

[2247] Anonymous. Book reviews. Scientific American, 272(1):100–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Timpane:1995:EHC

[2248] John Timpane. Essay: How to convince a reluctant scientist. Scientific American, 272(1):104, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n1/pdf/scientificamerican0195-
104.pdf.

Timpane:1995:E

[2249] John Timpane. Essay. Scientific American, 272(1):104–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AEa

[2250] Anonymous. The analytical economist. Scientific American, 272(1):??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Pa

[2251] Anonymous. Profile. Scientific American, 272(1):??, January 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBa

[2252] Anonymous. Technology and business. Scientific American, 272(1):??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Ea

[2253] Anonymous. Errata. Scientific American, 272(2):6, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-6a.pdf.

Anonymous:1995:LEb

[2254] Anonymous. Letters to the Editors. Scientific American, 272(2):6, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n2/pdf/scientificamerican0295-6.pdf.

Anonymous:1995:YAY

[2255] Anonymous. 50 and 100 years ago: 50 and 100 years ago. Scientific American, 272(2):8, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
8.pdf.

Yam:1995:SCD

[2256] Philip Yam. Science and the citizen: Dangerous sex. Scientific American, 272(2):10–12, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
10.pdf.

Beardsley:1995:PAT

[2257] Tim Beardsley. Putting Alzheimer’s to the tests. Scientific American, 272(2):12–13, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
12.pdf.

Schneider:1995:GWS

[2258] David Schneider. Global warming is still a hot topic. Scientific American, 272(2):13–14, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
13.pdf.

Stix:1995:BD

[2259] Gary Stix. Broken dreamtime. Scientific American, 272(2):14–15, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n2/pdf/scientificamerican0295-14.pdf.

Beardsley:1995:CAa

[2260] Tim Beardsley. Commanding attention. Scientific American, 272(2):16, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n2/pdf/scientificamerican0295-16b.pdf.

Leutwyler:1995:FFA

[2261] Kristin Leutwyler. Finessing Fermat, again. Scientific American, 272(2):16–18, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
16a.pdf.

Schneider:1995:SHE

[2262] David Schneider. Seeing how the Earth moved. Scientific American, 272(2):18, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
18b.pdf.

Schneider:1995:GEF

[2263] David Schneider. It’s getting easier to find a date. Scientific American, 272(2):18–20, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
18a.pdf.

Mirsky:1995:NPY

[2264] Steve Mirsky. Nothing personal, you’re just not my type. Scientific American, 272(2):20–21, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
20.pdf.

Nemecek:1995:LF

[2265] Sasha Nemecek. Out of the lab and into the fire. Scientific American, 272(2):21–24, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
21.pdf.

Vames:1995:NRI

[2266] Steven Vames. No, really, it was this big. Scientific American, 272(2):24, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n2/pdf/scientificamerican0295-24.pdf.

McCloskey:1995:AEO

[2267] Donald N. McCloskey. The analytical economist: Once upon a time there was a theory. Scientific American, 272(2):25, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-25.pdf.

Beardsley:1995:TBB

[2268] Tim Beardsley. Technology and business: a budgetary storm is brewing. Scientific American, 272(2):26, February 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-26.pdf.

Browning:1995:AOA

[2269] John Browning. Agents and other animals. Scientific American, 272(2):28–29, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
28.pdf.

Wallich:1995:CWC

[2270] Paul Wallich. The chilling wind of copyright law? Scientific American, 272(2):30, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
30.pdf.

Mukerjee:1995:PYN

[2271] Madhusree Mukerjee. Profile: Yoichiro Nambu: Strings and gluons — the Seer saw them all. Scientific American, 272(2):37–39, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-37.pdf.

Dasgupta:1995:PPL

[2272] Partha S. Dasgupta. Population, poverty and the local environment. Scientific American, 272(2):40–45, February 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-40.pdf.

Putterman:1995:SSL

[2273] Seth J. Putterman. Sonoluminescence: Sound into light. Scientific American, 272(2):46–51 (Intl. ed. 32–37), February 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-46.pdf.

Tjian:1995:MMC

[2274] Robert Tjian. Molecular machines that control genes. Scientific American, 272(2):54–61, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
54.pdf.

Jamison:1995:MDI

[2275] Kay Redfield Jamison. Manic-depressive illness and creativity. Scientific American, 272(2):62–67, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
62.pdf.

Elitzur:1995:MS

[2276] Moshe Elitzur. Masers in the sky. Scientific American, 272(2):68–74, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n2/pdf/scientificamerican0295-68.pdf.

Stix:1995:TSMa

[2277] Gary Stix. Trends in semiconductor manufacturing: Toward “point one”. Scientific American, 272(2):72–??, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hoberman:1995:HST

[2278] John M. Hoberman and Charles E. Yesalis. The history of synthetic testosterone. Scientific American, 272(2):76–81, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-76.pdf.

Hiller:1995:PLB

[2279] Robert A. Hiller and Bradley P. Barber. Producing light from a bubble of air. Scientific American, 272(2):78–??, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Larson:1995:MCS

[2280] Roger L. Larson. The Mid-Cretaceous Superplume Episode. Scientific American, 272(2):82–86 (Intl. ed. 66–70), February 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-82.pdf.

Stix:1995:TPO

[2281] Gary Stix. Toward “Point one”. Scientific American, 272(2):90–95 (Intl. ed. 72–77), February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1995:TSMb

[2282] Gary Stix. Trends in semiconductor manufacturing: Toward “point one”. Scientific American, 272(2):90–95, February 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n2/pdf/scientificamerican0295-90.pdf.

Hiller:1995:ASP

[2283] Robert A. Hiller and Bradley P. Barber. The amateur scientist: Producing light from a bubble of air. Scientific American, 272(2):96–98, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n2/pdf/scientificamerican0295-96.pdf.

Anonymous:1995:ASa

[2284] Anonymous. The amateur scientist. Scientific American, 272(2):96–??, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kamin:1995:BRB

[2285] Leon J. Kamin. Book review: Behind the curve. Scientific American, 272(2):99–103, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
99.pdf.

Anonymous:1995:BRb

[2286] Anonymous. Book review. Scientific American, 272(2):99–??, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1995:EST

[2287] Anne Eisenberg. Essay: Scientists and their CD-ROMs. Scientific American, 272(2):104, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n2/pdf/scientificamerican0295-
104.pdf.

Anonymous:1995:Eb

[2288] Anonymous. Essay. Scientific American, 272(2):104–??, February 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEc

[2289] Anonymous. Letters to the editors. Scientific American, 272(3):7–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEd

[2290] Anonymous. Letters to the Editors. Scientific American, 272(3):10, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-10.pdf.

Anonymous:1995:YAb

[2291] Anonymous. 50 and 100 years ago. Scientific American, 272(3):10–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIb

[2292] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 272(3):12, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
12.pdf.

Beardsley:1995:SCE

[2293] Tim Beardsley. Science and the citizen: Endangered: One Endangered Species Act. Scientific American, 272(3):18–20, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-18.pdf.

Anonymous:1995:SCb

[2294] Anonymous. Science and the citizen. Scientific American, 272(3):18–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nemecek:1995:SAS

[2295] Sasha Nemecek. Science and art on stage. Scientific American, 272(3):20–21, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
20.pdf.

Nemecek:1995:YEM

[2296] Sasha Nemecek. Not yet elemental, my dear Seaborg. Scientific American, 272(3):21, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
21.pdf.

Yam:1995:TT

[2297] Philip Yam. Talking trash. Scientific American, 272(3):24, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-24.pdf.

Beardsley:1995:PPP

[2298] Tim Beardsley. Pass the plutonium, please. Scientific American, 272(3):28–29, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
28.pdf.

Mukerjee:1995:NE

[2299] Madhusree Mukerjee. Nuclear empowerment. Scientific American, 272(3):29, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
29.pdf.

Schneider:1995:SWO

[2300] David Schneider. Swing wide of that one. Scientific American, 272(3): 30, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-30b.pdf.

Powell:1995:CW

[2301] Corey S. Powell. The chaos within. Scientific American, 272(3):30–32, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-30a.pdf.

Vames:1995:SSS

[2302] Steven Vames. Scientists’ sense of snow. Scientific American, 272(3):32, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-32b.pdf.

Stix:1995:BE

[2303] Gary Stix. Ban that embargo. Scientific American, 272(3):32–34, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-32a.pdf.

Mukerjee:1995:LCD

[2304] Madhusree Mukerjee. Liquid crystals on display. Scientific American, 272(3):34, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
34b.pdf.

Gibbs:1995:TTB

[2305] W. Wayt Gibbs. Treatment that tightens the belt. Scientific American, 272(3):34–35, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
34a.pdf.

Wallich:1995:AEM

[2306] Paul Wallich. The analytical economist: a mystery inside a riddle inside an enigma. Scientific American, 272(3):37, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-37.pdf.

Stix:1995:TBR

[2307] Gary Stix. Technology and business: a recombinant feast. Scientific American, 272(3):38–40, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
38.pdf.

Browning:1995:GUB

[2308] John Browning. GIF us a break. Scientific American, 272(3):40, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-40.pdf.

Schneider:1995:MGB

[2309] David Schneider. MRI goes back to the future. Scientific American, 272(3):42, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
42.pdf.

Browning:1995:PF

[2310] John Browning. Phone fight. Scientific American, 272(3):43, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-43.pdf.

Horgan:1995:PFH

[2311] John Horgan. Profile: Fred Hoyle: The return of the maverick. Scientific American, 272(3):46–47, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
46.pdf.

Kessler:1995:FEV

[2312] David A. Kessler and Karyn L. Feiden. Faster evaluation of vital drugs. Scientific American, 272(3):48–54, March 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-48.pdf.

Brumer:1995:LCC

[2313] Paul Brumer and Moshe Shapiro. Laser control of chemical reactions. Scientific American, 272(3):56–63, March 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-56.pdf.

Triantafyllou:1995:ESM

[2314] Michael S. Triantafyllou and George S. Triantafyllou. An efficient swimming machine. Scientific American, 272(3):64–70, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-64.pdf.

Cavenee:1995:GBC

[2315] Webster K. Cavenee and Raymond L. White. The genetic basis of cancer. Scientific American, 272(3):72–79, March 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-72.pdf.

Gibbs:1995:TBSa

[2316] W. Wayt Gibbs. Trends in behavioral science: Seeking the criminal element. Scientific American, 272(3):76–??, March 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

DeWaal:1995:BSS

[2317] Frans B. M. De Waal. Bonobo sex and society. Scientific American, 272(3):82–88, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
82.pdf.

Anonymous:1995:TTA

[2318] Anonymous. Turning the tables around. Scientific American, 272(3):84– ??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Birge:1995:PBCa

[2319] Robert R. Birge. Protein-based computers. Scientific American, 272(3):90–95, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
90.pdf.

Runnels:1995:EDA

[2320] Curtis N. Runnels. Environmental degradation in Ancient Greece. Scientific American, 272(3):96–99, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
96.pdf.

Gibbs:1995:TBSb

[2321] W. Wayt Gibbs. Trends in behavioral science: Seeking the criminal element. Scientific American, 272(3):100–107, March 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n3/pdf/scientificamerican0395-100.pdf.

Gibbs:1995:SCE

[2322] W. Wayt Gibbs. Seeking the criminal element. Scientific American, 272 (3):100–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1995:MRT

[2323] Ian Stewart. Mathematical recreations: Turning the tables around. Scientific American, 272(3):108–110, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
108.pdf.

Anonymous:1995:MRb

[2324] Anonymous. Mathematical recreations. Scientific American, 272(3):108– ??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1995:BBR

[2325] Philip Morrison. Brief: Book reviews. Scientific American, 272(3):111– 114, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n3/pdf/scientificamerican0395-111.pdf.

Morrison:1995:BR

[2326] Philip Morrison. Book reviews. Scientific American, 272(3):111–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shackley:1995:ERR

[2327] Steven Shackley. Essay: Relics, rights and regulations. Scientific American, 272(3):115, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n3/pdf/scientificamerican0395-
115.pdf.

Anonymous:1995:Pb

[2328] Anonymous. Profile. Scientific American, 272(3):115–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shakley:1995:MLH

[2329] Essay Shakley. A misguided law hurts Native Americans and museums. Scientific American, 272(3):115–??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AEb

[2330] Anonymous. The analytical economist. Scientific American, 272(3):??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Ta

[2331] Anonymous. Technology. Scientific American, 272(3):??, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEe

[2332] Anonymous. Letters to the Editors. Scientific American, 272(4):10, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-10.pdf.

Anonymous:1995:LEf

[2333] Anonymous. Letters to the editors. Scientific American, 272(4):10– ??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIc

[2334] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 272(4):12, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
12.pdf.

Anonymous:1995:YAc

[2335] Anonymous. 50 and 100 years ago. Scientific American, 272(4):12– ??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1995:SCB

[2336] Glenn Zorpette. Science and the citizen: Bracing for the next big one. Scientific American, 272(4):14–16, April 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-14.pdf.

Anonymous:1995:SCc

[2337] Anonymous. Science and the citizen. Scientific American, 272(4):14– ??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Derr:1995:ER

[2338] Mark Derr. The end of the road. Scientific American, 272(4):16–20, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-16.pdf.

Nemecek:1995:SSD

[2339] Sasha Nemecek. Sun spotting the difference. Scientific American, 272(4): 20, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-20b.pdf.

Powell:1995:AD

[2340] Corey S. Powell. Astronomers in the dark. Scientific American, 272(4):20–22, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
20a.pdf.

Vames:1995:RVS

[2341] Steven Vames. A ringside view of stars. Scientific American, 272(4):22, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-22.pdf.

Mukerjee:1995:CS

[2342] Madhusree Mukerjee. Coming out in the sciences. Scientific American, 272(4):24, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
24a.pdf.

Nemecek:1995:GNK

[2343] Sasha Nemecek. I get no kick from CH_{3}CH_{2}OH. Scientific American, 272(4):24–26, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
24b.pdf.

Beardsley:1995:FAT

[2344] Tim Beardsley. Fighting all the time. Scientific American, 272(4):26–29, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-26.pdf.

Schneider:1995:GM

[2345] David Schneider. Down and out in the Gulf of Mexico. Scientific American, 272(4):29, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
29.pdf.

Yam:1995:CW

[2346] Philip Yam. Catching that wave. Scientific American, 272(4):30–31, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-30.pdf.

McCloskey:1995:AEI

[2347] Donald N. McCloskey. The analytical economist: The insignificance of statistical significance. Scientific American, 272(4):32–33, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-32.pdf.

Gibbs:1995:TBCa

[2348] W. Wayt Gibbs. Technology and business: Coming in for a landing. Scientific American, 272(4):34–35, April 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-34.pdf.

Wallich:1995:WBF

[2349] Paul Wallich. A widget’s best friend. Scientific American, 272(4):35, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-35.pdf.

Stix:1995:SB

[2350] Gary Stix. Simply, the best. Scientific American, 272(4):36, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-36a.pdf.

Beardsley:1995:GPD

[2351] Tim Beardsley. Genes in the not so public domain. Scientific American, 272(4):36–37, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
36b.pdf.

Stix:1995:VC

[2352] Gary Stix. Voting for a cure. Scientific American, 272(4):37, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-37.pdf.

Mukerjee:1995:FBW

[2353] Madhusree Mukerjee. The fishy business of waste. Scientific American, 272(4):40, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
40.pdf.

Schneider:1995:CI

[2354] David Schneider. Changing the image. Scientific American, 272(4):42, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-42.pdf.

Gibbs:1995:EF

[2355] W. Wayt Gibbs. Engineering the future. Scientific American, 272(4):44– 48, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n4/pdf/scientificamerican0495-44.pdf.

Blaustein:1995:PDA

[2356] Andrew R. Blaustein and David B. Wake. The puzzle of declining amphibian populations. Scientific American, 272(4):52–57, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-52.pdf.

Jokipii:1995:QLH

[2357] J. R. Jokipii and Frank B. McDonald. Quest for the limits of the heliosphere. Scientific American, 272(4):58–63, April 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-58.pdf.

Abu-Mostafa:1995:MLHa

[2358] Yaser S. Abu-Mostafa. Machines that learn from hints. Scientific American, 272(4):64–69, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
64.pdf.

Greenspan:1995:UGC

[2359] Ralph J. Greenspan. Understanding the genetic construction of behavior. Scientific American, 272(4):72–78, April 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-72.pdf.

Leutwyler:1995:THCa

[2360] Kristin Leutwyler. Trends in health care: The price of prevention. Scientific American, 272(4):98–??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chiavinglese:1995:CRJ

[2361] Ricardo Chiavinglese. Computerized restoration of juvenile art. Scientific American, 272(4):104–??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schwartz:1995:SPA

[2362] Lillian Schwartz. Science in pictures: The art historian’s computer. Scientific American, 272(4):106–111, April 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-106.pdf.

Schwartz:1995:AHC

[2363] Lillian Schwartz. The art historian’s computer. Scientific American, 272(4):106–?? (Intl. ed. 80–??), April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Moore:1995:BHI

[2364] A. W. Moore. A brief history of infinity. Scientific American, 272(4): 112–116 (Intl. ed. 86–??), April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
112.pdf.

Stone:1995:TPM

[2365] Elizabeth C. Stone and Paul Zimansky. The tapestry of power in a Mesopotamian city. Scientific American, 272(4):118–123, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-118.pdf.

Leutwyler:1995:THCb

[2366] Kristin Leutwyler. Trends in health care: The price of prevention. Scientific American, 272(4):124–129, April 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-124.pdf.

Leutwyler:1995:PP

[2367] Kristin Leutwyler. The price of prevention. Scientific American, 272 (4):124–??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chiavinglese:1995:ASC

[2368] Ricardo Chiav’inglese. The amateur scientist: Computerized restoration of juvenile art. Scientific American, 272(4):130–132, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n4/pdf/scientificamerican0495-130.pdf.

Anonymous:1995:ASb

[2369] Anonymous. The amateur scientist. Scientific American, 272(4):130– ??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schacter:1995:BRM

[2370] Daniel L. Schacter. Book reviews: Memory wars. Scientific American, 272(4):135–139, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
135.pdf.

Schacter:1995:BR

[2371] Daniel L. Schacter. Book review. Scientific American, 272(4):135–??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bernstein:1995:EPP

[2372] Jeremy Bernstein. Essay: The poor person’s guide to the bell curve. Scientific American, 272(4):140, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n4/pdf/scientificamerican0495-
140.pdf.

Bernstein:1995:E

[2373] Jeremy Bernstein. Essay. Scientific American, 272(4):140–??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AEc

[2374] Anonymous. The analytical economist. Scientific American, 272(4): ??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Pc

[2375] Anonymous. Profile. Scientific American, 272(4):??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBb

[2376] Anonymous. Technology and business. Scientific American, 272(4):??, April 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEg

[2377] Anonymous. Letters to the Editors. Scientific American, 272(5):8, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-8.pdf.

Anonymous:1995:YAId

[2378] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 272(5):10, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
10.pdf.

Gibbs:1995:SCI

[2379] W. Wayt Gibbs. Science and the citizen: Information have-nots. Scientific American, 272(5):12–14, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
12.pdf.

Anonymous:1995:AEd

[2380] Anonymous. The analytical economist. Scientific American, 272(5):12B– ??, May 1, 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCd

[2381] Anonymous. Science and the citizen. Scientific American, 272(5):12B– ??, May 1, 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBc

[2382] Anonymous. Technology and business. Scientific American, 272(5):12B– ??, May 1, 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1995:SOT

[2383] David Schneider. The sound of one tree breathing. Scientific American, 272(5):14, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
14.pdf.

Beardsley:1995:CWD

[2384] Tim Beardsley. The Cold War’s dirty secrets. Scientific American, 272 (5):16, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-16a.pdf.

Mukerjee:1995:TS

[2385] Madhusree Mukerjee. Tribal struggle. Scientific American, 272(5):16– 20, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-16b.pdf.

Vames:1995:SS

[2386] Steven Vames. Sponging off shrimp. Scientific American, 272(5):18, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-18.pdf.

Horgan:1995:TPT

[2387] John Horgan. Top price for the top quark. Scientific American, 272(5): 20–21, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-20a.pdf.

Stix:1995:WW

[2388] Gary Stix. Why worry? Scientific American, 272(5):20–21, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-20b.pdf.

Zorpett:1995:MSL

[2389] Glenn Zorpett. So many salmon, but so little. Scientific American, 272(5):21–22, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
21.pdf.

Beardsley:1995:LD

[2390] Tim Beardsley. Life’s a draw. Scientific American, 272(5):23–24, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-23.pdf.

Gibbs:1995:NTW

[2391] W. Wayt Gibbs. The naughtiest teens in the world. Scientific American, 272(5):24, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
24a.pdf.

Yam:1995:TLD

[2392] Philip Yam. As they lay dying. Scientific American, 272(5):24–25, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-24b.pdf.

Stix:1995:ME

[2393] Gary Stix and Paul Wallich. Miracles for export. Scientific American, 272(5):26–27, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
26.pdf.

Stix:1995:LBP

[2394] Gary Stix. Lithography becomes political pork. Scientific American, 272 (5):30, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-30.pdf.

Wallich:1995:RR

[2395] Paul Wallich. A rogue’s routing. Scientific American, 272(5):31, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-31.pdf.

Schneider:1995:FDL

[2396] David Schneider. Feature from the Dark Lagoon. Scientific American, 272(5):32, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
32a.pdf.

Stix:1995:PMB

[2397] Gary Stix. Putting the mass back in transit. Scientific American, 272(5): 32–33, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-32b.pdf.

Paterson:1995:EG

[2398] David Paterson. Electric genes. Scientific American, 272(5):33–34, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-33.pdf.

Browning:1995:EL

[2399] John Browning. Europeans on-line. Scientific American, 272(5):35, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-35.pdf.

Horgan:1995:JIJ

[2400] John Horgan. Josephson’s inner junction. Scientific American, 272(5): 40–41, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-40.pdf.

Tomalia:1995:DMa

[2401] Donald A. Tomalia. Dendrimer molecules. Scientific American, 272(5): 42–??, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bartecchi:1995:GTE

[2402] Carl E. Bartecchi, Thomas D. MacKenzie, and Robert W. Schrier. The Global Tobacco Epidemic. Scientific American, 272(5):44–51, May 1, 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-44.pdf.

Schmitt:1995:OSFa

[2403] Raymond W. Schmitt, Jr. The ocean’s salt fingers. Scientific American, 272(5):50–??, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Piran:1995:BNS

[2404] Tsvi Piran. Binary neutron stars. Scientific American, 272(5):52–61, May 1, 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-52.pdf.

Tomalia:1995:DMb

[2405] Donald A. Tomalia. Dendrimer molecules. Scientific American, 272(5): 62–66, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-62.pdf.

Schmitt:1995:OSFb

[2406] Raymond W. Schmitt, Jr. The ocean’s salt fingers. Scientific American, 272(5):70–75, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
70.pdf.

Litke:1995:SMD

[2407] Alan M. Litke and Andreas S. Schwarz. The silicon microstrip detector. Scientific American, 272(5):76–81, May 1, 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-76.pdf.

Holloway:1995:TAPa

[2408] Marguerite Holloway. Trends in archaeology: The preservation of past. Scientific American, 272(5):78–??, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:FF

[2409] Anonymous. Fibonacci forgeries. Scientific American, 272(5):82–??, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SIA

[2410] Anonymous. Special investigation: The atomic intrigues of Niels Bohr. Scientific American, 272(5):83, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
83.pdf.

Bethe:1995:DBS

[2411] Hans A. Bethe, Kurt Gottfried, and Roald Z. Sagdeev. Did Bohr share nuclear secrets? Scientific American, 272(5):84–90, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-84.pdf.

Bernstein:1995:WDH

[2412] Jeremy Bernstein. What did Heisenberg tell Bohr about the bomb? Scientific American, 272(5):92–97, May 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-92.pdf.

Holloway:1995:TAPb

[2413] Marguerite Holloway. Trends in archaeology: The preservation of past. Scientific American, 272(5):98–101, May 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-98.pdf.

Holloway:1995:PP

[2414] Marguerite Holloway. The preservation of the past. Scientific American, 272(5):98–??, May 1, 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1995:MRF

[2415] Ian Stewart. Mathematical recreations: Fibonacci forgeries. Scientific American, 272(5):102–105, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
102.pdf.

Anonymous:1995:Ra

[2416] Anonymous. Reviews. Scientific American, 272(5):106–111, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n5/pdf/scientificamerican0595-106.pdf.

Mitchell:1995:EPP

[2417] William J. Mitchell. Essay: The parable of the pizza parlor. Scientific American, 272(5):112, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
112.pdf.

Anonymous:1995:LEh

[2418] Anonymous. Letters to the Editors. Scientific American, 272(6):12, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-12.pdf.

Anonymous:1995:YAIe

[2419] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 272(6):14, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
14.pdf.

Anonymous:1995:YAd

[2420] Anonymous. 50 and 100 years ago. Scientific American, 272(6):14– ??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mukerjee:1995:SCP

[2421] Madhusree Mukerjee. Science and the citizen: Persistently toxic. Scientific American, 272(6):16–18, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
16.pdf.

Anonymous:1995:SCe

[2422] Anonymous. Science and the citizen. Scientific American, 272(6):16– ??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1995:AAa

[2423] W. Wayt Gibbs. Ambivalent anniversaries. Scientific American, 272(6): 18, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-18.pdf.

Powell:1995:CWC

[2424] Corey S. Powell. Crisis? what crisis? Scientific American, 272(6):19–22, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-19.pdf.

Stock:1995:FFT

[2425] Christina Stock. Flying in the face of tradition. Scientific American, 272(6):22, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
22.pdf.

Leutwyler:1995:DDS

[2426] Kristin Leutwyler. Depression’s double standard. Scientific American, 272(6):23–26, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
23.pdf.

Powell:1995:DH

[2427] Corey S. Powell. Dinosaurs in the halls. Scientific American, 272(6):26, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-26.pdf.

Beardsley:1995:DA

[2428] Tim Beardsley. Death by analysis. Scientific American, 272(6):27–28, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-27.pdf.

Zorpette:1995:IBS

[2429] Glenn Zorpette. The importance of being sneaky. Scientific American, 272(6):28–30, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
28.pdf.

Horgan:1995:BLT

[2430] John Horgan. Better late than never. Scientific American, 272(6):32, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-32a.pdf.

Horgan:1995:EI

[2431] John Horgan. The eyes have it. Scientific American, 272(6):32, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-32b.pdf.

Beardsley:1995:RR

[2432] Tim Beardsley. Rio redux. Scientific American, 272(6):36, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-36a.pdf.

Wallich:1995:YPT

[2433] Paul Wallich. Yesterday the peso, tomorrow the dollar? Scientific American, 272(6):36–38, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
36b.pdf.

Stix:1995:SIM

[2434] Gary Stix. Suck it to me. Scientific American, 272(6):38–39, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-38.pdf.

Gibbs:1995:TBCb

[2435] W. Wayt Gibbs. Turning back the clock. Scientific American, 272(6):40, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-40.pdf.

Browning:1995:TUA

[2436] John Browning. Tell us another story, please, bill. Scientific American, 272(6):44, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
44.pdf.

Gibbs:1995:CF

[2437] W. Wayt Gibbs. Computing with fire. Scientific American, 272(6):46, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-46b.pdf.

Nemecek:1995:H

[2438] Sasha Nemecek. Have a heart. Scientific American, 272(6):46, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-46a.pdf.

Nemecek:1995:EBN

[2439] Sasha Nemecek. Employment blues: Nothing to do with being green. Scientific American, 272(6):48, June 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-48b.pdf.

Paterson:1995:SS

[2440] David Paterson. Setting a standard. Scientific American, 272(6):48, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v272/n6/pdf/scientificamerican0695-48a.pdf.

Holloway:1995:TI

[2441] Marguerite Holloway. Turning the inside out. Scientific American, 272(6):49–50, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
49.pdf.

Pearce:1995:DE

[2442] David Pearce, Neil Adger, David Maddison, and Dominic Moran. Debt and the environment. Scientific American, 272(6):52–56, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-52.pdf.

Boebinger:1995:BWR

[2443] Greg Boebinger, Al Passner, and Joze Bevk. Building worldrecord magnets. Scientific American, 272(6):58–66, June 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-58.pdf.

Hotez:1995:HI

[2444] Peter J. Hotez and David I. Pritchard. Hookworm infection. Scientific American, 272(6):68–74, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
68.pdf.

Horgan:1995:TCSa

[2445] John Horgan. Trends in complexity studies: From complexity to perplexity. Scientific American, 272(6):74–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nowak:1995:AMH

[2446] Martin A. Nowak, Robert M. May, and Karl Sigmund. The arithmetics of mutual help. Scientific American, 272(6):76–81, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-76.pdf.

Lloyd:1995:CBP

[2447] Alan L. Lloyd. Computing bouts of the Prisoner’s Dilemma. Scientific American, 272(6):80–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Haselberger:1995:SPD

[2448] Lothar Haselberger. Science in pictures: Deciphering a Roman blueprint. Scientific American, 272(6):84–89, June 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-84.pdf.

Haselberger:1995:DRB

[2449] Lothar Haselberger. Deciphering a Roman blueprint. Scientific American, 272(6):84–?? (Intl. ed. 56–??), June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Austin:1995:HN

[2450] Sam M. Austin and George F. Bertsch. Halo nuclei. Scientific American, 272(6):90–95, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
90.pdf.

Pfennig:1995:KR

[2451] David W. Pfennig and Paul W. Sherman. Kin recognition. Scientific American, 272(6):98–103, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n6/pdf/scientificamerican0695-
98.pdf.

Horgan:1995:TCSb

[2452] John Horgan. Trends in complexity studies: From complexity to perplexity. Scientific American, 272(6):104–109, June 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-104.pdf.

Horgan:1995:CP

[2453] John Horgan. From complexity to perplexity. Scientific American, 272 (6):104–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lloyd:1995:ASC

[2454] Alun L. Lloyd. The amateur scientist: Computing bouts of the Prisoner’s Dilemma. Scientific American, 272(6):110–115, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-110.pdf.

Anonymous:1995:ASc

[2455] Anonymous. Amateur scientist. Scientific American, 272(6):110–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Rb

[2456] Anonymous. Reviews. Scientific American, 272(6):116–119, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-116.pdf.

Anonymous:1995:Rc

[2457] Anonymous. Reviews. Scientific American, 272(6):116–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gomory:1995:EKU

[2458] Ralph E. Gomory. Essay: The known, the unknown and the unknowable. Scientific American, 272(6):120, June 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v272/
n6/pdf/scientificamerican0695-120.pdf.

Comory:1995:E

[2459] R. E. Comory. Essay. Scientific American, 272(6):120–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gomory:1995:E

[2460] Ralph E. Gomory. Essay. Scientific American, 272(6):120–??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AEe

[2461] Anonymous. The analytical economist. Scientific American, 272(6): ??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Pd

[2462] Anonymous. Profile. Scientific American, 272(6):??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBd

[2463] Anonymous. Technology and business. Scientific American, 272(6):??, June 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEi

[2464] Anonymous. Letters to the Editors. Scientific American, 273(1):8, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-8.pdf.

Anonymous:1995:YAe

[2465] Anonymous. 50 and 100 years ago. Scientific American, 273(1):10, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-10.pdf.

Anonymous:1995:YAf

[2466] Anonymous. 50 and 100 years ago. Scientific American, 273(1):10– ??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1995:SCD

[2467] Tim Beardsley. Science and the citizen: Darwin denied. Scientific American, 273(1):12–14, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
12.pdf.

Anonymous:1995:SCf

[2468] Anonymous. Science and the citizen. Scientific American, 273(1):12– ??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1995:L

[2469] David Schneider. On the level. Scientific American, 273(1):14, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n1/pdf/scientificamerican0795-14.pdf.

Horgan:1995:BBH

[2470] John Horgan. Bashing black holes. Scientific American, 273(1):16– 17, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-16.pdf.

Nemecek:1995:AG

[2471] Sasha Nemecek. Artistic genes. Scientific American, 273(1):17, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-17a.pdf.

Mirsky:1995:FB

[2472] Steve Mirsky. Fright of the bumblebee. Scientific American, 273(1):17– 18, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-17b.pdf.

Morga:1995:WI

[2473] John Morga. The Waterfall Illusion. Scientific American, 273(1):18, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-18a.pdf.

Yam:1995:AC

[2474] Philip Yam. In the atomic corral. Scientific American, 273(1):18, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-18b.pdf.

Mukerjee:1995:TA

[2475] Madhusree Mukerjee. Toxins abounding. Scientific American, 273(1): 22, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-22a.pdf.

Zorpette:1995:JSI

[2476] Glenn Zorpette. Just scratch it. Scientific American, 273(1):22–23, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-22b.pdf.

Leutwyler:1995:IJ

[2477] Kristin Leutwyler. An inside job. Scientific American, 273(1):24, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-24a.pdf.

Schneider:1995:RMS

[2478] David Schneider. A river (of mud) still runs through it. Scientific American, 273(1):24, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
24b.pdf.

McCloskey:1995:AEC

[2479] Donald N. McCloskey. The analytical economist: Computation outstrips analysis. Scientific American, 273(1):26, July 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n1/pdf/scientificamerican0795-26.pdf.

Beardsley:1995:MM

[2480] Tim Beardsley. It’s melting, it’s melting. Scientific American, 273(1): 28, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-28b.pdf.

Beardsley:1995:TBE

[2481] Tim Beardsley. Technology and business: Environmental secrets. Scientific American, 273(1):28, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
28a.pdf.

Zorpette:1995:FI

[2482] Glenn Zorpette. Food indigo. Scientific American, 273(1):29, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n1/pdf/scientificamerican0795-29a.pdf.

Nemecek:1995:WSG

[2483] Sasha Nemecek. When smog gets in your eyes. Scientific American, 273(1):29–30, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
29b.pdf.

Garfinkel:1995:PP

[2484] Simson Garfinkel. A prime patent. Scientific American, 273(1):30, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-30.pdf.

Browning:1995:FM

[2485] John Browning. Follow the money. Scientific American, 273(1):31– 32, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-31.pdf.

Beardsley:1995:CS

[2486] Tim Beardsley. Ceramic superwire. Scientific American, 273(1):32– 33, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-32.pdf.

Yam:1995:PJR

[2487] Philip Yam. Profile: James Randi: a skeptically inquiring mind. Scientific American, 273(1):34–35, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
34.pdf.

Wayne:1995:PRW

[2488] Robert K. Wayne and John L. Gittleman. The problematic red wolf. Scientific American, 273(1):36–39, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
36.pdf.

Schafrik:1995:PG

[2489] Robert E. Schafrik and Sara E. Church. Protecting the greenback. Scientific American, 273(1):40–46, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
40.pdf.

Lacy:1995:TDT

[2490] Paul E. Lacy. Treating diabetes with transplanted cells. Scientific American, 273(1):50–58, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
50.pdf.

Robison:1995:LOM

[2491] Bruce H. Robison. Light in the ocean’s midwaters. Scientific American, 273(1):60–60A, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
60.pdf.

Chevedden:1995:T

[2492] Paul E. Chevedden, Les Eigenbrod, Vernard Foley, Werner Soedel, and Vernard Foley. The trebuchet. Scientific American, 273(1):66– 71, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-66.pdf.

Rigden:1995:ROB

[2493] John S. J. Rigden. Robert Oppenheimer: Before the war. Scientific American, 273(1):68–??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kammen:1995:CDW

[2494] Daniel M. Kammen. Cookstoves for the developing world. Scientific American, 273(1):72–75 (Intl. ed. 64–??), July 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n1/pdf/scientificamerican0795-72.pdf.

Yarn:1995:TMS

[2495] Philip Yarn. Trends in materials science: Plastics get wired. Scientific American, 273(1):74–??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rigden:1995:JRO

[2496] John S. Rigden. J. Robert Oppenheimer: Before the War. Scientific American, 273(1):76–81, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
76.pdf.

Anonymous:1995:EFB

[2497] Anonymous. Election fever in Blockvotia. Scientific American, 273(1): 80–??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1995:PGW

[2498] Philip Yam. Plastics get wired. Scientific American, 273(1):82–87 (Intl. ed. 74–79), July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1995:TMS

[2499] Philip Yam. Trends in materials science: Plastics get wired. Scientific American, 273(1):82–87, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
82.pdf.

Stewart:1995:MRE

[2500] Ian Stewart. Mathematical recreations: Election fever in Blockvotia. Scientific American, 273(1):88–89, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
88.pdf.

Anonymous:1995:MRc

[2501] Anonymous. Mathematical recreations. Scientific American, 273(1):88– ??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:RBB

[2502] Anonymous. Reviews: Beauty in the beast. Scientific American, 273(1): 90–94, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-90.pdf.

Anonymous:1995:Rd

[2503] Anonymous. Reviews. Scientific American, 273(1):90–??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1995:ERR

[2504] Anne Eisenberg. Essay: Rest and relaxation on the net. Scientific American, 273(1):95, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n1/pdf/scientificamerican0795-
95.pdf.

Anonymous:1995:Ec

[2505] Anonymous. Essay. Scientific American, 273(1):95–??, July 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eisenberg:1995:E

[2506] Anne Eisenberg. Essay. Scientific American, 273(1):95–??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AEf

[2507] Anonymous. The analytical economist. Scientific American, 273(1): ??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Pf

[2508] Anonymous. Profile. Scientific American, 273(1):??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBe

[2509] Anonymous. Technology and business. Scientific American, 273(1):??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Ed

[2510] Anonymous. Erratum. Scientific American, 273(2):8, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-8b.pdf.

Anonymous:1995:LEj

[2511] Anonymous. Letters to the Editors. Scientific American, 273(2):8, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-8a.pdf.

Anonymous:1995:DI

[2512] Anonymous. Defining the infinite. Scientific American, 273(2):8–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEL

[2513] Anonymous. Letters to the editor: Losing land. Scientific American, 273(2):8–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCg

[2514] Anonymous. Science on crime. Scientific American, 273(2):8–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIf

[2515] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50 and 100 years ago. Scientific American, 273(2):10, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n2/pdf/scientificamerican0895-
10.pdf.

Anonymous:1995:CPF

[2516] Anonymous. 1945: Car phones foretold. Scientific American, 273(2):10– ??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:MS

[2517] Anonymous. 1895: The molasses sidewalk. Scientific American, 273 (2):10–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAg

[2518] Anonymous. 50 and 100 years ago. Scientific American, 273(2):10–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Tashman:1995:SCM

[2519] Billy Tashman. Science and the citizen: Misreading dyslexia. Scientific American, 273(2):14–16, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n2/pdf/scientificamerican0895-
14.pdf.

Anonymous:1995:AEV

[2520] Anonymous. The analytical economist: The velocity of money. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:BAT

[2521] Anonymous. For the brain, it’s all in the timing. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:BS

[2522] Anonymous. Biosphere 2, science 0. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:FNL

[2523] Anonymous. Fallout of the nuclear legacy. Scientific American, 273(2): 14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:HCF

[2524] Anonymous. How to catch a fly ball. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:MGH

[2525] Anonymous. Miming gas hydrates in Japan. Scientific American, 273 (2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCDa

[2526] Anonymous. Science and the citizen: Dyslexia misdiagnosed? Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCh

[2527] Anonymous. SIDS controversy. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SSM

[2528] Anonymous. A spoonful of sugar makes the medicine. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBO

[2529] Anonymous. Technology and business: Optical disks aim to erase videotape — permanently. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:UIC

[2530] Anonymous. The ultimate ice cube. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gould:1995:P

[2531] Stephen Jay Gould. Profile. Scientific American, 273(2):14–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1995:AT

[2532] John Horgan. It’s all in the timing. Scientific American, 273(2):16, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-16a.pdf.

Schneider:1995:BP

[2533] David Schneider. Blast from the past. Scientific American, 273(2):16, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-16b.pdf.

Yam:1995:CC

[2534] Philip Yam. Coming in from the cold. Scientific American, 273(2):18–19, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-18.pdf.

Mukerjee:1995:Q

[2535] Madhusree Mukerjee. Quasimodal. Scientific American, 273(2):20–21, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-20b.pdf.

Powell:1995:WK

[2536] Corey S. Powell. What the Keck? Scientific American, 273(2):20–21, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-20a.pdf.

Stock:1995:HT

[2537] Christina Stock. High tidings. Scientific American, 273(2):21–22, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-21.pdf.

Horga:1995:MS

[2538] John Horga. The mystery of SIDS. Scientific American, 273(2):22–24, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-22.pdf.

Beardsley:1995:E

[2539] Tim Beardsley. Down to Earth. Scientific American, 273(2):24–26, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-24.pdf.

Horgan:1995:HCF

[2540] John Horgan. How to catch a fly ball. Scientific American, 273(2):26, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-26a.pdf.

Nemecek:1995:PB

[2541] Sasha Nemecek. Putting bombs away. Scientific American, 273(2):26–28, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-26b.pdf.

Beardsley:1995:TT

[2542] Tim Beardsley. Testing’s toll. Scientific American, 273(2):28, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-28.pdf.

Wallich:1995:FC

[2543] Paul Wallich. Fast cash. Scientific American, 273(2):30, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-30.pdf.

Gibbs:1995:BB

[2544] W. Wayt Gibbs. Beyond binary. Scientific American, 273(2):31–34, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-31.pdf.

Beardsley:1995:SSa

[2545] Tim Beardsley. Sweet success. Scientific American, 273(2):34, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-34.pdf.

Browning:1995:PH

[2546] John Browning. I.T., phone home. Scientific American, 273(2):35–36, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-35.pdf.

DeKoker:1995:G

[2547] Brenda DeKoker. Going down. Scientific American, 273(2):36, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-36a.pdf.

Powell:1995:MD

[2548] Corey S. Powell. Metal detectors. Scientific American, 273(2):36–37, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-36b.pdf.

Morgan:1995:ECI

[2549] John Morgan. Escaping in a cloud of ink. Scientific American, 273(2):37– 41, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-37.pdf.

Morrison:1995:RNW

[2550] Philip Morrison. Recollections of a nuclear war. Scientific American, 273(2):42–46, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n2/pdf/scientificamerican0895-
42.pdf.

Davies-Jones:1995:T

[2551] Robert Davies-Jones. Tornadoes. Scientific American, 273(2):48–57, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-48.pdf.

Nowak:1995:HHD

[2552] Martin A. Nowak and Andrew J. McMichael. How HIV defeats the immune system. Scientific American, 273(2):58–65, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-58.pdf.

Moss:1995:BBN

[2553] Frank Moss and Kurt Wiesenfeld. The benefits of background noise. Scientific American, 273(2):66–69, August 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-66.pdf.

Moon:1995:PDI

[2554] Richard E. Moon, Richard D. Vann, and Peter B. Bennett. The physiology of decompression illness. Scientific American, 273(2):70–77, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-70.pdf.

Gibbs:1995:TSCa

[2555] W. Wayt Gibbs. Trends in scientific communication: Lost science in the third world. Scientific American, 273(2):76–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Narins:1995:FC

[2556] Peter M. Narins. Frog communication. Scientific American, 273(2):78– 83, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n2/pdf/scientificamerican0895-78.pdf.

Levy:1995:CSL

[2557] David H. Levy, Eugene M. Shoemaker, and Carolyn S. Shoemaker. Comet Shoemaker–Levy 9 meets Jupiter. Scientific American, 273(2): 84–91 (Intl. ed. 68–75), August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n2/pdf/scientificamerican0895-
84.pdf.

Garver:1995:DSN

[2558] Wayne Garver and Frank Moss. Detecting signals with noise. Scientific American, 273(2):84–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1995:TSCb

[2559] W. Wayt Gibbs. Trends in scientific communication: Lost science in the Third World. Scientific American, 273(2):92–99, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-92.pdf.

Gibbs:1995:LST

[2560] W. Wayt Gibbs. Lost science in the third world. Scientific American, 273(2):92–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Garver:1995:ASD

[2561] Wayne Garver and Frank Moss. The amateur scientist: Detecting signals with noise. Scientific American, 273(2):100–103, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-100.pdf.

Anonymous:1995:ASU

[2562] Anonymous. The amateur scientist: Using controlled static to pump up the volume. Scientific American, 273(2):100–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McGavin:1995:RSL

[2563] George C. McGavin and Vern L. Bullough. Reviews: Six legs good. Scientific American, 273(2):104–107, August 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n2/pdf/scientificamerican0895-104.pdf.

Anonymous:1995:RLB

[2564] Anonymous. Reviews: Love those bugs. Scientific American, 273(2):104– ??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCi

[2565] Anonymous. Sex counts. Scientific American, 273(2):104–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SSV

[2566] Anonymous. Seven samurai versus the cosmos. Scientific American, 273 (2):104–??, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zuger:1995:EHC

[2567] Abigail Zuger. Essay: The high cost of living. Scientific American, 273(2):108, August 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n2/pdf/scientificamerican0895-
108.pdf.

Anonymous:1995:LEk

[2568] Anonymous. Letters to the Editors. Scientific American, 273(3):10, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-10.pdf.

Anonymous:1995:BNE

[2569] Anonymous. Bohr and nuclear espionage. Scientific American, 273(3): 10–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEm

[2570] Anonymous. Letters to the editors. Scientific American, 273(3):10–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:URV

[2571] Anonymous. UV radiation and vanishing frogs. Scientific American, 273 (3):10–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:VML

[2572] Anonymous. Da Vinci and Mona Lisa. Scientific American, 273(3):10–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIg

[2573] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 273(3):12–17, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-12.pdf.

Anonymous:1995:YAh

[2574] Anonymous. 50, 100 and 150 years ago. Scientific American, 273(3): 12–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1995:SCT

[2575] W. Wayt Gibbs. Science and the citizen: The treaty that worked — almost. Scientific American, 273(3):18–20, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-18.pdf.

Anonymous:1995:AEg

[2576] Anonymous. The analytical economist. Scientific American, 273(3): 18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:BAS

[2577] Anonymous. Become an arms smuggler: Carry software overseas. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:BLD

[2578] Anonymous. Boston’s lobsters dredge up trouble. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:BMO

[2579] Anonymous. Black-market ozone wreckers. Scientific American, 273(3): 18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:DCP

[2580] Anonymous. DNA computing: Plugging pieces into proteins. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:FRF

[2581] Anonymous. Flying robots fight for glory. Scientific American, 273(3): 18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:NFF

[2582] Anonymous. No funds to find Ebola. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:OSF

[2583] Anonymous. Oil spills in the former Soviet Union. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Pg

[2584] Anonymous. Profile. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:PFI

[2585] Anonymous. Psychologists find inspiration in chaos. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCj

[2586] Anonymous. Science and the citizen. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBf

[2587] Anonymous. Technology and business. Scientific American, 273(3):18– ??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:WPP

[2588] Anonymous. Worries over a patent plague gene therapy. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Draper:1995:C

[2589] Tom Draper. Cover. Scientific American, 273(3):18–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1995:MDA

[2590] W. Wayt Gibbs. The most dangerous animal. Scientific American, 273(3):20, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
20b.pdf.

Leutwyler:1995:CD

[2591] Kristin Leutwyler. Calculating with DNA. Scientific American, 273(3):20–22, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
20a.pdf.

Stock:1995:ER

[2592] Christina Stock. Ectoplasm reigns. Scientific American, 273(3):22–24, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-22.pdf.

Stix:1995:BCS

[2593] Gary Stix. Boot camp for surgeons. Scientific American, 273(3):24, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-24b.pdf.

Nemecek:1995:NLA

[2594] Sasha Nemecek. New letters for alphabet soup. Scientific American, 273(3):24–28, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
24a.pdf.

Horgan:1995:CF

[2595] John Horgan. Complexifying freud. Scientific American, 273(3):28–30, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-28.pdf.

Nemecek:1995:OGP

[2596] Sasha Nemecek. One good pest deserves another. Scientific American, 273(3):30–31, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
30.pdf.

Zorpette:1995:LS

[2597] Glenn Zorpette. Lobster stew. Scientific American, 273(3):31–33, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-31.pdf.

Leutwyler:1995:HS

[2598] Kristin Leutwyler. Hide-and-seek. Scientific American, 273(3):33–36, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-33.pdf.

Derr:1995:CS

[2599] Mark Derr. Common scents. Scientific American, 273(3):36–37, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-36.pdf.

Mirsky:1995:SG

[2600] Steve Mirsky. Silence of the genes. Scientific American, 273(3):40–42, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-40.pdf.

Schneider:1995:LS

[2601] David Schneider. Like a sieve. Scientific American, 273(3):42, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-42.pdf.

McCloskey:1995:AEG

[2602] Donald N. McCloskey. The analytical economist: The Gulliver Effect. Scientific American, 273(3):44, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-44.pdf.

Beardsley:1995:TBP

[2603] Tim Beardsley. Technology and business: Patently obvious. Scientific American, 273(3):45, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
45.pdf.

Schneider:1995:MMM

[2604] David Schneider. Magnificent men (mostly) and their flying machines. Scientific American, 273(3):46, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-46b.pdf.

Zorpette:1995:AP

[2605] Glenn Zorpette. Arrest that passenger. Scientific American, 273 (3):46–48, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
46a.pdf.

Leutwyler:1995:PNN

[2606] Kristin Leutwyler. Profile: Nicholas Negroponte: The guru of cyberspace. Scientific American, 273(3):50–52, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-50.pdf.

Rennie:1995:IUT

[2607] John Rennie. Introduction: The uncertainties of technological innovation. Scientific American, 273(3):56–58, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-56.pdf.

Chan:1995:AONa

[2608] Vincent W. S. Chan. All-optical networks. Scientific American, 273(3): 56–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1995:UTI

[2609] John Rennie. The uncertainties of technological innovation. Scientific American, 273(3):56–?? (Intl. ed. 42–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:IT

[2610] Anonymous. Information technologies. Scientific American, 273 (3):60–94, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
60.pdf.

Patterson:1995:M

[2611] David A. Patterson. Microprocessors in 2020. Scientific American, 273 (3):62–?? (Intl. ed. 48–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zysman:1995:WN

[2612] George I. Zysman. Wireless networks. Scientific American, 273(3):68–?? (Intl. ed. 52–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Laurel:1995:VR

[2613] Brenda Laurel. Virtual reality. Scientific American, 273(3):70–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Daggatt:1995:SDW

[2614] Russell Daggatt. Satellites for a developing world. Scientific American, 273(3):71–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chan:1995:AONb

[2615] Vincent W. S. Chan. All-optical networks. Scientific American, 273(3): 72–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eastham:1995:HSRa

[2616] Tony R. Eastham. High-speed rail: Another golden age? Scientific American, 273(3):74–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lenat:1995:AI

[2617] Douglas B. Lenat. Artificial intelligence. Scientific American, 273(3): 80–82 (Intl. ed. 62–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Maes:1995:IS

[2618] Pattie Maes. Intelligent software. Scientific American, 273(3):84–86 (Intl. ed. 66–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Laurel:1995:CVR

[2619] Brenda Laurel. Commentary: Virtual reality. Scientific American, 273 (3):90–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cervero:1995:WGA

[2620] Robert Cervero. Why go anywhere? Scientific American, 273(3):92–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Daggatt:1995:CSD

[2621] Russell Daggatt. Commentary: Satellites for a developing world. Scientific American, 273(3):94–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Tc

[2622] Anonymous. Transportation. Scientific American, 273(3):98–120, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-98.pdf.

Dyson:1995:TCS

[2623] Freeman J. Dyson. Transportation: 21st-Century spacecraft. Scientific American, 273(3):114–116, 116A, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-144.pdf.

Anonymous:1995:Td

[2624] Anonymous. Transportation. Scientific American, 273(3):98–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Eastham:1995:HSRb

[2625] Tony R. Eastham. High-speed rail. another golden age? Scientific American, 273(3):100–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zetsche:1995:ACC

[2626] Dieter Zetsche. The automobile: Clean and customized. Scientific American, 273(3):102–106 (Intl. ed. 76–??), September 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Covert:1995:ECA

[2627] Eugene E. Covert. Evolution of the commercial airliner. Scientific American, 273(3):110–113 (Intl. ed. 82–??), September 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Caplan:1995:IF

[2628] Arthur Caplan. An improved future? Scientific American, 273(3):110– ??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dyson:1995:CSb

[2629] Freeman J. Dyson. 21st-Century spacecraft. Scientific American, 273(3): 114–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Whitesides:1995:SAMa

[2630] George M. Whitesides. Self-assembling materials. Scientific American, 273(3):114–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cervero:1995:CWG

[2631] Robert Cervero. Commentary: Why go anywhere? Scientific American, 273(3):118–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:M

[2632] Anonymous. Medicine. Scientific American, 273(3):122–143, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-122.pdf.

Anderson:1995:GT

[2633] W. French Anderson. Gene therapy. Scientific American, 273(3):124–?? (Intl. ed. 96–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chu:1995:HTSa

[2634] Paul C. W. Chu. High-temperature superconductors. Scientific American, 273(3):128–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Langer:1995:AO

[2635] Robert Langer and Joseph P. Vacanti. Artificial organs. Scientific American, 273(3):130–133 (Intl. ed. 100–??), September 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Engelberger:1995:RC

[2636] Joseph F. Engelberger. Robotics in the 21st Century. Scientific American, 273(3):132–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Alexander:1995:FC

[2637] Nancy J. Alexander. Future contraceptives. Scientific American, 273(3): 136–?? (Intl. ed. 104–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Caplan:1995:CIF

[2638] Arthur Caplan. Commentary: An improved future? Scientific American, 273(3):142–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:MMM

[2639] Anonymous. Machines, materials and manufacturing. Scientific American, 273(3):144–166, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
144.pdf.

Whitesides:1995:SAMb

[2640] George M. Whitesides. Self-assembling materials. Scientific American, 273(3):146–149, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gabriel:1995:EMM

[2641] Kaigham J. Gabriel. Engineering microscopic machines. Scientific American, 273(3):150–153 (Intl. ed. 118–??), September 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

VonLersner:1995:OEE

[2642] Heinrich Von Lersner. Outline for an ecological economy. Scientific American, 273(3):153–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rogers:1995:IM

[2643] Craig A. Rogers. Intelligent materials. Scientific American, 273(3):154– 157 (Intl. ed. 122–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chu:1995:HTSb

[2644] Paul C. W. Chu. High-temperature superconductors. Scientific American, 273(3):162–165, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:GDR

[2645] Anonymous. The Great Drain Robbery. Scientific American, 273(3): 166–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Engelberger:1995:CRC

[2646] Joseph F. Engelberger. Commentary: Robotics in the 21st century. Scientific American, 273(3):166–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:EE

[2647] Anonymous. Energy and environment. Scientific American, 273 (3):168–188, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
168.pdf.

Hoagland:1995:SE

[2648] William Hoagland. Solar energy. Scientific American, 273(3):170–173 (Intl. ed. 136–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Furth:1995:F

[2649] Harold P. Furth. Fusion. Scientific American, 273(3):174–176 (Intl. ed. 140–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Frosch:1995:IEC

[2650] Robert A. Frosch. The industrial ecology of the 21st Century. Scientific American, 273(3):178–?? (Intl. ed. 144–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Plucknett:1995:TSA

[2651] Donald L. Plucknett and Donald L. Winkelmann. Technology for a sustainable agriculture. Scientific American, 273(3):182–?? (Intl. ed. 148– ??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

vonLersner:1995:COE

[2652] Heinrich von Lersner. Commentary: Outline for an ecological economy. Scientific American, 273(3):188–??, September 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LNTa

[2653] Anonymous. Living with new technologies. Scientific American, 273(3):190–205, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
190.pdf.

Anonymous:1995:CLN

[2654] Anonymous. Commentaries: Living with new technologies. Scientific American, 273(3):190–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LNTb

[2655] Anonymous. Living with new technologies. Scientific American, 273(3): 190–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Prabhakar:1995:TI

[2656] Arati Prabhakar. Technology infrastructure. Scientific American, 273(3): 192–?? (Intl. ed. 158–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Norman:1995:DF

[2657] Donald A. Norman. Designing the future. Scientific American, 273(3): 194–?? (Intl. ed. 159–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lanham:1995:DL

[2658] Richard A. Lanham. Digital literacy. Scientific American, 273(3):198–?? (Intl. ed. 160–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Varian:1995:IE

[2659] Hal R. Varian. The information economy. Scientific American, 273(3): 200–?? (Intl. ed. 161–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zuboff:1995:ENW

[2660] Shoshana Zuboff. The Emperor’s new workplace. Scientific American, 273(3):202–?? (Intl. ed. 162–??), September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lucky:1995:WTA

[2661] Robert W. Lucky. What technology alone cannot do. Scientific American, 273(3):204–?? (Intl. ed. 164–??), September 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1995:MRG

[2662] Ian Stewart. Mathematical recreations: The Great Drain Robbery. Scientific American, 273(3):206–207, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n3/pdf/scientificamerican0995-206.pdf.

Anonymous:1995:MRd

[2663] Anonymous. Mathematical recreations. Scientific American, 273(3): 206–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Re

[2664] Anonymous. Reviews. Scientific American, 273(3):210–215, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n3/pdf/scientificamerican0995-210.pdf.

Anonymous:1995:DD

[2665] Anonymous. Dare to be digital. Scientific American, 273(3):210–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:IBP

[2666] Anonymous. Internet business primers. Scientific American, 273(3): 210–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Rf

[2667] Anonymous. Reviews. Scientific American, 273(3):210–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:WTD

[2668] Anonymous. The way things don’t work on CD-ROMs. Scientific American, 273(3):210–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Penny:1995:EPL

[2669] Simon Penny. Essay: The pursuit of the living machine. Scientific American, 273(3):216, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n3/pdf/scientificamerican0995-
216.pdf.

Anonymous:1995:ESP

[2670] Anonymous. Essay: Simon penny. Scientific American, 273(3):216–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Penny:1995:E

[2671] S. Penny. Essay. Scientific American, 273(3):216–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LEn

[2672] Anonymous. Letters to the Editors. Scientific American, 273(4):10, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-10.pdf.

Anonymous:1995:CC

[2673] Anonymous. Controversy over complexity. Scientific American, 273(4): 10–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SRC

[2674] Anonymous. Smoking: The real costs. Scientific American, 273(4):10–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIh

[2675] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 273(4):12, October 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-12.pdf.

Anonymous:1995:A

[2676] Anonymous. 1845: The autopilot. Scientific American, 273(4):12–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:PRP

[2677] Anonymous. 1895: Pasteur, R.I.P. Scientific American, 273(4):12–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TS

[2678] Anonymous. 1945: Transmitters in the sky. Scientific American, 273 (4):12–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAj

[2679] Anonymous. 50, 100 and 150 years ago. Scientific American, 273(4): 12–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mukerjee:1995:SCH

[2680] Madhusree Mukerjee. Science and the citizen: Hidden scars. Scientific American, 273(4):14–15, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
14.pdf.

Anonymous:1995:AEW

[2681] Anonymous. The analytical economist: A world of debt. Scientific American, 273(4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:CPV

[2682] Anonymous. Chicken pox vaccine in question. Scientific American, 273 (4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:ECa

[2683] Anonymous. Electromagnetic contamination. Scientific American, 273 (4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:ESA

[2684] Anonymous. The endangered species act. Scientific American, 273(4): 14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:HOB

[2685] Anonymous. Human origins on better footing. Scientific American, 273 (4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:IBN

[2686] Anonymous. The ice beyond Neptune. Scientific American, 273(4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:PCG

[2687] Anonymous. Profile: Cosmologist George F. R. Ellis lives on a divided Earth. Scientific American, 273(4):14–??, October 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SAB

[2688] Anonymous. Sexual abuse on the brain. Scientific American, 273(4): 14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SCk

[2689] Anonymous. Science and the citizen. Scientific American, 273(4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SDB

[2690] Anonymous. Sex, death and a backflip. Scientific American, 273(4): 14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SUE

[2691] Anonymous. Secret U.S. export: Sulfates. Scientific American, 273(4): 14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBG

[2692] Anonymous. Technology and business: Good-bye, Commerce Department. Scientific American, 273(4):14–??, October 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:UL

[2693] Anonymous. Undersea laboratory. Scientific American, 273(4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:VC

[2694] Anonymous. Viewing cyberspace. Scientific American, 273(4):14–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wright:1995:SFT

[2695] Karen Wright. Seeing the forest for the trees. Scientific American, 273(4):21–24, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
21.pdf.

Horgan:1995:BN

[2696] John Horgan. Beyond Neptune. Scientific American, 273(4):24–26, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-24.pdf.

Powell:1995:WB

[2697] Corey S. Powell. Why so blue? Scientific American, 273(4):26, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-26a.pdf.

Schneider:1995:HT

[2698] David Schneider. High tension. Scientific American, 273(4):26–31, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-26b.pdf.

Zorpette:1995:ROV

[2699] Glenn Zorpette. 3 rms, ocean view. Scientific American, 273(4):30, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-30.pdf.

Carpi:1995:PP

[2700] John Carpi. A pox on the pox. Scientific American, 273(4):32, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-32b.pdf.

Doyle:1995:DCB

[2701] Rodger Doyle. Deaths caused by breast cancer, by county. Scientific American, 273(4):32, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
32D.pdf.

Mukerjee:1995:GYA

[2702] Madhusree Mukerjee. Giving your all. Scientific American, 273(4):32, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-32a.pdf.

Beardsley:1995:EA

[2703] Tim Beardsley. Endangered again. Scientific American, 273(4):34, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-34a.pdf.

Nemecek:1995:TF

[2704] Sasha Nemecek. A tight fit. Scientific American, 273(4):34–36, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-34b.pdf.

Beardsley:1995:FWM

[2705] Tim Beardsley. These feet were made for walking — and? Scientific American, 273(4):36–37, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
36b.pdf.

Zorpette:1995:MCT

[2706] Glenn Zorpette. More coral trouble. Scientific American, 273(4):36–37, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-36a.pdf.

Wallich:1995:AEB

[2707] Paul Wallich. The analytical economist: Are band-aids enough for Third World debt? Scientific American, 273(4):38, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-38.pdf.

Beardsley:1995:TBS

[2708] Tim Beardsley. Technology and business: Slash and burn. Scientific American, 273(4):39, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
39.pdf.

DeKoker:1995:AT

[2709] Brenda DeKoker. An acid test. Scientific American, 273(4):40, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-40a.pdf.

Yam:1995:WF

[2710] Philip Yam. Writing on the fringe. Scientific American, 273(4):40, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-40b.pdf.

Browning:1995:TNP

[2711] John Browning. Theres no place like cyberspace. Scientific American, 273(4):44–46, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
44.pdf.

Gibbs:1995:SW

[2712] W. Wayt Gibbs. Soft wear. Scientific American, 273(4):46, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-46b.pdf.

Wallich:1995:TG

[2713] Paul Wallich. A tiny Gutenberg. Scientific American, 273(4):46, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-46a.pdf.

Gibbs:1995:PGF

[2714] W. Wayt Gibbs. Profile: George F. R. Ellis: Thinking globally, acting universally. Scientific American, 273(4):50–55, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-50.pdf.

LeGuenno:1995:EV

[2715] Bernard Le Guenno. Emerging viruses. Scientific American, 273(4): 56–64 (Intl. ed. 30–??), October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
56.pdf.

Boss:1995:CYS

[2716] Alan P. Boss. Companions to young stars. Scientific American, 273(4):134–139, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
134.pdf.

Lloyd:1995:QMC

[2717] Seth Lloyd. Quantum-mechanical computers. Scientific American, 273(4):140–145, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
140.pdf.

Loizeaux:1995:DI

[2718] J. Mark Loizeaux and Douglas K. Loizeaux. Demolition by implosion. Scientific American, 273(4):146–153 (Intl. ed. 120–??), October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-146.pdf.

Horgan:1995:TSS

[2719] John Horgan. Trends in social science: The new social Darwinists. Scientific American, 273(4):150–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Axel:1995:MLS

[2720] Richard Axel. The molecular logic of smell. Scientific American, 273(4):154–159, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
154.pdf.

Anonymous:1995:NEC

[2721] Anonymous. The never-ending chess game. Scientific American, 273 (4):158–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Baldwin:1995:SPL

[2722] Neil Baldwin. Science in pictures: The laboratory notebooks of Thomas Edison. Scientific American, 273(4):160–163, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-160.pdf.

Baldwin:1995:LNT

[2723] Neil Baldwin. The laboratory notebooks of Thomas Edison. Scientific American, 273(4):160–?? (Intl. ed. 138–??), October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Davis:1995:TSS

[2724] Devra Lee Davis and H. Leon Bradlow. Trends in social science: Can environmental estrogens cause breast cancer? Scientific American, 273(4):166–172, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
166.pdf.

Davis:1995:CEE

[2725] Devra Lee Davis and H. Leon Bradlow. Can environmental estrogens cause breast cancer? Scientific American, 273(4):166–?? (Intl. ed. 144– ??), October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1995:NSD

[2726] John Horgan. The new social Darwinists. Scientific American, 273(4):174–181, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
174.pdf.

Stewart:1995:MRN

[2727] Ian Stewart. Mathematical recreations: The never-ending chess game. Scientific American, 273(4):182–183, October 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n4/pdf/scientificamerican1095-182.pdf.

Anonymous:1995:MRHa

[2728] Anonymous. Mathematical recreations: How to play a never-ending game of chess. Scientific American, 273(4):182–??, October 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Rg

[2729] Anonymous. Reviews. Scientific American, 273(4):184–190, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n4/pdf/scientificamerican1095-184.pdf.

Anonymous:1995:RAF

[2730] Anonymous. Reviews: Apollo 13 in film, books and reality. Scientific American, 273(4):184–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SM

[2731] Anonymous. Sea monsters. Scientific American, 273(4):184–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holton:1995:ECE

[2732] Gerald Holton. Essay: The controversy over the end of science. Scientific American, 273(4):191, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n4/pdf/scientificamerican1095-
191.pdf.

Holton:1995:EES

[2733] Gerald Holton. Essay: The end of science is nowhere in sight. Scientific American, 273(4):191–??, October 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1995:LEa

[2734] John Rennie. Letter from the Editors. Scientific American, 273(5):6, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-6.pdf.

Anonymous:1995:LEo

[2735] Anonymous. Letters to the Editors. Scientific American, 273(5):8, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-8.pdf.

Anonymous:1995:CT

[2736] Anonymous. The counterfeiting threat. Scientific American, 273(5):8–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:HW

[2737] Anonymous. Harvard’s women. Scientific American, 273(5):8–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:RWN

[2738] Anonymous. Red wolves: A new species? Scientific American, 273 (5):8–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIi

[2739] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 273(5):10, November 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n5/pdf/scientificamerican1195-10.pdf.

Anonymous:1995:DW

[2740] Anonymous. 1945: DDT warning. Scientific American, 273(5):10–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LB

[2741] Anonymous. 1895: Loss of the bison. Scientific American, 273(5):10–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBh

[2742] Anonymous. 1845: Telegraph balloons. Scientific American, 273(5): 10–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAk

[2743] Anonymous. 50, 100 and 150 years ago. Scientific American, 273(5): 10–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1995:SCG

[2744] John Horgan. Science and the citizen: Get smart, take a test. Scientific American, 273(5):12–14, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
12.pdf.

Anonymous:1995:AAS

[2745] Anonymous. Algae against sewage. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AET

[2746] Anonymous. The analytical economist: Taxes and the female workforce. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AM

[2747] Anonymous. Antarctic meltdown. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AO

[2748] Anonymous. Attractive odors. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:FCF

[2749] Anonymous. Fiberglass and cancer fears. Scientific American, 273(5): 12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:GGU

[2750] Anonymous. “gay genes” under new scrutiny. Scientific American, 273 (5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:LNS

[2751] Anonymous. Linking nerves to silicon. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:MHD

[2752] Anonymous. Mapping heart disease. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:PKR

[2753] Anonymous. Profile: Kay Redfield Jamison talks of moods and madness. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:RI

[2754] Anonymous. Rising IQ. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBCa

[2755] Anonymous. Technology and business: Congress tackles technology without advice. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TR

[2756] Anonymous. Thalidomide rehabilitated. Scientific American, 273(5): 12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:VM

[2757] Anonymous. Volcano music. Scientific American, 273(5):12–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1995:AP

[2758] David Schneider. Attracted to the pole. Scientific American, 273(5):14, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-14.pdf.

Yam:1995:FMG

[2759] Philip Yam. Fiber that may not be good for you. Scientific American, 273(5):15–17, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
15.pdf.

Horgan:1995:BT

[2760] John Horgan. The big thaw. Scientific American, 273(5):17–20, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-17.pdf.

Mirsky:1995:NI

[2761] Steve Mirsky. The noses have it. Scientific American, 273(5):20, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-20b.pdf.

Nemecek:1995:THJ

[2762] Sasha Nemecek. Transforming Hyde into Jekyll. Scientific American, 273(5):20–22, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
20a.pdf.

Doyle:1995:MDC

[2763] Rodger Doyle. Male deaths caused by heart disease, by county. Scientific American, 273(5):22, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
22.pdf.

Perrin:1995:NEF

[2764] Nicola Perrin. A never-ending feast. Scientific American, 273(5):24, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-24b.pdf.

Wallich:1995:P

[2765] Paul Wallich. Out of place. Scientific American, 273(5):24, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-24a.pdf.

Horgan:1995:GGR

[2766] John Horgan. Gay genes, revisited. Scientific American, 273(5):26, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-26a.pdf.

Beardsley:1995:SSb

[2767] Tim Beardsley. Solar secrets. Scientific American, 273(5):26–29, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-26b.pdf.

Schneider:1995:CM

[2768] David Schneider. Country music. Scientific American, 273(5):28, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-28.pdf.

Wallich:1995:SWM

[2769] Paul Wallich. Some women are more equal than others. Scientific American, 273(5):29, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
29.pdf.

Zorpette:1995:LH

[2770] Glenn Zorpette. Luddites on the Hill. Scientific American, 273(5):30, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-30.pdf.

Saga:1995:WIF

[2771] Carl Saga. The world’s imperiled fish. Scientific American, 273(5):30–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1995:CAb

[2772] Tim Beardsley. Clearing the air. Scientific American, 273(5):32, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-32a.pdf.

Yam:1995:MMM

[2773] Philip Yam. Mind meets machine, sort of. Scientific American, 273(5):32, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-32b.pdf.

Wallich:1995:MV

[2774] Paul Wallich. Meta-virus. Scientific American, 273(5):34, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-34.pdf.

Browning:1995:RP

[2775] John Browning. Rights of privacy. Scientific American, 273(5):36–42, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-36.pdf.

Gibbs:1995:STP

[2776] W. Wayt Gibbs. Sewage treatment plants. Scientific American, 273(5):42, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
42.pdf.

Leutwyler:1995:CTM

[2777] Kristin Leutwyler. Coming through madness. Scientific American, 273(5):44–45, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
44.pdf.

Safina:1995:WIF

[2778] Carl Safina. The world’s imperiled fish. Scientific American, 273(5):46–53, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
46.pdf.

Streit:1995:BIS

[2779] Wolfgang J. Streit and Carol A. Kincaid-Colton. The brain’s immune system. Scientific American, 273(5):54–61, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n5/pdf/scientificamerican1195-54.pdf.

Broecker:1995:CC

[2780] Wallace S. Broecker. Chaotic climate. Scientific American, 273(5):62–68, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-62.pdf.

Psaltis:1995:HM

[2781] Demetri Psaltis and Fai Mok. Holographic memories. Scientific American, 273(5):70–76 (Intl. ed. 52–58), November 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n5/pdf/scientificamerican1195-70.pdf.

Milner:1995:CDL

[2782] Richard Milner. Charles Darwin: The last portrait. Scientific American, 273(5):78, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
78.pdf.

Dawkins:1995:GUFa

[2783] Richard Dawkins. God’s utility function. Scientific American, 273(5):80–85, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
80.pdf.

Carlson:1995:MWH

[2784] Shawn Carlson. Measuring the wind with hot metal. Scientific American, 273(5):80–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dawkins:1995:GUFb

[2785] Richard Dawkins. God’s utility function. Scientific American, 273(5): 80–?? (Intl. ed. 62–??), November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:WTS

[2786] Anonymous. Ways to tile space with knots. Scientific American, 273 (5):82–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Farmelo:1995:DXR

[2787] Graham Farmelo. The discovery of X-rays. Scientific American, 273(5):86–91, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
86.pdf.

Beek:1995:SJ

[2788] Peter J. Beek and Arthur Lewbel. The science of juggling. Scientific American, 273(5):92–97, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
92.pdf.

Carlson:1995:ASMa

[2789] Shawn Carlson. The amateur scientist: Measuring the wind with hot metal. Scientific American, 273(5):98–99, November 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n5/pdf/scientificamerican1195-98.pdf.

Anonymous:1995:ASMa

[2790] Anonymous. The amateur scientist: Measuring wind speed in tight places. Scientific American, 273(5):98–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1995:MRW

[2791] Ian Stewart. Mathematical recreations: Ways to tile space with knots. Scientific American, 273(5):100–101, November 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n5/pdf/scientificamerican1195-100.pdf.

Anonymous:1995:MRHb

[2792] Anonymous. Mathematical recreations: How to fill space with knots and doughnuts. Scientific American, 273(5):100–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sheaffer:1995:RCR

[2793] Robert Sheaffer, Mark Monmonier, and James S. Trefil. Reviews and commentaries: Reviews. Scientific American, 273(5):102–107, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-102.pdf.

Anonymous:1995:ACR

[2794] Anonymous. Atlases on CD-ROM. Scientific American, 273(5):102–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:ECE

[2795] Anonymous. Extremely close encounters. Scientific American, 273(5): 102–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:MWB

[2796] Anonymous. Morrison’s “Wonders” and Burke’s “Connections”. Scientific American, 273(5):102–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:RCa

[2797] Anonymous. Reviews and commentaries. Scientific American, 273(5): 102–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SFb

[2798] Anonymous. Science-in-fiction. Scientific American, 273(5):102–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1995:CNA

[2799] Philip Morrison. Commentaries: On neutrino astronomy. Scientific American, 273(5):108–109, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
108.pdf.

Burke:1995:SR

[2800] James Burke. The Silk Road. Scientific American, 273(5):109–110, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n5/pdf/scientificamerican1195-109.pdf.

Eisenberg:1995:EDB

[2801] Anne Eisenberg. Essay: Doing business on the net. Scientific American, 273(5):111, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n5/pdf/scientificamerican1195-
111.pdf.

Eisenberg:1995:EEC

[2802] Anne Eisenberg. Essay: Electronic commerce could drop the net on personal privacy. Scientific American, 273(5):111–??, November 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1995:LEb

[2803] John Rennie. Letter from the Editors. Scientific American, 273(6):8, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-8.pdf.

Anonymous:1995:LEp

[2804] Anonymous. Letters to the Editors. Scientific American, 273(6):10, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-10.pdf.

Anonymous:1995:CD

[2805] Anonymous. Creationism in disguise. Scientific American, 273(6):10–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:OS

[2806] Anonymous. Overlooked science. Scientific American, 273(6):10–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TND

[2807] Anonymous. The trebuchet next door. Scientific American, 273(6):10–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAIj

[2808] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 273(6):12, December 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-12.pdf.

Anonymous:1995:APP

[2809] Anonymous. 1945: Atomic power prediction. Scientific American, 273 (6):12–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:Ef

[2810] Anonymous. 1895: The electric hen. Scientific American, 273(6):12–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:UC

[2811] Anonymous. 1845: Uncountable comets. Scientific American, 273(6): 12–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:YAl

[2812] Anonymous. 50, 100 and 150 years ago. Scientific American, 273(6): 12–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1995:SCT

[2813] David Schneider. Science and the citizen: Tempest on the high sea. Scientific American, 273(6):14–16, December 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-14.pdf.

Anonymous:1995:AEI

[2814] Anonymous. The analytical economist: Indexing inflation. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:CI

[2815] Anonymous. Crowning the IgNobility. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:DGW

[2816] Anonymous. Digesting global warming. Scientific American, 273(6): 14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:DS

[2817] Anonymous. Danger at sea. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:FPP

[2818] Anonymous. The FAA puts planes in free flight. Scientific American, 273 (6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:GPC

[2819] Anonymous. The genetics (and politics) of crime. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:GRT

[2820] Anonymous. Golfers road test hydrogen cars. Scientific American, 273 (6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:OFD

[2821] Anonymous. Oily Federal deals. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:PMG

[2822] Anonymous. Profile: Martin Gardner, alias Dr. Matrix, the mathematical gamester. Scientific American, 273(6):14–??, December 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:RKU

[2823] Anonymous. Rebellious kids in utero. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SL

[2824] Anonymous. Sign language. Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:TBS

[2825] Anonymous. Technology and business: Star Wars is back: So what? Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:UWO

[2826] Anonymous. Uh, where’s the outlet? Scientific American, 273(6):14–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1995:PP

[2827] David Schneider. Plug and play. Scientific American, 273(6):16, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-16.pdf.

Gibbs:1995:GE

[2828] W. Wayt Gibbs. Great expectations. Scientific American, 273(6):18, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-18b.pdf.

Horgan:1995:SB

[2829] John Horgan. A sign is born. Scientific American, 273(6):18–19, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-18a.pdf.

Mirsky:1995:HSH

[2830] Steve Mirsky. Home, sweet home. Scientific American, 273(6):19, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-19b.pdf.

Beardsley:1995:CP

[2831] Tim Beardsley. Crime and punishment. Scientific American, 273(6):19–22, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
19a.pdf.

Doyle:1995:S

[2832] Rodger Doyle. Suicide. Scientific American, 273(6):22, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-22.pdf.

Mirsky:1995:YMA

[2833] Steve Mirsky. You may already be a Wiener. Scientific American, 273(6):24, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
24a.pdf.

Stykes:1995:OIW

[2834] Mervin Stykes. And those other Ig winners are. … Scientific American, 273(6):24, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
24b.pdf.

Mukerjee:1995:AF

[2835] Madhusree Mukerjee. About face. Scientific American, 273(6):25, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-25b.pdf.

Horgan:1995:SW

[2836] John Horgan. The struggle within. Scientific American, 273(6):25–26, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-25a.pdf.

Zorpette:1995:BHS

[2837] Glenn Zorpette. Buy high, sell low. Scientific American, 273(6):28, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-28a.pdf.

Gibbs:1995:SLI

[2838] W. Wayt Gibbs. Some like it hot. Scientific American, 273(6):28–30, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-28b.pdf.

Wallich:1995:AEC

[2839] Paul Wallich. The analytical economist: The confusing price index. Scientific American, 273(6):30, December 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-30.pdf.

Dupont:1995:TBG

[2840] Daniel Dupont. Technology and business: The GOP strikes back. Scientific American, 273(6):32, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
32.pdf.

Gibbs:1995:AAb

[2841] W. Wayt Gibbs. Artificial art. Scientific American, 273(6):34, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-34b.pdf.

Gibbs:1995:FAF

[2842] W. Wayt Gibbs. Free-for-all flights. Scientific American, 273(6):34–37, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-34a.pdf.

Stix:1995:NL

[2843] Gary Stix. Nice legs. Scientific American, 273(6):36–37, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-36.pdf.

Schneider:1995:PG

[2844] David Schneider. Putting greens. Scientific American, 273(6):37–38, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v273/n6/pdf/scientificamerican1295-37.pdf.

Yam:1995:PMG

[2845] Philip Yam. Profile: Martin Gardner: The mathematical gamester. Scientific American, 273(6):38–41, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
38.pdf.

Johnson:1995:GM

[2846] Torrence V. Johnson. The Galileo mission. Scientific American, 273(6):44–51, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
44.pdf.

Welsh:1995:CF

[2847] Michael J. Welsh and Alan E. Smith. Cystic fibrosis. Scientific American, 273(6):52–59, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
52.pdf.

Heiniger:1995:SPL

[2848] Paolo Heiniger. Science in pictures: The Leaning Tower of Pisa. Scientific American, 273(6):62–67, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
62.pdf.

Heiniger:1995:LTP

[2849] Paolo Heiniger. The Leaning Tower of Pisa. Scientific American, 273 (6):62–?? (Intl. ed. 44–??), December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hyndman:1995:GEP

[2850] Roy D. Hyndman. Giant earthquakes of the Pacific Northwest. Scientific American, 273(6):68–75, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
68.pdf.

Stix:1995:TDTa

[2851] Gary Stix. Trends in defense technology: Fighting future wars. Scientific American, 273(6):74–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Newman:1995:HBM

[2852] Jack Newman. How breast milk protects newborns. Scientific American, 273(6):76–79, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
76.pdf.

Chalmers:1995:PCE

[2853] David J. Chalmers. The puzzle of conscious experience. Scientific American, 273(6):80–86, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
80.pdf.

Carlson:1995:MMS

[2854] Shawn Carlson. Measuring the metabolism of small organisms. Scientific American, 273(6):84–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:AP

[2855] Anonymous. The anthropomorphic principle. Scientific American, 273 (6):86–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beth:1995:CCI

[2856] Thomas Beth. Confidential communication on the Internet. Scientific American, 273(6):88–91 (Intl. ed. 70–73), December 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-88.pdf.

Stix:1995:TDTb

[2857] Gary Stix. Trends in defense technology: Fighting future wars. Scientific American, 273(6):92–98, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
92.pdf.

Stix:1995:FFW

[2858] Gary Stix. Fighting future wars. Scientific American, 273(6):92–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carlson:1995:ASMb

[2859] Shawn Carlson. The amateur scientist: Measuring the metabolism of small organisms. Scientific American, 273(6):102–103, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-102.pdf.

Anonymous:1995:ASMb

[2860] Anonymous. The amateur scientist: Measuring micrometabolism: How fast does a beetle breathe? Scientific American, 273(6):102–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1995:MRA

[2861] Ian Stewart. Mathematical recreations: The anthropomorphic principle. Scientific American, 273(6):104–106, December 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v273/
n6/pdf/scientificamerican1295-104.pdf.

Anonymous:1995:MRM

[2862] Anonymous. Mathematical recreations: Murphy’s Law demystified: Why toast falls butter-side down. Scientific American, 273(6):104–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1995:RCB

[2863] Philip Morrison and Phylis Morrison. Reviews and commentaries: The Scientific American Young Readers Book Awards. Scientific American, 273(6):108–114, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
108.pdf.

Anonymous:1995:CSS

[2864] Anonymous. Connections: Springs, steel and W.C.’s. Scientific American, 273(6):108–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:RCb

[2865] Anonymous. Reviews and commentaries. Scientific American, 273(6): 108–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:SAY

[2866] Anonymous. The Scientific American Young Readers Book Awards. Scientific American, 273(6):108–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1995:CTB

[2867] James Burke. Commentaries: Top to bottom. Scientific American, 273(6):115–116, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
115.pdf.

Anonymous:1995:AIA

[2868] Anonymous. Annual index: Annual index 1995. Scientific American, 273(6):117–119, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
117.pdf.

Boyk:1995:EEP

[2869] James Boyk. Essay: The endangered piano technician. Scientific American, 273(6):120, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v273/n6/pdf/scientificamerican1295-
120.pdf.

Anonymous:1995:EJB

[2870] Anonymous. Essay: James Boyk: Some of the most virtuoso piano talent never perform onstage. Scientific American, 273(6):120–??, December 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1996:LEa

[2871] John Rennie. Letter from the Editor. Scientific American, 274(1):6, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-6.pdf.

Anonymous:1996:LEF

[2872] Anonymous. Letters to the editors: Fly the crowded skies. how much energy?. the dilemmas of AIDS. Scientific American, 274(1):6–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0196issue/0196toc.html.

Anonymous:1996:YAM

[2873] Anonymous. 50, 100 and 150 years ago: 1946: Making high-octane gasoline 1896: The missing link in Java. 1846: Mesmerizing crime. Scientific American, 274(1):7–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Anonymous:1996:SCCa

[2874] Anonymous. Science and the citizen: Culture and mental illness. RNA and the origin of life. space junk. quantum erasers. resistant microbes. the studs of science. new planets. the analytical economist — gutting social research. Scientific American, 274(1):8–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0196issue/0196toc.html.

Anonymous:1996:Ea

[2875] Anonymous. Erratum. Scientific American, 274(1):10, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n1/pdf/scientificamerican0196-10b.pdf.

Anonymous:1996:LEa

[2876] Anonymous. Letters to the Editors. Scientific American, 274(1):10, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-10a.pdf.

Anonymous:1996:YAIa years ago. Scientific American, 274(1):12, January 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n1/pdf/scientificamerican0196-12.pdf.

Stix:1996:SCL

[2878] Gary Stix. Science and the citizen: Listening to culture. Scientific American, 274(1):16–21, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
16.pdf.

Mukerjee:1996:CTI

[2879] Madhusree Mukerjee. Changing their image. Scientific American, 274(1): 21, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-21.pdf.

Powell:1996:SD

[2880] Corey S. Powell. Star dreck. Scientific American, 274(1):22, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-22a.pdf.

Powell:1996:SP

[2881] Corey S. Powell. Strange places. Scientific American, 274(1):22–23, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-22b.pdf.

Nemecek:1996:VP

[2882] Sasha Nemecek. Virtual pollution. Scientific American, 274(1):24, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-24.pdf.

Mirsky:1996:WGY

[2883] Steve Mirsky. Into the wild green yonder. Scientific American, 274(1):25, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-25.pdf.

Beardsley:1996:RR

[2884] Tim Beardsley. Resisting resistance. Scientific American, 274(1):26, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-26.pdf.

Williams:1996:RTNa

[2885] Phil Williams and Paul N. Woessner. The real threat of nuclear smuggling. Scientific American, 274(1):26–??, January 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0196issue/0196toc.html.

Doyle:1996:CCM

[2886] Rodger Doyle. Colorectal cancer mortality among men. Scientific American, 274(1):27, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
27a.pdf.

Horgan:1996:WAR

[2887] John Horgan. The world according to RNA. Scientific American, 274(1):27–30, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
27b.pdf.

Yam:1996:RQE

[2888] Philip Yam. Rubbed out with the quantum eraser. Scientific American, 274(1):30–31, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
30.pdf.

Nemecek:1996:RRW

[2889] Sasha Nemecek. Return of the red wolf. Scientific American, 274(1):31– 32, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-31.pdf.

Wallich:1996:AEF

[2890] Paul Wallich. The analytical economist: Flying blind. Scientific American, 274(1):32, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
32.pdf.

Zorpette:1996:TBR

[2891] Glenn Zorpette. Technology and business: Return of the breeder. Scientific American, 274(1):33, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
33.pdf.

Browning:1996:MFS

[2892] John Browning. Making free software pay. Scientific American, 274(1): 35, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-35a.pdf.

Seife:1996:F

[2893] Charles Seife. Freewheeling. Scientific American, 274(1):35, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-35b.pdf.

Abate:1996:MH

[2894] Tom Abate. The midnight hour. Scientific American, 274(1):36–37, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-36b.pdf.

Gibbs:1996:LM

[2895] W. Wayt Gibbs. Light over matter. Scientific American, 274(1):36–37, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-36a.pdf.

Gibbs:1996:PSF

[2896] W. Wayt Gibbs. Playing Slartibartfast with fractals. Scientific American, 274(1):36–37, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
36c.pdf.

Landau:1996:PJR

[2897] Susan Landau. Profile: Joseph Rotblat: From fission research to a prize for peace. Scientific American, 274(1):38–39, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n1/pdf/scientificamerican0196-38.pdf.

Williams:1996:RTNb

[2898] Phil Williams and Paul N. Woessner. The real threat of nuclear smuggling. Scientific American, 274(1):40–44, January 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n1/pdf/scientificamerican0196-40.pdf.

Hutcheson:1996:TESa

[2899] G. Dan Hutcheson and Jerry D. Hutcheson. Technology and economics in the semiconductor industry. Scientific American, 274(1):40–46, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0196issue/0196toc.html.

Weindruch:1996:CRA

[2900] Richard Weindruch. Caloric restriction and aging. Scientific American, 274(1):46–52, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
46.pdf;   http://www.sciam.com/0196issue/0196toc.html.

Hutcheson:1996:TESb

[2901] G. Dan Hutcheson and Jerry D. Hutcheson. Technology and economics in the semiconductor industry. Scientific American, 274(1):54–62, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-54.pdf.

Grillner:1996:NNV

[2902] Sten Grillner. Neural networks for vertebrate locomotion. Scientific American, 274(1):64–69, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
64.pdf;   http://www.sciam.com/0196issue/0196toc.html.

Malle:1996:CRR

[2903] Karl Geert Malle. Cleaning up the river Rhine. Scientific American, 274(1):70–75, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
70.pdf;   http://www.sciam.com/0196issue/0196toc.html.

Mukerjee:1996:TTPa

[2904] Madhusree Mukerjee. Trends in theoretical physics: Explaining everything. Scientific American, 274(1):72–??, January 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0196issue/0196toc.html.

Taylor:1996:ECC

[2905] S. Ross Taylor and Scott M. McLennan. The evolution of continental crust. Scientific American, 274(1):76–81, January 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n1/pdf/scientificamerican0196-76.pdf;   http://www.sciam.com/
0196issue/0196toc.html.

Anonymous:1996:ASH

[2906] Anonymous. The amateur scientist: How to record and collect the sounds of nature. Scientific American, 274(1):80–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0196issue/0196toc.html.

Anonymous:1996:RNS

[2907] Anonymous. Recording nature’s sounds. Scientific American, 274(1):80– ??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/0196issue/0196toc.html.

Schmidt:1996:SPW

[2908] Michael J. Schmidt. Science in pictures: Working elephants. Scientific American, 274(1):82–87, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
82.pdf.

Anonymous:1996:MRSa

[2909] Anonymous. Mathematical recreations: The slippery puzzle under Mother Worm’s Blanket. Scientific American, 274(1):82–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0196issue/0196toc.html.

Schmidt:1996:WE

[2910] Michael J. Schmidt. Working elephants. Scientific American, 274(1): 82–?? (Intl. ed. 66–??), January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Anonymous:1996:RCW

[2911] Anonymous. Reviews and commentaries: The why of sex. hypertext. Scientific American, 274(1):84–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Burke:1996:CHCa

[2912] James Burke. Connections: Hydraulics and cornflakes. Scientific American, 274(1):84–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Morrison:1996:WCN

[2913] Philip Morrison. Wonders: A century of new physics. Scientific American, 274(1):84–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Mukerjee:1996:TTPb

[2914] Madhusree Mukerjee. Trends in theoretical physics: Explaining everything. Scientific American, 274(1):88–94, January 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n1/pdf/scientificamerican0196-88.pdf.

Mukerjee:1996:EE

[2915] Madhusree Mukerjee. Explaining everything. Scientific American, 274 (1):88–?? (Intl. ed. 72–??), January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Carlson:1996:ASR

[2916] Shawn Carlson. The amateur scientist: Recording nature’s sounds. Scientific American, 274(1):96–97, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
96.pdf.

Anonymous:1996:AS

[2917] Anonymous. Amateur scientist. Scientific American, 274(1):96–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0196issue/0196toc.html.

deDuve:1996:ELW

[2918] Christian de Duve. The evolution of life was not so unlikely after all. Scientific American, 274(1):96–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0196issue/0196toc.html.

Stewart:1996:MRMa

[2919] Ian Stewart. Mathematical recreations: Mother worm’s blanket. Scientific American, 274(1):98–99, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
98.pdf.

Anonymous:1996:RCa

[2920] Anonymous. Reviews and commentaries. Scientific American, 274(1):102–105, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
102.pdf.

Morrison:1996:HSC

[2921] Philip Morrison. Head start on the 20th Century. Scientific American, 274(1):106–107, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
106.pdf.

Burke:1996:BT

[2922] James Burke. Breakfast thoughts. Scientific American, 274(1):107–108, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n1/pdf/scientificamerican0196-107.pdf.

deDuve:1996:ECC

[2923] Christian de Duve. Essay: The constraints of chance. Scientific American, 274(1):112, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n1/pdf/scientificamerican0196-
112.pdf.

Rennie:1996:LEb

[2924] John Rennie. Letter from the Editor. Scientific American, 274(2):8, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-8.pdf.

Anonymous:1996:Eb

[2925] Anonymous. Erratum. Scientific American, 274(2):10, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n2/pdf/scientificamerican0296-10b.pdf.

Anonymous:1996:LEb

[2926] Anonymous. Letters to the Editors. Scientific American, 274(2):10, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-10a.pdf.

Anonymous:1996:LES

[2927] Anonymous. Letters to the editors: Sex, evolution and psychology. identifying estrogens. physics survives. Scientific American, 274(2):10–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/0296issue/0296toc.html.

Anonymous:1996:YAIb

[2928] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 274(2):12, February 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n2/pdf/scientificamerican0296-12.pdf.

Anonymous:1996:YAN

[2929] Anonymous. 50, 100 and 150 years ago: 1946: No atomic cars. 1896: Spanish meteorite. 1846: Fire extinguisher. Scientific American, 274(2):

12–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/0296issue/0296toc.
html.

Stix:1996:SCK

[2930] Gary Stix. Science and the citizen: Keeping vaccines cold. Scientific American, 274(2):14–16, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
14.pdf.

Anonymous:1996:AEC

[2931] Anonymous. The analytical economist: Communism’s new capitalist clothes. Scientific American, 274(2):14–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0296issue/0296toc.html.

Anonymous:1996:PDCa

[2932] Anonymous. Profile: Daniel C. Dennett explains consciousness and unleashes Darwin. Scientific American, 274(2):14–??, February 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0296issue/0296toc.html.

Anonymous:1996:SCDa

[2933] Anonymous. Science and the citizen: Delivering vaccines. cooling budgets at the South Pole. schizophrenia. virus amok down under. how many doctors in the house? why France really wants nuclear tests. the minority majority. radar into the past. water fluoridation. Scientific American, 274(2):14–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Anonymous:1996:TBI

[2934] Anonymous. Technology and business: Insurers fret over climate change the evolving Internet. history lesson for Bill Gates. Scientific American, 274(2):14–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Yam:1996:BKP

[2935] Philip Yam. Bose knows P.R. Scientific American, 274(2):16, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-16.pdf.

Horgan:1996:FA

[2936] John Horgan. Frozen assets. Scientific American, 274(2):18–19, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-18a.pdf.

Powell:1996:RR

[2937] Corey S. Powell. Radar range. Scientific American, 274(2):18–19, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-18b.pdf.

Doyle:1996:F

[2938] Rodger Doyle. Fluoridation. Scientific American, 274(2):20, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-20a.pdf.

Schneider:1996:DG

[2939] David Schneider. The doctor glut. Scientific American, 274(2):20–21, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-20b.pdf.

Stix:1996:RM

[2940] Gary Stix. The rainbow majority. Scientific American, 274(2):22, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-22b.pdf.

Leutwyler:1996:SR

[2941] Kristin Leutwyler. Schizophrenia revisited. Scientific American, 274(2):22–23, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
22a.pdf.

Zorpette:1996:GB

[2942] Glenn Zorpette. Going out with a bang. Scientific American, 274(2):23, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-23.pdf.

Mirsky:1996:SM

[2943] Steve Mirsky. Sinister movements. Scientific American, 274(2):24, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-24b.pdf.

Seife:1996:HS

[2944] Charles Seife. A harebrained scheme. Scientific American, 274(2):24–26, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-24a.pdf.

Wallich:1996:REE

[2945] Paul Wallich. Reaching an economic event horizon. Scientific American, 274(2):26, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
26.pdf.

Stix:1996:GP

[2946] Gary Stix. Green policies. Scientific American, 274(2):27–28, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-27.pdf.

Gibbs:1996:NWS

[2947] W. Wayt Gibbs. A new way to spell relief: V-e-n-o-m. Scientific American, 274(2):28–30, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
28.pdf.

Browning:1996:T

[2948] John Browning. Teleconomics. Scientific American, 274(2):30–32b, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-30.pdf.

Wallich:1996:MRR

[2949] Paul Wallich. More rules of the road. Scientific American, 274(2):31, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-31.pdf.

Horgan:1996:BGA

[2950] John Horgan. Bill Gates’s apocryphal history. Scientific American, 274(2):32–32b, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
32.pdf.

Beardsley:1996:DDI

[2951] Tim Beardsley. Dennett’s dangerous idea. Scientific American, 274(2):34–35, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
34.pdf.

Brown:1996:MPI

[2952] J. Larry Brown and Ernesto Pollitt. Malnutrition, poverty and intellectual development. Scientific American, 274(2):38–43, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-38.pdf;   http:/
/www.sciam.com/0296issue/0296toc.html.

Herring:1996:GPS

[2953] Thomas A. Herring. The Global Positioning System. Scientific American, 274(2):44–50 (Intl. ed. 32–38), February 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n2/pdf/scientificamerican0296-44.pdf.

Buckingham:1996:SUB

[2954] Michael J. Buckingham, John R. Potter, and Chad L. Epifanio. Seeing underwater with background noise. Scientific American, 274(2): 86–90 (Intl. ed. 40–44), February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
86.pdf;   http://www.sciam.com/0296issue/0296toc.html.

Greider:1996:TTC

[2955] Carol W. Greider and Elizabeth H. Blackburn. Telomeres, telomerase and cancer. Scientific American, 274(2):92–97, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-92.pdf;   http:/
/www.sciam.com/0296issue/0296toc.html.

Veilleux:1996:CGE

[2956] Sylvain Veilleux, Gerald Cecil, and Jonathan Bland-Hawthorn. Colossal galactic explosions. Scientific American, 274(2):98–103, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-98.pdf;   http:/
/www.sciam.com/0296issue/0296toc.html.

Blaser:1996:BBU

[2957] Martin J. Blaser. The bacteria behind ulcers. Scientific American, 274(2):104–107, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
104.pdf;   http://www.sciam.com/0296issue/0296toc.html.

Schiebinger:1996:LPa

[2958] Londa Schiebinger. The loves of the plants. Scientific American, 274(2):110–115, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
110.pdf.

Anonymous:1996:GSL

[2959] Anonymous. Growing seedlings at less than 1 G. Scientific American, 274(2):110–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Schiebinger:1996:LPb

[2960] Londa Schiebinger. The loves of the plants. Scientific American, 274 (2):110–?? (Intl. ed. 98–??), February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Anonymous:1996:PPI

[2961] Anonymous. Proof of purchase on the Internet. Scientific American, 274(2):112–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Weingarten:1996:QC

[2962] Donald H. Weingarten. Quarks by computer. Scientific American, 274(2):116–120, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
116.pdf.

Carlson:1996:ASG

[2963] Shawn Carlson. The amateur scientist: Growing seedlings at less than 1 G. Scientific American, 274(2):122–123, February 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n2/pdf/scientificamerican0296-122.pdf.

Anonymous:1996:ASSa

[2964] Anonymous. The amateur scientist: Simulating how plants would grow on Mars. Scientific American, 274(2):122–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0296issue/0296toc.html.

Stewart:1996:MRPa

[2965] Ian Stewart. Mathematical recreations: Proof of purchase on the Internet. Scientific American, 274(2):124–125, February 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n2/pdf/scientificamerican0296-124.pdf.

Anonymous:1996:MRZ

[2966] Anonymous. Mathematical recreations: Zero-based transactions: they know that you know that they know. Scientific American, 274(2):124–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/0296issue/0296toc.html.

Anonymous:1996:RCb

[2967] Anonymous. Reviews and commentaries. Scientific American, 274 (2):126–129, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
126.pdf.

Anonymous:1996:RCSa

[2968] Anonymous. Reviews and commentaries: Star guides. ecology in error? Scientific American, 274(2):126–??, February 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0296issue/0296toc.html.

Burke:1996:CFTa

[2969] James Burke. Connections: Fairy tales and photoelectricity. Scientific American, 274(2):126–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Morrison:1996:WBB

[2970] Philip Morrison and Phylis Morrison. Wonders: Binary beauty. Scientific American, 274(2):126–??, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0296issue/0296toc.html.

Morrison:1996:PBN

[2971] Philip Morrison and Phylis Morrison. The physics of binary numbers. Scientific American, 274(2):130–131, February 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n2/pdf/scientificamerican0296-130.pdf.

Burke:1996:ST

[2972] James Burke. Show time! Scientific American, 274(2):132–133, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n2/pdf/scientificamerican0296-132.pdf.

Randi:1996:EIM

[2973] James Randi. Essay: Investigating miracles, Italian-style. Scientific American, 274(2):136, February 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n2/pdf/scientificamerican0296-
136.pdf.

Randi:1996:EWM

[2974] James Randi. Essay: These weeping Madonnas are less than miraculous. Scientific American, 274(2):136–??, February 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0296issue/0296toc.html.

Rennie:1996:LEc

[2975] John Rennie. Letter from the Editor. Scientific American, 274(3):6, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-6.pdf.

Anonymous:1996:LEc

[2976] Anonymous. Letters to the Editors. Scientific American, 274(3):8, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-8.pdf.

Anonymous:1996:LER

[2977] Anonymous. Letters to the editors: Rising IQs. life’s purpose. alien abductions and freud. Scientific American, 274(3):8–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0396issue/0396toc.html.

Anonymous:1996:YAIc

[2978] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 274(3):10, March 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-10.pdf.

Anonymous:1996:YAX

[2979] Anonymous. 50, 100 and 150 years ago: 1945: X-rays in factories. 1896: A pioneer of flight. 1846: Bigfoot or tall tale? Scientific American, 274(3):10–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0396issue/0396toc.html.

Wallich:1996:SCR

[2980] Paul Wallich and Madhusree Mukerjee. Science and the citizen: Regulating the body business. Scientific American, 274(3):12–14, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-12.pdf.

Anonymous:1996:AEW

[2981] Anonymous. The analytical economist: Women’s real economic prospects. Scientific American, 274(3):12–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0396issue/0396toc.html.

Anonymous:1996:PAL

[2982] Anonymous. Profile: Albert Libchaber brings order to chaos studies. Scientific American, 274(3):12–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0396issue/0396toc.html.

Anonymous:1996:SCHa

[2983] Anonymous. Science and the citizen: How much for the liver?. NASA and nausea. helium shortage. Brazil’s lost desert. thirsty moths. viruses trace neurons. cosmic rays. Escher for the ear. getting Washington’s goats. Scientific American, 274(3):12–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0396issue/0396toc.html.

Anonymous:1996:TBSa

[2984] Anonymous. Technology and business: The scoop on plutonium processing military prototypes in Bosnia. public-key encryption at risk. Scientific American, 274(3):12–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0396issue/0396toc.html.

Yam:1996:EE

[2985] Philip Yam. Escher for the ear. Scientific American, 274(3):14, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-14.pdf.

Schneider:1996:PS

[2986] David Schneider. Plane scary. Scientific American, 274(3):16, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-16b.pdf.

Powell:1996:W

[2987] Corey S. Powell. Out of this world. Scientific American, 274(3):16–18, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-16a.pdf.

Leutwyler:1996:VT

[2988] Kristin Leutwyler. Viral tracers. Scientific American, 274(3):18, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-18.pdf.

Schneider:1996:RFC

[2989] David Schneider. Rain forest crunch. Scientific American, 274(3):19, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-19.pdf.

Nemecek:1996:RET

[2990] Sasha Nemecek. Rescuing an endangered tree. Scientific American, 274(3):22, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
22b.pdf.

Wright:1996:GG

[2991] Karen Wright. Getting the goats. Scientific American, 274(3):22–26, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-22a.pdf.

Doyle:1996:SMM

[2992] Rodger Doyle. Stroke mortality in men ages 35 to 74. Scientific American, 274(3):26, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
26.pdf.

Mirsky:1996:AHZ

[2993] Steve Mirsky. It’s all happening at the zoo(logy meeting). Scientific American, 274(3):28, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
28b.pdf.

Powell:1996:NLM

[2994] Corey S. Powell. No light matter. Scientific American, 274(3):28–30, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-28a.pdf.

Wu:1996:PSP

[2995] Kai Wu. Pass the salt, please. Scientific American, 274(3):30, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-30.pdf.

Wallich:1996:HIA

[2996] Paul Wallich. Having it all. Scientific American, 274(3):31, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-31.pdf.

Zorpette:1996:HP

[2997] Glenn Zorpette. Hot pork. Scientific American, 274(3):32–33, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-32.pdf.

Dupont:1996:SS

[2998] Daniel G. Dupont. Smart shopping. Scientific American, 274(3):34, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-34a.pdf.

Gibbs:1996:TS

[2999] W. Wayt Gibbs. Tough stuff. Scientific American, 274(3):34, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-34b.pdf.

Seife:1996:BT

[3000] Charles Seife. Bad timing. Scientific American, 274(3):36, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-36a.pdf.

Mukerjee:1996:SWS

[3001] Madhusree Mukerjee. Seeing the world in a snowflake. Scientific American, 274(3):36–42, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
36b.pdf.

Rabinovitch:1996:UPC

[3002] Jonas Rabinovitch and Josef Leitmann. Urban planning in Curitiba. Scientific American, 274(3):46–53, March 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-46.pdf;   http://www.sciam.com/
0396issue/0396toc.html.

Gehrels:1996:CCA

[3003] Tom Gehrels. Collisions with comets and asteroids. Scientific American, 274(3):54–58, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
54.pdf;   http://www.sciam.com/0396issue/0396toc.html.

Caldwell:1996:AAE

[3004] John C. Caldwell and Pat Caldwell. The African AIDS epidemic. Scientific American, 274(3):62–68, March 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-62.pdf;   http://www.sciam.com/
0396issue/0396toc.html.

Kivelson:1996:EFQ

[3005] S. Kivelson, Dung-Hai Lee, and Shou-Cheng Zhang. Electrons in flatland [quantum Hall effect]. Scientific American, 274(3):64–69, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rothman:1996:BVL

[3006] James E. Rothman and Lelio Orci. Budding vesicles in living cells. Scientific American, 274(3):70–75, March 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-70.pdf;   http://www.sciam.com/
0396issue/0396toc.html.

Brugioni:1996:SPA

[3007] Dino A. Brugioni. Science in pictures: The art and science of photoreconnaissance. Scientific American, 274(3):78–85, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-78.pdf.

Brugioni:1996:ASP

[3008] Dino A. Brugioni. The art and science of photoreconnaissance. Scientific American, 274(3):78–?? (Intl. ed. 56–??), March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0396issue/0396toc.html.

Anonymous:1996:ECB

[3009] Anonymous. Exploring chemical bonds. Scientific American, 274(3):82– ??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/0396issue/0396toc.html.

Anonymous:1996:PQQ

[3010] Anonymous. Playing with quads and quazars. Scientific American, 274(3):84–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0396issue/0396toc.html.

Kivelson:1996:EF

[3011] Steven Kivelson, Dung-Hai Lee, and Shou-Cheng Zhang. Electrons in Flatland. Scientific American, 274(3):86–91, March 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-86.pdf;   http://www.sciam.com/
0396issue/0396toc.html.

Rutzler:1996:CMS

[3012] Klaus Rützler and Ilka C. Feller. Caribbean mangrove swamps. Scientific American, 274(3):94–99, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
94.pdf;   http://www.sciam.com/0396issue/0396toc.html.

Beardsley:1996:THG

[3013] Tim Beardsley. Trends in human genetics: Vital data. Scientific American, 274(3):100–105, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
100.pdf;   http://www.sciam.com/0396issue/0396toc.html.

Carlson:1996:ASE

[3014] Shawn Carlson. The amateur scientist: Exploring chemical bonds. Scientific American, 274(3):106–107, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
106.pdf.

Anonymous:1996:ASMa

[3015] Anonymous. The amateur scientist: Measuring the strength of chemical bonds. Scientific American, 274(3):106–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0396issue/0396toc.html.

Stewart:1996:MRPb

[3016] Ian Stewart. Mathematical recreations: Playing with quads and quazars. Scientific American, 274(3):108–109, March 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-108.pdf.

Anonymous:1996:MRSb

[3017] Anonymous. Mathematical recreations: Squaring off in a board game of Quads. Scientific American, 274(3):108–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0396issue/0396toc.html.

Anonymous:1996:RCc

[3018] Anonymous. Reviews and commentaries. Scientific American, 274(3):110–114, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
110.pdf.

Anonymous:1996:RCSb

[3019] Anonymous. Reviews and commentaries: Star trek physics. surviving the future. Scientific American, 274(3):110–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0396issue/0396toc.html.

Burke:1996:CCC

[3020] James Burke. Connections: Code and commerce. Scientific American, 274(3):110–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0396issue/0396toc.html.

Morrison:1996:WEA

[3021] Philip Morrison and Phylis Morrison. Wonders: Earth’s asymmetries. Scientific American, 274(3):110–??, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/0396issue/0396toc.html.

Morrison:1996:SSB

[3022] Phylis Morrison and Philip Morrison. Symmetry and symmetry-breaking in South Africa. Scientific American, 274(3):114–116, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n3/pdf/scientificamerican0396-114.pdf.

Burke:1996:SC

[3023] James Burke. Satisfied customers. Scientific American, 274(3):116–117, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n3/pdf/scientificamerican0396-116.pdf.

Eisenberg:1996:EPD

[3024] Anne Eisenberg. Essay: Privacy and data collection on the net. Scientific American, 274(3):120, March 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n3/pdf/scientificamerican0396-
120.pdf.

Eisenberg:1996:EDM

[3025] Anne Eisenberg. Essay: Data mining and privacy invasion on the Net. Scientific American, 274(3):120–??, March 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0396issue/0396toc.html.

Rennie:1996:ECS

[3026] John Rennie. From the Editors: Changing to stay the same. Scientific American, 274(4):6, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
6.pdf.

Anonymous:1996:Ec

[3027] Anonymous. Errata. Scientific American, 274(4):10, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n4/pdf/scientificamerican0496-10b.pdf.

Anonymous:1996:LEd

[3028] Anonymous. Letters to the Editors. Scientific American, 274(4):10, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-10a.pdf.

Anonymous:1996:YAId

[3029] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 274(4):12, April 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n4/pdf/scientificamerican0496-12.pdf.

Maranto:1996:NAE

[3030] Gina Maranto. News and analysis: Embryo overpopulation. Scientific American, 274(4):16–18, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
16.pdf.

Anonymous:1996:FFE

[3031] Anonymous. In focus: Frozen embryos face an uncertain tomorrow. Scientific American, 274(4):16–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Yam:1996:AKN

[3032] Philip Yam. Attack of the killer neutrinos. Scientific American, 274(4): 20, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-20b.pdf.

Beardsley:1996:F

[3033] Tim Beardsley. Out of food? Scientific American, 274(4):20–22, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-20a.pdf.

Anonymous:1996:SCHb

[3034] Anonymous. Science and the citizen: Human ancestors outside Africa. Polly wants a student. killer neutrinos. antihydrogen. Scientific American, 274(4):20–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Leutwyler:1996:Ba

[3035] Kristin Leutwyler. In brief. Scientific American, 274(4):22–26, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-22a.pdf.

Mukerjee:1996:BP

[3036] Madhusree Mukerjee. The berry and the parasite. Scientific American, 274(4):22–26, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
22b.pdf.

Doyle:1996:LBW

[3037] Rodger Doyle. Low-birth-weight babies. Scientific American, 274(4):24– 25, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-24.pdf.

Yam:1996:SA

[3038] Philip Yam. A smattering of antimatter. Scientific American, 274(4):26– 28, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-26.pdf.

Mukerjee:1996:IP

[3039] Madhusree Mukerjee. Interview with a parrot. Scientific American, 274(4):28, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
28.pdf.

Horgan:1996:ET

[3040] John Horgan. The not so enormous theorem. Scientific American, 274(4): 29, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-29a.pdf.

Schneider:1996:RR

[3041] David Schneider. Rocking rocks. Scientific American, 274(4):29–32, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-29b.pdf.

Mukerjee:1996:EF

[3042] Madhusree Mukerjee. Engendering faces. Scientific American, 274(4): 32, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-32.pdf.

Browning:1996:PID

[3043] John Browning. The perils of an irregular deregulation. Scientific American, 274(4):33, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
33.pdf.

Anonymous:1996:CVCa

[3044] Anonymous. Cyber view: Censorship and the telecoms bill. Scientific American, 274(4):33–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Gregory:1996:WNN

[3045] Gregory. When novelty is not new. Scientific American, 274(4):34, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-34b.pdf.

Gibbs:1996:BEW

[3046] W. Wayt Gibbs. Battling the enemy within. Scientific American, 274(4):34–36, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
34a.pdf.

Anonymous:1996:TBF

[3047] Anonymous. Technology and business: Federal software inefficiencies; litigating the science of implants. Scientific American, 274(4):34–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0496issue/0496toc.html.

Nemecek:1996:AD

[3048] Sasha Nemecek. Augmenting discord. Scientific American, 274(4):36– 38, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-36.pdf.

Harby:1996:DE

[3049] Karla Harby. A discerning eye. Scientific American, 274(4):38, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-38.pdf.

Holloway:1996:ETE

[3050] Marguerite Holloway. Evolutionary theories for everyday life. Scientific American, 274(4):40–42, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
40.pdf.

Anonymous:1996:PBM

[3051] Anonymous. Profile: Biologist Margie Profet argues why sickness makes sense. Scientific American, 274(4):40–??, April 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0496issue/0496toc.html.

Shcherbak:1996:CNL

[3052] Yuri M. Shcherbak. Confronting the nuclear legacy — Part I: Ten years of the Chornobyl Era. Scientific American, 274(4):44–49, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n4/pdf/scientificamerican0496-44.pdf.

Shcherbak:1996:TYC

[3053] Yuri M. Shcherbak. Ten years of the Chornobyl era. Scientific American, 274(4):44–?? (Intl. ed. 32–??), April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

deDuve:1996:BCCa

[3054] Christian de Duve. The birth of complex cells. Scientific American, 274(4):50–57, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
50.pdf.

DeDave:1996:BCC

[3055] Christian De Dave. Birth of complex cells. Scientific American, 274(4): 50–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/0496issue/0496toc.
html.

deDuve:1996:BCCb

[3056] Christian de Duve. The birth of complex cells. Scientific American, 274(4):50–?? (Intl. ed. 38–??), April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Angel:1996:SLOa

[3057] J. Roger P. Angel and Neville J. Woolf. Searching for life on other planets. Scientific American, 274(4):60–66, April 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n4/pdf/scientificamerican0496-60.pdf.

Roger:1996:SLO

[3058] J. Roger, P. Angel, and Neville J. Woolf. Searching for life on other planets. Scientific American, 274(4):60–?? (Intl. ed. 46–??), April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0496issue/0496toc.html.

Pentland:1996:SR

[3059] Alex P. Pentland. Smart rooms. Scientific American, 274(4):68–76, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-68.pdf;   http:/
/www.sciam.com/0496issue/0496toc.html.

Musto:1996:AAH

[3060] David F. Musto. Alcohol in American history. Scientific American, 274(4):78–83, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
78.pdf;   http://www.sciam.com/0496issue/0496toc.html.

Stix:1996:TNWa

[3061] Gary Stix. Trends in nanotechnology: Waiting for breakthroughs. Scientific American, 274(4):78–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Grimaldi:1996:SPC

[3062] David A. Grimaldi. Science in pictures: Captured in amber. Scientific American, 274(4):84–91, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
84.pdf.

Anonymous:1996:NBS

[3063] Anonymous. The new backyard seismology. Scientific American, 274(4): 84–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/0496issue/0496toc.
html.

Grimaldi:1996:CA

[3064] David A. Grimaldi. Captured in amber. Scientific American, 274 (4):84–?? (Intl. ed. 70–??), April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Anonymous:1996:HFM

[3065] Anonymous. How fair is Monopoly? Scientific American, 274(4):86– ??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/0496issue/0496toc.html.

Stix:1996:TNWb

[3066] Gary Stix. Trends in nanotechnology: Waiting for breakthroughs. Scientific American, 274(4):94–99, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
94.pdf.

Stix:1996:WB

[3067] Gary Stix. Waiting for breakthroughs. Scientific American, 274(4):94– ??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/0496issue/0496toc.html.

Carlson:1996:ASN

[3068] Shawn Carlson. The amateur scientist: The new backyard seismology. Scientific American, 274(4):100–102, April 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n4/pdf/scientificamerican0496-100.pdf.

Anonymous:1996:MEY

[3069] Anonymous. Monitoring earthquakes in your backyard. Scientific American, 274(4):100–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Stewart:1996:MRH

[3070] Ian Stewart. Mathematical recreations: How fair is Monopoly? Scientific American, 274(4):104–105, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
104.pdf.

Anonymous:1996:PSW

[3071] Anonymous. Probability shows why all’s fair in monopoly. Scientific American, 274(4):104–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Anonymous:1996:RCD

[3072] Anonymous. Reviews and Commentaries: Do dinosaur books savage paleontology?. numskull numbers. monkeying with science. Scientific American, 274(4):108–111, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
108.pdf;   http://www.sciam.com/0496issue/0496toc.html.

Burke:1996:CHCb

[3073] James Burke. Connections: From hot coffee to evolution. Scientific American, 274(4):108–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Morrison:1996:WAE

[3074] Philip Morrison and Phylis Morrison. Wonders: Apartheid’s electronic legacy. Scientific American, 274(4):108–??, April 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0496issue/0496toc.html.

Morrison:1996:SC

[3075] Phylis Morrison and Philip Morrison. Symmetries and civilizations. Scientific American, 274(4):112–114, April 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n4/pdf/scientificamerican0496-112.pdf.

Burke:1996:WN

[3076] James Burke. What’s in a name? Scientific American, 274(4):113–115, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n4/pdf/scientificamerican0496-113.pdf.

Jacobs:1996:WKU

[3077] Martin I. Jacobs. Working knowledge: Unzipping Velcro. Scientific American, 274(4):116, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n4/pdf/scientificamerican0496-
116.pdf.

Anonymous:1996:WPZ

[3078] Anonymous. What puts the zip in this fastener. Scientific American, 274(4):116–??, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0496issue/0496toc.html.

Rennie:1996:E

[3079] John Rennie. From the Editors. Scientific American, 274(5):4, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-4.pdf.

Hanson:1996:LE

[3080] Harold P. Hanson. Letters to the Editors. Scientific American, 274(5): 6, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-6.pdf.

Anonymous:1996:YAIe

[3081] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 274(5):8, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
8.pdf.

Anonymous:1996:FPS

[3082] Anonymous. In focus: Physicians still do not honor living wills. Scientific American, 274(5):8–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0596issue/0596currentissue.html.

Anonymous:1996:SCM

[3083] Anonymous. Science and the citizen: Medical trials in question. the future chess champion. biodiversity and productivity. what pigs think. Scientific American, 274(5):10–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horgan:1996:NAR

[3084] John Horgan. News and analysis: Right to die. Scientific American, 274(5):12–14, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
12.pdf.

Horgan:1996:PNM

[3085] John Horgan. Plotting the next move. Scientific American, 274(5): 16, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-16b.pdf.

Leutwyler:1996:XMS

[3086] Kristin Leutwyler. X marks the spots. Scientific American, 274(5):16– 18, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-16a.pdf.

Leutwyler:1996:Bb

[3087] Kristin Leutwyler. In brief. Scientific American, 274(5):18–22, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-18.pdf.

Anonymous:1996:CVB

[3088] Anonymous. Cyber view: Broadcasting on a narrow medium. Scientific American, 274(5):18–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1996:TBT

[3089] Anonymous. Technology and business: A tailless airplane. fake muscles, real bones. wandering genes. Scientific American, 274(5):19–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Seife:1996:RE

[3090] Charles Seife. Relatively expensive. Scientific American, 274(5):20, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-20b.pdf.

Wallich:1996:BAA

[3091] Paul Wallich. Not so blind, after all. Scientific American, 274(5):20– 22, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-20a.pdf.

Mirsky:1996:PBS

[3092] Steve Mirsky. Pork barrel science. Scientific American, 274(5):22, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-22.pdf.

Mukerjee:1996:SWY

[3093] Madhusree Mukerjee. Sowing where you reap. Scientific American, 274(5):23–25, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
23.pdf.

Doyle:1996:CDE

[3094] Rodger Doyle. Carbon dioxide emissions. Scientific American, 274(5): 24, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-24.pdf.

Anonymous:1996:PDN

[3095] Anonymous. Profile: Distinguished naturalist Miriam Rothschild defies categorization. Scientific American, 274(5):24–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1996:MSM

[3096] David Schneider. The more species, the merrier. Scientific American, 274 (5):26, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-26b.pdf.

Zorpette:1996:ESA

[3097] Glenn Zorpette. Electric smile-aid. Scientific American, 274(5):26, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-26a.pdf.

Strada:1996:HLMa

[3098] Gino Strada. The horror of land mines. Scientific American, 274(5):26– ??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Browning:1996:TAI

[3099] John Browning. Television arrives on the Internet. Scientific American, 274(5):28, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
28.pdf.

Gibbs:1996:SE

[3100] W. Wayt Gibbs. Systematic errors. Scientific American, 274(5):30– 31, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-30.pdf.

Eisenberg:1996:RNa

[3101] Anne Eisenberg. Recently netted. … Scientific American, 274(5):32, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-32a.pdf.

Gibbs:1996:PD

[3102] W. Wayt Gibbs. On permanent displays. Scientific American, 274(5): 32, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-32b.pdf.

Luu:1996:KBa

[3103] Jane X. Luu and David C. Jewitt. The Kuiper belt. Scientific American, 274(5):32–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1996:AN

[3104] Tim Beardsley. Advantage: Nature. Scientific American, 274(5):33, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-33a.pdf.

Gibbs:1996:MM

[3105] W. Wayt Gibbs. Make a muscle. Scientific American, 274(5):33, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-33b.pdf.

Yam:1996:RR

[3106] Philip Yam. Remote repair. Scientific American, 274(5):34, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-34a.pdf.

Zorpette:1996:WI

[3107] Glenn Zorpette. Winging it. Scientific American, 274(5):34, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-34b.pdf.

Holloway:1996:NHF

[3108] Marguerite Holloway. A natural history of fleas and butterflies. Scientific American, 274(5):36–38, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
36.pdf.

Strada:1996:HLMb

[3109] Gino Strada. The horror of land mines. Scientific American, 274(5):40– 45, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-40.pdf.

Perera:1996:UNCa

[3110] Frederica P. Perera. Uncovering new clues to cancer risk. Scientific American, 274(5):40–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Luu:1996:KBb

[3111] Jane X. Luu and David C. Jewitt. The Kuiper Belt. Scientific American, 274(5):46–52, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
46.pdf.

Perera:1996:UNCb

[3112] Frederica P. Perera. Uncovering new clues to cancer risk. Scientific American, 274(5):54–62, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
54.pdf.

Myers:1996:PHa

[3113] David G. Myers and Ed Diener. The pursuit of happiness. Scientific American, 274(5):54–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beland:1996:BWSa

[3114] Pierre Beland. The beluga whales of the St. Lawrence river. Scientific American, 274(5):58–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Birman:1996:SRN

[3115] Kenneth P. Birman and Robbert van Renesse. Software for reliable networks. Scientific American, 274(5):64–69, May 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-64.pdf.

Hale:1996:LTAa

[3116] John R. Hale. The lost technology of Ancient Greek rowing. Scientific American, 274(5):66–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Myers:1996:PHb

[3117] David G. Myers and Ed Diener. The pursuit of happiness. Scientific American, 274(5):70–72, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
70.pdf.

Zorpette:1996:HNW

[3118] Glenn Zorpette. Hanford’s nuclear wasteland. Scientific American, 274 (5):72–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beland:1996:BWSb

[3119] Pierre Béland. The beluga whales of the St. Lawrence River. Scientific American, 274(5):74–81, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
74.pdf.

Hale:1996:LTAb

[3120] John R. Hale. The lost technology of Ancient Greek rowing. Scientific American, 274(5):82–85, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
82.pdf.

Anonymous:1996:ASD

[3121] Anonymous. The amateur scientist: Detecting low-frequency electromagnetic waves. Scientific American, 274(5):82–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1996:MRF

[3122] Anonymous. MATHEMATICAL RECREATIONS: Fractal sculpture turns cubes into flowing spirals. Scientific American, 274(5):86–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1996:SAS

[3123] Anonymous. The sculptures of Alan St George. Scientific American, 274(5):86–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1996:CNL

[3124] Glenn Zorpette. Confronting the nuclear legacy — Part II: Hanford’s nuclear wasteland. Scientific American, 274(5):88–97, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-88.pdf.

Anonymous:1996:RCF

[3125] Anonymous. Reviews and Commentaries: Four books make complexity less confusing. the bomb on CD-ROM. endangered flora. Darwin goes to the movies. Scientific American, 274(5):88–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1996:CPW

[3126] James Burke. Connections: From phonetic writing to stained brains. Scientific American, 274(5):88–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1996:WFI

[3127] Philip Morrison. Wonders: Finding invisible planets. Scientific American, 274(5):88–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1996:WKWa

[3128] Anonymous. Working knowledge: Why elevators are safe. Scientific American, 274(5):96–??, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carlson:1996:ASDa

[3129] Shawn Carlson. The amateur scientist: Detecting natural electromagnetic waves. Scientific American, 274(5):98–101, May 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-98.pdf.

Stewart:1996:MRSa

[3130] Ian Stewart. Mathematical recreations: The sculptures of Alan St. George. Scientific American, 274(5):102–103, May 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n5/pdf/scientificamerican0596-102.pdf.

Anonymous:1996:RCe

[3131] Anonymous. Reviews and commentaries. Scientific American, 274(5):104–108, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
104.pdf.

Morrison:1996:PT

[3132] Philip Morrison. Planet-tude. Scientific American, 274(5):109–111, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-109.pdf.

Burke:1996:HS

[3133] James Burke. Highbrow stuff. Scientific American, 274(5):110–111, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n5/pdf/scientificamerican0596-110.pdf.

Sheeran:1996:WKV

[3134] William Sheeran. Working knowledge: Vertical safety. Scientific American, 274(5):112, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n5/pdf/scientificamerican0596-
112.pdf.

Rennie:1996:ESS

[3135] John Rennie. From the Editor: The Space Station’s disappointing odyssey. Scientific American, 274(6):6, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-6.pdf.

Anonymous:1996:Ed

[3136] Anonymous. Errata. Scientific American, 274(6):8, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-8b.pdf.

Anonymous:1996:LEe

[3137] Anonymous. Letters to the Editors. Scientific American, 274(6):8, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-8a.pdf.

Anonymous:1996:FCP

[3138] Anonymous. In focus: China’s plans for “peaceful” atomic tests inspire unease. Scientific American, 274(6):9–??, June 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0696issue/0696currentissue.html.

Anonymous:1996:YAIf

[3139] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 274(6):10, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-10.pdf.

Anonymous:1996:SCDb

[3140] Anonymous. Science and the citizen: Downtown on the farm. life on the Unabomber’s list. Galileo at Jupiter. lizard kings. forecasting Alzheimer’s. Scientific American, 274(6):12–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0696issue/0696currentissue.html.

Horgan:1996:NAP

[3141] John Horgan. News and analysis: “peaceful” nuclear explosions. Scientific American, 274(6):14–16, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
14.pdf.

Gibbs:1996:SHO

[3142] W. Wayt Gibbs. Star-hopping by the outhouse. Scientific American, 274(6):18, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
18b.pdf.

Stix:1996:U

[3143] Gary Stix. Urbaculture. Scientific American, 274(6):18–20, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-18a.pdf.

Anonymous:1996:Ba

[3144] Anonymous. In brief. Scientific American, 274(6):20–24, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-20.pdf.

Anonymous:1996:CVK

[3145] Anonymous. Cyber view: Keeping databases under lock and key. Scientific American, 274(6):21–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Zorpette:1996:FDW

[3146] Glenn Zorpette. Forecast: Dry and windy. Scientific American, 274(6): 22, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-22.pdf.

Anonymous:1996:TBMa

[3147] Anonymous. TECHNOLOGY AND BUSINES’: Mind reading for movement. cruising by balloon. super-ultrasound. Scientific American, 274(6):22–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Howard:1996:AD

[3148] Ken Howard. Artful dating. Scientific American, 274(6):24, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-24.pdf.

Eisenberg:1996:APK

[3149] Anne Eisenberg. Among the papers in Kaczynski’s cabin. Scientific American, 274(6):25, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
25.pdf.

Mirsky:1996:LK

[3150] Steve Mirsky. The lizard kings. Scientific American, 274(6):26, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-26a.pdf.

Wallich:1996:SW

[3151] Paul Wallich. Senile words. Scientific American, 274(6):26–28, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-26b.pdf.

Doyle:1996:ACR

[3152] Rodger Doyle. AIDS cases reported, 1994–1995. Scientific American, 274(6):28, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
28.pdf.

Browning:1996:PFL

[3153] John Browning. Playing facts and loose. Scientific American, 274(6):30– 32, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-30.pdf.

Anonymous:1996:PDCb

[3154] Anonymous. Profile: Demographic criminologist James Alan Fox braces for a crime wave. Scientific American, 274(6):30–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0696issue/0696currentissue.html.

Gibbs:1996:UNP

[3155] W. Wayt Gibbs. Ultrasound’s new phase. Scientific American, 274(6):32– 34, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-32.pdf.

Randazzese:1996:SSa

[3156] Lucien P. Randazzese. Semiconductor subsidies. Scientific American, 274(6):32–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Gibbs:1996:MR

[3157] W. Wayt Gibbs. Mind readings. Scientific American, 274(6):34–36, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-34.pdf.

Schneider:1996:GW

[3158] David Schneider. Good wood. Scientific American, 274(6):36–38, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-36.pdf.

Beardsley:1996:TSSa

[3159] Tim Beardsley. Trends in space science: Science in the sky. Scientific American, 274(6):36–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Beardsley:1996:TT

[3160] Tim Beardsley. Testing, testing. Scientific American, 274(6):38–39, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-38.pdf.

Dupont:1996:X

[3161] Daniel G. Dupont. Up, up and away. Scientific American, 274(6):39, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-39.pdf.

Yam:1996:CCC

[3162] Philip Yam. Catching a coming crime wave. Scientific American, 274(6):40–44, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
40.pdf.

Kearney:1996:TOAa

[3163] Jay T. Kearney. Training the Olympic athlete. Scientific American, 274(6):44–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Randazzese:1996:SSb

[3164] Lucien P. Randazzese. Semiconductor subsidies. Scientific American, 274(6):46–49, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
46.pdf.

Kearney:1996:TOAb

[3165] Jay T. Kearney. Training the Olympic athlete. Scientific American, 274(6):50–63, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
50.pdf.

Whipple:1996:CNW

[3166] Chris G. Whipple. Can nuclear waste be stored safely at Yucca Mountain? Scientific American, 274(6):56–64, June 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0696issue/0696currentissue.html.

Beardsley:1996:TSSb

[3167] Tim Beardsley. Trends in space science: Science in the sky. Scientific American, 274(6):64–70, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
64.pdf.

Whipple:1996:CNL

[3168] Chris G. Whipple. Confronting the nuclear legacy: Can nuclear waste be stored safely at Yucca Mountain? Scientific American, 274(6):72–79, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-72.pdf.

Paul:1996:ACR

[3169] Gregory S. Paul. The art of Charles R. Knight. Scientific American, 274(6):74–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Bernstein:1996:RFB

[3170] Jeremy Bernstein. The reluctant father of black holes [Einstein]. Scientific American, 274(6):80–85, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL https://www.math.

utah.edu/pub/tex/bib/einstein.bib;   http://www.nature.com/
scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
80.pdf;   http://www.sciam.com/0696issue/0696currentissue.html.

Nicolaou:1996:TNWa

[3171] K. C. Nicolaou, Rodney K. Guy, and Pierre Potier. Taxoids: New weapons against cancer. Scientific American, 274(6):84–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0696issue/0696currentissue.html.

Paul:1996:SPA

[3172] Gregory S. Paul. Science in pictures: The art of Charles R. Knight. Scientific American, 274(6):86–93, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-86.pdf.

Anonymous:1996:ASP

[3173] Anonymous. The amateur scientist: A professional-quality balance without a heavy price tag. Scientific American, 274(6):90–??, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0696issue/0696currentissue.html.

Anonymous:1996:HME

[3174] Anonymous. Homemade microgram electrobalances. Scientific American, 274(6):90–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Anonymous:1996:MRO

[3175] Anonymous. MATHEMATICAL RECREATIONS: An overlooked numerical series spins out things of beauty. Scientific American, 274 (6):92–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0696issue/
0696currentissue.html.

Anonymous:1996:TNN

[3176] Anonymous. Tales of a neglected number. Scientific American, 274 (6):92–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0696issue/
0696currentissue.html.

Nicolaou:1996:TNWb

[3177] K. C. Nicolaou, Rodney K. Guy, and Pierre Potier. Taxoids: New weapons against cancer. Scientific American, 274(6):94–98, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-94.pdf.

Anonymous:1996:RCAa

[3178] Anonymous. Reviews and Commentaries: Amazonian riches. the haunting lure of pseudoscience. on-line onstage. chemistry’s yin and yang. Scientific American, 274(6):94–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Burke:1996:CGF

[3179] James Burke. Connections: From galvanized Frankenstein to atomized gasoline. Scientific American, 274(6):94–??, June 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0696issue/0696currentissue.html.

Morrison:1996:WIS

[3180] Philip Morrison. Wonders: The intense science of creating high pressures. Scientific American, 274(6):94–??, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Carlson:1996:ASH

[3181] Shawn Carlson. The amateur scientist: Homemade microgram electrobalances. Scientific American, 274(6):100–101, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-100.pdf.

Stewart:1996:MRT

[3182] Ian Stewart. Mathematical recreations: Tales of a neglected number. Scientific American, 274(6):102–103, June 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v274/
n6/pdf/scientificamerican0696-102.pdf.

Anonymous:1996:RCg

[3183] Anonymous. Reviews and commentaries. Scientific American, 274(6):106–112, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
106.pdf.

Morrison:1996:UP

[3184] Philip Morrison. Under pressure. Scientific American, 274(6):113–115, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-113.pdf.

Burke:1996:G

[3185] James Burke. Out of gas. Scientific American, 274(6):114–115, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v274/n6/pdf/scientificamerican0696-114.pdf.

Anonymous:1996:WKF

[3186] Anonymous. WORKING KNOWLEDGE: The fleet inflation of air bags. Scientific American, 274(6):116, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0696issue/0696currentissue.html.

Resh:1996:WKA

[3187] Robert E. Resh. Working knowledge: Air bags. Scientific American, 274(6):116, June 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v274/n6/pdf/scientificamerican0696-
116.pdf.

Rennie:1996:EGF

[3188] John Rennie. From the Editor: Glimpses of the familiar but unknown. Scientific American, 275(1):6, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
6.pdf.

Anonymous:1996:LEg

[3189] Anonymous. Letters to the Editors. Scientific American, 275(1):8, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-8.pdf.

Anonymous:1996:YAIg

[3190] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 275(1):10, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
10.pdf.

Anonymous:1996:FRS

[3191] Anonymous. IN FOCUS: It’s rise-and-shine time for sleep research. Scientific American, 275(1):10–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Beardsley:1996:NAW

[3192] Tim Beardsley. News and analysis: Waking up. Scientific American, 275(1):14–18, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
14.pdf.

Anonymous:1996:SCS

[3193] Anonymous. Science and the citizen: Supersymmetry. linear A not just Greek to archaeologists. measuring the quality of life. arguments over aquaculture. Scientific American, 275(1):14–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0796issue/0796currentissue.html.

Schneider:1996:PL

[3194] David Schneider. Pot luck. Scientific American, 275(1):20, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-20.pdf.

Anonymous:1996:CVCb

[3195] Anonymous. Cyber view: Couch potatoes pull up their roots. Scientific American, 275(1):20–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Yam:1996:MM

[3196] Philip Yam. Mirror, mirror. Scientific American, 275(1):21, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-21a.pdf.

Anonymous:1996:Bb

[3197] Anonymous. In brief. Scientific American, 275(1):21–26, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-21b.pdf.

Anonymous:1996:TBMb

[3198] Anonymous. Technology and business: Micro-optics on a microchip. A telescope better than Hubble. shining leaves. going critical. Scientific American, 275(1):21–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Mukerjee:1996:PG

[3199] Madhusree Mukerjee. Pink gold. Scientific American, 275(1):24–26, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-24.pdf.

Horgan:1996:HC

[3200] John Horgan. Headshrinker convention. Scientific American, 275(1): 26, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-26.pdf.

Anonymous:1996:PCP

[3201] Anonymous. Profile: Contraceptives pioneer G. P. Talwar takes on nature and critics. Scientific American, 275(1):26–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0796issue/0796currentissue.html.

Doyle:1996:CQL

[3202] Rodger Doyle. The changing quality of life. Scientific American, 275(1): 28, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-28.pdf.

Mirsky:1996:WT

[3203] Steve Mirsky. Wonderful town. Scientific American, 275(1):29, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-29b.pdf.

Horgan:1996:GT

[3204] John Horgan. Group think. Scientific American, 275(1):29–30, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-29a.pdf.

Schneider:1996:TB

[3205] David Schneider. Triassic bug. Scientific American, 275(1):30, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-30.pdf.

Browning:1996:NSN

[3206] John Browning. New stars for the new media. Scientific American, 275 (1):31, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-31.pdf.

Stix:1996:LW

[3207] Gary Stix. Light work. Scientific American, 275(1):32–34, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-32.pdf.

Beardsley:1996:TV

[3208] Tim Beardsley. Treble vision. Scientific American, 275(1):34, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-34.pdf.

Wallich:1996:PL

[3209] Paul Wallich. Panacea lost? Scientific American, 275(1):35–36, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-35.pdf.

Zorpette:1996:ULB

[3210] Glenn Zorpette. The underwater lightness of being. Scientific American, 275(1):36, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
36b.pdf.

Zorpette:1996:KTB

[3211] Glenn Zorpette. Keeping the “tiger” at bay. Scientific American, 275(1): 36–37, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-36a.pdf.

Nemecek:1996:LIT

[3212] Sasha Nemecek. Leaf it to them. Scientific American, 275(1):37, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-37.pdf.

Mukerjee:1996:PEV

[3213] Madhusree Mukerjee. Pushing the envelope for vaccines. Scientific American, 275(1):38–40, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
38.pdf.

Leffell:1996:SSCa

[3214] David J. Leffell and Douglas E. Brash. Sunlight and skin cancer. Scientific American, 275(1):38–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Bell:1996:FCD

[3215] Alan E. Bell. The future of compact disk technology: Next-generation compact discs. Scientific American, 275(1):42–46, July 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-42.pdf.

Bell:1996:NGCa

[3216] Alan E. Bell. Next generation compact discs. Scientific American, 275(1):42–46 (Intl. ed. 28–??), July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Hawking:1996:NSTa

[3217] Stephen W. Hawking and Roger Penrose. The nature of space and time. Scientific American, 275(1):44–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Gunshor:1996:BLC

[3218] Robert L. Gunshor and Arto V. Nurmikko. Blue-laser CD technology. Scientific American, 275(1):48–51, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
48.pdf;   http://www.sciam.com/0796issue/0796currentissue.html.

AguileraHellweg:1996:HWS

[3219] Max Aguilera Hellweg. The hidden world of surgery. Scientific American, 275(1):50–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Leffell:1996:SSCb

[3220] David J. Leffell and Douglas E. Brash. Sunlight and skin cancer. Scientific American, 275(1):52–59, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
52.pdf.

Erwin:1996:MMEa

[3221] Douglas H. Erwin. The mother of mass extinctions. Scientific American, 275(1):56–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Hawking:1996:NSTb

[3222] Stephen W. Hawking and Roger Penrose. The nature of space and time. Scientific American, 275(1):60–65, July 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-60.pdf;   http://www.sciam.com/
0796issue/0796currentissue.html.

Okerson:1996:WODa

[3223] Ann Okerson. Who owns digital works? Scientific American, 275 (1):64–68, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0796issue/
0796currentissue.html.

Aguilera-Hellweg:1996:SPH

[3224] Max Aguilera-Hellweg. Science in pictures: The hidden world of surgery. Scientific American, 275(1):66–71, July 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-66.pdf.

Zill:1996:ESWa

[3225] Sasha N. Zill and Ernst-August Seyfarth. Exoskeletal sensors for walking. Scientific American, 275(1):70–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Erwin:1996:MMEb

[3226] Douglas H. Erwin. The mother of Mass Extinctions. Scientific American, 275(1):72–78, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
72.pdf.

Carlson:1996:ASCa

[3227] Shawn Carlson. The amateur scientist: Covert observations of nesting sparrows. Scientific American, 275(1):76–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0796issue/0796currentissue.html.

Anonymous:1996:MRK

[3228] Anonymous. Mathematical recreations: Knotty arithmetic unravels shoelaces. Scientific American, 275(1):78–??, July 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0796issue/0796currentissue.html.

Okerson:1996:WODb

[3229] Ann Okerson. Who owns digital works? Scientific American, 275(1):80– 84, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-80.pdf.

Anonymous:1996:RCO

[3230] Anonymous. Reviews and Commentaries: An odd but fitting museum of humanity. the life of Linus Pauling. double takes on digital photography. Scientific American, 275(1):82–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Burke:1996:CPD

[3231] James Burke. Connections: From pneumatic dredgers to grand opera. Scientific American, 275(1):82–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Morrison:1996:WSL

[3232] Philip Morrison. Wonders: Sunless life on the seafloor. Scientific American, 275(1):82–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Zill:1996:ESWb

[3233] Sasha N. Zill and Ernst-August Seyfarth. Exoskeletal sensors for walking. Scientific American, 275(1):86–90, July 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-86.pdf.

Carlson:1996:ASCb

[3234] Shawn Carlson. The amateur scientist: Covert observations of nesting sparrows. Scientific American, 275(1):92–93, July 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n1/pdf/scientificamerican0796-92.pdf.

Anonymous:1996:WKI

[3235] Anonymous. Working knowledge: Illuminating halogen lights. Scientific American, 275(1):92–??, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Stewart:1996:MRA

[3236] Ian Stewart. Mathematical recreations: Arithmetic and old lace. Scientific American, 275(1):94–97, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n1/pdf/scientificamerican0796-
94.pdf.

Anonymous:1996:RCh

[3237] Anonymous. Reviews and commentaries. Scientific American, 275(1):98– 103, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-98.pdf.

Morrison:1996:G

[3238] Philip Morrison. Gutless. Scientific American, 275(1):104–107, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-104.pdf.

Burke:1996:FG

[3239] James Burke. Folies de grandeur. Scientific American, 275(1):106– 107, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-106.pdf.

McGowan:1996:WK

[3240] Terry McGowan. Working knowledge. Scientific American, 275(1):108, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n1/pdf/scientificamerican0796-108.pdf.

Rennie:1996:EAP

[3241] John Rennie. From the Editors: Aliens at play. Scientific American, 275(2):6, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
6.pdf.

Anonymous:1996:LEh

[3242] Anonymous. Letters to the Editors. Scientific American, 275(2):8, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-8.pdf.

Anonymous:1996:YAIh

[3243] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 275(2):10, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-10.pdf.

Leutwyler:1996:NAP

[3244] Kristin Leutwyler. News and analysis: Paying attention. Scientific American, 275(2):12–14, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
12.pdf.

Anonymous:1996:ADD

[3245] Anonymous. Attention-deficit disorder comes under close scrutiny. Scientific American, 275(2):12–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

Stix:1996:PE

[3246] Gary Stix. Public eye. Scientific American, 275(2):18–20, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-18.pdf.

Anonymous:1996:SCW

[3247] Anonymous. Science and the citizen: Watchdogging governments from orbit. Zebra mussels: the good, the bad and the ugly. Space weather. Fertility transplant. Scientific American, 275(2):18–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0896issue/0896currentissue.html.

Wallich:1996:WF

[3248] Paul Wallich. The wall falls. Scientific American, 275(2):20–22, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-20b.pdf.

Leutwyler:1996:Bc

[3249] Kristin Leutwyler. In brief. Scientific American, 275(2):20–24, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-20a.pdf.

Zorpette:1996:MMC

[3250] Glenn Zorpette. Mussel mayhem, continued. Scientific American, 275(2):22–23, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
22.pdf.

Sinha:1996:NKH

[3251] Gunjan Sinha. A new king and his tiny minion. Scientific American, 275(2):24, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
24b.pdf.

Powell:1996:BV

[3252] Corey S. Powell. The blustery void. Scientific American, 275(2):24–25, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-24a.pdf.

Mirsky:1996:PST

[3253] Steve Mirsky. Patient, smell thyself. Scientific American, 275(2):25, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-25.pdf.

Doyle:1996:LA

[3254] Rodger Doyle. Lethal accidents, 1979–1992. Scientific American, 275(2): 26, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-26a.pdf.

Beardsley:1996:MFR

[3255] Tim Beardsley. Mouse to father rat? Scientific American, 275(2):26–28, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-26b.pdf.

Schneider:1996:IGX

[3256] David Schneider. Insects of Generation X. Scientific American, 275(2): 28, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-28.pdf.

Browning:1996:NC

[3257] John Browning. The notworking computer. Scientific American, 275(2):30–31, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
30.pdf.

Anonymous:1996:CVNa

[3258] Anonymous. Cyber view: Networking the wrong way. Scientific American, 275(2):30–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

Eisenberg:1996:RNb

[3259] Anne Eisenberg. Recently netted. … Scientific American, 275(2):31, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-31.pdf.

Yam:1996:MB

[3260] Philip Yam. Magnet on the brain. Scientific American, 275(2):32, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-32.pdf.

Anonymous:1996:TBSb

[3261] Anonymous. Technology and business: Steering drugs magnetically. gene patent overload. digital fingerprinting. Scientific American, 275(2):32–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

Aharonian:1996:TMT

[3262] Gregory Aharonian. Too much for too little. Scientific American, 275(2): 33, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-33a.pdf.

Sinha:1996:RT

[3263] Gunjan Sinha. The right touch. Scientific American, 275(2):33, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-33b.pdf.

Gibbs:1996:OSS

[3264] W. Wayt Gibbs. One small step. Scientific American, 275(2):33–35, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-33c.pdf.

Nemecek:1996:CMC

[3265] Sasha Nemecek. Coat of many colors. Scientific American, 275(2):34–35, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-34.pdf.

Beardsley:1996:EBS

[3266] Tim Beardsley. Exploring the business of space. Scientific American, 275(2):36–38, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
36.pdf.

Anonymous:1996:PSE

[3267] Anonymous. Profile: Space entrepreneur Shelley A. Harrison sees commerce taking flight. Scientific American, 275(2):36–??, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0896issue/0896currentissue.html.

NussleinVolhard:1996:GOE

[3268] Christiane Nüsslein Volhard. Gradients that organize embryo development. Scientific American, 275(2):38–??, August 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0896issue/0896currentissue.html.

Fancher:1996:SC

[3269] Carol H. Fancher. Smart cards. Scientific American, 275(2):40–45, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-40.pdf;   http:/
/www.sciam.com/0896issue/0896currentissue.html.

Nesme-Ribes:1996:SD

[3270] Elizabeth Nesme-Ribes, Sallie L. Baliunas, and Dmitry Sokoloff. The stellar dynamo. Scientific American, 275(2):46–52, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-46.pdf;   http:/
/www.sciam.com/0896issue/0896currentissue.html.

Kirtley:1996:PHTa

[3271] J. R. Kirtley and C. C. Tsuei. Probing high-temperature superconductivity. Scientific American, 275(2):50–55, August 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0896issue/0896currentissue.html.

Nusslein-Volhard:1996:GOE

[3272] Christiane Nüsslein-Volhard. Gradients that organize embryo development. Scientific American, 275(2):54–61, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-54.pdf;   http://www.sciam.com/
0896issue/0896currentissue.html.

Mack:1996:SPS

[3273] Walter N. Mack and Elizabeth A. Leistikow. Science in pictures: Sands of the world. Scientific American, 275(2):62–67, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-62.pdf.

Mack:1996:SW

[3274] Walter N. Mack and Elizabeth A. Leistikow. Sands of the world. Scientific American, 275(2):62–?? (Intl. ed. 44–??), August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0896issue/0896currentissue.html.

Kirtley:1996:PHTb

[3275] John R. Kirtley and Chang C. Tsuei. Probing high temperature superconductivity. Scientific American, 275(2):68–73, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-68.pdf;   http:/
/www.sciam.com/0896issue/0896currentissue.html.

Gibbs:1996:TMGa

[3276] W. Wayt Gibbs. Trends in medicine: Gaining on fat. Scientific American, 275(2):70–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

DeKeersmaecker:1996:MLB

[3277] Jacques De Keersmaecker. The mystery of Lambic beer. Scientific American, 275(2):74–80, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
74.pdf;   http://www.sciam.com/0896issue/0896currentissue.html.

Anonymous:1996:DMS

[3278] Anonymous. Detecting micron-size movements. Scientific American, 275(2):78–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

Anonymous:1996:SLD

[3279] Anonymous. Shedding a little darkness. Scientific American, 275(2): 81–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0896issue/
0896currentissue.html.

Marten:1996:RBD

[3280] Ken Marten, Karim Shariff, Suchi Psarakos, and Don J. White. Ring bubbles of dolphins. Scientific American, 275(2):82–87, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-82.pdf;   http:/
/www.sciam.com/0896issue/0896currentissue.html.

Gibbs:1996:TMGb

[3281] W. Wayt Gibbs. Trends in medicine: Gaining on fat. Scientific American, 275(2):88–94, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
88.pdf.

Gibbs:1996:GFO

[3282] W. Wayt Gibbs. Gaining on fat: Obesity plagues the industrial world. don’t blame sloth or gluttony. Scientific American, 275(2):88–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0896issue/0896currentissue.
html.

Carlson:1996:ASDb

[3283] Shawn Carlson. The amateur scientist: Detecting micron-size movements. Scientific American, 275(2):96–99, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-96.pdf.

Anonymous:1996:ASMb

[3284] Anonymous. The amateur scientist: A micromotion detector counts insect heartbeats. Scientific American, 275(2):96–??, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0896issue/0896currentissue.html.

Stewart:1996:MRSb

[3285] Ian Stewart. Mathematical recreations: Shedding a little darkness. Scientific American, 275(2):100–103, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-100.pdf.

Anonymous:1996:MRSc

[3286] Anonymous. Mathematical recreations: Searching for shadows in rooms made of mirrors. Scientific American, 275(2):100–??, August 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0896issue/0896currentissue.html.

Davis:1996:RCC

[3287] Ben Davis. Reviews and commentaries: The culture machine: Science and art on the Web. Scientific American, 275(2):104–108, August 1996.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n2/pdf/scientificamerican0896-104.pdf.

Anonymous:1996:RCCa

[3288] Anonymous. Reviews and Commentaries: Can art thrive on the Net?. colors of the reef. digital storm chasing. Feynman found. Scientific American, 275(2):104–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

Burke:1996:CSE

[3289] James Burke. Connections: From the shape of the Earth to the Scottish rebellion. Scientific American, 275(2):104–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0896issue/0896currentissue.html.

Morrison:1996:WUA

[3290] Philip Morrison and Phylis Morrison. Wonders: Urban agriculture teaches old lessons of the soil. Scientific American, 275(2):104–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0896issue/0896currentissue.
html.

Morrison:1996:GBC

[3291] Phylis Morrison and Philip Morrison. Great books and city crops. Scientific American, 275(2):109–111, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n2/pdf/scientificamerican0896-
109.pdf.

Burke:1996:LB

[3292] James Burke. A lot of boloney. Scientific American, 275(2):110–111, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-110.pdf.

Heidel:1996:WK

[3293] Raymond Heidel. Working knowledge. Scientific American, 275(2):112, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n2/pdf/scientificamerican0896-112.pdf.

Anonymous:1996:WKJ

[3294] Anonymous. Working knowledge: Jackpot! inside a slot machine. Scientific American, 275(2):112–??, August 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0896issue/0896currentissue.html.

Rennie:1996:ERH

[3295] John Rennie. From the Editors: Reasons for hope. Scientific American, 275(3):6, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
6.pdf.

Anonymous:1996:LEi

[3296] Anonymous. Letters to the Editors. Scientific American, 275(3):10, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-10.pdf.

Anonymous:1996:YAIi

[3297] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 275(3):14, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-14.pdf.

Beardsley:1996:NAH

[3298] Tim Beardsley. News and analysis: HIV’s Achilles’ Heel. Scientific American, 275(3):16–18, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
16.pdf.

Anonymous:1996:NDN

[3299] Anonymous. New drugs, in new combinations, offer relief from AIDS. Scientific American, 275(3):16–??, September 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0996issue/0996currentissue.html.

Zorpette:1996:DAF

[3300] Glenn Zorpette. A day at the Armageddon Factory. Scientific American, 275(3):20, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
20b.pdf.

Powell:1996:CP

[3301] Corey S. Powell and Madhusree Mukerjee. Cosmic puffery. Scientific American, 275(3):20–22, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
20a.pdf.

Anonymous:1996:SCCb

[3302] Anonymous. Science and the citizen: Cosmological theory begins to deflate. pollution relief. taking apart the bomb. Scientific American, 275(3):20–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Leutwyler:1996:Bd

[3303] Kristin Leutwyler. In brief. Scientific American, 275(3):22–26, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-22.pdf.

Nemecek:1996:MM

[3304] Sasha Nemecek. Mysterious maladies. Scientific American, 275(3):24–26, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-24.pdf.

Sinha:1996:SS

[3305] Gunjan Sinha. Smog from space. Scientific American, 275(3):26–30, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-26.pdf.

Mirsky:1996:PSI

[3306] Steve Mirsky. Put a sock on it. Scientific American, 275(3):30, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-30.pdf.

Doyle:1996:WBC

[3307] Rodger Doyle. World birth-control use. Scientific American, 275(3):34, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-34.pdf.

Browning:1996:ILC

[3308] John Browning. The Internet is learning to censor itself. Scientific American, 275(3):38, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
38.pdf.

Anonymous:1996:CVCc

[3309] Anonymous. Cyber view: Can PICS police the Internet? Scientific American, 275(3):38–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Schneider:1996:AN

[3310] David Schneider. Alarming nets. Scientific American, 275(3):40–42, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-40.pdf.

Anonymous:1996:TBAa

[3311] Anonymous. Technology and business: Artificial blood starts circulating. fishermen sound off for porpoises. encryption chaos continues. Scientific American, 275(3):40–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Gibbs:1996:NCO

[3312] W. Wayt Gibbs. New chip off the old block. Scientific American, 275(3):42–44, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
42.pdf.

Beardsley:1996:YEO

[3313] Tim Beardsley. For your eyes only? Scientific American, 275(3):44, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-44a.pdf.

Gibbs:1996:ABQ

[3314] W. Wayt Gibbs. Artificial blood quickens. Scientific American, 275(3):44–48, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
44b.pdf.

Eisenberg:1996:RNc

[3315] Anne Eisenberg. Recently netted. … Scientific American, 275(3):48, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-48.pdf.

Gibbs:1996:PVR

[3316] W. Wayt Gibbs. Profile: T. V. Raman: Envisioning speech. Scientific American, 275(3):52, 54, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Gibbs:1996:ES

[3317] W. Wayt Gibbs. Envisioning speech. Scientific American, 275(3):52–54, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-52.pdf.

Kase:1996:WCC

[3318] Lori Miller Kase. Why community cancer clusters are often ignored. Scientific American, 275(3):55–??, September 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0996issue/0996currentissue.html.

Rennie:1996:IMH

[3319] John Rennie and Ricki Rusting. Introduction: Making headway against cancer. Scientific American, 275(3):56–59, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-56.pdf.

Rennie:1996:MHA

[3320] John Rennie and Ricki Rusting. Making headway against cancer. Scientific American, 275(3):56–?? (Intl. ed. 28–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0996issue/0996currentissue.html.

Anonymous:1996:FU

[3321] Anonymous. Fundamental understandings. Scientific American, 275(3):61, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
61.pdf.

Weinberg:1996:FUH

[3322] Robert A. Weinberg. Fundamental understandings: How cancer arises. Scientific American, 275(3):62–70, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-62.pdf.

Weinberg:1996:HCA

[3323] Robert A. Weinberg. How cancer arises. Scientific American, 275(3): 62–?? (Intl. ed. 32–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Ruoslahti:1996:HCS

[3324] Erkki Ruoslahti. How cancer spreads. Scientific American, 275(3):72–77, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-72.pdf;   http:/
/www.sciam.com/0996issue/0996currentissue.html.

Giger:1996:ATIa

[3325] Maryellen L. Giger and Charles A. Pelizzari. Advances in tumor imaging. Scientific American, 275(3):76–??, September 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0996issue/0996currentissue.html.

Anonymous:1996:CP

[3326] Anonymous. Causes and prevention. Scientific American, 275(3):79, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-79.pdf.

Trichopoulos:1996:CPW

[3327] Dimitrios Trichopoulos, Frederick P. Li, and David J. Hunter. Causes and prevention: What causes cancer? Scientific American, 275(3):80–87, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-80.pdf.

Trichopoulos:1996:WCC

[3328] Dimitrios Trichopoulos, Frederick P. Li, and David J. Hunter. What causes cancer? Scientific American, 275(3):80–?? (Intl. ed. 50– ??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0996issue/
0996currentissue.html.

MillerKase:1996:WCC

[3329] Lori Miller Kase. Why community cancer clusters are often ignored. Scientific American, 275(3):85–??, September 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0996issue/0996currentissue.html.

Willett:1996:SMC

[3330] Walter C. Willett, Graham A. Colditz, and Nancy E. Mueller. Strategies for minimizing cancer risk. Scientific American, 275(3):88–95, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-88.pdf;   http:/
/www.sciam.com/0996issue/0996currentissue.html.

Greenwald:1996:CC

[3331] Peter Greenwald. Chemoprevention of cancer. Scientific American, 275(3):96–99, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
96.pdf;   http://www.sciam.com/0996issue/0996currentissue.html.

Davidson:1996:CCH

[3332] Nancy E. Davidson. Current controversy: Is hormone replacement therapy a risk? Scientific American, 275(3):101, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-101.pdf.

Davidson:1996:HRT

[3333] Nancy E. Davidson. Is hormone replacement therapy a risk? Scientific American, 275(3):101–?? (Intl. ed. 68–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0996issue/0996currentissue.html.

Anonymous:1996:TED

[3334] Anonymous. Toward earlier detection. Scientific American, 275(3):103, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-103.pdf;   http:
//www.sciam.com/0996issue/0996currentissue.html.

Sidransky:1996:TED

[3335] David Sidransky. Toward earlier detection: Advances in cancer detection. Scientific American, 275(3):104–109, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-104.pdf.

Sidransky:1996:ACD

[3336] David Sidransky. Advances in cancer detection. Scientific American, 275 (3):104–?? (Intl. ed. 70–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Stix:1996:GTP

[3337] Gary Stix. Is genetic testing premature? Scientific American, 275(3): 107–?? (Intl. ed. 73–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Giger:1996:ATIb

[3338] Maryellen L. Giger and Charles A. Pelizzari. Advances in tumor imaging. Scientific American, 275(3):110–112, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-110.pdf;   http://www.sciam.com/
0996issue/0996currentissue.html.

Maranto:1996:SWT

[3339] Gina Maranto. Should women in their 40s have mammograms? Scientific American, 275(3):113, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
113.pdf;   http://www.sciam.com/0996issue/0996currentissue.
html.

Hanks:1996:DSP

[3340] Gerald E. Hanks and Peter T. Scardino. Does screening for prostate cancer make sense? Scientific American, 275(3):114–115, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-114.pdf;   http:
//www.sciam.com/0996issue/0996currentissue.html.

Folkman:1996:FCT

[3341] Judah Folkman. Fighting cancer by at tacking its blood supply. Scientific American, 275(3):116–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Anonymous:1996:ICT

[3342] Anonymous. Improving conventional therapy. Scientific American, 275(3):117, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
117.pdf;   http://www.sciam.com/0996issue/0996currentissue.
html.

Hellman:1996:ICT

[3343] Samuel Hellman and Everett E. Vokes. Imroving conventional therapy: Advancing current treatments for cancer. Scientific American, 275(3):118–123, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
118.pdf.

Hellman:1996:ACT

[3344] Samuel Hellman and Everett E. Vokes. Advancing current treatments for cancer. Scientific American, 275(3):118–?? (Intl. ed. 84–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0996issue/0996currentissue.
html.

Antman:1996:WBM

[3345] Karen Antman. When are bone marrow transplants considered? Scientific American, 275(3):124–125, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
124.pdf;   http://www.sciam.com/0996issue/0996currentissue.
html.

Editors:1996:FST

[3346] The Editors. Fact sheet: Twelve major cancers. Scientific American, 275(3):126–132, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
126.pdf.

Anonymous:1996:TMC

[3347] Anonymous. Twelve major cancers. Scientific American, 275(3):126– ??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0996issue/
0996currentissue.html.

Anonymous:1996:PEP

[3348] Anonymous. The pleasures of exploring ponds. Scientific American, 275(3):132–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Anonymous:1996:IF

[3349] Anonymous. The interrogator’s fallacy. Scientific American, 275(3): 134–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0996issue/
0996currentissue.html.

Anonymous:1996:TF

[3350] Anonymous. Therapies of the future. Scientific American, 275(3):135, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-135.pdf;   http:
//www.sciam.com/0996issue/0996currentissue.html.

Old:1996:TFI

[3351] Lloyd J. Old. Therapies of the future: Immunotherapy for cancer. Scientific American, 275(3):136–143, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-136.pdf.

Old:1996:IC

[3352] Lloyd J. Old. Immunotherapy for cancer. Scientific American, 275(3): 136–?? (Intl. ed. 102–??), September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Oliff:1996:NMT

[3353] Allen Oliff, Jackson B. Gibbs, and Frank McCormick. New molecular targets for cancer therapy. Scientific American, 275(3):144–149, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-144.pdf;   http:
//www.sciam.com/0996issue/0996currentissue.html.

Folkman:1996:FCA

[3354] Judah Folkman. Fighting cancer by attacking its blood supply. Scientific American, 275(3):150–154, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-150.pdf;   http://www.sciam.com/
0996issue/0996currentissue.html.

Anonymous:1996:LC

[3355] Anonymous. Living with cancer. Scientific American, 275(3):157, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-157.pdf;   http:
//www.sciam.com/0996issue/0996currentissue.html.

Holland:1996:LCC

[3356] Jimmie C. Holland. Living with cancer: Cancer’s psychological challenges. Scientific American, 275(3):158–161, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-158.pdf.

Holland:1996:CPC

[3357] Jimmie C. Holland. Cancer’s psychological challenges. Scientific American, 275(3):158–?? (Intl. ed. 122–??), September 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0996issue/0996currentissue.html.

Aulas:1996:ACT

[3358] Jean-Jacques Aulas. Alternative cancer treatments. Scientific American, 275(3):162–163, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
162.pdf;   http://www.sciam.com/0996issue/0996currentissue.
html.

Foley:1996:CPC

[3359] Kathleen M. Foley. Controlling the pain of cancer. Scientific American, 275(3):164–165, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
164.pdf;   http://www.sciam.com/0996issue/0996currentissue.
html.

Gibbs:1996:WOI

[3360] W. Wayt Gibbs. What are obstacles to ideal care? Scientific American, 275(3):166, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
166.pdf;   http://www.sciam.com/0996issue/0996currentissue.
html.

Editors:1996:FMI

[3361] The Editors. Finding more information. Scientific American, 275(3):167, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-167.pdf.

Anonymous:1996:FMI

[3362] Anonymous. Finding more information. Scientific American, 275(3): 167–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0996issue/
0996currentissue.html.

Carlson:1996:ASP

[3363] Shawn Carlson. The amateur scientist: The pleasures of exploring ponds. Scientific American, 275(3):169–171, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-169.pdf.

Anonymous:1996:ASSb

[3364] Anonymous. The Amateur Scientist: Small ponds hold plenty of wildlife for the backyard naturalist. Scientific American, 275(3):169–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0996issue/0996currentissue.
html.

Stewart:1996:MRI

[3365] Ian Stewart. Mathematical recreations: The interrogator’s fallacy. Scientific American, 275(3):172–175, September 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n3/pdf/scientificamerican0996-172.pdf.

Anonymous:1996:MRG

[3366] Anonymous. Mathematical Recreations: Guilty or innocent? calculate the odds that a confession is true. Scientific American, 275(3):172–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Anonymous:1996:RCi

[3367] Anonymous. Reviews and commentaries. Scientific American, 275 (3):176–180, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
176.pdf.

Anonymous:1996:RCAb

[3368] Anonymous. Reviews and Commentaries: The animal origins of human morality. the alphabet takes wing. specious thinking on pollution’s dangers. Scientific American, 275(3):176–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0996issue/0996currentissue.html.

Burke:1996:CBV

[3369] James Burke. Connections: From bottled veggies to a pointillist picnic. Scientific American, 275(3):176–??, September 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0996issue/0996currentissue.html.

Gingerich:1996:WSV

[3370] Owen Gingerich. Wonders: The scientific value of prediction is overrated. Scientific American, 275(3):176–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0996issue/0996currentissue.html.

Gingerich:1996:NVN

[3371] Owen Gingerich. Neptune, Velikovsky and the name of the game. Scientific American, 275(3):181–183, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
181.pdf.

Burke:1996:I

[3372] James Burke. Impressions. Scientific American, 275(3):182–183, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n3/pdf/scientificamerican0996-182.pdf.

Aschkenasy:1996:WK

[3373] Herbert Aschkenasy. Working knowledge. Scientific American, 275(3):184, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n3/pdf/scientificamerican0996-
184.pdf.

Anonymous:1996:WKH

[3374] Anonymous. Working Knowledge: How to freeze-dry anything. Scientific American, 275(3):184–??, September 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0996issue/0996currentissue.html.

Rennie:1996:EMM

[3375] John Rennie. From the Editors: Microbes from Mars? Maybe. Scientific American, 275(4):8, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
8.pdf.

Anonymous:1996:LEj

[3376] Anonymous. Letters to the Editors. Scientific American, 275(4):10–12, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-10.pdf.

Anonymous:1996:Ee

[3377] Anonymous. Erratum. Scientific American, 275(4):12, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-12.pdf.

Anonymous:1996:YAIj

[3378] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 275(4):16, October 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-16.pdf.

Gibbs:1996:NAB

[3379] W. Wayt Gibbs and Corey S. Powell. News and analysis: Bugs in the data? Scientific American, 275(4):20–22, October 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-20.pdf.

Anonymous:1996:FLM

[3380] Anonymous. In focus: Life on Mars: How does the evidence hold up? Scientific American, 275(4):20–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Mukerjee:1996:FF

[3381] Madhusree Mukerjee. Fish fight. Scientific American, 275(4):24, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-24a.pdf.

Zorpette:1996:BBB

[3382] Glenn Zorpette. Building a better T-bone. Scientific American, 275(4): 24, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-24b.pdf.

Anonymous:1996:SCE

[3383] Anonymous. Science and the citizen: Earth’s wrong-way core. Tasmanian tiger hunt. hurricanes may be gone with the wind. Scientific American, 275(4):24–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Leutwyler:1996:Be

[3384] Kristin Leutwyler and Gunjan Sinha. In brief. Scientific American, 275(4):26, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
26.pdf.

Drollette:1996:SCB

[3385] Daniel Drollette. A spinning crystal ball. Scientific American, 275(4):28– 33, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-28.pdf.

Schneider:1996:TT

[3386] David Schneider. On the tail of the tiger. Scientific American, 275(4):32– 33, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-32.pdf.

Doyle:1996:SEC

[3387] Rodger Doyle. Soil erosion of cropland in the U.S., 1982 to 1992. Scientific American, 275(4):34, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
34b.pdf.

Schneider:1996:HH

[3388] David Schneider. Hurricane hullabaloo? Scientific American, 275(4):34, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-34a.pdf.

Mirsky:1996:JSN

[3389] Steve Mirsky. Just say NO. Scientific American, 275(4):36, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-36b.pdf.

Yam:1996:UH

[3390] Philip Yam. Unicorn hunts? Scientific American, 275(4):36–37, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-36a.pdf.

Browning:1996:TON

[3391] John Browning. Television by any other name. Scientific American, 275(4):40, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
40.pdf.

Anonymous:1996:CVNb

[3392] Anonymous. Cyber view: Networking in Europe. Scientific American, 275(4):40–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Zorpette:1996:XRM

[3393] Glenn Zorpette. “X” (rays) mark the tumor. Scientific American, 275(4):44–45, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
44.pdf.

Anonymous:1996:TBP

[3394] Anonymous. Technology and business: Programming computers genetically. defense work helps cancer researchers. Scientific American, 275(4):44–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Beardsley:1996:PC

[3395] Tim Beardsley. Picking on cotton. Scientific American, 275(4):45–48, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-45.pdf.

Gibbs:1996:PPO

[3396] W. Wayt Gibbs. Programming with primordial ooze. Scientific American, 275(4):48–50, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
48.pdf.

Krim:1996:FASa

[3397] J. Krim. Friction at the atomic scale. Scientific American, 275 (4):48–51, 54–56, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Eisenberg:1996:RNd

[3398] Anne Eisenberg. Recently netted. … Scientific American, 275(4):50, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-50.pdf.

Stix:1996:PMO

[3399] Gary Stix. Probing medicine’s outer reaches. Scientific American, 275(4):52–56, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
52.pdf.

Anonymous:1996:PWB

[3400] Anonymous. Profile: Wayne B. Jonas redirects the Office of Alternative Medicine. Scientific American, 275(4):52–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1096issue/1096currentissue.html.

Bassuk:1996:SMW

[3401] Ellen L. Bassuk, Angela Browne, and John C. Buckner. Single mothers and welfare: Contrary to popular stereotypes, few lowincome single mothers are teenagers or second-generation welfare recipients. Scientific American, 275(4):60–67, October 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-60.pdf;   http://www.sciam.com/
1096issue/1096currentissue.html.

Fredrickson:1996:MDI

[3402] James K. Fredrickson and Tullis C. Onstott. Microbes deep inside the Earth. Scientific American, 275(4):68–73, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-68.pdf;   http://www.sciam.com/
1096issue/1096currentissue.html.

Miller:1996:HUH

[3403] Steven Miller. How an underwater habitat benefits marine science. Scientific American, 275(4):68–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Krim:1996:FASb

[3404] Jacqueline Krim. Friction at the atomic scale. Scientific American, 275 (4):74–80 (Intl. ed. 48–??), October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
74.pdf;   http://www.sciam.com/1096issue/1096currentissue.html.

Lusted:1996:CCN

[3405] Hugh S. Lusted and R. Benjamin Knapp. Controlling computers with neural signals. Scientific American, 275(4):82–87 (Intl. ed. 58–63), October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-82.pdf;   http:/
/www.sciam.com/1096issue/1096currentissue.html.

Holloway:1996:TESa

[3406] Marguerite Holloway. Trends in ecology: Sounding out science. Scientific American, 275(4):82–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Edmunds:1996:TDU

[3407] Peter J. Edmunds. Ten days under the sea. Scientific American, 275(4):88–95, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
88.pdf;   http://www.sciam.com/1096issue/1096currentissue.html.

Anonymous:1996:WV

[3408] Anonymous. Working in a vacuum. Scientific American, 275(4): 90–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1096issue/
1096currentissue.html.

Anonymous:1996:MR

[3409] Anonymous. Monopoly revisited. Scientific American, 275(4):92–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/1096issue/1096currentissue.
html.

Milner:1996:CDA

[3410] Richard Milner. Charles Darwin and Associates, Ghostbusters. Scientific American, 275(4):96–101, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
96.pdf;   http://www.sciam.com/1096issue/1096currentissue.html.

Casti:1996:CSL

[3411] John L. Casti. Confronting science’s logical limits. Scientific American, 275(4):102–105, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
102.pdf;   http://www.sciam.com/1096issue/1096currentissue.
html.

Holloway:1996:TESb

[3412] Marguerite Holloway. Trends in ecology: Sounding out science. Scientific American, 275(4):106–112, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
106.pdf.

Holloway:1996:SS

[3413] Marguerite Holloway. Sounding out science. Scientific American, 275(4):106–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1096issue/1096currentissue.html.

Carlson:1996:ASW

[3414] Shawn Carlson. The amateur scientist: Working in a vacuum. Scientific American, 275(4):114–115, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
114.pdf.

Anonymous:1996:HVC

[3415] Anonymous. Homemade vacuum chambers fill a void for experimenters. Scientific American, 275(4):114–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/1096issue/1096currentissue.html.

Stewart:1996:MRMb

[3416] Ian Stewart. Mathematical recreations: Monopoly revisited. Scientific American, 275(4):116–119, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
116.pdf.

Anonymous:1996:GDJ

[3417] Anonymous. Go directly to jail: revisiting the fairness of Monopoly. Scientific American, 275(4):116–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/1096issue/1096currentissue.html.

Anonymous:1996:RCj

[3418] Anonymous. Reviews and commentaries. Scientific American, 275(4):120–124, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n4/pdf/scientificamerican1096-
120.pdf.

Anonymous:1996:WTG

[3419] Anonymous. Why things go wrong. the making of a physician. fusion on film in Chain Reaction. Scientific American, 275(4):120–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/1096issue/1096currentissue.
html.

Burke:1996:CFTb

[3420] James Burke. Connections: Fingerprinting, tanks and the average Joe. Scientific American, 275(4):120–??, October 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1096issue/1096currentissue.html.

Morrison:1996:WOW

[3421] Philip Morrison. Wonders: The Old World, the New World and humanity. Scientific American, 275(4):120–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1096issue/1096currentissue.html.

Morrison:1996:NWO

[3422] Philip Morrison. New World and Old: a matter of magnitude. Scientific American, 275(4):125–127, October 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n4/pdf/scientificamerican1096-125.pdf.

Burke:1996:MYM

[3423] James Burke. Making your mark. Scientific American, 275(4):126–127, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-126.pdf.

Holt:1996:WK

[3424] Chip Holt. Working knowledge. Scientific American, 275(4):128, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n4/pdf/scientificamerican1096-128.pdf.

Anonymous:1996:HPW

[3425] Anonymous. How photocopiers work. Scientific American, 275(4): 128–??, October 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1096issue/
1096currentissue.html.

Rennie:1996:EHQ

[3426] John Rennie. From the Editor: An honest quantum con job. Scientific American, 275(5):6, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
6.pdf.

Anonymous:1996:LEk

[3427] Anonymous. Letters to the Editors. Scientific American, 275(5):8, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-8.pdf.

Anonymous:1996:YAIk

[3428] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 275(5):10, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-10.pdf.

Gibbs:1996:NAP

[3429] W. Wayt Gibbs. News and analysis: The price of silence. Scientific American, 275(5):15–16, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
15.pdf.

Anonymous:1996:FAR

[3430] Anonymous. In focus: Academic researchers choke on industrial gag rules. Scientific American, 275(5):15–??, November 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1196issue/1196currentissue.html.

Beardsley:1996:BMS

[3431] Tim Beardsley. Bring me a shrubbery. Scientific American, 275(5):20, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-20b.pdf.

Schneider:1996:HJ

[3432] David Schneider. Hot Jupiters. Scientific American, 275(5):20–22, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-20a.pdf.

Anonymous:1996:SCHc

[3433] Anonymous. Science and the citizen: How giant planets get close to stars… self-organizing particles. … gossip as grooming. Scientific American, 275(5):20–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Leutwyler:1996:Bf

[3434] Kristin Leutwyler. In brief. Scientific American, 275(5):22–28, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-22.pdf.

Horgan:1996:MS

[3435] John Horgan. Multicultural studies. Scientific American, 275(5):24–28, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-24.pdf.

Mukerjee:1996:SB

[3436] Madhusree Mukerjee. Science with brass. Scientific American, 275(5):28–36, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
28.pdf.

Doyle:1996:GFC

[3437] Rodger Doyle. Global forest cover. Scientific American, 275(5):32, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-32.pdf.

Stix:1996:DS

[3438] Gary Stix. Different strokes. Scientific American, 275(5):36–37, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-36.pdf.

Mirsky:1996:PK

[3439] Steve Mirsky. On presidents and king. Scientific American, 275(5):37, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-37.pdf.

Browning:1996:RH

[3440] John Browning. The Rosetta Hack. Scientific American, 275(5):38, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-38.pdf.

Anonymous:1996:CVS

[3441] Anonymous. Cyber view: Still waiting for automatic translators. Scientific American, 275(5):38–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Gibbs:1996:PC

[3442] W. Wayt Gibbs. Pressure to change. Scientific American, 275(5):40–41, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-40.pdf.

Anonymous:1996:TBAb

[3443] Anonymous. Technology and business: Artificial heart… supercritical CO2 for superstrength… molding drug molecules. Scientific American, 275(5):40–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Dupont:1996:NCW

[3444] Daniel G. Dupont and Richard Lardner. Needles in a Cold War haystack. Scientific American, 275(5):41–44, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-41.pdf.

Sinha:1996:PI

[3445] Gunjan Sinha. Pump it up. Scientific American, 275(5):44–45, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-44.pdf.

Wallich:1996:MM

[3446] Paul Wallich. Molecular molds. Scientific American, 275(5):45–46, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-45.pdf.

Stix:1996:PWT

[3447] Gary Stix. Pictures worth a thousand cameras. Scientific American, 275(5):46, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
46.pdf.

Beardsley:1996:SCS

[3448] Tim Beardsley. Starting with a clean slate. Scientific American, 275(5):50–52, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
50.pdf.

Anonymous:1996:PGT

[3449] Anonymous. Profile: Geneticist Thereza Imanishi-Kari, exonerated at last. Scientific American, 275(5):50–??, November 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1196issue/1196currentissue.html.

Sperling:1996:CEV

[3450] Daniel Sperling. The case for electric vehicles. Scientific American, 275(5):54–59, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
54.pdf;   http://www.sciam.com/1196issue/1196currentissue.html.

Beck:1996:EIS

[3451] Gregory Beck and Gail S. Habicht. The evolution of the immune system: Immunity and the invertebrates. Scientific American, 275(5):60–66, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-60.pdf.

Beck:1996:II

[3452] Gregory Beck and Gail S. Habicht. Immunity and the invertebrates. Scientific American, 275(5):60–?? (Intl. ed. 42–??), November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1196issue/1196currentissue.html.

Litman:1996:SOV

[3453] Gary W. Litman. Sharks and the origins of vertebrate immunity. Scientific American, 275(5):67–71, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-67.pdf;   http://www.sciam.com/
1196issue/1196currentissue.html.

Kwiat:1996:QSD

[3454] Paul Kwiat, Harald Weinfurter, and Anton Zeilinger. Quantum seeing in the dark. Scientific American, 275(5):72–78, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-72.pdf;   http:/
/www.sciam.com/1196issue/1196currentissue.html.

Kargel:1996:GCC

[3455] Jeffrey S. Kargel and Robert G. Strom. Global climatic change on Mars. Scientific American, 275(5):80–88, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-80.pdf;   http://www.sciam.com/
1196issue/1196currentissue.html.

Prosterman:1996:CCF

[3456] Roy L. Prosterman, Tim Hanstad, and Li Ping. Can China feed itself? Scientific American, 275(5):90–96, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-90.pdf;   http://www.sciam.com/
1196issue/1196currentissue.html.

Anonymous:1996:MAA

[3457] Anonymous. Much ado about nothing. Scientific American, 275(5): 94–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1196issue/
1196currentissue.html.

Anonymous:1996:GCD

[3458] Anonymous. A guide to computer dating. Scientific American, 275(5):96–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Shaywitz:1996:D

[3459] Sally E. Shaywitz. Dyslexia. Scientific American, 275(5):98–104, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n5/pdf/scientificamerican1196-98.pdf;   http:/
/www.sciam.com/1196issue/1196currentissue.html.

Solomon:1996:RAS

[3460] Anne Solomon. Rock art in Southern Africa. Scientific American, 275(5):106–113, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
106.pdf;   http://www.sciam.com/1196issue/1196currentissue.
html.

Carlson:1996:ASM

[3461] Shawn Carlson. The amateur scientist: Much ado about nothing. Scientific American, 275(5):114–115, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-114.pdf.

Anonymous:1996:ASF

[3462] Anonymous. The amateur scientist: Fun with a jar full of nothing. Scientific American, 275(5):114–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Stewart:1996:MRG

[3463] Ian Stewart. Mathematical recreations: a guide to computer dating. Scientific American, 275(5):116–118, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n5/pdf/scientificamerican1196-116.pdf.

Anonymous:1996:MRW

[3464] Anonymous. Mathematical recreations: Will January 1, 1,000,000 A.D., be a Tuesday? Scientific American, 275(5):116–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1196issue/1196currentissue.html.

Anonymous:1996:RCk

[3465] Anonymous. Reviews and commentaries. Scientific American, 275 (5):119–123, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
119.pdf.

Anonymous:1996:RCR

[3466] Anonymous. Reviews and commentaries: Ranking the CD-ROM encyclopedias. … the physics of God, and vice versa… the morose Island of Dr. Brando. Scientific American, 275(5):119–??, November 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1196issue/1196currentissue.html.

Burke:1996:CBB

[3467] James Burke. Connections: Bagging buzz bombs and snapping stars. Scientific American, 275(5):119–??, November 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1196issue/1196currentissue.html.

Morrison:1996:WTG

[3468] Philip Morrison. Wonders: On the trail of giant squid. Scientific American, 275(5):119–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Morrison:1996:GAG

[3469] Philip Morrison. Giant against giant in the dark. Scientific American, 275(5):124–126, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
124.pdf.

Burke:1996:WGA

[3470] James Burke. What goes around comes around. Scientific American, 275(5):125–126, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
125.pdf.

Aust:1996:WKT

[3471] Edgar W. Aust. Working knowledge: Television ratings. Scientific American, 275(5):127, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n5/pdf/scientificamerican1196-
127.pdf.

Anonymous:1996:WKWb

[3472] Anonymous. Working knowledge: Watching you watch television. Scientific American, 275(5):127–??, November 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1196issue/1196currentissue.html.

Rennie:1996:EER

[3473] John Rennie. From the Editors: Experienced readers for young minds. Scientific American, 275(6):6, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
6.pdf.

Anonymous:1996:Ef

[3474] Anonymous. Erratum. Scientific American, 275(6):10, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n6/pdf/scientificamerican1296-10a.pdf.

Anonymous:1996:LEl

[3475] Anonymous. Letters to the Editors. Scientific American, 275(6):10, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-10.pdf.

Anonymous:1996:YAIl

[3476] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 275(6):14, December 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n6/pdf/scientificamerican1296-14.pdf.

Beardsley:1996:NAD

[3477] Tim Beardsley. News and analysis: Deadly enigma. Scientific American, 275(6):16–18, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
16.pdf.

Anonymous:1996:SAM

[3478] Anonymous. Safeguarding against “mad cow disease” grows more maddening. Scientific American, 275(6):16–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1296issue/1296currentissue.html.

Gibbs:1996:JM

[3479] W. Wayt Gibbs. Jungle medicine. Scientific American, 275(6):20, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-20b.pdf.

Zorpette:1996:D

[3480] Glenn Zorpette. Down the drain. Scientific American, 275(6):20–24, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-20a.pdf.

Anonymous:1996:RDN

[3481] Anonymous. Russia dumps its nuclear waste. guppy love. unmeltable ice. the Ig Nobels for 1996. Scientific American, 275(6):20–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/1296issue/1296currentissue.
html.

Leutwyler:1996:Bg

[3482] Kristin Leutwyler. In brief. Scientific American, 275(6):24–30, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-24.pdf.

Horgan:1996:BTB

[3483] John Horgan. Beyond the test ban. Scientific American, 275(6):26–30, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-26.pdf.

Doyle:1996:DCA

[3484] Rodger Doyle. Deaths caused by alcohol. Scientific American, 275(6):30–31, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
30b.pdf.

DeKoker:1996:SSC

[3485] Brenda DeKoker. Sex and the spinal cord. Scientific American, 275(6):30–32, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
30a.pdf.

Mirsky:1996:VGD

[3486] Steve Mirsky. The victors go despoiled. Scientific American, 275(6):32, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-32.pdf.

Mukerjee:1996:HM

[3487] Madhusree Mukerjee. Hard to melt. Scientific American, 275(6):33–36, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-33.pdf.

Beardsley:1996:FF

[3488] Tim Beardsley. Fear and fecundity. Scientific American, 275(6):36, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-36.pdf.

Gibbs:1996:SCG

[3489] W. Wayt Gibbs. Snap, crunch or GigaPOP? Scientific American, 275(6):38–40, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
38.pdf.

Anonymous:1996:LEM

[3490] Anonymous. A less equal, more dependable net. Scientific American, 275(6):38–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1296issue/1296currentissue.html.

Zorpette:1996:SNM

[3491] Glenn Zorpette. The sale of a new machine. Scientific American, 275(6):40–44, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
40.pdf.

Anonymous:1996:SCL

[3492] Anonymous. Scientific computing’s last stand. electric polymers welding with a match. Scientific American, 275(6):40–??, December 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1296issue/1296currentissue.html.

Gibbs:1996:WWB

[3493] W. Wayt Gibbs. Where the wind blows. Scientific American, 275(6):44, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-44.pdf.

Garcia:1996:WM

[3494] Erica Garcia. Welding with a match. Scientific American, 275(6):46, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-46a.pdf.

Leutwyler:1996:PP

[3495] Kristin Leutwyler. Plastic power. Scientific American, 275(6):46–48, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-46b.pdf.

Eisenberg:1996:RNe

[3496] Anne Eisenberg. Recently netted. … Scientific American, 275(6):48, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-48.pdf.

Duke:1996:CSHa

[3497] Richard C. Duke, David M. Ojcius, and John Ding-E Young. Cell suicide in health and disease. Scientific American, 275(6):48–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1296issue/1296currentissue.html.

Holloway:1996:MWW

[3498] Marguerite Holloway. The man who would conquer malaria. Scientific American, 275(6):52–56, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
52.pdf.

Anonymous:1996:MEP

[3499] Anonymous. Manuel Elkin Patarroyo tests his malaria vaccine, despite controversy. Scientific American, 275(6):52–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1296issue/1296currentissue.html.

Cole:1996:SBW

[3500] Leonard A. Cole. The specter of biological weapons. Scientific American, 275(6):60–65, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
60.pdf;   http://www.sciam.com/1296issue/1296currentissue.html.

Hogan:1996:PDB

[3501] Craig J. Hogan. Primordial deuterium and the Big Bang. Scientific American, 275(6):68–73 (Intl. ed. 36–??), December 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n6/pdf/scientificamerican1296-68.pdf;   http://www.sciam.com/
1296issue/1296currentissue.html.

Siegel:1996:CNM

[3502] Richard W. Siegel. Creating nanophase materials. Scientific American, 275(6):74–79 (Intl. ed. 42–??), December 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n6/pdf/scientificamerican1296-74.pdf;   http://www.sciam.com/
1296issue/1296currentissue.html.

Duke:1996:CSHb

[3503] Richard C. Duke, David M. Ojcius, and John Ding-E Young. Cell suicide in health and disease. Scientific American, 275(6):80–87, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n6/pdf/scientificamerican1296-80.pdf.

Anonymous:1996:DBS

[3504] Anonymous. Dissecting the brain with sound. Scientific American, 275(6):80–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1296issue/1296currentissue.html.

Ojcius:1996:CSH

[3505] Richard C. Duke David M. Ojcius and John Ding-E Young. Cell suicide in health and disease. Scientific American, 275(6):80–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1296issue/1296currentissue.html.

Anonymous:1996:CM

[3506] Anonymous. Cows in the maze. Scientific American, 275(6):84– ??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1296issue/
1296currentissue.html.

Hedin:1996:ADA

[3507] Lars O. Hedin and Gene E. Likens. Atmospheric dust and acid rain. Scientific American, 275(6):88–92 (Intl. ed. 56–??), December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-88.pdf;   http:/
/www.sciam.com/1296issue/1296currentissue.html.

Webb:1996:CR

[3508] Barbara Webb. A cricket robot. Scientific American, 275(6):94–99 (Intl. ed. 62–??), December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
94.pdf;   http://www.sciam.com/1296issue/1296currentissue.html.

McDowell:1996:DLA

[3509] Andrea G. McDowell. Daily life in Ancient Egypt. Scientific American, 275(6):100–105, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
100.pdf;   http://www.sciam.com/1296issue/1296currentissue.
html.

Horgan:1996:TPW

[3510] John Horgan. Trends in psychology: Why Freud isn’t dead. Scientific American, 275(6):106–111, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
106.pdf;   http://www.sciam.com/1296issue/1296currentissue.
html.

Carlson:1996:ASDc

[3511] Shawn Carlson. The amateur scientist: Dissecting the brain with sound. Scientific American, 275(6):112–115, December 1996. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v275/
n6/pdf/scientificamerican1296-112.pdf.

Anonymous:1996:ASE

[3512] Anonymous. The amateur scientist: Experiment on your own brain (safely) with a new CD. Scientific American, 275(6):112–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/1296issue/1296currentissue.
html.

Stewart:1996:MRC

[3513] Ian Stewart. Mathematical recreations: Cows in the maze. Scientific American, 275(6):116–118, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
116.pdf.

Anonymous:1996:MRM

[3514] Anonymous. Mathematical recreations: Moo-ving through the logical maze of Where Are the Cows? Scientific American, 275(6):116– ??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1296issue/
1296currentissue.html.

Morrison:1996:RCB

[3515] Philip Morrison and Phylis Morrison. Reviews and commentaries: The Scientific American Young Readers Book Awards. Scientific American, 275(6):120–125, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
120.pdf.

Burke:1996:CHCc

[3516] James Burke. Connections: Hot cocoa, German gymnastics and Lucky Lindy. Scientific American, 275(6):120–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1296issue/1296currentissue.html.

Morrison:1996:SAY

[3517] Phylis Morrison and Philip Morrison. The Scientific American Young Readers Book Awards. Scientific American, 275(6):120–??, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/1296issue/1296currentissue.
html.

Burke:1996:SD

[3518] James Burke. Sweet dreams. Scientific American, 275(6):126–127, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-126.pdf.

Anonymous:1996:AI

[3519] Anonymous. Annual index 1996. Scientific American, 275(6):129–131, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v275/n6/pdf/scientificamerican1296-129.pdf.

Rogers:1996:WKP

[3520] Wesley G. Rogers. Working knowledge: Pop tops. Scientific American, 275(6):132, December 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v275/n6/pdf/scientificamerican1296-
132.pdf.

Anonymous:1996:WKC

[3521] Anonymous. Working knowledge: Can-do thinking behind the pop top. Scientific American, 275(6):132–??, December 1996. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1296issue/1296currentissue.html.

Rennie:1997:EDR

[3522] John Rennie. From the Editor: Defending reason reasonably. Scientific American, 276(1):6, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n1/pdf/scientificamerican0197-
6.pdf.

Anonymous:1997:LEa

[3523] Anonymous. Letters to the Editors. Scientific American, 276(1):8, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-8.pdf.

Anonymous:1997:YAIa

[3524] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 276(1):10, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-10.pdf.

Rennie:1997:SBN

[3525] John Rennie, Paul Wallich, and Philip Yam. Special briefing: The 1996 Nobel Prizes in science. Scientific American, 276(1):14–18, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-14.pdf.

Anonymous:1997:NPS

[3526] Anonymous. The 1996 Nobel Prizes for Science: An overview of the outstanding work that took the honors in physics, chemistry, medicine and economics. Scientific American, 276(1):14–??, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0197issue/0197currentissue.html.

Schneider:1997:NAF

[3527] David Schneider and W. Wayt Gibbs. News and analysis: Flight of fancy. Scientific American, 276(1):22–24, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-22.pdf.

Anonymous:1997:FSD

[3528] Anonymous. In focus: Submarine deep flight readies for the seafloor, but some ask why. Scientific American, 276(1):22–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0197issue/0197currentissue.html.

Mukerjee:1997:SA

[3529] Madhusree Mukerjee. Suburban amber. Scientific American, 276(1):26, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-26b.pdf.

Beardsley:1997:SR

[3530] Tim Beardsley. Steps to recovery. Scientific American, 276(1):26–28, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-26a.pdf.

Anonymous:1997:SCSa

[3531] Anonymous. Science and the citizen: Stars in the blink of an eye… volcano forecasting. … regrowing neurons. Scientific American, 276(1):26–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0197issue/0197currentissue.html.

Anonymous:1997:B

[3532] Anonymous. In brief. Scientific American, 276(1):28–30, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-28.pdf.

Powell:1997:AT

[3533] Corey S. Powell. All in the timing. Scientific American, 276(1):29–30, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-29.pdf.

Schneider:1997:ABB

[3534] David Schneider. Awaiting the Big Bang? Scientific American, 276(1):30–32, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n1/pdf/scientificamerican0197-
30.pdf.

Mirsky:1997:CF

[3535] Steve Mirsky. Chewing the fat. Scientific American, 276(1):31, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-31.pdf.

Doyle:1997:TM

[3536] Rodger Doyle. Threatened mammals. Scientific American, 276(1):32, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-32.pdf.

Zorpette:1997:DTH

[3537] Glenn Zorpette. Do try this @home. Scientific American, 276(1):33–36, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-33.pdf.

Anonymous:1997:PCM

[3538] Anonymous. Profile: Cable-modem maven Milo Medin. Scientific American, 276(1):33–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0197issue/0197currentissue.html.

Beardsley:1997:FC

[3539] Tim Beardsley. Fetal checkup. Scientific American, 276(1):38, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-38.pdf.

Anonymous:1997:TBE

[3540] Anonymous. Technology and business: Easier prenatal tests. … squeezing more into data lines … micro-fan. Scientific American, 276(1):38–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0197issue/0197currentissue.html.

Beardsley:1997:CC

[3541] Tim Beardsley. Chilling chips. Scientific American, 276(1):40, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-40a.pdf.

Beardsley:1997:MGP

[3542] Tim Beardsley. More gallons per mile. Scientific American, 276(1):40–41, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-40b.pdf.

Gibbs:1997:BU

[3543] W. Wayt Gibbs. Bandwidth, unlimited. Scientific American, 276(1):41, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-41.pdf.

Browning:1997:NM

[3544] John Browning. No more 9 to 5. Scientific American, 276(1):42, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-42.pdf.

Anonymous:1997:CVRa

[3545] Anonymous. Cyber view: Recomputing work. Scientific American, 276(1):42–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0197issue/0197currentissue.html.

Cronin:1997:CRE

[3546] James W. Cronin, Thomas K. Gaisser, and Simon P. Swordy. Cosmic rays at the energy frontier. Scientific American, 276(1):44–49, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-44.pdf;   http:/
/www.sciam.com/1997/0107issue/0197swordy.html.

Youdim:1997:UPD

[3547] Moussa B. H. Youdim and Peter Riederer. Understanding Parkinson’s disease. Scientific American, 276(1):52–59, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-52.pdf;   http://www.sciam.com/
0197issue/0197currentissue.html.

Moin:1997:TTS

[3548] Parviz Moin and John Kim. Tackling turbulence with supercomputers. Scientific American, 276(1):62–68, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-62.pdf;   http://www.sciam.com/
1997/0107issue/0197moin.html.

Velander:1997:TLD

[3549] William H. Velander, Henryk Lubon, and William N. Drohan. Transgenic livestock as drug factories. Scientific American, 276(1):70–74, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-70.pdf;   http:/
/www.sciam.com/0197issue/0197currentissue.html.

Kennedy:1997:HBD

[3550] John M. Kennedy. How the blind draw. Scientific American, 276(1):76– 81, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-76.pdf;   http:/
/www.sciam.com/0197issue/0197currentissue.html.

Collier:1997:EFG

[3551] Michael P. Collier, Robert H. Webb, and Edmund D. Andrews. Experimental flooding in Grand Canyon. Scientific American, 276(1):82–89, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-82.pdf;   http:/
/www.sciam.com/0197issue/0197currentissue.html.

Dannen:1997:ESR

[3552] Gene Dannen. The Einstein–Szilard refrigerators. Scientific American, 276(1):90–95, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n1/pdf/scientificamerican0197-
90.pdf;   http://www.sciam.com/0197issue/0197currentissue.html.

Anonymous:1997:TSS

[3553] Anonymous. Trends in society: Science versus antiscience? Scientific American, 276(1):96–101, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n1/pdf/scientificamerican0197-
96.pdf.

Stix:1997:SVA

[3554] Gary Stix, Sasha Nemecek, and Philip Yam. Science versus antiscience? Scientific American, 276(1):96–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/0197issue/0197currentissue.html.

Carlson:1997:ASCa

[3555] Shawn Carlson. The amateur scientist: Catch a comet by its tail. Scientific American, 276(1):102–105, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-102.pdf.

Anonymous:1997:ASS

[3556] Anonymous. The amateur scientist: Sky watchers can make history by studying Comet Hale–Bopp. Scientific American, 276(1):102–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0197issue/0197currentissue.
html.

Stewart:1997:MRA

[3557] Ian Stewart. Mathematical recreations: Alphamagic squares. Scientific American, 276(1):106–109, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v276/n1/pdf/scientificamerican0197-
106.pdf.

Anonymous:1997:MRPa

[3558] Anonymous. Mathematical recreations: Presto chango! an alphabetical twist on math’s “magic squares.”. Scientific American, 276(1):106–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0197issue/0197currentissue.html.

Anonymous:1997:RCRa

[3559] Anonymous. Reviews and commentaries: Reviews and commentaries. Scientific American, 276(1):110–115, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n1/pdf/scientificamerican0197-110.pdf.

Anonymous:1997:RCTa

[3560] Anonymous. Reviews and commentaries: Two books on aging need to mature… space station in 3-D. … human evolution. … happy birthday, HAL. Scientific American, 276(1):110–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0197issue/0197currentissue.html.

Burke:1997:CBB

[3561] James Burke. Connections: Butterflies, birth control and vitamin B1. Scientific American, 276(1):110–??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/0197issue/0197currentissue.html.

Morrison:1997:WUR

[3562] Philip Morrison. Wonders: The unpredictable reshuffling of the planets. Scientific American, 276(1):110–??, January 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0197issue/0197currentissue.html.

Morrison:1997:DPS

[3563] Philip Morrison. Doing the Poincaré Shuffle. Scientific American, 276(1):116–118, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n1/pdf/scientificamerican0197-
116.pdf.

Burke:1997:BF

[3564] James Burke. A bit of a flutter. Scientific American, 276(1):117–118, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-117.pdf.

Brown:1997:WK

[3565] Rich Brown. Working knowledge. Scientific American, 276(1):119, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n1/pdf/scientificamerican0197-119.pdf.

Anonymous:1997:WKC

[3566] Anonymous. Working knowledge: The cold, hard facts about man-made snow. Scientific American, 276(1):119–??, January 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0197issue/0197currentissue.html.

Stix:1997:TS

[3567] Gary Stix, Sasha Nemecek, and Philip Yam. Trends in society. Scientific American, 276(1):??, January 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:EAQ

[3568] John Rennie. From the Editors: The animal question. Scientific American, 276(2):4, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
4.pdf.

Anonymous:1997:LEb

[3569] Anonymous. Letters to the Editors. Scientific American, 276(2):8, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-8.pdf.

Anonymous:1997:YAIb

[3570] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 276(2):10, February 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n2/pdf/scientificamerican0297-10.pdf.

Beardsley:1997:NAN

[3571] Tim Beardsley. News and analysis: The next star trek. Scientific American, 276(2):12–14, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
12.pdf.

Anonymous:1997:FUB

[3572] Anonymous. In focus: The U.S. is not so boldly going to the final frontier. Scientific American, 276(2):12–??, February 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0297issue/0297currentissue.html.

Beardsley:1997:AA

[3573] Tim Beardsley. Agent angst. Scientific American, 276(2):16, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-16b.pdf.

Horgan:1997:WCF

[3574] John Horgan. Which came first? Scientific American, 276(2):16, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-16a.pdf.

Anonymous:1997:SCF

[3575] Anonymous. Science and the citizen: Fiber-optic sponge… quasars. birds and dinosaurs… pneumonia… moose-suit science. Scientific American, 276(2):16–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Leutwyler:1997:Ba

[3576] Kristin Leutwyler. In brief. Scientific American, 276(2):18–26, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-18.pdf.

Poppe:1997:PM

[3577] Christoph Pöppe and Madhusree Mukerjee. Prize mistake. Scientific American, 276(2):22, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
22.pdf.

Sabbagh:1997:LB

[3578] Karl Sabbagh. Late bloomer. Scientific American, 276(2):23–24, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-23.pdf.

Holloway:1997:SR

[3579] Marguerite Holloway. Soaking up the rays. Scientific American, 276(2):24–26, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
24.pdf.

Beardsley:1997:GG

[3580] Tim Beardsley. Galactic gushers. Scientific American, 276(2):26–28, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-26.pdf.

Mirsky:1997:DOS

[3581] Steve Mirsky. Dropping one for science. Scientific American, 276(2):28, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-28.pdf.

Doyle:1997:UDP

[3582] Rodger Doyle. U.S. deaths from pneumonia. Scientific American, 276(2):29, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
29.pdf.

Holloway:1997:WA

[3583] Marguerite Holloway. Into the wilds of Africa. Scientific American, 276(2):30–32, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
30.pdf.

Anonymous:1997:PEP

[3584] Anonymous. Profile: Ecologist Patricia D. Moehlman defends the misunderstood jackal. Scientific American, 276(2):30–??, February 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0297issue/0297currentissue.html.

Zorpette:1997:HP

[3585] Glenn Zorpette. It’s a helicopter! It’s a plane! Scientific American, 276(2):34, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
34.pdf.

Anonymous:1997:TBM

[3586] Anonymous. Technology and business: Making grammar compute… polyester on the vine… A radical commuter copter. Scientific American, 276(2):34–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Moore:1997:NS

[3587] Samuel K. Moore. Natural synthetics. Scientific American, 276(2):36–37, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-36.pdf.

Gibbs:1997:MS

[3588] W. Wayt Gibbs. Making sense. Scientific American, 276(2):37–38, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-37.pdf.

Nemecek:1997:BB

[3589] Sasha Nemecek. Beating bacteria. Scientific American, 276(2):38–39, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-38.pdf.

Garcia:1997:SB

[3590] Erica Garcia. Scoring with buckyballs. Scientific American, 276(2):39, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-39.pdf.

Browning:1997:UD

[3591] John Browning. Universal disservice. Scientific American, 276(2):40, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-40.pdf.

Anonymous:1997:CVH

[3592] Anonymous. Cyber view: How not to wire the poor. Scientific American, 276(2):40–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Landry:1997:ICA

[3593] Donald W. Landry. Immunotherapy for cocaine addiction. Scientific American, 276(2):42–45, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
42.pdf;   http://www.sciam.com/0297issue/0297currentissue.html.

Massonnet:1997:SRI

[3594] Didier Massonnet. Satellite radar interferometry. Scientific American, 276(2):46–53, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
46.pdf;   http://www.sciam.com/1997/0207issue/0297massonnet.
html.

Bothun:1997:GGa

[3595] Gregory D. Bothun. The ghostliest galaxies. Scientific American, 276(2):56–61, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
56.pdf.

Baldwin:1997:LKE

[3596] Neil Baldwin. The lesser known Edison. Scientific American, 276(2):62–67, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
62.pdf;   http://www.sciam.com/0297issue/0297currentissue.html.

Losick:1997:WHB

[3597] Richard Losick and Dale Kaiser. Why and how bacteria communicate. Scientific American, 276(2):68–73, February 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n2/pdf/scientificamerican0297-68.pdf;   http://www.sciam.com/
0297issue/0297currentissue.html.

Crandall:1997:CLN

[3598] Richard E. Crandall. The challenge of large numbers. Scientific American, 276(2):74–78, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
74.pdf;   http://www.sciam.com/0297issue/0297currentissue.html.

Rowan:1997:FBE

[3599] Andrew N. Rowan. Forum: The benefits and ethics of animal research. Scientific American, 276(2):79, February 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n2/pdf/scientificamerican0297-79.pdf.

Anonymous:1997:FBE

[3600] Anonymous. Forum: The benefits and ethics of animal research. Scientific American, 276(2):79–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Barnard:1997:ARW

[3601] Neal D. Barnard and Stephen R. Kaufman. Animal research is wasteful and misleading. Scientific American, 276(2):80–82, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-80.pdf;   http:/
/www.sciam.com/1997/0207issue/0297barnard.html.

Bothun:1997:GGb

[3602] Gregory D. Bothun. The ghostliest galaxies. Scientific American, 276 (2):80–?? (Intl. ed. 56–61), February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Botting:1997:ARV

[3603] Jack H. Botting and Adrian R. Morrison. Animal research is vital to medicine. Scientific American, 276(2):83–85, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-83.pdf;   http:/
/www.sciam.com/1997/0207issue/0297botting.html.

Mukerjee:1997:TAR

[3604] Madhusree Mukerjee. Trends in animal research. Scientific American, 276(2):86–93, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
86.pdf;   http://www.sciam.com/1997/0207issue/0297trends.html.

Carlson:1997:ASP

[3605] Shawn Carlson. The amateur scientist: a picture-perfect comet. Scientific American, 276(2):94–95, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
94.pdf.

Anonymous:1997:ASCa

[3606] Anonymous. The amateur scientist: Capturing Hale–Bopp, the comet of the century. Scientific American, 276(2):94–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0297issue/0297currentissue.html.

Stewart:1997:MRCa

[3607] Ian Stewart. Mathematical recreations: Crystallography of a golf ball. Scientific American, 276(2):96–98, February 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n2/pdf/scientificamerican0297-96.pdf.

Anonymous:1997:MRDa

[3608] Anonymous. Mathematical recreations: The dimpled symmetries of golf balls. Scientific American, 276(2):96–??, February 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0297issue/0297currentissue.html.

Ferris:1997:RCH

[3609] Timothy Ferris and Jeffrey L. Meikle. Reviews and commentaries: Heaven on Earth. Scientific American, 276(2):100–104, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n2/pdf/scientificamerican0297-100.pdf.

Anonymous:1997:RCP

[3610] Anonymous. Reviews and commentaries: A professional CD-ROM tool for amateur astronomers. … history of plastic… Da Vinci on disk. Scientific American, 276(2):100–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Arthur:1997:WRE

[3611] W. Brian Arthur. Wonders: The rocketing evolution of technology. Scientific American, 276(2):100–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Burke:1997:COS

[3612] James Burke. Connections: Oh, say, can you see… where this song came from? Scientific American, 276(2):100–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0297issue/0297currentissue.html.

Arthur:1997:HFT

[3613] W. Brian Arthur. How fast is technology evolving? Scientific American, 276(2):105–107, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n2/pdf/scientificamerican0297-
105.pdf.

Burke:1997:WF

[3614] James Burke. Waving the flag. Scientific American, 276(2):106–107, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-106.pdf.

Pierce:1997:WK

[3615] Michael A. Pierce. Working knowledge. Scientific American, 276(2):108, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n2/pdf/scientificamerican0297-108.pdf.

Anonymous:1997:WKF

[3616] Anonymous. Working knowledge: Fixing a knee from the inside. Scientific American, 276(2):108–??, February 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0297issue/0297currentissue.html.

Rennie:1997:ECI

[3617] John Rennie. From the Editors: Civilizing the Internet. Scientific American, 276(3):6, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
6.pdf.

Anonymous:1997:LEc

[3618] Anonymous. Letters to the Editors. Scientific American, 276(3):10, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-10.pdf.

Anonymous:1997:YAIc

[3619] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 276(3):12, March 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n3/pdf/scientificamerican0397-12.pdf.

Gibbs:1997:NAC

[3620] W. Wayt Gibbs. News and analysis: Computer bombs. Scientific American, 276(3):14–16, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
14.pdf.

Anonymous:1997:FUS

[3621] Anonymous. In focus: University supercomputers compete to simulate nuclear weapons. Scientific American, 276(3):14–??, March 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0397issue/0397currentissue.html.

Nemecek:1997:AL

[3622] Sasha Nemecek. Amphibians on-line. Scientific American, 276(3):18, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-18b.pdf.

Leutwyler:1997:SP

[3623] Kristin Leutwyler. Suicide prevention. Scientific American, 276(3):18– 20, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-18a.pdf.

Anonymous:1997:SCLa

[3624] Anonymous. Science and the citizen: The linguistics and politics of “Ebonics”. the neurobiology of suicide. seafloor storage for radioactive waste. Scientific American, 276(3):18–??, March 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0397issue/0397currentissue.html.

Schneider:1997:MB

[3625] David Schneider. Not in my backyard. Scientific American, 276(3):20– 22, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-20b.pdf.

Leutwyler:1997:Bb

[3626] Kristin Leutwyler. In brief. Scientific American, 276(3):20–24, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-20a.pdf.

Garcia:1997:WDT

[3627] Erica Garcia. Where do turtles go? Scientific American, 276(3):24–25, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-24.pdf.

Mirsky:1997:BB

[3628] Steve Mirsky. Body blow. Scientific American, 276(3):25, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n3/pdf/scientificamerican0397-25b.pdf.

Gibbs:1997:ML

[3629] W. Wayt Gibbs. A matter of language. Scientific American, 276(3):25– 27, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-25a.pdf.

Doyle:1997:GFP

[3630] Rodger Doyle. Global fertility and population. Scientific American, 276(3):26, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
26.pdf.

Horgan:1997:JA

[3631] John Horgan. Juggling act. Scientific American, 276(3):28–30, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-28.pdf.

Anonymous:1997:PRG

[3632] Anonymous. Profile: Ronald L. Graham of AT&T Labs-Research has his (juggling) hands full. Scientific American, 276(3):28–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0397issue/0397currentissue.html.

Beardsley:1997:MMO

[3633] Tim Beardsley. Memories are made of. … Scientific American, 276(3):32–33, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
32.pdf.

Anonymous:1997:TBC

[3634] Anonymous. Technology and business: Chinese quandary over biotech and eugenics. memory drugs? food-poisoning sensor. Scientific American, 276(3):32–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0397issue/0397currentissue.html.

Beardsley:1997:CS

[3635] Tim Beardsley. China syndrome. Scientific American, 276(3):33–34, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-33.pdf.

Powell:1997:MC

[3636] Corey S. Powell. A mirror, cheaply. Scientific American, 276(3):34, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-34.pdf.

Zorpette:1997:BRT

[3637] Glenn Zorpette. Better red than dead. Scientific American, 276(3):36, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-36.pdf.

Eisenberg:1997:WM

[3638] Anne Eisenberg. Where the money is. Scientific American, 276(3):37, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-37.pdf.

Anonymous:1997:CVF

[3639] Anonymous. Cyber view: Fishing for money, some services cast a small Net. Scientific American, 276(3):37–??, March 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0397issue/0397currentissue.html.

Lang:1997:SRS

[3640] Kenneth R. Lang. SOHO reveals the secrets of the Sun. Scientific American, 276(3):40–47, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
40.pdf;   http://www.sciam.com/0397issue/0397currentissue.html.

Anonymous:1997:IBO

[3641] Anonymous. The Internet: Bringing order from chaos. Scientific American, 276(3):49–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0397issue/0397currentissue.html.

Editors:1997:SRI

[3642] The Editors. Special report: The Internet: Bringing order from chaos. Scientific American, 276(3):50–51, March 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n3/pdf/scientificamerican0397-50.pdf.

Lynch:1997:SI

[3643] Clifford Lynch. Searching the Internet. Scientific American, 276(3):52– 56, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-52.pdf.

Stix:1997:FPW

[3644] Gary Stix. Finding pictures on the Web. Scientific American, 276(3):54– 55, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-54.pdf.

Lesk:1997:GD

[3645] Michael Lesk. Going digital. Scientific American, 276(3):58–60, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-58.pdf.

Resnick:1997:FII

[3646] Paul Resnick. Filtering information on the Internet. Scientific American, 276(3):62–64, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
62.pdf.

Hearst:1997:ISW

[3647] Marti A. Hearst. Interfaces for searching the Web. Scientific American, 276(3):68–72, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
68.pdf.

Raman:1997:WM

[3648] T. V. Raman. Websurfing without a monitor. Scientific American, 276(3):73, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
73.pdf;   http://www.sciam.com/0397issue/0397currentissue.html.

Notes the importance of structure in online documents, and decries the trend in vendor extensions of HTML to unstructured visual markup.

Oudet:1997:MI

[3649] Bruno Oudet. Multilingualism on the Internet. Scientific American, 276(3):77–78, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
77.pdf;   http://www.sciam.com/0397issue/0397currentissue.html.

Discusses the problems of multilingualism, with brief mention of ISO10646 and Unicode character sets, and of automated language translation systems in current use.

Stefik:1997:TS

[3650] Mark Stefik. Trusted systems. Scientific American, 276(3):78–81, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-78.pdf.

Kahle:1997:PI

[3651] Brewster Kahle. Preserving the Internet. Scientific American, 276(3):82– 83, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-82.pdf.

Kleinman:1997:PGC

[3652] Arthur Kleinman and Alex Cohen. Psychiatry’s global challenge. Scientific American, 276(3):86–89, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
86.pdf;   http://www.sciam.com/0397issue/0397currentissue.html.

Haseltine:1997:DGN

[3653] William A. Haseltine. Discovering genes for new medicines. Scientific American, 276(3):92–98, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
92.pdf;   http://www.sciam.com/0397issue/0397currentissue.html.

Ouboter:1997:HKO

[3654] Rudolf de Bruyn Ouboter. Heike Kamerlingh Onnes’s discovery of superconductivity. Scientific American, 276(3):98–103, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n3/pdf/scientificamerican0397-98.pdf.

deBruynOuboter:1997:HKO

[3655] Rudolf de Bruyn Ouboter. Heike Kamerlingh Onnes’s discovery of superconductivity. Scientific American, 276(3):98–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0397issue/0397currentissue.html.

Seymour:1997:PWT

[3656] Roger S. Seymour. Plants that warm themselves. Scientific American, 276(3):104–109, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
104.pdf;   http://www.sciam.com/0397issue/0397currentissue.
html.

Schneider:1997:TCRa

[3657] David Schneider. Trends in climate research: The rising seas. Scientific American, 276(3):112–117, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
112.pdf.

Schneider:1997:RS

[3658] David Schneider. The rising seas. Scientific American, 276(3):112– ?? (Intl. ed. 96–101), March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0397issue/0397currentissue.html.

Stewart:1997:MRJ

[3659] Ian Stewart. Mathematical recreations: Juniper green. Scientific American, 276(3):118–120, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
118.pdf.

Anonymous:1997:MRPb

[3660] Anonymous. Mathematical recreations: Prime-time fun playing Juniper Green. Scientific American, 276(3):118–??, March 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0397issue/0397currentissue.html.

Carlson:1997:ASA

[3661] Shawn Carlson. The amateur scientist: Algorithm of the gods. Scientific American, 276(3):121–123, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
121.pdf.

Anonymous:1997:ASTa

[3662] Anonymous. The amateur scientist: Telescopes and a problem for traveling salesmen in space. Scientific American, 276(3):121–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0397issue/0397currentissue.html.

Anonymous:1997:RCa

[3663] Anonymous. Reviews and commentaries. Scientific American, 276(3):124–128, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
124.pdf.

Anonymous:1997:RCVa

[3664] Anonymous. Reviews and commentaries: Vision works both ways. who gets credit for the big bang?. beetles on the brain. Scientific American, 276(3):124–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0397issue/0397currentissue.html.

Burke:1997:CCD

[3665] James Burke. Connections: From a Copernican disclaimer to trench warfare. Scientific American, 276(3):124–??, March 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0397issue/0397currentissue.html.

Morrison:1997:WSM

[3666] Philip Morrison and Phylis Morrison. Wonders: Sculpting with molecules and atoms. Scientific American, 276(3):124–??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0397issue/0397currentissue.html.

Morrison:1997:MCM

[3667] Philip Morrison and Phylis Morrison. Molecular crayons and mustard seed avalanches. Scientific American, 276(3):129–131, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n3/pdf/scientificamerican0397-129.pdf.

Burke:1997:RS

[3668] James Burke. Revolutionary stuff. Scientific American, 276(3):130–131, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n3/pdf/scientificamerican0397-130.pdf.

Scott:1997:WKS

[3669] Brian Scott. Working knowledge: Stopping bullets. Scientific American, 276(3):132, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n3/pdf/scientificamerican0397-
132.pdf.

Anonymous:1997:WKB

[3670] Anonymous. Working knowledge: Bullet-resistant vests: dressed to the nine millimeters. Scientific American, 276(3):132–??, March 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0397issue/0397currentissue.html.

Anonymous:1997:SRa

[3671] Anonymous. Special report. Scientific American, 276(3):??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1997/0307issue/0397intro.html.

Schneider:1997:TCRb

[3672] David Schneider. Trends in climate research: The rising seas. Scientific American, 276(3):??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:ESW

[3673] John Rennie. From the Editor: Seriously, we’re not kidding. Scientific American, 276(4):8, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
8.pdf.

Anonymous:1997:LEd

[3674] Anonymous. Letters to the Editors. Scientific American, 276(4):10, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-10.pdf.

Anonymous:1997:YAId

[3675] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 276(4):12, April 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n4/pdf/scientificamerican0497-12.pdf.

Stix:1997:NAS

[3676] Gary Stix. News and analysis: Small (lending) is beautiful. Scientific American, 276(4):16–20, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
16.pdf.

Anonymous:1997:FMH

[3677] Anonymous. In focus: Microcreditors help to stem poverty. Scientific American, 276(4):16–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0497issue/0497currentissue.html.

Schneider:1997:HS

[3678] David Schneider. Hot-spotting. Scientific American, 276(4):22–24, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-22.pdf.

Anonymous:1997:SCH

[3679] Anonymous. Science and the citizen: Hot spots; Europa; minimizing stroke damage; splicing saffron; what’s your EQ?; electric car ride. Scientific American, 276(4):22–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0497issue/0497currentissue.html.

Garcia:1997:SS

[3680] Erica Garcia. Salivating for saffron. Scientific American, 276(4):24–25, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-24b.pdf.

Leutwyler:1997:Bc

[3681] Kristin Leutwyler. In brief. Scientific American, 276(4):24–28, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-24a.pdf.

Mirsky:1997:SE

[3682] Steve Mirsky. Separate but EQ. Scientific American, 276(4):25, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-25b.pdf.

Zorpette:1997:GCI

[3683] Glenn Zorpette. A golf cart, it isn’t. Scientific American, 276(4):25–26, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-25a.pdf.

Leutwyler:1997:SS

[3684] Kristin Leutwyler. Stopping strokes. Scientific American, 276(4):26–28, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-26.pdf.

Doyle:1997:APU

[3685] Rodger Doyle. Air pollution in the U.S. Scientific American, 276(4):27, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-27.pdf.

Powell:1997:GE

[3686] Corey S. Powell. The greening of Europa. Scientific American, 276(4):28– 29, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-28.pdf.

Gibbs:1997:SZ

[3687] W. Wayt Gibbs. From Satan to Zen. Scientific American, 276(4):32–34, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-32.pdf.

Anonymous:1997:PDF

[3688] Anonymous. Profile: Dan Farmer, computer security expert, hacks up the Web. Scientific American, 276(4):32–??, April 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0497issue/0497currentissue.html.

Dupont:1997:PN

[3689] Daniel G. Dupont. Playing nice. Scientific American, 276(4):35–38, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-35.pdf.

Anonymous:1997:TBD

[3690] Anonymous. Technology and business: Defense cost sharing; drugnet: catching cholesterol; flexible batteries. Scientific American, 276(4):35–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0497issue/0497currentissue.
html.

Gibbs:1997:DDS

[3691] W. Wayt Gibbs. Dredging the digestive system. Scientific American, 276(4):38, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
38.pdf.

Beardsley:1997:PP

[3692] Tim Beardsley. Plastic power. Scientific American, 276(4):39–40, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-39.pdf.

Stix:1997:CYB

[3693] Gary Stix. Check your bags. Scientific American, 276(4):40, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n4/pdf/scientificamerican0497-40.pdf.

Wallich:1997:CUC

[3694] Paul Wallich. Cracking the U.S. code. Scientific American, 276(4):42, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-42.pdf.

Anonymous:1997:CVC

[3695] Anonymous. Cyber view: Is a code cracker a concealed weapon? Scientific American, 276(4):42–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0497issue/0497currentissue.html.

Rice:1997:CSM

[3696] Richard E. Rice, Raymond E. Gullison, and John W. Reid. Can sustainable management save tropical forests? Scientific American, 276(4):44– 49, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-44.pdf;   http:/
/www.sciam.com/0497issue/0497currentissue.html.

Susskind:1997:BHI

[3697] Leonard Susskind. Black holes and the information paradox. Scientific American, 276(4):52–57, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
52.pdf;   http://www.sciam.com/0497issue/0497currentissue.html.

Tattersall:1997:AAA

[3698] Ian Tattersall. Out of Africa again… and again? Scientific American, 276(4):60–67, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
60.pdf;   http://www.sciam.com/0497issue/0497currentissue.html.

Plunkett:1997:CCN

[3699] Matthew J. Plunkett and Jonathan A. Ellman. Combinatorial chemistry and new drugs. Scientific American, 276(4):68–73, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-68.pdf;   http:/
/www.sciam.com/0497issue/0497currentissue.html.

Pinter:1997:HEB

[3700] Nicholas Pinter and Mark T. Brandon. How erosion builds mountains. Scientific American, 276(4):74–79, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
74.pdf;   http://www.sciam.com/0497issue/0497currentissue.html.

Madigan:1997:E

[3701] Michael T. Madigan and Barry L. Marrs. Extremophiles. Scientific American, 276(4):82–87, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
82.pdf;   http://www.sciam.com/1997/0007issue/0497marrs.html.

Matthews:1997:SML

[3702] Robert A. J. Matthews. The science of Murphy’s Law. Scientific American, 276(4):88–91, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
88.pdf;   http://www.sciam.com/0497issue/0497currentissue.html.

Evans:1997:JVM

[3703] Arthur B. Evans and Ron Miller. Jules Verne, misunderstood visionary. Scientific American, 276(4):92–97, April 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n4/pdf/scientificamerican0497-92.pdf;   http://www.sciam.com/
1997/0007issue/0497evans.html.

Carlson:1997:ASJ

[3704] Shawn Carlson. The amateur scientist: The joys of armchair ornithology. Scientific American, 276(4):100–101, April 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n4/pdf/scientificamerican0497-100.pdf.

Anonymous:1997:ASA

[3705] Anonymous. The amateur scientist: Armchair ornithology is easy, but beware: it can be addictive. Scientific American, 276(4):100–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0497issue/0497currentissue.
html.

Stewart:1997:MRK

[3706] Ian Stewart. Mathematical recreations: Knight’s tours. Scientific American, 276(4):102–104, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
102.pdf.

Anonymous:1997:MRTa

[3707] Anonymous. Mathematical recreations: Taking the knight’s tour of a chessboard. Scientific American, 276(4):102–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0497issue/0497currentissue.html.

Anonymous:1997:RCb

[3708] Anonymous. Reviews and commentaries. Scientific American, 276(4):108–112, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
108.pdf.

Anonymous:1997:RCU

[3709] Anonymous. Reviews and commentaries: An updated history of cryptology “forgotten genius” Nikola Tesla; archaeological eyewitnesses. Scientific American, 276(4):108–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0497issue/0497currentissue.html.

Burke:1997:CRO

[3710] James Burke. Connections: The romantic overtones of oil exploration. Scientific American, 276(4):108–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/0497issue/0497currentissue.html.

Morrison:1997:WSS

[3711] Philip Morrison. Wonders: Scents and sensibility. Scientific American, 276(4):108–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0497issue/0497currentissue.html.

Morrison:1997:SG

[3712] Philip Morrison. The silicon gourmet. Scientific American, 276(4):113– 115, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-113.pdf.

Burke:1997:FF

[3713] James Burke. Feathered friends. Scientific American, 276(4):114–115, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n4/pdf/scientificamerican0497-114.pdf.

Bunker:1997:WKS

[3714] Merton Bunker, Jr. Working knowledge: Smoke detectors. Scientific American, 276(4):116, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n4/pdf/scientificamerican0497-
116.pdf.

Anonymous:1997:WKW

[3715] Anonymous. Working knowledge: When there’s smoke, these fire. Scientific American, 276(4):116–??, April 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0497issue/0497currentissue.html.

Rennie:1997:EMC

[3716] John Rennie. From the Editor: The misunderstood clone. Scientific American, 276(5):8, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
8.pdf.

Anonymous:1997:LEe

[3717] Anonymous. Letters to the Editors. Scientific American, 276(5):10, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-10.pdf.

Anonymous:1997:YAIe

[3718] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 276(5):12, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
12.pdf.

Beardsley:1997:NASa

[3719] Tim Beardsley. News and analysis: The start of something big? Scientific American, 276(5):15–16, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
15.pdf.

Anonymous:1997:FBB

[3720] Anonymous. In focus: Baa baa, cloned sheep, have you any worth? The ethics and conundrums of Dolly. Scientific American, 276(5):15–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0597issue/0597currentissue.
html.

Powell:1997:VW

[3721] Corey S. Powell. Vanishing world. Scientific American, 276(5):17–18, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-17.pdf.

Anonymous:1997:SCLb

[3722] Anonymous. Science and the citizen: Left-handed meteorite; disappearing planets?; faking a memory; scent trek. Scientific American, 276(5):17–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Horgan:1997:SC

[3723] John Horgan. The sinister Cosmos. Scientific American, 276(5):18– 24, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-18b.pdf.

Leutwyler:1997:Bd

[3724] Kristin Leutwyler. In brief. Scientific American, 276(5):18–28, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-18a.pdf.

Doyle:1997:FIW

[3725] Rodger Doyle. Female illiteracy worldwide. Scientific American, 276(5): 20, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-20.pdf.

Beardsley:1997:TGB

[3726] Tim Beardsley. As time goes by. … Scientific American, 276(5):24– 28, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-24.pdf.

Stix:1997:SP

[3727] Gary Stix. Scent prospectors. Scientific American, 276(5):28–30, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-28.pdf.

Mirsky:1997:CT

[3728] Steve Mirsky. Coffee talk. Scientific American, 276(5):30, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-30.pdf.

Zorpette:1997:ECP

[3729] Glenn Zorpette. Electric cars and pterosaurs are my business. Scientific American, 276(5):32–36, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
32.pdf.

Anonymous:1997:PEC

[3730] Anonymous. Profile: Electric-car designer Alan Cocconi gets a charge out of beating Detroit. Scientific American, 276(5):32–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0597issue/0597currentissue.html.

Stix:1997:UW

[3731] Gary Stix. Under the wire. Scientific American, 276(5):38–40, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-38.pdf.

Anonymous:1997:TBH

[3732] Anonymous. Technology and business: How wires trip up chips; selling electricity to utilities; anti-Parkinson’s implant. Scientific American, 276(5):38–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Nemecek:1997:SA

[3733] Sasha Nemecek. Strange attractors. Scientific American, 276(5):40, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-40.pdf.

Schneider:1997:PP

[3734] David Schneider. Power to the people. Scientific American, 276(5): 44, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-44a.pdf.

Gibbs:1997:TT

[3735] W. Wayt Gibbs. Taming tremor. Scientific American, 276(5):44–45, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-44b.pdf.

Gibbs:1997:WWW

[3736] W. Wayt Gibbs. World wide widgets. Scientific American, 276(5): 48, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-48.pdf.

Anonymous:1997:CVM

[3737] Anonymous. Cyber view: Metacomputing: the Internet route to giant computers. Scientific American, 276(5):48–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0597issue/0597currentissue.html.

Packer:1997:DWF

[3738] Craig Packer and Anne E. Pusey. Divided we fall: Cooperation among lions. Scientific American, 276(5):52–59, May 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-52.pdf;   http://www.sciam.com/
0597issue/0597currentissue.html.

Helmreich:1997:MHE

[3739] Robert L. Helmreich. Managing human error in aviation. Scientific American, 276(5):62–67, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
62.pdf;   http://www.sciam.com/0597issue/0597currentissue.html.

Horwitz:1997:IHa

[3740] Alan F. Horwitz. Integrins and health. Scientific American, 276(5):68– 75, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-68.pdf.

Horwitz:1997:IHI

[3741] Alan F. Horwitz. Integrins and health: Integrins are a class of adhesion molecules that “glue” cells in place. surprisingly, at a fundamental level, they also regulate most functions of the body. Scientific American, 276(5):68–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Karl:1997:CC

[3742] Thomas R. Karl, Neville Nicholls, and Jonathan Gregory. The coming climate. Scientific American, 276(5):78–83, May 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-78.pdf;   http://www.sciam.com/
1997/0507issue/0597karl.html.

Ewen:1997:NCT

[3743] John A. Ewen. New chemical tools to create plastics. Scientific American, 276(5):86–91, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
86.pdf;   http://www.sciam.com/1997/0507issue/0597ewen.html.

Macchetto:1997:GYU

[3744] F. Duccio Macchetto and Mark Dickinson. Galaxies in the Young Universe. Scientific American, 276(5):92–99, May 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-92.pdf;   http://www.sciam.com/
0597issue/0597currentissue.html.

Horgan:1997:THC

[3745] John Horgan. Trends in health care: Seeking a better way to die. Scientific American, 276(5):100–105, May 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-100.pdf.

Horgan:1997:SBW

[3746] John Horgan. Seeking a better way to die. Scientific American, 276(5):100–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Carlson:1997:ASW

[3747] Shawn Carlson. The amateur scientist: When hazy skies are rising. Scientific American, 276(5):106–107, May 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-106.pdf.

Anonymous:1997:ASTb

[3748] Anonymous. The amateur scientist: Take crystal-clear readings of atmospheric haze. Scientific American, 276(5):106–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0597issue/0597currentissue.html.

Stewart:1997:MRB

[3749] Ian Stewart. Mathematical recreations: Big game hunting in primeland. Scientific American, 276(5):108–111, May 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n5/pdf/scientificamerican0597-108.pdf.

Anonymous:1997:MRH

[3750] Anonymous. Mathematical recreations: Hunting for giant primes. Scientific American, 276(5):108–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Anonymous:1997:RCc

[3751] Anonymous. Reviews and commentaries. Scientific American, 276(5):112–116, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
112.pdf.

Anonymous:1997:RCF

[3752] Anonymous. Reviews and commentaries: Four books make complexity less confusing; the bomb on the coffee table; Darwin goes to the movies. Scientific American, 276(5):112–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Burke:1997:CIM

[3753] James Burke. Connections: The “influence machine,” Mother Goose and the Rosetta Stone. Scientific American, 276(5):112–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0597issue/0597currentissue.html.

Morrison:1997:WWC

[3754] Philip Morrison. Wonders: Watery clues to life in space. Scientific American, 276(5):112–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Morrison:1997:HLE

[3755] Philip Morrison. HOH and life elsewhere. Scientific American, 276(5):117–119, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
117.pdf.

Burke:1997:SB

[3756] James Burke. Stones and bones. Scientific American, 276(5):118–119, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n5/pdf/scientificamerican0597-118.pdf.

Ryan:1997:WKA

[3757] Joseph Ryan, Jr. Working knowledge: Antishoplifting labels. Scientific American, 276(5):120, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n5/pdf/scientificamerican0597-
120.pdf.

Anonymous:1997:WKE

[3758] Anonymous. Working knowledge: Electronic labels fight shoplifting. Scientific American, 276(5):120–??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0597issue/0597currentissue.html.

Anonymous:1997:THC

[3759] Anonymous. Trends in health care. Scientific American, 276(5):??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Horwitz:1997:IHb

[3760] Alan F. Horwitz. Integrins and health. Scientific American, 276(5): ??, May 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:ERT

[3761] John Rennie. From the Editor: Recognizing technological genius. Scientific American, 276(6):8, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
8.pdf.

Anonymous:1997:LEAa

[3762] Anonymous. Letters to the Editors: Animal experimentation debate, round two. Scientific American, 276(6):10, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-10.pdf.

Anonymous:1997:YAIf years ago. Scientific American, 276(6):12, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-12.pdf.

Editors:1997:SBN

[3764] The Editors. Special briefing: The 1997 National Medal of Technology. Scientific American, 276(6):16, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-16.pdf.

Anonymous:1997:NMT

[3765] Anonymous. The National Medal of Technology. Scientific American, 276(6):16–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0697issue/0697currentissue.html.

Horgan:1997:NAP

[3766] John Horgan. News and analysis: Pinning down inflation. Scientific American, 276(6):17–18, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
17.pdf.

Anonymous:1997:FCI

[3767] Anonymous. In focus: Is cosmic inflation overblown? testing the means to test the theory. Scientific American, 276(6):17–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0697issue/0697currentissue.html.

Beardsley:1997:WE

[3768] Tim Beardsley. War without end? Scientific American, 276(6):20–22, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-20.pdf.

Anonymous:1997:SCB

[3769] Anonymous. Science and the citizen: Banning land mines; nitrogen overload; fruit flies flummoxed. Scientific American, 276(6):20–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0697issue/0697currentissue.html.

Mukerjee:1997:BB

[3770] Madhusree Mukerjee. Bee blight. Scientific American, 276(6):22–24, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-22b.pdf.

Leutwyler:1997:Be

[3771] Kristin Leutwyler. In brief. Scientific American, 276(6):22–26, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-22a.pdf.

Beardsley:1997:WNT

[3772] Tim Beardsley. When nutrients turn noxious. Scientific American, 276(6):24–26, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
24.pdf.

Horgan:1997:SFV

[3773] John Horgan. Sex, flies and videotape. Scientific American, 276(6):26– 31, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-26.pdf.

Doyle:1997:LCU

[3774] Rodger Doyle. Lung cancer in U.S. males. Scientific American, 276(6): 28, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-28.pdf.

Mirsky:1997:SF

[3775] Steve Mirsky. Small fry. Scientific American, 276(6):30, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-30.pdf.

Schneider:1997:SH

[3776] David Schneider. Scanning the horizon. Scientific American, 276(6):32– 34, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-32.pdf.

Anonymous:1997:PIF

[3777] Anonymous. Profile: The indomitable father of MRI, Raymond V. Damadian. Scientific American, 276(6):32–??, June 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0697issue/0697currentissue.html.

Zorpette:1997:SB

[3778] Glenn Zorpette. Seller beware. Scientific American, 276(6):36–38, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-36.pdf.

Anonymous:1997:TBI

[3779] Anonymous. Technology and business: Inventing islands; nuclear sellout; arthritis treatment roundup; the U.S.’s flying saucer. Scientific American, 276(6):36–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0697issue/0697currentissue.html.

Stix:1997:FG

[3780] Gary Stix. Floating giants. Scientific American, 276(6):38, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-38.pdf.

Zorpette:1997:SS

[3781] Glenn Zorpette. Spying saucer. Scientific American, 276(6):40, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-40b.pdf.

Mukerjee:1997:MM

[3782] Madhusree Mukerjee. Medical mismatch. Scientific American, 276(6):40– 41, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-40a.pdf.

Nemecek:1997:AA

[3783] Sasha Nemecek. Attacking arthritis. Scientific American, 276(6):41, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-41.pdf.

Eisenberg:1997:DI

[3784] Anne Eisenberg. Disliking the Internet. Scientific American, 276(6):44, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-44.pdf.

Anonymous:1997:CVDa

[3785] Anonymous. Cyber view: Demonizing the Internet. Scientific American, 276(6):44–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0697issue/0697currentissue.html.

Schwarzbach:1997:INP

[3786] David A. Schwarzbach. Iran’s nuclear puzzle. Scientific American, 276(6):62–65, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
62.pdf;   http://www.sciam.com/0697issue/0697currentissue.html.

Villasenor:1997:CC

[3787] John Villasenor and William H. Mangione-Smith. Configurable computing. Scientific American, 276(6):66–71, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-66.pdf;   http://www.sciam.com/
1997/0607issue/0697villasenor.html.

Leakey:1997:EHF

[3788] Meave Leakey and Alan Walker. Early hominid fossils from Africa. Scientific American, 276(6):74–79, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
74.pdf;   http://www.sciam.com/0697issue/0697currentissue.html.

Pratson:1997:SPP

[3789] Lincoln F. Pratson and William F. Haxby. Science in pictures: Panoramas of the seafloor. Scientific American, 276(6):82–87, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-82.pdf.

Pratson:1997:PS

[3790] Lincoln F. Pratson and William F. Haxby. Panoramas of the seafloor. Scientific American, 276(6):82–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1997/0607issue/0697pratson.html.

Forsyth:1997:SDP

[3791] David Forsyth, Jitendra Malik, and Robert Wilensky. Searching for digital pictures. Scientific American, 276(6):88–93, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-88.pdf;   http:/
/www.sciam.com/0697issue/0697currentissue.html.

Friedmann:1997:SRO

[3792] Theodore Friedmann. Special report: Overcoming the obstacles to gene therapy. Scientific American, 276(6):95–101, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-95.pdf.

Anonymous:1997:MGT

[3793] Anonymous. Making gene therapy work. Scientific American, 276 (6):95–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/0697issue/
0697currentissue.html.

Felgner:1997:NSG

[3794] Philip L. Felgner. Nonviral strategies for gene therapy. Scientific American, 276(6):102–106, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
102.pdf.

Blaese:1997:GTC

[3795] R. Michael Blaese. Gene therapy for cancer. Scientific American, 276(6):111–115, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
111.pdf.

Ho:1997:GTN

[3796] Dora Y. Ho and Robert M. Sapolsky. Gene therapy for the nervous system. Scientific American, 276(6):116–120, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-116.pdf.

Mirsky:1997:WCM

[3797] Steve Mirsky and John Rennie. What cloning means for gene therapy. Scientific American, 276(6):121–123, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-121.pdf.

Yam:1997:TPB

[3798] Philip Yam. Trends in physics: Bringing Schrödinger’s Cat to life. Scientific American, 276(6):124–129, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
124.pdf.

Yam:1997:BSC

[3799] Philip Yam. Bringing Schrödinger’s cat to life. Scientific American, 276(6):124–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0697issue/0697currentissue.html.

Carlson:1997:ASGa

[3800] Shawn Carlson. The amateur scientist: Getting inside an ant’s head. Scientific American, 276(6):130–132, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-130.pdf.

Anonymous:1997:ASF

[3801] Anonymous. The amateur scientist: Fixing and mounting small insects for microscopic view. Scientific American, 276(6):130–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0697issue/0697currentissue.html.

Stewart:1997:MRSa

[3802] Ian Stewart. Mathematical recreations: The sifting sands of factorland. Scientific American, 276(6):134–137, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-134.pdf.

Anonymous:1997:MRS

[3803] Anonymous. Mathematical recreations: Sifting the sands of Factorland through a fine sieve. Scientific American, 276(6):134–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0697issue/0697currentissue.html.

Anonymous:1997:RCRb

[3804] Anonymous. Reviews and commentaries: Reviews and commentaries. Scientific American, 276(6):140–143, June 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v276/
n6/pdf/scientificamerican0697-140.pdf.

Anonymous:1997:RCN

[3805] Anonymous. Reviews and commentaries: The neurobiology of emotion; justifying the costs of science; cats illustrated. Scientific American, 276(6):140–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0697issue/0697currentissue.html.

Burke:1997:CGB

[3806] James Burke. Connections: From gaslight to B-29 bombers. Scientific American, 276(6):140–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0697issue/0697currentissue.html.

Morrison:1997:WPT

[3807] Philip Morrison. Wonders: Playful technology enwraps islands and coastlines. Scientific American, 276(6):140–??, June 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0697issue/0697currentissue.html.

Morrison:1997:WSA

[3808] Philip Morrison and Phylis Morrison. Wrapping up science and art. Scientific American, 276(6):144–146, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
144.pdf.

Burke:1997:ND

[3809] James Burke. Notice the difference? Scientific American, 276(6):145– 146, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v276/n6/pdf/scientificamerican0697-145.pdf.

Katz:1997:WKD

[3810] Saul N. Katz. Working knowledge: Decaffeinating coffee. Scientific American, 276(6):148, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v276/n6/pdf/scientificamerican0697-
148.pdf.

Anonymous:1997:WKS

[3811] Anonymous. Working knowledge: The supercritical extraction process of decaffeinating coffee. Scientific American, 276(6):148–??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0697issue/0697currentissue.html.

Anonymous:1997:SP

[3812] Anonymous. Science in pictures. Scientific American, 276(6):??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:SRb

[3813] Anonymous. Special report. Scientific American, 276(6):??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:TP

[3814] Anonymous. Trends in physics. Scientific American, 276(6):??, June 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:EFP

[3815] John Rennie. From the Editor: The future and past of China. Scientific American, 277(1):6, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
6.pdf.

Anonymous:1997:Ea

[3816] Anonymous. Errata. Scientific American, 277(1):8, July 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-8b.pdf.

Anonymous:1997:LEf

[3817] Anonymous. Letters to the Editors. Scientific American, 277(1):8, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-8a.pdf.

Anonymous:1997:YAIg

[3818] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 277(1):10, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
10.pdf.

Fluendy:1997:NAS

[3819] Simon Fluendy. News and analysis: Strategic investments. Scientific American, 277(1):15–16, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
15.pdf.

Anonymous:1997:NAF

[3820] Anonymous. News and analysis: In focus: China makes Hong Kong into a high-tech center, but scientists worry about repression. Scientific American, 277(1):15–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Yam:1997:RP

[3821] Philip Yam. Rights of passage. Scientific American, 277(1):16–18, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-16.pdf.

Horgan:1997:TS

[3822] John Horgan. Twist and shout. Scientific American, 277(1):19, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-19.pdf.

Anonymous:1997:SCD

[3823] Anonymous. Science and the citizen: Doubts on a directional universe Rogue parrots; earlier ancestor of humans and apes? Scientific American, 277(1):20–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Wong:1997:MM

[3824] Kate Wong. Moroto morass. Scientific American, 277(1):21–22, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-21.pdf.

Zorpette:1997:PP

[3825] Glenn Zorpette. Parrots and plunder. Scientific American, 277(1):22– 25, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-22b.pdf.

Leutwyler:1997:Bf

[3826] Kristin Leutwyler. In brief. Scientific American, 277(1):22–26, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-22a.pdf.

Mirsky:1997:ENT

[3827] Steve Mirsky. The Emperor’s new toilet paper. Scientific American, 277 (1):24, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-24.pdf.

Doyle:1997:AI

[3828] Rodger Doyle. Access to the Internet. Scientific American, 277(1): 26, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-26.pdf.

Leutwyler:1997:WWR

[3829] Kristin Leutwyler. What will really be. Scientific American, 277(1):28– 29, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-28.pdf.

Anonymous:1997:PMD

[3830] Anonymous. Profile: Michael L. Dertouzos of M.I.T. embraces poets and programmers. Scientific American, 277(1):28–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0797issue/0797currentissue.html.

Werner:1997:DS

[3831] Louis Werner. Dam safety. Scientific American, 277(1):31–32, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-31.pdf.

Anonymous:1997:TBF

[3832] Anonymous. Technology and business: Floating tunnels; computer border guards; lasers against angina; rising fears at Aswan. Scientific American, 277(1):31–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Stix:1997:TV

[3833] Gary Stix. Tunnel visions. Scientific American, 277(1):32, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-32.pdf.

Gibbs:1997:IGS

[3834] W. Wayt Gibbs. The infinitesimal gets smaller. Scientific American, 277 (1):34, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-34b.pdf.

Gibbs:1997:HH

[3835] W. Wayt Gibbs. Helping heartache. Scientific American, 277(1):34– 36, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-34a.pdf.

Stix:1997:MM

[3836] Gary Stix. Mug machine. Scientific American, 277(1):36, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-36.pdf.

Lewis:1997:WBC

[3837] Ted Lewis. www.batmobile.car. Scientific American, 277(1):38, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-38.pdf.

Anonymous:1997:CVW

[3838] Anonymous. Cyber view: The Web on wheels. Scientific American, 277(1):38–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Agnew:1997:CBT

[3839] Neville Agnew and Fan Jinshi. China’s Buddhist treasures at Dunhuang. Scientific American, 277(1):40–45, July 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-40.pdf;   http://www.sciam.com/
0797issue/0797currentissue.html.

Lanza:1997:X

[3840] Robert P. Lanza, David K. C. Cooper, and William L. Chick. Xenotransplantation. Scientific American, 277(1):54–59, July 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-54.pdf;   http://www.sciam.com/
0797issue/0797currentissue.html.

Doyle:1997:SFF

[3841] Brian E. Doyle. Strong fabrics for fast sails. Scientific American, 277(1): 60–67, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-60.pdf;   http:/
/www.sciam.com/0797issue/0797currentissue.html.

Fishman:1997:GRB

[3842] Gerald J. Fishman and Dieter H. Hartmann. Gamma-ray bursts. Scientific American, 277(1):46–51, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-

46.pdf;   http://www.sciam.com/0797issue/0797currentissue.html;
 http://www.sciam.com/1997/0707issue/0797fishman.html.

Alleman:1997:AR

[3843] James E. Alleman and Brooke T. Mossman. Asbestos revisited. Scientific American, 277(1):70–75, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
70.pdf;   http://www.sciam.com/0797issue/0797currentissue.html.

Smil:1997:GPN

[3844] Vaclav Smil. Global population and the nitrogen cycle. Scientific American, 277(1):76–81, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
76.pdf;   http://www.sciam.com/0797issue/0797currentissue.html.

Gibbs:1997:TCT

[3845] W. Wayt Gibbs. Trends in computing: Taking computers to task. Scientific American, 277(1):82–89, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
82.pdf;   http://www.sciam.com/1997/0707issue/0797trends.html.

Carlson:1997:ASH

[3846] Shawn Carlson. The amateur scientist: How-to’s of butterfly rookeries. Scientific American, 277(1):90–92, July 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-90.pdf.

Anonymous:1997:ASCb

[3847] Anonymous. The amateur scientist: Catching, raising and collecting butterflies. Scientific American, 277(1):90–??, July 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0797issue/0797currentissue.html.

Stewart:1997:MRSb

[3848] Ian Stewart. Mathematical recreations: Squaring the square. Scientific American, 277(1):94–96, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
94.pdf.

Anonymous:1997:MRTb

[3849] Anonymous. Mathematical recreations: Tiling a square, a rectangle or a Möbius strip. Scientific American, 277(1):94–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0797issue/0797currentissue.html.

Anonymous:1997:RCd

[3850] Anonymous. Reviews and commentaries. Scientific American, 277(1):98– 101, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n1/pdf/scientificamerican0797-98.pdf.

Anonymous:1997:RCTb

[3851] Anonymous. Reviews and commentaries: Tom Stoppard’s Arcadia and chaotic love; submarine reporting; the NASA atlas of the solar system. Scientific American, 277(1):98–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Burke:1997:CBF

[3852] James Burke. Connections: Of Ben Franklin, galvanic frogs and the antimalaria machine. Scientific American, 277(1):98–??, July 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0797issue/0797currentissue.html.

Morrison:1997:WMP

[3853] Philip Morrison. Wonders: 20,000 megabits per second under the sea. Scientific American, 277(1):98–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Morrison:1997:BG

[3854] Philip Morrison and Phylis Morrison. Bandwidth galore. Scientific American, 277(1):102–104, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n1/pdf/scientificamerican0797-
102.pdf.

Burke:1997:CS

[3855] James Burke. Cool stuff. Scientific American, 277(1):103–104, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-103.pdf.

Matthews:1997:WKG

[3856] Mike Matthews and Robert Myer. Working knowledge: Guitareffects pedals. Scientific American, 277(1):105, July 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n1/pdf/scientificamerican0797-105.pdf.

Anonymous:1997:WKHa

[3857] Anonymous. Working knowledge: How my guitar gently weeps. Scientific American, 277(1):105–??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0797issue/0797currentissue.html.

Anonymous:1997:TC

[3858] Anonymous. Trends in computing. Scientific American, 277(1):??, July 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:ECE

[3859] John Rennie. From the Editor: Current events. Scientific American, 277(2):6, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
6.pdf.

Anonymous:1997:LEg

[3860] Anonymous. Letters to the Editors. Scientific American, 277(2):8, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-8.pdf.

Anonymous:1997:YAIh

[3861] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 277(2):10, August 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n2/pdf/scientificamerican0897-10.pdf.

Beardsley:1997:NASb

[3862] Tim Beardsley. News and analysis: Science’s survival strategy. Scientific American, 277(2):13–16, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
13.pdf.

Anonymous:1997:FFC

[3863] Anonymous. In focus: As Federal cutbacks squeeze budgets, scientists set priorities. Scientific American, 277(2):13–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0897issue/0897currentissue.html.

Schneider:1997:BN

[3864] David Schneider. A blue note. Scientific American, 277(2):18, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-18.pdf.

Anonymous:1997:SCSb

[3865] Anonymous. Science and the citizen: Slushballs from space; music of the deep; turning on soil; biodiversity be dammed; jungle gym science. Scientific American, 277(2):18–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0897issue/0897currentissue.html.

Powell:1997:RSS

[3866] Corey S. Powell. Return of the space snowballs. Scientific American, 277(2):19–20, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
19a.pdf.

Leutwyler:1997:Bg

[3867] Kristin Leutwyler. In brief. Scientific American, 277(2):19–22, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-19b.pdf.

Nemecek:1997:FMD

[3868] Sasha Nemecek. Frankly, my dear, I don’t want a dam. Scientific American, 277(2):20–22, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
20.pdf.

Stix:1997:PTS

[3869] Gary Stix. Play time and space. Scientific American, 277(2):22–26, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-22.pdf.

Mirsky:1997:SI

[3870] Steve Mirsky. Space invaders. Scientific American, 277(2):24, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-24.pdf.

Doyle:1997:PRU

[3871] Rodger Doyle. Plants at risk in the U.S. Scientific American, 277(2):26, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-26.pdf.

Stix:1997:DPB

[3872] Gary Stix. Dark prophet of biogenetics. Scientific American, 277(2):28– 32, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-28.pdf.

Anonymous:1997:PAC

[3873] Anonymous. Profile: The activist career of Jeremy Rifkin, nemesis of biotechnology. Scientific American, 277(2):28–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0897issue/0897currentissue.html.

Gibbs:1997:CCa

[3874] W. Wayt Gibbs. Command and control. Scientific American, 277(2):33– 34, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-33.pdf.

Anonymous:1997:TBSa

[3875] Anonymous. Technology and business: Swifter, smarter software delivery cancer cells get a bad cold; microgears from powder. Scientific American, 277(2):33–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0897issue/0897currentissue.html.

Beardsley:1997:MLW

[3876] Tim Beardsley. Making light work. Scientific American, 277(2):34, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-34.pdf.

Zacks:1997:GDD

[3877] Rebecca Zacks. Getting the dirt on dirt. Scientific American, 277(2):35, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-35b.pdf.

DeKoker:1997:FL

[3878] Brenda DeKoker. Farming with lint. Scientific American, 277(2):35–36, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-35a.pdf.

Gibbs:1997:CCb

[3879] W. Wayt Gibbs. A cold for cancer. Scientific American, 277(2):36–37, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-36.pdf.

Wallich:1997:PDA

[3880] Paul Wallich. Parental discretion advised. Scientific American, 277(2): 38, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-38.pdf.

Anonymous:1997:CVDb

[3881] Anonymous. Cyber view: Dirty tricks of antiporn add-ons to Web browsers. Scientific American, 277(2):38–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0897issue/0897currentissue.html.

Wallace:1997:MDA

[3882] Douglas C. Wallace. Mitochondrial DNA in aging and disease. Scientific American, 277(2):40–47, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
40.pdf;   http://www.sciam.com/0897issue/0897currentissue.html.

Diels:1997:IESa

[3883] Jean-Claude Diels, Ralph Bernstein, Karl E. Stahlkopf, and Xin Miao Zhao. Investigating electricity in the sky: Lightning control with lasers. Scientific American, 277(2):50–55, August 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n2/pdf/scientificamerican0897-50.pdf.

Diels:1997:LCL

[3884] J.-C. Diels, R. Bernstein, K. E. Stahlkopf, and Xin Miao Zhao. Lightning control with lasers. Scientific American, 277(2):50–55, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0897issue/0897currentissue.html.

Mende:1997:LBE

[3885] Stephen B. Mende, Davis D. Sentman, and Eugene M. Wescott. Lightning between Earth and Space. Scientific American, 277(2):56–59, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-56.pdf;   http:/
/www.sciam.com/1997/0807issue/0897mende.html.

El-Baz:1997:SAA

[3886] Farouk El-Baz. Space age archaeology. Scientific American, 277(2):60– 65, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-60.pdf;   http:/
/www.sciam.com/0897issue/0897currentissue.html.

Gwynne:1997:GG

[3887] Darryl T. Gwynne. Glandular gifts. Scientific American, 277(2):66–71, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-66.pdf;   http:/
/www.sciam.com/0897issue/0897currentissue.html.

Gorelik:1997:TSL

[3888] Gennady Gorelik. The top-secret life of Lev Landau. Scientific American, 277(2):72–77, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
72.pdf;   http://www.sciam.com/0897issue/0897currentissue.html.

Anonymous:1997:TNM

[3889] Anonymous. Trends in neuroscience: The machinery of thought. Scientific American, 277(2):78–83, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
78.pdf.

Carlson:1997:ASGb

[3890] Shawn Carlson. The amateur scientist: Getting a charge out of rain. Scientific American, 277(2):84–85, August 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n2/pdf/scientificamerican0897-84.pdf.

Anonymous:1997:ASMa

[3891] Anonymous. The amateur scientist: Measuring the electrical charge on raindrops. Scientific American, 277(2):84–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0897issue/0897currentissue.html.

Stewart:1997:MREa

[3892] Ian Stewart. Mathematical recreations: Empires on the Moon. Scientific American, 277(2):86–88, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
86.pdf.

Anonymous:1997:MRB

[3893] Anonymous. Mathematical recreations: Blue moon: coloring maps on multiple planets. Scientific American, 277(2):86–??, August 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0897issue/0897currentissue.html.

Anonymous:1997:RCe

[3894] Anonymous. Reviews and commentaries. Scientific American, 277(2):89– 92, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-89.pdf.

Anonymous:1997:RCVb

[3895] Anonymous. Reviews and commentaries: Virtual reality re-creates ancient Rome, the pyramids, imperial China and a Los Angeles that never was. Scientific American, 277(2):89–??, August 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0897issue/0897currentissue.html.

Burke:1997:CDZ

[3896] James Burke. Connections: Of divine zoos, riveting ships and Morse code. Scientific American, 277(2):89–??, August 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0897issue/0897currentissue.html.

Morrison:1997:WEA

[3897] Philip Morrison. Wonders: The earliest atom splitters. Scientific American, 277(2):89–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0897issue/0897currentissue.html.

Morrison:1997:SC

[3898] Philip Morrison. 1997: Subatomic centenary. Scientific American, 277(2):93–95, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
93.pdf.

Burke:1997:LM

[3899] James Burke. Lucky he missed. Scientific American, 277(2):94–95, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n2/pdf/scientificamerican0897-94.pdf.

MacChesney:1997:WKO

[3900] John MacChesney. Working knowledge: Optical fibers. Scientific American, 277(2):96, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n2/pdf/scientificamerican0897-
96.pdf.

Anonymous:1997:WKHb

[3901] Anonymous. Working knowledge: How fiber-optic cables carry light. Scientific American, 277(2):96–??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0897issue/0897currentissue.html.

Anonymous:1997:TN

[3902] Anonymous. Trends in neuroscience. Scientific American, 277(2):??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1997:MT

[3903] Tim Beardsley. The machinery of thought. Scientific American, 277 (2):??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1997/0807issue/
0897trends.html.

Diels:1997:IESb

[3904] Jean-Claude Diels, Ralph Bernstein, Karl E. Stahlkopf, and Xin Miao Zhao. Investigating electricity in the sky. Scientific American, 277(2): ??, August 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:EMH

[3905] John Rennie. From the Editor: Making (up) history. Scientific American, 277(3):6, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
6.pdf.

Anonymous:1997:LEh

[3906] Anonymous. Letters to the Editors. Scientific American, 277(3):8, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-8.pdf.

Anonymous:1997:YAIi

[3907] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 277(3):12, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n3/pdf/scientificamerican0997-12.pdf.

Beardsley:1997:NAE

[3908] Tim Beardsley. News and analysis: Evolution evolving. Scientific American, 277(3):15–18, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
15.pdf.

Anonymous:1997:FBM

[3909] Anonymous. In focus: Bacteria mutating “deliberately” once again pit Darwin against Lamarck. Scientific American, 277(3):15–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0997issue/0997currentissue.
html.

Zacks:1997:LA

[3910] Rebecca Zacks. Looking for alternatives. Scientific American, 277(3):20, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-20.pdf.

Anonymous:1997:SCW

[3911] Anonymous. Science and the citizen: Was this chemist’s death necessary?… caveman music. … shocking stars. … birds in peril. … volunteering under pressure. Scientific American, 277(3):20–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/0997issue/0997currentissue.
html.

Powell:1997:CB

[3912] Corey S. Powell. Crash and burn. Scientific American, 277(3):22, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-22a.pdf.

Zorpette:1997:FB

[3913] Glenn Zorpette. Flying and the bends. Scientific American, 277(3):22–24, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
22b.pdf.

Editors:1997:JYC

[3914] The Editors. Jacques-Yves Cousteau, 1910–1997. Scientific American, 277(3):24, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
24.pdf.

Mirsky:1997:AG

[3915] Steve Mirsky. An axis to grind. Scientific American, 277(3):26, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-26b.pdf.

Leutwyler:1997:Bh

[3916] Kristin Leutwyler. In brief. Scientific American, 277(3):26–28, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-26a.pdf.

Wong:1997:NN

[3917] Kate Wong. Neanderthal notes. Scientific American, 277(3):28–30, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-28.pdf.

Doyle:1997:TB

[3918] Rodger Doyle. Threatened birds. Scientific American, 277(3):30, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-30.pdf.

Mukerjee:1997:WBW

[3919] Madhusree Mukerjee. A whistle-blower’s wars. Scientific American, 277(3):32–33, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
32.pdf.

Anonymous:1997:PES

[3920] Anonymous. Profile: In exile with science: whistle-blower Jan Moor-Jankowski. Scientific American, 277(3):32–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0997issue/0997currentissue.html.

Schneider:1997:BRC

[3921] David Schneider. Buck Rogers, CEO. Scientific American, 277(3):34–36, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-34.pdf.

Anonymous:1997:TBSb

[3922] Anonymous. Technology and business: Shoot for the moon, then sell it… new blue lasers… hormone treatments …. micromechanical radios. Scientific American, 277(3):34–??, September 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0997issue/0997currentissue.html.

Zorpette:1997:TB

[3923] Glenn Zorpette. True blue. Scientific American, 277(3):36, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-36.pdf.

Nemecek:1997:HH

[3924] Sasha Nemecek. Hold the hormones? Scientific American, 277(3):38–40, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-38.pdf.

Stix:1997:BI

[3925] Gary Stix. Beam it up. Scientific American, 277(3):40–41, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-40.pdf.

Gibbs:1997:F

[3926] W. Wayt Gibbs. Not so fast. Scientific American, 277(3):42, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-42.pdf.

Anonymous:1997:CVAa

[3927] Anonymous. Cyber view: Automated highways may outsmart themselves. Scientific American, 277(3):42–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0997issue/0997currentissue.html.

OBrien:1997:SAR

[3928] Stephen J. O’Brien and Michael Dean. In search of AIDS-Resistance genes. Scientific American, 277(3):44–51, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n3/pdf/scientificamerican0997-44.pdf;   http://www.sciam.com/
1997/0907issue/0997obrien.html.

Liss:1997:DTQ

[3929] Tony M. Liss and Paul L. Tipton. The discovery of the top quark. Scientific American, 277(3):54–59, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n3/pdf/scientificamerican0997-54.pdf;   http://www.sciam.com/
1997/0907issue/0997tipton.html.

Bayley:1997:BDCa

[3930] Hagan Bayley. Building doors into cells. Scientific American, 277(3):62–67, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
62.pdf.

Glasheen:1997:RW

[3931] James W. Glasheen and Thomas A. McMahon. Running on water. Scientific American, 277(3):68–69, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n3/pdf/scientificamerican0997-68.pdf;   http://www.sciam.com/
0997issue/0997currentissue.html.

Loftus:1997:CFM

[3932] Elizabeth F. Loftus. Creating false memories. Scientific American, 277(3):70–75, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
70.pdf;   http://www.sciam.com/0997issue/0997currentissue.html.

Smith:1997:LPA

[3933] Michael E. Smith. Life in the provinces of the Aztec Empire. Scientific American, 277(3):76–83, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
76.pdf;   http://www.sciam.com/0997issue/0997currentissue.html.

Nori:1997:BS

[3934] Franco Nori, Paul Sholtz, and Michael Bretz. Booming sand. Scientific American, 277(3):84–89, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
84.pdf;   http://www.sciam.com/0997issue/0997currentissue.html.

Carlson:1997:ASU

[3935] Shawn Carlson. The amateur scientist: Unraveling the secrets of monarchs. Scientific American, 277(3):90–91, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n3/pdf/scientificamerican0997-90.pdf.

Anonymous:1997:ASMb

[3936] Anonymous. The amateur scientist: Monitoring the monarchs. Scientific American, 277(3):90–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0997issue/0997currentissue.html.

Stewart:1997:MREb

[3937] Ian Stewart. Mathematical recreations: Empires and electronics. Scientific American, 277(3):92–94, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n3/pdf/scientificamerican0997-92.pdf.

Anonymous:1997:MRE

[3938] Anonymous. Mathematical recreations: Emperors of electronics graph a circuit board. Scientific American, 277(3):92–??, September 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0997issue/0997currentissue.html.

Anonymous:1997:RCg

[3939] Anonymous. Reviews and commentaries. Scientific American, 277(3):95–98, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
95.pdf.

Anonymous:1997:RCSa

[3940] Anonymous. Reviews and commentaries: Still debating what killed the dinosaurs. . . . . . . make a joyful noise: music and the mind. . . . . . creature feature. Scientific American, 277(3):95–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0997issue/0997currentissue.html.

Burke:1997:CLSa

[3941] James Burke. Connections: The long, sticky road to headache relief. Scientific American, 277(3):95–??, September 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0997issue/0997currentissue.html.

Morrison:1997:WST

[3942] Philip Morrison. Wonders: Seeing through trickery’s illusions. Scientific American, 277(3):95–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0997issue/0997currentissue.html.

Morrison:1997:I

[3943] Philip Morrison and Phylis Morrison. Illusions. Scientific American, 277(3):99–101, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
99.pdf.

Burke:1997:TTA

[3944] James Burke. Take two acronyms. Scientific American, 277(3):100–101, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n3/pdf/scientificamerican0997-100.pdf.

Nathan:1997:WKB

[3945] Alan M. Nathan. Working knowledge: Baseball pitches. Scientific American, 277(3):102–103, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n3/pdf/scientificamerican0997-
102.pdf.

Anonymous:1997:WKA

[3946] Anonymous. Working knowledge: Aerodynamics on the mound. Scientific American, 277(3):102–??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/0997issue/0997currentissue.html.

Bayley:1997:BDCb

[3947] Hagan Bayley. Building doors into cells. Scientific American, 277(3):??, September 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1997:EWG

[3948] John Rennie. From the Editors: The way to go. Scientific American, 277(4):6, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
6.pdf.

Anonymous:1997:LEi

[3949] Anonymous. Letters to the Editors. Scientific American, 277(4):10, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-10.pdf.

Anonymous:1997:YAY

[3950] Anonymous. 50, 100 and 150 years ago: 50, 100 and 150 years ago. Scientific American, 277(4):12, October 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-12.pdf.

Stix:1997:NAG

[3951] Gary Stix. News and analysis: Growing a new field. Scientific American, 277(4):15–16, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
15.pdf.

Drollette:1997:NH

[3952] Dan Drollette. The next hop. Scientific American, 277(4):20, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-20.pdf.

Holloway:1997:FS

[3953] Marguerite Holloway. Field and stream. Scientific American, 277(4):24– 28, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-24b.pdf.

Leutwyler:1997:Bi

[3954] Kristin Leutwyler. In brief. Scientific American, 277(4):24–32, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-24a.pdf.

Nemecek:1997:GKW

[3955] Sasha Nemecek. Gotta know when to fold ’em. Scientific American, 277(4):28–32, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
28.pdf.

Powell:1997:SC

[3956] Corey S. Powell. Science in court. Scientific American, 277(4):32–34, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-32.pdf.

Zacks:1997:WTT

[3957] Rebecca Zacks. What are they thinking? Scientific American, 277(4):34, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-34.pdf.

Mirsky:1997:SS

[3958] Steve Mirsky. He shoots, he scars. Scientific American, 277(4):38, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-38b.pdf.

Beardsley:1997:MM

[3959] Tim Beardsley. Matter over mind. Scientific American, 277(4):38–40, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-38a.pdf.

Doyle:1997:COP

[3960] Rodger Doyle. Chronic obstructive pulmonary disease. Scientific American, 277(4):40, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
40.pdf.

Holloway:1997:GFP

[3961] Marguerite Holloway. Gombe’s famous primate. Scientific American, 277(4):42–44, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
42.pdf.

Gibbs:1997:CW

[3962] W. Wayt Gibbs. Change in the wind. Scientific American, 277(4):46, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-46a.pdf.

Gibbs:1997:HMM

[3963] W. Wayt Gibbs. Heavy metal meets its match. Scientific American, 277(4):46–48, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
46b.pdf.

Gibbs:1997:CM

[3964] W. Wayt Gibbs. Charging to market. Scientific American, 277(4):48–49, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-48.pdf.

Gibbs:1997:SS

[3965] W. Wayt Gibbs. A sense of synesthesia. Scientific American, 277(4):49, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-49.pdf.

Grossman:1997:MYD

[3966] Wendy M. Grossman. Master of your domain. Scientific American, 277(4):52, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
52.pdf.

Gibbs:1997:TPP

[3967] W. Wayt Gibbs. Transportation’s perennial problems. Scientific American, 277(4):54–57, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
54.pdf;   http://www.sciam.com/1997/1007issue/1097gibbs.html.

Schafer:1997:PFG

[3968] Andreas Schafer and David Victor. The past and future of global mobility. Scientific American, 277(4):58–61, October 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-58.pdf;   http://www.sciam.com/
1997/1007issue/1097schafer.html.

Rennie:1997:VWN

[3969] John Rennie. 13 vehicles that went nowhere. Scientific American, 277(4):64–67, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
64.pdf;   http://www.sciam.com/1997/1007issue/1097rennie.html.

Wouk:1997:HEV

[3970] Victor Wouk. Hybrid electric vehicles. Scientific American, 277(4):70–74, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-70.pdf;   http:/
/www.sciam.com/1997/1007issue/1097wouk.html.

Rosen:1997:FHV

[3971] Harold A. Rosen and Deborah R. Castleman. Flywheels in hybrid vehicles. Scientific American, 277(4):75–77, October 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-75.pdf;   http://www.sciam.com/
1997/1007issue/1097rosen.html.

Rillings:1997:AH

[3972] James H. Rillings. Automated highways. Scientific American, 277(4):80– 85, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-80.pdf;   http:/
/www.sciam.com/1997/1007issue/1097rillings.html.

Howard:1997:UTC

[3973] Kenneth R. Howard. Unjamming traffic with computers. Scientific American, 277(4):86–88, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
86.pdf;   http://www.sciam.com/1997/1007issue/1097howard.html.

Mokhtarian:1997:NTC

[3974] Patricia L. Mokhtarian. Now that travel can be virtual, will congestion virtually disappear? Scientific American, 277(4):93, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-93.pdf;   http:/
/www.sciam.com/1997/1007issue/1097mokhtarian.html.

Stix:1997:DM

[3975] Gary Stix. Driving to Mach 1. Scientific American, 277(4):94–97, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-94.pdf;   http:/
/www.sciam.com/1997/1007issue/1097stix_car.html.

Leutwyler:1997:SVN

[3976] Kristin Leutwyler. Speed versus need. Scientific American, 277(4):98–99, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-98.pdf;   http:/
/www.sciam.com/1997/1007issue/1097leutwyler.html.

Raoul:1997:HHS

[3977] Jean-Claude Raoul. How high-speed trains make tracks. Scientific American, 277(4):100–105, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
100.pdf;   http://www.sciam.com/1997/1007issue/1097raoul.html.

Perl:1997:FTW

[3978] Anthony Perl and James A. Dunn, Jr. Fast trains: Why the U.S. lags. Scientific American, 277(4):106–108, October 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-106.pdf;   http://www.sciam.com/
1997/1007issue/1097perl.html.

Stix:1997:MRO

[3979] Gary Stix. Maglev: Racing to oblivion? Scientific American, 277(4):109, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-109.pdf;   http:
//www.sciam.com/1997/1007issue/1097stix_maglev.html.

Mark:1997:SB

[3980] Hans Mark. Straight up into the blue. Scientific American, 277(4):110–115, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
110.pdf;   http://www.sciam.com/1997/1007issue/1097mark.html.

Carlson:1997:LI

[3981] Shawn Carlson. The lure of Icarus. Scientific American, 277(4):116–119, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-116.pdf;   http:
//www.sciam.com/1997/1007issue/1097carlson.html.

Mattingly:1997:SRS

[3982] T. K. Mattingly. A simpler ride into space. Scientific American, 277(4):120–125, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
120.pdf;   http://www.sciam.com/1997/1007issue/1097mattingly.
html.

Giles:1997:FSF

[3983] David L. Giles. Faster ships for the future. Scientific American, 277(4):126–131, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
126.pdf;   http://www.sciam.com/1997/1007issue/1097giles.html.

Hawkes:1997:MGS

[3984] Graham S. Hawkes. Microsubs go to sea. Scientific American, 277(4):132–135, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
132.pdf;   http://www.sciam.com/1997/1007issue/1097hawkes.html.

Lacob:1997:EM

[3985] Miriam Lacob. Elevators on the move. Scientific American, 277(4):136–137, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
136.pdf;   http://www.sciam.com/1997/1007issue/1097lacob.html.

Carlson:1997:ASR

[3986] Shawn Carlson. The amateur scientist: Recording the sounds of life. Scientific American, 277(4):138–139, October 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-138.pdf.

Stewart:1997:MRT

[3987] Ian Stewart. Mathematical recreations: Two-way jigsaw puzzles. Scientific American, 277(4):140–145, October 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-140.pdf.

Boellstorff:1997:RCQ

[3988] Tom Boellstorff and Lawrence Cohen. Reviews and commentaries: Queer science indeed. Scientific American, 277(4):146–148, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n4/pdf/scientificamerican1097-146.pdf.

Morrison:1997:AC

[3989] Philip Morrison. Air-cooled. Scientific American, 277(4):149–150, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-149.pdf.

Burke:1997:BSH

[3990] James Burke. The buck stops here. Scientific American, 277(4):151–152, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-151.pdf.

Yam:1997:CG

[3991] Philip Yam. Clean genes. Scientific American, 277(4):153–154, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n4/pdf/scientificamerican1097-153.pdf.

Rainey:1997:WKF

[3992] William S. Rainey. Working knowledge: Fish ladders. Scientific American, 277(4):156, October 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n4/pdf/scientificamerican1097-
156.pdf.

Rennie:1997:EPI

[3993] John Rennie. From the Editors: Prove it. Scientific American, 277(5):8, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-8.pdf.

Anonymous:1997:LEj

[3994] Anonymous. Letters to the Editors. Scientific American, 277(5):10, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-10.pdf.

Anonymous:1997:YAIj

[3995] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 277(5):12, November 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n5/pdf/scientificamerican1197-12.pdf.

Beardsley:1997:NAD

[3996] Tim Beardsley. News and analysis: Death in the deep. Scientific American, 277(5):17–20, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
17.pdf.

Anonymous:1997:FDZ

[3997] Anonymous. In focus: The dead zone: an expanse of the Gulf of Mexico is weirdly barren. Scientific American, 277(5):17–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1197issue/1197currentissue.html.

Mukerjee:1997:GS

[3998] Madhusree Mukerjee. Girth of a star. Scientific American, 277(5):22, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-22.pdf.

Anonymous:1997:SCE

[3999] Anonymous. Science and the citizen: Eat (and sequence) your vegetables… sizing up a neutron star. … the evil weevil… swap two Darwins for an Einstein. Scientific American, 277(5):22–??, November 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1197issue/1197currentissue.html.

Beardsley:1997:FGP

[4000] Tim Beardsley. The Food Genome Project. Scientific American, 277(5):24, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
24b.pdf.

Leutwyler:1997:Bj

[4001] Kristin Leutwyler. In brief. Scientific American, 277(5):24–32, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-24a.pdf.

Mirsky:1997:BP

[4002] Steve Mirsky. The big picture. Scientific American, 277(5):28, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-28.pdf.

Zorpette:1997:RD

[4003] Glenn Zorpette. A real dive. Scientific American, 277(5):32–36, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-32.pdf.

Beardsley:1997:BN

[4004] Tim Beardsley. Biological noncontrol. Scientific American, 277(5):36–38, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-36.pdf.

Doyle:1997:ASD

[4005] Rodger Doyle. Access to safe drinking water. Scientific American, 277(5):38, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
38.pdf.

Nemecek:1997:ROL

[4006] Sasha Nemecek. Rescuing the ozone layer. Scientific American, 277(5):40–43, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
40.pdf.

Anonymous:1997:PNC

[4007] Anonymous. Profile: Nobel chemist Mario Molina still faces skeptics over CFCs and ozone loss. Scientific American, 277(5):40–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1197issue/1197currentissue.html.

Gibbs:1997:P

[4008] W. Wayt Gibbs. Plantibodies. Scientific American, 277(5):44, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-44.pdf.

Anonymous:1997:TBA

[4009] Anonymous. Technology and business: Antibodies by the bushel… tracking underground oil. … rampaging robots. Scientific American, 277(5):44–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1197issue/1197currentissue.html.

Gibbs:1997:CCc

[4010] W. Wayt Gibbs. From chips to cubes. Scientific American, 277(5):45, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-45a.pdf.

Martineau:1997:OD

[4011] Kimberly Martineau. Oil in 4-D. Scientific American, 277(5):45–47, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-45b.pdf.

Holloway:1997:PND

[4012] Marguerite Holloway. “please, no double-sticky tape”. Scientific American, 277(5):47–48, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
47.pdf.

Gibbs:1997:B

[4013] W. Wayt Gibbs. Boom. Scientific American, 277(5):48, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n5/pdf/scientificamerican1197-48.pdf.

Lewis:1997:WDN

[4014] Ted Lewis. We don’t need no regulation. Scientific American, 277(5):52, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-52.pdf.

Anonymous:1997:CVRb

[4015] Anonymous. Cyber view: Regulating the privacy of Internet commerce. Scientific American, 277(5):52–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1197issue/1197currentissue.html.

Nelson:1997:MFP

[4016] Robert M. Nelson. Mercury: The forgotten planet. Scientific American, 277(5):54–67, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
54.pdf;   http://www.sciam.com/1197issue/1197currentissue.html.

Singh:1997:FLS

[4017] Simon Singh and Kenneth A. Ribet. Fermat’s last stand. Scientific American, 277(5):68–73, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
68.pdf;   http://www.sciam.com/1197issue/1197currentissue.html.

Blair:1997:TNW

[4018] Bruce G. Blair, Harold A. Feiveson, and Frank N. von Hippel. Taking nuclear weapons off hair-trigger alert. Scientific American, 277(5):74–81, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-74.pdf;   http:/
/www.sciam.com/1997/1107issue/1197vonhippel.html.

Beckage:1997:PWS

[4019] Nancy E. Beckage. The parasitic wasp’s secret weapon. Scientific American, 277(5):82–87, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
82.pdf;   http://www.sciam.com/1197issue/1197currentissue.html.

Kephart:1997:FCV

[4020] Jeffrey O. Kephart, Gregory B. Sorkin, David M. Chess, and Steve R. White. Fighting computer viruses. Scientific American, 277(5):88–93, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-88.pdf;   http:/
/www.sciam.com/1997/1107issue/1197kephart.html.

Ndoro:1997:GZ

[4021] Webber Ndoro. Great Zimbabwe. Scientific American, 277(5):94–99, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-94.pdf;   http:/
/www.sciam.com/1197issue/1197currentissue.html.

Ronald:1997:MRD

[4022] Pamela C. Ronald. Making rice disease-resistant. Scientific American, 277(5):100–105, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
100.pdf;   http://www.sciam.com/1197issue/1197currentissue.
html.

Carlson:1997:ASCb

[4023] Shawn Carlson. The amateur scientist: Caught in a wind tunnel. Scientific American, 277(5):106–109, November 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n5/pdf/scientificamerican1197-106.pdf.

Anonymous:1997:ASMc

[4024] Anonymous. The amateur scientist: Make your own wind tunnel. Scientific American, 277(5):106–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1197issue/1197currentissue.html.

Stewart:1997:MRL

[4025] Ian Stewart. Mathematical recreations: The lore and lure of dice. Scientific American, 277(5):110–113, November 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n5/pdf/scientificamerican1197-110.pdf.

Anonymous:1997:MRDb

[4026] Anonymous. Mathematical recreations: Dicey odds when shooting craps. Scientific American, 277(5):110–??, November 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1197issue/1197currentissue.html.

Anonymous:1997:RCh

[4027] Anonymous. Reviews and commentaries. Scientific American, 277 (5):114–119, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n5/pdf/scientificamerican1197-
114.pdf.

Anonymous:1997:RCW

[4028] Anonymous. Reviews and commentaries: Where postmodernist critics and pathological scientists go wrong… sudden infant death and murder. Scientific American, 277(5):114–??, November 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1197issue/1197currentissue.html.

Burke:1997:CLSb

[4029] James Burke. Connections: Lilac statistics and the angel of mercy. Scientific American, 277(5):114–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1197issue/1197currentissue.html.

Gingerich:1997:WHV

[4030] Owen Gingerich. Wonders: The high value of magnificent fakes. Scientific American, 277(5):114–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1197issue/1197currentissue.html.

Gingerich:1997:PF

[4031] Owen Gingerich. In praise of fakes. Scientific American, 277(5):120–121, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-120.pdf.

Burke:1997:HB

[4032] James Burke. Healthy blooms. Scientific American, 277(5):122–123, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-122.pdf.

Musa:1997:WK

[4033] Samuel Musa. Working knowledge. Scientific American, 277(5):124, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n5/pdf/scientificamerican1197-124.pdf.

Anonymous:1997:WKL

[4034] Anonymous. Working knowledge: Liquid crystals on display. Scientific American, 277(5):124–??, November 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1197issue/1197currentissue.html.

Rennie:1997:EBE

[4035] John Rennie. From the Editors: Building excitement. Scientific American, 277(6):6, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
6.pdf.

Anonymous:1997:Eb

[4036] Anonymous. Errata. Scientific American, 277(6):8, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n6/pdf/scientificamerican1297-8b.pdf.

Anonymous:1997:LEk

[4037] Anonymous. Letters to the Editors. Scientific American, 277(6):8, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-8a.pdf.

Anonymous:1997:YAIk

[4038] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 277(6):10, December 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n6/pdf/scientificamerican1297-10.pdf.

Beardsley:1997:NAB

[4039] Tim Beardsley. News and analysis: The big shrink. Scientific American, 277(6):15–16, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
15.pdf.

Anonymous:1997:FCC

[4040] Anonymous. In focus: Chipmakers cast ultraviolet in a new light — and other tricks. Scientific American, 277(6):15–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1297issue/1297currentissue.html.

Vergoth:1997:NBA

[4041] Karin Vergoth. No bones about it. Scientific American, 277(6):18, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-18.pdf.

Anonymous:1997:SCSc

[4042] Anonymous. Science and the citizen: For sale: T. rex, slightly used… cell phone confusion… burning biomass and bacteria… the universe shows its age. Scientific American, 277(6):18–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1297issue/1297currentissue.html.

Beardsley:1997:SAa

[4043] Tim Beardsley. Say that again? Scientific American, 277(6):20, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-20b.pdf.

Leutwyler:1997:Bk

[4044] Kristin Leutwyler. In brief. Scientific American, 277(6):20–24, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-20a.pdf.

Mirsky:1997:FSF

[4045] Steve Mirsky. Full of sound and furry. Scientific American, 277(6):22, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-22.pdf.

Beardsley:1997:SAb

[4046] Tim Beardsley. Smoke alarm. Scientific American, 277(6):24–28, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-24.pdf.

Mukerjee:1997:SAP

[4047] Madhusree Mukerjee. Slaying the “age paradox”. Scientific American, 277(6):28–29, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
28.pdf.

Brack:1997:MCMa

[4048] Matthias Brack. Metal clusters and magic numbers. Scientific American, 277(6):30–35, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Doyle:1997:FFR

[4049] Rodger Doyle. Freshwater fish at risk in the U.S. Scientific American, 277(6):32, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
32.pdf.

Holloway:1997:MW

[4050] Marguerite Holloway. Molding the Web. Scientific American, 277(6):34–36, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
34.pdf.

Anonymous:1997:PWW

[4051] Anonymous. Profile: World Wide Webspinner Tim Berners-Lee. Scientific American, 277(6):34–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1297issue/1297currentissue.html.

Stix:1997:SWS

[4052] Gary Stix. Shock-wave showdown in the Old West. Scientific American, 277(6):38–40, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
38.pdf.

Anonymous:1997:TBSc

[4053] Anonymous. Technology and business: Speed record: 763-mph jetmobile beats Scientific American’s own linear-accelerated go-cart… trees against pollution. Scientific American, 277(6):38–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1297issue/1297currentissue.html.

Gibbs:1997:NE

[4054] W. Wayt Gibbs. Newton 1, Einstein 0. Scientific American, 277(6):40–42, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
40.pdf.

Gibbs:1997:XRS

[4055] W. Wayt Gibbs. X-ray sound. Scientific American, 277(6):42–46, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-42.pdf.

Schmiedeskamp:1997:PPP

[4056] Mia Schmiedeskamp. Pollution-purging poplars. Scientific American, 277(6):46, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
46.pdf.

Joseph:1997:LAG

[4057] Patrick Joseph. On-line advertising goes one-on-one. Scientific American, 277(6):48, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
48.pdf.

Anonymous:1997:CVAb

[4058] Anonymous. Cyber view: Advertisers find new ways to target Web surfers. Scientific American, 277(6):48–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1297issue/1297currentissue.html.

Brack:1997:MCMb

[4059] Matthias Brack. Metal clusters and magic numbers. Scientific American, 277(6):50–55, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
50.pdf.

Braek:1997:MCM

[4060] Matthias Braek. Metal clusters and magic numbers. Scientific American, 277(6):50–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1297issue/1297currentissue.html.

Gibson:1997:CRL

[4061] Everett K. Gibson, Jr., David S. McKay, Kathie Thomas-Keprta, and Christopher S. Romanek. The case for relic life on Mars. Scientific American, 277(6):56–65, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
56.pdf;   http://www.sciam.com/1997/1107issue/1297gibson.html.

Lenhoff:1997:WSB

[4062] Howard M. Lenhoff, Paul P. Wang, Frank Greenberg, and Ursula Bellugi. Williams Syndrome and the brain. Scientific American, 277(6):68–73, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-68.pdf;   http:/
/www.sciam.com/1297issue/1297currentissue.html.

Thomas:1997:TDA

[4063] David A. Thomas and James O. Farlow. Tracking a dinosaur attack weapon. Scientific American, 277(6):74–79, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n6/pdf/scientificamerican1297-74.pdf;   http://www.sciam.com/
1297issue/1297currentissue.html.

Yam:1997:TPE

[4064] Philip Yam. Trends in physics: Exploiting zero-point energy. Scientific American, 277(6):82–85, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
82.pdf;   http://www.sciam.com/1997/1107issue/1297yam.html.

Editors:1997:SRB

[4065] The Editors. Special report: Building the biggest. Scientific American, 277(6):87, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
87.pdf.

Anonymous:1997:BB

[4066] Anonymous. Building the biggest. Scientific American, 277(6):87– ??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1297issue/
1297currentissue.html.

Kashima:1997:LSB

[4067] Satoshi Kashima and Makoto Kitagawa. The longest suspension bridge. Scientific American, 277(6):88–92, December 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n6/pdf/scientificamerican1297-88.pdf;   http://www.sciam.com/
1297issue/1297currentissue.html.

Pelli:1997:WTB

[4068] Cesar Pelli, Charles Thornton, and Leonard Joseph. The world’s tallest buildings. Scientific American, 277(6):92–101, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-92.pdf;   http:/
/www.sciam.com/1297issue/1297currentissue.html.

Kosowatz:1997:BNG

[4069] John J. Kosowatz. Building a new gateway to China. Scientific American, 277(6):102–111, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
102.pdf;   http://www.sciam.com/1297issue/1297currentissue.
html.

Mitchell:1997:DWS

[4070] William J. Mitchell. Do we still need skyscrapers? Scientific American, 277(6):112–113, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
112.pdf;   http://www.sciam.com/1297issue/1297currentissue.
html.

Carlson:1997:AST

[4071] Shawn Carlson. The amateur scientist: Taking back the final frontier. Scientific American, 277(6):114–116, December 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n6/pdf/scientificamerican1297-114.pdf.

Anonymous:1997:ASY

[4072] Anonymous. The amateur scientist: So you want to be a rocketeer. Scientific American, 277(6):114–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1297issue/1297currentissue.html.

Stewart:1997:MRCb

[4073] Ian Stewart. Mathematical recreations: Cat’s cradle calculus challenge. Scientific American, 277(6):118–121, December 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v277/
n6/pdf/scientificamerican1297-118.pdf.

Anonymous:1997:MRC

[4074] Anonymous. Mathematical recreations: Is cat’s cradle child’s play? knot to a mathematician. Scientific American, 277(6):118–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1297issue/1297currentissue.html.

Morrison:1997:RCB

[4075] Philip Morrison and Phylis Morrison. Reviews and commentaries: The Scientific American Young Readers Book Awards. Scientific American, 277(6):122–126, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
122.pdf.

Anonymous:1997:RCSb

[4076] Anonymous. Reviews and commentaries: The Scientific American Young Readers Book Awards. Philip and Phylis Morrison survey the best on science for children and teens. Scientific American, 277(6):122–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.sciam.com/1297issue/1297currentissue.
html.

Burke:1997:CCB

[4077] James Burke. Connections: Cold beer and the Red Baron. Scientific American, 277(6):122–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1297issue/1297currentissue.html.

Burke:1997:T

[4078] James Burke. On track. Scientific American, 277(6):127–128, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-127.pdf.

Anonymous:1997:AI

[4079] Anonymous. Annual index. Scientific American, 277(6):129–131, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v277/n6/pdf/scientificamerican1297-129.pdf.

Reicherter:1997:WKP

[4080] Joel Reicherter. Working knowledge: The polygraph. Scientific American, 277(6):132, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v277/n6/pdf/scientificamerican1297-
132.pdf.

Anonymous:1997:WKT

[4081] Anonymous. Working knowledge: Tell the truth: how polygraphs detect lies. Scientific American, 277(6):132–??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1297issue/1297currentissue.html.

Anonymous:1997:BNG

[4082] Anonymous. Building a New Gateway to China by John J. Kosowatz. Scientific American, 277(6):??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:DWS

[4083] Anonymous. Do We Still Need Skyscrapers? by William J. Mitchell. Scientific American, 277(6):??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:LSB

[4084] Anonymous. The longest suspension bridge. Scientific American, 277 (6):??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:SRc

[4085] Anonymous. Special report. Scientific American, 277(6):??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:WTB

[4086] Anonymous. The World’s Tallest Buildings by Cesar Pelli, Charles Thornton and Leonard Joseph. Scientific American, 277(6):??, December 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sternberg:1997:MBI

[4087] Esther M. Sternberg and Philip W. Gold. The mind-body interaction in disease. Scientific American, ????(Special edition):8–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Crick:1997:PC

[4088] Francis Crick and Christof Koch. The problem of consciousness. Scientific American, ????(Special edition):18–??, ???? 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chalmers:1997:PCE

[4089] David J. Chalmers. The puzzle of conscious experience. Scientific American, ????(Special edition):30–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Myers:1997:PH

[4090] David G. Myers and Ed Diener. The pursuit of happiness. Scientific American, ????(Special edition):40–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jamison:1997:MDI

[4091] Kay Redfield Jamison. Manic-depressive illness and creativity. Scientific American, ????(Special edition):44–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leutwyler:1997:DDS

[4092] Kristin Leutwyler. Depression’s double standard. Scientific American, ????(Special edition):53–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Winson:1997:MD

[4093] Jonathan Winson. The meaning of dreams. Scientific American, ???? (Special edition):58–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

LeDoux:1997:EMB

[4094] Joseph E. LeDoux. Emotion, memory and the brain. Scientific American, ????(Special edition):68–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kalin:1997:NF

[4095] Ned H. Kalin. The neurobiology of fear. Scientific American, ???? (Special edition):76–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Melzack:1997:PL

[4096] Ronald Melzack. Phantom limbs. Scientific American, ????(Special edition):84–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Frith:1997:A

[4097] Uta Frith. Autism. Scientific American, ????(Special edition):92–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1997:SCE

[4098] W. Wayt Gibbs. Seeking the criminal element. Scientific American, ???? (Special edition):102–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1997:FYH

[4099] Glenn Zorpette. Fifty years of heroes and epiphanies: Introducing an epic of raw technology and human triumph. Scientific American, Special Edition(1):7–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Riordan:1997:BE

[4100] Michael Riordan and Lillian Hoddeson. Birth of an era. Scientific American, Special Edition(1):10–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rockett:1997:T

[4101] Frank H. Rockett. The transistor. Scientific American, Special Edition (1):18–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:CT

[4102] Anonymous. Computers from transistors. Scientific American, Special Edition(1):24–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Corcoran:1997:DD

[4103] Elizabeth Corcoran and Glenn Zorpette. Diminishing dimensions. Scientific American, Special Edition(1):24C–??, ???? 1997. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Baliga:1997:HST

[4104] B. Jayant Baliga. How the super-transistor works. Scientific American, Special Edition(1):34–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Riezenman:1997:WTR

[4105] Michael J. Riezenman. Where tubes rule. Scientific American, Special Edition(1):44–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Keyes:1997:FT

[4106] Robert W. Keyes. The future of the transistor. Scientific American, Special Edition(1):46–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Barrett:1997:SSM

[4107] Craig R. Barrett. From sand to silicon: Manufacturing an integrated circuit. Scientific American, Special Edition(1):56–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1997:LM

[4108] W. Wayt Gibbs. The law of more. Scientific American, Special Edition (1):62–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hutches:1997:TES

[4109] G. Dan Hutches and Jerry D. Hutcheson. Technology and economics in the semiconductor industry. Scientific American, Special Edition(1):66– ??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:TPO

[4110] Anonymous. Toward “Point One”. Scientific American, Special Edition (1):74–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Goldstein:1997:TT

[4111] Alan Goldstein. Tackling tyranny. Scientific American, Special Edition (1):80–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Amato:1997:SM

[4112] Ivan Amato. The semiconducting menagerie. Scientific American, Special Edition(1):82–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Patterson:1997:M

[4113] David A. Patterson. Microprocessors in 2020. Scientific American, Special Edition(1):86–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1997:PGW

[4114] Philip Yam. Plastics get wired. Scientific American, Special Edition(1): 90–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lloyd:1997:QMC

[4115] Seth Lloyd. Quantum-mechanical computers. Scientific American, Special Edition(1):98–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Friedlander:1997:FP

[4116] Brad Friedlander and Martyn Roetter. The future of the PC. Scientific American, Special Edition(1):106–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1997:FFA

[4117] Anonymous. Fast facts about the transistor. Scientific American, Special Edition(1):112–??, ???? 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1998:ESD

[4118] John Rennie. From the Editors: a stab in the dark. Scientific American, 278(1):6, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
6.pdf.

Anonymous:1998:LEa

[4119] Anonymous. Letters to the Editors. Scientific American, 278(1):8, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-8.pdf;   http://
www.sciam.com/1998/0198issue/0198currentissue.html.

Anonymous:1998:YAIa years ago. Scientific American, 278(1):10, January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-10.pdf.

Anonymous:1998:YAa

[4121] Anonymous. 50, 100 and 150 years ago. Scientific American, 278(1): 10–??, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0198issue/
0198currentissue.html.

Beardsley:1998:NPS

[4122] Tim Beardsley, Sasha Nemecek, Gary Stix, and Philip Yam. The 1997 Nobel Prizes for Science: A look at the contributions and controversies of the winning work. Scientific American, 278(1):14–16, 18 (Intl. ed. 8–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0198issue/
0198currentissue.html.

Stix:1998:SBN

[4123] Gary Stix. Special briefing: The 1997 Nobel Prizes in science. Scientific American, 278(1):14–18, January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-14.pdf.

Schneider:1998:NAB

[4124] David Schneider. News and analysis: Burying the problem. Scientific American, 278(1):21–22, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
21.pdf.

Schneider:1998:FPC

[4125] David Schneider. In focus: Pumping CO_{2} out of the air could help fight the greenhouse effect. Scientific American, 278(1):21–?? (Intl. ed. 13–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0198issue/
0198currentissue.html.

Gibbs:1998:ES

[4126] W. Wayt Gibbs. Extreme science. Scientific American, 278(1):24–28, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-24.pdf.

Anonymous:1998:SCRa

[4127] Anonymous. Science and the citizen: Reassessing Neanderthal DNA. how stress hurts brains. meat TNTenderizer. Scientific American, 278 (1):24–?? (Intl. ed. 17–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0198issue/0198currentissue.html.

Leutwyler:1998:DS

[4128] Kristin Leutwyler. Don’t stress. Scientific American, 278(1):28–30, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-28.pdf.

Anonymous:1998:CVM

[4129] Anonymous. Cyber view: Making fashion compute. Scientific American, 278(1):29–??, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0198issue/
0198currentissue.html.

Wong:1998:AQ

[4130] Kate Wong. Ancestral quandary. Scientific American, 278(1):30–32, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-30b.pdf.

Anonymous:1998:Ba

[4131] Anonymous. In brief. Scientific American, 278(1):30–34, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-30a.pdf.

Mirsky:1998:TB

[4132] Steve Mirsky. Tender is the bite. Scientific American, 278(1):34, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-34.pdf.

Doyle:1998:WPT

[4133] Rodger Doyle. Women in politics throughout the world. Scientific American, 278(1):35, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
35.pdf.

Gibbs:1998:NMN

[4134] W. Wayt Gibbs. From naked men to a new-world order. Scientific American, 278(1):38–40, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
38.pdf.

Gibbs:1998:PCL

[4135] W. Wayt Gibbs. Profile: Claude Levi-Strauss, anthropologist. Scientific American, 278(1):38, 40 (Intl. ed. 24–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0198issue/0198currentissue.html.

Zorpette:1998:H

[4136] Glenn Zorpette. Off with its head! Scientific American, 278(1):41, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-41.pdf.

Anonymous:1998:TBC

[4137] Anonymous. Technology and business: Carbon adds zip to silicon. cloning for organs. roaches at the wheel. Scientific American, 278(1): 41, 44–45 (Intl. ed. 26–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0198issue/0198currentissue.html.

Dupont:1998:LS

[4138] Daniel G. Dupont. Laser show. Scientific American, 278(1):44, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-44a.pdf.

Hayashi:1998:NST

[4139] Alden M. Hayashi. New silicon tricks. Scientific American, 278(1):44–45, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-44b.pdf.

Yam:1998:RW

[4140] Philip Yam. Roaches at the wheel. Scientific American, 278(1):45, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-45.pdf.

Grossman:1998:WYC

[4141] Wendy M. Grossman. Wearing your computer. Scientific American, 278(1):46, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
46.pdf.

Grossman:1998:CVM

[4142] Wendy M. Grossman. Cyber view: Making fashion compute. Scientific American, 278(1):46–??, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0198issue/0198currentissue.html.

Ingber:1998:AL

[4143] Donald E. Ingber. The architecture of life. Scientific American, 278(1): 48–57 (Intl. ed. 30–39), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
48.pdf;   http://www.sciam.com/1998/0108issue/0198ingber.html.

Krauss:1998:CA

[4144] Lawrence M. Krauss. Cosmological antigravity. Scientific American, 278 (1):52–59, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hollister:1998:BRW

[4145] Charles D. Hollister and Steven Nadis. Burial of radioactive waste under the seabed. Scientific American, 278(1):60–65 (Intl. ed. 40–45), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-60.pdf;   http:/
/www.sciam.com/1998/0198issue/0198currentissue.html.

Miller:1998:BGS

[4146] Robert V. Miller. Bacterial gene swapping in nature. Scientific American, 278(1):66–71 (Intl. ed. 46–??), January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-66.pdf.

Tallamy:1998:CCA

[4147] Douglas W. Tallamy and Ken Preston-Mafham. Child care among the insects. Scientific American, 278(1):72–77, January 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Smith:1998:UM

[4148] Edward J. Smith and Richard G. Marsden. The Ulysses mission. Scientific American, 278(1):74–79 (Intl. ed. 52–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-74.pdf;   http:/
/www.sciam.com/1998/0198issue/0198currentissue.html.

Laver:1998:DFV

[4149] W. Graeme Laver, Norbert Bischofberger, and Robert G. Webster. Disarming flu viruses. Scientific American, 278(1):78–87, January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1999/0199issue/0199laver.html.

Sime:1998:LMD

[4150] Ruth Lewin Sime. Lise Meitner and the discovery of nuclear fission. Scientific American, 278(1):80–85 (Intl. ed. 58–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-80.pdf.

Anonymous:1998:MRBa

[4151] Anonymous. Mathematical recreations: Bubbles make complex math easy. Scientific American, 278(1):82–??, January 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0198issue/0198currentissue.html.

Maris:1998:PU

[4152] Humphrey Maris. Picosecond ultrasonics. Scientific American, 278(1): 86–89 (Intl. ed. 64–67), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
86.pdf;   http://www.sciam.com/1998/0198issue/0198currentissue.
html.

Brown:1998:PE

[4153] Walter A. Brown. The Placebo Effect. Scientific American, 278(1): 90–95 (Intl. ed. 68–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
90.pdf.

Anonymous:1998:WKHa

[4154] Anonymous. Working knowledge: Holograms: giving pictures depth. Scientific American, 278(1):92–??, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0198issue/0198currentissue.html.

Moxon:1998:DMA

[4155] E. Richard Moxon and Christopher Wills. DNA microsatellites: Agents of evolution? Scientific American, 278(1):94–99, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Foley:1998:LIW

[4156] Vernard Foley. Leonardo and the invention of the wheellock. Scientific American, 278(1):96–101 (Intl. ed. 74–??), January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-96.pdf;   http://www.sciam.com/
1998/0108issue/0198foley.html.

Carlson:1998:ASK

[4157] Shawn Carlson. The amateur scientist: a kitchen centrifuge. Scientific American, 278(1):102–103 (Intl. ed. 80–??), January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-102.pdf;   http://www.sciam.com/
1998/0198issue/0198currentissue.html.

Stewart:1998:MRBa

[4158] Ian Stewart. Mathematical recreations: Bubbles make complex math easy. Scientific American, 278(1):104–107, January 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0198issue/0198currentissue.html.

Stewart:1998:MRD

[4159] Ian Stewart. Mathematical recreations: Double bubble, toil and trouble. Scientific American, 278(1):104–107, January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-104.pdf.

Logsdon:1998:RCR

[4160] John M. Logsdon. Reviews and commentaries: The race into space. Scientific American, 278(1):108–110, January 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n1/pdf/scientificamerican0198-108.pdf;   http://www.sciam.com/
1998/0198issue/0198currentissue.html.

Wallich:1998:RCP

[4161] Paul Wallich. Reviews and commentaries: Preserving the world. Scientific American, 278(1):110, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0198issue/0198currentissue.html.

Morrison:1998:Wa

[4162] Philip Morrison and Phylis Morrison. Wildfire. Scientific American, 278(1):112, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n1/pdf/scientificamerican0198-
112.pdf.

Morrison:1998:WW

[4163] Philip Morrison and Phylis Morrison. Wonders: Wildfire. Scientific American, 278(1):112, 114 (86–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/1998/0198issue/0198currentissue.html.

Burke:1998:AT

[4164] James Burke. Anybody out there? Scientific American, 278(1):113–114, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-113.pdf.

Burke:1998:CAT

[4165] James Burke. Connections: Anybody out there? Scientific American, 278(1):113–114 (Intl. ed. 86–??), January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0198issue/0198currentissue.html.

Benton:1998:WK

[4166] Stephen A. Benton. Working knowledge. Scientific American, 278(1):115, January 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n1/pdf/scientificamerican0198-115.pdf.

Benton:1998:WKH

[4167] Stephen A. Benton. Working knowledge: Holograms: giving pictures depth. Scientific American, 278(1):115, January 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0198issue/0198currentissue.html.

Rennie:1998:ESB

[4168] John Rennie. From the Editor: Saving at the blood bank. Scientific American, 278(2):6, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
6.pdf.

Anonymous:1998:FBN

[4169] Anonymous. In focus: Brookhaven National Laboratory recovers from a public-relations meltdown. Scientific American, 278(2):7– ??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Anonymous:1998:Ea

[4170] Anonymous. Errata. Scientific American, 278(2):8, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-8b.pdf.

Anonymous:1998:LEb

[4171] Anonymous. Letters to the Editors. Scientific American, 278(2):8, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-8a.pdf;   http:/
/www.sciam.com/1998/0298issue/0298currentissue.html.

Anonymous:1998:SCP

[4172] Anonymous. Science and the citizen: PCB research in limbo… the ecology of plastic flamingos… flea-size supernovae. Scientific American, 278 (2):11–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Anonymous:1998:YAIb

[4173] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 278(2):12, February 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-12.pdf.

Anonymous:1998:YAb

[4174] Anonymous. 50, 100 and 150 years ago. Scientific American, 278(2): 12–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Mukerjee:1998:FBB

[4175] Madhusree Mukerjee. In focus: Brookhaven brouhaha: The Laboratory tries to recover from a public-relations fallout of radioactive leaks and chemical dumping. Scientific American, 278(2):15–16, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0298issue/0298currentissue.html.

Mukerjee:1998:NAB

[4176] Madhusree Mukerjee. News and analysis: Brookhaven brouhaha. Scientific American, 278(2):15–16, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
15.pdf.

Anonymous:1998:BNL

[4177] Anonymous. Brookhaven National Laboratory recovers from a publicrelations meltdown. Scientific American, 278(2):15–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0298issue/0298currentissue.html.

Gibbs:1998:SCP

[4178] W. Wayt Gibbs. Science and the citizen: PCB research in limbo… the ecology of plastic flamingos… flea-size supernovae. Scientific American, 278(2):17, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Gibbs:1998:PS

[4179] W. Wayt Gibbs. Playing with stars. Scientific American, 278(2):17, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-17.pdf.

Anonymous:1998:PFCa

[4180] Anonymous. Profile: Francis S. Collins, leading the U.S. Human Genome Project. Scientific American, 278(2):18–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0298issue/0298currentissue.html.

Mirsky:1998:W

[4181] Steve Mirsky. Whatchamacallit. Scientific American, 278(2):20, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-20b.pdf.

GordonSaugerties:1998:PP

[4182] Jim GordonSaugerties and N. Y. Politics and pcb. Scientific American, 278(2):20–22, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
20a.pdf.

Leutwyler:1998:Ba

[4183] Kristin Leutwyler. In brief. Scientific American, 278(2):22–24, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-22.pdf.

Anonymous:1998:TBHa

[4184] Anonymous. Technology and business: Herbal medicine under scrutiny. polishing for flatter, faster chips… trapping light. Scientific American, 278(2):22–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0298issue/0298currentissue.html.

Mirsky:1998:PB

[4185] Steve Mirsky. The painted bird. Scientific American, 278(2):23–24, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-23.pdf.

Yam:1998:BB

[4186] Philip Yam. Boom box. Scientific American, 278(2):24–26, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-24.pdf.

Doyle:1998:DEC

[4187] Rodger Doyle. Deaths from excessive cold and excessive heat. Scientific American, 278(2):26, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
26.pdf.

Beardsley:1998:PFC

[4188] Tim Beardsley. Profile: Francis S. Collins, leading the U.S. Human Genome Project. Scientific American, 278(2):28–29, February 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0298issue/0298currentissue.html.

Beardsley:1998:WSR

[4189] Tim Beardsley. Where science and religion meet. Scientific American, 278(2):28–29, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
28.pdf.

Anonymous:1998:PFCb

[4190] Anonymous. Profile: Francis S. Collins, leading the U.S. Human Genome Project. Scientific American, 278(2):28–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0298issue/0298currentissue.html.

Hayashi:1998:TBH

[4191] Alden M. Hayashi. Technology and business: Herbal medicine under scrutiny. polishing for flatter, faster chips… trapping light. Scientific American, 278(2):30, 32, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0298issue/0298currentissue.html.

Stix:1998:PM

[4192] Gary Stix. Plant matters. Scientific American, 278(2):30–32, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-30.pdf.

Hayashi:1998:LTN

[4193] Alden M. Hayashi. Let there be no light. Scientific American, 278(2):32, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-32.pdf.

Gibbs:1998:NBG

[4194] W. Wayt Gibbs. Natural-born guinea pigs. Scientific American, 278(2):34, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
34a.pdf.

Stix:1998:FSF

[4195] Gary Stix. Faster, smaller, flatter. Scientific American, 278(2):34–36, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-34b.pdf.

Stix:1998:ES

[4196] Gary Stix. Is the end in sight? Scientific American, 278(2):36, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-36.pdf.

Lewis:1998:CVM

[4197] Ted Lewis. Cyber view: Is Microsoft a monopoly? Scientific American, 278(2):37–37 (Intl. ed. 26–??), February 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0298issue/0298currentissue.html;   http:
//www.sciam.com/1998/0298issue/0298cyber.html.

Lewis:1998:MNM

[4198] Ted Lewis. Is Microsoft a natural monopoly? Scientific American, 278(2):37, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
37.pdf.

Padian:1998:OBT

[4199] Kevin Padian and Luis M. Chiappe. The origin of birds and their flight. Scientific American, 278(2):38–47 (Intl. ed. 28–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-38.pdf.

Richelson:1998:SB

[4200] Jeffrey T. Richelson. Scientists in black. Scientific American, 278(2): 48–55 (Intl. ed. 38–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
48.pdf.

Hale:1998:VL

[4201] John R. Hale. The Viking longship. Scientific American, 278(2):56– 63 (Intl. ed. 46–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
56.pdf;   http://www.sciam.com/1998/0208issue/0298hale.html.

Duff:1998:TFK

[4202] Michael J. Duff. The theory formerly known as strings. Scientific American, 278(2):64–69 (Intl. ed. 54–??), February 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-64.pdf.

Nucci:1998:SBS

[4203] Mary L. Nucci and Abraham Abuchowski. The search for blood substitutes. Scientific American, 278(2):72–77 (Intl. ed. 60–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-72.pdf.

Alley:1998:GIC

[4204] Richard B. Alley and Michael L. Bender. Greenland ice cores: Frozen in time. Scientific American, 278(2):80–85 (Intl. ed. 66–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-80.pdf.

Anonymous:1998:RCR

[4205] Anonymous. Reviews and commentaries: Rachel Carson appreciated… how science and myth made Star Wars successful. Scientific American, 278(2):83–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0298issue/0298currentissue.html.

Burke:1998:CGSa

[4206] James Burke. Connections: Green silk dresses, the speed of light and botanical gardens. Scientific American, 278(2):83–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0298issue/0298currentissue.html.

Morrison:1998:WPS

[4207] Philip Morrison. Wonders: Putting the stars in their places. Scientific American, 278(2):83–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0298issue/0298currentissue.html.

Ott:1998:EET

[4208] Wayne R. Ott and John W. Roberts. Everyday exposure to toxic pollutants. Scientific American, 278(2):86–91 (Intl. ed. 72–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-86.pdf;   http:/
/www.sciam.com/1998/0208issue/0298ott.html.

Anonymous:1998:WKS

[4209] Anonymous. Working knowledge: Stop! how hydraulic brakes work. Scientific American, 278(2):88–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/1998/0298issue/0298currentissue.html.

Carlson:1998:ASBa

[4210] Shawn Carlson. The amateur scientist: Bird-watching by the numbers. Scientific American, 278(2):92–93 (Intl. ed. 78–??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-92.pdf;   http:/
/www.sciam.com/1998/0298issue/0298currentissue.html.

Anonymous:1998:ASa

[4211] Anonymous. Amateur scientist. Scientific American, 278(2):92– ??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Stewart:1998:MRGa

[4212] Ian Stewart. Mathematical recreations: Geometry puts the squeeze on sardines. Scientific American, 278(2):94–95, 97 (Intl. ed. 80– ??), February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Stewart:1998:MRT

[4213] Ian Stewart. Mathematical recreations: Tight tins for round sardines. Scientific American, 278(2):94–97, February 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n2/pdf/scientificamerican0298-94.pdf.

Anonymous:1998:RCa

[4214] Anonymous. Reviews and commentaries. Scientific American, 278(2):98–99, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
98.pdf.

Holloway:1998:RCR

[4215] Marguerite Holloway. Reviews and commentaries: Rachel Carson appreciated… how science and myth made Star Wars successful. Scientific American, 278(2):98–99, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0298issue/0298currentissue.html.

Anonymous:1998:RCb

[4216] Anonymous. Reviews and commentaries. Scientific American, 278(2): 98–??, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Morrison:1998:SM

[4217] Philip Morrison. The star mapper. Scientific American, 278(2):100–102, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-100.pdf.

Morrison:1998:WSM

[4218] Philip Morrison. Wonders: The star mapper. Scientific American, 278 (2):100, 102–103, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0298issue/0298currentissue.html.

Burke:1998:CGSb

[4219] James Burke. Connections: Green silk dresses. Scientific American, 278 (2):101–102, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0298issue/
0298currentissue.html.

Burke:1998:LC

[4220] James Burke. Local color. Scientific American, 278(2):101–103, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n2/pdf/scientificamerican0298-101.pdf.

Stokes:1998:WKH

[4221] Stanley L. Stokes. Working knowledge: Hydraulic brakes. Scientific American, 278(2):104, February 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n2/pdf/scientificamerican0298-
104.pdf.

Stokes:1998:WKS

[4222] Stanley L. Stokes. Working knowledge: Stop! how hydraulic brakes work. Scientific American, 278(2):104, February 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0298issue/0298currentissue.html.

Rennie:1998:ESI

[4223] John Rennie. From the Editor: Screen idols and real stars. Scientific American, 278(3):6, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
6.pdf.

Anonymous:1998:FPT

[4224] Anonymous. In focus: PCs and TVs get hitched, while their industries feud. Scientific American, 278(3):7–??, March 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0398issue/0398currentissue.html.

Anonymous:1998:Eb

[4225] Anonymous. Erratum. Scientific American, 278(3):10, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-10b.pdf.

Anonymous:1998:LEc

[4226] Anonymous. Letters to the Editors. Scientific American, 278(3):10, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-10a.pdf;   http:
//www.sciam.com/1998/0398issue/0398currentissue.html.

Anonymous:1998:SCS

[4227] Anonymous. Science and the citizen: A second cosmic background radiation. muscles over mind. vanishing languages. Scientific American, 278(3):10–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Anonymous:1998:YAIc

[4228] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 278(3):12, March 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-12.pdf.

Anonymous:1998:YAc

[4229] Anonymous. 50, 100 and 150 years ago. Scientific American, 278 (3):12–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Eisenberg:1998:FDD

[4230] Anne Eisenberg. In focus: Digital dilemma: The upcoming digital format for television lends itself to computer use, but disagreements about transmission standards could affect the melding of the two. Scientific American, 278(3):15–16, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0398issue/0398currentissue.html.

Eisenberg:1998:NAD

[4231] Anne Eisenberg. News and analysis: Digital dilemma. Scientific American, 278(3):15–16, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
15.pdf.

Anonymous:1998:Fa

[4232] Anonymous. In focus. Scientific American, 278(3):15–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0398issue/0398currentissue.html.

Anonymous:1998:PPAa

[4233] Anonymous. Profile: Physicist Alan Sokal worries about relativism, not relativity. Scientific American, 278(3):17–??, March 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0398issue/0398currentissue.html.

Musser:1998:GD

[4234] George Musser. Glow in the dark. Scientific American, 278(3):18, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-18.pdf.

Musser:1998:SCG

[4235] George Musser. Science and the citizen: Glow in the dark: A second cosmic background radiation permeates the sky. Scientific American, 278(3):18, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Anonymous:1998:TBT

[4236] Anonymous. Technology and business: Turning off studies of power-line effects on health. airborne composites. A biosolar cell. Scientific American, 278(3):19–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0398issue/0398currentissue.html.

Harby:1998:UR

[4237] Karla Harby. Unsound reasoning. Scientific American, 278(3):20–21, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-20b.pdf.

Anonymous:1998:Bb

[4238] Anonymous. In brief. Scientific American, 278(3):20–25, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-20a.pdf.

Mirsky:1998:UM

[4239] Steve Mirsky. Urine the money. Scientific American, 278(3):21, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-21.pdf.

Zorpette:1998:YSB

[4240] Gelenn Zorpette. You see brawny; I see scrawny. Scientific American, 278(3):24–25, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
24.pdf.

Beardsley:1998:CR

[4241] Tim Beardsley. Catching the rays. Scientific American, 278(3):25–26, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-25.pdf.

Doyle:1998:LDD

[4242] Rodger Doyle. Languages, disappearing and dead. Scientific American, 278(3):26, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
26.pdf.

Doyle:1998:NLD

[4243] Rodger Doyle. By the numbers: Languages, disappearing and dead. Scientific American, 278(3):26, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0398issue/0398currentissue.html.

Mukerjee:1998:PUE

[4244] Madhusree Mukerjee. Profile: Undressing the Emperor: Physicist and Social Text prankster Alan Sokal fires another salvo at thinkers in the humanities. Scientific American, 278(3):30, 32, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0398issue/0398currentissue.html.

Mukerjee:1998:UE

[4245] Madhusree Mukerjee. Undressing the emperor. Scientific American, 278(3):30–32, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
30.pdf.

Anonymous:1998:Pa

[4246] Anonymous. Profile. Scientific American, 278(3):30–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0398issue/0398currentissue.html.

Stix:1998:CB

[4247] Gary Stix. Closing the book. Scientific American, 278(3):33–34, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-33.pdf.

Stix:1998:TBC

[4248] Gary Stix. Technology and business: Closing the book: Are powerline fields a dead issue? Scientific American, 278(3):33–34, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0398issue/0398currentissue.html.

Scott:1998:CS

[4249] Phil Scott. Composite sketch. Scientific American, 278(3):34–35, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-34.pdf.

Dekoker:1998:SC

[4250] Brenda Dekoker. Scene of the crime. Scientific American, 278(3):35–36, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-35.pdf.

Stix:1998:FIB

[4251] Gary Stix. Fake it before you make it. Scientific American, 278(3):36, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-36.pdf.

Grossman:1998:CVD

[4252] Wendy M. Grossman. Cyber view: Downloading as a crime. Scientific American, 278(3):37, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0398issue/0398currentissue.html.

Grossman:1998:DC

[4253] Wendy M. Grossman. Downloading as a crime. Scientific American, 278(3):37, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
37.pdf.

Anonymous:1998:CVP

[4254] Anonymous. Cyber view: Piracy and profits: software restrictions bruise the public. Scientific American, 278(3):37–?? (Intl. ed. 23– ??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Cornell:1998:BEC

[4255] Eric A. Cornell and Carl E. Wieman. The Bose–Einstein condensate. Scientific American, 278(3):40–45 (Intl. ed. 26–31), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-40.pdf;   http:/
/www.sciam.com/1998/0398issue/0398currentissue.html.

Gourley:1998:M

[4256] Paul L. Gourley. Microlasers. Scientific American, 278(3):40–45, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Levy:1998:CAR

[4257] Stuart B. Levy. The challenge of antibiotic resistance. Scientific American, 278(3):46–53 (Intl. ed. 32–??), March 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-46.pdf;   http://www.sciam.com/
1998/0308issue/0398levy.html.

Gourley:1998:N

[4258] Paul L. Gourley. Nanolasers. Scientific American, 278(3):56–61, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-56.pdf;   http:/
/www.sciam.com/1998/0398issue/0398currentissue.html.

Hodgins:1998:AHM

[4259] Jessica K. Hodgins. Animating human motion. Scientific American, 278 (3):64–69 (Intl. ed. 46–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
64.pdf;   http://www.sciam.com/1998/0308issue/0398hodgins.html.

Brazaitis:1998:CT

[4260] Peter Brazaitis, Myrna E. Watanabe, and George Amato. The caiman trade. Scientific American, 278(3):70–76 (Intl. ed. 52–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-70.pdf.

Anonymous:1998:SRPa

[4261] Anonymous. Special report: Preventing the next oil crun. Scientific American, 278(3):77, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
77.pdf.

Anonymous:1998:PNO

[4262] Anonymous. Preventing the next oil crunch. Scientific American, 278 (3):77–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Anonymous:1998:SRPb

[4263] Anonymous. Special report: Preventing the next oil crunch. Scientific American, 278(3):77–?? (Intl. ed. 59–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0398issue/0398currentissue.html.

Campbell:1998:ECO

[4264] Colin J. Campbell and Jean H. Laherrère. The end of cheap oil. Scientific American, 278(3):78–83 (Intl. ed. 60–??), March 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-78.pdf.

Burke:1998:CTW

[4265] James Burke. Connections: The Trojan War, modern armies and the pas de deux. Scientific American, 278(3):82–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0398issue/0398currentissue.html.

Morrison:1998:WIR

[4266] Philip Morrison. Wonders: The iron rainbow. Scientific American, 278 (3):82–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

George:1998:MO

[4267] Richard L. George. Mining for oil. Scientific American, 278(3):84– 85 (Intl. ed. 66–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
84.pdf.

Anderson:1998:OPC

[4268] Roger N. Anderson. Oil production in the 21st Century. Scientific American, 278(3):86–91 (Intl. ed. 68–??), March 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-86.pdf.

Anonymous:1998:WKY

[4269] Anonymous. Working knowledge: “you have mail”: how the Internet delivers. Scientific American, 278(3):88–??, March 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0398issue/0398currentissue.html.

Fouda:1998:LFN

[4270] Safaa A. Fouda. Liquid fuels from natural gas. Scientific American, 278 (3):92–95 (Intl. ed. 74–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
92.pdf.

Carlson:1998:ASE

[4271] Shawn Carlson. The amateur scientist: Exploring a micropond. Scientific American, 278(3):96–98 (Intl. ed. 78–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0398issue/0398currentissue.html.

Carlson:1998:ASP

[4272] Shawn Carlson. The amateur scientist: The pleasures of pond scum. Scientific American, 278(3):96–98, March 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n3/pdf/scientificamerican0398-96.pdf.

Anonymous:1998:ASb

[4273] Anonymous. Amateur scientist. Scientific American, 278(3):96–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Stewart:1998:MRGb

[4274] Ian Stewart. Mathematical recreations: Glass Klein bottles, Möbius strips and stranger surfaces. Scientific American, 278(3):100–101 (Intl. ed. 80–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
100.pdf;   http://www.sciam.com/1998/0398issue/0398currentissue.
html.

Anonymous:1998:MRa

[4275] Anonymous. Mathematical recreations. Scientific American, 278(3): 100–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Anonymous:1998:RCc

[4276] Anonymous. Reviews and commentaries. Scientific American, 278(3):102–103, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
102.pdf.

Johnson:1998:RCW

[4277] Donald Johnson. Reviews and commentaries: Why humans are human and Neanderthals were not: the evolution of our unique nature. Scientific American, 278(3):102–103 (Intl. ed. 82–??), March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0398issue/0398currentissue.html.

Burke:1998:RC

[4278] James Burke. Reviews and commentaries. Scientific American, 278 (3):102–??, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Morrison:1998:PA

[4279] Philip Morrison and Phylis Morrison. Painters’ atoms. Scientific American, 278(3):104–106, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-
104.pdf.

Morrison:1998:WPA

[4280] Philip Morrison and Phylis Morrison. Wonders: Painters’ atoms. Scientific American, 278(3):104, 106–107, March 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0398issue/0398currentissue.html.

Burke:1998:CTD

[4281] James Burke. Connections: Turkish delight. Scientific American, 278 (3):105–106, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0398issue/
0398currentissue.html.

Burke:1998:TD

[4282] James Burke. Turkish delight. Scientific American, 278(3):105–107, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n3/pdf/scientificamerican0398-105.pdf.

Hoffman:1998:WKI

[4283] Paul Hoffman. Working knowledge: Internet electronic mail. Scientific American, 278(3):108, March 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n3/pdf/scientificamerican0398-

108.pdf;   http://www.sciam.com/1998/0398issue/0398currentissue.
html.

Rennie:1998:EHB

[4284] John Rennie. From the Editor: How Scientific American works. Scientific American, 278(4):6, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
6.pdf.

Anonymous:1998:FRA

[4285] Anonymous. In focus: Research on an AIDS vaccine continues despite discouragements. Scientific American, 278(4):9–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0498issue/0498currentissue.html.

Anonymous:1998:LEd

[4286] Anonymous. Letters to the Editors. Scientific American, 278(4):10, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-10.pdf;   http:/
/www.sciam.com/1998/0498issue/0498currentissue.html.

Anonymous:1998:SCEa

[4287] Anonymous. Science and the citizen: Earth invades Mars, kills microfossils. shaky ground for earthquake predictions. knees and jet lag. Scientific American, 278(4):13–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Anonymous:1998:YAId

[4288] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 278(4):14, April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n4/pdf/scientificamerican0498-14.pdf.

Anonymous:1998:YAd

[4289] Anonymous. 50, 100 and 150 years ago. Scientific American, 278 (4):14–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Beardsley:1998:FLB

[4290] Tim Beardsley. In focus: Lives in the balance: Researchers plan to modify HIV and try it as a live AIDS vaccine. Scientific American, 278(4):17–18, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Beardsley:1998:NAL

[4291] Tim Beardsley. News and analysis: Lives in the balance. Scientific American, 278(4):17–18, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
17.pdf.

Anonymous:1998:Fb

[4292] Anonymous. In focus. Scientific American, 278(4):17–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0498issue/0498currentissue.html.

Gibbs:1998:E

[4293] W. Wayt Gibbs. Endangered. Scientific American, 278(4):19–20, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-19.pdf.

Anonymous:1998:SCEb

[4294] Anonymous. Science and the citizen: Earth invades Mars, kills microfossils. shaky ground for earthquake predictions. knees and jet lag. Scientific American, 278(4):19–24, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Hopkin:1998:CS

[4295] Karen Hopkin. Clock setting. Scientific American, 278(4):20–22, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-20.pdf.

Anonymous:1998:PSSa

[4296] Anonymous. Profile: Sociologist Sherry Turkle looks at human-machine relations. Scientific American, 278(4):20–??, April 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0498issue/0498currentissue.html.

Mirsky:1998:CR

[4297] Steve Mirsky. Comic relief. Scientific American, 278(4):22–23, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-22b.pdf.

Leutwyler:1998:Bb

[4298] Kristin Leutwyler. In brief. Scientific American, 278(4):22–24, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-22a.pdf.

Anonymous:1998:TBM

[4299] Anonymous. Technology and business: Medications expire; profits don’t. new micro-fuel cell. fertilizing the ocean. faster modems on the way. Scientific American, 278(4):22–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Schneider:1998:SG

[4300] David Schneider. On shaky ground. Scientific American, 278(4):23–24, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-23.pdf.

Doyle:1998:FDU

[4301] Rodger Doyle. Forest density in the U.S. Scientific American, 278(4):28, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-28.pdf.

Holloway:1998:EC

[4302] Marguerite Holloway. An ethnologist in cyberspace. Scientific American, 278(4):29–30, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
29.pdf.

Holloway:1998:PEC

[4303] Marguerite Holloway. Profile: An ethnologist in cyberspace. Scientific American, 278(4):29–30, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Anonymous:1998:Pb

[4304] Anonymous. Profile. Scientific American, 278(4):29–??, April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0498issue/0498currentissue.html.

Nemecek:1998:WDO

[4305] Sasha Nemecek. Not what the doctor ordered. Scientific American, 278(4):31–32, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
31.pdf.

Anonymous:1998:TB

[4306] Anonymous. Technology and business. Scientific American, 278(4): 31–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Hayashi:1998:TEB

[4307] Alden M. Hayashi. Taking on the energizer bunny. Scientific American, 278(4):32, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
32.pdf.

Nadis:1998:FS

[4308] Steve Nadis. Fertilizing the sea. Scientific American, 278(4):33, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-33.pdf.

Zorpette:1998:NFP

[4309] Glenn Zorpette. A new fat pipe. Scientific American, 278(4):34, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-34.pdf.

Eisenberg:1998:CVU

[4310] Anne Eisenberg. Cyber view: The Unabomber and the bland decade. Scientific American, 278(4):35, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Eisenberg:1998:UBD

[4311] Anne Eisenberg. The Unabomber and the bland decade. Scientific American, 278(4):35, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
35.pdf.

Anonymous:1998:CVU

[4312] Anonymous. Cyber view: The Unabomber, the Internet and radicalism. Scientific American, 278(4):35–?? (Intl. ed. 27–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0498issue/0498currentissue.html.

Tarle:1998:CA

[4313] Gregory Tarlé and Simon P. Swordy. Cosmic antimatter. Scientific American, 278(4):36–41 (Intl. ed. 30–??), April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n4/pdf/scientificamerican0498-36.pdf.

Halstead:1998:PPS

[4314] Lauro S. Halstead. Post-polio syndrome. Scientific American, 278(4): 42–47 (Intl. ed. 36–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
42.pdf.

Ford:1998:EV

[4315] Brian J. Ford. The earliest views. Scientific American, 278(4):50–53 (Intl. ed. 42–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Ford:1998:SPE

[4316] Brian J. Ford. Science in pictures: The earliest views. Scientific American, 278(4):50–53, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
50.pdf.

Dugatkin:1998:HFC

[4317] Lee Alan Dugatkin and Jean-Guy J. Godin. How females choose their mates. Scientific American, 278(4):56–61 (Intl. ed. 46–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-56.pdf;   http:/
/www.sciam.com/1998/0408issue/0498dugatkin.html.

Berns:1998:LST

[4318] Michael W. Berns. Laser scissors and tweezers. Scientific American, 278 (4):62–67 (Intl. ed. 52–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
62.pdf;   http://www.sciam.com/1998/0408issue/0498berns.html.

Anonymous:1998:SRW

[4319] Anonymous. Special report: Wireless technologies. Scientific American, 278(4):69, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
69.pdf.

Anonymous:1998:WTD

[4320] Anonymous. Wireless technologies: Demand for cellular phones and wireless modems is skyrocketing — and they represent only a few of the applications for the new. Scientific American, 278(4):69–69 (Intl. ed. 59–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Anonymous:1998:WT

[4321] Anonymous. Wireless technologies. Scientific American, 278(4): 69–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Evans:1998:NSP

[4322] John V. Evans. New satellites for personal communications. Scientific American, 278(4):70–77 (Intl. ed. 60–??), April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n4/pdf/scientificamerican0498-70.pdf.

Pelton:1998:TC

[4323] Joseph N. Pelton. Telecommunications for the 21st Century. Scientific American, 278(4):80–85 (Intl. ed. 68–??), April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n4/pdf/scientificamerican0498-80.pdf.

Hills:1998:TWN

[4324] Alex Hills. Terrestrial wireless networks. Scientific American, 278(4): 86–91 (Intl. ed. 74–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
86.pdf.

Anonymous:1998:MRR

[4325] Anonymous. Mathematical recreations: Repealing the law of averages. Scientific American, 278(4):87–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Burke:1998:CCL

[4326] James Burke. Commentaries: Connections: Lady Liberty, liquid hydrogen and a quartz sandwich. Scientific American, 278(4): 89–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Morrison:1998:CWS

[4327] Philip Morrison. Commentaries: Wonders: The serious pursuit of precision timekeeping. Scientific American, 278(4):89–??, April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0498issue/0498currentissue.html.

Stutzman:1998:MBW

[4328] Warren L. Stutzman and Carl B. Dietrich, Jr. Moving beyond wireless voice systems. Scientific American, 278(4):92–93 (Intl. ed. 80–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-92.pdf.

Hughes:1998:SSR

[4329] David R. Hughes and Dewayne Hendricks. Spread-spectrum radio. Scientific American, 278(4):94–96 (Intl. ed. 82–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-94.pdf;   http:/
/www.sciam.com/1998/0498issue/0498currentissue.html.

Carlson:1998:ASM

[4330] Shawn Carlson. The amateur scientist: Making experiments out of thin air. Scientific American, 278(4):98–100, April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n4/pdf/scientificamerican0498-98.pdf.

Anonymous:1998:ASc

[4331] Anonymous. Amateur scientist. Scientific American, 278(4):98–??, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Carlson:1998:ASC

[4332] Shawn Carlson. The amateur scientist: Controlling atmospheric pressure. Scientific American, 278(4):98, 100 (Intl. ed. 85–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0498issue/0498amsci.html;   http:
//www.sciam.com/1998/0498issue/0498currentissue.html.

Stewart:1998:MRRa

[4333] Ian Stewart. Mathematical recreations: Repealing the law of averages. Scientific American, 278(4):102–104, April 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n4/pdf/scientificamerican0498-102.pdf;   http://www.sciam.com/
1998/0498issue/0498currentissue.html.

Morrison:1998:WTE

[4334] Philip Morrison. Wonders: The timekeeping ELF. Scientific American, 278(4):105, 107, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0498issue/
0498currentissue.html.

Morrison:1998:CTE

[4335] Philip Morrison. Commentaries: The timekeeping ELF. Scientific American, 278(4):105–107, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n4/pdf/scientificamerican0498-
105.pdf.

Burke:1998:CSP

[4336] James Burke. Connections: Sheer poetry: It came to be known as a Langevin sandwich: a layer of quartz between two layers of steel. Scientific American, 278(4):106–107, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0498issue/0498currentissue.html.

Burke:1998:SP

[4337] James Burke. Sheer poetry. Scientific American, 278(4):106–107, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-106.pdf.

Goldberg:1998:WK

[4338] Rube Goldberg. Working knowledge. Scientific American, 278(4):108, April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n4/pdf/scientificamerican0498-108.pdf.

Goldberg:1998:WKR

[4339] Rube Goldberg. Working knowledge: Rube Goldberg’s automatic napkin. Scientific American, 278(4):108–108 (Intl. ed. 92–??), April 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0498issue/0498currentissue.html.

Rennie:1998:EOO

[4340] John Rennie. From the Editor: Outsmarting our genes. Scientific American, 278(5):8, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
8.pdf.

Anonymous:1998:Ec

[4341] Anonymous. From the editors. Scientific American, 278(5):8–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Anonymous:1998:Ed

[4342] Anonymous. Errata. Scientific American, 278(5):10, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-10b.pdf.

Anonymous:1998:LEe

[4343] Anonymous. Letters to the Editors. Scientific American, 278(5):10, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-10a.pdf;   http:
//www.sciam.com/1998/0598issue/0598currentissue.html.

Anonymous:1998:YAIe

[4344] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 278(5):12, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
12.pdf.

Anonymous:1998:YAe

[4345] Anonymous. 50, 100 and 150 years ago. Scientific American, 278 (5):12–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Gibbs:1998:NAT

[4346] W. Wayt Gibbs. News and analysis: Taking aim at tumors. Scientific American, 278(5):17–20, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
17.pdf.

Anonymous:1998:Fc

[4347] Anonymous. In focus. Scientific American, 278(5):17–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0598issue/0598currentissue.html.

Anonymous:1998:FND

[4348] Anonymous. In focus: Nuclear detonations improve radiotherapies against cancer. Scientific American, 278(5):17–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0598issue/0598currentissue.html.

Yam:1998:MDI

[4349] Philip Yam. Making a deep impact. Scientific American, 278(5):22– 24, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-22.pdf.

Anonymous:1998:SCH

[4350] Anonymous. Science and the citizen: How not to save the world from asteroids. the future gets old. blindingly fast beetles. Scientific American, 278(5):22–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Gibbs:1998:SM

[4351] W. Wayt Gibbs. Snow men. Scientific American, 278(5):24–26, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-24b.pdf.

Anonymous:1998:Bc

[4352] Anonymous. In brief. Scientific American, 278(5):24–28, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-24a.pdf.

Schneider:1998:WWE

[4353] David Schneider. Water, water everywhere. Scientific American, 278(5): 26–29, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-26.pdf.

Mirsky:1998:NYS

[4354] Steve Mirsky. Now you see it, now you don’t. Scientific American, 278(5): 27, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-27.pdf.

McKinsey:1998:DW

[4355] Krista McKinsey. Dances of worms. Scientific American, 278(5):28, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-28.pdf.

Doyle:1998:FO

[4356] Rodger Doyle. The future of the old. Scientific American, 278(5):29, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-29.pdf.

Beardsley:1998:RNG

[4357] Tim Beardsley. Rebottling the nuclear genie. Scientific American, 278(5): 34–35, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-34.pdf.

Anonymous:1998:PTB

[4358] Anonymous. Profile: Thomas B. Cochran, nuclear activist, fights bombs with information. Scientific American, 278(5):34–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0598issue/0598currentissue.html.

Randal:1998:WLM

[4359] Judith Randal. When less is more. Scientific American, 278(5):36– 37, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-36.pdf.

Anonymous:1998:TBD

[4360] Anonymous. Technology and business: Downsizing organs. saving a biopesticide. electronic tongue. Scientific American, 278(5): 36–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Stix:1998:RF

[4361] Gary Stix. Resistance fighting. Scientific American, 278(5):38, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-38.pdf.

Zorpette:1998:SL

[4362] Glenn Zorpette. Seeing the light. Scientific American, 278(5):39–40, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-39.pdf.

Gibbs:1998:TL

[4363] W. Wayt Gibbs. A tongue for love. Scientific American, 278(5):40– 44, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-40.pdf.

Alpert:1998:CI

[4364] Mark Alpert. A cool idea. Scientific American, 278(5):44, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-44.pdf.

Grossman:1998:BI

[4365] Wendy M. Grossman. Bringing down the Internet. Scientific American, 278(5):45, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
45.pdf.

Anonymous:1998:CVHa

[4366] Anonymous. Cyber view: How to kill the Internet. Scientific American, 278(5):45–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Lucid:1998:SMM

[4367] Shannon W. Lucid. Six months on Mir. Scientific American, 278(5):46– 55, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-46.pdf;   http:/
/www.sciam.com/1998/0508issue/0598lucid.html.

Bennet-Clark:1998:HCM

[4368] Henry C. Bennet-Clark. How cicadas make their noise. Scientific American, 278(5):58–61, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
58.pdf;   http://www.sciam.com/1998/0598issue/0598currentissue.
html.

Plomin:1998:GCA

[4369] Robert Plomin and John C. DeFries. The genetics of cognitive abilities and disabilities. Scientific American, 278(5):62–69, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-62.pdf;   http:/
/www.sciam.com/1998/0598issue/0598currentissue.html.

Sobel:1998:NST

[4370] Alan Sobel. The new shape of television: Television’s bright new technology. Scientific American, 278(5):70–77, May 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-70.pdf.

Sobel:1998:TBN

[4371] Alan Sobel. Television’s bright new technology. Scientific American, 278(5):70–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Lim:1998:DTH

[4372] Jae S. Lim. Digital television: Here at last. Scientific American, 278(5): 78–83, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-78.pdf;   http:/
/www.sciam.com/1998/0598issue/0598currentissue.html.

Fukagawa:1998:JTG

[4373] Tony Rothmanwith the cooperation of Hidetoshi Fukagawa. Japanese temple geometry. Scientific American, 278(5):84–91, May 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-84.pdf.

Rothman:1998:JTG

[4374] Tony Rothman and Hidetoshi Fukagawa. Japanese temple geometry. Scientific American, 278(5):84–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0508issue/0598rothman.html.

Stix:1998:TEC

[4375] Gary Stix. Trends in economics: a calculus of risk. Scientific American, 278(5):92–97, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
92.pdf.

Stix:1998:CR

[4376] Gary Stix. A calculus of risk. Scientific American, 278(5):92– ??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Carlson:1998:ASSa

[4377] Shawn Carlson. The amateur scientist: Sensing subtle tsunamis. Scientific American, 278(5):98–99, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
98.pdf.

Anonymous:1998:ASM

[4378] Anonymous. The amateur scientist: Measuring atmospheric tsunamis. Scientific American, 278(5):98–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0598issue/0598currentissue.html.

Stewart:1998:MRCa

[4379] Ian Stewart. Mathematical recreations: Cementing relationships. Scientific American, 278(5):100–103, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
100.pdf.

Anonymous:1998:MRc

[4380] Anonymous. Mathematical recreations. Scientific American, 278(5): 100–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Anonymous:1998:WPP

[4381] Anonymous. What piles of powder explain about astronomy. Scientific American, 278(5):100–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0598issue/0598currentissue.html.

Anonymous:1998:RCd

[4382] Anonymous. Reviews and commentaries. Scientific American, 278(5):104–107, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
104.pdf.

Anonymous:1998:RCe

[4383] Anonymous. Reviews and commentaries. Scientific American, 278 (5):104–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0598issue/
0598currentissue.html.

Burke:1998:CSH

[4384] James Burke. Connections: The sound heard ’round the world. Scientific American, 278(5):104–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0598issue/0598currentissue.html.

Darwin:1998:EFO

[4385] Charles Darwin. The evolution of the frown and the origin of smiles. Scientific American, 278(5):104–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0598issue/0598currentissue.html.

Morrison:1998:DBR

[4386] Philip Morrison. Double bass redoubled. Scientific American, 278(5):109–111, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n5/pdf/scientificamerican0598-
109.pdf.

Morrison:1998:WDB

[4387] Philip Morrison. Wonders: Double bass redoubled. Scientific American, 278(5):109, 111 (Intl. ed. 104–??), May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0598issue/0598currentissue.html.

Burke:1998:Ca

[4388] James Burke. Cheers. Scientific American, 278(5):110–111, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n5/pdf/scientificamerican0598-110.pdf.

Dragon:1998:WK

[4389] Elizabeth A. Dragon. Working knowledge. Scientific American, 278(5): 112, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n5/pdf/scientificamerican0598-112.pdf.

Anonymous:1998:WKM

[4390] Anonymous. Working knowledge: Making DNA haystacks with PCR. Scientific American, 278(5):112–??, May 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0598issue/0598currentissue.html.

Rennie:1998:EJF

[4391] John Rennie. From the Editor: Not just a fish story. Scientific American, 278(6):6, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
6.pdf.

Anonymous:1998:Ee

[4392] Anonymous. From the editors. Scientific American, 278(6):6–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0698issue/
0698currentissue.html.

Anonymous:1998:Ef

[4393] Anonymous. Errata. Scientific American, 278(6):8, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n6/pdf/scientificamerican0698-8A.pdf.

Anonymous:1998:LEf

[4394] Anonymous. Letters to the Editors. Scientific American, 278(6):8–8A, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-8.pdf;   http://
www.sciam.com/1998/0698issue/0698currentissue.html.

Anonymous:1998:YAIf

[4395] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 278(6):9, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
9.pdf.

Anonymous:1998:YAf

[4396] Anonymous. 50, 100 and 150 years ago. Scientific American, 278 (6):9–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0698issue/
0698currentissue.html.

Beardsley:1998:NAC

[4397] Tim Beardsley. News and analysis: Culturing new life. Scientific American, 278(6):11–12, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
11.pdf.

Anonymous:1998:FCS

[4398] Anonymous. In focus: Cultured stem cells might repair damaged brains, blood and more. Scientific American, 278(6):11–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0698issue/0698currentissue.html.

Anonymous:1998:NA

[4399] Anonymous. News and analysis. Scientific American, 278(6):11– ??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0698issue/
0698currentissue.html.

McKinsey:1998:LL

[4400] Krista McKinsey. Lupus in limbo. Scientific American, 278(6):13, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-13.pdf.

Anonymous:1998:SCW

[4401] Anonymous. Science and the citizen: Wolves in limbo… greenhouse methane is disappearing — why?… Tamoxifen and breast cancer… crabby mates. Scientific American, 278(6):13–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0698issue/0698currentissue.html.

Schneider:1998:GNG

[4402] David Schneider. Good news for the greenhouse. Scientific American, 278(6):14, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
14a.pdf.

Zorpette:1998:MMT

[4403] Glenn Zorpette. Mesozoic mystery tour. Scientific American, 278(6):14– 18, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-14b.pdf.

Leutwyler:1998:Bc

[4404] Kristin Leutwyler. In brief. Scientific American, 278(6):18–28, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-18.pdf.

Nelson:1998:BB

[4405] Roxanne Nelson. The blight is back. Scientific American, 278(6):20–26, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-20.pdf.

Doyle:1998:UW

[4406] Rodger Doyle. U.S. wetlands. Scientific American, 278(6):24–25, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-24.pdf.

Gibbs:1998:PP

[4407] W. Wayt Gibbs. The prevention pill. Scientific American, 278(6):26–27, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-26.pdf.

Mirsky:1998:GTU

[4408] Steve Mirsky. Greatness thrust upon it. Scientific American, 278(6):28, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-28.pdf.

Stix:1998:WBL

[4409] Gary Stix. Where the bodies lie. Scientific American, 278(6):30–32, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-30.pdf.

Anonymous:1998:PEJ

[4410] Anonymous. Profile: Epidemiologist Joel Schwartz warns of dangerous particles in the air. Scientific American, 278(6):30–??, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0698issue/0698currentissue.html.

Anonymous:1998:Pc

[4411] Anonymous. Profile. Scientific American, 278(6):30–??, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0698issue/0698currentissue.html.

Dupont:1998:SW

[4412] Daniel G. Dupont. Star warned. Scientific American, 278(6):34, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-34.pdf.

Anonymous:1998:TBF

[4413] Anonymous. Technology and business: A fast Y2K bug fix. confidentiality without encryption… Earthcam.com. Scientific American, 278 (6):34–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0698issue/
0698currentissue.html.

McKinsey:1998:RM

[4414] Krista McKinsey. Running on MMT? Scientific American, 278(6):35, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-35.pdf.

Hayashi:1998:MBZ

[4415] Alden M. Hayashi. Millennium bug zapper. Scientific American, 278(6): 36, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-36a.pdf.

Eisenberg:1998:CY

[4416] Anne Eisenberg. Confidentially yours. Scientific American, 278(6):36– 37, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-36b.pdf.

Beardsley:1998:ES

[4417] Tim Beardsley. Eye in the sky. Scientific American, 278(6):37, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n6/pdf/scientificamerican0698-37.pdf.

Gibbs:1998:CVW

[4418] W. Wayt Gibbs. Cyber view: The Web learns to read. Scientific American, 278(6):40, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0698issue/0698currentissue.html. Discusses HTML, MathML, Mu-

sicML, ChemML, and XML.

Gibbs:1998:WLR

[4419] W. Wayt Gibbs. The Web learns to read. Scientific American, 278(6): 40, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-40.pdf.

Anonymous:1998:CV

[4420] Anonymous. Cyber view. Scientific American, 278(6):40–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0698issue/0698currentissue.html.

Anonymous:1998:CVN

[4421] Anonymous. Cyber view: A new language for the Web. Scientific American, 278(6):40–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0698issue/0698currentissue.html.

Nemeroff:1998:ND

[4422] Charles B. Nemeroff. The neurobiology of depression. Scientific American, 278(6):42–49, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
42.pdf;   http://www.sciam.com/1998/0608issue/0698nemeroff.
html.

Disney:1998:NLQ

[4423] Michael Disney. A new look at quasars. Scientific American, 278(6):52– 57, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-52.pdf;   http:/
/www.sciam.com/1998/0698issue/0698currentissue.html.

Boyd:1998:SAE

[4424] Claude E. Boyd and Jason W. Clay. Shrimp aquaculture and the environment. Scientific American, 278(6):58–65, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n6/pdf/scientificamerican0698-58.pdf;   http://www.sciam.com/
1998/0698issue/0698currentissue.html.

Gershenfeld:1998:QCM

[4425] Neil Gershenfeld and Isaac L. Chuang. Quantum computing with molecules. Scientific American, 278(6):66–71, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n6/pdf/scientificamerican0698-66.pdf;   http://www.sciam.com/
1998/0608issue/0698gershenfeld.html.

Bell:1998:GG

[4426] Robin E. Bell. Gravity gradiometry. Scientific American, 278(6):74–79, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-74.pdf;   http:/
/www.sciam.com/1998/0698issue/0698currentissue.html.

Vallee:1998:AWW

[4427] Bert L. Vallee. Alcohol in the Western World. Scientific American, 278(6):80–85, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
80.pdf;   http://www.sciam.com/1998/0698issue/0698currentissue.
html.

Eisenberg:1998:DSL

[4428] Mickey S. Eisenberg. Defibrillation: The spark of life. Scientific American, 278(6):86–90, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
86.pdf;   http://www.sciam.com/1998/0698issue/0698currentissue.
html.

Bartecchi:1998:IYD

[4429] Carl E. Bartecchi. If you don’t have a defibrillator. Scientific American, 278(6):91, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
91.pdf.

Carlson:1998:ASW

[4430] Shawn Carlson. The amateur scientist: Waiter, there’s a hair in my hygrometer. Scientific American, 278(6):92–94, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v278/
n6/pdf/scientificamerican0698-92.pdf.

Anonymous:1998:ASW

[4431] Anonymous. The amateur scientist: What humidity does to hair. Scientific American, 278(6):92–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0698issue/0698currentissue.html.

Stewart:1998:MRW

[4432] Ian Stewart. Mathematical recreations: What a coincidence! Scientific American, 278(6):95–96, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
95.pdf.

Anonymous:1998:MRH

[4433] Anonymous. Mathematical recreations: Happy birthday to you… and you. Scientific American, 278(6):95–??, June 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0698issue/0698currentissue.html.

Anonymous:1998:RCf

[4434] Anonymous. Reviews and commentaries. Scientific American, 278(6):97– 98, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-97.pdf.

Anonymous:1998:RCg

[4435] Anonymous. Reviews and commentaries. Scientific American, 278 (6):97–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0698issue/
0698currentissue.html.

Burke:1998:CDR

[4436] James Burke. Connections: Doppler’s rising pitch, laughing gas and handwriting analysis. Scientific American, 278(6):97–??, June 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0698issue/0698currentissue.html.

Morrison:1998:WNN

[4437] Philip Morrison. Wonders: Natural nuclear reactors. Scientific American, 278(6):97–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0698issue/0698currentissue.html.

Wilson:1998:RCW

[4438] O. Wilson. Reviews and commentaries: E. O. Wilson of sociobiology fame tries to unify the sciences and humanities. Scientific American, 278(6):97–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0698issue/
0698currentissue.html.

Morrison:1998:WFB

[4439] Philip Morrison. Where fiction became ancient fact. Scientific American, 278(6):99–101, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
99.pdf.

Burke:1998:SS

[4440] James Burke. Scribble, scribble. Scientific American, 278(6):100–101, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v278/n6/pdf/scientificamerican0698-100.pdf.

McCreary:1998:WKD

[4441] Michael D. McCreary. Working knowledge: Digital cameras. Scientific American, 278(6):102, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v278/n6/pdf/scientificamerican0698-
102.pdf.

Anonymous:1998:WKP

[4442] Anonymous. Working knowledge: Photos without film: how digital cameras work. Scientific American, 278(6):102–??, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0698issue/0698currentissue.html.

Rennie:1998:EAO

[4443] John Rennie. From the Editors: All for one. Scientific American, 279 (1):8, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-8.pdf.

Anonymous:1998:Eg

[4444] Anonymous. From the editors. Scientific American, 279(1):8–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/0798issue/
0798currentissue.html.

Anonymous:1998:Eh

[4445] Anonymous. Errata. Scientific American, 279(1):10, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-10a.pdf.

Anonymous:1998:LEg

[4446] Anonymous. Letters to the Editors. Scientific American, 279(1):10, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-10.pdf;   http:/
/www.sciam.com/1998/0798issue/0798currentissue.html.

Anonymous:1998:YAIg

[4447] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 279(1):12, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
12.pdf.

Anonymous:1998:YAg

[4448] Anonymous. 50, 100 and 150 years ago. Scientific American, 279 (1):12–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0798issue/
0798currentissue.html.

Gibbs:1998:NAS

[4449] W. Wayt Gibbs. News and analysis: Statistical uncertainty. Scientific American, 279(1):17–18, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
17.pdf.

Anonymous:1998:FCC

[4450] Anonymous. In focus: Census without consensus: a political fight over how to count. Scientific American, 279(1):17–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0798issue/0798currentissue.html.

Anonymous:1998:Fd

[4451] Anonymous. In focus. Scientific American, 279(1):17–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0798issue/0798currentissue.html.

Musser:1998:IDL

[4452] George Musser. Inflation is dead; long live inflation. Scientific American, 279(1):19–20, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
19.pdf.

Anonymous:1998:SCI

[4453] Anonymous. Science and the citizen: Inflation chases the expanding universe. virtual anthropology. miscarried males. the Earth drags space. arms imports. Scientific American, 279(1):19–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0798issue/0798currentissue.html.

Leutwyler:1998:Bd

[4454] Kristin Leutwyler. In brief. Scientific American, 279(1):20–28, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-20.pdf.

Wong:1998:F

[4455] Kate Wong. Face off. Scientific American, 279(1):21–22, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-21.pdf.

Alpert:1998:WAB

[4456] Mark Alpert. Where have all the boys gone? Scientific American, 279(1): 22–24, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-22.pdf.

Ariza:1998:ED

[4457] Luis Miguel Ariza. Einstein’s drag. Scientific American, 279(1):24– 25, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-24.pdf.

Mirsky:1998:GOM

[4458] Steve Mirsky. Gorilla in our midst. Scientific American, 279(1):28, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-28.pdf.

Doyle:1998:AT

[4459] Rodger Doyle. The arms trade. Scientific American, 279(1):29, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-29.pdf.

Gibbs:1998:BTW

[4460] W. Wayt Gibbs. Big Tobacco’s worst nightmare. Scientific American, 279(1):30–31, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
30.pdf.

Anonymous:1998:Pd

[4461] Anonymous. Profile. Scientific American, 279(1):30–??, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0798issue/0798currentissue.html.

Anonymous:1998:PSSb

[4462] Anonymous. Profile: Scientist Stanton A. Glantz spoke, and the tobacco industry fumed. Scientific American, 279(1):30–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0798issue/0798currentissue.html.

Beardsley:1998:HC

[4463] Tim Beardsley. Hot coolants. Scientific American, 279(1):32, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-32.pdf.

Anonymous:1998:TBB

[4464] Anonymous. Technology and business: Black market in CFCs. America’s Cup racers sail into the lab. computers that feel your pain. Richter-scale models. fusion plasma spirals. Scientific American, 279 (1):32–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0798issue/
0798currentissue.html.

Leutwyler:1998:TT

[4465] Kristin Leutwyler. Tempest in a teacup. Scientific American, 279(1): 34, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-34a.pdf.

Beardsley:1998:GR

[4466] Tim Beardsley. Getting real? Scientific American, 279(1):34–36, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-34b.pdf.

Schneider:1998:TS

[4467] David Schneider. Tectonics in a sandbox. Scientific American, 279(1): 36, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-36.pdf.

Gibbs:1998:NTF

[4468] W. Wayt Gibbs. A new twist in fusion. Scientific American, 279(1): 37, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-37b.pdf.

Hayashi:1998:LC

[4469] Alden M. Hayashi. Lost in cyberspace. Scientific American, 279(1): 37, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-37a.pdf.

Golombek:1998:MPM

[4470] Matthew P. Golombek. The Mars Pathfinder Mission. Scientific American, 279(1):40–49, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
40.pdf;   http://www.sciam.com/1998/0798issue/0798currentissue.
html.

Gazzaniga:1998:SBR

[4471] Michael S. Gazzaniga. The split brain revisited. Scientific American, 279(1):50–55, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
50.pdf;   http://www.sciam.com/1998/0798issue/0798currentissue.
html.

Feld:1998:SAL

[4472] Michael S. Feld and Kyungwon An. The single-atom laser. Scientific American, 279(1):56–59, 62–63, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
56.pdf;   http://www.sciam.com/1998/0798issue/0798currentissue.
html.

Rutowski:1998:MSB

[4473] Ronald L. Rutowski. Mating strategies in butterflies. Scientific American, 279(1):64–69, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
64.pdf;   http://www.sciam.com/1998/0708issue/0798rutowski.
html.

Tobin:1998:LF

[4474] William Tobin. Léon Foucault. Scientific American, 279(1):70–77, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-70.pdf;   http:/
/www.sciam.com/1998/0798issue/0798currentissue.html.

Editors:1998:SRD

[4475] The Editors. Special report: Defeating AIDS: What will it take? Scientific American, 279(1):81, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
81.pdf.

Anonymous:1998:DAW

[4476] Anonymous. Defeating AIDS: What will it take?: Infections with HIV, the human immunodeficiency virus that causes AIDS, continue to sweep the world. cures and vaccines remain elusive, although the search goes on. the good news is that safer behaviors and — for those with access to proper care — better drug treatments and tests can save or extend lives. these leading investigators describe the state of the fight against HIV today and the prospects for winning. Scientific American, 279(1):81–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0798issue/
0798currentissue.html.

Mann:1998:HGP

[4477] Jonathan M. Mann and Daniel J. M. Tarantola. HIV 1998: The global picture. Scientific American, 279(1):82–83, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-82.pdf;   http://www.sciam.com/
1998/0798issue/0798currentissue.html.

Bartlett:1998:IHT

[4478] John G. Bartlett and Richard D. Moore. Improving HIV therapy. Scientific American, 279(1):84–87, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
84.pdf;   http://www.sciam.com/1998/0708issue/0798bartlett.
html.

Richman:1998:HDR

[4479] Douglas D. Richman. How drug resistance arises. Scientific American, 279(1):88–89, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
88.pdf;   http://www.sciam.com/1998/0708issue/0798bartlettbox1.
html.

Mellors:1998:VLT

[4480] John W. Mellors. Viral-load tests provide valuable answers. Scientific American, 279(1):90–93, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
90.pdf;   http://www.sciam.com/1998/0708issue/0798bartlettbox2.
html.

Wilfert:1998:WCH

[4481] Catherine M. Wilfert and Ross E. McKinney, Jr. When children harbor HIV. Scientific American, 279(1):94–95, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-94.pdf;   http://www.sciam.com/
1998/0798issue/0798currentissue.html.

Coates:1998:PHI

[4482] Thomas J. Coates and Chris Collins. Preventing HIV infection. Scientific American, 279(1):96–97, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
96.pdf;   http://www.sciam.com/1998/0798issue/0798currentissue.
html.

Baltimore:1998:HVP

[4483] David Baltimore and Carole Heilman. HIV vaccines: Prospects and challenges. Scientific American, 279(1):98–103, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-98.pdf;   http://www.sciam.com/
1998/0798issue/0798currentissue.html.

Buchbinder:1998:AIA

[4484] Susan Buchbinder. Avoiding infection after HIV exposure. Scientific American, 279(1):104–105, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
104.pdf;   http://www.sciam.com/1998/0798issue/0798currentissue.
html.

Beardsley:1998:CHE

[4485] Tim Beardsley. Coping with HIV’s ethical dilemmas. Scientific American, 279(1):106–107, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
106.pdf;   http://www.sciam.com/1998/0798issue/0798currentissue.
html.

Carlson:1998:ASY

[4486] Shawn Carlson. The amateur scientist: a year for the oceans. Scientific American, 279(1):108–109, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
108.pdf.

Anonymous:1998:ASR

[4487] Anonymous. The amateur scientist: Recreational divers lend a fin to marine biologists. Scientific American, 279(1):108–??, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0798issue/0798currentissue.html.

Stewart:1998:MRBb

[4488] Ian Stewart. Mathematical recreations: The Bellows Conjecture. Scientific American, 279(1):110–111, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
110.pdf.

Anonymous:1998:MRF

[4489] Anonymous. Mathematical recreations: Finding exceptions to an “inflexible” rule of geometry. Scientific American, 279(1):110–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0798issue/0798currentissue.html.

Anonymous:1998:RCh

[4490] Anonymous. Reviews and commentaries. Scientific American, 279(1):113–114, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
113.pdf.

Anonymous:1998:RCi

[4491] Anonymous. Reviews and commentaries. Scientific American, 279 (1):113–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0798issue/
0798currentissue.html.

Burke:1998:CIW

[4492] James Burke. Connections: From Izaak Walton to Isaac Newton. Scientific American, 279(1):113–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0798issue/0798currentissue.html.

Dennett:1998:RCB

[4493] Daniel C. Dennett. Reviews and commentaries: In brainchildren, Daniel C. Dennett argues that philosophers of the mind need to loosen up. Scientific American, 279(1):113–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0798issue/0798currentissue.html.

Morrison:1998:WTS

[4494] Philip Morrison and Phylis Morrison. Wonders: Taking the sum of all human knowledge. Scientific American, 279(1):113–??, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/0798issue/0798currentissue.html.

Morrison:1998:SHK

[4495] Philip Morrison and Phylis Morrison. The sum of human knowledge? Scientific American, 279(1):115–117, July 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n1/pdf/scientificamerican0798-115.pdf.

Burke:1998:HS

[4496] James Burke. Heavy stuff. Scientific American, 279(1):116–117, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n1/pdf/scientificamerican0798-116.pdf.

Gerpheide:1998:WKT

[4497] George Gerpheide. Working knowledge: Touchpad pointing device. Scientific American, 279(1):118, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n1/pdf/scientificamerican0798-
118.pdf.

Anonymous:1998:WKC

[4498] Anonymous. Working knowledge: Computer touchpads get the point. Scientific American, 279(1):118–??, July 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0798issue/0798currentissue.html.

Anonymous:1998:YAh

[4499] Anonymous. 50, 100 and 150 years ago. Scientific American, 279 (2):6–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0898issue/
0898currentissue.html.

Rennie:1998:EBY

[4500] John Rennie. From the Editors: Believing your eyes. Scientific American, 279(2):7, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
7.pdf.

Anonymous:1998:FHI

[4501] Anonymous. In focus: How India and Pakistan got the bomb. Scientific American, 279(2):7–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Anonymous:1998:LEh

[4502] Anonymous. Letters to the Editors. Scientific American, 279(2):8–9, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-8.pdf.

Anonymous:1998:Ei

[4503] Anonymous. Errata. Scientific American, 279(2):9, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-9.pdf.

Anonymous:1998:SCN

[4504] Anonymous. Science and the citizen: Neutrino mass detected. the second-biggest bang? the hungry among us. science and theology. mutant athletes. Scientific American, 279(2):9–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0898issue/0898currentissue.html.

Anonymous:1998:YAIh years ago. Scientific American, 279(2):10, August 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-10.pdf.

Anonymous:1998:PJC

[4506] Anonymous. Profile: J. Craig Venter of the Institute for Genomic Research races to sequence human DNA. Scientific American, 279(2): 14–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0898issue/
0898currentissue.html.

Zorpette:1998:NAD

[4507] Glenn Zorpette and Tim Beardsley. News and analysis: Deadly secrets. Scientific American, 279(2):15–16, August 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-15.pdf.

Gibbs:1998:MD

[4508] W. Wayt Gibbs. A massive discovery. Scientific American, 279(2):18–20, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-18.pdf.

Anonymous:1998:NAR

[4509] Anonymous. News and analysis: In a regulatory soup. Scientific American, 279(2):18–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0898issue/0898currentissue.html.

Anonymous:1998:TBHb

[4510] Anonymous. Technology and business: How trade treaties weaken environmental laws. the euro bugs computers. unions and productivity. Scientific American, 279(2):18–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Gibbs:1998:BP

[4511] W. Wayt Gibbs. Beyond physics. Scientific American, 279(2):20–22, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-20.pdf.

Anonymous:1998:CVW

[4512] Anonymous. Cyber view: Women exit computer science. Scientific American, 279(2):21–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Mirsky:1998:HG

[4513] Steve Mirsky. Hoop genes. Scientific American, 279(2):22, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-22b.pdf.

Leutwyler:1998:Be

[4514] Kristin Leutwyler. In brief. Scientific American, 279(2):22–26, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-22a.pdf.

Yonas:1998:FZPa

[4515] Gerold Yonas. Fusion and the Z-pinch. Scientific American, 279(2): 22–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0808issue/
0898yonas.html.

Nemecek:1998:FT

[4516] Sasha Nemecek. Food for thought. Scientific American, 279(2):23–24, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-23.pdf.

Musser:1998:BLB

[4517] George Musser. Bright lights, big mystery. Scientific American, 279(2):24–26, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
24.pdf.

Doyle:1998:AR

[4518] Rodger Doyle. Amphibians at risk. Scientific American, 279(2):27, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-27.pdf.

Deyo:1998:LBPa

[4519] Richard A. Deyo. Low-back pain. Scientific American, 279(2):28–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1998:ERG

[4520] Tim Beardsley. An express route to the genome? Scientific American, 279(2):30–32, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
30.pdf.

Holloway:1998:TR

[4521] Marguerite Holloway. Trade rules. Scientific American, 279(2):33–35, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-33.pdf.

Hayashi:1998:OCP

[4522] Alden M. Hayashi. The other computer problem. Scientific American, 279(2):34–36, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
34.pdf.

Adleman:1998:CDa

[4523] Leonard M. Adleman. Computing with DNA. Scientific American, 279 (2):34–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wallich:1998:LUL

[4524] Paul Wallich. Look for the union label. Scientific American, 279(2):36– 38, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-36.pdf.

Grossman:1998:AD

[4525] Wendy M. Grossman. Access denied. Scientific American, 279(2):38, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-38.pdf.

Yonas:1998:FZPb

[4526] Gerold Yonas. Fusion and the Z pinch. Scientific American, 279(2):40– 47, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-40.pdf.

Johnson:1998:MCDa

[4527] Nicholas L. Johnson. Monitoring and controlling debris in space: Dead satellites, dismembered rockets, drifting bolts, flecks of paint and more dangerous junk circle the …. Scientific American, 279(2): 42–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0898issue/
0898currentissue.html.

Deyo:1998:LBPb

[4528] Richard A. Deyo. Low-back pain. Scientific American, 279(2):48–53, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-48.pdf.

Gardner:1998:QCRa

[4529] Martin Gardner. A quarter-century of recreational mathematics. Scientific American, 279(2):48–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0808issue/0898math.html.

Adleman:1998:CDb

[4530] Leonard M. Adleman. Computing with DNA. Scientific American, 279(2):54–61, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
54.pdf.

Glenn:1998:ICSa

[4531] Edward P. Glenn, J. Jed Brown, and James W. O’Leary. Irrigating crops with seawater: Seawater is lethal to conventional crops, yet salt-tolerant plants—suitable for food, animal forage and oilseed—can. … Scientific American, 279(2):56–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0808issue/0898glenn.html.

Johnson:1998:MCDb

[4532] Nicholas L. Johnson. Monitoring and controlling debris in space. Scientific American, 279(2):62–67, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
62.pdf.

Trautman:1998:Ma

[4533] Rachael Trautman, Brendan J. Griffin, and David Scharf. Microdiamonds. Scientific American, 279(2):62–??, August 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gardner:1998:QCRb

[4534] Martin Gardner. A quarter-century of recreational mathematics. Scientific American, 279(2):68–75, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
68.pdf.

Foster:1998:PYFa

[4535] Kenneth R. Foster, Mary F. Jenkins, and Anna Coxe Toogood. The Philadelphia Yellow Fever Epidemic of 1793. Scientific American, 279 (2):68–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0898issue/
0898currentissue.html.

Anonymous:1998:ASD

[4536] Anonymous. The amateur scientist: Dream a little stream with me. Scientific American, 279(2):74–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Glenn:1998:ICSb

[4537] Edward P. Glenn, J. Jed Brown, and James W. O’Leary. Irrigating crops with seawater. Scientific American, 279(2):76–81, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-76.pdf.

Anonymous:1998:MRI

[4538] Anonymous. Mathematical recreations: If he knows that she knows that he knows. Scientific American, 279(2):76–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0898issue/0898currentissue.html.

Anonymous:1998:RCS

[4539] Anonymous. Reviews and commentaries: Secrets behind the revival of the U.S. industrial economy. the bottomline value of biodiversity. Scientific American, 279(2):78–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Burke:1998:CCM

[4540] James Burke. Connections: Clockmakers, mapmakers and Wanamaker’s. Scientific American, 279(2):78–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

sciam.com/1998/0898issue/0898currentissue.html.

Morrison:1998:WMM

[4541] Philip Morrison and Phylis Morrison. Wonders: Mirror molecules. Scientific American, 279(2):78–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Trautman:1998:Mb

[4542] Rachael Trautman, Brendan J. Griffin, and David Scharf. Microdiamonds. Scientific American, 279(2):82–87, August 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-82.pdf.

Anonymous:1998:WKHb

[4543] Anonymous. Working knowledge: How to breathe underwater. Scientific American, 279(2):84–??, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0898issue/0898currentissue.html.

Foster:1998:PYFb

[4544] Kenneth R. Foster, Mary F. Jenkins, and Anna Coxe Toogood. The Philadelphia Yellow Fever Epidemic of 1793. Scientific American, 279(2):88–93, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
88.pdf.

Carlson:1998:ASBb

[4545] Shawn Carlson. The amateur scientist: Building a consciousness of streams. Scientific American, 279(2):94–95, August 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-94.pdf.

Stewart:1998:MRM

[4546] Ian Stewart. Mathematical recreations: Monks, blobs and common knowledge. Scientific American, 279(2):96–97, August 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n2/pdf/scientificamerican0898-96.pdf.

Anonymous:1998:RCj

[4547] Anonymous. Reviews and commentaries. Scientific American, 279(2):98–100, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
98.pdf.

Morrison:1998:LR

[4548] Philip Morrison and Phylis Morrison. The Left and the Right. Scientific American, 279(2):101–103, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
101.pdf.

Burke:1998:TT

[4549] James Burke. Tick tock. Scientific American, 279(2):102–103, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n2/pdf/scientificamerican0898-102.pdf.

Bowden:1998:WKS

[4550] William A. Bowden. Working knowledge: Scuba regulators. Scientific American, 279(2):104, August 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n2/pdf/scientificamerican0898-
104.pdf.

Rennie:1998:EGA

[4551] John Rennie. From the Editors: Go ahead, walk in the mud. Scientific American, 279(3):7, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
7.pdf.

Anonymous:1998:LEi

[4552] Anonymous. Letters to the Editors. Scientific American, 279(3):8–9, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-8.pdf.

Anonymous:1998:FVC

[4553] Anonymous. In focus: Violence in the classroom proves hard to prevent. Scientific American, 279(3):8–??, September 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0998issue/0998currentissue.html.

Anonymous:1998:SCRb

[4554] Anonymous. Science and the citizen: Resolving the universe. medicinal marijuana? A river of acid. damnable weather. the burning season. Scientific American, 279(3):11–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0998issue/0998currentissue.html.

Anonymous:1998:YAIi

[4555] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 279(3):12, September 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n3/pdf/scientificamerican0998-12.pdf.

Nemecek:1998:NAF

[4556] Sasha Nemecek. News and analysis: Forestalling violence. Scientific American, 279(3):15–16, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
15.pdf.

Stix:1998:HR

[4557] Gary Stix. Herb remedy. Scientific American, 279(3):18–19, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-18.pdf.

Mukerjee:1998:SWD

[4558] Madhusree Mukerjee. Stalking the wild dugong. Scientific American, 279(3):19–20, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
19.pdf.

Anonymous:1998:PRL

[4559] Anonymous. Profile: Rolf Landauer seeks the physical limits of computation. Scientific American, 279(3):20–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0998issue/0998currentissue.html.

Musser:1998:FSU

[4560] George Musser. The flip side of the Universe. Scientific American, 279(3):22, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
22b.pdf.

Leutwyler:1998:Bf

[4561] Kristin Leutwyler. In brief. Scientific American, 279(3):22–26, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-22a.pdf.

Anonymous:1998:TBE

[4562] Anonymous. Technology and business: Electronic paper. the first cancer vaccine arrives. holographic memories that hold up. Scientific American, 279(3):22–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0998issue/0998currentissue.html.

Mirsky:1998:TPF

[4563] Steve Mirsky. Tomorrow, partly froggy. Scientific American, 279(3):24, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-24.pdf.

Anonymous:1998:CVHb

[4564] Anonymous. Cyber view: Headaches from inscrutable computers. Scientific American, 279(3):25–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0998issue/0998currentissue.html.

Nemecek:1998:WA

[4565] Sasha Nemecek. A world aflame. Scientific American, 279(3):26, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-26a.pdf.

Ariza:1998:RV

[4566] Luis Miguel Ariza. River of vitriol. Scientific American, 279(3):26–28, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-26b.pdf.

Agnew:1998:PLFa

[4567] Neville Agnew and Martha Demas. Preserving the Laetoli footprints. Scientific American, 279(3):26–??, September 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0998issue/0998currentissue.html.

Doyle:1998:EGW

[4568] Rodger Doyle. Ethnic groups in the world. Scientific American, 279(3):30, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
30.pdf.

Stix:1998:RBE

[4569] Gary Stix. Riding the back of electrons. Scientific American, 279(3):32–34, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
32.pdf.

Gibbs:1998:RP

[4570] W. Wayt Gibbs. The reinvention of paper. Scientific American, 279(3):36–40, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
36.pdf.

White:1998:WHBa

[4571] Ronald J. White. Weightlessness and the human body. Scientific American, 279(3):38–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0908issue/0998white.html.

Gibbs:1998:HC

[4572] W. Wayt Gibbs. Healing cancer. Scientific American, 279(3):40–41, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-40.pdf.

Beardsley:1998:DH

[4573] Tim Beardsley. The dope on holography. Scientific American, 279(3):41, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-41.pdf.

Grossman:1998:OT

[4574] Wendy M. Grossman. Opaque transparency. Scientific American, 279(3):42, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
42.pdf.

Agnew:1998:PLFb

[4575] Neville Agnew and Martha Demas. Preserving the Laetoli footprints. Scientific American, 279(3):44–55, September 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n3/pdf/scientificamerican0998-44.pdf.

Barkley:1998:ADHa

[4576] Russell A. Barkley. Attention-deficit hyperactivity disorder. Scientific American, 279(3):44–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0908issue/0998barkley.html.

Armbruster:1998:EMMa

[4577] Paul Armbruster and Fritz Peter Hessberger. Elementary matters: Making new elements. Scientific American, 279(3):50–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Scerri:1998:EPSa

[4578] Eric R. Scerri. The evolution of the periodic system. Scientific American, 279(3):56–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1998/

0998issue/0998currentissue.html.

White:1998:WHBb

[4579] Ronald J. White. Weightlessness and the human body. Scientific American, 279(3):58–63, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
58.pdf.

Weissman:1998:OCa

[4580] Paul R. Weissman. The Oort cloud. Scientific American, 279(3):62– ??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/0998issue/
0998currentissue.html.

Barkley:1998:ADHb

[4581] Russell A. Barkley. Attention-deficit hyperactivity disorder. Scientific American, 279(3):66–71, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
66.pdf.

Coutts:1998:Ta

[4582] Timothy J. Coutts and Mark C. Fitzgerald. Thermophotovoltaics. Scientific American, 279(3):68–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0998issue/0998currentissue.html.

Armbruster:1998:EMMb

[4583] Peter Armbruster and Fritz Peter Hessberger. Elementary matters: Making new elements. Scientific American, 279(3):72–77, September 1998.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n3/pdf/scientificamerican0998-72.pdf.

Anonymous:1998:ASE

[4584] Anonymous. The amateur scientist: Extracting DNA in your own kitchen. Scientific American, 279(3):74–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0998issue/0998currentissue.html.

Anonymous:1998:MRBb

[4585] Anonymous. Mathematical recreations: Building the pyramids on schedule. Scientific American, 279(3):76–??, September 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/0998issue/0998currentissue.html.

Scerri:1998:EPSb

[4586] Eric R. Scerri. The evolution of the periodic system. Scientific American, 279(3):78–83, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
78.pdf.

Burke:1998:CIS

[4587] James Burke. Connections: Insurance, stamps and stolen marbles. Scientific American, 279(3):80–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0998issue/0998currentissue.html.

Morrison:1998:RCM

[4588] Philip Morrison and Phylis Morrison. Reviews and commentaries: A math book for everybody. wonders: Alien constellations. Scientific American, 279(3):80–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0998issue/0998currentissue.html.

Weissman:1998:OCb

[4589] Paul R. Weissman. The Oort Cloud. Scientific American, 279(3):84–89, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-84.pdf.

Anonymous:1998:WKHc

[4590] Anonymous. Working knowledge: How CD players turn light into sound. Scientific American, 279(3):86–??, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/0998issue/0998currentissue.html.

Coutts:1998:Tb

[4591] Timothy J. Coutts and Mark C. Fitzgerald. Thermophotovoltaics. Scientific American, 279(3):90–95, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
90.pdf.

Carlson:1998:ASSb

[4592] Shawn Carlson. The amateur scientist: Spooling the stuff of life. Scientific American, 279(3):96–97, September 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n3/pdf/scientificamerican0998-96.pdf.

Stewart:1998:MRCb

[4593] Ian Stewart. Mathematical recreations: Counting the pyramid builders. Scientific American, 279(3):98–100, September 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n3/pdf/scientificamerican0998-98.pdf.

Albers:1998:RCM

[4594] Donald J. Albers. Reviews and commentaries: Mathematics for the many. Scientific American, 279(3):102–103, September 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n3/pdf/scientificamerican0998-102.pdf.

Morrison:1998:Wb

[4595] Philip Morrison and Phylis Morrison. Wonders. Scientific American, 279(3):106, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n3/pdf/scientificamerican0998-
106.pdf.

Burke:1998:Cb

[4596] James Burke. Connections. Scientific American, 279(3):107–108, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-107.pdf.

Pohlmann:1998:WK

[4597] Ken C. Pohlmann. Working knowledge. Scientific American, 279(3):109, September 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n3/pdf/scientificamerican0998-109.pdf.

Anonymous:1998:CSIa

[4598] Anonymous. Celebrating the sea — an introduction — the oceans revealed. Scientific American, 279(3(special)):7–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:AO

[4599] Anonymous. Atlantic Ocean. Scientific American, 279(3(special)):8– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:POa

[4600] Anonymous. Pacific Ocean. Scientific American, 279(3(special)):10–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:IO

[4601] Anonymous. Indian Ocean. Scientific American, 279(3(special)):12–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:POb

[4602] Anonymous. Polar oceans. Scientific American, 279(3(special)):14–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kasting:1998:OWE

[4603] James F. Kasting. The origins of water on Earth — nearly three quarters of our planet is covered by oceans because Earth retained the water that rained down from space. Scientific American, 279(3(special)):16– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1998:BNG

[4604] Glenn Zorpette. Bikini’s nuclear ghosts — nuclear testing ravaged Bikini Atoll during the 1940s and 1950. today it’s a wreck-diver’s dream. Scientific American, 279(3(special)):24–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Schneider:1998:RS

[4605] David Schneider. The rising seas. Scientific American, 279(3(special)): 28–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Coakley:1998:FDB

[4606] Bernard J. Coakley. Forty days in the Bell of the Beast — conducting research on an attack submarine under the Arctic ice pack isn’t just a job — it’s an adventure. Scientific American, 279(3(special)):36–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Webster:1998:OW

[4607] Peter J. Webster and Judith A. Curry. The oceans and weather. Scientific American, 279(3(special)):38–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Edmunds:1998:TDU

[4608] Peter J. Edmunds. Ten days under the sea. Scientific American, 279 (3(special)):44–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nixon:1998:ESD

[4609] Scott W. Nixon. Enriching the sea to death. Scientific American, 279 (3(special)):48–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Marshall:1998:WRF

[4610] Justin Marshall. Why are reef fish so colorful? — fish residing around corals are living rainbows. but the biological utility of those hues is complex. Scientific American, 279(3(special)):54–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Safina:1998:WIF

[4611] Carl Safina. The world’s imperiled fish — the closure of prime fishing grounds and the declining yield of capture fisheries around the world demonstrate that people have sorely overtaxed a precious living resource: marine fish. Scientific American, 279(3(special)):58–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:PPAb

[4612] Anonymous. The promise and perils of aquaculture. Scientific American, 279(3(special)):64–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nonis:1998:FZS

[4613] Anton Nonis. Fishing the “zone” in Sri Lanka — in theory, coastal nations control fishing out to 200 nautical miles. but the reality can be quite different. Scientific American, 279(3(special)):70–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anderson:1998:SMB

[4614] R. Charles Anderson. Sharks mean business. Scientific American, 279 (3(special)):72–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nybakken:1998:LO

[4615] James W. Nybakken and Steven K. Webster. Life in the ocean. Scientific American, 279(3(special)):74–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dellinger:1998:AWT

[4616] Thomas Dellinger. The Atlantic’s wandering turtles — keeping track of these wide-ranging creatures is challenging — but not impossible — with satellite technology. Scientific American, 279(3(special)):88–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Binns:1998:MWB

[4617] Raymond A. Binns and David L. Dekker. The mineral wealth of the Bismarck Sea — as entrepreneurs consider mining valuable metals from the floor of the ocean near Papua New Guinea, scientists weigh both the economic potential and the threat to deep-sea life. Scientific American, 279(3(special)):92–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Malahoff:1998:IB

[4618] Alexander Malahoff. An island is born. Scientific American, 279 (3(special)):98–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jacobson:1998:EOL

[4619] Jon L. Jacobson and Alison Rieser. The evolution of ocean law. Scientific American, 279(3(special)):100–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Menduno:1998:EOP

[4620] Michael Menduno. Exploring the ocean planet. Scientific American, 279 (3(special)):106–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1998:ELH

[4621] John Rennie. From the Editors: Learning from the hackers. Scientific American, 279(4):6, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
6.pdf.

Anonymous:1998:LEj

[4622] Anonymous. Letters to the Editors. Scientific American, 279(4):8–9, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-8.pdf.

Anonymous:1998:YAIj

[4623] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 279(4):12, October 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n4/pdf/scientificamerican1098-12.pdf.

Stix:1998:FPD

[4624] Gary Stix. In focus: Personalizing drugs through genetics. Scientific American, 279(4):17–18 (Intl. ed. 10–??), October 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/1098issue/1098currentissue.html.

Stix:1998:NAP

[4625] Gary Stix. News and analysis: Personal pills. Scientific American, 279(4):17–18, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
17.pdf.

Beardsley:1998:HN

[4626] Tim Beardsley. In the heat of the night. Scientific American, 279(4):20, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-20.pdf.

Beardsley:1998:SCT

[4627] Tim Beardsley. Science and the citizen: Testing quantum gravity. globalwarming trees. did life’s molecules get a start in space? Scientific American, 279(4):20–20 (Intl. ed. 12–??), October 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/1098issue/1098currentissue.html.

Schneider:1998:PL

[4628] David Schneider. Polarized life. Scientific American, 279(4):24, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-24a.pdf.

Musser:1998:SI

[4629] George Musser. String instruments. Scientific American, 279(4):24–28, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-24b.pdf.

Leutwyler:1998:Bg

[4630] Kristin Leutwyler. In brief. Scientific American, 279(4):28–32, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-28.pdf.

Beardsley:1998:AD

[4631] Tim Beardsley. After the deluge. Scientific American, 279(4):30, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-30.pdf.

Doyle:1998:WDA

[4632] Rodger Doyle. Where the doctors aren’t. Scientific American, 279(4):31, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-31.pdf.

Mirsky:1998:WWW

[4633] Steve Mirsky. Whale weight watching. Scientific American, 279(4):32, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-32.pdf.

Ezzel:1998:PJF

[4634] Carol Ezzel. Profile: Judah Folkman kills tumors by choking off their blood supply. Scientific American, 279(4):33–34 (Intl. ed. 21– ??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1098issue/
1098currentissue.html.

Ezzell:1998:STT

[4635] Carol Ezzell. Starving tumors of their lifeblood. Scientific American, 279(4):33–34, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
33.pdf.

Nelson:1998:BT

[4636] Roxanne Nelson. Bloodless testing. Scientific American, 279(4):38–39, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-38.pdf.

Nelson:1998:TBW

[4637] Roxanne Nelson, Daniel G. Dupont, David Schneider, and Gary Stix. Technology and business: Will insurers snub new diabetes monitors? acoustic spotlight. new strategies for the urban jungle. better prosthetic hand. Scientific American, 279(4):38–41, 44 (Intl. ed. 23– ??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1098issue/
1098currentissue.html.

Dupont:1998:ICV

[4638] Daniel G. Dupont. Inner-city violence. Scientific American, 279(4):39– 40, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-39.pdf.

Schneider:1998:AS

[4639] David Schneider. In the audio spotlight. Scientific American, 279(4):40– 41, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-40.pdf.

Stix:1998:PT

[4640] Gary Stix. Phantom touch. Scientific American, 279(4):41–44, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-41.pdf.

Grossman:1998:CVA

[4641] Wendy Grossman. Cyber view: Armageddon and the Year 2000 bug. Scientific American, 279(4):48–48 (Intl. ed. 27–??), October 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/1098issue/1098currentissue.html.

Grossman:1998:YEW

[4642] Wendy M. Grossman. Y2K: The end of the world as we know it. Scientific American, 279(4):48, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
48.pdf.

Kraan-Korteweg:1998:GBM

[4643] Renée C. Kraan-Korteweg and Ofer Lahav. Galaxies behind the Milky Way. Scientific American, 279(4):50–57 (Intl. ed. 28–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-50.pdf;   http:/
/www.sciam.com/1998/1008issue/1098laham.html.

Jordan:1998:DE

[4644] V. Craig Jordan. Designer estrogens. Scientific American, 279(4):60– 67 (Intl. ed. 36–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
60.pdf;   http://www.sciam.com/1998/1098issue/1098currentissue.
html.

Martini:1998:SSH

[4645] Frederic H. Martini. Secrets of the slime hag. Scientific American, 279(4): 70–75 (Intl. ed. 44–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
70.pdf;   http://www.sciam.com/1998/1098issue/1098currentissue.
html.

Quinn:1998:ABM

[4646] Helen R. Quinn and Michael S. Witherell. The asymmetry between matter and antimatter. Scientific American, 279(4):76–81 (Intl. ed. 50–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-76.pdf;   http:/
/www.sciam.com/1998/1098issue/1098currentissue.html. See er-

rata [4839].

Ben-Jacob:1998:AM

[4647] Eshel Ben-Jacob and Herbert Levine. The artistry of microorganisms. Scientific American, 279(4):82–87 (Intl. ed. 56–??), October 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/1008issue/1098levine.html.

Ben-Jacob:1998:SPA

[4648] Eshel Ben-Jacob and Herbert Levine. Science in pictures: The artistry of microorganisms. Scientific American, 279(4):82–87, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n4/pdf/scientificamerican1098-82.pdf.

Moyer:1998:SNA

[4649] Albert E. Moyer. Simon Newcomb: Astronomer with an attitude. Scientific American, 279(4):88–93 (Intl. ed. 62–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-88.pdf;   http:/
/www.sciam.com/1998/1098issue/1098currentissue.html.

Anonymous:1998:CSIb

[4650] Anonymous. Computer security and the Internet. Scientific American, 279(4):95–95 (Intl. ed. 69–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/1098issue/1098currentissue.html.

Editors:1998:SRC

[4651] The Editors. Special report: Computer security and the Internet. Scientific American, 279(4):95, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
95.pdf.

Meinel:1998:HHB

[4652] Carolyn P. Meinel. How hackers break in … and how they are caught. Scientific American, 279(4):98–105 (Intl. ed. 70–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-98.pdf;   http:/
/www.sciam.com/1998/1098issue/1098currentissue.html.

Cheswick:1998:CRF

[4653] William Cheswick. Case report: Firewalls. Scientific American, 279(4):106–107, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
106.pdf.

Cheswick:1998:HCS

[4654] William Cheswick and Steven M. Bellovin. How computer security works: Firewalls. Scientific American, 279(4):106–107 (Intl. ed. 78– ??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1098issue/
1098currentissue.html.

Ford:1998:DC

[4655] Warwick Ford. Digital certificates. Scientific American, 279(4):108, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-108.pdf.

Ford:1998:HCS

[4656] Warwick Ford. How computer security works: Digital certificates. Scientific American, 279(4):108–108 (Intl. ed. 78–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/1098issue/1098currentissue.html.

Gosling:1998:HCS

[4657] James Gosling. How computer security works: The Java sandbox. Scientific American, 279(4):109–109 (Intl. ed. 78–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/1098issue/1098currentissue.html.

Gosling:1998:JS

[4658] James Gosling. The Java sandbox. Scientific American, 279(4):109, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-109.pdf.

Zimmermann:1998:CI

[4659] Philip R. Zimmermann. Cryptography for the Internet. Scientific American, 279(4):110–115 (Intl. ed. 82–??), October 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n4/pdf/scientificamerican1098-110.pdf;   http://www.sciam.com/
1998/1098issue/1098currentissue.html.

Rivest:1998:CAR

[4660] Ronald L. Rivest. The case against regulating encryption technology. Scientific American, 279(4):116–117 (Intl. ed. 88–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-116.pdf;   http:
//www.sciam.com/1998/1098issue/1098currentissue.html.

Carlson:1998:ASH

[4661] Shawn Carlson. The amateur scientist: Home movies of an Invisible World. Scientific American, 279(4):118–121, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n4/pdf/scientificamerican1098-118.pdf;   http://www.sciam.com/
1998/1098issue/1098currentissue.html.

Stewart:1998:MRP

[4662] Ian Stewart. Mathematical recreations: Playing with chocolate. Scientific American, 279(4):122–124, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
122.pdf;   http://www.sciam.com/1998/1098issue/1098currentissue.
html.

Tavris:1998:RCP

[4663] Carol Tavris. Reviews and commentaries: The paradox of gender. Scientific American, 279(4):122, 124 (Intl. ed. 93–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/1098issue/1098currentissue.html.

Anonymous:1998:RCk

[4664] Anonymous. Reviews and commentaries. Scientific American, 279(4):126–128, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
126.pdf.

Anonymous:1998:ERa

[4665] Anonymous. The Editors recommend. Scientific American, 279 (4):128–130, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
128.pdf.

Morrison:1998:FGN

[4666] Philip Morrison and Phylis Morrison. Field guide to the new biology lab. Scientific American, 279(4):131–133, October 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n4/pdf/scientificamerican1098-131.pdf.

Morrison:1998:WFG

[4667] Philip Morrison and Phylis Morrison. Wonders: Field guide to the new biology lab. Scientific American, 279(4):131, 133 (Intl. ed. 95– ??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1098issue/
1098currentissue.html.

Burke:1998:CDT

[4668] James Burke. Connections: Does this take you back? Scientific American, 279(4):132–133 (Intl. ed. 95–??), October 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/1098issue/1098currentissue.html.

Burke:1998:DTY

[4669] James Burke. Does this take you back? Scientific American, 279(4):132–133, October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n4/pdf/scientificamerican1098-
132.pdf.

Boegner:1998:WKM

[4670] Ray F. Boegner. Working knowledge: Motion-picture projectors. Scientific American, 279(4):134–135 (Intl. ed. 102–??), October 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n4/pdf/scientificamerican1098-134.pdf;   http:
//www.sciam.com/1998/1098issue/1098currentissue.html.

Rennie:1998:EWW

[4671] John Rennie. From the Editor: Who and what we lost. Scientific American, 279(5):1, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
1.pdf.

Anonymous:1998:LEk

[4672] Anonymous. Letters to the Editors. Scientific American, 279(5):8–10, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-8.pdf.

Anonymous:1998:Ej

[4673] Anonymous. Errata. Scientific American, 279(5):10, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n5/pdf/scientificamerican1198-10.pdf.

Anonymous:1998:YAj

[4674] Anonymous. 50, 100 and 150 years ago. Scientific American, 279(5): 14–14 (Intl. ed. 8–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/1198issue/1198currentissue.html.

Anonymous:1998:YAIk

[4675] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 279(5):14, November 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n5/pdf/scientificamerican1198-14.pdf.

Gibbs:1998:FBR

[4676] W. Wayt Gibbs. In focus: A breakthrough report on regenerating human neurons. Scientific American, 279(5):19–20 (Intl. ed. 9– ??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1198issue/
1198currentissue.html.

Gibbs:1998:NAD

[4677] W. Wayt Gibbs. News and analysis: Dogma overturned. Scientific American, 279(5):19–20, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
19.pdf.

Musser:1998:IC

[4678] George Musser. Inconstant constants. Scientific American, 279(5):24–28, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-24.pdf.

Musser:1998:SCL

[4679] George Musser, Carol Ezzel, Gary Stix, Steve Mirsky, and Tim Beardsley. Science and the citizen: Lab bacteria become drug resistant. Inconstant nature? Alternatives to amniocentesis. Voting on the environment. Scientific American, 279(5):24, 28–29, 32, 34–37 (Intl. ed. 13–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1198issue/
1198currentissue.html.

Ezzell:1998:D

[4680] Carol Ezzell. Down detection. Scientific American, 279(5):28–29, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-28.pdf.

Stix:1998:CC

[4681] Gary Stix. The E. Coli are coming. Scientific American, 279(5):29, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-29.pdf.

Mirsky:1998:LL

[4682] Steve Mirsky. Lucky Laima. Scientific American, 279(5):32, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-32b.pdf.

Anonymous:1998:Bd

[4683] Anonymous. In brief. Scientific American, 279(5):32–36, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n5/pdf/scientificamerican1198-32a.pdf.

Beardsley:1998:TTC

[4684] Tim Beardsley. Tool time on cactus hill. Scientific American, 279(5):34–36, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
34.pdf.

Doyle:1998:HCV

[4685] Rodger Doyle. How Congress voted on the environment. Scientific American, 279(5):36–37, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
36.pdf.

Gibbs:1998:MMT

[4686] W. Wayt Gibbs. Monstrous moonshine is true. Scientific American, 279(5):40–41, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
40.pdf.

Gibbs:1998:PMR

[4687] W. Wayt Gibbs. Profile: Mathematician Richard Borcherds and the Monstrous Moonshine. Scientific American, 279(5):40–41 (Intl. ed. 21– ??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1198issue/
1198currentissue.html.

Zorpette:1998:PB

[4688] Glenn Zorpette. Patent blunder. Scientific American, 279(5):42, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-42.pdf.

Zorpette:1998:TBU

[4689] Glenn Zorpette, Steven J. Frank, Alden M. Hayashi, and Gary Stix. Technology and business: U.S. patents and terrorist weapons. microexplosions. first antisense drug approved. Bill Gates, you’re in the Navy now. Scientific American, 279(5):42, 46, 50–51 (Intl. ed. 25– ??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1198issue/
1198currentissue.html.

Hayashi:1998:RSS

[4690] Alden M. Hayashi. Rough sailing for smart ships. Scientific American, 279(5):46, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
46a.pdf.

Stix:1998:SG

[4691] Gary Stix. Shutting down a gene. Scientific American, 279(5):46–50, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-46b.pdf.

Stix:1998:LB

[4692] Gary Stix. Little bangs. Scientific American, 279(5):50–51, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-50.pdf.

Wallich:1998:CVH

[4693] Paul Wallich. Cyber view: This is not a hoax. Scientific American, 279 (5):54–54 (Intl. ed. 29–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/1198issue/1198currentissue.html.

Wallich:1998:H

[4694] Paul Wallich. This is not a hoax! Scientific American, 279(5):54, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-54.pdf.

MacDonald:1998:NOS

[4695] Ian R. MacDonald. Natural oil spills. Scientific American, 279(5):56– 61 (Intl. ed. 30–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
56.pdf;   http://www.sciam.com/1998/1198issue/1198currentissue.
html.

Wynn:1998:MHD

[4696] Jeffrey C. Wynn and Eugene M. Shoemaker. The meteorite hunters: The day the sands caught fire. Scientific American, 279(5):64–71 (Intl. ed. 36–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1108issue/

1198wynn.html. Discusses the Nineteenth Century meteorite impact in the Empty Quarter (Rub’ al-Khali) of Saudi Arabia.

Wynn:1998:MHP

[4697] Jeffrey C. Wynn and Eugene M. Shoemaker. The meteorite hunters: Part I: The day the sands caught fire. Scientific American, 279(5):64–71, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-64.pdf.

Gibbs:1998:MHP

[4698] W. Wayt Gibbs. The meteorite hunters: Part II: The search for Greenland’s mysterious meteor. Scientific American, 279(5):72–79, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-72.pdf.

Gibbs:1998:MHS

[4699] W. Wayt Gibbs. The meteorite hunters: The search for Greenland’s mysterious meteor. Scientific American, 279(5):72–79 (Intl. ed. 44–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/explorations/1998/
080398meteor/index.html.

Close:1998:G

[4700] Frank E. Close and Philip R. Page. Glueballs. Scientific American, 279 (5):80–85 (Intl. ed. 52–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
80.pdf;   http://www.sciam.com/1998/1198issue/1198currentissue.
html.

Nesse:1998:EOD

[4701] Randolph M. Nesse and George C. Williams. Evolution and the origins of disease. Scientific American, 279(5):86–93 (Intl. ed. 58–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-86.pdf;   http:/
/www.sciam.com/1998/1108issue/1198nesse.html.

Preston-Mafham:1998:MSS

[4702] Ken Preston-Mafham and Rod Preston-Mafham. Mating strategies of spiders. Scientific American, 279(5):94–99 (Intl. ed. 68– ??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1198issue/
1198currentissue.html.

Preston-Mafham:1998:SPM

[4703] Ken Preston-Mafham and Rod Preston-Mafham. Science in pictures: Mating strategies of spiders. Scientific American, 279(5):94–99, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-94.pdf.

Gerstein:1998:SWM

[4704] Mark Gerstein and Michael Levitt. Simulating water and the molecules of life. Scientific American, 279(5):100–105 (Intl. ed. 74–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/

journal/v279/n5/pdf/scientificamerican1198-100.pdf;   http:
//www.sciam.com/1998/1198issue/1198currentissue.html.

Livingston:1998:YMM

[4705] James D. Livingston. 100 years of magnetic memories. Scientific American, 279(5):106–111 (Intl. ed. 80–??), November 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n5/pdf/scientificamerican1198-106.pdf;   http://www.sciam.com/
1998/1198issue/1198currentissue.html.

Carlson:1998:ASFa

[4706] Shawn Carlson. The amateur scientist: Floating a challenge. Scientific American, 279(5):112–114, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
112.pdf.

Carlson:1998:ASFb

[4707] Shawn Carlson. The amateur scientist: Floating a challenge. Scientific American, 279(5):112, 114 (Intl. ed. 86–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1998/1198issue/1198currentissue.html.

Stewart:1998:MRRb

[4708] Ian Stewart. Mathematical recreations: Resurrection shuffle. Scientific American, 279(5):116–117 (Intl. ed. 88–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n5/pdf/scientificamerican1198-116.pdf;   http://www.sciam.com/
1998/1198issue/1198currentissue.html.

Anonymous:1998:RCl

[4709] Anonymous. Reviews and commentaries. Scientific American, 279 (5):118–120, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
118.pdf.

Livingston:1998:TBM

[4710] James D. Livingston. Three books on materials science explain where civilization would be without, well, stuff. Scientific American, 279(5): 118–120 (Intl. ed. 90–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/1198issue/1198currentissue.html.

Anonymous:1998:ERb

[4711] Anonymous. The Editors recommend. Scientific American, 279 (5):120–121, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
120.pdf.

Anonymous:1998:ERN

[4712] Anonymous. The editors recommend: New, noteworthy books on science. Scientific American, 279(5):120–121 (Intl. ed. 92–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/1998/1198issue/1198currentissue.html.

Burke:1998:CO

[4713] James Burke. Connections: Oops. Scientific American, 279(5):122– 123(Intl. ed. 94–??), November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/1198issue/1198currentissue.html.

Burke:1998:O

[4714] James Burke. Oops. Scientific American, 279(5):122–123, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n5/pdf/scientificamerican1198-122.pdf.

Morrison:1998:NPC

[4715] Philip Morrison. Numbers: Prime or choice? Scientific American, 279(5):123–124, November 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n5/pdf/scientificamerican1198-
124.pdf.

Morrison:1998:WNP

[4716] Philip Morrison. Wonders: Numbers: Prime or choice? Scientific American, 279(5):124, 123(Intl. ed. 96–??), November 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/1998/1198issue/1198currentissue.html.

Rennie:1998:ECO

[4717] John Rennie. From the Editor: Counting our blessings. Scientific American, 279(6):8, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
8.pdf.

Anonymous:1998:LEl

[4718] Anonymous. Letters to the Editors. Scientific American, 279(6):10–12, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-10.pdf.

Anonymous:1998:YAIl

[4719] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 279(6):14, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-14.pdf.

Anonymous:1998:YAk

[4720] Anonymous. 50, 100 and 150 years ago. Scientific American, 279 (6):14, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Stix:1998:NAN

[4721] Gary Stix. News and analysis: Nothing but light. Scientific American, 279(6):17–20, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
17.pdf.

Stix:1998:FNL

[4722] Gary Stix. In focus: Nothing but light: Hunger for bandwidth drives alloptical technology to market. Scientific American, 279(6):17, 20 (Intl. ed. 17–??), December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Zorpette:1998:AA

[4723] Glenn Zorpette. Andro angst. Scientific American, 279(6):22–26, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-22.pdf.

Zorpette:1998:SCM

[4724] Glenn Zorpette, George Musser, Carol Ezzell, Rodger Doyle, Madhusree Mukerjee, Kristin Leutwyler, and Steve Mirsky. Science and the citizen: A mysterious new force pulling space probes? Probably not… “Fat hormone” linked to other ills … Drugs for jocks… Unexpected declines in the birth rate. Scientific American, 279(6):22, 26, 28, 30–35 (Intl. ed. 22–??), December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Musser:1998:PGL

[4725] George Musser. Pioneering gas leak? Scientific American, 279(6):26–28, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-26b.pdf.

Anonymous:1998:Be

[4726] Anonymous. In brief. Scientific American, 279(6):26–34, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-26a.pdf.

Ezzell:1998:LL

[4727] Carol Ezzell. Leaping leptin. Scientific American, 279(6):28–30, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-28.pdf.

Doyle:1998:HRT

[4728] Rodger Doyle. Human rights throughout the world. Scientific American, 279(6):30–31, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
30.pdf.

Mukerjee:1998:PS

[4729] Madhusree Mukerjee. The population slide. Scientific American, 279(6):32–33, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
32.pdf.

Mirsky:1998:LS

[4730] Steve Mirsky. A leg to stand on. Scientific American, 279(6):34, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-34.pdf.

Beardsley:1998:STS

[4731] Tim Beardsley. Smashing through science’s glass ceiling. Scientific American, 279(6):36–38, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
36.pdf.

Beardsley:1998:PRR

[4732] Tim Beardsley. Profile: Rita R. Colwell, the new head of the National Science Foundation. Scientific American, 279(6):36, 38 (Intl. ed. 36–??), December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Stix:1998:ST

[4733] Gary Stix. Shading the twinkle. Scientific American, 279(6):40, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-40a.pdf.

Schneider:1998:WNB

[4734] David Schneider. Where no brush can reach. Scientific American, 279(6):40–44, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
40b.pdf.

Stix:1998:TBA

[4735] Gary Stix, David Schneider, Wendy M. Grossman, and W. Wayt Gibbs. Technology and business: Antirust bacteria… Chaotic laser computer… Jamming the monitor spies. Scientific American, 279(6):40, 44, 47(Intl. ed. 40–??), December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Grossman:1998:BT

[4736] Wendy M. Grossman. Beating the tempest. Scientific American, 279(6):44–45, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
44.pdf.

Gibbs:1998:CC

[4737] W. Wayt Gibbs. Computing with chaos. Scientific American, 279(6):45–47, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
45.pdf.

Gibbs:1998:CVH

[4738] W. Wayt Gibbs. Cyber view: Hello, is this the Web? Scientific American, 279(6):48, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
48.pdf;   http://www.sciam.com/1998/1298issue/1298currentissue.
html.

Henry:1998:EGC

[4739] J. Patrick Henry, Ulrich G. Briel, and Hans Böhringer. The evolution of galaxy clusters. Scientific American, 279(6):52–57, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-52.pdf;   http:/
/www.sciam.com/1998/1298issue/1298currentissue.html.

Wilmut:1998:CM

[4740] Ian Wilmut. Cloning for medicine. Scientific American, 279(6):58–63, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-58.pdf;   http:/
/www.sciam.com/1998/1298issue/1298wilmut.html.

Altarelli:1998:MUX

[4741] Massimo Altarelli, Fred Schlachter, and Jane Cross. Making ultrabright X-rays: Radiation a billion times brighter than the sun’s is illuminating a host of scientific and technical phenomena. Scientific American, 279(6):66–73, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-66.pdf;   http://www.sciam.com/
1998/1298issue/1298currentissue.html.

Garnick:1998:CPC

[4742] Marc B. Garnick and William R. Fair. Combating prostate cancer. Scientific American, 279(6):74–83, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-74.pdf;   http://www.sciam.com/
1998/1298issue/1298currentissue.html.

Groves:1998:SPL

[4743] Paul Groves and photographs by Paul Sutherland. Science in pictures: Leafy sea dragons. Scientific American, 279(6):84–89, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-84.pdf;   http:/
/www.sciam.com/1998/1298issue/1298currentissue.html.  Pho-

tographs by Paul Sutherland.

OBrochta:1998:BBB

[4744] David A. O’Brochta and Peter W. Atkinson. Building the better bug. Scientific American, 279(6):90–95, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-90.pdf;   http://www.sciam.com/
1998/1298issue/1298currentissue.html.

Brown:1998:PWJ

[4745] Laurie M. Brown and Yoichiro Nambu. Physicists in wartime Japan. Scientific American, 279(6):96–103, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-96.pdf;   http://www.sciam.com/
1998/1298issue/1298currentissue.html.

Jones:1998:SS

[4746] Capers Jones. Sizing up software. Scientific American, 279(6):104–109, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-104.pdf;   http:
//www.sciam.com/1998/1298issue/1298jones.html.

Carlson:1998:ASSc

[4747] Shawn Carlson. The amateur scientist: Sorting molecules with electricity. Scientific American, 279(6):110–111, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-110.pdf;   http://www.sciam.com/
1998/1298issue/1298currentissue.html.

Stewart:1998:MRY

[4748] Ian Stewart. Mathematical recreations: Your half’s bigger than my half! Scientific American, 279(6):112–114, December 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v279/
n6/pdf/scientificamerican1298-112.pdf;   http://www.sciam.com/
1998/1298issue/1298currentissue.html.

Anonymous:1998:RCm

[4749] Anonymous. Reviews and commentaries. Scientific American, 279 (6):116–121, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v279/n6/pdf/scientificamerican1298-
116.pdf.

Morrison:1998:SAY

[4750] Philip Morrison and Phylis Morrison. The Scientific American Young Readers Book Awards. Scientific American, 279(6):116–121, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Burke:1998:CTA

[4751] James Burke. Connections: Tea, anyone? Scientific American, 279(6): 122–123, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Burke:1998:TA

[4752] James Burke. Tea, anyone? Scientific American, 279(6):122–123, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-122.pdf.

Anonymous:1998:AI

[4753] Anonymous. Annual index 1998. Scientific American, 279(6):124–126, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-124.pdf;   http:
//www.sciam.com/1998/1298issue/1298currentissue.html.

Poynter:1998:WK

[4754] Dan Poynter. Working knowledge. Scientific American, 279(6):127, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v279/n6/pdf/scientificamerican1298-127.pdf.

Poynter:1998:WKP

[4755] Dan Poynter. Working knowledge: Parachutes. Scientific American, 279 (6):127, December 1998. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1998/1298issue/
1298currentissue.html.

Marcy:1998:GPO

[4756] Geoffrey W. Marcy and R. Paul Butler. Giant planets orbiting farway stars: The first-detected planets around other suns are already overthrowing traditional theories about how solar systems form. Scientific American, ??(1):10–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jakosky:1998:SLO

[4757] Bruce M. Jakosky. Searching for life in our solar system: The more that is learned about our neighboring planets and moons, the more hospitable some of them look as havens for life, today or in the distant past. Scientific American, ??(1):16–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Angel:1998:SLO

[4758] Roger Angel and Neville J. Woolf. Searching for life in other solar systems: Worlds supporting life have characteristics that new generations of telescopes and other instruments should …. Scientific American, ?? (1):22–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:PTP

[4759] Anonymous. Planetary tour: A pictorial guide to the diverse, myriad worlds of our solar system — from gas giants to wandering pebbles — and their many peculiarities. Scientific American, ??(1):26–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:Ma

[4760] Anonymous. Mercury. Scientific American, ??(1):28–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:V

[4761] Anonymous. Venus. Scientific American, ??(1):30–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:Ek

[4762] Anonymous. Earth. Scientific American, ??(1):32–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:Mb

[4763] Anonymous. Mars. Scientific American, ??(1):34–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:S

[4764] Anonymous. Saturn. Scientific American, ??(1):38–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:U

[4765] Anonymous. Uranus. Scientific American, ??(1):40–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:N

[4766] Anonymous. Neptune. Scientific American, ??(1):42–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:Pe

[4767] Anonymous. Pluto. Scientific American, ??(1):44–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:CA

[4768] Anonymous. Comets and asteroids. Scientific American, ??(1):46–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Starrfield:1998:VCM

[4769] Sumner Starrfield and Steven N. Shore. V1974 Cygni 1992: The most important nova of the century. Scientific American, ??(1):56–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cronin:1998:CRE

[4770] James W. Cronin, Thomas K. Gaisser, and Simon P. Swordy. Cosmic rays at the energy frontier: Atomic particles packing the wallop of a pitcher’s fastball strike Earth’s atmosphere every day. Scientific American, ??(1): 62–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Veilleux:1998:CGE

[4771] Sylvain Veilleux, Gerald Cecil, and Jonathan Bland-Hawthorn. Colossal galactic explosions: At the heart of many galaxies rages a violent inferno, powered either by an ultramassive black hole or a burst of stellar birth. Scientific American, ??(1):74–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Peebles:1998:EUC

[4772] P. James E. Peebles, David N. Schramm, Edwin L. Turner, and Richard G. Kron. The evolution of the universe: Cosmologists have pieced together much about how the universe as we know it grew from a fireball instants after the big bang. yet unanswered questions remain. Scientific American, ??(1):86–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Freedman:1998:ERS

[4773] Wendy L. Freedman. The expansion rate and size of the universe: How fast the universe is expanding and what its diameter might be fundamentally limit cosmological theories. new observations yield better estimates of both. Scientific American, ??(1):92–??, ???? 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Linde:1998:SRI

[4774] Andrei Linde. The self-reproducing inflationary universe: Our universe may be just one infinitesimal part of a “multiverse” in which branching bubbles of spacetime contain different physical realities. Scientific American, ??(1):98–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rubin:1998:DMU

[4775] Vera Rubin. Dark matter in the universe: Somewhere in space hide masses of “dark matter” that hold together galaxies and galactic clusters. its nature and quantity rule the rate of the universe. Scientific American, ??(1):106–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1998:SAG

[4776] Tim Beardsley. A Scientific Armada: A guide to upcoming space missions. Scientific American, ??(1):112–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:IWH

[4777] Anonymous. The importance of women’s health: An introduction from the editors. plus: An interview with three experts in the field of women’s health: Phyllis Greenberger, M. S. W., Wanda K. Jones, Dr.P. H., and Vivian W. Pinn, M.D. Scientific American, ??(2):6–??, ???? 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:TYF

[4778] Anonymous. Tens and 20s: Your first pelvic exam, smoking, emergency contraception. Scientific American, ??(2):14–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leutwyler:1998:DTE

[4779] Kristin Leutwyler. Dying to be thin: Eating disorders cripple — literally — millions of young women. Scientific American, ??(2):16–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sheftell:1998:MH

[4780] Fred D. Sheftell. Migraine headaches. Scientific American, ??(2):20– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Grady:1998:HVR

[4781] Denise Grady. Help for victims of rape: Confronting painful memories can help victims cope with the trauma. Scientific American, ??(2):21– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1998:IAY

[4782] Marguerite Holloway. The importance of addressing young men’s health. Scientific American, ??(2):22–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Koutsky:1998:WWN

[4783] Laura A. Koutsky. What women need to know about sexually transmitted diseases: Left undiagnosed, STDs can be deadly. Scientific American, ??(2):22–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McKinsey:1998:AYA

[4784] Krista McKinsey. Arm yourself against STDs. Scientific American, ?? (2):22–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hede:1998:FES

[4785] Karyn Hede. Focus on education: Single-sex classrooms; girls, math and science. Scientific American, ??(2):28–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leibenluft:1998:WMW

[4786] Ellen Leibenluft. Why are so many women depressed?: Women may be more sensitive to some changes in the environment. Scientific American, ??(2):30–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Strauss:1998:FOW

[4787] Evelyn Strauss. The female orgasm: Why are so many women missing out? Scientific American, ??(2):38–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:PTM

[4788] Anonymous. 30s and 40s: Prenatal testing, mammograms, maintaining strong bones. Scientific American, ??(2):42–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Faustman:1998:WBA

[4789] Denise Faustman. When the body attacks itself: Autoimmune diseases afflict women much more frequently. Scientific American, ??(2):44–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hopkin:1998:ADC

[4790] Karen Hopkin. Are autoimmune disorders colorblind? Scientific American, ??(2):44–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rosenwaks:1998:I

[4791] Zev Rosenwaks and Mark V. Sauer. Infertility. Scientific American, ??(2):51–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shaffer:1998:EMC

[4792] Marjorie Shaffer. Endometriosis: A major cause of infertility in women. Scientific American, ??(2):51–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1998:EAR

[4793] Tim Beardsley. The ethics of assisted reproduction: Medicine can do a lot to help people become parents — sometimes, maybe too much. Scientific American, ??(2):55–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Arthur:1998:GMH

[4794] Stephanie J. Arthur. Get moving: How much exercise is enough? Scientific American, ??(2):59–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Brown:1998:FPP

[4795] Kathryn Sergeant Brown and Denise Grady. Focus on pregnancy: Preeclampsia; birth timing; lessening pain during labor. Scientific American, ??(2):60–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

King:1998:GBO

[4796] Mary-Claire King. The genetics of breast and ovarian cancer. Scientific American, ??(2):64–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mellman:1998:CVA

[4797] Lisa A. Mellman. The consequences of violence against women: Violence begets violence. Scientific American, ??(2):68–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wilsnack:1998:WA

[4798] Sharon Wilsnack. Women and alcohol. Scientific American, ??(2):72– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Silver:1998:BDO

[4799] Lisa Silver. Bad day at the office?: Job stress affects women and men differently. Scientific American, ??(2):73–??, ???? 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:FHD

[4800] Anonymous. 50s and 60s: Folate and heart disease, screening for cancer. Scientific American, ??(2):76–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wise:1998:MBC

[4801] Phyllis M. Wise. Menopause and the brain: Chemical changes in the brain may signal the onset of menopause. Scientific American, ??(2):78– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lobo:1998:HRT

[4802] Rogerio A. Lobo and Graham A. Colditz. Hormone replacement therapy. Scientific American, ??(2):82–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:SBC

[4803] Anonymous. Smoking and breast cancer: Lung cancer: Why women’s risks are higher. Scientific American, ??(2):86–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shields:1998:SBC

[4804] Peter G. Shields and Christine B. Ambrosone. Smoking and breast cancer: Cigarettes may cause more cases than all the breast cancer genes combined. Scientific American, ??(2):86–??, ???? 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fackelmann:1998:WSF

[4805] Kathleen Fackelmann. What’s in store for the future. Scientific American, ??(2):90–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hill:1998:HDS

[4806] Martha N. Hill. Heart disease and stroke. Scientific American, ??(2):90– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ezzell:1998:FCI

[4807] Carol Ezzell. Fat chances: Is it okay to be plump? Scientific American, ??(2):94–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:VTO

[4808] Anonymous. 70s and up: Vaccines, testing for osteoporosis, choosing a pharmacy. Scientific American, ??(2):98–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Perls:1998:WWL

[4809] Thomas T. Perls and Ruth C. Fretts. Why women live longer than men: What gives women the extra years? Scientific American, ??(2):100– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lindsay:1998:O

[4810] Robert Lindsay and Donald P. McDonnell. Osteoporosis. Scientific American, ??(2):104–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Khachaturian:1998:MRA

[4811] Zaven S. Khachaturian. At more risk for Alzheimer’s?: Looking at how genes and gender interact in Alzheimer’s disease. Scientific American, ??(2):110–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Appell:1998:UI

[4812] Rodney A. Appell. Urinary incontinence. Scientific American, ??(2):114– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Maranto:1998:HBO

[4813] Gina Maranto. Having a ball: Older women share tips on enjoying a long and healthy life. Scientific American, ??(2):116–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:YHE

[4814] Anonymous. To your health: the editors. Scientific American, ??(2):120– ??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Yam:1998:IC

[4815] Philip Yam. Intelligence considered. Scientific American, Special Edition (4):6–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sternberg:1998:HII

[4816] Robert J. Sternberg. How intelligent is intelligence testing? Scientific American, Special Edition(4):12–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gardner:1998:MI

[4817] Howard Gardner. A multiplicity of intelligences. Scientific American, Special Edition(4):18–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gottfredson:1998:GIF

[4818] Linda S. Gottfredson. The general intelligence factor. Scientific American, Special Edition(4):24–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1998:WDB

[4819] Tim Beardsley. For whom did the bell curve toll? Scientific American, Special Edition(4):30–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Winner:1998:UTG

[4820] Ellen Winner. Uncommon talents: Gifted children, prodigies and savants. Scientific American, Special Edition(4):32–??, ???? 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1998:SSD

[4821] Marguerite Holloway. Seeking “smart” drugs. Scientific American, Special Edition(4):38–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Calvin:1998:EI

[4822] William H. Calvin. The emergence of intelligence. Scientific American, Special Edition(4):44–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gould:1998:RA

[4823] James L. Gould and Carol Grant Gould. Reasoning in animals. Scientific American, Special Edition(4):52–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pepperberg:1998:TAL

[4824] Irene M. Pepperberg. Talking with Alex: Logic and speech in parrots. Scientific American, Special Edition(4):60–??, ???? 1998. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gallup:1998:CAE

[4825] Gordon Gallup, Jr. Can animals empathize? Scientific American, Special Edition(4):66–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Povinelli:1998:CAE

[4826] Daniel J. Povinelli. Can animals empathize? maybe not. Scientific American, Special Edition(4):66–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ford:1998:CWR

[4827] Kenneth M. Ford and Patrick J. Hayes. On computational wings: Rethinking the goals of artificial intelligence. Scientific American, Special Edition(4):78–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ginsberg:1998:CGR

[4828] Matthew L. Ginsberg. Computers, games and the real world. Scientific American, Special Edition(4):84–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pentland:1998:WI

[4829] Alex P. Pentland. Wearable intelligence. Scientific American, Special Edition(4):90–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lemarchand:1998:TIL

[4830] Guillermo A. Lemarchand. Is there intelligent life out there? Scientific American, Special Edition(4):96–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1998:TMS

[4831] Anonymous. Table of major SETI projects. Scientific American, Special Edition(4):100–??, ???? 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Musser:1999:EGC

[4832] George Musser. From the Editors: Getting complicated. Scientific American, 280(1):6, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
6.pdf.

Anonymous:1999:Ea

[4833] Anonymous. From the Editor. Scientific American, 280(1):6–??, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:LEa

[4834] Anonymous. Letters to the editors. Scientific American, 280(1):10–11, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:LEb

[4835] Anonymous. Letters to the Editors. Scientific American, 280(1):10–11, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-10.pdf.

Anonymous:1999:Eb

[4836] Anonymous. Errata. Scientific American, 280(1):11, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-11.pdf.

Anonymous:1999:YA

[4837] Anonymous. 50, 100 and 150 years ago. Scientific American, 280(1):14, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:YAIa

[4838] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 280(1):14, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-14.pdf.

Quinn:1999:EAB

[4839] Helen R. Quinn and Michael S. Witherell. Errata: “The Asymmetry between Matter and Antimatter”. Scientific American, 280(1):14–14 (Intl. ed. 11–11), January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). See [4646].

Anonymous:1999:SBN

[4840] Anonymous. Special briefing: The 1998 Nobel Prizes in science. Scientific American, 280(1):16–19, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-16.pdf.

Gibbs:1999:NPS

[4841] W. Wayt Gibbs, Sasha Nemecek, and Gary Stix. The 1998 Nobel Prizes in Science. Scientific American, 280(1):16–19, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hayashi:1999:FCH

[4842] Alden M. Hayashi. In focus: Computer “hacktivists” fight for human rights with the Internet. Scientific American, 280(1):21–22, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hayashi:1999:NAN

[4843] Alden M. Hayashi. News and analysis: The net effect. Scientific American, 280(1):21–22, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
21.pdf.

Mukerjee:1999:AA

[4844] Madhusree Mukerjee. Out of Africa, into Asia. Scientific American, 280(1):24, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
24.pdf.

Mukerjee:1999:SCM

[4845] Madhusree Mukerjee, Kate Wong, Philip Yam, Mark Alpert, Steve Mirsky, George Musser, and Rodger Doyle. Science and the citizen: More genetic support for an African Eve. Unexpected cosmic rays. Whence whales? Space geriatrics. privacy in the workplace. Scientific American, 280(1):24, 26, 30, 32, 34, 36, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Wong:1999:CC

[4846] Kate Wong. Cetacean creation. Scientific American, 280(1):26–30, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-26b.pdf.

Yam:1999:Ba

[4847] Philip Yam. In brief. Scientific American, 280(1):26–32, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-26a.pdf.

Alpert:1999:JGE

[4848] Mark Alpert. John Glenn’s excellent adventure. Scientific American, 280(1):30–32, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
30.pdf.

Musser:1999:CP

[4849] George Musser. Cosmic power. Scientific American, 280(1):32–36, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-32.pdf.

Mirsky:1999:TM

[4850] Steve Mirsky. Taste matters. Scientific American, 280(1):34, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-34.pdf.

Doyle:1999:PW

[4851] Rodger Doyle. Privacy in the workplace. Scientific American, 280(1):36, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-36.pdf.

Anonymous:1999:PJR

[4852] Anonymous. Profile: James R. Flynn ponders the strange rise in IQ scores. Scientific American, 280(1):37–38, January 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1999:FE

[4853] Marguerite Holloway. Flynn’s Effect. Scientific American, 280(1):37–38, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-37.pdf.

Beardsley:1999:HLY

[4854] Tim Beardsley. Here’s looking at you. Scientific American, 280(1):39–40, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-39.pdf.

Beardsley:1999:TBC

[4855] Tim Beardsley, Sasha Nemecek, Philip Yam, and Peta Firth. Technology and business: Conscious cog. A racial and economic health gap. Food frenzies. Scientific American, 280(1):39–43, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nemecek:1999:UH

[4856] Sasha Nemecek. Unequal health. Scientific American, 280(1):40–41, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-40.pdf.

Firth:1999:CF

[4857] Peta Firth. Consuming fears. Scientific American, 280(1):41–42, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-41.pdf.

Yam:1999:BEI

[4858] Philip Yam. Brace for (educational) impact. Scientific American, 280(1): 42, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-42.pdf.

Dupont:1999:CVU

[4859] Daniel G. Dupont. Cyber view: The U.S. military defends itself against the World Wide Web. Scientific American, 280(1):44, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dupont:1999:S

[4860] Daniel G. Dupont. Out of site. Scientific American, 280(1):44, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-44.pdf.

Editors:1999:SRR

[4861] The Editors. Special report: Revolution in cosmology. Scientific American, 280(1):45, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
45.pdf.

Hogan:1999:SST

[4862] Craig J. Hogan, Robert P. Kirshner, and Nicholas B. Suntzeff. Surveying space-time with supernovae. Scientific American, 280(1):46– 51 (Intl. ed. 45–??), January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
46.pdf.

Krauss:1999:CA

[4863] Lawrence M. Krauss. Cosmological antigravity. Scientific American, 280(1):52–59, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
52.pdf.

Bucher:1999:ILD

[4864] Martin A. Bucher and David N. Spergel. Inflation in a low-density Universe. Scientific American, 280(1):62–69, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-62.pdf.

Tallamy:1999:CCA

[4865] Douglas W. Tallamy. Child care among the insects. Scientific American, 280(1):72–77, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
72.pdf.

Laver:1999:DFV

[4866] W. Graeme Laver, Norbert Bischofberger, and Robert G. Webster. Disarming flu viruses. Scientific American, 280(1):78–87, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-78.pdf.

deJager:1999:YMB

[4867] Peter de Jager. Y2K: So many bugs … so little time. Scientific American, 280(1):88–93, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
88.pdf;   http://www.sciam.com/1999/0199issue/0199dejager.html.

Moxon:1999:DMA

[4868] E. Richard Moxon and Christopher Wills. DNA microsatellites: Agents of evolution? Scientific American, 280(1):94–99, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-94.pdf.

Zorpette:1999:ESS

[4869] Glenn Zorpette. Expeditions: To save a salmon. Scientific American, 280 (1):100–105, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1999:SS

[4870] Glenn Zorpette. To save a salmon. Scientific American, 280(1):100–105, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-100.pdf.

Moxon:1999:ESS

[4871] E. Richard Moxon and Christopher Wills. Expeditions: To save a salmon. Scientific American, 280(1):100–??, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carlson:1999:ASTa

[4872] Shawn Carlson. The amateur scientist: Taking the Earth’s magnetic pulse. Scientific American, 280(1):106–108, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-106.pdf.

Carlson:1999:ASMa

[4873] Shawn Carlson. The amateur scientist: Monitoring the Earth’s magnetism. Scientific American, 280(1):106, 108 (Intl. ed. 106–??), January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1999:MRDa

[4874] Ian Stewart. Mathematical recreations: Division without envy. Scientific American, 280(1):110–111, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
110.pdf.

Panosian:1999:RCB

[4875] Claire Panosian. Reviews and commentaries: Blood money: a sanguine history of transfusions, contamination and commerce. Scientific American, 280(1):112–113, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Panosian:1999:RCL

[4876] Claire Panosian. Reviews and commentaries: Life and death is in the blood. Scientific American, 280(1):112–113, January 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n1/pdf/scientificamerican0199-112.pdf.

Anonymous:1999:ERa

[4877] Anonymous. The Editors recommend. Scientific American, 280 (1):113–114, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
113.pdf.

Anonymous:1999:ERBa

[4878] Anonymous. The editors recommend: Books on animal intelligence, mummies, invisible computers and more. Scientific American, 280(1): 113–114, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:NW

[4879] James Burke. And now the weather. Scientific American, 280 (1):115–117, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
115.pdf.

Burke:1999:CNW

[4880] James Burke. Connections: And now the weather. Scientific American, 280(1):115, 117 (Intl. ed. 115–??), January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:CPR

[4881] James Burke. Connections: Pendulums, radioactivity and the Suez Canal. Scientific American, 280(1):115–??, January 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1999:MC

[4882] Philip Morrison and Phylis Morrison. Mound and crater. Scientific American, 280(1):116–117, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v280/n1/pdf/scientificamerican0199-
116.pdf.

Morrison:1999:WMC

[4883] Philip Morrison and Phylis Morrison. Wonders: Mound and crater. Scientific American, 280(1):116–117, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mohr:1999:WK

[4884] Michael Mohr. Working knowledge. Scientific American, 280(1):118–119, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n1/pdf/scientificamerican0199-118.pdf.

Mohr:1999:WKG

[4885] Michael Mohr. Working knowledge: Grand piano. Scientific American, 280(1):118–119, January 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1999:EWG

[4886] John Rennie. From the Editor: Worm gets the early bird. Scientific American, 280(2):8, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
8.pdf.

Anonymous:1999:LEc

[4887] Anonymous. Letters to the Editors. Scientific American, 280(2):10–12, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-10.pdf.

Anonymous:1999:Ec

[4888] Anonymous. Erratum. Scientific American, 280(2):12, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n2/pdf/scientificamerican0299-12.pdf.

Anonymous:1999:YAIb

[4889] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 280(2):14, February 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n2/pdf/scientificamerican0299-14.pdf.

Anonymous:1999:KSE

[4890] Anonymous. Key space explorations of the next decade. Scientific American, 280(2):18–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Strauss:1999:NAP

[4891] Evelyn Strauss. News and analysis: Pretesting tumors. Scientific American, 280(2):19–22, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
19.pdf.

Anonymous:1999:FPT

[4892] Anonymous. In focus: Pretesting tumor therapies remains controversial. Scientific American, 280(2):19–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1999:ISS

[4893] Tim Beardsley. The International Space Station: A work in progress. Scientific American, 280(2):20–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Musser:1999:TMD

[4894] George Musser. Taming Maxwell’s Demon. Scientific American, 280(2):24, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
24.pdf.

Anonymous:1999:SCM

[4895] Anonymous. Science and the citizen: Microrotors and Maxwell’s demon. Suppressing anti-nuke protesters. Ants against elephants. Scientific American, 280(2):24–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Slakey:1999:RVH

[4896] Francis Slakey. Robots vs. humans: Who should explore space? Scientific American, 280(2):24–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Spudis:1999:RVH

[4897] Paul D. Spudis. Robots vs. humans: Who should explore space? Scientific American, 280(2):24–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hayashi:1999:AFA

[4898] Alden M. Hayashi. Attack of the fire ants. Scientific American, 280(2):26–28, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
26b.pdf.

Yam:1999:Bb

[4899] Philip Yam. In brief. Scientific American, 280(2):26–30, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n2/pdf/scientificamerican0299-26a.pdf.

Anonymous:1999:NAI

[4900] Anonymous. News and analysis: Invading fire ants. Scientific American, 280(2):26–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Mukerjee:1999:BF

[4901] Madhusree Mukerjee. Blast fallout. Scientific American, 280(2):28–30, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-28.pdf.

Mirsky:1999:OT

[4902] Steve Mirsky. This is only a test. Scientific American, 280(2):29, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-29.pdf.

Gibbs:1999:CP

[4903] W. Wayt Gibbs. On cemetery pond. Scientific American, 280(2):30–33, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-30.pdf.

Doyle:1999:TTC

[4904] Rodger Doyle. Tranquillity, turmoil and chaos for international business. Scientific American, 280(2):32, February 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n2/pdf/scientificamerican0299-32.pdf.

Mirsky:1999:BP

[4905] Steve Mirsky. Bones to pick. Scientific American, 280(2):36–38, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-36.pdf.

Anonymous:1999:PPD

[4906] Anonymous. Profile: Paleoanthropologist Dennis Stanford shreds a mammoth. Scientific American, 280(2):36–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1999:C

[4907] W. Wayt Gibbs. Not cleaning up. Scientific American, 280(2):39–40, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-39.pdf.

Anonymous:1999:TBW

[4908] Anonymous. Technology and business: Why pollution cleanups stall. liquid air for young lungs. RNA vaccines. Scientific American, 280(2): 39–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1999:BFL

[4909] W. Wayt Gibbs. Breath of fresh liquid. Scientific American, 280(2):40–41, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
40.pdf.

Zorpette:1999:WNM

[4910] Glenn Zorpette. What’s next for Mars. Scientific American, 280(2): 40–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1999:LTI

[4911] Tim Beardsley. A laser in tune with itself. Scientific American, 280(2):41–42, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
41.pdf.

Beardsley:1999:II

[4912] Tim Beardsley. Innovative immunity. Scientific American, 280(2):42–44, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-42.pdf.

Grossman:1999:PP

[4913] Wendy M. Grossman. Private parts. Scientific American, 280(2):44–45, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-44.pdf.

Anonymous:1999:CVL

[4914] Anonymous. Cyber view: On-line privacy guarantees pit the U.S. against Europe. Scientific American, 280(2):44–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kahabka:1999:SXR

[4915] Peter Kahabka, Edward P. J. van den Heuvel, and Saul A. Rappaport. Supersoft X-ray stars and supernovae. Scientific American, 280(2):46–53, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-46.pdf.

Zubrin:1999:SHM

[4916] Robert Zubrin. Sending humans to Mars. Scientific American, 280(2): 46–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

McKay:1999:BLM

[4917] Christopher P. McKay. Bringing life to Mars. Scientific American, 280 (2):52–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Cooper:1999:PHA

[4918] Richard S. Cooper, Charles N. Rotimi, and Ryk Ward. The puzzle of hypertension in African–Americans. Scientific American, 280(2):56–62, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-56.pdf;   http:/
/www.sciam.com/1999/0209issue/0299cooper.html.

C:1999:HBP

[4919] R. S. C., C. N. R., and R. W. High blood pressure and the slave trade. Scientific American, 280(2):63, February 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n2/pdf/scientificamerican0299-63.pdf.

Stiassny:1999:CRL

[4920] Melanie L. J. Stiassny and Axel Meyer. Cichlids of the Rift lakes. Scientific American, 280(2):64–69, February 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n2/pdf/scientificamerican0299-64.pdf.

Mandelbrot:1999:MWW

[4921] Benoit B. Mandelbrot. A multifractal walk down Wall Street. Scientific American, 280(2):70–73, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
70.pdf.

Riddle:1999:HLD

[4922] Robert D. Riddle and Clifford J. Tabin. How limbs develop. Scientific American, 280(2):74–79, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
74.pdf.

Beardsley:1999:WGS

[4923] Tim Beardsley. The way to go in space. Scientific American, 280(2):80–97, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
80.pdf;   http://www.sciam.com/1999/0209issue/0299beardsley.
html.

McClinton:1999:ABE

[4924] Charles R. McClinton. Air-breathing engines. Scientific American, 280(2):84–85, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
84.pdf;   http://www.sciam.com/1999/0209issue/0299beardsleybox2.
html.

Forward:1999:ST

[4925] Robert L. Forward and Robert P. Hoyt. Space tethers. Scientific American, 280(2):86–87, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
86.pdf;   http://www.sciam.com/1999/0209issue/0299beardsleybox3.
html.

Myrabo:1999:HLa

[4926] Leik N. Myrabo. Highways of light. Scientific American, 280(2):88–89, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-88.pdf.

Ferris:1999:ISC

[4927] Timothy Ferris. Interstellar spaceflight: Can we travel to other stars? Scientific American, 280(2):88–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Myrabo:1999:HLb

[4928] Leik N. Myrabo. Highways of light. Scientific American, 280(2):88–?? (Intl. ed. 66–??), February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1999/

0209issue/0299beardsleybox4.html.

Harris:1999:LS

[4929] Henry M. Harris. Light sails. Scientific American, 280(2):90–91, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-90.pdf;   http:/
/www.sciam.com/1999/0209issue/0299beardsleybox5.html.

Powell:1999:CNRa

[4930] James R. Powell. Compact nuclear rockets. Scientific American, 280(2):92–93, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
92.pdf.

Alpert:1999:MMS

[4931] Mark Alpert. Making money in space. Scientific American, 280(2):92–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powell:1999:CNRb

[4932] James R. Powell. Compact nuclear rockets. Scientific American, 280 (2):92–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1999/0209issue/
0299beardsleybox6.html.

Leifer:1999:RS

[4933] Stephanie D. Leifer. Reaching for the stars. Scientific American, 280(2):94–95, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
94.pdf;   http://www.sciam.com/1999/0209issue/0299beardsleybox7.
html.

Evans:1999:NSP

[4934] John V. Evans. New satellites for personal communications. Scientific American, 280(2):96–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carlson:1999:ASTb

[4935] Shawn Carlson. The amateur scientist: Tackling the triple point. Scientific American, 280(2):98–99, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
98.pdf.

Anonymous:1999:ASC

[4936] Anonymous. The amateur scientist: Capturing the three phases of water in one bottle. Scientific American, 280(2):98–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:1999:MRO

[4937] Ian Stewart. Mathematical recreations: Origami tessellations. Scientific American, 280(2):100–101, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
100.pdf.

Anonymous:1999:MRO

[4938] Anonymous. Mathematical recreations: Origami gets practical. Scientific American, 280(2):100–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lewis:1999:TWS

[4939] John S. Lewis. Tapping the waters of space. Scientific American, 280 (2):100–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:RCa

[4940] Anonymous. Reviews and commentaries. Scientific American, 280 (2):102–104, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
102.pdf.

Anonymous:1999:RCO

[4941] Anonymous. Reviews and commentaries: Once upon a number: John Allen Paulos finds the mathematics in entertaining stories, and the stories in entertaining math. Scientific American, 280(2):102–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:ERBb

[4942] Anonymous. The editors recommend: Books on robots, extraterrestrial intelligence and quantum physics. Scientific American, 280(2):103–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1999:NF

[4943] Philip Morrison and Phylis Morrison. Noah’s Flood? Scientific American, 280(2):105–107, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n2/pdf/scientificamerican0299-
105.pdf.

Morrison:1999:WNF

[4944] Philip Morrison and Phylis Morrison. Wonders: Noah’s flood revealed. Scientific American, 280(2):105–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:LLN

[4945] James Burke. A light little number. Scientific American, 280(2):106–107, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-106.pdf.

Burke:1999:CBBa

[4946] James Burke. Connections: From Bordeaux to balloons. Scientific American, 280(2):106–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Shapiro:1999:WKG

[4947] Howard I. Shapiro, Jay P. Shapiro, and Lawrence K. Shapiro. Working knowledge: Giant crawler cranes. Scientific American, 280(2):108, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n2/pdf/scientificamerican0299-108.pdf.

Anonymous:1999:WKH

[4948] Anonymous. Working knowledge: How construction cranes stay upright. Scientific American, 280(2):108–??, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1999:EEI

[4949] John Rennie. From the Editors: The elite inventions. Scientific American, 280(3):8, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
8.pdf.

Anonymous:1999:LEd

[4950] Anonymous. Letters to the Editors. Scientific American, 280(3):12–14, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-12.pdf.

Anonymous:1999:EE

[4951] Anonymous. From the Editors: Erratum. Scientific American, 280(3): 14, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-14.pdf.

Anonymous:1999:YAIc

[4952] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 280(3):16, March 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-16.pdf.

Gibbs:1999:NAR

[4953] W. Wayt Gibbs. News and analysis: R X for B and C. Scientific American, 280(3):17–20, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
17.pdf.

Anonymous:1999:FND

[4954] Anonymous. In focus: New drugs and vaccines push back against hepatitis B and C. Scientific American, 280(3):17–??, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hopkin:1999:DSM

[4955] Karen Hopkin. Death to sperm mitochondria. Scientific American, 280(3):21, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
21a.pdf.

Hayashi:1999:OS

[4956] Alden M. Hayashi. On the origins of subspecies. Scientific American, 280(3):21–24, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
21b.pdf.

Yam:1999:Bc

[4957] Philip Yam. In brief. Scientific American, 280(3):24–28, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-24.pdf.

Mirsky:1999:FFF

[4958] Steve Mirsky. Feathers, flight and faith. Scientific American, 280(3):26, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-26.pdf.

Schmiedeskamp:1999:CAN

[4959] Mia Schmiedeskamp. Crimes against nature. Scientific American, 280(3):28–30, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
28.pdf.

Musser:1999:RW

[4960] George Musser. Revenge of the wimps. Scientific American, 280(3):30– 34, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-30.pdf.

Doyle:1999:DAS

[4961] Rodger Doyle. Divorce, American-style. Scientific American, 280(3):34, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-34.pdf.

Beardsley:1999:HU

[4962] Tim Beardsley. Humans unite! Scientific American, 280(3):35–36, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-35.pdf.

Holloway:1999:OW

[4963] Marguerite Holloway. Oil in water. Scientific American, 280(3):38, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-38.pdf.

Nemecek:1999:GG

[4964] Sasha Nemecek. Green is good. Scientific American, 280(3):39–40, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-39.pdf.

Scott:1999:OSS

[4965] Phil Scott. This old Space Station. Scientific American, 280(3):40–42, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-40.pdf.

Chase:1999:RDG

[4966] Brandon D. Chase. Rushing the double-gate. Scientific American, 280(3):42–43, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
42.pdf.

Wallich:1999:CVB

[4967] Paul Wallich. Cyber view: The best things in cyberspace are free. Scientific American, 280(3):44, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-

44.pdf. Discusses the U.S. Justice Department lawsuit against Microsoft, and the Linux phenomenon.

Anonymous:1999:NMT

[4968] Anonymous. The 1998 National Medal of Technology. Scientific American, 280(3):45–49, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
45.pdf;   http://www.sciam.com/1998/0308issue/0399medals.html.

Winners: Denton A. Cooley (cardiovascular surgeon); Robert T. Fraley, Robert B. Horsch, Ernest G. Jaworski, and Stephen G. Rogers (biotechnologists); Kenneth L. Thompson and Dennis M. Ritchie (computer scientists); Biogen, Inc. (biopharmaceuticals); Bristol-Myers Squibb Company (pharmaceuticals).

Bullock:1999:GCC

[4969] Mark A. Bullock and David H. Grinspoon. Global climate change on Venus. Scientific American, 280(3):50–57, March 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-50.pdf.

Mukerjee:1999:LBB

[4970] Madhusree Mukerjee. A little Big Bang. Scientific American, 280(3):60– 65, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-60.pdf.

Smith:1999:TB

[4971] Roger Smith. The timing of birth. Scientific American, 280(3):68–75, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-68.pdf.

Smith:1999:VHE

[4972] Bradley R. Smith. Visualizing human embryos. Scientific American, 280(3):76–81, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
76.pdf.

Smith:1999:SPV

[4973] Bradley R. Smith. Science in pictures: Visualizing human embryos. Scientific American, 280(3):76–??, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1998/0308issue/0399smith.html.

Ciofi:1999:KD

[4974] Claudio Ciofi. The Komodo dragon. Scientific American, 280(3):84–91, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-84.pdf;   http:/
/www.sciam.com/1998/0308issue/0399ciofi.html.

Thomke:1999:CM

[4975] Stefan Thomke, Michael Holzner, and Touraj Gholami. The crash in the machine. Scientific American, 280(3):92–97, March 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-92.pdf.

Gorelik:1999:MAS

[4976] Gennady Gorelik. The metamorphosis of Andrei Sakharov. Scientific American, 280(3):98–101, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
98.pdf.

Carlson:1999:ASHa

[4977] Shawn Carlson. The amateur scientist: a homemade high-precision thermometer. Scientific American, 280(3):102–103, March 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-102.pdf.

Stewart:1999:MRSa

[4978] Ian Stewart. Mathematical recreations: The synchronicity of firefly flashing. Scientific American, 280(3):104–106, March 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-104.pdf.

Dawkins:1999:RCO

[4979] Richard Dawkins. Reviews and commentaries: One man’s rainbow. Scientific American, 280(3):107–109, March 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n3/pdf/scientificamerican0399-107.pdf.

Dawkins:1999:RC

[4980] Richard Dawkins. Reviews and commentaries. Scientific American, 280 (3):107–??, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:ERc

[4981] Anonymous. The Editors recommend. Scientific American, 280(3):109– 110, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-109.pdf.

Morrison:1999:WRS

[4982] Philip Morrison and Phylis Morrison. Walk, run — and skip. Scientific American, 280(3):111–113, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n3/pdf/scientificamerican0399-
111.pdf.

Morrison:1999:WPY

[4983] Philip Morrison and Phylis Morrison. Wonders: Pace yourself. Scientific American, 280(3):111–??, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:LMY

[4984] James Burke. Lend me your ear. Scientific American, 280(3):112–113, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-112.pdf.

Burke:1999:CSS

[4985] James Burke. Connections: Scarlet, the Scots and scurvy. Scientific American, 280(3):112–??, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zamboni:1999:WK

[4986] Richard F. Zamboni. Working knowledge. Scientific American, 280(3): 114, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n3/pdf/scientificamerican0399-114.pdf.

Anonymous:1999:I

[4987] Anonymous. Introduction. Scientific American, 280(4):4–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hopkin:1999:CCD

[4988] Karen Hopkin. Couture cures: This drug’s for you. Scientific American, 280(4):6–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1999:EAR

[4989] John Rennie. From the Editor: Attack of the robo-editor. Scientific American, 280(4):8, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
8.pdf.

Anonymous:1999:LEe

[4990] Anonymous. Letters to the Editors. Scientific American, 280(4):10–12, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-10.pdf.

Anonymous:1999:Ed

[4991] Anonymous. Errata. Scientific American, 280(4):12, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-12.pdf.

Anonymous:1999:YAId

[4992] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 280(4):16, April 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-16.pdf.

Beardsley:1999:NAF

[4993] Tim Beardsley. News and analysis: Facing an ill wind. Scientific American, 280(4):19–20, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
19.pdf.

Anonymous:1999:FUG

[4994] Anonymous. In focus: U.S. girds for the battle against bioweapons. Scientific American, 280(4):19–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1999:PP

[4995] Glenn Zorpette. A probing prelude. Scientific American, 280(4):22, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-22.pdf.

Musser:1999:HBY

[4996] George Musser. A hundred billion years of solitude. Scientific American, 280(4):24, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
24.pdf.

Mirsky:1999:DR

[4997] Steve Mirsky. Diamond reflections. Scientific American, 280(4):26, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-26b.pdf.

Anonymous:1999:Ba

[4998] Anonymous. In brief. Scientific American, 280(4):26–32, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-26a.pdf.

Nemecek:1999:W

[4999] Sasha Nemecek. Whiteout. Scientific American, 280(4):30–32, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-30.pdf.

Beardsley:1999:MG

[5000] Tim Beardsley. Mutations galore. Scientific American, 280(4):32–36, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-32.pdf.

Hopkin:1999:MM

[5001] Karen Hopkin. Making Methuselah. Scientific American, 280(4):32– ??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Doyle:1999:HCC

[5002] Rodger Doyle. Health care costs. Scientific American, 280(4):36, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-36.pdf.

Brown:1999:YRN

[5003] Kathryn S. Brown. Are you ready for a new sensation? Scientific American, 280(4):38–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Alpert:1999:JFG

[5004] Mark Alpert. Not just fun and games. Scientific American, 280(4):40– 42, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-40.pdf.

Anonymous:1999:PMJ

[5005] Anonymous. Profile: Mathematician John H. Conway, inventor of the game of Life. Scientific American, 280(4):40–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1999:WS

[5006] Glenn Zorpette. Waiting for the supercar. Scientific American, 280(4):46–47, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
46.pdf.

Drollette:1999:TBS

[5007] Dan Drollette. Taking ballistics by storm. Scientific American, 280(4): 50, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-50.pdf.

Scott:1999:BL

[5008] Phil Scott. A bug’s lift. Scientific American, 280(4):51–54, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-51.pdf.

Stixon:1999:JFB

[5009] Gary Stixon. Japan fields a big-league light gatherer. Scientific American, 280(4):54, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
54.pdf.

Gibbs:1999:WW

[5010] W. Wayt Gibbs. Watch the watchers. Scientific American, 280(4):55, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-55.pdf.

Kurzweil:1999:CMM

[5011] Ray Kurzweil. The coming merging of mind and machine. Scientific American, 280(4):56–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:SRP

[5012] Anonymous. Special report: The promise of tissue engineering. Scientific American, 280(4):59, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
59.pdf.

Editors:1999:PTE

[5013] The Editors. The promise of tissue engineering. Scientific American, 280(4):59–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1999/0409issue/
0499ezzell.html.

Mooney:1999:GNO

[5014] David J. Mooney and Antonios G. Mikos. Growing new organs. Scientific American, 280(4):60–65, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
60.pdf;   http://www.sciam.com/1999/0409issue/0499mooney.html.

Hamer:1999:TGB

[5015] Dean Hamer. Tweaking the genetics of behavior. Scientific American, 280(4):62–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pedersen:1999:ESCa

[5016] Roger A. Pedersen. Embryonic stem cells for medicine. Scientific American, 280(4):68–73, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
68.pdf.

Gray:1999:WRB

[5017] Stephen Gray. When off-the-rack becomes off-the-net. Scientific American, 280(4):68–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pedersen:1999:ESCb

[5018] Roger A. Pedersen. Embryonic stem cells for medicine. Scientific American, 280(4):68–?? (Intl. ed. 18–??), April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Brown:1999:SS

[5019] Kathryn S. Brown. Smart stuff. Scientific American, 280(4):72–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nadis:1999:WWD

[5020] Steve Nadis. What the well-dressed warrior will wear. Scientific American, 280(4):74–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lysaght:1999:ECTa

[5021] Michael J. Lysaght and Patrick Aebischer. Encapsulated cells as therapy. Scientific American, 280(4):76–82, April 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-76.pdf.

Agnew:1999:WWO

[5022] Bruce Agnew. Will we be one nation, indivisible? Scientific American, 280(4):76–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lysaght:1999:ECTb

[5023] Michael J. Lysaght and Patrick Aebischer. Encapsulated cells as therapy. Scientific American, 280(4):76–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Green:1999:C

[5024] Ronald M. Green. I, clone. Scientific American, 280(4):80–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:SFT

[5025] Anonymous. Skin: The first tissue-engineered products. Scientific American, 280(4):83–85, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
83.pdf.

Parenteau:1999:SFT

[5026] Nancy Parenteau and Gail Naughton. Skin: The first tissue-engineered products. Scientific American, 280(4):83–??, April 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Joseph:1999:LT

[5027] Patrick Joseph. Living in technology. Scientific American, 280(4):84– ??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Langer:1999:TEC

[5028] Robert S. Langer and Joseph P. Vacanti. Tissue engineering: The challenges ahead. Scientific American, 280(4):86–89, April 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-86.pdf.

Kling:1999:FF

[5029] Jim Kling. Future feast. Scientific American, 280(4):88–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Luminet:1999:SF

[5030] Jean-Pierre Luminet, Glenn D. Starkman, and Jeffrey R. Weeks. Is space finite? Scientific American, 280(4):90–97, April 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-90.pdf;   http://www.sciam.com/
1999/0409issue/0499weeks.html.

Kling:1999:NM

[5031] Jim Kling. The new metropolis. Scientific American, 280(4):92–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Copeland:1999:ATF

[5032] B. Jack Copeland and Diane Proudfoot. Alan Turing’s forgotten ideas in computer science. Scientific American, 280(4):98–103, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-98.pdf.

Mirsky:1999:FS

[5033] Steve Mirsky. Future schlock. Scientific American, 280(4):100–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stix:1999:ENE

[5034] Gary Stix. Expeditions: A new eye opens on the cosmos. Scientific American, 280(4):104–111, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
104.pdf.

Vreeland:1999:RCC

[5035] James M. Vreeland, Jr. The revival of colored cotton. Scientific American, 280(4):112–118, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
112.pdf.

Carlson:1999:ASDa

[5036] Shawn Carlson. The amateur scientist: Detecting “hot” clouds. Scientific American, 280(4):120–122, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
120.pdf.

Stewart:1999:MRT

[5037] Ian Stewart. Mathematical recreations: Tangling with topology. Scientific American, 280(4):123–124, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
123.pdf.

Anonymous:1999:RCb

[5038] Anonymous. Reviews and commentaries. Scientific American, 280(4):125–127, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
125.pdf.

Anonymous:1999:ERd

[5039] Anonymous. The Editors recommend. Scientific American, 280(4):127– 128, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n4/pdf/scientificamerican0499-127.pdf.

Morrison:1999:CA

[5040] Philip Morrison and Phylis Morrison. Circumnavigators all. Scientific American, 280(4):129–131, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
129.pdf.

Morrison:1999:WQT

[5041] Philip Morrison and Phylis Morrison. Wonders: Quick trips around the world. Scientific American, 280(4):129–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:Za

[5042] James Burke. Zzzzzzz. Scientific American, 280(4):130–131, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n4/pdf/scientificamerican0499-130.pdf.

Burke:1999:CPL

[5043] James Burke. Connections: Phrenologists and lunatics. Scientific American, 280(4):130–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bloomfield:1999:WKF

[5044] Louis A. Bloomfield. Working knowledge: The flight of the Frisbee. Scientific American, 280(4):132, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n4/pdf/scientificamerican0499-
132.pdf.

Anonymous:1999:WKF

[5045] Anonymous. Working knowledge: The flight of the Frisbee. Scientific American, 280(4):132–??, April 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rennie:1999:ESW

[5046] John Rennie. From the Editors: Survivors and a winner. Scientific American, 280(5):6, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
6.pdf.

Anonymous:1999:LEf

[5047] Anonymous. Letters to the Editors. Scientific American, 280(5):8– 10, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-8.pdf.

Anonymous:1999:YAIe

[5048] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 280(5):12, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
12.pdf.

Gibbs:1999:FMA

[5049] W. Wayt Gibbs. In focus: The magnetic attraction. Scientific American, 280(5):17–18, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). This brief article describes work by Richard M. Lienau and Laurence P. Sadwick (University of Utah electrical engineer) on magnetic memory chips capable of holding data for years without external power, operating at a speed of five to ten times that of current DRAM technology, and costing about the same as conventional DRAM.

Gibbs:1999:NAM

[5050] W. Wayt Gibbs. News and analysis: The magnetic attraction. Scientific American, 280(5):17–18, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
17.pdf.

Anonymous:1999:FPM

[5051] Anonymous. In focus: Persistence of memory: magnetic RAM shakes up electronics industry. Scientific American, 280(5):17–??, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Musser:1999:HCS

[5052] George Musser. Here come the suns. Scientific American, 280(5):20, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-20.pdf.

Mirsky:1999:SD

[5053] Steve Mirsky. Supply and demand. Scientific American, 280(5):22, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-22b.pdf.

Anonymous:1999:Bb

[5054] Anonymous. In brief. Scientific American, 280(5):22–26, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n5/pdf/scientificamerican0599-22a.pdf.

Nelson:1999:PP

[5055] Roxanne Nelson. Prevented prevention. Scientific American, 280(5):24– 26, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-24.pdf.

Netting:1999:WE

[5056] Jessa Netting. Wink of an eye. Scientific American, 280(5):26–28, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-26.pdf.

Gibbs:1999:FD

[5057] W. Wayt Gibbs. Field of dreams. Scientific American, 280(5):28–30, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-28.pdf.

Anonymous:1999:LUU

[5058] Anonymous. Labor unions under stress. Scientific American, 280(5): 30, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-30.pdf.

Beardsley:1999:MCE

[5059] Tim Beardsley. A medical crusader for editorial freedom. Scientific American, 280(5):32–33, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
32.pdf.

Firth:1999:LBT

[5060] Peta Firth. Leaving a bad taste. Scientific American, 280(5):34–35, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-34.pdf.

Collins:1999:QS

[5061] Graham P. Collins. Quantum sculpting. Scientific American, 280(5):35– 36, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-35.pdf.

Ezzell:1999:AR

[5062] Carol Ezzell. Adam’s rib? Scientific American, 280(5):36–37, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n5/pdf/scientificamerican0599-36.pdf.

Gibbs:1999:M

[5063] W. Wayt Gibbs. Micromicrophones. Scientific American, 280(5):37– 38, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-37.pdf.

Grossman:1999:PSM

[5064] Wendy M. Grossman. Putting the squeeze on music. Scientific American, 280(5):38–39, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
38.pdf.

Lasota:1999:UBH

[5065] Jean-Pierre Lasota. Unmasking black holes. Scientific American, 280(5): 40–47, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-40.pdf.

Kempermann:1999:NNC

[5066] Gerd Kempermann and Fred H. Gage. New nerve cells for the adult brain. Scientific American, 280(5):48–53, May 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n5/pdf/scientificamerican0599-48.pdf.

Gonzalez:1999:T

[5067] Frank I. González. Tsunami! Scientific American, 280(5):56–65, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-56.pdf;   http:/
/www.sciam.com/1999/0599issue/0599gonzalez.html.

Wroe:1999:KKO

[5068] Stephen Wroe. Killer kangaroos and other murderous marsupials. Scientific American, 280(5):68–74, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
68.pdf.

Kim:1999:AFC

[5069] Eugene Eric Kim and Betty Alexandra Toole. Ada and the first computer. Scientific American, 280(5):76–81, May 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n5/pdf/scientificamerican0599-76.pdf. The collaboration between

Ada, countess of Lovelace, and computer pioneer Charles Babbage resulted in a landmark publication that described how to program the world’s first computer.

Venkateswar:1999:AI

[5070] Sita Venkateswar. The Andaman Islanders. Scientific American, 280(5): 82–88, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-82.pdf.

Bosak:1999:XSG

[5071] Jon Bosak and Tim Bray. XML and the second-generation Web. Scientific American, 280(5):89–93, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
89.pdf;   http://www.sciam.com/1999/0599issue/0599bosak.html.

Carlson:1999:ASHb

[5072] Shawn Carlson. The amateur scientist: Hot views of the microscopic world. Scientific American, 280(5):94–97, May 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n5/pdf/scientificamerican0599-94.pdf.

Stewart:1999:MRP

[5073] Ian Stewart. Mathematical recreations: a puzzle for pirates. Scientific American, 280(5):98–99, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
98.pdf.

Smuts:1999:RCS

[5074] Barbara Smuts. Reviews and commentaries: Sanctifying the Cosmos. Scientific American, 280(5):100–102, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
100.pdf.

Anonymous:1999:ERe

[5075] Anonymous. The Editors recommend. Scientific American, 280(5):103– 104, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-103.pdf.

Morrison:1999:GC

[5076] Philip Morrison and Phylis Morrison. Galileo’s cicada. Scientific American, 280(5):105–107, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n5/pdf/scientificamerican0599-
105.pdf.

Morrison:1999:WPG

[5077] Philip Morrison and Phylis Morrison. Wonders: A parable by Galileo. Scientific American, 280(5):105–??, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:EC

[5078] James Burke. Entente cordiale. Scientific American, 280(5):106–107, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-106.pdf.

Burke:1999:CMM

[5079] James Burke. Connections: Mach, magnetism and minting. Scientific American, 280(5):106–??, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Garavito:1999:WK

[5080] R. Michael Garavito. Working knowledge. Scientific American, 280(5): 108, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n5/pdf/scientificamerican0599-108.pdf.

Rennie:1999:EAB

[5081] John Rennie. From the Editors: The adventures of bat men. Scientific American, 280(6):4, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
4.pdf.

Anonymous:1999:LEg

[5082] Anonymous. Letters to the Editors. Scientific American, 280(6):6–8, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-6.pdf.

Anonymous:1999:YAIf

[5083] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 280(6):10, June 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n6/pdf/scientificamerican0699-10.pdf.

Yam:1999:NAE

[5084] Philip Yam. News and analysis: Explosive reactions. Scientific American, 280(6):13–14, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
13.pdf.

Anonymous:1999:FON

[5085] Anonymous. In focus: Overreaction to nuclear secret leaks may hamper science. Scientific American, 280(6):13–??, June 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1999:DS

[5086] W. Wayt Gibbs. A diabetes switch? Scientific American, 280(6):16, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-16.pdf.

Mirsky:1999:SF

[5087] Steve Mirsky. Semper fly. Scientific American, 280(6):18, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n6/pdf/scientificamerican0699-18b.pdf.

Anonymous:1999:Bc

[5088] Anonymous. In brief. Scientific American, 280(6):18–22, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n6/pdf/scientificamerican0699-18a.pdf.

Alpert:1999:LSC

[5089] Mark Alpert. The little spacecraft that could. Scientific American, 280(6):20–22, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
20.pdf.

Ezzell:1999:ICD

[5090] Carol Ezzell. It came from the deep. Scientific American, 280(6):22–24, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-22.pdf.

Stix:1999:RNW

[5091] Gary Stix. Rediscovering the New World. Scientific American, 280(6): 24, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-24a.pdf.

Beardsley:1999:GW

[5092] Tim Beardsley. Getting wired. Scientific American, 280(6):24–25, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-24b.pdf.

Doyle:1999:IIU

[5093] Rodger Doyle. Income inequality in the U.S. Scientific American, 280(6):25–26, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
25.pdf.

Nemecek:1999:WGH

[5094] Sasha Nemecek. When good health is good business. Scientific American, 280(6):28–30, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
28.pdf.

Williams:1999:TT

[5095] Wendy Williams. Turtle tragedy. Scientific American, 280(6):32–33, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-32.pdf.

Beardsley:1999:SBL

[5096] Tim Beardsley. Seeing the breath of life. Scientific American, 280(6):33– 34, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-33.pdf.

Dupont:1999:RSW

[5097] Daniel G. Dupont. The real Star Wars. Scientific American, 280(6):36, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-36.pdf.

Wallich:1999:YCW

[5098] Paul Wallich. Your 0.002 cent’s worth. Scientific American, 280(6):37, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-37.pdf.

Landy:1999:MU

[5099] Stephen D. Landy. Mapping the Universe. Scientific American, 280(6):38–45, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
38.pdf;   http://www.sciam.com/1999/0699issue/0699landy.html.

Belmonte:1999:HBT

[5100] Juan Carlos Izpisúa Belmonte. How the body tells left from right. Scientific American, 280(6):46–51, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
46.pdf.

Project:1999:HW

[5101] Members of the Clever Project. Hypersearching the Web. Scientific American, 280(6):54–60, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
54.pdf;   http://www.sciam.com/1999/0699issue/0699raghavan.
html.

Grimson:1999:IGS

[5102] W. Eric L. Grimson, Ron Kikinis, Ferenc A. Jolesz, and Peter McL. Black. Image-guided surgery. Scientific American, 280(6):62–69, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-62.pdf.

Rogers:1999:BWA

[5103] Paul Rogers, Simon Whitby, and Malcolm Dando. Biological warfare against crops. Scientific American, 280(6):70–75, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n6/pdf/scientificamerican0699-70.pdf.

Dawson:1999:GLL

[5104] John W. Dawson, Jr. Gödel and the limits of logic. Scientific American, 280(6):76–81, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
76.pdf.

Zorpette:1999:ECG

[5105] Glenn Zorpette. Expeditions: Chasing the ghost bat. Scientific American, 280(6):82–89, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
82.pdf.

Zorpette:1999:CGB

[5106] Glenn Zorpette. Chasing the ghost bat. Scientific American, 280(6):82– ??, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carlson:1999:ASE

[5107] Shawn Carlson. The amateur scientist: Expert secrets for preserving plants. Scientific American, 280(6):90–92, June 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n6/pdf/scientificamerican0699-90.pdf.

Stewart:1999:MRCa

[5108] Ian Stewart. Mathematical recreations: Crossed lines in the brick factory. Scientific American, 280(6):94–96, June 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v280/
n6/pdf/scientificamerican0699-94.pdf.

Hamer:1999:RCF

[5109] Dean Hamer. Reviews and commentaries: of flies and men. Scientific American, 280(6):98–99, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
98.pdf.

Anonymous:1999:ERf

[5110] Anonymous. The Editors recommend. Scientific American, 280(6):99– 100, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-99.pdf.

Morrison:1999:FCT

[5111] Philip Morrison and Phylis Morrison. First comes the thunder. Scientific American, 280(6):101–103, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
101.pdf.

Burke:1999:RHV

[5112] James Burke. Room with (half) a view. Scientific American, 280(6):102– 103, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v280/n6/pdf/scientificamerican0699-102.pdf.

Burke:1999:CRB

[5113] James Burke. Connections: The railroad bridge to the “planets”. Scientific American, 280(6):102–??, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bloomfield:1999:WKI

[5114] Louis A. Bloomfield. Working knowledge: Instant glue. Scientific American, 280(6):104, June 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v280/n6/pdf/scientificamerican0699-
104.pdf.

Rennie:1999:EGG

[5115] John Rennie. From the Editor: a generation of genius. Scientific American, 281(1):6, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
6.pdf.

Anonymous:1999:LEh

[5116] Anonymous. Letters to the Editors. Scientific American, 281(1):8– 10, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-8.pdf.

Anonymous:1999:Ef

[5117] Anonymous. Erratum. Scientific American, 281(1):10, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-10.pdf.

Anonymous:1999:YAIg

[5118] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 281(1):12, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
12.pdf.

Stix:1999:NAS

[5119] Gary Stix. News and analysis: Seeking common ground. Scientific American, 281(1):15–18, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
15.pdf.

Musser:1999:MSW

[5120] George Musser. Make science, not war. Scientific American, 281(1): 22, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-22.pdf.

Nemecek:1999:DC

[5121] Sasha Nemecek. Discerning CERN. Scientific American, 281(1):24, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-24.pdf.

Doyle:1999:CD

[5122] Rodger Doyle. Christian differences. Scientific American, 281(1):26, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-26.pdf.

Ezzell:1999:DV

[5123] Carol Ezzell. Death of a vaccine? Scientific American, 281(1):27, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-27.pdf.

Mirsky:1999:SWF

[5124] Steve Mirsky. Soyuz wanna fly in space. Scientific American, 281(1): 28, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-28b.pdf.

Anonymous:1999:Bd

[5125] Anonymous. In brief. Scientific American, 281(1):28–30, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-28a.pdf.

Beardsley:1999:SCC

[5126] Tim Beardsley. Stem cells come of age. Scientific American, 281(1):30– 31, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-30.pdf.

Hayashi:1999:PB

[5127] Alden M. Hayashi. Pinker and the brain. Scientific American, 281(1):32– 34, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-32.pdf.

Stix:1999:PC

[5128] Gary Stix. Parsing cells. Scientific American, 281(1):35–36, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-35.pdf.

Scott:1999:LS

[5129] Phil Scott. Lots in space. Scientific American, 281(1):36–38, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-36.pdf.

Musser:1999:PF

[5130] George Musser. Practical fractals. Scientific American, 281(1):38, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-38.pdf.

Editors:1999:BPL

[5131] The Editors. Bonus poster: Looking back at Apollo. Scientific American, 281(1):40, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
40A.pdf.

Gibbs:1999:HM

[5132] W. Wayt Gibbs. Holey magic. Scientific American, 281(1):40, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-40.pdf.

Anonymous:1999:LBA

[5133] Anonymous. Looking back at Apollo. Scientific American, 281(1):40– ??, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Grossman:1999:NAL

[5134] Wendy M. Grossman. News and analysis: On-line U. Scientific American, 281(1):41, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
41.pdf.

Bernstein:1999:LFF

[5135] Max P. Bernstein, Scott A. Sandford, and Louis J. Allamandola. Life’s far-flung raw materials. Scientific American, 281(1):42–49, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-42.pdf;   http:/
/www.sciam.com/1999/0799issue/0799bernstein.html.

Weiner:1999:GV

[5136] David B. Weiner and Ronald C. Kennedy. Genetic vaccines. Scientific American, 281(1):50–57, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
50.pdf;   http://www.sciam.com/1999/0799issue/0799weiner.html.

Rith:1999:MNS

[5137] Klaus Rith and Andreas Schäfer. The mystery of nucleon spin. Scientific American, 281(1):58–63, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
58.pdf.

Foster:1999:SPE

[5138] Karen Polinger Foster. Science in pictures: The earliest zoos and gardens. Scientific American, 281(1):64–71, July 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-64.pdf.

Foster:1999:EZG

[5139] Karen Polinger Foster. The earliest zoos and gardens. Scientific American, 281(1):64–??, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Editors:1999:FFC

[5140] The Editors. The future of fuel cells. Scientific American, 281(1):72– 73, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-72.pdf.

Anonymous:1999:FFC

[5141] Anonymous. The future of fuel cells. Scientific American, 281(1):72– ??, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Appleby:1999:EEV

[5142] A. John Appleby. The electrochemical engine for vehicles. Scientific American, 281(1):74–79, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
74.pdf.

Lloyd:1999:PPY

[5143] Alan C. Lloyd. The power plant in your basement. Scientific American, 281(1):80–86, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
80.pdf.

Dyer:1999:RBP

[5144] Christopher K. Dyer. Replacing the battery in portable electronics. Scientific American, 281(1):88–93, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
88.pdf.

Carlson:1999:ASDb

[5145] Shawn Carlson. The amateur scientist: Detecting the Earth’s electricity. Scientific American, 281(1):94–95, July 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n1/pdf/scientificamerican0799-94.pdf.

Stewart:1999:MRA

[5146] Ian Stewart. Mathematical recreations: The art of elegant tiling. Scientific American, 281(1):96–98, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
96.pdf.

Anonymous:1999:RCc

[5147] Anonymous. Reviews and commentaries. Scientific American, 281(1):100–102, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
100.pdf.

Morrison:1999:HCR

[5148] Philip Morrison. The hidden cosmic ruckus. Scientific American, 281(1):104–106, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
104.pdf.

Morrison:1999:WIR

[5149] Philip Morrison. Wonders: The icy rubble of creation. Scientific American, 281(1):104–??, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:FN

[5150] James Burke. A few notes. Scientific American, 281(1):105–106, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n1/pdf/scientificamerican0799-105.pdf.

Burke:1999:CBBb

[5151] James Burke. Connections: Binoculars, barometers and Bell’s brainstorm. Scientific American, 281(1):105–??, July 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zambelli:1999:WKA

[5152] George R. Zambelli, Sr. Working knowledge: Aerial fireworks. Scientific American, 281(1):108, July 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n1/pdf/scientificamerican0799-
108.pdf.

Rennie:1999:EDW

[5153] John Rennie. From the Editor: The detectives wore white. Scientific American, 281(2):4, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
4.pdf.

Anonymous:1999:LEi

[5154] Anonymous. Letters to the Editors. Scientific American, 281(2):6–8, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-6.pdf.

Anonymous:1999:YAIh

[5155] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 281(2):10, August 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-10.pdf.

Wong:1999:NAA

[5156] Kate Wong. News and analysis: Is out of Africa going out the door? Scientific American, 281(2):13–14, August 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-13.pdf.

Anonymous:1999:FDM

[5157] Anonymous. In focus: Did modern humans evolve only in Africa? new results cast doubts. Scientific American, 281(2):13–??, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Alpert:1999:RE

[5158] Mark Alpert. Replumbing the Everglades. Scientific American, 281(2):16–18, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
16.pdf.

Simpson:1999:MW

[5159] Sarah Simpson. Making waves. Scientific American, 281(2):18, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-18.pdf.

Zorpette:1999:PF

[5160] Glenn Zorpette. A pound of flesh. Scientific American, 281(2):20, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-20.pdf.

Mirsky:1999:TOB

[5161] Steve Mirsky. Thinking outside the box. Scientific American, 281(2):22, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-22b.pdf.

Anonymous:1999:Be

[5162] Anonymous. In brief. Scientific American, 281(2):22–24, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-22a.pdf.

Reed:1999:PL

[5163] Christina Reed. A patch for love. Scientific American, 281(2):23, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-23.pdf.

Collins:1999:QCa

[5164] Graham P. Collins. Qubit chip. Scientific American, 281(2):24, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-24.pdf.

Doyle:1999:BBU

[5165] Rodger Doyle. Behind bars in the U.S. and Europe. Scientific American, 281(2):25, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
25.pdf.

Stix:1999:SS

[5166] Gary Stix. Of survival and science. Scientific American, 281(2):26–27, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-26.pdf.

Stix:1999:BE

[5167] Gary Stix. The butterfly effect. Scientific American, 281(2):28–29, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-28.pdf.

Stix:1999:TGT

[5168] Gary Stix. Too good to be true. Scientific American, 281(2):29–31, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-29.pdf.

Goodman:1999:CP

[5169] Brenda DeKoker Goodman. Calculating pie. Scientific American, 281(2): 30, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-30.pdf.

Netting:1999:VGS

[5170] Jessa Netting. Viral gene screen. Scientific American, 281(2):31–32, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-31.pdf.

Wallich:1999:HSMa

[5171] Paul Wallich. How to steal millions in chump change. Scientific American, 281(2):32–33, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
32.pdf.

Lewis:1999:WNM

[5172] George N. Lewis, Theodore A. Postol, and John Pike. Why National Missile Defense won’t work. Scientific American, 281(2):36–41, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-36.pdf.

Burkholder:1999:LPP

[5173] JoAnn M. Burkholder. The lurking perils of Pfiesteria. Scientific American, 281(2):42–49, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
42.pdf.

Dertouzos:1999:OPF

[5174] Michael L. Dertouzos. The Oxygen Project: The future of computing. Scientific American, 281(2):52–55, August 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-52.pdf.

Anonymous:1999:SRO

[5175] Anonymous. Special report: The Oxygen Project. Scientific American, 281(2):52–??, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dertouzos:1999:FC

[5176] Michael L. Dertouzos. The future of computing. Scientific American, 281(2):52–??, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://www.sciam.com/1999/0809issue/
0899dertouzos.html.

Zue:1999:TYC

[5177] Victor Zue. Talking with your computer. Scientific American, 281(2):56– 57, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-56.pdf;   http:/
/www.sciam.com/1999/0809issue/0899zue.html.

Guttag:1999:CC

[5178] John V. Guttag. Communications chameleons. Scientific American, 281(2):58–59, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
58.pdf;   http://www.sciam.com/1999/0809issue/0899guttag.html.

Agarwal:1999:RC

[5179] Anant Agarwal. Raw computation. Scientific American, 281(2):60–63, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-60.pdf;   http:/
/www.sciam.com/1999/0809issue/0899agarwal.html.

Kearns:1999:DMN

[5180] Edward Kearns, Takaaki Kajita, and Yoji Totsuka. Detecting massive neutrinos. Scientific American, 281(2):64–71, August 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-64.pdf.

Damon:1999:MDC

[5181] William Damon. The moral development of children. Scientific American, 281(2):72–78, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
72.pdf.

Gibbs:1999:ETV

[5182] W. Wayt Gibbs. Expeditions: Trailing a virus. Scientific American, 281(2):80–87, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
80.pdf.

Gibbs:1999:TV

[5183] W. Wayt Gibbs. Trailing a virus. Scientific American, 281(2):80–??, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Carlson:1999:ASS

[5184] Shawn Carlson. The amateur scientist: Sun of a gun. Scientific American, 281(2):88–89, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
88.pdf.

Stewart:1999:MRSb

[5185] Ian Stewart. Mathematical recreations: Sierpinski’s ubiquitous gasket. Scientific American, 281(2):90–91, August 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n2/pdf/scientificamerican0899-90.pdf.

Anonymous:1999:RCd

[5186] Anonymous. Reviews and commentaries. Scientific American, 281(2):92– 95, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-92.pdf.

Morrison:1999:DA

[5187] Philip Morrison and Phylis Morrison. Dining on ammonia. Scientific American, 281(2):96–98, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
96.pdf.

Burke:1999:SI

[5188] James Burke. Sound ideas. Scientific American, 281(2):97–98, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n2/pdf/scientificamerican0899-97.pdf.

Bloomfield:1999:WKA

[5189] Louis A. Bloomfield. Working knowledge: Air conditioners. Scientific American, 281(2):100, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n2/pdf/scientificamerican0899-
100.pdf.

Wallich:1999:HSMb

[5190] Paul Wallich. How to steal millions in chump change. Scientific American, 281(2):??, August 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1999/ 0899issue/0899cyber.html. Describes an off-shore credit card fraud scheme that in late 1998 accounted for as much as 4%of Visa chargebacks, and stole as much as US$45M from bogus US$19.95 transactions.

Rennie:1999:EFB

[5191] John Rennie. From the Editor: Follow the bouncing planet. Scientific American, 281(3):4, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
4.pdf.

Anonymous:1999:LEj

[5192] Anonymous. Letters to the Editors. Scientific American, 281(3):6–8, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-6.pdf.

Anonymous:1999:YAIi

[5193] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 281(3):10, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-10.pdf.

Alpert:1999:NAF

[5194] Mark Alpert. News and analysis: The fallout from Cassini. Scientific American, 281(3):13–14, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
13.pdf.

Anonymous:1999:FCP

[5195] Anonymous. In focus: The Cassini probe’s flyby of Earth prompts antinuke protests. Scientific American, 281(3):13–??, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1999:CC

[5196] Marguerite Holloway. Congo City. Scientific American, 281(3):17–18, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-17.pdf.

Reed:1999:SS

[5197] Christina Reed. Silicone safe. Scientific American, 281(3):18, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-18.pdf.

Beardsley:1999:CI

[5198] Tim Beardsley. Calculating immunity. Scientific American, 281 (3):20–21, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
20.pdf.

Anonymous:1999:Bf

[5199] Anonymous. In brief. Scientific American, 281(3):21–26, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-21.pdf.

Scott:1999:SZ

[5200] Phil Scott. Strike zone. Scientific American, 281(3):22, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-22.pdf.

Mirsky:1999:SAD

[5201] Steve Mirsky. Strife after death. Scientific American, 281(3):24, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-24.pdf.

Musser:1999:SCB

[5202] George Musser. Skewing the cosmic bell curve. Scientific American, 281(3):26–29, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
26.pdf.

Doyle:1999:UI

[5203] Rodger Doyle. U.S. immigration. Scientific American, 281(3):28–29, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-28.pdf.

Beardsley:1999:DPK

[5204] Tim Beardsley. Defender of the plant kingdom. Scientific American, 281(3):30–31, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
30.pdf.

Dupont:1999:PS

[5205] Daniel G. Dupont. In plane sight. Scientific American, 281(3):34, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-34.pdf.

Beardsley:1999:ERS

[5206] Tim Beardsley. Enter robots, slowly. Scientific American, 281(3):36–37, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-36.pdf.

Nelson:1999:MS

[5207] Roxanne Nelson. Not making scents. Scientific American, 281(3):38–39, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-38.pdf.

Grossman:1999:WPC

[5208] Wendy M. Grossman. When publishing could mean perishing. Scientific American, 281(3):40, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
40.pdf.

Erickson:1999:BLT

[5209] Gregory M. Erickson. Breathing life into Tyrannosaurus rex. Scientific American, 281(3):42–49, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
42.pdf.

Abler:1999:TT

[5210] William L. Abler. The teeth of the tyrannosaurs. Scientific American, 281(3):50–51, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
50.pdf.

Simpson:1999:BED

[5211] George Gaylord Simpson. Book excerpt: The dechronization of Sam Magruder. Scientific American, 281(3):52–55, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-52.pdf.

Simpson:1999:DSM

[5212] George Gaylord Simpson. The dechronization of Sam Magruder. Scientific American, 281(3):52–??, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Malhotra:1999:MP

[5213] Renu Malhotra. Migrating planets. Scientific American, 281(3):56–63, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-56.pdf;   http:/
/www.sciam.com/1999/0909issue/0999malhotra.html.

McDonald:1999:RDS

[5214] John W. McDonald and the Research Consortium of the Christopher Reeve Paralysis Foundation. Repairing the damaged spinal cord. Scientific American, 281(3):64–73, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-64.pdf.

Paine:1999:CAV

[5215] Christopher E. Paine. A case against virtual nuclear testing. Scientific American, 281(3):74–79, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
74.pdf.

Levin:1999:TST

[5216] Theodore C. Levin and Michael E. Edgerton. The throat singers of Tuva. Scientific American, 281(3):80–87, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-80.pdf;   http://www.sciam.com/
1999/0909issue/0999levin.html.

Larson:1999:SRA

[5217] Edward J. Larson and Larry Witham. Scientists and religion in America. Scientific American, 281(3):88–93, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-88.pdf.

Carlson:1999:ASC

[5218] Shawn Carlson. The amateur scientist: Counting atmospheric ions. Scientific American, 281(3):96–97, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-96.pdf.

Stewart:1999:MRDb

[5219] Ian Stewart. Mathematical recreations: Dances with dodecahedra. Scientific American, 281(3):98–100, September 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-98.pdf.

Anonymous:1999:RCe

[5220] Anonymous. Reviews and commentaries. Scientific American, 281 (3):101–103, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n3/pdf/scientificamerican0999-
101.pdf.

Burke:1999:M

[5221] James Burke. Or maybe not. Scientific American, 281(3):104–105, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n3/pdf/scientificamerican0999-104.pdf.

Burke:1999:CFFa

[5222] James Burke. Connections: Fallacies, forgeries and continental drift. Scientific American, 281(3):104–??, September 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Silberman:1999:WKR

[5223] Cyril Silberman and Earl Santee. Working knowledge: Retractable stadium roofs. Scientific American, 281(3):106, September 1999.

CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n3/pdf/scientificamerican0999-106.pdf.

Alda:1999:ECS

[5224] Alan Alda. From the Editors: Curiosity and Schrödinger’s Cat. Scientific American, 281(4):10, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
10.pdf.

Anonymous:1999:LEk

[5225] Anonymous. Letters to the Editors. Scientific American, 281(4):14, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-14a.pdf.

Anonymous:1999:Eg

[5226] Anonymous. Erratum. Scientific American, 281(4):14–16, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-14b.pdf.

Anonymous:1999:YAIj

[5227] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 281(4):18, October 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-18.pdf.

Beardsley:1999:NAT

[5228] Tim Beardsley. News and analysis: Truth or consequences. Scientific American, 281(4):21–24, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
21.pdf.

Anonymous:1999:FPP

[5229] Anonymous. In focus: Polygraph perils for scientists. Scientific American, 281(4):21–??, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Williams:1999:PF

[5230] Wendy Williams. Pirate fear. Scientific American, 281(4):26–30, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-26.pdf.

Collins:1999:QDV

[5231] Graham P. Collins. Quantum déjà vu. Scientific American, 281(4):30–32, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-30.pdf.

Anonymous:1999:Bg

[5232] Anonymous. In brief. Scientific American, 281(4):32–34, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-32a.pdf.

Anonymous:1999:FF

[5233] Anonymous. Fat in the fire. Scientific American, 281(4):32–34, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-32b.pdf.

Musser:1999:BB

[5234] George Musser. Boom or bust? Scientific American, 281(4):34–36, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-34.pdf.

Simpson:1999:DS

[5235] Sarah Simpson. Deserting the Sahara. Scientific American, 281(4):36–38, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-36.pdf.

Mirsky:1999:OTO

[5236] Steve Mirsky. It’s not oeuvre till it’s oeuvre. Scientific American, 281(4): 38, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-38.pdf.

Doyle:1999:MWC

[5237] Rodger Doyle. Men, women and college. Scientific American, 281(4):40, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-40.pdf.

Teller:1999:IHA

[5238] Edward Teller. Infamy and honor at the Atomic Café. Scientific American, 281(4):42–44, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
42.pdf.

Menduno:1999:DIY

[5239] Michael Menduno. Do-it-yourself financial planning. Scientific American, 281(4):46, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
46.pdf.

Holmes:1999:D

[5240] Liz Holmes. In the drink. Scientific American, 281(4):47–48, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-47.pdf.

Beardsley:1999:BG

[5241] Tim Beardsley. To boldly grow. … Scientific American, 281(4):48–49, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-48B.pdf.

Wallich:1999:EM

[5242] Paul Wallich. To err is mechanical. Scientific American, 281(4):50, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-50.pdf.

Pappalardo:1999:HOE

[5243] Robert T. Pappalardo, James W. Head, and Ronald Greeley. The hidden ocean of Europa. Scientific American, 281(4):54–63, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-54.pdf.

Eberhart:1999:WTB

[5244] Mark E. Eberhart. Why things break. Scientific American, 281(4):66–73, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-66.pdf.

Agnew:1999:PNL

[5245] Neville Agnew and Shin Maekawa. Preserving Nefertari’s legacy. Scientific American, 281(4):74–79, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
74.pdf.

DiBisceglie:1999:UCH

[5246] Adrian M. Di Bisceglie and Bruce R. Bacon. The unmet challenges of hepatitis C. Scientific American, 281(4):80–85, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-80.pdf.

Gibbs:1999:TEF

[5247] W. Wayt Gibbs and Douglas Fox. Trends in education: The false crisis in science education. Scientific American, 281(4):86–93, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-86.pdf;   http:/
/www.sciam.com/1999/1009issue/1099gibbs.html.

Clark:1999:SRH

[5248] David D. Clark. Special report:: High-speed data races home. Scientific American, 281(4):94–99, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
94.pdf;   http://www.sciam.com/1999/1009issue/1099clark.html.

Medin:1999:IC

[5249] Milo Medin and Jay Rolls. The Internet via cable. Scientific American, 281(4):100–101, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
100.pdf.

Medin:1999:SRI

[5250] Milo Medin and Jay Rolls. Special report: The Internet via cable. Scientific American, 281(4):100–101, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1999/1009issue/1099medin.html.

Hawley:1999:DBP

[5251] George T. Hawley. Dsl: Broadband by phone. Scientific American, 281(4):102–103, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
102.pdf.

Hawley:1999:SRD

[5252] George T. Hawley. Special report: DSL: Broadband by phone. Scientific American, 281(4):102–103, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1999/1009issue/1099hawley.html.

Shumate:1999:BB

[5253] Paul W. Shumate, Jr. The broadest broadband. Scientific American, 281(4):104–105, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
104.pdf;   http://www.sciam.com/1999/1009issue/1099shumate.
html.

Norcross:1999:SSH

[5254] Robert P. Norcross. Satellites: The strategic high ground. Scientific American, 281(4):106–107, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
106.pdf.

Norcross:1999:SRS

[5255] Robert P. Norcross. Special report: Satellites: The strategic high ground. Scientific American, 281(4):107–108, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1999/1009issue/1099norcross.html.

Skoro:1999:LBW

[5256] John Skoro. LMDS: Broadband wireless access. Scientific American, 281(4):108–109, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
108.pdf.

Bane:1999:LEP

[5257] P. William Bane and Stephen P. Bradley. The light at the end of the pipe. Scientific American, 281(4):110–115, October 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-110.pdf.

Stewart:1999:MRCb

[5258] Ian Stewart. Mathematical recreations: Cone with a twist. Scientific American, 281(4):116–117, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
116.pdf.

Carlson:1999:ASMb

[5259] Shawn Carlson. The amateur scientist: Modeling the atomic Universe. Scientific American, 281(4):118–119, October 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n4/pdf/scientificamerican1099-118.pdf.

Robinson:1999:RC

[5260] Leif J. Robinson. Reviews and commentaries. Scientific American, 281(4):120–123, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
120.pdf.

Anonymous:1999:ERg

[5261] Anonymous. The Editors recommend. Scientific American, 281(4):123, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-123.pdf.

Rennie:1999:TER

[5262] John Rennie. A total eclipse of reason. Scientific American, 281(4):124, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-124.pdf.

Rennie:1999:CEJ

[5263] John Rennie. Commentary: Eclipsed judgment in Kansas. Scientific American, 281(4):124–??, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Morrison:1999:NDS

[5264] Philip Morrison and Phylis Morrison. Nitrogen: The dark side. Scientific American, 281(4):125, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
125.pdf.

Morrison:1999:WNE

[5265] Philip Morrison. Wonders: Nitrogen’s explosive nature. Scientific American, 281(4):125–??, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:MD

[5266] James Burke. A matter of degree. Scientific American, 281(4):126–127, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n4/pdf/scientificamerican1099-126.pdf.

Burke:1999:CSE

[5267] James Burke. Connections: Stretching equatorial rubber to the North Pole. Scientific American, 281(4):126–??, October 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bloomfield:1999:WKL

[5268] Louis A. Bloomfield. Working knowledge: Laser printers. Scientific American, 281(4):128, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n4/pdf/scientificamerican1099-
128.pdf.

Rennie:1999:EWW

[5269] John Rennie. From the Editors: Who wants to live forever? Scientific American, 281(5):6, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
6.pdf.

Anonymous:1999:LEl

[5270] Anonymous. Letters to the Editors. Scientific American, 281(5):8–10, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-8.pdf.

Anonymous:1999:Eh

[5271] Anonymous. Erratum. Scientific American, 281(5):10, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-10.pdf.

Anonymous:1999:FAS

[5272] Anonymous. In focus: Asthma’s strange rise in the inner city. Scientific American, 281(5):13–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:YAIk

[5273] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 281(5):14, November 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-14.pdf.

Alpert:1999:NAI

[5274] Mark Alpert. News and analysis: The invisible epidemic. Scientific American, 281(5):19–20, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
19.pdf.

Appell:1999:SS

[5275] David Appell. Speaking up for science. Scientific American, 281(5):22–24, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
22.pdf.

Mirsky:1999:FF

[5276] Steve Mirsky. Floral fiend. Scientific American, 281(5):24, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-24.pdf.

Zorpette:1999:TW

[5277] Glenn Zorpette. A taste of weightlessness. Scientific American, 281(5):26, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
26.pdf.

Mirsky:1999:F

[5278] Steve Mirsky. Down in front. Scientific American, 281(5):28, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-28a.pdf.

Yam:1999:Bd

[5279] Philip Yam. In brief. Scientific American, 281(5):28–30, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-28b.pdf.

Howard:1999:MMP

[5280] Ken Howard. Mickey Mouse, Ph.D. Scientific American, 281(5):30–37, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-30.pdf.

Krauss:1999:FLUa

[5281] Lawrence M. Krauss and Glenn D. Starkman. The fate of life in the universe. Scientific American, 281(5):36–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Collins:1999:QCb

[5282] Graham P. Collins. Quantum claustrophobia. Scientific American, 281(5):37–39, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
37.pdf.

Anonymous:1999:CF

[5283] Anonymous. Campaign finance. Scientific American, 281(5):38, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-38.pdf.

Holloway:1999:AS

[5284] Marguerite Holloway. The ascent of scent. Scientific American, 281(5):42–44, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
42.pdf.

Logothetis:1999:VWCa

[5285] Nikos K. Logothetis. Vision: A window on consciousness. Scientific American, 281(5):44–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Klotzko:1999:PP

[5286] Arlene Judith Klotzko. Pork progress. Scientific American, 281(5):48, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-48.pdf.

Zucker:1999:MM

[5287] Mimi Zucker. Mind over matter. Scientific American, 281(5):51–52, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-51.pdf.

Bojin:1999:LBS

[5288] Marie-Hélène Bojin and Gary Stix. Little big science. Scientific American, 281(5):52–55, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
52.pdf.

Suess:1999:FIa

[5289] Erwin Suess, Gerhard Bohrmann, Jens Greinert, and Erwin Lausch. Flammable ice. Scientific American, 281(5):52–??, November 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Grossman:1999:NWR

[5290] Wendy M. Grossman. No way to run a network. Scientific American, 281(5):56, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
56.pdf.

Krauss:1999:FLUb

[5291] Lawrence M. Krauss and Glenn D. Starkman. The fate of life in the Universe. Scientific American, 281(5):58–65, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-58.pdf.

Topoff:1999:SMQa

[5292] Howard Topoff. Slave-making queens. Scientific American, 281(5):60–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fink:1999:TRAa

[5293] Mathias Fink. Time-reversed acoustics. Scientific American, 281(5): 67–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Logothetis:1999:VWCb

[5294] Nikos K. Logothetis. Vision: a window on consciousness. Scientific American, 281(5):68–75, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
68.pdf.

Suess:1999:FIb

[5295] Erwin Suess, Gerhard Bohrmann, Jens Greinert, and Erwin Lausch. Flammable ice. Scientific American, 281(5):76–83, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-76.pdf.

Hagenlocher:1999:ZCa

[5296] Klaus G. Hagenlocher. A Zeppelin for the 21st century. Scientific American, 281(5):80–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

1999/1109issue/1199hagenlocher.html.

Topoff:1999:SMQb

[5297] Howard Topoff. Slave-making queens. Scientific American, 281(5):84–90, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-84.pdf.

Scott:1999:BFRa

[5298] Phil Scott. The balloon that flew round the world. Scientific American, 281(5):86–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/1999/

1109issue/1199scott.html.

Yunus:1999:GBa

[5299] Muhammad Yunus. The Grameen Bank. Scientific American, 281(5): 90–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Fink:1999:TRAb

[5300] Mathias Fink. Time-reversed acoustics. Scientific American, 281 (5):91–97, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
91.pdf.

Smith:1999:FS

[5301] I. Steve Smith, Jr. and James A. Cutts. Floating in space. Scientific American, 281(5):98–103 (Intl. ed. 74–??), November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/1999/1109issue/1199smith.html.

Smith:1999:SAB

[5302] I. Steve Smith, Jr. and James A. Cutts. Space-age ballooning: Floating in space. Scientific American, 281(5):98–103, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-98.pdf.

Hagenlocher:1999:ZCb

[5303] Klaus G. Hagenlocher. A Zeppelin for the 21st Century. Scientific American, 281(5):104–109, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
104.pdf.

Morrison:1999:WLBa

[5304] Philip Morrison and Phylis Morrison. Wonders: Leonardo’s bronze horse. Scientific American, 281(5):105–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:CFFb

[5305] James Burke. Connections: From Frankenstein to feminists. Scientific American, 281(5):106–??, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Scott:1999:BFRb

[5306] Phil Scott. The balloon that flew round the world. Scientific American, 281(5):110–113, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
110.pdf.

Yunus:1999:GBb

[5307] Muhammad Yunus. The Grameen Bank. Scientific American, 281 (5):114–119, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
114.pdf.

Carlson:1999:ASF

[5308] Shawn Carlson. The amateur scientist: Falling into chaos. Scientific American, 281(5):120–121, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
120.pdf.

Stewart:1999:MRM

[5309] Ian Stewart. Mathematical recreations: Most-perfect magic squares. Scientific American, 281(5):122–123, November 1999. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
122.pdf.

Metzinger:1999:RCH

[5310] Thomas Metzinger. Reviews and commentaries: The hint half guessed. Scientific American, 281(5):125–126, November 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n5/pdf/scientificamerican1199-125.pdf.

Anonymous:1999:ERh

[5311] Anonymous. The Editors recommend. Scientific American, 281 (5):127–128, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
127.pdf.

Morrison:1999:SR

[5312] Philip Morrison and Phylis Morrison. The surefire résumé. Scientific American, 281(5):129–131, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
129.pdf.

Burke:1999:VU

[5313] James Burke. Various, unrequited. Scientific American, 281(5):130–131, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n5/pdf/scientificamerican1199-130.pdf.

Burrell:1999:WKT

[5314] Kenneth H. Burrell. Working knowledge: Toothpaste. Scientific American, 281(5):132, November 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n5/pdf/scientificamerican1199-
132.pdf.

Rennie:1999:EOT

[5315] John Rennie. From the Editors: Overtaking tomorrow. Scientific American, 281(6):12, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
12.pdf.

Anonymous:1999:LEm

[5316] Anonymous. Letters to the Editors. Scientific American, 281(6):14–16, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-14.pdf.

Anonymous:1999:YAIl

[5317] Anonymous. 50, 100 and 150 years ago: Innovation and discovery as chronicled in Scientific American: 50, 100 and 150 years ago. Scientific American, 281(6):18, December 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-18.pdf.

Holloway:1999:NAA

[5318] Marguerite Holloway. News and analysis: The aborted crime wave? Scientific American, 281(6):23–24, December 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-23.pdf.

Anonymous:1999:FDL

[5319] Anonymous. In focus: Did legalizing abortion really cause a drop in crime? Scientific American, 281(6):23–??, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nemecek:1999:PP

[5320] Sasha Nemecek. A plan for Panama. Scientific American, 281(6):26, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-26.pdf.

LeVay:1999:ST

[5321] Simon LeVay. Stress test. Scientific American, 281(6):28–30, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-28.pdf.

Alpert:1999:AD

[5322] Mark Alpert. Apocalypse deferred. Scientific American, 281(6):30, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-30.pdf.

Mirsky:1999:NU

[5323] Steve Mirsky. Notes from the underground. Scientific American, 281(6):32, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
32b.pdf.

Anonymous:1999:Bh

[5324] Anonymous. In brief. Scientific American, 281(6):32–34, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-32a.pdf.

Anonymous:1999:CI

[5325] Anonymous. Cave inn. Scientific American, 281(6):34–37, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-34.pdf.

Doyle:1999:DM

[5326] Rodger Doyle. The decline of marriage. Scientific American, 281(6):36–37, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
36.pdf.

Lewis:1999:DT

[5327] Julie Lewis. Driven up a tree. Scientific American, 281(6):40–42, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-40.pdf.

Beardsley:1999:FMS

[5328] Tim Beardsley. Fly me to the stars. Scientific American, 281(6):46, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-46.pdf.

Beardsley:1999:WC

[5329] Tim Beardsley. But where are the cupholders? Scientific American, 281(6):50, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
50a.pdf.

Stix:1999:BC

[5330] Gary Stix. Bad connections. Scientific American, 281(6):50–54, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-50b.pdf.

Pescovitz:1999:CF

[5331] David Pescovitz. Cable-free. Scientific American, 281(6):55–56, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-55.pdf.

Beardsley:1999:NS

[5332] Tim Beardsley. No secrets. Scientific American, 281(6):58, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-58.pdf.

Dupont:1999:SB

[5333] Daniel G. Dupont. Seen before. Scientific American, 281(6):58–59, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-56.pdf.

Pescovitz:1999:YLH

[5334] David Pescovitz. To your on-line health. Scientific American, 281(6):60, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-60.pdf.

Maddox:1999:USC

[5335] Sir John Maddox. The unexpected science to come. Scientific American, 281(6):62–67, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
62.pdf;   http://www.sciam.com/1999/1209issue/1299maddox.html.

Weinberg:1999:UP

[5336] Steven Weinberg. A unified physics by 2050? Scientific American, 281(6):68–75, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
68.pdf;   http://www.sciam.com/1999/1209issue/1299weinberg.
html.

Rees:1999:EOU

[5337] Martin Rees. Exploring our Universe and others. Scientific American, 281(6):78–83, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
78.pdf.

Collins:1999:DCL

[5338] Francis S. Collins and Karin G. Jegalian. Deciphering the code of life. Scientific American, 281(6):86–91, December 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-86.pdf.

deWaal:1999:ENV

[5339] Frans B. M. de Waal. The end of nature versus nurture. Scientific American, 281(6):94–99, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
94.pdf.

Karl:1999:HICa

[5340] Thomas R. Karl and Kevin E. Trenberth. The human impact on climate. Scientific American, 281(6):100–105, December 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-100.pdf.

Rose:1999:CHA

[5341] Michael R. Rose. Can human aging be postponed? Scientific American, 281(6):106–111, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
106.pdf.

Damasio:1999:HBCa

[5342] Antonio R. Damasio. How the brain creates the mind. Scientific American, 281(6):112–117, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
112.pdf.

Domasio:1999:HBC

[5343] Antonio R. Domasio. How the brain creates the mind. Scientific American, 281(6):112–??, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Tarter:1999:TLE

[5344] Jill C. Tarter and Christopher F. Chyba. Is there life elsewhere in the Universe? Scientific American, 281(6):118–123, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-118.pdf.

Moravec:1999:RR

[5345] Hans Moravec. Rise of the robots. Scientific American, 281(6):124–135, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-124.pdf.

Stewart:1999:MRDc

[5346] Ian Stewart. Mathematical recreations: Defend the Roman Empire! Scientific American, 281(6):136–138, December 1999. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-136.pdf.

Fisher:1999:RCM

[5347] Helen Fisher. Reviews and commentaries: Mother Nature is an old lady with bad habits. Scientific American, 281(6):140–142, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v281/
n6/pdf/scientificamerican1299-140.pdf.

Anonymous:1999:ERi

[5348] Anonymous. The Editors recommend. Scientific American, 281 (6):142–145, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
142.pdf.

Morrison:1999:CP

[5349] Philip Morrison and Phylis Morrison. A century of physics. Scientific American, 281(6):146–148, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
146.pdf.

Morrison:1999:WLBb

[5350] Philip Morrison and Phylis Morrison. Wonders: Looking back at the Century of Physics. Scientific American, 281(6):146–??, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Burke:1999:Zb

[5351] James Burke. The O Zone. Scientific American, 281(6):147–148, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-147.pdf.

Burke:1999:CEC

[5352] James Burke. Connections: Explosive cotton, elephant teeth and electromagnetic fields. Scientific American, 281(6):147–??, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:AI

[5353] Anonymous. Annual index 1999. Scientific American, 281(6):149–151, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v281/n6/pdf/scientificamerican1299-149.pdf.

Bloomfield:1999:WKW

[5354] Louis A. Bloomfield. Working knowledge: Water filters. Scientific American, 281(6):152, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v281/n6/pdf/scientificamerican1299-
152.pdf.

Anonymous:1999:EMS

[5355] Anonymous. End of the Millennium Special Issue. Scientific American, 281(6):??, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Damasio:1999:HBCb

[5356] Antonio R. Damasio. How the brain creates the mind. Scientific American, 281(6):??, December 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Karl:1999:HICb

[5357] Thomas R. Karl and Kevin E. Trenberth. The human impact on climate. Scientific American, 281(6):??, December 1999. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

White:1999:HT

[5358] Robert J. White. Head transplants. Scientific American, ??(2):24–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Zorpette:1999:MA

[5359] Glenn Zorpette. Muscular again. Scientific American, ??(2):27–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Strauss:1999:FF

[5360] Evelyn Strauss. Feeling the future. Scientific American, ??(2):44–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Pescovitz:1999:GRC

[5361] David Pescovitz. Getting real in cyberspace. Scientific American, ??(2): 48–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Small:1999:NM

[5362] Meredith F. Small. Nosing out a mate. Scientific American, ??(2):52– ??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powledge:1999:UBB

[5363] Tabitha M. Powledge. The ultimate baby bottle. Scientific American, ??(2):96–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1999:EEP

[5364] Anonymous. Engineering at the edge of the possible. Scientific American, ??(3):8–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Kosowatz:1999:MM

[5365] John J. Kosowatz. Mighty monolith. Scientific American, ??(3):14– ??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nemecek:1999:SAR

[5366] Sasha Nemecek. Some assembly required. Scientific American, ??(3):24– ??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Freeman:1999:BGS

[5367] Eva Freeman. Building gargantuan software. Scientific American, ??(3): 28–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Beardsley:1999:LS

[5368] Tim Beardsley. Life in space. Scientific American, ??(3):32–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Voss:1999:SW

[5369] David Voss. A small world. Scientific American, ??(3):34–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Holloway:1999:BBB

[5370] Marguerite Holloway. Bringing back the barrier. Scientific American, ??(3):38–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Musser:1999:SWM

[5371] George Musser. Seven wonders of modern astronomy. Scientific American, ??(3):42–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Netting:1999:BCF

[5372] Jessa Netting. A bridge to a composite future. Scientific American, ??(3):50–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Nemecek:1999:SSR

[5373] Sasha Nemecek. Subterranean speed record. Scientific American, ??(3): 52–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gibbs:1999:BB

[5374] W. Wayt Gibbs. Blitzing bits. Scientific American, ??(3):56–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Howard:1999:HTR

[5375] Ken Howard. Harder than rocket science. Scientific American, ??(3):62– ??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hayashi:1999:SL

[5376] Alden M. Hayashi. The sky’s the limit. Scientific American, ??(3):66– ??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Roesset:1999:BS

[5377] Jose M. Roesset. To the bottom of the sea. Scientific American, ??(3): 73–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Hopkin:1999:DG

[5378] Karen Hopkin. Designer genomes. Scientific American, ??(3):78–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Lundhus:1999:BBS

[5379] Peter Lundhus. Bridging borders in Scandinavia. Scientific American, ??(3):82–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Alpert:1999:GPN

[5380] Mark Alpert. The greatest projects never built. Scientific American, ??(3):90–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Petroski:1999:HEE

[5381] Henry Petroski. The hubris of extreme engineering. Scientific American, ??(3):94–??, ???? 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Barnes:1991:CGb

[5382] Joshua Barnes, Lars Hernquist, and François Schweizer. Colliding galaxies. Scientific American [International Edition], 265(2):40–?? (Intl. ed. 26–33), August 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bishop:1993:RHTb

[5383] David J. Bishop, Peter L. Gammel, and David A. Huse. Resistance in high-temperature superconductors. Scientific American [International Edition], 268(2):48–?? (Intl. ed. 24–31), February 1993. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Birge:1995:PBCb

[5384] R. R. Birge. Protein-based computers. Scientific American [International Edition], 272(3):66–71, March 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Abu-Mostafa:1995:MLHb

[5385] Yaser S. Abu-Mostafa. Machines that learn from hints. Scientific American [International Edition], 272(4):68–73, April 1995. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Angel:1996:SLOb

[5386] J. R. P. Angel and N. J. Woolf. Searching for life on other planets. Scientific American [International Edition], 274(4):46–52, April 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bell:1996:NGCb

[5387] A. E. Bell. Next-generation compact discs. Scientific American [International Edition], 275(1):28–32, July 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.sciam.com/

0796issue/0796currentissue.html.

Beardsley:1997:TNM

[5388] Tim Beardsley. Trends in neuroscience: the machinery of thought. Scientific American [International Edition], 277(2):78–83, August 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.sciam.com/0897issue/0897currentissue.html.

Girard:1990:LL

[5389] J.-Y. Girard. La logique linéaire. Pour La Science, Édition Française de Scientific American, 150(????):74–85, April 1990. ISSN 0153-4092.